A Complete Bibliography of Publications in Isis, 1990–1999

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 0.08

Title word cross-reference

c [1275]. AΠOP HMA [2901]. BOT ANIKON [2901]. ΠEP IT ΩNΠEΠONΘΩNT OΠΩN [1716]. ⊂ [431]. ⊃ [431].

-1708 [2436]. -4 [3189].

/Max [3367, 1215].

0Die [1766].

1 [1169, 2655, 2935, 566, 1131, 1939]. 1.7 [1001]. 1.7-7 [1001]. 10 [2649, 2983]. 100 [323]. 129 [1808]. 1333 [1938]. 1336 [2425]. 1345 [2250, 920]. 1400 [3429]. 1420 [2078]. 1450 [1797]. 1483 [348]. 150-Year [2452]. 1500 [29]. 1530 [30]. 1543 [441]. 1550 [2160, 3491, 1246]. 1570 [1998]. 1597 [3531]. 1600 [3326, 2734, 440, 151, 347]. 1610 [1724]. 1610/11 [1651]. 1620 [2652]. 1626 [2003]. 1632 [2000]. 1650 [1377]. 1653 [2901].

1654 [2346]. 1657 [732]. 1659 [2816]. 1662 [357]. 1676 [1379, 452]. 1683 [1531]. 1685 [838]. 1687 [1976]. 1690 [2661]. 1696 [1531]. 1699 [835]. 1700 [34, 2491, 3315, 2975]. 1701 [2512]. 1715 [1820]. 1718 [2167]. 1727 [1193, 42]. 1730 [1733]. 1740 [2899]. 1742 [260]. 1750 [3140, 1479, 1560, 3142, 1286, 1566, 2746, 3141, 2351, 1385, 3404]. 1753 [456]. 1770 [460, 3152]. 1773 [3342]. 1777 [1483]. 1783 [2749]. 1785 [3057]. 1789 [461]. 1789/90 [461]. 1791 [3146]. 1792 [1734]. 1795 [2174, 165]. 1799 [561, 3442]. 17de [2814]. 1800 [2356, 326, 2412, 44, 923, 1928, 2902, 2101, 932, 245, 3590]. 1801 [935]. 1802 [2275]. 1803 [167]. 1804 [1829]. 1810 [3144]. 1811 [2906]. 1812 [1301]. 1813 [3259]. 1815 [3443]. 1820 [2989, 1080, 1480]. 1825 [2443]. 1826 [1482]. 1828 [364]. 1830 [49, 2091]. 1831 [2982]. 1832 [3552]. 1837 [540, 1394]. 1843 [2526, 463, 3258]. 1846 [1945]. 1847 [2984]. 1847/48 [2984]. 1848 [1123, 2176]. 1850 [2491, 1708, 2100, 1027, 2827, 269, 2095, 1824, 2014, 1153, 2520, 23]. 1851 [185, 1753]. 1853 [2991]. 1855 [852, 2603]. 1857 [852, 1543, 1748]. 1859 [938, 3127, 3353]. 1860 [1748, 3163, 2751, 3145, 53, 2239, 274, 46, 1126]. 1861 [2449]. 1862 [2756, 2912, 3263, 1487]. 1864 [939]. 1870 [942, 1495, 3162, 844]. 1871 [2184]. 1873 [372, 2280]. 1875 [3270]. 1876 [1834, 1947]. 1877 [2369]. 1880 [2368, 2454, 2910, 3357]. 1881 [943]. 1889 [2759, 2371]. 1892 [3077]. 1893 [1699]. 1895 [2458]. 1899 [88, 1401]. 1900 [1309, 1399, 1535, 2576, 177, 759, 1228, 722, 1933, 1743, 1263, 2364, 3540]. 1902 [3348, 470]. 1905 [2669]. 1906 [3073, 1130, 670]. 1909 [768, 904]. 1910 [52, 3268]. 1910s [3464]. 1910s-1970s [3464]. 1911 [2117, 2278]. 1913 [2675]. 1914 [2032, 627, 2019, 2824, 2757, 1209, 2460, 1946, 2470, 2357, 2117, 3161, 2450, 171, 1312, 3544, 945, 1599, 762, 2182]. 1915 [2695, 1128, 180]. 1917 [476, 2766]. 1918 [1262, 2773, 1580, 466]. 1919 [2604, 273]. 1920 [1668, 376, 1494, 2761, 695, 230, 1757]. 1921 [713, 2366]. 1922 [2015, 292]. 1923 [865]. 1924 [2119]. 1925 [3180, 2459]. 1926 [212, 2677, 1214, 3277]. 1927 [3375]. 1929 [947, 2026, 765]. 1930 [1851, 912, 2776, 596]. 1933 [2534]. 1934 [3548, 3287]. 1935 [3078, 3474, 1590]. 1937 [291]. 1938 [1596, 3120]. 1939 [2114, 2674, 2285, 2993, 1677, 225, 212]. 1939-1945 [225]. 1940 [1706, 130, 585, 2616, 675, 1843, 3098, 3450, 1670, 22, 1235, 1508, 1233, 2699, 2292]. 1941 [3463]. 1943 [1215]. 1944 [3001]. 1945 [968, 1490, 2383, 951, 70, 1143, 2125, 595, 2288, 1458, 2296, 2414, 225, 2465, 2413, 3480, 1501, 1766, 1952, 2204, 676, 712]. 1946 [2540, 1777]. 1947 [766, 225]. 1947-1977 [225]. 1948 [2532, 2535]. 1949 [1402, 2291, 873]. 1950 [2491, 1084, 1591, 2181, 1742, 3071, 1267, 2033, 2314]. 1950s [2566, 225]. 1951 [1053, 2542]. 1953 [679]. 1954 [3585]. 1955 [3577, 2106, 2132, 877]. 1956 [2206]. 1957 [1859]. 1960 [589, 1868]. 1960s [3562]. 1961 [2386, 2609]. 1963 [2200, 2607, 2130]. 1964 [584, 2211]. 1965 [1043, 289, 3385, 1046]. 1967 [481]. 1968 [1227]. 1969 [1848]. 1970 [629, 1219, 2682, 816, 875, 3584, 2728, 1234, 2876]. 1970s [3464, 3562]. 1971 [598, 1867]. 1972 [2467]. 1973 [3079]. 1974 [139, 2213]. 1975 [1329]. 1977 [225, 597]. 1978 [2201]. 1980 [1412, 378]. 1982 [481, 2781]. 1984 [160]. 1985 [2193, 2547, 1054, 242]. 1986 [293, 3292]. 1987 [1087, 858, 1509]. 1988 [101, 605, 1597, 1278, 1393, 1268]. 1989 [221, 312, 618, 620, 1767]. 1990 [409, 614, 898, 706, 500, 2287, 583, 1779, 3011, 1979]. 1991 [705, 707, 896, 2385, 1415]. 1992 [2313, 1258, 1941, 1626]. 1993 [1614, 1792, 1793, 2483, 2453, 3252, 2416, 1892]. 1994 [2946, 1889, 1972, 1890, 2057, 1973, 1791]. 1995 [2312, 2311, 2634, 2401, 3402, 3094, 2646, 3370, 2283]. 1996 [2482, 2626, 2625, 3030, 2954, 2624, 2627, 2687, 2670]. 1997 [2955, 2715, 3500, 2956, 2947]. 1998 [3302, 3303, 3311, 3312, 3047]. 1c [348].

2 [1280, 1476, 2192, 1359, 2970, 3490, 1877]. 2.1 [2649]. 2.1-10 [2649]. 2003 [140]. 2089 [566]. 2089-1 [566]. 20th [902]. 220 [1808]. 250-jährigen [471].

3 [2655, 1359, 1366]. 3.4-8 [1092]. 300 [160]. 34 [961].

4 [3189]. 41 [2939]. 48 [2984]. 49 [580].

5 [3246]. 5.-9 [1993]. 56 [182].

6 [3250]. ’600 [645]. 60er [3586].

7 [1001]. ’700 [564]. 7A [2347].

90 [461]. 96 [3171].

= [1737, 2733].

âge [825, 210]. être [821].

A. [807]. A.D [2412, 245, 1797]. A.D. [1907]. Aaboe [1462, 531]. Aafke [1913]. aan [361]. Aaron [1673, 1047]. Årsbok [514]. Aaserud [949]. Abbasid [1275]. Abbo [1808]. Abbott [75]. Abbreviation [2426]. Abd [1810]. abditis [2587]. Abdus [191]. Abe [2199]. Abed [1006]. Abel [3326, 1497]. Abeles [2456]. Abendland [989]. Abir [2723]. Abir-Am [2723]. Able [3064]. Abnormal [1673]. Aboin [1374]. Abortion [3079, 2156, 3232, 481, 481, 481]. Abortive [96]. Abraham [248, 1998, 579, 67, 1408, 2640, 2605, 2196]. Abram [861]. Abrégé [1559]. Abridged [3436]. Absolute [622]. absoluten [2755]. absolutischer [3341]. Absolutism [1616]. Abtreibungsmittel [2081]. Abu [2426]. Abul [3525, 1810]. Abu’l-Fath [1810]. Abundance [171]. Abuse [2701]. Abzug [3006]. ac [2587]. Academia [184, 590]. Academiae [504]. Academic [3310, 876, 1954, 3151, 3557, 1757, 95, 2315, 64]. Academiciens [2747]. Academie [3020, 2746, 2618, 745]. Academies [2397, 1052]. Academy [1626, 3104, 835, 927, 3073, 2264, 2602, 618, 2357, 816, 1695, 2307, 3576, 2225]. Acar [1245]. accademici [2590]. Access [1778]. Accomplishments [2670]. According [3240, 1015]. Account [709, 3573, 1087, 225, 379, 480]. Accounts [2653]. Accuracy [1055]. Accused [1958]. Aceves [2444]. Achenbaum [2697]. Achievements [3315]. Achinstein [1059]. Achté [513]. achttiende [1297]. Acid [1135]. Ackerman [2078, 762]. Acoustic [2026]. Acoustics [3050, 2792]. across [1776, 1169, 3282]. Act [670]. Acta [504, 523]. Actes [3460]. Action [2552, 1656, 2467]. Activism [2796]. Activist [2201]. Activities [2619, 3008]. Activity [1162, 2591, 3211]. Actors [3008, 970]. actuelle [2247]. ad [2903, 523]. Adam [3527, 3417, 603, 538, 2342]. Adams [2748, 1243, 2961, 2080, 2499, 957, 1243, 588, 2126]. Adamson [148, 2431]. Adaptation [610]. Adas [419]. Addiction [1967, 2927, 1967]. Addinall [1206]. Additions [2788]. Adelard [2426]. Adele [1656, 1506]. Adelmo [1002]. Adept [3502]. Adjusting [1768]. Administrateurs [1931]. Administration [2130, 1987]. Adolf [1240, 779]. Adolph [2830]. Adrian [2287, 1169, 3152, 467]. Adrianus [2310]. Adrien [729]. Adrienne [2032, 1015]. Advanced [680]. Advancement [2851, 1773]. Advent [628]. Adventurer [1731]. Adventures [2830]. Adversaries [2969]. Advertising [1414]. Advice [1584]. Advisers [1330, 1584]. Advising [1584]. Advisors [672]. Advisory [225]. Advocated [1922]. AEG [2923]. Aegidii [2249]. Aequaties [658]. Aerial [85]. Aeronautical [197, 373]. Aesthetic [3522, 692, 753]. Aesthetics [2858, 3223]. Aether [147]. Aevi [1637, 2253]. Affair [1017, 2227, 2136, 1147, 1189]. Affect [3307]. Affinitat [1927]. Affinities [1076]. Affirmative [2467]. Afflictions [1743]. Afield [2367]. Afines [525]. Africa [2876, 593, 2696]. African [3190, 3392, 3180, 912, 301, 2128]. African-American [912]. Aftalion [1317]. after [3451, 2056, 2473, 3566, 3371, 3470]. Again [1971]. Against [915, 342, 2924, 2845, 3013, 2310, 2693, 3267, 3388, 1602]. Agassiz [1202]. Age [601, 3419, 455, 1048, 49, 3205, 3463, 1080, 3235, 2698, 1299, 1414, 923, 814, 733, 198, 2563, 2299, 908, 3362, 2657, 3286, 1482, 22, 2288, 2245, 3345, 2439, 2976, 2159, 2174, 245, 1769, 672, 2247, 3062, 2021, 1081, 3477, 575, 628, 363, 1732, 225, 966, 1836, 3459, 2878, 1145]. Agency [881, 2550, 3220]. Agenda [1590, 1926]. Agent [2053, 729]. Ages [2806, 1726, 1280, 826, 3228, 3123, 1636, 3335, 807, 1636, 2138, 148, 726, 1632, 1719, 1812, 2807, 901, 732]. Aggregationes [1641]. Agnes [2549]. Agneta [1366]. Agreement [1922]. Agricola [1801, 827]. Agricultural [2565, 3156, 2049, 3510]. Agriculturalists [291]. Agriculturchemie [266]. Agriculture [2073, 135]. Agrimi [638, 1994]. Agrippa [1918]. Agro [2972]. Agro-Industries [2972]. Aguilar [432]. Agustín [523]. Aharon [1683]. ahjar [2071]. Ahmad [510, 3525]. Ahmet [1452]. AIDS [1141, 2796, 2804, 2782, 195]. Aigentler [2457]. Aim [1253]. Aiming [2854]. Ainley [1859]. Air [373]. Airplane [3480, 3006]. Airs [2970]. Airshipmen [1235]. Aiton [705, 3055, 359, 2903]. Akademgorodok [3172]. Akademie [1830, 837]. akademii [887]. akademische [1830]. Akerman [1702, 1562]. Akhundov [203]. Akustik [3145]. Al [3242, 730, 1631, 2645, 1828, 2834, 2443, 2526, 2490, 1005, 728, 431, 2340, 2877, 3154, 1991, 1551, 1717, 2097, 2070, 1810, 1809, 510, 2808, 641, 3525, 1633, 641, 1718, 2973, 2071, 1809, 33, 3242, 1630]. al- [431]. al-ahjar [2071]. al-Andalus [1633]. al-Andalusi [1551]. al-Baitar [2070]. al-Biruni [1717]. Al-chemia [730]. al-Din [1991]. Al-Farabi [1630]. al-hajj [1718]. al-hakim [33]. al-Hassan [510]. al-hay [1991]. al-Haytham [728, 1809]. al-Hiyal [1005]. al-Hwarizmi [2973]. Al-Jabr [641]. al-Jazzar [2808]. al-Khazin [2877]. al-Khazini [1810]. Al-Khwarizmi [1631, 641]. al-Quhi [2877, 1809]. al-Rahman [1810]. al-Safa [1717]. al-Samh [2877]. al-Sihha [2070]. Al-Sijzi [3242, 3242]. al-Tifashi [3525]. al-Tusi [1991]. Ala [431]. Alabama [858]. Alain [2111]. Alamos [2125, 1409, 1502]. Alan [1904, 1751, 29, 3073, 990, 788, 1744, 817, 1961, 2212, 3484, 2266, 2015, 2345, 2280, 974, 2267, 808, 756, 1949, 205, 2276, 2129, 631, 1272, 1628, 17]. Albarracín [523]. Albert [160, 1617, 1649, 1351, 1454, 2178, 768, 2460, 831, 827, 1314, 2117, 2766, 2070, 822, 3351, 914, 1651, 2202, 1026, 1198, 1686, 768, 2460, 2117, 2766]. Albertini [566]. Alberto [1288, 352, 2750, 3093, 1601]. Alberts [439]. Albertus [1715, 2502, 1642]. Albinus [459, 459]. albores [2516]. Albrecht [1200, 1274, 1766, 3149, 350]. Albright [3168]. Albritton [422]. Albumasar [310]. Albury [506]. Alchemical [3502, 552, 1190]. Alchemie [1724, 1535]. alchemisches [1721]. Alchemist [2169, 2071]. Alchemistische [1724, 1640]. Alchemists [1997]. Alchemomedizinische [3531]. Alchemy [2663, 1380, 1527, 2810, 116, 2168, 1094, 1068, 2341]. alchimia [1101]. alchimie [823, 1729]. alchimique [824]. alchimista [1915]. alchimistes [823, 824]. Alchimy [1557]. Alchon [1294]. Alcoa [293]. Alder [3443]. Aldous [2692]. Aldrich [1544, 3073, 2993]. Aldrovandi [743]. Alec [1835, 1592]. Alejandro [2103]. Aleksey [2385]. Alessandra [2271]. Alessandro [3545, 743, 2893, 2421]. Alestalo [513]. Aletta [2995]. Alex [1049, 764]. Alexander [1603, 2120, 430, 1830, 1001, 2020, 2835, 3546, 431, 3392, 1516, 144, 1852, 365, 2185, 1829, 1170, 1163, 365, 3454, 1852]. Alexander-von-Humboldt-Kolloquiums [2185]. Alexanderson [1598]. Alexandria [430, 339, 1361, 550, 2419]. Alfarabi [1006]. Alfonsi [1995]. Alfonso [3053]. Alfred [1943, 2734, 341, 2858, 297, 1033, 1404, 2614, 2844, 803, 1943, 343, 766]. Alfredo [1284, 3444]. Algebra [1712, 3137, 2795, 3143, 1531]. algèbre [150, 1042, 349]. algebrica [2270]. Algeria [3495]. algorismus [1631]. Algra [2417]. Alice [2944, 927, 573]. Alien [2932]. Alier [239]. alimentos [1911]. Alioto [1700]. Alistair [2595, 2626]. Alker [2849]. alla [2980, 1002, 1912]. Allan [3472, 575, 583, 2042, 944, 2472, 2041, 1703]. Allard [1631]. Alle [2526, 2589, 1194]. Allen [3511, 1376, 2501]. Allgemeinen [163, 1314]. Allgemeiner [1531]. Alliances [3464]. Allianz [1841]. Allied [110, 1053, 3559, 1952]. Allied-German [1952]. Allies [967]. Allison [2660]. Alloys [25]. Alltagspsychologie [1322]. Almaden [3444]. Almagest [1360]. Almanac [2073, 2425]. Almeida [1125]. Almost [2986]. Alnoor [2246].

Alois [475, 297]. Alok [1551]. Alonso [186]. Alpha [1256, 2544]. Alsace [1268]. Alsted [1375]. Altarpiece [444]. Alter [230]. Alteration [3507]. Alternating [1234]. Alternative [2030]. Alternativen [663]. alteste [2973]. ältesten [1120]. Althoff [1463]. Althusser [686]. Altmeister [755]. Altruism [1672]. Alvar [1654]. Alvarez [2408, 2157, 2933]. Alvaro [752, 2824]. Alvin [2205]. Always [2211]. Alwyn [225]. Alwyne [119]. Am [2202, 2281, 1398, 2378, 2723, 2046]. Amasiaci [1014]. Amateur [1391]. Ambiguous [1827]. Ambix. [116]. Ambrose [1887]. Ambrosius [2901]. Amedeo [2756, 2756]. America [2021, 2799, 3477, 1956, 3410, 3364, 2156, 3326, 62, 2193, 3009, 2030, 1219, 484, 817, 3156, 2470, 2467, 3098, 2234, 1352, 3554, 1541, 942, 177, 2609, 3363, 2459, 3286, 412, 1175, 3161, 2612, 3237, 225, 1203, 1509, 1750, 2802, 477, 1233, 1086, 1858, 844, 2479, 3206, 1983, 1786, 804, 1947, 612, 217, 805, 3113, 1970, 321, 396, 3160, 3453, 767, 1666, 474, 2389, 1561, 64, 1148, 3293, 2061, 1577, 3289, 1326, 1140, 1231, 772, 2466, 2823, 2543, 2255, 668, 2113, 2037, 2298, 1391, 2126, 3550, 2022, 1648, 54, 2383, 1843]. American [138, 2695, 1951, 498, 1128, 2468, 291, 2384, 2611, 3414, 52, 3274, 716, 3451, 1617, 3267, 2114, 2100, 66, 3424, 3378, 1382, 3481, 781, 465, 3108, 855, 885, 947, 1954, 70, 2454, 2020, 674, 591, 2759, 2169, 2781, 2993, 2282, 1838, 1591, 3382, 2377, 484, 1400, 573, 2936, 1414, 2780, 1038, 178, 1574, 3190, 3578, 878, 3365, 2845, 952, 1204, 1538, 376, 2210, 1585, 2186, 2299, 912, 383, 292, 126, 1497, 53, 681, 1494, 1665, 69, 695, 303, 1598, 3513, 3362, 3377, 722, 1619, 3584, 1067, 21, 3452, 1263, 3162, 815, 2720, 3347, 86, 274, 225, 1590, 1039]. American [1979, 3351, 1763, 2844, 246, 375, 73, 3264, 2929, 3480, 592, 698, 2758, 3366, 1578, 1769, 1329, 2041, 1318, 2992, 779, 1528, 1696, 3307, 3027, 2224, 405, 1693, 1432, 2226, 611, 2370, 629, 1807, 2192, 2934, 2367, 1668, 2576, 1689, 1416]. americana [2158, 2157]. americanas [2158, 746]. Americanists [591]. Americans [2193, 2136, 537]. Amerika [1829]. ameriki [887]. Amérique [718]. Ameriques [1649]. Ames [1924]. Amico [1102]. Amiens [1916]. Aminoff [1745]. Amirdovlat [1014]. Ammonius [1091]. among [967, 684, 702, 1697, 1940]. Amos [3402]. Amosa [1655]. Amouretti [2324]. Ampere [2717]. Amrine [2752]. Amsterdam [3061]. Amsterdamska [392]. Amy [3410, 2102]. Ana [2983]. Anaclastics [219]. Analisi [2270]. Analogien [940]. Analogy [1764]. analyse [3062]. Analysen [2131]. Analyses [3075]. Analysis [1325, 1205, 900, 878, 2196, 311, 3329, 3120, 2480, 2551]. Analyst [1386]. Analytical [2906]. Analytics [1001, 2649, 1988]. analytique [1296, 2906, 349]. analytische [2984]. Anand [3473]. Ananieva [1245]. Anaritiu [2075]. Anaritius [2075]. anarquismo [2824]. Anastasios [687, 2960]. anastatica [2901]. Anatol [654]. Anatomia [2346, 1199]. Anatomical [2510, 1539, 3052]. Anatomist [185]. Anatomy [2346, 3399, 3392, 1784, 1394, 459]. Ancestry [3450]. Ancien [751, 2747]. ancienne [820]. Ancient [544, 2240, 633, 1461, 723, 2732, 2156, 2969, 905, 1712, 1357, 1363, 1364, 1687, 1075, 28, 3333, 249, 3328, 1625, 1987, 32, 531, 1545, 1895, 373, 3519, 819, 2329, 2579, 990, 24, 2886, 2887, 3521]. Ancients [3052]. Andalus [1633]. Andalusi [1551, 2334]. anderen [758]. Anders [225]. Andersen [515]. Anderson [2283, 3372, 1802, 1402, 933]. Andersson [2390]. Andre [2032, 2438, 2519, 763, 2132, 1631, 1811, 2008, 161, 35, 1032, 245, 2717]. Andre-Marie [2717]. Andrea [925, 2586, 346]. Andreas [2506, 3385, 1778, 1658, 1766, 755, 2257]. Andreas-Holger [1658]. Andree [444]. Andrees [1442]. Andreu [3386]. Andrew [2346, 241, 754, 3543, 1066, 1457, 3007, 855, 3358, 353, 2837, 1978, 2894, 628, 2412, 840, 1906, 1870, 1704, 2735, 1725, 1677, 1743, 1285, 2550, 2697, 3052, 1895, 1207, 275]. Andrewes [3252]. Andrews [1947]. Androgynie [3427]. Anesthesia [1087]. Anfange [1658, 1283]. Anfangen [1587, 712]. Ang [1712]. Angel [95]. Angela [371, 230]. Angelical [1388]. Angelika [2585]. Angelo [973]. Angleterre [2514]. Anglican [46]. Anglicus [1554, 1012]. Anglo [2611, 1382, 855, 1418, 1309, 2279, 3307, 2974, 2074, 1992]. Anglo-Australian [1418]. Anglo-Indian [1309, 2279]. Anglo-Norman [2974, 2074]. Anglo-Saxon [1992]. anglosajón [134]. Anguel [81]. Angus [1143]. anhand [1322]. Anhang [461]. Anima [1092]. Animadversions [1077]. Animal [960, 2937, 350, 2364, 2029, 1827, 603, 2328, 490]. animalibus [1913]. Animals [805, 1913, 1472, 2069, 1459]. Animate [3148]. Anita [1936]. Anke [3586]. Ankupfung [2238]. Anlass [2185]. Ann [1388, 2751, 3278, 1499, 2344, 3051, 1665, 1820, 2664, 3227, 2606, 2648]. Anna [768, 2117, 3532, 1722, 381]. Annales [513, 2446]. Annali [521]. Annals [895, 103]. Anne [2801, 2922, 110, 1466, 3524, 1412, 2655, 2278, 3328, 1337, 1933, 3344, 2776, 2351, 1361, 2326, 212, 2162, 1222, 2814]. Anne-Marie [1466, 2326, 212]. Annemarie [1622]. Annette [2095, 2766]. Annibale [2084]. Annick [2436]. Annie [2743, 834, 2995]. Annotated [2942, 835, 1776, 2885, 2787, 300, 885, 3488, 1518, 2322, 2178, 1502, 795, 2049, 1249, 604, 603, 1963, 1428, 3393, 1248, 1780, 2706, 3076, 491, 698, 1426, 699]. Annotations [83, 428]. Annual [3303, 1614, 221, 2311, 1972, 1869, 896, 2627, 1258, 2947, 3017, 500]. annum [1246]. Annus [3330]. ano [1374]. anonymer [822]. Ansatze [1313]. Ansell [3483]. Answer [1866]. Ant [1672]. Antagonist [185]. ante [1246]. Antebellum [1970, 2022]. antecedentes [868]. Antecedents [3300, 3247, 729]. Anthology [1943, 3513, 2609, 2606]. Anthony [3409, 1111, 2579, 714, 3259, 326, 2738, 1752, 923, 1687, 1700, 532, 1372, 1815, 2159, 3519]. Anthropological [3388, 684, 65]. Anthropologie [1937, 1129, 1982, 1982]. Anthropologist [574, 2961, 777, 3470]. Anthropologists [383, 1338]. Anthropology [1490, 2880, 976, 3585, 1849, 2876, 2014, 2045, 1338, 2408, 1849, 120, 1850]. Anti [2227, 1216, 1854]. Anti-Newtonian [2227]. Anti-Realist [1216]. Anti-Science [1854]. Antibody [3093]. antica [1365]. Antiguo [3154]. antiguos [1633]. Antike [3049, 1993, 3327, 2415]. antiken [3241, 341].

Antimonliteratur [394]. Antipodeans [2228]. Antique [3163, 1004]. antiquiora [1473]. Antiquite [2324, 3427, 1093, 210]. Antiquity [1196, 917, 710, 88, 27, 1907, 1590, 314, 1003, 548, 2730, 1627]. Antoine [2596, 2441, 3146, 1753, 1026, 41, 3255]. Antoine-Laurent [2441, 41]. Antoinette [1568]. Anton [1747, 2992, 1225]. Antoni [2896, 1331, 835, 2512, 3455, 283, 1917, 3523, 2511]. Antonino [2098, 2259]. Antonio [2434, 2587, 321, 2027, 2080, 2741, 2341]. Antonius [2077, 2077]. Antry [731]. Anxiety [2041]. Any [3064]. AO [3256]. Aobo [1625]. Apariencia [1197]. Ape [368]. Ape-Man [368]. Apelt [1393]. Aphrodisias [1001]. aplicada [1723]. Apokin [910]. Apollo [583, 2130, 2388, 2387]. Apollonius [1113]. apologétique [1571]. Apologia [3208]. Aporetic [2497]. Apothekenwesens [1142]. Apotheker [3360]. Apparatus [3158, 3393]. apparence [687, 821]. apparences [2508]. Appel [2175]. Appendix [3199, 3219]. Apple [2936, 1175]. Applied [1968, 596]. apport [931]. Appreciations [3565]. Apprentice [3426]. Apprentices [1482]. Apprenticeships [493]. Approach [2471, 3130, 1243, 1901, 790, 3329, 1420]. Approaches [220, 168]. Approaching [2800]. Appropriating [193]. apres [2567]. apres-Darwin [2567]. April [2542, 3312, 2369, 1687, 2624, 614]. Aquilecchia [2338, 2895]. Aquinas [1007]. Arab [3525]. arabe [642, 914]. arabes [150]. Arabic [1913, 918, 339, 2151, 2245, 2244, 510, 2808, 3330, 1006, 2398, 2481, 510]. Arabic-Latin [1913]. Arabick [2006]. Arabisch [510, 1360]. Arabisch-Islamischen [510]. arabisch-mittelalterliche [1360]. arabischen [2070]. arabischer [822]. Arabo [431]. Arabo-Byzantine [431]. Arabs [905]. Aragon [2250]. aragonés [154]. Aramis [2642]. Arana [1197]. Arboretum [1607, 2284]. Arc. [1146]. arcano [2272]. Archaeological [1242, 935, 630, 1041, 1243]. Archaeology [694, 591, 3508, 2968, 2728, 1549, 2929, 3471, 3506, 3386, 2232]. Archaic [1987]. Archeologique [825]. Archetype [1254]. Archibald [1822]. Archimedes [702]. Archipelago [330]. Architectural [734, 2080, 2792]. Architecture [3286, 2522, 2080]. Archiv [518]. Archival [2377]. Archive [112]. Archives [87, 696, 209, 606, 888, 213, 503, 3526, 119]. Archivo [525]. Arctic [1699]. Arduino [2097]. Area [1502, 888]. Arend [1766]. Arens [462]. Argaman [249]. Argentina [2686]. Argentine [1611]. Argonne [3171]. Arguelles [1197]. Argument [1891, 1169, 808]. Argumentationstheorie [2153]. Arguments [283]. Arias [2158]. Arien [1707]. Ariew [2727, 1531, 1282]. Aristocratic [2710]. Aristotele [1556]. Aristoteles [3246, 1463, 439]. Aristotelian [2571, 1179, 1006]. Aristotelians [1182, 2087]. aristotelisch [439]. aristotelisch-scholastische [439]. Aristotelischen [3246, 2742]. Aristotelismus [2742]. Aristotilis [3431]. Aristotle [342, 1001, 1091, 1547, 1913, 918, 1988, 724, 1269, 147, 1358, 3130, 2649, 1092, 2065, 649, 343, 1818, 547, 2418, 1552, 2498, 1270, 2497, 546, 2244]. Arithmetic [1712, 2501]. arithmetica [2089]. arithmetice [1096].

Arithmetick [3143]. arithmétique [150, 2580]. Arizona [3261]. Arleen [2184]. Armagh [906]. Armand [255]. Armando [449]. Armelle [2888]. Armin [1013, 1046]. armonia [811]. Arms [199, 968, 3443, 225]. Army [2826, 225, 1256]. Arnal [1084]. Arnald [438]. Arnaldi [438]. Arnauld [836, 1020, 737]. Arno [1895]. Arnold [2284, 107, 911, 943, 332, 2279, 1458]. Arnold-Forster [1458]. Aron [2032]. Aronowitz [3182]. Arquiola [1828]. arquitectura [921]. Arrheniu [475]. Arrhenius [475, 3080, 2836]. Arrizabalaga [3337, 868]. Arrowsmith [1448]. Arroyo [2825]. Ars [2563, 1916]. Art [287, 575, 1886, 550, 574, 1464, 1436, 1645, 2690, 2964, 903, 2339, 3343, 1293, 2819, 1919, 2976, 2524, 246, 349, 3523, 870, 574, 3160, 3410]. arte [152]. Artes [1993]. Artesfaculteit [361]. Artful [2009]. Arthmetique [2010]. Arthur [362, 121, 3078, 1893, 693, 1035, 3328, 2990, 1950, 697]. arti [1267]. Articles [1536, 225]. Artifact [1243]. Artifacts [2410]. Artificial [2912, 3124, 71, 975, 3188]. artis [1096]. Artist [574, 18, 1787]. Arts [835, 3453, 3368, 917, 2132, 1492, 406, 1891]. Artsen [2693]. Artsen-Onderlinge [2693]. Artur [1852]. Arturo [2408]. Artzney [453]. Artzney- [453]. Arznei [2081, 1266]. Arzneibuch [1638]. Arzneimittelforschung [3360]. Arzneimitteltherapie [2884]. Arzneimittelversorgung [3150]. Arzt [2523, 1930, 1813]. Arzte [2382]. Arztelexikon [2415]. Arztes [461]. arztliche [1481]. Asa [178]. Ascent [3407]. Asceticism [1364]. Asclepiades [1090]. Asclepio [523]. Asger [531]. Ashmore [882]. Asian [1452, 2145]. Ask [2603]. Askey [767]. Aspecten [228]. Aspectos [1723]. Aspects [488, 3378, 2036, 481, 1919]. Aspekt [2507]. Aspekte [2742]. Aspekten [1313]. Aspirations [3315]. Aspiring [3502]. Aspray [1319]. Assassination [2740]. Assault [3380]. Assembly [2125]. Assessment [2641]. Assessments [2173]. Assimilation [3309]. Assis [2519]. assises [1063]. Assisi [434]. Associated [968]. Association [1607, 506, 2034]. Associations [1309]. Assyria [3163]. Assyrian [3163]. Astrain [3445]. astralable [2248]. astres [2079]. Astro [3253]. Astro-Technik [3253]. Astrolabe [2248]. astrolabes [1814, 3423]. Astrologer [1563]. Astrologi [254]. astrologia [1101]. astrologica [1246]. astrologie [3189, 2079]. astrologue [2333]. Astrology [1641, 2398, 2426, 556, 2078, 2152, 1642, 338, 3432, 2344]. Astronome [2166]. Astronomen [2016, 3441]. Astronomer [172, 1298, 562]. Astronomers [602, 2579]. Astronomia [1819, 1365]. astronomiae [919, 1642]. Astronomica [354, 1246]. astronomicae [2903]. Astronomical [378, 600, 247, 1462, 431, 31, 19]. Astronomie [145, 2254, 3253, 40, 256, 2515]. Astronomies [3243]. astronomique [2581]. astronomische [2336]. astronomischen [3253]. astronomisches [1013]. astronomo [2355]. Astronomy [2684, 906, 3050, 2529, 2787, 3107, 24, 905, 610, 2072, 2151, 431, 5, 2968, 1991, 2245, 1586, 739, 2638, 892, 905, 2329, 1548, 1986, 3007, 2082, 158, 3514, 2599, 2729, 1650, 3323, 2731, 114, 1137, 2732, 905, 3228, 533, 1909]. astrophile [2079]. Astrophysics [739]. Asylum [3235, 1148]. Asymmetries [188]. atelier [825]. Athena [2524]. Atherton [1117]. Atilla [1005]. Atkins [632]. Atkinson [1087]. Atlanta [2627]. Atlantic [2007, 3221]. Atlantis [3172]. Atlas [3442, 793]. Atmospheric [888]. Atmungstheorie [3570]. Atoll [2040]. Atom [189, 3413, 225, 1867, 672, 2041]. Atombombe [1766]. atomi [3139, 860]. Atomic [2206, 3563, 3477, 582, 3320, 968, 2924, 1409, 1502, 56, 1768, 2040, 2213, 225, 966, 1584, 2925, 968, 1529, 967, 3170, 225, 2832]. atomica [2756]. atomismo [1925]. Atomphysiker [2378]. Atoms [2246, 582, 2118, 1383]. Atomtheorien [2332]. Atran [976]. Attack [3183]. Attempt [1257, 3228]. Attention [3130]. atti [2445, 327]. Attitudes [5, 53, 434]. attossicato [3268]. Attraction [1534]. Attractions [2817]. Attributed [552]. attualità [124]. Atul [200]. Audience [3537]. Audrey [87, 1541]. Audubon [3451, 1666]. Aufklärung [461, 1150, 2827, 40]. augmentée [914]. August [3302, 160, 3311, 2946, 223, 2955, 2816, 1841, 2954, 2624, 614, 1890, 3546]. Auguste [2906, 2667, 1026, 1939]. Augustin [2102, 1040]. Augustin-Louis [2102]. Augustine [807]. Augustus [1041]. AURA [3173]. Ausbau [325]. Ausbreitung [442]. Ausdifferenzierung [1115]. Ausgewahlte [2843, 929]. Auspices [835, 968]. Austen [1575]. Austin [1294]. Austoker [679]. Australasian [506]. Australia [580, 571, 3072, 1133, 385, 569, 1942]. Australian [1981, 3356, 1418, 684, 3072, 2729, 506, 1085]. Australis [3123]. australischen [3551]. Austrian [1332]. Auswahl [394]. Auswahl-Bibliographie [394]. Authorities [2969]. Authority [1851, 1053, 3293, 1831, 792, 1263, 3223, 227, 3037, 445]. Authorizing [1099]. Authors [395]. Authorship [2563, 2162]. Autobiographical [465, 379]. Autobiographische [365]. autobiographischen [2123]. Autobiography [197, 2318, 2854, 187]. autographe [1466]. Automatic [445]. Automatisierung [1599]. Automobile [3300, 2037]. Autopsie [2750]. autor [154]. Autoren [2920]. autoritaren [2843]. Autour [2658]. autres [2247, 2508, 2528]. Autumn [2233]. aux [1847, 1026, 1361, 2188, 423]. avant [2326]. Avelino [2249]. Aveni [2579, 532, 3519]. aventura [2103]. Averroe [1095]. Averroes [1276, 1095]. Avery [3559]. Aviation [3463, 2773, 3006, 3381]. Avicenna [2331, 727, 1910, 2331]. Avogadro [2756, 2518]. Avoiding [2404]. Awarded [3199]. Awareness [125]. Aweigh [1597]. Axel [3413, 288, 2413]. Axelrod [1904]. axiomatischer [1916]. Axtell [3124]. Aydin [2483]. Ayers [3387]. azafea [432]. Azarquiel [432]. Azogue [3444]. Aztec [993].

B [1147, 580, 1395, 2021, 622, 2240, 1673, 1996, 29, 1007, 909, 435, 3473, 1956, 2751, 3108, 1091, 2898, 1491, 3508, 2981, 67, 1936, 1761, 339, 2501, 1418, 849, 2197, 334, 1947, 145, 3580, 511, 1087, 2369, 3176, 441, 1539, 1452, 2145, 3233, 1120, 2694, 393, 2576, 1140, 1405, 2609, 854, 2540, 356, 151, 957, 2012, 766, 1675, 2702, 1275, 156, 531, 966, 1391, 1039, 636, 2126, 1198, 1207, 641, 2003, 2878, 2116, 1006, 779, 3199]. B. [1791, 1862]. B.C [3131, 2412, 425, 1797]. Baasner [40]. Baatz [816]. Babbage [2480, 617]. Babcock [383]. Babies [2695, 947, 2761]. Babylonian [3520, 1462, 610]. babylonischen [998]. Bache [2020, 1163]. Bachelard [1065]. Bachir [1124]. Bachon [1557]. Back [2, 93, 215, 306, 399, 402, 495, 608, 700, 797, 800, 890, 983, 1071, 1156, 1160, 1251, 1340, 1430, 1521, 1525, 1609, 1692, 1783, 1880, 1884, 1965, 2051, 2134, 2218, 2222, 2305, 2394, 2477, 2557, 2561, 2628, 2708, 2790, 2862, 2866, 2948, 3022, 3110, 3193, 3200, 3201, 3304, 3396, 3493, 3592, 1078, 2770]. Background [1833, 1187, 2358]. Backyard [225]. Bacon [1371, 3528, 1557, 924, 1775, 436, 3170, 3135, 1652, 2164, 2740, 3335, 2003, 3250, 3435, 3225]. Baconian [2104]. Baconianism [1536]. Baconschen [2046]. Bacterial [958, 2310]. Bacteriological [1078]. Bacteriology [469]. Bad [3085, 1959, 1619]. Badash [2607]. Baden [2184, 46]. Baer [1747]. Baeumker [1634]. Bagheri [3242]. Bahn [3506]. Baigrie [2964]. Bailes [2288]. Bailey [2706, 3371, 631]. Bailin [2524]. Bains [3186]. Baird [1667, 2670]. Baitar [2070]. Bajja [2244]. Bakelites [2703]. Baker [2582, 2201, 3585, 2775]. Bakken [1519]. Balance [1081]. Balancing [3176]. Baldini [2000]. Baldwin [2938]. Balfour [2711]. Balibar [1847]. Ball [3383, 1957, 225]. Ballard [1301]. Ballester [868, 2249]. Balloon [387]. Ballooning [474]. Balogh [1329]. Balpe [2182]. Baltimore [3504]. Balzac [2361]. Balzer [577]. Bamber [3127]. Bamberg [348]. Bambergensis [1638]. Bamberger [348]. Bamboo [2035]. Ban [2607]. Band [2356]. Bandelier [2830, 2830]. Bang [2120, 2769]. Bank [1405]. Banks [1740, 2598, 393]. Banne [3060]. Banner [363]. Banu [1005, 2877]. Bar [2389]. Barbara [1028, 121, 1382, 2108, 1596, 3348, 2813, 383, 1094, 3037, 3102, 872, 481, 2009, 2706, 1926, 2943]. Barber [884, 1478]. Barberos [3445]. Barbieri [3350, 1779]. Barbour [622, 2197]. Barcelona [2825]. Barcia [2656]. Bardinet [820, 2149]. Barfoot [2603]. Bargaining [2686, 1596]. Barger [1491]. Barigazzi [1002]. Barkan [1594]. Barker [1906, 1282, 2727]. Barlow [1395]. Barnaba [2355]. Barnabas [641]. Barnard [2372]. Barnes [1001, 2551, 316, 1988, 2109]. Barnett [3320]. Barocke [3060]. Baron [3483, 159]. Baroncini [2001]. Baroque [2007]. Barr [3505]. Barreau [3411]. Barrett [1395]. Barriers [1771]. Barroso [2354]. Barrotta [2277]. Barrow [3451, 1257, 2985, 741]. Barrow-Green [2985]. Barry [2691, 316, 3514, 2551]. Barsanti [2443]. Barsky [3086]. Bart [1853]. Barthelemy [1657, 729]. Bartholin [1822]. Bartholomaeus [1554]. Bartholomew [3548, 3548, 421]. Barton [1134, 2213, 1804]. Bartram [1483]. Baruch [2366]. Bas [1421, 880]. Basalla [243]. Bascom [1047]. Base [308]. Based [1301]. Basement [3359]. Bases [1778]. Basic [3561]. Basil [1477]. Basim [510]. Basrian [2246]. Bassi [2978, 613, 1342]. Basso [2963]. Bastian [2382]. Bastin [1392]. Bataafsch [994]. Batchelor [3462]. Batchen [3275]. Bate [3244, 1936]. Bateman [2774]. Batens [317]. Bates [1596]. Bath [2426]. Bathurst [1081, 1081]. Batschelet [885]. Battail [320]. Batten [756]. Battigelli [3532]. Battista [2744, 1102]. Battle [1820]. Battling [612]. Baudhyayana [28]. Baumann [3360]. Baumeister [1398]. Baumer [3360, 1222]. Baustatik [1587]. Bauten [3367]. Bautista [3442].

Baxter [2759]. Bayerischen [1337, 3441]. Bayerisches [3256]. Bayern [3256]. Bayertz [196]. Bayes [1422]. Bayesian [790, 1422]. Bayle [1820]. Baynton [2845]. Be [37, 2662, 3064, 3164, 2202, 2693]. Beagle [366]. Beagon [1714]. Beamtimes [192]. Beasts [1370]. Beat [2381, 3009, 3183]. Beate [2978]. Beatrice [2526]. Beatrix [3360]. Beattie [3384]. Beatty [418]. Beaucamp [2353]. Beauchamp [2966]. Beaujouan [1633]. Beaulieu [255]. Beaune [349]. Beauperthuy [184]. Beautiful [1433]. Beauty [2226, 2857, 3227]. Became [3471]. Becher [3179, 3179]. Bechler [1290]. Bechtel [1876]. Bechthold [1777]. Beck [768, 2117]. Becker [2885, 471]. Beckova [1655]. Becomes [3496]. Becoming [2414]. Becquemont [3556, 1570]. Bedani [454]. Beddoes [2096]. Bede [2870]. Bedeutung [1322, 2336]. Bedi [606]. Bedini [567, 1019]. Bedlam [3358]. Bedroom [2038]. Bedside [1050, 3291, 3357]. Beeckman [1289]. Beef [2212]. Been [1796]. Beer [2719]. Before [583, 2467, 2137, 2250, 2470, 3004, 1507, 2566, 95, 2375, 2612, 3071, 3497, 52, 74, 595, 1849, 2735, 635, 741, 1858]. Beginn [2281]. Beginning [25, 13]. Beginnings [1308, 1433, 285, 3008, 1797, 3578]. Begluckende [1830]. Begriffs [1322]. Beguet [1312]. Behavior [194, 3390]. Behavioral [2549, 121]. Behaviorism [285]. Behaviour [2286]. Behaviourism [2116]. behind [1581, 3383, 1144, 2856]. Beigun [1926]. Being [2524, 61]. Beispiel [2914, 1856, 2378]. Beispielwerk [1916]. Beissinger [276]. Beisswanger [3150]. Beitrag [1923, 736, 1274, 266, 549, 1640, 2070, 3570, 2359, 3091]. Beiträge [81]. Bekanntnisse [365]. Bela [1581]. Belfast [986]. Belgian [2257]. Belgie [761]. Belgique [2247]. Belhoste [2182]. Belief [753, 3102]. Beliefs [2817, 1771, 2724]. Believing [1315]. Bell [3454, 1511, 231, 2721]. Bellarmine [1016]. Bellissima [3129]. Belloli [346]. Bellone [811]. Belloni [771]. Belna [2911]. Belousov [654]. Beltran [2741, 2434]. Ben [686, 1057, 920, 1917]. Ben-David [1057]. Benavides [2158]. Bench [3291]. Benchimol [2027]. Bendegem [317]. Bender [619, 13]. Beneath [1028]. Benedetti [2893]. Benedetto [157]. Benedict [201, 1200, 2178, 1359, 2413, 2829]. Benefiel [2540]. Benfey [1134, 1317, 995]. Bengt [279]. Benivieni [2587]. Benjafield [2575]. Benjamin [2684, 2839, 1371, 1680, 2748, 1323, 199, 1768, 1307, 2706, 2127, 2943, 568, 2622, 934]. Bennett [843, 2651, 1802, 906, 3322, 3076]. Benninga [1135]. Benoit [3062]. Bensaude [2879]. Bensaude-Vincent [2879]. Benson [3536, 1318]. Bentley [86, 1600]. Benz [748]. Benzenhofer [1937, 3134, 730]. Beon [3459]. Berdoulay [388]. Berengario [1188]. Beretta [2596, 1801, 2274]. Berg [2847]. Bergakademie [2185]. Bergbaukunde [1120]. Bergdolt [1813]. Berge [1499, 358]. Bergevin [1351]. Berggren [531, 2731]. Bergman [67]. Bergon [1917]. Bericht [2356]. Berichte [517]. Bering [260]. Berkel [1646, 3279]. Berkeley [1483, 1386, 176, 457, 1131, 456, 3211, 176, 1117, 1736]. Berlin [2984, 3360, 712, 2766, 2923, 755, 2648, 2648]. Berliner [1830]. Berlinische [89]. Berlins [2914]. Berman [1184]. Bermannus [827]. Bernadette [2879].

Bernandino [505]. Bernard [2363, 2010, 928, 3583, 993, 2885, 2275, 1176, 2242, 2748, 884, 1399, 463, 1486, 3413, 549, 1109, 1297, 2437, 459, 920, 1663, 1729, 3069, 2319, 2983, 379, 2928]. Bernard-Pierre [3583]. Bernardo [1912]. Bernd [1113, 3179]. Berndt [2380]. Bernette [762]. Bernhard [2332, 267, 1215, 348]. Bernhardt [451, 925]. Bernier [1559, 1559, 423]. Bernoulli [2979, 3534, 2438, 2008, 2979]. Bernstein [2461]. beroep [1733, 761]. Berrios [3046, 2779]. Bert [516, 1281, 368, 127]. Bertelli [1555]. Berthold [2359, 2252]. Bertier [2227]. Bertoloni [1656]. Bertram [2843]. Bertrand [967]. Berucksichtigung [3253]. Berzelius [1739, 320]. Beschlüsse [1123]. beschreibende [162]. Beschreibung [1313]. besichtigt [1304]. bespiegeling [3068]. Bessel [2185, 2185]. Bessey [1574]. Besso [1697]. Best [803, 3525]. Bestiary [1370]. Betancourt [1040]. Betrachtungen [174]. Better [3266, 225]. Bettina [1856]. Betty [3389, 1380, 2268]. Bettyann [2848]. Between [2399, 447, 141, 1061, 1594, 623, 1081, 3310, 1200, 1647, 2060, 238, 2795, 2765, 1669, 1758, 2319, 65, 1015, 3353, 676, 779, 2650, 2452, 3291]. Beukers [996]. Beverley [2087]. Beverly [459]. Bevilacqua [327]. bevordering [994]. bewegende [174]. Bewegung [439, 374]. Beyerlein [2359]. Beyond [1382, 2951, 1438, 2672, 1241, 189, 1586]. Beziehungen [475]. Bezilla [135]. bi [2732]. Bialas [354]. Bianchi [643, 2901]. Bianchini [1013]. Biannual [506]. Bibel [63]. Bible [1536, 3379, 3316, 1016]. Biblical [2236, 249, 1989, 1206]. Bibliografia [1779]. bibliograficos [1690]. Bibliographia [2491, 2491, 2491]. Bibliographic [2622, 2178, 1778]. Bibliographical [393, 3191]. Bibliographie [3298, 2987, 3367, 2013, 2353, 162, 394, 255, 210, 3542, 1069]. Bibliographie/Charles [2987]. Bibliographies [3017]. bibliographische [90]. bibliographischer [51]. Bibliography [2788, 605, 2392, 1776, 885, 1518, 1502, 1, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 83, 3393, 1133, 1607, 3018, 3017, 3590, 698, 1877, 794, 2942, 2787, 3488, 2322, 2049, 1249, 604, 603, 1963, 2257, 1781, 1067, 1780, 2706, 1426, 2987, 602]. Bibliometria [2491]. Bibliometrische [3367]. Biblioteca [1690, 1920, 1556]. Bibliotheca [2596, 1688, 1428]. Bibliothek [1923]. BibliothÈQue [1649, 1466, 2333]. Bicentenary [232]. Bicentennial [2399, 20, 231]. Bichat [1032]. bichos [2027]. Bicycles [2703]. Bienkowski [1655, 512]. Biermann [365, 1830, 1029]. Biewer [2502]. Biezunski [2028]. Big [2684, 2120, 2769, 1763, 1328]. Bigalke [673]. Bigelow [1152]. Biggemann [1375]. Biggs [3286]. Bignon [837]. Bijdragen [127]. Bijker [2703]. Bikini [2040]. Bild [3367, 1808]. Bilder [1941]. Bildung [1856, 1303]. Bilharzia [1229]. Billings [2021]. Billy [3153]. Bilstein [583]. Bimber [2641]. bin [1005]. Bindocci [1963]. Bingen [434, 2889, 2889]. Bio [51, 3590, 3191]. Bio- [51]. Bio-Bibliographical [3191]. Bio-Bibliography [3590]. Biobibliographic [1878]. Biobibliographies [1998]. Biobibliography [491]. Biochemical [3590, 1031]. biochemischen [3360].

Biochemist [62, 2403]. Biochemistry [1706, 698]. Biodiversity [3466]. Bioethics [1050]. Biogeography [1434]. Biografie [2335, 3279]. Biographical [1904, 846, 1617, 750, 393, 1133, 395, 2706, 225, 2987, 2048, 2943, 3108, 695]. Biographie [2353, 659, 3568, 3441]. Biographien [3118]. Biographies [3018, 3488]. biographique [2132, 212]. biographisch [90]. biographisch-bibliographische [90]. biographischem [2987]. biographischen [1313]. Biography [1581, 1859, 3271, 2108, 1045, 1049, 1106, 863, 1853, 334, 2841, 1285, 225, 373, 954, 1939, 1323, 848, 770, 2691, 359, 2392, 662, 3269, 464, 2430, 2533, 794]. Biologen [90, 1410]. Biological [3096, 115, 2273, 2441, 968, 1709, 2771, 2210, 1504, 323, 2972, 2711, 1139, 590, 589]. Biologie [851, 3091, 423, 3103, 1222]. Biologieunterricht [2647]. biologique [1023]. Biologisch [3367]. Biologisch-Medizinischen [3367]. biologische [940]. Biologisierung [758]. Biologist [2024, 2202]. Biology [3104, 411, 1128, 2309, 1594, 3464, 2917, 3389, 2471, 612, 2840, 2902, 2724, 808, 3450, 704, 2972, 1435, 2243, 3026, 3221, 2793, 3104, 118, 281, 1606, 76, 2608, 118, 3389, 1221, 2916, 864, 1593, 2838, 625, 1592, 2451, 1962, 1318, 3465, 3104, 2364, 2450]. Biomedical [1891, 2846, 2392, 805, 794]. Biomedicine [2782, 3580]. Biophysics [698]. Biosphere [3518]. Biotechnics [1265]. Biotechnology [715, 590, 1135, 2209]. Birds [3451, 1666]. Birefringent [1822]. Birke [3505]. Birken [171]. Birkhan [1640]. Birnie [2521]. Birth [2931, 3314, 1035, 3119, 956, 639, 1863, 3459]. Births [1726]. Biruni [1717]. bis [325, 160, 2356, 585, 3531, 2411, 3375, 989, 3327, 730, 712, 2415, 3385, 1941, 2185, 175, 1855, 1283, 1587, 812]. Bishop [731, 457]. Bitbol [2743, 2463, 834, 3002]. Bitbol-Hesperies [2743]. Bithynia [1090]. Bjerknes [193]. Bjork [1862]. Bkouche [1484]. Black [2695, 3334, 2321, 2891, 593, 349, 2296, 3231]. Blackmore [2689]. Blacks [289, 886]. Blackwell [2063, 1922, 1016]. Blædel [57]. Blaine [3101, 2783]. Blair [3051]. Blaise [1804, 2431, 357]. Blakely [3359]. Blanc [1738]. Blanckaert [1875]. Blanpied [2542]. Blay [522, 1296, 2004, 3412]. Bleaching [1428]. Blessings [2061]. Blockbuch [348]. Blocks [3374]. Blood [3224, 3289, 1439]. Bloor [2551]. Blower [1958]. Blue [249, 3296]. Blueprint [3573]. Blues [3296]. Blum [473, 1665]. Blumberg [672]. Blumenbach [3028]. Blumenfeld [639]. Blumenfeld-Kosinski [639]. Blustein [1578]. Blut [196]. Blutezeit [3338]. Boag [1217]. Boas [1193, 2610]. Bob [2983]. Bobzien [1001]. Boccardi [2763]. Bocking [2930]. Bodanszky [1223]. Bodde [1187]. Bodet [824]. Bodies [2066, 2267, 2142]. Bodily [3579]. Bodin [3051]. Body [2494, 3482, 2782, 2777, 3399, 2737, 1539, 3230, 2985, 3491, 3485, 1575, 2025, 652, 2513, 3231, 2361, 2279, 572, 1468, 3392, 237]. Body-Mind [3485]. Boece [2580]. Boehlich [943]. Boehme [1285]. Boeuffle [427]. Bogdan [512]. Bohm [3498, 3167]. Bohme [2046, 360]. Bohr [3276, 78, 2834, 190, 860, 2377, 1216, 1408, 2291, 2118, 278, 3564, 949, 57, 1804].

Boissonnade [2906]. Bok [1853]. Bolletino [521]. Bolling [2195]. Bologna [566, 613, 1194]. Bolotin [3130]. Bolsheviks [3582]. Bolts [274]. Boltzmann [3352, 2689, 1130, 755, 2457, 1130]. Boltzmann-Henriette [2457]. Boltzmanns [755]. Bolzano [463, 2983, 1486]. Bomb [2206, 3563, 987, 968, 1502, 2540, 225, 1584, 1581, 2853, 967, 1409, 225]. Bombing [2383]. Bombs [2924, 225]. Bonah [3075]. Bonati [2271]. Bondeson [3048]. Bonduelle [2762]. Bone [2848]. Bones [3354, 3359]. Bonfioli [2271]. Bonjorn [1917]. Bonker [2585]. Bonker-Vallon [2585]. Bonner [1310, 2414]. Bonnet [166]. Bonnie [1578]. Bono [3249, 2429, 1102]. Bonta [2367, 1352]. Boockmann [1474, 354]. Book [248, 392, 57, 1626, 627, 528, 1706, 3346, 2726, 3089, 2251, 1490, 3361, 3104, 2684, 2931, 513, 3039, 1981, 140, 244, 161, 1336, 539, 971, 2363, 202, 2214, 2896, 1911, 3181, 25, 130, 709, 819, 2329, 2546, 2914, 3178, 3408, 2284, 2453, 2548, 2179, 2368, 525, 868, 1024, 2552, 1308, 30, 250, 84, 719, 1564, 2764, 3094, 601, 486, 2343, 236, 2180, 3019, 138, 298, 1147, 1232, 2148, 2695, 3317, 1894, 2433, 562, 740, 1563, 3445, 1043, 1213, 1484, 2018, 559, 687, 1008, 1874, 1891, 2348, 2659, 2743, 3339, 3527, 2509, 1413, 2086, 3475, 2661, 286, 1322]. Book [1857, 1951, 2059, 2843, 789, 3054, 2293, 2678, 1177, 1468, 3334, 201, 264, 1128, 2692, 1581, 2206, 2461, 2639, 715, 418, 1061, 2010, 2597, 3418, 1288, 2000, 3502, 1583, 3352, 1982, 488, 1316, 2237, 1603, 2570, 3249, 734, 2294, 2527, 3049, 2913, 1859, 3284, 330, 902, 818, 39, 2346, 2468, 327, 2647, 291, 1325, 2384, 2120, 3262, 2494, 3555, 132, 713, 3096, 1718, 1661, 2979, 1999, 2645, 2596, 3419, 281, 997, 2989, 3512, 882, 3255, 3545, 3460, 769, 2189, 1923, 2681, 3182, 124, 1313, 2531, 1017, 2383, 858, 2797, 197, 580, 2611, 2962, 3414, 1018, 2812]. Book [87, 1395, 2839, 3581, 2370, 1671, 82, 2289, 1135, 2427, 52, 752, 856, 1036, 1324, 1594, 1690, 960, 1904, 2069, 325, 848, 2021, 3042, 554, 1295, 1371, 2408, 870, 2032, 2604, 2799, 3274, 3537, 338, 430, 544, 1548, 1713, 1986, 2386, 3549, 50, 939, 869, 3520, 622, 1656, 1638, 345, 2725, 170, 587, 2879, 3447, 455, 105, 241, 754, 1028, 1115, 1481, 1660, 2177, 2353, 2754, 3043, 629, 1055, 1623, 1751, 3079, 906, 1662, 2240, 2677, 2853, 3272, 1331, 2434, 2314, 396, 567, 770, 2019, 2301, 2700, 2944, 3276, 736, 1606, 1673, 1708, 568, 328, 1048, 1795]. Book [3252, 487, 690, 362, 1306, 213, 1317, 2846, 1806, 3299, 115, 302, 846, 1483, 2788, 2464, 2549, 49, 367, 623, 1262, 3466, 2973, 3503, 716, 961, 437, 2075, 2150, 571, 2521, 851, 2691, 68, 2493, 1715, 148, 58, 517, 1210, 3367, 605, 2580, 16, 160, 233, 359, 1738, 1830, 3464, 3573, 388, 852, 1081, 1537, 1943, 2449, 2636, 3451, 2356, 2752, 3543, 2915, 1084, 1294, 1734, 1828, 1534, 1576, 2392, 78, 1132, 874, 1440, 1597, 1678, 3271, 1707, 2587, 2653, 2829, 3516, 1152, 778, 1851, 2287, 2801, 2834, 2920, 3563, 2848, 2824, 3267]. Book [1679, 2622, 2833, 3547, 3350, 3577, 3160, 3246, 482, 3479, 1374, 650, 1873, 321, 526, 1543, 3281, 2376, 535, 3140, 981, 3000, 2591, 1635, 2436, 3097, 2303, 2912, 37, 718, 1388, 2062, 2114, 2768, 940, 3415, 180, 663, 2574, 2930, 633, 2100, 913, 3472, 555, 1066, 1189, 951, 2382, 107, 352, 420, 838, 928, 2516, 2942, 3116, 3356, 1807, 3463, 726, 1424, 1461, 1796, 3477, 570, 1910, 2334, 2302, 2321, 240, 516, 744, 835, 927, 1457, 66, 655, 866, 577, 1765, 1864, 3424, 2674, 2828, 3453, 447, 1649, 1702, 1998, 2773, 3240, 3583, 2104, 2429]. Book [1996, 3378, 911, 2233, 121, 2115, 2209, 978, 1332, 810, 970, 2081, 3295, 2035, 993, 2806, 2891, 1927, 2099, 3020, 579, 809, 2252, 3482, 959, 172, 602, 632, 887, 910, 1212, 3242, 1776, 2027, 2157, 2491, 2600, 2982, 15, 3050, 2554, 2381, 2463, 2919, 3562, 3014, 2884, 29, 342, 1007, 1366, 2216, 3349, 3163, 80, 287, 2532, 270, 1080, 1382, 1540, 2405, 3409, 767, 1419, 1485, 1975, 2890, 2905, 207, 290, 909, 992, 1438, 1536, 2170, 2885, 2899, 2958, 1200, 3369, 3481, 1000, 200, 2108, 791, 2260, 2937, 2529, 3082, 2416, 149, 530, 3553, 575, 3073, 417, 833, 1111]. Book [1193, 1287, 2514, 3315, 1321, 2800, 748, 1045, 1774, 1866, 2680, 3003, 1685, 3080, 443, 1641, 1716, 3189, 208, 694, 2658, 1396, 1049, 1579, 1451, 2744, 2578, 2443, 2526, 2590, 391, 3148, 20, 378, 2579, 3007, 2909, 435, 1567, 3558, 1731, 3473, 585, 990, 1956, 2385, 3405, 2646, 1105, 1373, 2082, 2506, 781, 1480, 2275, 2750, 1705, 11, 1605, 2718, 3297, 1051, 2787, 2751, 187, 723, 1680, 1934, 465, 3108, 76, 1060, 582, 876, 2205, 2888, 855, 3235, 1407, 2380, 1230, 1191, 1553, 3448, 253, 1274, 1993, 2247, 2582, 3243, 547, 1001, 1091, 1362, 1547, 2462, 2094, 3165]. Book [2285, 3107, 2608, 2642, 2874, 3053, 239, 2922, 647, 3236, 3354, 1171, 1604, 1878, 2119, 134, 561, 300, 475, 654, 696, 885, 948, 3169, 3488, 3153, 520, 2510, 429, 2996, 1913, 3358, 3373, 2782, 1657, 662, 820, 24, 947, 1201, 1310, 2110, 2777, 3131, 1740, 2455, 110, 1205, 74, 316, 1954, 2406, 536, 673, 480, 2732, 1644, 853, 958, 1334, 3320, 651, 1387, 1518, 1666, 2442, 3410, 118, 3269, 3416, 129, 1800, 70, 3364, 190, 1176, 2242, 370, 3078, 3370, 209, 492, 714, 926, 1099, 1427, 2230, 2739, 967, 669, 1143, 1460, 1596, 2757, 1469, 2445, 322]. Book [1900, 3381, 550, 918, 972, 3159, 1555, 1721, 2825, 2641, 353, 540, 645, 743, 1101, 1284, 1477, 1556, 1726, 1960, 2269, 2355, 2978, 3522, 2156, 266, 2106, 934, 157, 521, 783, 973, 1513, 1771, 1922, 2001, 2084, 2589, 3340, 3437, 261, 1832, 1896, 336, 2454, 2277, 2283, 2322, 1595, 2020, 2748, 2967, 3368, 3232, 574, 3529, 1369, 1464, 3241, 2687, 2745, 691, 1686, 3326, 2662, 106, 1580, 2273, 1439, 775, 881, 1764, 1871, 112, 1296, 2102, 2183, 2354, 2903, 2984, 3534, 3086, 2250, 2654, 2969, 433, 509, 642, 1093, 1184, 1997, 2163]. Book [111, 279, 358, 514, 2036, 2199, 2975, 3531, 2525, 2441, 62, 2837, 3010, 1517, 683, 788, 977, 1151, 1241, 1347, 3270, 1280, 2418, 1323, 1247, 451, 2508, 2571, 2172, 1125, 2906, 479, 1377, 2173, 3406, 3294, 3528, 1479, 2264, 2719, 674, 1983, 1730, 2085, 3440, 3566, 1570, 1218, 1441, 2316, 2193, 2723, 600, 843, 991, 3009, 1988, 2798, 3145, 210, 333, 735, 1557, 2504, 2030, 194, 2619, 2917, 3389, 884, 2553, 189, 2898, 3058, 2860, 3183, 272, 474, 1053, 413, 2098, 1699, 331, 2779, 158, 905, 1106, 1491, 1688, 1816, 1920, 2336, 3088, 2026, 591]. Book [2759, 1034, 2933, 524, 1004, 2175, 2272, 2490, 3442, 1712, 36, 344, 439, 1185, 1552, 3298, 3429, 1351, 2342, 3348, 2125, 2389, 3427, 247, 1466, 1667, 1893, 3118, 975, 1119, 1436, 1569, 1801, 2096, 2169, 2998, 1453, 2963, 1399, 1684, 2569, 1561, 1723, 2158, 2880, 771, 860, 3001, 3508, 1953, 3166, 3454, 3168, 64, 1560, 2338, 3245, 542, 1411, 1817, 724, 1269, 2620, 2781, 3061, 2595, 2926, 3146, 159, 3128, 2698, 2993, 1013, 1103, 1196, 1474, 2007, 1221, 3175, 139, 231, 915, 1182, 2332, 2734, 3244, 463, 584, 1486, 1737, 1833, 2282]. Book [2679, 199, 826, 1012, 3158, 2981, 3413, 48, 235, 941, 1454, 1589, 1837, 3314, 1860, 67, 167, 259, 656, 1664, 2447, 2637, 1733, 2088, 834, 1476, 2507, 2593, 2660, 1838, 3379, 751, 1209, 3068, 1646, 1558, 3062, 3184, 879, 3318, 38, 446, 1021, 1118, 1299, 1899, 1978, 3259, 3536, 2894, 3087, 688, 1242, 1633, 1402, 3038, 968, 2851, 3443, 317, 1058, 1744, 2318, 228, 326, 1219, 1936, 2738, 1591, 1761, 147, 1358, 564, 935, 2178, 3066, 3380, 3582, 339, 825, 917, 1003, 1645, 2155, 2835, 3142, 2689, 1745, 1834, 2091, 2785, 3234, 749, 841, 2093, 2438, 2501]. Book [3511, 426, 549, 1546, 10, 504, 628, 1027, 2412, 2602, 3151, 193, 243, 2827, 3382, 3587, 1187, 823, 195, 3124, 3420, 3383, 1107, 1709, 2897, 3254, 768, 2023, 2460, 2895, 664, 2471, 185, 382, 837, 867, 1148, 1819, 1844, 1924, 2965, 693, 2924, 471, 710, 1109, 2166, 153, 258, 450, 830, 1488, 2187, 2805, 3517, 2826, 2690, 1872, 2079, 1356, 489, 2271, 2377, 3372, 1805, 2581, 227, 3360, 3005, 1216, 380, 793, 1409, 1502, 2964, 2123, 366, 1169, 2017, 2542, 1149, 341, 3489, 583, 671, 863, 1044, 1418, 1450, 1853, 2235, 2372, 2411]. Book [2530, 3083, 3407, 969, 2616, 1112, 3336, 267, 618, 861, 994, 1130, 1945, 2365, 1174, 3559, 1808, 484, 181, 512, 1655, 738, 1005, 3136, 2146, 3101, 329, 1400, 849, 2432, 2650, 2959, 573, 3388, 576, 269, 976, 1748, 1368, 2160, 174, 133, 464, 534, 621, 900, 1753, 1804, 1113, 56, 1762, 2181, 3538, 1146, 1514, 460, 2068, 2783, 1300, 2936, 2335, 2701, 3228, 817, 2042, 2197, 334, 1839, 2778, 2941, 1622, 2495, 3121, 1462, 1932, 2771, 3394, 372, 569, 1096, 3565, 390, 1773, 594, 795, 285, 1056, 1414, 2371, 3260, 3322, 675, 3156, 1946, 2281, 773, 1223, 2192, 1108]. Book [1929, 2594, 3057, 3551, 995, 1947, 2410, 2536, 2577, 2961, 3331, 2769, 386, 1095, 1276, 1908, 364, 2016, 2448, 2780, 3293, 3585, 361, 840, 1297, 1357, 1727, 1980, 684, 2423, 2503, 1898, 2188, 3266, 3557, 2547, 3006, 1365, 2095, 150, 924, 1616, 2063, 2858, 3504, 2934, 1134, 44, 1038, 925, 1961, 2004, 2437, 2977, 2470, 297, 1921, 2337, 89, 324, 1588, 1752, 2080, 2122, 3278, 3568, 593, 996, 145, 438, 548, 595, 728, 1011, 3524, 3579, 3580, 268, 780, 1183, 1397, 511, 1217, 836, 923, 1087, 1559, 2916, 2212, 3288, 2523, 2198, 2519, 1442, 2816, 85, 1768, 2767]. Book [864, 1742, 1412, 2749, 178, 1574, 154, 1224, 2290, 2652, 454, 351, 1363, 1470, 1571, 2586, 432, 2072, 2151, 1082, 163, 2849, 2498, 3130, 814, 2881, 3422, 3275, 3441, 3456, 2755, 507, 2446, 2537, 2836, 3546, 1840, 273, 2049, 2467, 346, 431, 533, 3152, 3190, 3392, 3300, 1239, 284, 3064, 1408, 206, 1097, 3561, 1447, 1843, 1957, 2061, 2200, 2369, 3176, 103, 1364, 2528, 3056, 733, 1687, 441, 725, 1178, 626, 3578, 1935, 2440, 1062, 3486, 1328, 1506, 1593, 2317, 2838, 3313, 179, 2367, 3090, 3167, 2804, 1186, 363, 3033, 12, 2124, 2291, 2320, 2918, 3461, 1775, 440, 2207]. Book [34, 442, 1123, 1835, 2013, 2184, 3075, 2921, 3008, 2668, 3588, 1825, 1421, 2737, 962, 1499, 878, 907, 1849, 2876, 237, 459, 980, 1539, 3273, 277, 1668, 3365, 470, 1035, 1208, 1824, 1897, 2671, 1220, 3374, 553, 831, 2077, 2241, 2651, 1124, 152, 1620, 3247, 2682, 551, 1279, 1463, 2845, 3230, 1173, 35, 920, 904, 3179, 2066, 3013, 952, 1204, 1577, 1659, 1930, 2840, 3403, 2073, 1244, 1102, 1621, 1779, 2811, 3375, 3467, 109, 294, 813, 1302, 1582, 1249, 2598, 2666, 198, 604, 1089, 3542, 1510, 1538, 376, 1050, 1448, 1459, 2210, 3355, 1535, 3469, 1504, 337, 919].

Book [2144, 2261, 3438, 1270, 938, 2492, 2607, 3172, 3411, 2512, 1444, 1700, 3173, 72, 1959, 2043, 2130, 3004, 3085, 1042, 2034, 2196, 3125, 2431, 2747, 468, 1127, 1496, 2357, 226, 903, 1720, 3316, 1144, 1618, 2040, 2213, 3290, 3150, 427, 436, 1632, 1452, 2145, 1906, 2067, 2149, 2327, 2420, 2886, 3239, 1914, 1286, 3256, 2474, 883, 3045, 1446, 1841, 2538, 785, 1585, 1519, 3127, 45, 3072, 2186, 2555, 3483, 2988, 1423, 1772, 3417, 2038, 2299, 3098, 816, 2046, 3139, 14, 2011, 77, 832, 3065, 3436, 3484, 18, 648, 1375, 2813, 572, 2419, 686, 2118, 3002, 123, 2339, 1085]. Book [1417, 3123, 1565, 2655, 304, 387, 2505, 1215, 2300, 3560, 3016, 1630, 211, 257, 349, 871, 989, 1192, 1319, 1869, 1928, 1984, 2089, 2215, 2266, 2347, 2592, 3371, 91, 589, 1226, 2024, 3390, 2165, 128, 931, 1798, 28, 1197, 2340, 2588, 3035, 3180, 2425, 2997, 2006, 2344, 3338, 288, 2786, 886, 912, 2842, 3506, 2902, 2724, 120, 383, 777, 3015, 3333, 2015, 1378, 2935, 2278, 2330, 1415, 2044, 2107, 2665, 3324, 3386, 2236, 828, 2345, 292, 3119, 3343, 1846, 1937, 2234, 3455, 126, 2774, 2847, 263, 1032, 183, 603, 1353, 2985, 3457, 3186, 790, 283, 519]. Book [63, 229, 469, 956, 1033, 2994, 3093, 3289, 3325, 1293, 3233, 3117, 503, 2927, 3147, 1516, 144, 1277, 3514, 3491, 950, 3051, 565, 1172, 1349, 1739, 2584, 1025, 249, 750, 1352, 3074, 1497, 1507, 371, 1335, 2497, 2649, 2911, 3032, 296, 598, 1114, 782, 1237, 1305, 625, 414, 827, 1120, 523, 731, 2249, 2331, 908, 1326, 19, 2694, 2101, 1500, 453, 2280, 3011, 3328, 320, 974, 1624, 2485, 1386, 2267, 808, 666, 53, 3327, 3530, 2599, 3164, 2974, 552, 2426, 3554, 2254, 360, 558, 1290, 1383, 42, 1023, 1963, 88, 176, 232, 393, 659, 787]. Book [1330, 1870, 2489, 3337, 323, 2670, 2841, 1030, 275, 681, 1494, 1704, 2576, 2675, 2109, 478, 505, 1140, 1405, 1665, 2064, 2328, 448, 998, 2360, 51, 1719, 1916, 2417, 2735, 476, 758, 1225, 1304, 2875, 3106, 3458, 59, 319, 1542, 1755, 1831, 3449, 445, 1104, 1282, 1376, 1566, 1725, 764, 1759, 2761, 3099, 3144, 2583, 2809, 3507, 69, 289, 695, 1541, 2822, 280, 2756, 2676, 230, 3485, 2763, 1938, 3515, 2722, 1625, 2663, 3041, 1337, 2939, 3155, 301, 2729, 278, 2153, 792, 1515, 3470, 355, 747, 1379, 1647, 1820, 2572, 2071, 2877, 3564, 3253, 875, 942]. Book [2960, 2142, 3376, 303, 1059, 1214, 606, 2854, 83, 162, 2815, 620, 2201, 643, 730, 1068, 1094, 1380, 1918, 2161, 295, 389, 1683, 2390, 2704, 2857, 1746, 763, 1598, 2358, 2105, 3044, 3513, 2830, 177, 1037, 2609, 3363, 1136, 1868, 2856, 3092, 556, 1019, 1976, 2078, 2152, 2248, 2968, 203, 845, 1231, 1443, 3362, 1554, 1989, 2257, 2424, 2893, 3332, 522, 772, 854, 1757, 2143, 2466, 90, 660, 759, 1489, 1747, 3450, 2887, 2459, 2904, 1592, 1271, 1562, 1640, 1812, 1245, 1674, 2111, 2772, 117, 3539, 619, 712, 318, 784, 2540, 3170, 136, 1141, 1199, 1637, 776, 3037, 2499]. Book [27, 423, 760, 1670, 2132, 356, 457, 711, 1384, 1925, 2002, 2657, 3387, 3229, 2154, 141, 1259, 1754, 2014, 3548, 717, 434, 2541, 1456, 2060, 1354, 847, 3138, 2511, 3059, 3434, 2274, 3129, 742, 3342, 2980, 3377, 1309, 1663, 2938, 2518, 3268, 1915, 943, 1228, 1677, 1845, 151, 638, 922, 1010, 1732, 3412, 3137, 3286, 3036, 2807, 3444, 238, 419, 508, 657, 756, 774, 1054, 1265, 1781, 1852, 1907, 2573, 2669, 3122, 1949, 3287, 532, 1117, 1360, 824, 1729, 3510, 397, 3476, 1813, 3248, 3154, 2486, 3081, 722, 727, 2422, 2972, 2823, 3329, 192, 2473, 2543, 2667, 2375, 412]. Book [1482, 1933, 2746, 720, 640, 2045, 2379, 3102, 2090, 2255, 2444, 2517, 1348, 1465, 1631, 1991, 680, 872, 205, 2083, 2031, 3261, 1401, 1619, 3468, 2742, 2664, 3047, 186, 271, 493, 560, 1307, 1575, 1931, 2121, 2362, 2820, 2999, 81, 2784, 3040, 1002, 2415, 2239, 3584, 1428, 2025, 3393, 1175, 335, 343, 546, 689, 394, 1190, 1818, 537, 957, 2705, 2087, 3135, 667, 862, 22, 1133, 1392, 1314, 1735, 1847, 2117, 1944, 2028, 2539, 2766, 3509, 1472, 2819, 3385, 2775, 1743, 55, 191, 1803, 2288, 3341, 1551, 1717, 2245, 142, 1202, 2112, 2323, 2990, 3571, 829, 1063, 1311]. Book [1568, 2012, 3344, 964, 1404, 1020, 1285, 2265, 2435, 2656, 1741, 2097, 3431, 2615, 644, 424, 685, 1083, 481, 3126, 1767, 3067, 155, 350, 639, 1370, 1467, 1607, 2901, 1652, 2164, 3425, 2195, 1129, 1941, 1067, 1264, 2765, 692, 1372, 2550, 2940, 1040, 1495, 1902, 21, 877, 2373, 234, 467, 766, 932, 1669, 1675, 1770, 2190, 2456, 2883, 3143, 3345, 365, 2185, 2534, 3586, 2070, 746, 916, 1153, 1359, 2502, 113, 347, 1139, 2859, 262, 3439, 332, 461, 566, 2613, 3521, 668, 1862, 2009, 2439, 3251, 1327, 2728, 3277, 1987, 1236, 175, 944, 1398, 1701, 166]. Book [395, 3478, 963, 1449, 1458, 2614, 2617, 2803, 1811, 3227, 1758, 2796, 3084, 3161, 182, 2366, 2458, 3452, 1014, 1810, 1263, 3162, 921, 1642, 1815, 1919, 2176, 2296, 2414, 2603, 40, 3541, 3265, 137, 260, 888, 815, 2612, 2697, 1848, 2475, 2551, 543, 665, 2409, 1533, 3071, 1682, 1749, 3237, 1639, 422, 3095, 3283, 850, 1394, 1829, 1940, 2103, 2364, 2451, 3069, 3556, 1333, 2567, 3103, 936, 2319, 2702, 3133, 2113, 2450, 1110, 2545, 2720, 2966, 3347, 1275, 1809, 2244, 2128, 1736, 1985, 171, 3157, 1909, 1243, 1917, 421, 857, 3177, 1865, 3258, 634, 86]. Book [3532, 1260, 2263, 2391, 3018, 3034, 3570, 32, 510, 2808, 3359, 3052, 156, 26, 146, 822, 2074, 2243, 2733, 3132, 2907, 3120, 2350, 2037, 1235, 1901, 43, 652, 765, 1248, 1799, 2349, 3428, 1281, 1636, 2359, 2983, 2727, 1780, 3141, 357, 531, 2986, 3535, 415, 3552, 3391, 3446, 3474, 1545, 2568, 2908, 3263, 188, 473, 965, 1131, 2640, 3105, 1029, 75, 3046, 2706, 108, 1977, 274, 2295, 2167, 2618, 225, 873, 966, 2855, 3100, 2585, 2776, 3017, 3487, 2469, 3575, 1826, 1389, 1836, 2688, 2298, 444, 1022, 1822, 1895, 2333]. Book [3433, 646, 1057, 1425, 1854, 541, 1052, 1478, 2976, 3526, 122, 255, 811, 1016, 1116, 1195, 1475, 2270, 2430, 2741, 47, 131, 527, 1590, 3076, 1498, 1312, 2605, 369, 1391, 2513, 1711, 2524, 1487, 2404, 1350, 1955, 859, 1150, 1403, 2238, 1121, 1855, 2832, 3567, 2672, 1823, 3238, 2276, 1039, 1995, 636, 1267, 1912, 1994, 3426, 315, 3174, 1238, 1390, 2407, 3282, 3544, 3285, 485, 1194, 1876, 588, 1203, 1573, 1689, 1979, 3351, 529, 135, 293, 682, 821, 914, 955, 1138, 1586, 1629, 1634, 1651, 1724, 1763, 1778, 1861, 1863, 1903, 2129, 2159, 2262, 2324, 2387]. Book [2488, 2770, 2810, 2844, 3187, 3231, 1672, 2126, 2208, 3550, 2740, 490, 599, 637, 729, 1009, 1092, 1658, 1710, 13, 491, 2361, 1549, 246, 1948, 1509, 3292, 1289, 2047, 2452, 2995, 3279, 1315, 2970, 483, 1088, 1814, 3569, 2932, 2174, 1950, 299, 3589, 3060, 1420, 1511, 1958, 506, 119, 125, 697, 1650, 3323, 1750, 2882, 3012, 2171, 2268, 3257, 3421, 631, 2022, 1584, 65, 368, 721, 953, 1338, 2203, 2610, 2696, 945, 375, 379, 3384, 2643, 168, 1122, 745, 2910, 165, 1246, 116, 679, 2351, 3490, 3296, 1069, 1601, 1229, 1508, 2325, 617, 1544]. Book [597, 3291, 2889, 739, 1098, 2638, 2202, 1292, 1273, 2500, 2892, 3335, 79, 786, 880, 1512, 425, 73, 2388, 2685, 428, 635, 1272, 1714, 2731, 1505, 251, 653, 1026, 1198, 1227, 2127, 2496, 2753, 3264, 245, 184, 2191, 2721, 3590, 2029, 563, 842, 2008, 1990, 1047, 1261, 3576, 1367, 31, 1268, 865, 515, 1064, 242, 2673, 1065, 3321, 2533, 2211, 2802, 3518, 282, 477, 624, 1233, 3505, 143, 1234, 1385, 1599, 1681, 1842, 2033, 2465, 3048, 1648, 340, 1361, 3430, 2699, 1471, 1892, 2168, 2831, 1207, 1760, 1827, 1211, 3191, 46, 127, 466, 661, 937, 979, 1031, 1206, 1572]. Book [2131, 2413, 2644, 3423, 2760, 1493, 2736, 2703, 2326, 204, 545, 641, 901, 1179, 1422, 538, 1180, 1628, 2730, 212, 348, 449, 3523, 2246, 114, 256, 458, 1015, 2258, 373, 630, 757, 954, 1041, 1406, 1455, 2929, 3353, 3471, 1320, 41, 3462, 1821, 2352, 2292, 946, 658, 1355, 590, 3357, 33, 732, 1381, 1643, 2003, 3250, 3435, 2279, 3480, 1728, 3055, 592, 3077, 252, 1992, 2162, 930, 1137, 698, 1426, 1877, 2971, 472, 578, 678, 1410, 1501, 1600, 1676, 1766, 1856, 1952, 2535, 2606, 2686, 2852, 2923, 3091, 3459, 1473, 1653, 1722, 2005, 2253, 3330]. Book [2758, 2991, 1416, 1503, 2878, 3280, 3366, 3574, 54, 762, 1086, 1578, 2817, 949, 1962, 3533, 1100, 1188, 581, 2039, 384, 557, 1283, 2256, 2341, 3134, 3404, 2116, 2286, 2374, 2544, 2575, 2762, 2987, 374, 518, 1266, 276, 1170, 2231, 1142, 1222, 196, 677, 2204, 2472, 1769, 2522, 2621, 586, 164, 61, 741, 1046, 1291, 2487, 2683, 3171, 2928, 254, 2259, 1278, 169, 385, 2378, 2232, 699, 794, 1875, 1181, 1797, 1145, 1329, 1867, 2041, 2925, 670, 1602, 1550, 1627, 2693, 2065, 737, 1492, 1532, 1802, 1939, 2048, 2147, 2182, 2520, 2717, 2821, 3540, 104]. Book [354, 755, 1298, 1756, 2457, 2818, 3063, 3149, 60, 1318, 2076, 23, 2315, 2992, 1926, 1126, 1301, 1303, 1858, 452, 933, 462, 753, 1154, 1240, 1393, 71, 173, 381, 3319, 1006, 1905, 1942, 265, 3572, 17, 649, 839, 929, 1531, 2814, 2900, 3070, 1850, 2297, 377, 2194, 3188, 2943, 3185, 2850, 2601, 596, 1703, 672, 1587, 2428, 3525, 416, 2403, 3465, 2648, 3519, 456, 1777, 1654, 2092, 2515, 3432, 812, 676, 761, 844, 2421, 779, 999, 1090, 1551, 1997, 1950, 430]. Booke [1248]. Bookkeeping [1987]. Books [2075, 602, 1776, 885, 1913, 492, 1247, 600, 1921, 728, 1293, 1566, 301, 1271, 1248, 225, 1634, 599, 699, 1519]. Boole [1124]. Boorse [225]. Booth [3259]. Bopp [392]. Borders [2536]. Borell [3311, 1139]. Borgato [2271]. Boris [654, 1953]. Bork [2201]. Born [1120, 639, 3223, 3314]. Borst [1895, 2761]. Bos [1718, 724, 3033, 2808]. Boschung [42]. Boscovich [564]. Bosković [650, 930]. Boslough [2770]. Bossier [1092]. Bossy [1189]. Boston [3295, 3473, 3380]. Bosworth [3473]. Botanica [2517, 2444, 1084, 261]. Botanical [1607, 2452, 1750]. botanicas [2803]. Botanico [2090]. Botanik [663]. Botanikers [3551]. Botanist [571, 178, 612, 3254]. Botanist-Cure [3254]. Botanists [2231]. Botanizers [1391]. Botany [927, 2751, 1343, 2665, 2243, 1574]. Bothwell [56]. Botsford [2786, 2391]. Bottin [2063]. Bottle [3554]. Bottom [3124]. Bottoni [1654]. Boudet [2333]. Boudoir [2307]. Boudri [2092]. Boulliau [2166]. Boulogne [1207]. Boulton [1087]. Bound [810]. Boundaries [2718, 1872]. Bouquiaux [2348]. Bourel [522]. Bourguet [270]. Boussingault [659]. Boutibonnes [2511]. Bouvet [837]. Bouvier [2178]. Boveri [1225, 2793]. Bovet [587]. Bovier [1109, 2437]. Bowden [750, 1041]. Bowdoin [1940]. Bowen [1628]. Bowers [282]. Bowker [2616]. Bowlby [662]. Bowler [1169, 367, 2574, 1840, 1302, 3450, 774, 2451, 2450]. Box [2321]. Boyer [1458, 3364]. Boyle [3502, 2170, 645, 2898, 3058, 1021, 1378, 1612, 1021, 1926]. Brabant [2032, 2604, 1368]. Bradford [3336, 2675, 779]. Bradley [2774]. Bradwardine [437, 2500]. Bragg [1617]. Brahe [739, 1253, 1106, 2428]. Brahes [2428]. Brain [664, 1898, 1579, 3324, 2374]. Brains [3238]. Brainstorming [590]. Bramwell [381]. Branch [2826, 1330]. Brandmuller [2445]. Brandt [259]. Branson [360]. Brantlinger [170]. brasileira [3181]. Braslow [3579]. Brass [1256, 697]. Braun [1865, 1599]. Braunschweig [2646, 3150, 2033, 3447, 3150, 1856]. Braunschweig-Wolfenbuttel [3150]. Bray [2971]. Brazil [957, 1230, 2686]. Brazilian [1353]. Break [3566]. Break-Through [3566]. Breaking [1492]. Breakthroughs [3032]. Breal [1308]. Breast [3554]. Breeding [2565, 675]. Breeze [1834]. Breidbach [656]. Breidert [456]. Breizman [1953]. Brenda [3041]. Brendan [3486]. Brennan [2833]. Brenner [687, 2960]. Brenni [2117]. Brent [1048, 2648]. Bret [2275, 3146]. Brethren [2454]. Breuer [1956]. Breve [1374]. brevis [2074]. Brewing [2105]. Brewster [686]. Brian [3334, 3020, 2199, 2964, 3407, 1056, 1173, 3515, 522, 2037, 2640, 3187, 1329, 2076, 1931]. Briant [2844]. Bricker [742]. Bridge [2823]. Bridgewater [2951]. Bridgman [1045]. Bridling [1636]. Brief [817, 1223]. Briefe [2506, 475, 3531, 837, 175, 1855, 1829]. Briefen [3080, 266, 2280, 1029]. Briefwechsel [2661, 2834, 3534, 1013, 364, 2016, 2291, 2265, 2185, 2008, 1572, 2987, 1531, 2457, 1531, 2195]. Briggs [924]. Brigham [3537]. Bright [3262, 3082]. Brigitte [1993, 2300, 1122]. Brilhac [2539]. Brind’Amour [2079]. Bring [620]. Bringmann [2965]. Brisson [1178, 1237, 3427]. Brissoni [449]. Bristol [1661]. Britain [1594, 2314, 302, 2674, 1480, 2285, 1479, 1174, 2876, 1249, 230, 1743, 932, 668, 3161, 2414, 2479, 574, 1458, 170, 2828, 1205, 1832, 941, 3260, 871, 956, 1025]. British [1490, 213, 482, 2996, 3348, 2781, 1834, 1027, 1849, 2682, 1928, 3343, 1136, 1309, 2819, 3143, 2451, 3258, 2907, 491, 2279, 23, 377, 406, 104]. Brittain [1598]. Britto [3181]. Britton [1462, 2068]. Broadcasting [292]. Broadsides [885]. Broadus [285]. Brock [1623, 958, 2749, 469, 3155]. Brocke [1215]. Brockliss [3043]. Brodbeck [2026]. Brodersen [266]. Brodie [2156]. Brodine [1169]. Broer [3060]. Broglies [2710]. Broke [1446]. Broks [3004]. Broman [3151]. Bromberg [1219]. Bronstein [2124]. Brooke [992, 104]. Brookes [481]. Brooks [1617, 76, 3557]. Brothers [869]. Brotherton [2269]. Brotons [2490]. Brown [3562, 3166, 898, 2640, 1203, 568, 1851, 683, 1514, 3012, 1098, 1227, 204, 1745]. Brown-Sequard [1745]. Browne [1748]. Browning [1502]. Brownrigg [1825]. Brozek [51, 3047, 3047]. Bru [2010, 2516, 3442]. Bruce [1795, 959, 2082, 2380, 1191, 2641, 158, 905, 2934, 1397, 1328, 1286, 1984, 1674, 1190, 1584, 630, 1416]. Bruch [758]. Brück [172, 172]. Brudern [1337]. Bruer [1604]. Bruges [729]. Bruins [516]. Brümmer [1123]. Bruna [913]. Brunelleschi [903]. Brunner [2700]. Bruno [1671, 555, 1189, 1796, 2642, 1560, 2338, 3245, 2895, 304, 974, 1312, 2182, 2621, 2585]. Brunsdon [1370]. Brunswik [2843]. Brush [2385]. Brute [2937]. Bryan [3074, 3161, 3165]. Bryant [1089]. Bryden [34]. Bryder [110, 679]. Bubonic [2829]. Bucciantini [1925, 157, 2589]. Bucciarelli [2302]. Buch [3246, 2502]. Buchanan [3041, 620]. Buchel [1914]. Buchhaupt [2462]. Buchler [1337]. Buchs [1996]. Buchwald [2117, 2569, 2023, 3260, 173]. Buckl [1996]. Buckley [285]. Buckminster [2938]. BuckyWorks [2938]. Bud [2209]. Buddhist [1364]. Budker [1953]. Buell [2758]. Buffalo [2453]. Buffon [1790, 651]. Buford [833]. Bui [3458]. Build [987, 1409]. Builders [734]. Building [2993, 308, 3374, 421, 1046, 3361, 3326]. Buildings [2522]. Buisseret [1649, 1702]. Bulbs [2703]. bulgarischen [81]. Bulger [1709]. Bulhof [1669]. Bulkeley [1138]. Bulletin [509, 825, 508, 110]. Bullough [2038]. Bundesrepublik [472]. Bunge [480]. Burbick [2113]. Burckhardt [536]. Burdens [195, 3420]. Bureau [960]. Bureaucrats [470]. burgerliche [2827]. Burghard [2923]. Burgundio [1716]. Burgundy [3433]. Burke [223]. Burkhardt [852, 2449, 269, 1748, 3263]. Burleigh [2204]. Burnby [493]. Burnet [1417]. Burnett [713, 2426, 3431]. Burnham [1619]. Burning [219, 3275]. Burns [225]. Burov [1852]. Bursey [948]. Burston [3468]. Burt [1147, 2026]. Burton [1260, 233]. Burtt [616]. Burwick [168]. Busard [1096]. Bush [3366, 987]. Bushmen [3180]. Business [2993, 675, 763, 2168]. Businessmen [1235]. Bust [1422]. Butantan [2027]. Butchart [3392]. Butlan [2070]. Butrica [3007]. Butt [2216]. Butte [2939]. Butts [2037]. Butzer [2150]. Bynum [110, 1896, 2180]. Byrhtferth [2582, 2582]. Byrth [1248]. Byzantine [431].

C [1232, 1484, 1891, 3096, 2681, 580, 571, 68, 3573, 1617, 2622, 3097, 744, 835, 579, 15, 767, 1451, 465, 2380, 807, 3243, 480, 783, 2322, 111, 194, 3259, 935, 3383, 863, 1418, 2616, 2536, 3331, 840, 924, 836, 2198, 2749, 1421, 551, 2845, 1577, 2666, 1089, 2512, 468, 2213, 1632, 1906, 304, 3324, 144, 1277, 2584, 782, 2809, 2854, 620, 2968, 1554, 318, 1141, 2473, 872, 1619, 2742, 2087, 191, 142, 3344, 964, 877, 422, 1613, 2733, 2037, 225, 1876, 1573, 2129, 2770, 1511, 697, 2268, 3421, 2202, 3335, 880, 661, 1628, 458, 2258, 1100, 2621]. C [2683, 1797, 2992, 2814, 2287, 29, 1764, 2335, 731, 3250, 1777]. c-Numbers [1764]. C.I.D.I [124]. ca [1080, 2260, 2739, 1561, 3375, 1566]. cabala [1101]. Cabale [2338]. Cabinet [3048]. Cadden [1636]. Cádiz [563]. Caen [745]. Caesar [3131]. Caesarean [639]. Cafarella [2963]. Caffrey [201]. Cahan [1945, 1945, 2365, 466]. Cajal [1760, 1760]. Calandrelli [2271]. Calaprice [2944]. Calcoen [1649]. calcul [1631, 271]. Calculating [1544]. Calculations [3117]. Calculus [900, 932, 159]. Caldani [2271]. Caldas [2175]. Calendar [2905, 1003, 2137, 2870, 3507, 3541, 2987, 2749]. calendario [1374]. Calendars [532]. Calenders [2329]. Calendrical [3117]. Calero [3440, 351]. Calestous [715]. California [3104, 180, 3073, 1218, 3483, 3096]. Callatay [3330]. Called [1924]. Callens [2820]. Calling [1210, 1843]. Callu [1808]. Calorimeter [497]. Caltech [1593]. Calvinism [986]. Cambrian [48]. Cambrian-Silurian [48]. Cambridge [2725, 3252, 29, 2416, 2172, 2593, 1558, 1217, 3506, 3387, 2711, 3487, 3323, 455]. Cambrosio [3093]. Came [2129]. Camel [3467]. Cameron [1536, 2626, 1992]. Camerota [1473]. Camille [118]. Campagna [1912]. Campaign [2845]. Campanella [736, 1922, 2432, 449]. Campbell [1281, 482, 151]. Campbell-Kelly [482]. Canada [1143, 1460, 56, 3098, 47, 595, 1133, 669, 2688]. Canadensis [1460, 527]. Canadian [3108, 3559, 1957, 198, 1133, 865]. Canal [2912, 3472]. Cancer [1232, 69, 3291, 3475, 3290, 2469, 2212, 2469]. Candid [967]. Candle [3406]. Cane [627]. Canedo [1197]. Canedo-Arguelles [1197]. Caneva [1664]. Canford [3163]. Canguilhem [331, 1893, 693, 1893]. Canham [1617, 3174, 3174]. Cannell [1833]. Cannizzaro [2756]. Canones [2503]. Canterzani [2271]. Cantor [486, 1366, 2911]. Cantwell [3187]. Cao [2918]. Caos [811]. Capacities [689]. Capel [262]. caphe [2272]. Capital [3161]. Capitalization [3027]. capitula [1366]. Capkova [1655]. Cappelletti [521]. Capponi [2590]. Captain [3123]. Captive [2853]. Capture [3320]. caracteristique [2508]. Cardinal [1584]. Cardiology [2934]. Cardot [1580]. Cardwell [464, 2488]. Care [1191, 70, 1742, 289, 412, 1458, 2545, 1681, 762, 3578, 2775, 384]. Career [1342, 405, 1192, 2373]. Careers [680, 2793, 299]. carence [1412]. Carey [235]. Carl [520, 574, 2183, 2984, 968, 1134, 2013, 952, 2675, 2060, 1029, 3149, 2140, 1930, 3458]. Carlensis [2077]. Carleton [2021]. Carlino [2586]. Carlo [3352, 3000, 1555, 1732]. Carlos [2342, 3244, 637]. Carlson [1399]. Carmélia [916]. Carnap [2303, 1601]. Carnegie [275]. Carnot [2098, 174]. Carol [1673, 2829, 2912, 2554, 1221]. Carola [2994]. Carole [2654]. Carolina [2673, 948]. Caroline [110, 1994]. Carolingian [2150]. Carolinum [3447, 2646]. Carolyn [331, 624]. Caron [1580]. Carothers [3285, 3569]. Caroti [2590, 1912, 2892]. Carozzi [1454, 2178, 1026, 1198, 1454, 1198]. Carpenter [2148]. Carpi [1188]. Carpintero [525]. Carra [339]. Carrara [1556]. carrés [2165]. Carrier [140]. Carrion [2090]. Carrithers [1411]. Carroll [2234, 2830, 2456]. Carson [2211, 1510]. Carsten [2842]. carteggi [2756]. Carteggio [743, 157, 2271, 2270]. Carter [2108, 2108, 393, 73]. Cartes [1649]. Cartesian [2571, 1377, 355, 1820, 737, 2564, 1192, 1384, 737]. cartesianesimo [1821]. Cartesianism [1730]. Cartesio [450]. Cartographers [1998]. Cartographic [1702, 85, 3414]. Cartography [1649, 85, 1452, 2145, 2813]. Carton [1673]. Cartwright [2765, 2777, 689]. Casa [868]. Casado [2915]. Case [2294, 2427, 1888, 2738, 2941, 3504, 2184, 376, 1075, 2236, 1704, 319, 1669, 2713, 2844, 499, 894, 1244, 2907]. Casebook [1825]. Cashdollar [668]. caso [853]. Caspar [2258]. Casper [2174, 165]. Cassedy [1086]. Cassidy [1617, 2461, 2376]. Castagnola [1101]. Castelli [157]. Castle [3473]. Cat [2765]. Catala [2255]. Catalog [2178, 795, 1428, 2885]. Catalogazione [1471]. Cataloging [1778]. Catalogo [3350]. Catalogue [2596, 492, 1247, 600, 2651, 750, 1248, 3076, 599, 1291, 1466, 1066, 383, 301, 1360, 699]. Catalyst [3283]. catalytique [1671]. Catastrofe [3444]. Catastrophes [3354]. Catastrophic [422]. Catastrophism [850]. Catching [2761]. Categories [547, 1551, 1091]. Cathars [2632]. Cathedrals [3336, 179]. Catherine [559, 1796, 2642, 2782, 2125, 2671, 2592, 931, 1037, 2263, 3357, 1145]. catholicae [1916]. Catholics [2632]. catholique [1571]. Catolico [1374]. Catoptrica [1367]. Catrien [2335]. Cattelani [3350]. Cattenius [1651]. Catterall [2935]. Cauchy [2102]. Caudill [3449]. Caufield [1145]. Caulhiaco [3248]. Causal [785]. Causality [1060, 77, 979]. Causation [1730]. Cause [1382, 1789, 2895]. Causes [3512]. causis [2583, 1359, 2587]. Cavaille [2895]. Cavazza [1194]. Caveing [1176]. Cavendish [3532, 3065]. CD [2908]. CD-ROM [2908]. Cecil [2522]. Cecile [3375]. Celebration [1133]. celeste [1728, 1653]. célestes [2247]. Celestial [346, 1756, 2913]. Celestino [2272, 2517]. Celia [2262]. Cell [2494, 2024]. Cells [2494, 698]. cellulaire [660]. Celse [3132]. Celtic [2069]. cendres [2895]. censura [1470]. Censure [3429]. Census [2257]. Centaurus [515]. Centenary [1674]. Centennial [2399]. Center [1607, 2387, 3099, 2129, 2036]. Centers [1860, 831, 19, 2544]. Centi [2526]. Central [1873, 3026]. centrales [653]. Centre [210, 1307]. Centrifugal [6]. Centuries [1427, 600, 2805, 3343, 1248, 623, 744, 748, 2662, 2602, 2025, 2819, 2341]. Century [2896, 902, 1661, 1904, 1028, 2177, 629, 1795, 1576, 2829, 3267, 3160, 1873, 927, 1998, 2104, 970, 602, 767, 2578, 2608, 2996, 1205, 3364, 492, 2230, 2156, 574, 2687, 2036, 2525, 2085, 1491, 64, 2595, 3158, 941, 2088, 3536, 2851, 2178, 3151, 1844, 3489, 1945, 3136, 1839, 2771, 3260, 3557, 595, 431, 1506, 2918, 1499, 2055, 1208, 1279, 35, 1577, 1249, 2666, 1504, 2420, 3072, 2997, 2006, 2344, 2935, 956, 3289, 1025, 2109, 1665, 448, 695, 2105, 1037, 2609, 1231, 2904, 1562, 760, 3387, 2823, 2087, 3344, 1285, 963, 2458, 1810, 1639, 2113, 3359, 156, 2349]. Century [2986, 2640, 3526, 1391, 2844, 2714, 745, 2500, 3430, 1211, 1206, 2736, 2279, 2817, 2925, 381, 1942, 1888, 2632, 2398, 2869, 1786, 3399, 217, 1078, 805, 3113, 407, 2137, 3112, 1139, 3496, 1165, 3398, 611, 3400, 3114, 3361, 2180, 2289, 1295, 2653, 2848, 2833, 290, 1191, 353, 2269, 2779, 1034, 917, 841, 181, 2197, 2577, 3579, 533, 3190, 103, 179, 1959, 2588, 2584, 371, 1755, 3059, 142, 1139, 2545, 483, 946, 592, 3366, 1492, 462, 173, 3070, 565, 493, 1532]. Century-of-Progress [1844]. Ceranski [2978]. Cercignani [3352]. cercle [931]. CERN [1046]. Certainty [2368, 2910, 664, 2492]. Ceruzzi [2215]. cerveau [2188]. Cesar [1356]. Cesare [2742, 2167]. Cevasco [3108]. Chabas [1917]. Chadarevian [3464]. Chadwell [225]. Chadwyck [1291]. Chadwyck-Healey [1291]. Chaga [3378]. Chaia [1548]. Chain [1358, 1329]. Chair [2603]. Chaix [3254]. chaleur [3067]. Challenge [961, 1582, 3371, 1376, 23]. Challenger [3176, 1749, 1682]. Chalmers [788].

Chambers [2017]. Chamla [3434]. Champion [1761]. Champlin [1836]. Chance [418, 187, 1804, 45, 2121, 971]. Chances [785]. Chandelier [2338]. Chandler [1518]. Chandos [568]. Chandra [863]. Chandrasekhar [863]. Change [1336, 2386, 1334, 1402, 1358, 45, 3040, 299, 2802, 336, 775, 3588, 246, 2703, 2264]. Changed [418]. Changing [2148, 1594, 1081, 3310, 2871, 990, 1494, 1867, 3185, 2941, 2547]. Channell [997, 604]. Chaos [2913, 2021, 2816, 2348, 2573, 810, 2816]. Chaos-Manuscript [2816]. Chapel [1134]. Chapman [2655, 596, 2021]. Chapter [1135]. Chapters [3406, 1865, 1990]. Character [574, 45, 2740, 3504]. Characters [756]. Charcot [2188, 2762]. Chard [373]. Charge [2037]. Charismatic [2295]. Charitable [751, 2666]. Charle [212]. Charles [3527, 1395, 1904, 1738, 852, 1943, 2449, 3453, 172, 3165, 2455, 479, 688, 935, 549, 2965, 1169, 269, 1748, 1804, 925, 1574, 2067, 3324, 2426, 2830, 2856, 203, 3047, 3431, 2373, 2456, 668, 3251, 1458, 2908, 3263, 636, 617, 1211, 2374, 2987, 2480, 803, 2634, 1169, 2136, 848, 3549, 1975, 662, 479, 366, 569, 2875, 3571, 2987]. Charlesworth [684]. Charlotte [3358, 2761, 2819]. Charlton [1092]. Charmasson [212]. Charpa [1393]. Chartier [522, 141]. Chartrain [1466]. Charts [85]. Chastity [2339]. Chattopadhyaya [28]. Chaucer [2810, 1009, 252]. Chauliac [3248]. Chaussees [1753]. Checklist [1566, 397]. chemia [730]. Chemical [2386, 213, 1317, 3299, 2275, 329, 2181, 2192, 572, 1376, 3155, 2274, 2422, 1190, 588, 1031, 1191, 83, 588]. Chemie [1316, 1927, 2252, 3003, 266, 3360, 1932, 1535, 3546, 1398]. Chemiker [3447, 3360, 2122]. Chemikerin [3567]. chemins [320]. Chemisch [3367]. Chemisch-Physikalisch-Technischen [3367]. chemische [1535]. chemischer [654]. Chemist [99, 773, 3544]. Chemistry [1480, 1878, 279, 1801, 1402, 329, 2181, 1134, 405, 565, 1349, 1739, 2267, 2822, 772, 2274, 98, 2422, 2373, 3544, 1047, 2341, 1134, 2293, 3255, 2879, 1623, 1751, 3508, 2595, 1223, 1350, 3569, 116, 1172, 1047]. Chemists [2871, 2192, 944]. Chen [3050, 480, 128, 128]. Chene [2571]. Cheng [128, 3050]. Cheng-Yih [128, 3050]. Chengxue [1186]. chercheur [1412]. Chernin [1852]. Chernyak [1033]. Cherry [2226, 2674]. Chertok [1744]. Cheryl [2537, 865, 2232]. Chester [641]. cheval [2338, 1356]. Cheyne [1118]. chez [2888, 1486, 834, 2911]. Chiapas [2648]. Chiara [638, 1994]. Chicago [1969, 2297]. Chickens [2565]. Chief [249]. Chiesa [2445]. Child [595, 2850]. Childbearing [53]. Childbed [2108]. Childbirth [1742, 3152, 2761]. Children [2828, 3288, 3333, 1858, 3099, 415]. chimica [3545]. Chimie [2446, 587]. chimique [2237]. chimiste [1671]. China [15, 480, 2732, 1402, 837, 329, 1798, 1625, 301, 2422, 2972, 2265, 3319, 3310, 3114, 2240, 2829, 1461, 129, 1712, 1187, 2971, 1767, 315, 3574]. Chinamissionars [837]. Chinese [291, 2035, 3050, 1800, 2969, 128, 3034, 1187]. chinesischen [837]. chinoise [931]. Chioniades [431]. Chiropractic [1693, 2030]. Chirurgia [3248, 2074, 2158]. Chladni [3145]. Choice [2597, 3247, 3480]. Choices [2486, 1584]. Cholera [2110, 2671]. Chomsky [3086]. Chr [822]. Chris [3100, 3426]. Christa [2251, 3065, 1240, 614]. Christel [2679]. Christfried [81]. Christiaan [257, 2347, 2814, 2092, 2814, 2814]. Christian [2353, 3543, 342, 2514, 915, 147, 1835, 3075, 1114, 2078, 434, 2972, 461, 1273, 2900, 5, 3401]. Christiane [3020, 1412, 1225, 1747, 2900]. Christians [1180]. Christianson [954, 1106, 2530, 3530]. Christie [2935, 722]. Christina [1562]. Christine [669, 44]. Christoph [1923, 2884, 2650, 2080, 1841, 2254, 162, 2415, 3319, 929]. Christophe [212]. Christopher [3281, 3097, 909, 2885, 1205, 2732, 1983, 3259, 795, 1144, 1417, 3229, 2720, 1238, 2361, 1338, 2762]. Christos [547]. Chronicle [3211, 3514]. Chronicled [2520, 3540]. Chronicles [1145]. Chronik [1215, 1941]. Chronological [991]. Chronologie [549]. Chronology [3019, 3300, 2555, 88, 123]. Chubin [717]. Church [2084, 3437, 2236, 906]. Churchill [2057]. chymiae [1134]. Chymiatria [1191]. chymische [453]. Ciancio [2905, 2750]. Ciardi [2756, 2518]. ciel [2812, 427, 2333, 3418, 3412]. cielo [1365]. Ciencia [1331, 524, 3181, 132, 1690, 321, 2490, 1723, 1735, 1261, 523, 2434, 2600, 1470, 2255, 1389]. Ciencias [525, 526, 524, 184, 1828, 1633, 321]. cientifica [2434, 752, 261, 262, 352, 829]. cientifico [2334]. cientificos [1470]. Cijfers [658]. Cimento [3001]. Cimino [521]. cinco [2803]. Cinematic [1432]. cinese [1365]. Cinquecento [1102, 1722]. Cipolla [1732]. Circle [1332, 1604, 2199, 2080, 1562, 1190, 2087, 1332]. Circuit [579]. circulacion [1654, 752]. Circulation [1439]. circuli [2089]. Circumstance [541]. cirujanos [3445]. Citation [501]. Cité [14]. Cities [3326, 2466]. Citizen [2012, 225]. Citizen-Patient [2012]. Citizenry [2950]. Citizenship [2176]. Citoyen [843]. Cittadino [759]. City [3303, 3295, 3473, 139, 3172, 289, 619, 13, 3077, 54, 704]. ciudad [1654]. Civic [3112]. Civico [3104]. Civil [803, 2685]. civile [2443]. Civilisation [329, 2422, 2972, 3319]. Civility [2230]. Civilization [3419, 2150, 911, 24, 128, 414, 3518]. Civilizing [1670]. Claassen [2232]. Clagett [819, 2329, 314]. Claire [2622, 2324]. Clap [2297]. Clara [3416, 2270]. Claren [1721]. Clarence [3183]. Clarissa [2391]. Clark [2555, 3099, 695, 1264, 1620, 3474]. Clarke [1506, 697, 2668, 3438, 355, 1022]. Clash [238]. Clasicos [1354, 2435]. Class [2468, 652]. Classen [1468]. Classic [1536, 1227, 3471]. Classical [1764, 1196, 2066, 1549, 1628]. Classification [1243, 1592]. Classify [181]. classique [3067]. Claude [148, 3189, 935, 1571, 1843, 3458, 522, 422, 1165, 3034, 914, 1875, 2363]. Claudette [2182]. Claudia [837, 1068, 3474]. Claudine [2132, 2147]. Claudio [2578, 340]. Claudius [1360, 914]. Claus [3256, 875]. Clavius [2650]. Clay [3300]. Cleary [2497]. Clelia [645]. Clemene [2986]. Clemens [648, 1634]. Clercq [3140]. Clerical [1273]. Clericuzio [2341]. Clerk [1487]. Cleveland [1607, 2783, 1607, 3101]. Clews [3015]. Cliff [128]. Clifford [112, 3066, 3470, 3238, 2008]. Climate [987]. Clinic [867]. Clinical [1597, 3473, 62, 1087, 1344, 3046, 3026, 566]. clinicien [1412]. Clio [1646, 504]. Clius [524]. Clock [1019]. Clocks [2329, 3057, 532, 3320]. clos [2004]. Close [1220]. Closed [2543]. Cloud [225, 2041]. Club [2461, 3211, 3025]. Clulee [447]. Clyde [1044]. Cnide [545]. Co [1984, 1054]. Co-Evolution [1984]. Coast [2020, 1163, 1229]. Coastal [2802]. Cobb [3465]. Cobban [29]. Cobras [2027]. Cochetti [1284]. Cochrane [141, 922]. Cockx [2257]. Cockx-Indestege [2257]. Cod [1640]. code [824]. Codell [2108]. Codes [1754]. Codex [1638, 1003, 2347, 2733]. Codex-Calendar [1003]. codice [1561]. Codices [2331]. codici [1912]. Codicological [435]. coelestium [1102]. Coffa [1601]. Coffin [1467]. Cognita [1648]. Cognition [1872, 977]. Cognitive [2785, 3138, 976, 2807]. Cohen [516, 1091, 108, 928, 934, 2748, 331, 3013, 414, 2566, 2319, 2147, 2814]. Cohn [2891]. Colaizzi [242]. Colapietro [2760]. Colbert [3453, 1406]. Cold [1954, 1502, 2542, 277, 1220, 1959, 1226, 2543, 2855, 3498, 894, 1959]. Cole [1604, 83, 1475]. Coleman [118]. Colette [1172]. Colin [3043, 213, 2416, 751, 790, 2822, 1472, 1942]. Collani [837]. Collapse [2650]. collecta [1637]. Collected [3276, 1673, 835, 2532, 768, 2460, 380, 2512, 2117, 2766, 473, 1291]. Collecting [2086, 2722, 2804]. Collection [601, 2885, 1270, 2152, 1428, 86, 697, 1458, 2814, 430, 795, 2276]. Collections [92, 214, 305, 398, 494, 607, 796, 889, 982, 1070, 1155, 1250, 1339, 1429, 1520, 1608, 1691, 1782, 1879, 1964, 2050, 2133, 2217, 2304, 2393, 2476, 2556, 2623, 2707, 2789, 2861, 2945, 3021, 3109, 3192, 3301, 3395, 3492, 3591, 87, 2633, 2377, 606, 1607, 888, 2621, 1649, 87, 696]. College [231, 2458, 135, 212, 2934]. Colleges [1528]. Collegialiteit [2693]. Collegiality [2693]. Collegiants [840]. Collegiate [308]. Collegio [2260]. Collegium [3447, 2646]. Collide [3159]. Colliding [776]. Collie [2521, 3076]. Collins [2044, 2510, 975, 2805, 3588]. Collisions [2475]. Colloids [279]. colloque [3460]. Colloquium [1278, 2199]. Colnort [824]. Colnort-Bodet [824]. Cologne [1134]. Colon [2681]. Colonial [2799, 1294, 1388, 2100, 2982, 897, 179, 3495, 2239, 1695, 1508, 2673, 3154, 2128]. Colonialism [1944, 2011]. Colonisation [1543]. Colonizing [2279]. Colored [2267]. Colors [1787]. Colour [3127]. Columbia [3204, 3204]. Columbus [1260, 1278]. combinationis [2183]. Come [3064]. Comes [3550]. Comet [3131, 2324]. comètes [2247]. Comets [3343, 991, 3119]. Coming [2021, 3205]. Comite [825]. Comiti [1897]. comme [3418, 2290]. Commemorate [1656, 2452]. Commentaire [2581, 1361, 2149]. Commentaries [3143, 2520, 3540, 1227, 1715, 826]. Commentarii [1274]. Commentary [2075, 2501, 3331, 836, 1279, 343, 2244, 2808, 2243, 734, 2380, 1386, 3133, 1367]. commentateurs [2877]. Commentationes [2903]. Commercial [3381, 1970, 1329]. Commercium [3534]. Commissaires [3538]. Commission [835, 1768, 2213, 225, 1529, 582]. Commissioned [2452]. Commitment [2283]. Committee [1967, 225, 679]. Common [1162, 207, 3448, 258, 2951, 3448, 1961, 3505]. Commonsense [77]. communautés [718]. Communicating [1127]. Communication [3408, 2326, 2343, 3589]. Communicationen [2992]. communion [3418]. Communist [55]. Communities [3112, 815]. Community [52, 1324, 927, 1230, 1786, 2282, 1591, 1148, 2534, 2351, 3471, 1604, 684, 3063, 672]. Compagni [1918]. Companion [328, 2172, 2593, 1558, 413]. Company [927, 1886, 386, 2693]. comparada [2027]. Comparative [559, 2549, 58, 980, 1459, 1757, 1202, 1394]. comparees [3075]. Comparisons [83]. Compendium [247, 436]. Competition [293]. compléments [914]. Complete [835, 2739, 935, 1677, 1948, 1816]. Completed [1146]. completes [2338, 2895, 2338]. Complex [910, 1954]. Complexity [1876]. Composed [1557]. composición [1040]. Composite [1488]. Compound [1289]. Compromise [1511]. Compton [583]. Computational [80]. Computer [910, 3118, 3371, 1868, 1895, 2874]. Computerized [1778]. Computers [2543, 872, 1778]. Computing [3370, 871, 2215, 1136, 1319]. Computistical [253]. Computus [1895]. Comte [2667, 1939, 2473]. Con [1374, 1872, 1102, 2347]. Conant [2878]. Concept [2839, 2632, 2085, 3166, 2916, 2127, 236, 1713]. Conception [691, 2163, 1712, 2448, 3275, 737, 385]. Conceptions [77, 1717, 1799, 2174, 2911, 203]. Concepts [84, 488, 1594, 774, 2417, 2231]. Conceptual [709, 905, 1224, 1290, 1289, 2918, 1512]. concerning [2964, 1720]. concetto [2526, 2241, 3434]. Concise [1087, 424]. conclusions [2495]. concursu [2508]. condamnation [3340]. Condit [2675]. conditions [1546]. Condon [894]. Condorcet [2010, 2597, 3538]. Condorcets [2747]. Conduct [2351]. Conference [3511]. Conferences [3104, 482]. conferma [973]. Confidence [2212]. Configurations [1892]. Confirmation [1422]. Conflict [448, 3100, 1206]. Conflicting [2226]. conflicto [2825]. Confrontation [870]. Congettura [1199]. Congregations [3340]. Congress [2641, 2159, 1750, 2621]. Congresscentrum [160]. Congressional [2641]. Congresso [327]. Conjunction [1388, 920]. Conjuring [2720]. connaissance [1237]. connaissances [1711]. Connection [1951, 986, 448]. Connections [3454]. Conner [2454, 3066]. conquêtes [1580]. Conquest [2179, 3163, 628, 1750]. conquise [1004]. conquista [2157]. Conrad [330, 2412, 1284]. Conscience [1509]. Consciousness [2672, 490]. Consensus [2294]. Consequences [1708, 1707, 2693]. Consequentia [3129]. Conser [2022]. Conservation [3448, 1664, 1404, 3283]. Conservatoire [2132]. consideracionibus [438]. consideratione [730]. consilia [1994]. Console [181]. Conspiracy [2214, 2822]. Constance [1468, 2063]. Constantine [1094]. Constitution [349]. construcao [3181]. construcción [432]. Constructing [941, 2762, 866, 2823]. Construction [2846, 1132, 11, 370, 3038, 193, 969, 3585, 346, 2118, 808, 1489, 3258, 499, 164]. Constructions [3392, 1905]. Constructivism [3313]. Construing [1968]. Consultations [566]. Consulting [1970]. Consumer [2212]. Consuming [3356, 3424, 171]. Contacts [2319]. Containing [1867]. contemporanea [1331, 122, 811]. contemporáneo [868]. Contemporaries [1007, 1642, 2610]. Contemporary [2930, 810, 3169, 129, 1330, 3102, 1545, 1198, 1280, 77, 296].

Contented [571]. contento [358]. Contested [1453]. Contesting [1831]. Context [488, 2309, 1543, 1112, 1494, 1695, 1253, 3433, 2951, 1797, 1727, 468, 3069, 108, 169]. Contexts [2718, 542, 815, 2806]. continenti [1589]. Continents [775, 776]. Contingency [2261, 2379]. continu [2528]. Continuing [278, 3363]. Continuity [807, 1282, 45]. Continuous [1545]. Contours [3045]. contra [3053]. Contraception [2156, 3232, 1087, 2156]. Contrasts [1031]. Contre [3433, 1875]. Contribution [149, 2283, 258, 2024]. Contributions [3461, 2257, 249, 3489]. contributo [2097]. Contributors [222, 502, 1079]. Contro [2589]. Control [3096, 199, 1005, 3006, 1244, 2357, 3223, 1087, 1860]. Controlling [281, 2694]. controverse [729]. controversia [253, 2445, 1589, 2763]. Controversial [1433, 3471]. Controversias [351]. Controversies [972, 296, 278]. Controversy [674, 1983, 941, 671, 1056, 2213, 2822, 3223, 2831, 1827, 2187, 3102, 48]. Convenzione [3457]. Conversations [405, 1109]. Convictions [2619]. Convolutions [900]. Conwy [2116]. Cook [2179, 1663, 2088, 3123]. Cooper [74]. Cooperation [2921, 95]. Cooperative [1162, 64]. Coopersmith [2677]. Cooter [491]. Copenhagen [2816, 2379]. Copernican [632, 710, 153, 1530]. Copernicana [2506, 973, 1920]. copernicanischen [442]. Copernicanism [2084]. copernicano [2741]. Coperniciens [3340]. Copernico [2445, 1920]. Copernicu [1816]. Copernicus [1894, 1105, 1373, 2506, 1816, 1013, 2650, 1015, 1888, 2308]. Coping [1347]. Copp [1704]. Copper [1495]. Coppermine [1444]. copula [3054]. Corbin [1908]. Corcos [2111]. Cord [1235]. Cordemoy [450]. cordiale [659]. Core [3405]. Corn [675]. Cornelia [471]. Cornelius [1641, 1918]. Cornette [522]. Cornut [3538]. Corollaries [915, 1362]. corollarium [2089]. corona [1828]. corpo [2586, 3138]. Corporate [3284]. corporum [2508, 1102]. Corps [3147, 2888, 339, 1259, 3495, 2826, 1042]. Corpus [552]. Corpuscularian [457]. Corrado [2744]. Correcting [3573]. Correction [411]. Correia [3416]. Correlate [620]. Correns [2140]. Correspondance [255]. Correspondance. [1808, 1808]. correspondant [1390]. Correspondence [2032, 2604, 1483, 2521, 852, 2449, 2834, 2303, 1212, 2905, 1200, 2275, 3146, 269, 1748, 2749, 1938, 1677, 2117, 3541, 3263, 953, 2987, 779, 1111, 836, 3076, 2202, 1747]. Correspondents [1786]. corrispondenza [3350]. corruptione [727, 2892]. Corsi [49, 46]. Cortada [211, 1868]. Cosandey [3509]. Cosmetic [3476]. Cosmic [724, 724, 426]. cosmicam [2903]. cosmique [3418]. Cosmographiae [435]. cosmologia [2811]. Cosmological [1717]. cosmologie [2573]. cosmologiques [1237]. Cosmology [1468, 544, 2768, 2119, 1621, 2497, 905, 671, 2650, 1621, 3119, 448, 1542, 634, 2246, 1800, 2638, 2831, 2730]. Cosmos [1976, 3083, 1372, 1657]. Cost [3176]. Costabel [312, 535]. Costandina [225]. costruzione [860]. Cotkin [655]. Cotterell [1173]. Cottingham [1111, 1558]. Couchoud [1711]. Coudert [2660]. Coudray [3346]. Coulston [935, 3251]. Coultaud [2227]. Coumet [522]. Council [1967, 580, 679]. Count [3578, 1802]. Countdown [225, 3587]. Counting [910]. Countries [744, 3470]. Country [3325]. Countryside [762]. Couples [2723]. Cournot [3157]. Cournotienne [3542]. Course [309, 2274]. Court [1286, 1190, 1527, 2360, 451]. Courtier [1616]. Courts [540, 3433]. Cover [2210]. Cover-Up [2210]. Cowling [574]. Coyle [459]. Coyne [2084, 207]. Cozzens [1517, 1244]. Crabs [2442]. Crabtree [3417, 603]. Cracking [2201]. Cracoviensia [1246]. Cradle [2799]. Crafting [3290]. Craftsmen [2002]. Craig [606, 2540]. Craige [1114]. Cram [1134]. Cranach [2549]. Cranefield [1229]. Cranium [1188]. Crater [2933]. Craven [1299]. Cravens [1858]. Crawford [2836, 680, 1676]. Cray [2856]. Crayfishes [2442]. Crazy [3379]. Crease [1871]. Created [2261]. Creath [2303]. Creating [2527, 2194]. Creation [2314, 3148, 1886, 674, 3379, 3382, 2023, 2017, 1418, 2838, 2196, 1043, 2809, 909, 1536]. Creationism [2609]. Creationists [1983, 1595]. Creative [3314, 978, 2723]. Creativity [714]. Creatures [2243]. Credibility [3260, 3408]. Cremieux [2539]. Cremieux-Brilhac [2539]. Cremo [3471]. Cremona [1360]. Cremonini [2742, 2259]. Crepel [2010]. Crest [1348]. creticis [1641]. Creveld [425]. Crew [74]. Crick [3374]. Crime [3079]. Crimean [182]. Crisciani [1994, 638]. Crises [3205, 2899]. crisi [1641]. Crisis [3198, 1045, 2193, 2894, 2474, 1138, 3154, 2043]. Criterion [428]. Critias [1178]. critica [1473]. Critical [1536, 3437, 2173, 258, 731, 2390, 1465, 2808, 3335, 1367, 1422, 1877, 2231, 2125]. Criticism [147, 2390]. Criticisms [2390]. Critics [714, 1845]. Critique [2289, 1790, 2662, 1744, 3293, 3064, 2566, 1138, 727, 3157, 914, 1240]. Croarken [871]. Croce [2368, 2910]. Crocker [1108]. Crohn [970]. Cróinín [253]. Croker [1227]. Crollius [3531]. Cromartie [2345]. Crombie [2626, 1891, 807]. Cromwell [3128]. cronicas [2158]. Cronin [1672]. Crosbie [846]. Crosby [2734]. Crosland [2446, 2357, 1624]. Cross [3296]. Crossgrove [2424]. Crossing [2697]. Crossroads [2040]. Croswell [2932]. Crowcroft [960]. Crowds [1401]. Crowe [3541, 710, 2235]. Crown [2250]. croyant [3460]. Crucial [3255]. Crusading [2373]. Cruz [3181, 2027, 225]. Cryptanalysis [2795]. Crystal [2785, 1097, 2358, 1865, 1822]. Crystallography [793]. cujus [2089]. Culinary [3425]. Cullen [2732]. Culpeper [1563]. Cult [2295]. Cultivating [2751, 2751]. Cultura [1556, 124, 1735]. Cultural [2309, 107, 1045, 2996, 574, 106, 2719, 3318, 1169, 2410, 2470, 2671, 1739, 1831, 1, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 560, 1167, 1404, 3237, 2720, 156, 3391, 1355, 3480, 1905, 141, 2475]. Culture [3408, 2433, 2086, 3284, 1295, 3537, 170, 1617, 3477, 287, 1080, 575, 1480, 2922, 336, 2020, 111, 2365, 1616, 3176, 441, 2845, 2299, 816, 3119, 1432, 3362, 1757, 2710, 1255, 2823, 2784, 2288, 3162, 2396, 2113, 171, 2268, 1273, 2168, 844, 3513, 2524, 3563, 2777, 2860, 2936, 1921, 2737, 2598, 3355, 1353, 1870, 1759, 69, 3513, 639, 1815, 652, 1281, 1636, 168, 2758, 1769, 1960, 1872]. Cultures [1776, 3243, 2939, 1530, 1093, 3394, 532, 3519]. cum [1134, 1728]. Cumaná [184]. Cumberland [1825]. Cummian [253]. Cummings [2453]. Cumulative [3017]. Cuneiform [247]. Cunningham [754, 3296, 2346, 2735, 1725, 3052, 757]. Cuppage [2179]. curandi [39]. Curar [3475]. Curators [844]. Cure [3254, 2882]. Cures [3579]. Curie [2639, 862]. Curiosities [2442, 3048]. Curious [3409, 3214, 3467]. Current [1306, 1, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 3017]. Currents [1164, 1234]. Curry [2962, 556]. Curtin [1034]. Curtis [905, 2759, 739]. Cusanus [2238]. Cushing [2257, 1132, 2379]. Cutcliffe [3513]. Cuthbertson [1207]. Cuvier [3354, 1518]. Cybernetics [1051]. Cyborgs [955]. Cycle [1462]. Cycles [3330]. Cyclohexans [654]. Cynthia [370, 1963, 347]. Cyril [1147]. Czartoryski [1816]. czasow [1655]. czeskich [1655].

D [1177, 2468, 2120, 713, 1999, 2549, 3543, 577, 2680, 1547, 958, 1800, 2106, 3294, 2098, 905, 1034, 542, 1833, 259, 1209, 3066, 3382, 1655, 2936, 3156, 2192, 3585, 1134, 1217, 1087, 3461, 1539, 337, 1906, 2266, 2935, 469, 53, 1023, 505, 3092, 203, 1141, 434, 1852, 2422, 3468, 493, 1175, 424, 155, 668, 1848, 2409, 2545, 1275, 188, 225, 1863, 1315, 697, 79, 2731, 252, 2923, 2991, 2522, 3070, 2253]. D.C [85, 1258, 888]. d’Abano [919]. DaCosta [1919]. daemonum [1371]. Daems [1637]. d’Afrodisia [2421]. Dagmar [3367]. Dagron [2338]. Daguerreotype [1491]. Dahan [2102]. Dahms [471]. Dáibhí [253]. d’Ailly [2078]. Daisie [490]. dal [2443, 2526]. Dalché [435]. Dale [3482, 3270, 73, 575]. d’Alembert [2093]. d’alentour [1633]. Dales [1277, 551, 1632, 2584]. d’Alexandrie [1176, 2242, 1361]. Dalitz [2532]. Dallas [2020, 1163]. Dallenbach [2923]. Dalley [3163]. Dalmedico [2102]. Dalrymple [1048]. Dalton [3393]. d’Alverny [2331]. Dam [2879, 1953]. Damerow [1987, 1289]. D’Amico [294]. Damigéron [146]. Dan [225]. Dangerous [590]. Daniel [2979, 587, 1306, 1287, 76, 1570, 3168, 2438, 195, 2073, 2327, 522, 3387, 2541, 3468, 22, 1862, 3556, 1039, 3280, 2039, 1170, 1756, 1531, 612]. Daniela [1556]. Daniele [3189, 3458, 2763]. Danielle [148]. Daniels [2972]. Danilo [789, 2443]. Danish [1626]. Danly [246]. Dannis [2594]. Dans [1649, 236, 1043, 687, 2812, 2247, 820, 322, 1093, 3427, 2079, 2188, 1496, 2330, 1063, 2333, 821, 914, 2496, 2182]. Dante [1097]. Danziger [866, 3469]. D’apres [1466, 2330, 914]. d’archivio [2259]. Darden [1334]. d’Aristote [1546]. d’aritmetica [3523]. Dark [3082]. Darke [2810]. Darkened [764]. Darlene [3482]. Darnell [777]. Darrigol [1764]. Darriulat [2165]. Darwin [1395, 3549, 50, 852, 2449, 662, 366, 1169, 269, 1748, 569, 284, 363, 2186, 2875, 1489, 1669, 1333, 2567, 2908, 3263, 1672, 2987, 803, 2139, 1395, 1169, 178, 1170, 1396, 1570, 1169, 177, 854, 2567, 1170]. Darwin. [2908]. Darwinian [367, 759, 3449]. Darwinism [952, 2840, 2024, 319, 3264, 986, 2139, 1840, 2024, 3550]. darwinisme [1570]. Darwinismus [663]. Darwinisten [758]. Darwins [2987]. Daryl [717]. Dassow [2525]. Daston [418]. d’astronomie [3524]. Data [2471, 304, 211, 1778]. Database [2739]. Dates [1462]. Dauben [2196].

Daughters [2233, 2751, 3005, 383]. d’Auguste [2667]. Daumas [1356]. d’Aurillac [1808]. David [709, 719, 789, 2206, 2461, 997, 2370, 3537, 544, 50, 328, 3466, 2313, 233, 778, 1617, 2801, 2376, 3424, 1649, 1702, 2891, 417, 2718, 3370, 1427, 1686, 881, 2525, 2798, 474, 2759, 247, 1569, 1453, 915, 2282, 941, 2438, 426, 583, 1044, 1853, 1945, 2365, 133, 2410, 3331, 684, 1134, 1087, 3130, 3578, 3167, 34, 980, 3230, 35, 920, 604, 1050, 1906, 2474, 1565, 3390, 3035, 2665, 1353, 1349, 1025, 2109, 3099, 289, 1515, 2071, 1231, 3548, 238, 3287, 2746, 2045, 872, 3509, 142, 3431, 1467, 1495, 332, 3541, 2551, 1909, 1917, 32, 156, 1281]. David [274, 3487, 1057, 1403, 485, 293, 745, 1273, 3335, 79, 31, 1233, 466, 757, 1455, 33, 2279, 1100, 1769, 1797, 1298, 2992, 1617, 3498, 3365]. Davies [1081, 1477, 387, 371, 230, 2753]. Davis [2898, 1872, 2236, 2609, 596, 3284, 1886, 3565, 2349, 2520, 3540]. Davy [1569]. Dawn [2676, 2770]. Dawne [3455]. Dawning [2919]. Dawson [2998, 2601, 166]. Day [2838, 2685, 225]. Days [3331, 2809]. DCL [1041]. DDT [2224]. Deacon [3121, 3015]. Deaconess [3295]. Dead [2792]. Deadly [1231]. Deaf [2323]. Deal [3293, 2929]. Dean [225, 416, 2066, 1741, 225]. Dean-Jones [2066]. Dear [2303, 808, 2570, 2303]. Death [2931, 2695, 2891, 67, 1838, 2210, 3045, 3345, 2726, 1034, 2091, 1742, 2204]. Debate [1564, 2054, 1832, 3222, 2771, 3376, 3363, 2310, 46, 1329, 71, 2872, 1056, 3580, 2328]. Debates [2043]. Debiprasad [28]. Deborah [2879, 605, 3322, 675, 626, 2999]. Debra [63]. Debru [2888]. Debunkers [1769]. Debus [1376, 3511]. débuts [256]. Debye [380, 380]. Decade [1133, 1321, 581]. Decades [783, 3470, 225]. DeCarvalho [1149]. December [2312, 1792, 2369, 2401, 1938, 1258, 1573]. Deception [2794]. Déchiffrer [270]. Decision [987, 3176, 1050, 2847]. Decision-Support [2847]. Decisive [3343, 2005]. Decker [1228, 2992]. Decline [1528, 2682, 952, 2076]. Declining [956]. Decomposition [1876]. Deconstruction [1344, 2792]. decouverte [3272, 587]. Dedekind [2911]. d’Edo [2436]. Dee [1066, 447, 2337]. Deep [2546, 1837, 961]. Deep-Sea [961]. Deepak [1543, 2669]. Deery [2692]. Dees [2631]. defauts [1259]. Defective [2695]. Defectology [1673]. DeFelipe [1760]. Defence [3485]. Defenders [1769, 2189, 923]. Defending [2201]. Defense [1696, 3297, 1922, 3092, 1602, 3236]. Deferent [32]. Deficiency [2801]. Defied [1080]. Defining [1832, 779]. Definition [1851, 703, 1801]. Defoe [3257]. Degani [3350]. Degenaar [2805]. Degeneration [956, 1760]. Degler [952]. degli [3139]. d’Egypte [935]. Dehue [2547]. Deichmann [1410]. Deified [1080]. d’Einstein [3460]. deJ [248]. Del [2645, 1911, 868, 3475, 3054, 752, 1828, 1374, 2526, 134, 2445, 2825, 1723, 1560, 1737, 564, 2155, 1183, 2586, 2241, 152, 1102, 2347, 2340, 643, 2893, 1354, 1915, 3444, 3154, 2444, 2517, 1311, 2435, 2803, 2741, 1912, 1471, 1821, 1722]. Delaporte [1893, 1035]. Delbruck [1221]. Deledalle [848]. Deliverance [2204]. dell [1199, 1194]. della [3054, 327, 3545, 124, 3350, 2890, 2443, 2526, 2750, 613, 2445, 1555, 645, 521, 1920, 771, 1589, 1183, 2347, 2763, 1199, 3434, 2097, 1826, 122, 811, 2270, 1498, 1471, 1643]. dell’affaire [2589]. dell’aria [2980]. dell’edizione [2901]. dell’illuminismo [2270]. dell’industria [3268]. dell’ipotesi [2518]. dell’unita [2270]. dell’Universita [1267]. dell’VIII [327]. Deluc [1026]. Demaitre [438]. Demeulenaere [3020]. Demeulenaere-Douyere [3020]. Demise [2224]. Democracy [218, 2201, 1330]. Democratic [1900, 3293, 1600]. Démocrite [2496]. Demographic [2091, 3292]. Demography [956]. Demonstration [2433, 2510, 2953]. demonstratione [1103]. Demonstrative [1547]. Demontage [2535]. Demystifying [999]. d’Encouragement [2633]. Denis [1537, 77, 2400]. Denise [1324]. Denkens [2742]. Dennis [2591, 748, 2571, 2904, 1741]. d’enquête [2495]. Dents [820]. Denver [3180]. Denys [146]. départementale [270]. Department [1047, 3210]. Depression [1036]. Derek [562, 386, 1134, 720]. Deriaz [1568]. deriva [1589]. Derived [2471]. Derk [1187]. Dermot [726]. dernier [1412]. Derrida [1064]. Dershowitz [3117]. Descartes [2743, 3527, 1111, 645, 905, 2507, 3062, 1775, 2261, 1820, 2572, 1116, 1289, 737, 3527, 1192, 3527, 3249, 2429, 834, 1558, 989, 747, 1475, 3339, 1288, 2591, 1287, 2658, 1377, 2085, 36, 2430, 2262, 1821]. Descent [1330]. Descriptio [435]. Description [78, 3061, 935]. Descriptions [83]. Descriptive [2779, 301, 1607, 699]. Desert [383]. desiderate [2256]. Desiderio [1793]. Design [2797, 2794, 96, 1134, 2823, 1253, 541]. Designing [2686, 2302]. Desire [2726, 3275, 171, 1636, 2307]. Desley [3015]. Desmond [1169, 3438, 355, 1750, 467]. d’Espagne [1633]. d’espèce [236]. destin [2147]. Destruction [2924]. Destructive [3575]. DESY [875]. Detailed [83]. Determination [1019]. Determinisme [1351]. Dethloff [2129]. d’Ethnographie [1129]. Detlev [2818]. detto [643]. d’Eudoxe [545]. Deutsch [2032, 2604, 3552, 730, 2502]. Deutsch-Franzosischen [3552]. Deutsche [875, 1766, 2424]. Deutschen [3385, 286, 2754, 1588, 3586, 40, 2413, 2987, 1142]. Deutscher [2122, 1946, 2535, 755]. deutsches [3003]. Deutschland [663, 2827, 2923, 374, 472, 196]. deutschsprachiger [2920]. Devastation [3153]. Developed [489]. Developing [3326, 2781]. Development [392, 244, 709, 488, 2828, 3473, 1230, 2441, 1247, 344, 968, 133, 2849, 3300, 179, 3008, 2607, 3483, 1497, 704, 932, 1289, 379, 2831, 2522, 586, 2520, 3540, 200, 1901, 1544]. Developments [2918]. développement [718]. Deviance [3176]. Device [2953]. Devices [910]. Devil [2636]. Devils [1371]. Devine [2262]. devineurs [3433]. dévoilé [824]. DeVorkin [474]. Devoted [107, 106, 509]. Devotion [3174]. DeVries [1644]. Dewachter [935]. Dewdney [3085]. Dewey [780]. DeWolf [968]. Dhanani [2246]. d’Hippocrate [549]. d’Histoire [1201, 503, 522, 1026]. d’historiographie [2247]. D’Hollander [1814]. Dhombres [900, 561, 522]. Diacu [2913]. Diagne [1124]. Diagnosis [1344, 3575]. Diagnostics [590]. Diagrams [905]. Dialectic [344]. Dialectical [298, 1243]. Dialogo [3053, 2741]. Dialogue [3502, 3437, 827]. Dialogues [724]. Diana [68, 1081, 1688, 815]. Diane [3580, 3176, 2694]. Diaries [3057]. Diario [2355]. Diary [1321, 3548, 225, 1863, 1301, 366]. Dias [1129]. Diätetikvorlesung [461]. Diaz [3386, 1390, 1040]. Diaz-Andreu [3386]. dibattito [2097, 1722]. d’Ibn [2070]. dichten [1313]. Dick [1866, 2771]. Dictionaries [406]. Dictionary [1904, 3108, 211, 303, 395, 3487, 1629, 1338, 2048, 2706, 2943, 84, 1153]. Dictionnaire [3189, 2132, 212]. Did [985, 2140, 284]. Didactic [1191]. Diderik [317]. Diderot [933]. diebus [1641]. Dieckmann [1225]. Diedrich [2016]. Diego [2947, 1888, 154]. Diekmann [2095]. Diemer [2521]. Dienel [1946]. Diephouse [2992]. dieser [174, 2013]. Diet [138]. Dieter [2884, 3145, 329, 3253, 1150, 2457, 1587]. Dietmar [2843]. Dietrich [2150, 517, 2122, 2070, 822]. Diez [3446]. Difference [2780, 3230, 1636, 2493, 2782]. Different [2381, 3528, 1503]. Differentialkalkul [2755]. Differenzierung [2413]. Difficult [2266]. Diffident [2170]. diffusione [251]. Digby [110, 2801, 2278]. Digging [2929]. Digital [2215]. digt [2428]. Dijksterhuis [1646, 3279]. Dilemma [70]. Dilemmas [2937, 2998, 1163]. Dilg [2256, 518]. Dilic [1245]. Dilke [725]. Diller [1941, 1406]. Dillingen [1819]. Dimension [692]. Dimensions [1709, 1425]. Diminishing [1516]. Dimitri [1906, 1550]. Din [1991]. Dingen [1996]. Dinner [3211]. Dinosaur [3572]. Dinosaurs [2990]. d’insegnare [1002]. Dion [3014]. Dionysus [692]. dioptrique [1809]. Dioramas [2548]. Dioscurides [822]. Dioskurides [2070]. Dioskurides-Erklarung [2070]. Diphtheria [376]. Diphtherieserum [2994]. Diptych [1385]. Dirac [2532, 770]. Direct [728]. Directions [11, 3317, 203]. Director [225]. Direktors [1941]. Dirk [3367, 153, 2714]. Dirt [3224, 1507]. Disabilities [1673, 2801]. Disability [3230]. Disablement [2693]. Disappearance [1344]. Disarmament [588]. Disasters [2967]. Disciple [2636]. Disciplinary [722]. Discipline [1210, 276, 1497, 2570]. Disciplines [2181]. Disciplining [2698]. Discontents [611, 1772]. Discontinuity [1984]. Discorso [2443]. discours [388, 2514]. Discourse [3014, 36, 1384, 2543, 1381, 808, 1531, 692]. Discourses [3230, 2391]. Discovered [2139]. Discoverer [2120, 1044, 368]. Discoveries [3353]. Discovering [2990, 1876]. Discovery [1981, 50, 622, 3573, 3522, 897, 67, 1450, 2042, 1244, 1683, 2083, 1612, 1940, 2932, 1750, 1648, 1134, 2159, 2745, 1687, 2643, 528, 1704]. Discrete [1545]. Discursive [792]. discurso [1374]. Discussion [2805, 489, 1728, 782]. Discussioni [3054]. Discussions [551]. Discvrso [2272]. Disease [2726, 1294, 2653, 3097, 3378, 2800, 3183, 1860, 3124, 3420, 593, 1459, 3045, 3289, 1507, 3337, 2109, 69, 1231, 1732, 1933, 1281, 1508, 2279, 332, 3182, 1707, 963, 2468, 376, 375]. Diseases [250, 2648]. Disenchanted [371]. Disinterestedness [1970]. Diskurs [1531]. Diskussion [1658]. d’Islam [149]. Disorder [810, 855, 2472]. Disorders [572, 3379, 3046]. Displacement [1306]. Display [1164]. disputata [2076]. Disputationes [648]. Disputations [2308]. Dispute [48]. Dissection [2737]. dissectione [2586]. Dissent [1969, 2844, 3086]. Dissertatio [1728]. dissertationis [358]. Distance [2397]. Distances [1015]. Distillateurs [824]. distillation [825]. distincta [1375]. Distinctiveness [375]. Distinguished [591, 3190]. District [1495]. Distrikte [3150]. Disunity [2718]. Disziplin [859]. Disziplinenbildung [1304]. Disziplingeschichte [325]. Diurnal [3057]. Diurno [2271]. Diversity [560]. Divide [1053]. Divination [3433]. Divine [96, 3244, 1582, 177, 155, 2261]. Divinorum [3244]. Division [75]. Divulging [1926]. Dixon [2897, 2339]. Djerassi [1134]. dlya [1014]. DMF [3189]. DNA [3573, 2690]. Do [803, 910, 1655]. Dobbs [2268, 1791, 1380]. Dobo [1032]. Dobson [3045]. Doby [2849]. Dobzhansky [2126]. Doctor [1028, 2088, 936, 2693, 2297, 2096]. doctorands [3298]. Doctoring [3358]. Doctors [2189, 1036, 1371, 3577, 2278, 2360, 481, 946, 30, 250, 1978, 334, 470, 1577, 3191, 677]. Doctrine [1320]. Doctrines [1717]. Documenta [2506]. documentarie [743]. Documentary [1706, 1017, 2689, 966]. documenti [2259]. Documents [2609, 953, 2685, 3340, 1026]. Dodgson [2456]. Does [2663]. Dog [2224]. Dogma [235]. Dogmatismo [2277]. Dogs [3581]. Doheny [3483]. Dohrn [1225, 1747]. Doing [2100, 2569, 1774]. Doktor [1660]. dokumentarische [3441]. Dokumenten [1762]. Dolby [3185]. Dollars [2929]. Dollo [2744]. Dolnikowski [2500]. Domenech [3475]. Domenico [1656, 2340]. Domestic [1126]. Domesticating [2466]. Dominance [419]. Domination [483]. Domingue [2011]. Dominguez [2249]. Dominic [548, 3490]. Dominick [1182]. dominicos [746]. Dominikaners [2253]. Dominique [1043, 3254, 1276, 522, 1046]. Don [1061, 769, 3581, 2593, 2117, 137]. Donahue [1650]. Donald [2931, 2348, 1055, 2240, 1806, 1873, 1105, 1469, 991, 3383, 3559, 464, 1134, 1038, 3561, 2431, 2064, 136, 1763, 2488, 3384, 3188]. Donawerth [3005]. Donder [1649]. Done [182]. Dongjuma [2515]. Donna [478, 3084, 955]. Donne [1817]. Donovan [98]. Don’t [2469]. Doolittle [3571]. Doom [2933]. door [41]. Doppler [1835]. Dora [2981, 1228, 2012, 3459]. Dorado [2517]. Dorinda [652]. Doris [2827, 476, 2258]. Dorman [1573]. Dorn [1082]. Dorothea [1245, 2087]. Dorothy [1706, 1483, 1027, 1400, 176, 3231, 2211, 590, 101]. d’Orphée [146]. Doskey [264]. Doubt [1382]. Doubts [2870]. Douglas [554, 2049, 2845, 1386, 787, 1404, 1736, 292]. Douyere [3020]. Dowbiggin [1208, 3098]. Down [2728]. Downing [3274]. Downwind [225]. Doxographical [1550]. Dozen [3025]. Dr [2521, 3003, 2749, 249, 3074, 1941]. Dr. [3477]. Draelants [2247]. Drago [2098]. Drake [1792, 649, 738, 831, 2815]. Drama [555, 3450, 634, 1406]. Drawing [3348, 3289]. Drawings [2080]. Dread [69]. Dream [126, 1948, 38, 933]. Dreams [3326, 1475, 590, 1866]. Drees [2702]. Dreiser [1861]. Dresden [3562]. Drew [479]. Dreyer [3241, 1916]. Drifting [775, 776]. Drinking [1619]. Dritten [2382, 2300, 2842, 1501, 1142]. drittes [175]. Drive [3009]. drogas [1911, 746]. Drogenkunde [1637]. droit [1897]. Dron [1852]. Dronamraju [2202]. Drosophila [2316]. Drossaart [1913]. Drouin [2326]. Drouin-Hans [2326]. Drug [1967, 2566]. Drugs [1619, 670]. Drum [2381]. Drunkenness [572]. drustva [930]. DSM [3379]. dualismo [450]. Duane [1890]. dubbele [1368]. dubitabilium [919]. Dublin [2463]. Dubois [368]. Dubose [1890]. Dubosson [3458]. Dubow [2696]. Duc [734]. Ducay [3053]. Duchenne [1207]. Duchesneau [1874, 3103, 660]. Ducker [571]. Duden [1028]. Dudley [1726]. Due [2963]. Duel [1956]. Duffin [1839, 3266, 1141]. Duffy [716]. Dufour [1390]. Dugald [1111, 190]. Duhem [1008, 3298, 2960, 2727, 1315, 687, 18, 687]. Duisburg [1821]. Duitse [2101]. Duke [3497]. Dukelow [2544]. Dukes [143]. Dummett [1485]. d’un [1043, 900, 1032]. Dunbar [3105, 1154]. Duncan [1081, 2449, 1748, 3263, 3055, 2595]. Dundes [631]. d’une [2363, 1546, 2485, 914]. Dunlap [1028, 477]. Dunsch [654]. Dunsink [20]. Dunstan [918]. Duplicity [714]. Dupree [178]. d’Upsal [320]. Dupuy [54, 1814]. Durbin [84]. durch [1830, 1113, 297, 1314]. Duren [767]. Dürer [350, 1472]. during [2275, 2125, 153, 3559, 2056, 179, 2245, 3496, 43]. Duris [3122, 1390, 1659]. Durkheim [2527]. Durling [1097, 1011, 1629, 1716]. Dürr [473]. Dust [1231]. Dutch [2745, 1377, 840, 996, 3496]. Dutch-Japanese [996]. Dutremblay [3538]. Duty [182]. Dutzend [175]. Duveen [1324, 2400]. Dwayne [2685]. Dyck [3541]. Dye [249]. Dyeing [1428]. Dyer [2712]. Dyestuffs [1752]. Dying [2224]. Dynamic [3310]. Dynamical [622]. Dynamics [3314, 2181, 1056, 3102, 3223, 622]. Dynamik [81]. dynamique [161, 388, 2508]. Dynamis [523]. Dynasties [1186]. Dyson [2301]. dziel [1655]. Dzielska [2419].

E. [616, 1889, 894, 669, 2451]. Eagle [370, 1442]. Eamon [1921]. Earlier [1978]. Earliest [1450, 1087]. Early [2086, 2054, 3249, 1295, 3537, 3043, 1434, 482, 1649, 2429, 1080, 1000, 465, 1191, 3243, 550, 2269, 3522, 1832, 2662, 2163, 1377, 1730, 735, 1837, 1209, 1021, 2894, 2602, 2542, 840, 1921, 1087, 2369, 1499, 3045, 349, 589, 1798, 603, 445, 1282, 1376, 3507, 1136, 556, 2002, 2025, 1612, 2819, 1987, 2545, 2263, 3218, 47, 1138, 1289, 2171, 1990, 1211, 557, 2544, 173, 498, 703, 2869, 2965, 1078, 5, 1164, 2138, 3401, 1986, 2100, 3050, 885, 326, 871, 2458, 2907, 1950, 2992]. Earman [1422]. Earth [291, 3299, 1702, 2416, 3405, 1310, 2798, 235, 969, 1357, 1538, 136, 3036, 1682, 422, 850, 1260, 1198, 1990, 1015, 8, 2228, 893, 3499, 1048, 117]. Earthly [2815]. East [2578, 2145, 32, 1229, 3077, 1987]. Eastern [2150]. Easton [1820]. Eastwood [905]. Easy [2267]. eaux [2970]. Eberhard [113]. ebraica [3434]. Ebstorfer [1278]. Eccentric [3238]. Echegaray [1311]. Echeverria [2508]. Echo [3383, 361]. echten [549]. Eckardt [3273]. Eckart [1215, 2415, 1501]. Eckert [1458, 472, 2378]. Eckhard [1278]. Eclectic [3557]. Eclectics [2114]. Eclipse [2368, 2649, 2870, 1255, 2009, 2910]. Eclipses [3036, 2247]. eclosion [3144]. Ecocide [1404]. Ecole [2102, 1753]. écoles [653]. ecologia [2915]. Ecological [3512, 1806, 2554, 2064, 2226, 2617, 239, 624]. Ecologist [1227, 1761]. Ecologists [960, 2930]. Ecology [2839, 1591, 1363, 3032, 759, 1404, 1227, 2127, 2930, 908, 1405, 381].

Econometric [857]. Economic [913, 3247, 3112, 872, 2121, 1987, 1549, 605]. Economics [539, 2321, 239, 2907, 539, 1030, 1516]. économique [271]. Economist [202]. Economy [2386, 1900, 593, 1869, 2064, 1765]. Ecosystem [2839, 1405, 2127]. écrits [2528]. Ecuador [1294]. ed [2277, 1199, 1267]. Eda [2823]. Eddington [2198]. Edelson [3374]. Edelstein [1428, 1973]. Eden [1835]. Edens [2239]. Edgar [2248]. Edge [1981, 577, 954]. Edgerton [1464, 2682]. Edges [1357]. Edgeworth [2121, 2121]. Edinburgh [1482, 2176, 2603]. Edison [372, 1947, 2369, 3178, 763]. edite [3415]. Editing [1435]. Edition [3246, 1641, 435, 1666, 167, 935, 731, 1386, 730, 2257, 1465, 1110, 2244, 2808, 1855, 3335, 1367, 2906, 727, 914, 648]. Editions [39]. Editor [2229, 3501, 3031, 2402, 615, 1974, 2635, 3226, 102, 1698, 806, 9, 899, 1794, 2957, 2141, 313, 410, 1168, 708, 2058, 224, 2716, 2484, 1346, 1615]. Editorial [802]. edizione [2347]. Edmondson [3173]. Edmund [1483, 176, 1064]. Edney [3258]. Edocere [638]. Eduard [654, 943, 1852]. Eduardo [2482]. Education [2287, 1310, 1945, 217, 1309, 1445, 1482, 2414, 2688, 1979, 579, 2690, 234, 2009]. Educator [1503]. Edward [2546, 1413, 2806, 290, 1373, 3235, 1555, 2745, 674, 2898, 2738, 382, 3372, 2372, 814, 980, 3374, 1519, 3483, 3117, 3074, 3106, 3449, 2609, 3363, 717, 1309, 2859, 225, 2407, 1549, 2882, 1760, 779, 672, 777, 2822]. Edwards [1103, 2543, 3250]. Edwin [3269, 1406, 2120, 2530]. Eells [1060]. eeuw [1297, 165, 2814, 658, 228]. Effect [3472, 1887, 3328, 3353, 2836]. Effects [2106, 2023, 572, 3393, 2925]. Efficacy [1800]. Efficiency [2794, 2480]. Efron [2189]. Egbert [1134]. Egg [3416]. Eggs [2565]. Eginhard [358]. Egon [2445]. Egorov [570]. Egypt [935, 633, 760, 962]. Egypte [935, 2149, 2330, 1711, 820]. Egyptian [819, 2329, 3521]. egyptiennes [1711]. Ehrenberg [85]. Ehrenfried [929]. Ehrlich [1433]. Ehud [249]. Eichstatt [2251]. Eigen [2360]. Eight [3383]. Eight-Ball [3383]. Eighteenth [1661, 1295, 1028, 2177, 623, 744, 748, 1427, 2269, 2662, 2595, 3536, 2178, 841, 3151, 2666, 1025, 2904, 2819, 3344, 2817, 3496, 565, 493]. Eighteenth-Century [1661, 1028, 2595, 3536, 2178, 3151, 2666, 1025, 2904, 3344, 2817, 565, 493]. Eighth [2838]. Eighty [2681]. Eighty-Year [2681]. Eileen [2976]. Eilhard [1398]. Eimas [1247]. einander [2013]. Einarson [1359]. einem [461]. Einfluss [1588]. Einfuhrung [2238, 145, 341]. Einhundert [654]. Einrichtung [325]. Einsamkeit [1304]. Einspruch [3479]. Einstein [1313, 2834, 768, 2460, 671, 2291, 2117, 2766, 1697, 2944, 769, 2376, 391, 2290, 278, 1847, 2028, 2605]. Einstein-Turm [1313]. Einsteins [1314]. Eiseley [1411, 954, 954]. Eisenhart [2057]. Eisenhower [582, 1582, 3183]. Eisenstaedt [2117, 2766]. Eisler [3123, 1472]. Ekbert [665]. Ekmeleddin [1005]. Elaboration [2952]. Elaine [554, 2322, 3328, 3421]. Elam [3548]. Elastizitatstheorie [1587]. Elazar [1594]. Elder [3383, 1714]. Electric [1399, 2023, 681, 3496, 2712, 2037]. Electrical [1887, 1957, 598, 606, 1598, 3496]. électricité [1580]. Electricity [748, 1164, 2466, 2966, 1827]. Electrification [2677]. Electrifying [1765, 1233]. Electrodynamics [1950, 2519, 2717]. Electromagnetic [1306, 2821]. Electromagnetism [1236]. Electronic [872, 3479]. Elegiac [616]. elektrischen [2033]. Elektrizitatsversorgung [2465]. Elektrizitatswirtschaft [2465]. Elektronen [875]. Elektronen-Synchrotron [875]. Elektronenmikroskops [2615]. Elektrotechnik [2757]. Elementarqualitaten [1463]. Elemente [1150]. Elementen [51]. elementis [1096]. Elements [3299, 2075, 3240, 2242, 1176, 2495, 2437, 3056, 2213, 1478]. Elena [1735, 352]. Eleonore [344]. élévateur [339]. Eleven [2177]. Eleventh [431, 2137]. Eleventh- [2137]. Eleventh-Century [431]. Elfenbeinturm [2747]. Eli [876, 1985, 597]. Elie [391]. Eliel [1134]. Elihu [1399]. Eliot [3168]. Elisabeth [1606, 2920, 2836, 1054, 1676]. Elisions [1076]. Elite [2035]. Elizabeth [2828, 2455, 1860, 195, 1709, 3336, 962, 1226, 2014, 1732, 3476, 815, 2298, 1391, 2844, 3012]. Elkana [108]. Elkhadem [1649, 2070]. Ellen [2110, 3320, 1709, 2049, 2299, 349, 2815, 65, 1578]. Ellenberger [3144, 3415]. Ellery [1060]. Elliot [2144]. Elliott [390, 686, 3390, 1264, 1962, 2555, 695, 2522]. ellipseos [2089]. Ellis [248, 1934, 696, 3571, 1209]. Ellmayrs [3256]. Ellrich [3536]. Ellsworth [1309]. Elly [2257]. Eloge [3302, 705, 2312, 1792, 2313, 3311, 101, 2946, 1889, 2482, 2626, 2625, 2634, 707, 223, 2400, 3312, 2955, 2715, 3500, 2401, 1793, 2954, 2956, 2624, 409, 614, 2483, 898, 1890, 312, 706, 2057, 1973, 1791, 3402]. Eloges [2747]. Elois [2648]. Else [2843]. Elshout [516, 996, 504]. Elsie [3015]. Elsky [1099]. Elsner [1113]. Elspeth [917]. Elton [960]. Elusive [2858]. Elvira [1828, 1505]. Elvire [1390]. Emancipation [3026]. Emanuel [876, 2313]. Emanuela [645]. Embassy [1189]. Emblazoned [2737]. Emblem [97]. Embodiment [3063]. Embodiments [3229]. Embracing [2869]. Embree [1239]. Embryo [918, 3376]. Embryology [1248, 1224]. Embryos [3399]. Emch [1568]. Emch-Deriaz [1568]. Emergence [1734, 1649, 958, 3566, 2282, 2053, 1141, 774, 171, 3569]. Emerges [2697]. Emerson [2372]. Emil [266, 267, 2280, 1572]. Emilie [346]. Emily [1192, 282]. émir [914]. Emmanuel [2275, 2510, 3380, 3147, 734]. Emotions [3307, 637]. Empedocles [2887]. Emphasis [2076]. Empire [418, 846, 1543, 3348, 2665, 759, 847, 332, 3258, 1229, 466, 1043, 250, 3348, 2499, 657, 2168, 130]. Empires [532]. Empirical [205]. Empiricism [1820, 883, 1333]. Empiricist [1421]. Empson [1817]. Empty [2799]. Emter [2920]. Enchantment [1960]. Enchiridion [2582]. Encounter [748, 3364, 1034, 2159, 2719]. Encounters [2913, 3414, 3142, 2159]. Encyclopaedia [818, 3394, 1375]. Encyclopedia [302, 3107, 2880, 1554, 1621]. Encyclopedias [2177, 406]. Encyclopedic [1154]. Encyclopédie [1571, 2177, 395]. Encyclopedists [395]. End [644, 225, 2855, 254]. Ende [2046, 2742, 822]. Endless [2541, 3366]. Ends [1310, 227, 136]. Energies [3424, 3518]. Energieubertragung [2033]. Energiewirtschaft [1599]. Energy [2148, 2206, 1859, 582, 239, 968, 56, 1768, 2213, 192, 965, 225, 1529, 1664, 170, 846, 2146]. Enfermedad [3154]. Engaging [2704]. Engel [260]. Engelen [1808]. Engelhard [1292]. Engine [667]. Engineer [1617, 1864, 1902, 2688, 1503, 3366, 197]. Engineering [281, 858, 2641, 2781, 1887, 604, 2060, 2823, 2688, 1039, 1509, 1005, 1598, 3443, 3176, 1456]. Engineers [2797, 2192, 1968, 198, 598, 945, 2302, 680]. Engines [3061]. England [30, 2287, 3295, 2899, 2751, 1012, 3152, 481, 765, 703, 2137, 1164, 1999, 1036, 1388, 555, 3448, 2739, 2654, 2096, 830, 268, 3045, 45, 2006, 764, 3507, 2822, 556, 2904, 646, 624, 1493, 2817, 2230]. Englebert [2438]. Englightenment [2269]. Englisch [1993]. English [1563, 39, 602, 29, 1099, 1012, 1118, 3142, 768, 3057, 44, 2337, 2598, 1565, 2278, 3491, 3164, 2360, 2152, 2087, 3133, 2810, 3335, 1367, 3430, 3353, 2398, 2808, 2346, 1477]. Englund [1987]. Enhancing [3576]. Enigma [3529, 2371, 166, 1642]. Enigmas [187]. Enjoying [1134]. Enlightenment [1564, 455, 1480, 1801, 2096, 3443, 2598, 2339, 3342, 2009, 2717, 1342, 43, 2433, 1387, 1469, 1726, 1119, 840, 560, 1568, 43, 1739, 3344]. Ennis [2458]. Enough [3285]. Enquête [2247]. Enquiries [2404]. Enquiry [2898]. Enrico [521, 2811, 3138, 811]. Ensayo [1040]. enseignement [657, 653, 2182]. ensenar [1374]. Ensign [2989]. Entangled [1405]. Entanglements [3143]. Entdeckung [2464, 1292]. Enter [1126]. Enterprise [1453, 565]. Enters [1191]. Entertainment [2009]. Entfernungen [2033]. Enthusiasm [629, 2662]. Entitled [2200]. Entomological [87]. Entomology [2454]. entre [687, 1571, 659, 729, 1026, 150]. Entropy [170, 76]. Entstehung [2018, 2757, 2679, 2465, 1587, 442]. Entwicklung [530, 3145, 174, 1588, 2755, 2994, 3253, 2359, 2033, 2923, 266]. Entwicklungslinien [1856]. Environment [180, 239, 1732, 245, 1867, 434, 2767, 1905]. Environmental [1806, 3097, 2574, 2930, 2738, 3325, 136, 2802, 2758, 381]. Environmentalism [1140, 2239]. Environmentalists [1904, 3108]. Envy [3476]. Epic [3529, 2932]. Epicurean [1383, 711]. epidemias [3154]. Epidemic [1707, 2653, 2538, 1933, 2279, 3153]. Epidemics [333, 3420, 962, 3077, 3124]. Epidemienbuch [2067]. Epidemienbüchern [549]. Epilogue [649]. Episodes [1453, 422]. Epistemic [3010]. Epistemological [2872, 1335]. épistémologique [545]. épistémologiques [1063]. Epistemology [789, 2262, 389]. Epistolicum [3534]. epistolier [2166]. Épitomé [2960, 1095]. epoque [2436, 270, 749]. Epsilon [1679]. Epstein [3124]. Equality [1245]. Equant [32]. Equations [349]. Equatorie [2162]. Equilibrium [913]. Equipment [3101]. Equivalence [1656]. Era [2368, 2100, 1080, 2205, 3235, 3024, 1739, 1067, 2910, 1087, 1413, 3539, 2721, 2673, 143, 1211]. Erasmus [363, 1822]. Erde [1635]. ère [535]. eresia [2277]. Eresus [1906, 2243]. Erfahrung [1531]. Erfinder [1842]. Erfindung [2827]. Erforschung [837]. Erganzungsband [2356]. Erhard [536]. Eric [3020, 2903, 1184, 589, 522, 1931, 3480, 705]. Erica [1384]. Erich [2119, 1542]. Erie [2912]. Erik [1371]. Erika [2620, 476]. Eriksson [514, 2007]. Erinnerung [2447]. Eriugena [726, 726]. Erkennen [1403].

Erkenntnis [1150, 259]. Erkenntnislehre [1914]. Erklarung [2070]. Erle [248]. erlebte [2447]. Erler [3049]. Ermenc [225]. ermete [2811]. Ermunterung [1830]. Erna [1373]. Ernan [1425]. Ernest [2967, 1134, 659, 522, 1677, 1243, 1865, 65, 1227]. Erneuerung [1993]. Ernst [2032, 2604, 748, 654, 2917, 1221, 1169, 442, 476, 758, 1747, 1941, 1544, 1098, 804, 1855]. eröffnete [453]. erreur [2820]. Errol [3573, 167]. Error [674, 2999]. erroribus [2183]. erste [2342]. ersten [2067, 1937]. Erstes [175]. erudit [2166]. Erudizione [2590]. Erwin [1313, 59, 586, 2463, 2033]. Erxleben [2353]. Erziehungssystem [1115]. escritores [746]. Esfeld [2977]. Esortazione [1002]. espaces [2573]. España [3445, 1331, 2824, 1723, 2158, 2444, 1389, 1024, 752, 261, 152, 2915, 868, 134]. español [921]. Española [1470, 3475, 524]. españolas [1911]. Espanya [1261]. Especially [1891]. espérance [2508]. Esperanza [3053]. esperienza [2001, 1199]. Espoirs [1580]. esprit [2658]. Essai [1042, 545]. Essay [1269, 879, 227, 133, 1035, 572, 3035, 666, 1236, 75, 737]. Essays [3104, 1308, 3094, 3512, 1543, 1800, 1241, 905, 1021, 426, 576, 3580, 284, 1849, 468, 2266, 1335, 1831, 1282, 1059, 191, 1675, 531, 1057, 1203, 2126, 679, 428, 661, 1802, 1850, 672, 1457, 649, 1531, 2463, 1536, 2504, 2365, 1270, 686, 977, 1817, 1112, 3090, 2727, 452]. Essence [2481]. essentiae [730]. Essential [1022]. est [2089]. Establishing [557]. Establishment [2446, 3497]. Establishments [2851]. estado [1389]. Estanislao [134]. Esteban [1723, 523, 1331, 1331]. Estes [3153, 633, 3345, 1153]. Estestvoznaniia [511]. Estrange [1802]. Estrella [2482]. estructura [262]. Estudio [1197]. Estudios [352, 1024]. eta [1471]. Etang [3538]. Etat [1931]. Etel [2686]. Eternally [946]. Eternity [342, 1007, 915, 551, 1720, 2500]. Ethel [2453]. ether [1847]. Etheridge [1860]. Ethical [1548, 1709, 968]. Ethics [719, 2659, 1021, 1978, 3176, 2868, 468, 1618, 2505, 381, 1978, 2190]. Ethik [758]. Ethnobiology [2648]. Ethnographer [1850]. Ethnographic [721]. Ethnography [2921]. Ethnomethodology [2552]. ethologie [2680]. Ethos [1057]. Etienne [843]. Etner [271]. étoiles [1653]. etonnant [3147]. Ettner [162]. etude [2436]. Etudes [320, 529]. Euclid [2075, 2731, 1367, 3401, 2731]. Euclide [1176, 2242]. euclidiennes [149]. Eugen [266, 3360]. Eugene [3006, 759, 1456, 1758, 2672, 734, 914, 368]. Eugeni [262]. Eugenic [273]. Eugenics [2546, 2695, 1951, 2611, 1143, 3580, 3098, 956, 1326, 957]. Eugenik [2914, 196]. Euler [2903, 3534, 2174, 165, 842]. Euleri [1738]. Eunice [3047]. Eureka [635]. Europa [2355, 1304, 1911]. Europaea [1819]. europäische [1115]. Europe [718, 1649, 322, 540, 3522, 1034, 326, 3336, 3075, 2921, 3008, 1632, 1068, 1812, 3081, 3544, 54, 905, 2053, 3026, 2189, 3243, 2163, 1752, 3337, 445, 2002, 3526]. European [1891, 264, 1325, 2408, 918, 600, 814, 3392, 2921, 1539, 1286, 3386, 1383, 2722, 151, 1348, 332, 1750, 3264, 3077, 1797, 2159]. europeo [251]. Euthanasia [2204]. Eva [2032, 2604, 1808, 1572, 212, 2683]. Eva-Maria [1808]. Eva-Marie [1572]. Evaluated [1146]. Evaluation [3426]. Evan [1739, 384]. Evangeliou [547]. Evans [1324, 2989, 1744, 1719, 1298, 779].

Evax [146]. Eve [3416, 1464, 3232]. Evelleen [1232]. Evelyn [1606, 2608, 762]. Evelynn [3037]. Event [1255]. Everdell [2997]. Everett [118]. Everyday [418]. Everyone [2044]. Everything [299, 3547]. Eviatar [1648]. Evidence [2308, 2859, 1382, 2236]. Evidencia [1654]. Evil [3236]. Evo [251]. Evolucionismo [2824]. évolutifs [1063]. Evolution [2684, 3549, 1434, 3310, 2035, 1983, 243, 390, 1087, 3403, 468, 387, 1984, 3390, 283, 218, 1541, 177, 203, 1489, 1592, 467, 2126, 2208, 125, 661, 1496, 709, 1713, 674, 194, 284, 907, 76, 468, 729]. Evolutionary [709, 1606, 49, 1162, 3160, 3389, 1169, 2017, 2916, 612, 1668, 3450, 1674, 7, 1784, 2190, 1170]. evolutionnisme [3555, 1570]. evropy [887]. ex [1637]. Exact [130, 1567, 112, 2266, 1670, 1167, 3251, 531]. exakten [1993, 654]. Examination [3050, 3528, 1422]. Excavating [3386]. Excavations [3163]. Excellence [135]. Excerpts [225]. Exchange [2163, 3318, 3526]. Excluded [2868]. Executed [1864]. exégétique [914]. exemple [653]. Exercise [946, 2511]. Exhibiting [2966]. Exhibition [21, 2651]. Exile [3498]. Exiles [3532]. Existence [1170]. Exons [1144]. Exotic [151, 1944]. Expand [1852]. Expansion [1670, 2239, 1318, 332]. Expansive [2876]. Expedicion [2517, 1561]. Expedition [2989, 961, 3348, 3180, 893]. Expeditions [1255]. Experience [2570, 1543, 429, 1838, 1834, 1027, 1087, 2323, 332, 1211, 557, 2514, 3284, 1686, 3106]. experiencia [2158]. Experiencing [3511]. Experiment [417, 1193, 3307, 317, 3011, 173, 2170, 926, 881, 2837, 881, 1703]. Experimental [1193, 99, 2316, 2631, 489, 612, 2999, 3026, 1505, 1703]. experimentale [1496]. Experimentation [66, 1871, 2612, 1703, 2937]. experimentele [2292]. experimentelle [394]. Experimenter [2712]. Experimenting [2873]. Experiments [309, 2227, 1513, 1924, 1514, 792, 2111, 2054, 1513, 1407, 1822]. Expert [75, 1329]. Expertise [2641]. Experts [975, 2299]. Explaining [2294, 2649]. Explanation [1891, 748, 879, 785, 77, 783, 783]. Explanations [2859]. explicacion [1374]. explication [2658]. Exploitation [678]. exploración [1084]. Exploration [583, 1357, 847, 2823, 1689, 3328, 1807, 63, 2685]. Explorations [1898, 1134, 640]. Explorer [1699]. Explorers [2729]. Exploring [323, 3425, 2321, 1171, 1289, 2685]. exposicion [2103]. Exposition [2897, 1182]. Expositions [1844]. Exposures [1145]. Expression [1207]. Exquisite [3093]. Extent [131]. Extinction [2043]. Extrabiblical [2236]. Extraordinary [773]. Extraterrestrial [2771]. Eye [1894, 2187, 3266, 3085, 3035, 1456]. Eye-Opening [3085]. Eyes [335]. Eyler [3078]. Ezawa [2537]. Ezell [304]. Ezio [2097]. Ezra [1098]. Ezrahi [1330].

F [3361, 2453, 2180, 1008, 2843, 734, 997, 1313, 52, 554, 1295, 2408, 1656, 58, 1617, 2516, 15, 2885, 2579, 3558, 3108, 110, 74, 1896, 3086, 111, 2441, 1517, 1688, 1667, 1103, 1737, 185, 1174, 849, 569, 2192, 996, 3167, 440, 1499, 604, 2419, 1565, 681, 2809, 2111, 1637, 3036, 532, 1991, 2990, 2456, 3521, 1862, 1749, 2706, 966, 131, 1039, 3174, 2844, 1229, 1273, 1261, 757, 2253, 2991, 518, 61, 1602, 1802, 60, 753, 2814, 2943, 672]. F.R.S [2905, 1825]. Faas [665].

Faasse [2452]. fabbrica [2586]. Fabian [358]. Fabienne [1580]. Fabio [327, 3129]. Fabri [1165]. Fabric [1539, 3059]. Fabrication [788]. Fabricius [2901]. Fabrics [2971]. Fabrikant [1212]. Fabrizio [3000]. Facade [287]. face [1201, 2361]. Faces [1380]. Facets [3039]. Fachgebietes [2465]. Fachglossar [1638]. Fachwissenschaft [1316]. Fachwortschatz [1637]. Facial [1207]. Faciles [1361]. Facolta [1821]. Facsimile [1293]. Facsimiles [1816]. Fact [3470, 2104]. Factories [2210]. Factors [3473]. Factory [3286, 872, 2630]. Facts [2314, 2410]. Fagan [3334]. Fagette [2929]. Fague [3459]. Faig [134]. Fail [3288]. Failure [1136]. Fairfield [1590]. Fairs [1844]. Faith [3039, 2601]. Falaquera [433]. Fall [1864, 2641, 2042, 1289, 73, 649]. Fallen [2899]. Fallout [3364]. Fallstudie [1932, 2614]. False [1020, 71]. Falzeder [2032, 2604]. Familientreffen [3586]. Family [3314, 867, 273]. Famous [1557]. fantasma [2518]. Fantasy [288]. Fantini [505]. Fantoli [2084]. Fanton [966]. Far [3218]. Fara [2817]. Farabi [1630]. Faraday [1938, 1532]. Farben [1053]. Farber [2190]. Farbzentrenforschung [585]. Fardon [721]. Farelly [1555]. Faria [793]. Farkas [424]. Farley [1981, 1229]. Farm [2461, 1679, 2200]. farmacia [2444]. Farn [1767]. Farnes [626]. Farquhar [3207, 3014]. Farrall [684]. Farrell [2700, 2156]. Fasol [1130]. Fasol-Boltzmann [1130]. Fatal [2501]. Fate [905]. Fath [1810]. Fathers [2748]. Fatigue [1413]. Fattori [1093]. Faubourg [3147]. Faulhaber [3338]. Faulkner [2003]. Fault [43]. Faust [2409]. Fauth [2123, 2123]. Fauvel [37]. favola [1589]. Fawcett [331]. Faye [1012, 1216]. FDR [1507]. Fe [1614]. Fear [225]. Featherstonhaugh [176]. febbre [2421]. Febrer [1911, 2158, 752]. febres [39]. February [705, 2626, 223, 372, 2715, 409]. Fechner [51]. Fechners [2107, 51]. Fed [137]. Federal [2212, 95, 95]. Federici [919]. Federico [853, 1654]. Fedorovic [861]. Fee [195, 3012]. Feeble [2115, 2699]. Feeble-Minded [2699]. Feeding [1681]. Feeling [2104]. Feher [237, 2941]. Fehlinterpretationen [1314]. Feider [791]. Feinberg [2995, 384]. Feingold [741]. Felber [2185]. Feldberg [2468]. Feldenkirchen [1842]. Feldhay [3437]. Feldman [1307]. Feldmessern [989]. Felipe [1561]. Felix [2282, 442]. Fellmann [3534]. Fellows [2171]. Fellowship [2641]. Fels [1537]. Felschow [1572]. Felt [3566]. Female [1528, 2486]. Feminine [1343]. Femininity [2846]. Feminism [3084, 2986, 3292]. Feminist [883, 2047]. feministischen [2620]. Femmes [1564]. Fences [703]. Fennicae [513]. Ferdinand [654, 1819, 2112, 175, 3551]. Ferejohn [1179]. Ferenc [1710]. Ferenczi [2032, 2604, 2032]. Ferguson [1456]. Feride [1245]. Fermat [2592, 3529]. Fermi [771, 225]. Fernand [1105, 1092, 1653]. Fernandez [2090, 1389]. Fernandez-Carrion [2090]. Fernando [868, 2375]. Fernsehempfanger [3385]. Ferranti [1887, 1957]. Ferranti-Packard [1957]. Ferrara [1267, 1920]. Ferrari [2893]. Ferraro [2340]. Ferrone [2269]. Ferroni [2443]. Ferry [2554]. Fertility [1087]. Feshbach [1404]. Fest [1463]. Festival [2080]. Festkörperphysik [585]. Festschrift [1941]. Fetal [3126]. Fettweis [1120]. Feuers [174]. Fever [2108, 1934, 3153, 795, 1035, 376, 963, 1229]. Fevered [3162]. Feyerabend [2390, 2318, 1241, 2318]. Feynman [2301, 2381]. fi [1718, 1991]. fiammiferi [3268]. Fiasco [1959]. Fichant [2508]. Fichman [1980]. Fichtner [1069]. Ficino [3249, 2429, 2501]. Fickers [3385]. Fiction [722, 543, 2524, 1576, 632, 3005, 1357]. Fictions [2524]. fidei [1916]. Field [858, 2197, 2918, 3008, 1352, 2002, 2990, 2821, 265, 3055, 3566, 2881]. Fields [2719, 1242, 3510, 2785, 1757]. Fieldwork [941, 2408, 3281]. Fifteenth [600]. Fifth [3334, 2042, 2420, 1330]. Fifth-Century [2420]. Fifty [3211, 2039, 959]. Fighting [3061]. Figueroa [2681]. figure [535]. Figures [2897, 3412]. Filho [2354]. filius [1640]. Film [397]. Films [1445]. filologo [2893]. filosofessa [613]. Filosofi [1643, 1915]. Filosofia [3545, 1912, 1095, 3054, 2000, 2890, 2590, 1197, 2763, 1826, 3129, 1183]. filosofica [1821]. filozofije [930]. Fin [2189, 856, 3026]. Fin-de-Siècle [2189, 3026]. Final [1321, 1866, 1789, 73, 1594]. finales [2444]. Financial [2693]. Findlen [2086]. Findley [2685]. Fine [3453, 2445, 1912]. Finger [2138, 1898]. Finishing [1428]. finitur [1134]. Fink [1935]. Finland [1765]. Finley [3581]. Finn [949, 2909]. Finney [1942]. Finnish [513]. Finocchiaro [298, 1017, 3436]. Fiocca [2271]. Fiona [2597]. Fire [2169, 3061, 2372, 2961, 3035, 3343]. Firenze [2590, 124]. Fires [3061]. First [2678, 1696, 2909, 1789, 2205, 62, 451, 2993, 3244, 1409, 583, 2197, 3156, 2577, 3579, 533, 2997, 1293, 176, 2274, 2458, 2776, 2770, 368, 2039]. Fisch [1488, 1305]. Fischer [1221, 267, 3441, 1843, 2248, 394, 1501]. Fisheries [2106, 2670]. Fishes [2442]. Fishing [2106]. fisica [3001, 1095, 811, 327, 860]. fisiche [2980]. fisiologia [1199]. Fissell [1661]. Fission [2607]. Fits [2267, 1324, 2267]. Fittkau [1941]. Fitzgerald [675]. Fitzhugh [594]. Five [1702, 1896, 583, 1990, 314]. Fix [840]. Fixed [905]. Fixer [2205]. Fixsternhimmels [998]. Flagella [2310]. Flail [3204]. Flamel [2897]. Flamm [2457]. Flanagan [434, 2889]. Flannery [1001]. Flashar [3049]. Flat [1260]. Flattening [1702]. flebotomia [438]. Fleck [2765]. Fleeting [1712]. Flegg [901]. Fleming [3094, 1887, 942, 888]. Fletcher [573, 2201, 1147]. Fleury [1808]. Flexner [579]. Flight [3482, 2637, 1442, 3587]. Floistad [1280]. Flood [37, 2236, 631]. Flora [2751, 1750]. Florence [1922, 182]. Florida [3178, 221, 2455]. Florilegium [1721]. Florin [2913]. Floris [516]. Flourens [1496]. Floyd [2111]. Fludd [644]. fluidos [3440]. Fluoridation [1056]. Flussig [1463]. Fly [2316]. Flying [869]. Foley [1924]. Folgen [2013]. Folgerns [2153]. Folin [62]. Folkerts [2973, 1778, 518]. Folklore [991]. folletos [2491]. Fondateurs [2877]. fondation [2618]. fondazione [2347]. Fondos [1690]. Fontaine [3538]. Fontana [1623]. Fontanon [2132]. Fontenelle [1109, 2437]. Food [2212, 3425, 670]. Footnote [3409]. Footnotes [2408]. Footsteps [1712, 3355]. Foppes [2515]. Forbes [757]. Forbidden [2845, 3471]. Force [3210, 1112, 2042, 3008, 831, 6, 2772]. Forces [3166]. Forderung [3367, 1215]. Foreest [1481]. Forelock [3259]. Forestry [330, 2972]. Forests [2214, 537]. Foretelling [2398]. Forge [450, 3438]. Forgers [714]. Forgery [953, 953]. Forgetfulness [2808]. Forging [1887, 2396, 225]. Forgotten [1332]. Form [2094, 2418, 3057, 2364, 168]. forma [2270]. formación [262]. formaler [859]. Forman [2851]. Formation [3025, 2698, 421, 3526, 2758, 2514]. Formative [3452]. forme [1826, 2001]. Formed [3017]. Former [1600]. Forms [2805]. Formulation [2542]. Forrest [282]. Forschung [2033, 81, 1410, 2256, 1215]. Forschungen [2153]. Forschungsstelle [2923]. Forster [1458]. Fortenbaugh [1906, 1550]. Fortescue [1583, 55]. Fortifications [2080]. Fortino [3457]. Fortis [2750]. fortuna [1560]. Fortunes [804]. Forum [3206]. Forward [3064]. Fossil [3354]. Fossilien [358]. Foster [1536, 1052, 2112]. Fothergill [3259]. Foucault [694, 3306, 485, 1242]. Fought [3081]. Found [3469]. Foundation [1593, 1415, 320]. Foundations [1548, 3276, 951, 2806, 616, 688, 1945, 976, 1621, 1240, 1227, 71, 2597, 2721, 1320]. Founders [3553, 1149, 1674]. Founding [2748]. Fountain [2766]. Four [1873, 2077, 3470, 1248, 1689, 1990, 732, 1676, 783]. Fournier [229, 127, 3059]. Fourteenth [1279, 1639, 2500]. Fourteenth-Century [1639, 2500]. Fourth [1984]. Fox [1606, 2622, 2608, 174, 3251, 1041, 195, 3040, 954]. Foxton [435]. Fraassen [1421, 880]. Fracastoro [2811]. Fracture [1188]. Fragments [237]. Frago [154]. Frame [905, 1492]. Framework [3485]. Franca [3350]. Francais [3189, 2182, 1201]. francaise [1043, 749, 914, 2102]. France [1484, 2812, 1324, 3272, 49, 270, 2514, 2285, 1580, 843, 1209, 3443, 2432, 900, 2652, 470, 1659, 263, 2014, 1054, 2823, 957, 2414, 3433, 1312, 745, 653, 245, 1234, 212, 3222, 2710, 3043, 2926, 751, 3557, 1208, 1897, 1504, 974, 2109, 1376, 1037, 1754, 271, 3344, 1129, 1758]. Frances [826]. Francesc [3523]. Francesca [2971, 1555]. Francesco [3350, 1101, 2063, 1779, 566, 1267, 1821]. francese [1288]. Francie [2155]. Francine [2456]. Francis [2179, 2346, 3164, 434, 2121, 2164, 3250, 3435, 3225, 2003, 948, 924, 3374, 3135, 1652, 2740, 2003]. franciscanos [746]. Francisco [3104, 554, 261, 2158, 351, 752, 2175]. Franciscus [1185]. Franck [1969]. Francois [1874, 3460, 3415, 1580, 3298, 1893, 1546, 1559, 1571, 2320, 3144, 1820, 2573, 2539, 3103, 1035, 320, 660, 271]. Francoise [3458, 1847, 981]. Frangsmyr [2440, 1739, 841, 618]. Frank [2839, 2177, 1461, 2880, 1733, 1220, 3173, 2655, 1831, 1104, 1938, 2002, 395, 697, 953, 2127, 757]. Frankenstein [1440, 3005, 3355]. Frankland [2822]. Franklin [396, 934, 2748, 1703, 2042]. Frankreich [2509]. Franques [3444]. Franz [2464, 864, 12, 3256, 2610, 2992]. Franzosischen [3552, 461]. Fraser [15, 3081]. Fraud [2227, 1511]. Frauen [2620]. Frauenleben [3278]. Frauenliga [3567]. Frazier [275, 381]. Fred [197, 1317, 542, 1333]. Freddy [2123]. Frederic [3255, 1466, 565, 1977, 2191]. Frederica [480]. Frederick [2853, 852, 2449, 2752, 1597, 3479, 691, 269, 1748, 2767, 1755, 3263, 168, 1181, 2992, 753]. Free [2898, 3216, 1289, 649, 1083]. Freedom [2659]. Freeland [2838]. Freeman [1395, 274, 2211, 2211, 2683]. Frege [2911, 1419, 1485]. Freiberg [2185]. Freiburg [3360]. Freiheit [1304]. French [1564, 1324, 455, 3267, 353, 2507, 3536, 2602, 181, 900, 1499, 2357, 3337, 2735, 1376, 1757, 1670, 680, 560, 1944, 3439, 652, 1978, 2886, 3495, 3337]. Frenkel [2843, 2835, 2715, 2124, 1852]. Frenkel-Brunswik [2843]. Frères [823]. Fresquet [1911, 752, 2158]. Freud [52, 2032, 2604, 1321, 2193, 879, 3417, 288, 854, 943, 1677, 1758, 499, 2700, 2700, 3582, 1146, 1228, 1845, 1948, 1069]. Freud-Bibliographie [1069]. Freudenthal [1289, 2418, 1639, 108, 1057, 529]. Freundlich [1313]. Freyer [2647]. Friar [2735]. Frias [3154, 2517]. Fribergensia [2185]. Friction [2053]. Fridolf [1011]. Friedberg [3573]. Friedel [21]. Friederike [1474, 354]. Friedlander [2553]. Friedlieb [654, 175]. Friedman [435, 1567, 193, 867, 117]. Friedmann [1852]. Friedrich [1481, 2356, 2183, 1474, 463, 167, 3546, 2013, 2185, 1029, 379, 2884, 3447, 2754, 2988, 3149]. Friend [178]. Friendly [1404]. Friends [1697]. Friendship [1697]. Fries [753, 1393]. Fries-Apelt-Schleiden [1393]. Fringes [2553]. Fritz [517, 266, 1757, 350, 3391, 3575, 2008, 518, 3080, 2122, 3568, 2533]. Frog [1827]. Frommel [2080]. Front [3, 4, 94, 216, 307, 403, 404, 496, 609, 701, 801, 891, 984, 1072, 1158, 1161, 1252, 1341, 1431, 1523, 1526, 1610, 1694, 1785, 1882, 1885, 1966, 2052, 2135, 2220, 2223, 2306, 2395, 2478, 2559, 2562, 2629, 2709, 2791, 2864, 2867, 2949, 3023, 3111, 3195, 3202, 3203, 3305, 3397, 3494, 3594, 1502]. Frontier [632, 3008, 3074, 2541, 3261, 73, 3366]. Frontiers [82, 2697, 1169, 3280]. Frost [260, 1234]. Frozen [2105]. FRS [1041]. Frugard [2074]. fruhen [1923, 1993, 3360]. Fruhgeschichte [2615]. Fruhhumanismus [1813]. frühmoderne [1115]. Fruhwerk [940]. Fruited [2802]. Fruton [3590, 1031]. Fryer [1557]. FSA [1041]. Fuat [510]. Fuchs [3003, 225, 2507]. fuentes [746]. Fujimoto [2537]. Fujimura [507, 1506, 3290]. fulgurante [1032]. Fuller [2938, 3184, 1772, 389, 2286]. Fulton [3207]. Function [1898]. Functioning [153]. Fundable [3092]. Fundament [654, 1398]. Fundamental [1336, 2198]. fundamental’nykh [1336]. Fundamentals [1673]. Funding [2130]. Funeral [3131]. Funkenstein [3402]. Funktion [2336]. funzioni [3138]. Furchtbare [2382]. furchtet [2978]. Furley [915, 426]. Furman [581]. Furst [1245, 1029]. Furst-Dilic [1245]. Furstenstaat [3341]. Furthest [3123]. Furukawa [902, 3569]. Fusion [1959, 1959, 1220]. Future [2398, 3064, 872, 1900, 2061]. Futures [2721]. Fye [2934]. Fysica [2515].

G [2453, 1859, 3284, 39, 3414, 52, 2386, 1048, 487, 160, 3516, 2622, 3479, 940, 1066, 570, 240, 66, 2385, 582, 3153, 918, 2355, 2745, 2183, 3534, 2969, 2193, 2723, 735, 3183, 1053, 3511, 2965, 1502, 1804, 361, 1087, 2849, 103, 1364, 3056, 626, 1720, 1906, 3484, 257, 28, 2006, 3506, 1335, 808, 1719, 1376, 2761, 1379, 141, 742, 922, 1781, 2999, 3584, 1133, 2323, 964, 2966, 634, 43, 637, 125, 2802, 1207, 1760, 2760, 630, 3462, 2292, 2003, 2878, 3366, 2575, 3171, 2065, 1802, 1905, 1531, 3185, 672, 999]. G. [161, 2173, 1953, 51, 1110]. G.-B [1207]. Gabriel [3415, 463, 1454, 457, 3122, 1026, 54]. Gabriele [2001, 3150]. Gad [2418, 1639, 108, 1057, 529]. Gage [232]. Gainesville [221]. Gal [1473]. Galápagos [50]. Galaxien [2411]. Galaxies [3082]. Galaxy [2119]. Gale [1411, 2530, 3530, 954]. Galen [3240, 1716, 1011, 1271, 1002, 1629, 1011]. galenistischen [1481]. Galeno [1556]. Galenus [1716]. Galien [2888]. Galilei [2445, 1103, 831, 3436, 1019, 2741, 1473, 1653, 2259]. galileiana [1925]. galileiani [3139]. Galileo [1017, 2590, 1922, 2084, 2589, 1103, 2650, 831, 3436, 1019, 2167, 2741, 1289, 1473, 1653, 649, 97, 807, 8, 311, 2713, 3499, 892, 649, 1018, 1171, 2976, 2084, 3437, 738, 1616, 3436, 1818, 1016, 2005]. Galison [2718, 1328]. Gallagher [962]. Gallart [3445]. Galliher [2844, 2844]. Galloway [627]. Galluzzi [521]. Galvani [1827]. Galvani-Volta [1827]. Gamble [2296]. Gambling [1619]. Game [1869]. Games [284, 3131]. Gamst [2992]. Gamwell [2470]. Gandt [1546]. Gans [35]. Gap [1134]. Garber [3387, 1531, 1287]. Garcia [791, 2249, 2090, 868, 154, 921]. Garden [1607]. Gardening [1467, 3274]. Gardens [1467]. Gardies [545]. Garding [3071]. Gareth [1091]. Garrett [223, 2593]. Garson [872]. Gary [694, 1553, 429, 3064, 1959, 2540, 356, 3161, 937, 1154]. Gas [2466, 2825]. Gascoigne [1418, 455, 2598, 3127]. Gassendi [2261, 1820, 1559, 2659]. gassendistes [1559]. Gasser [2188]. Gaston [1065]. Gastrointestinal [2648]. Gate [225]. Gatekeeper [3155]. Gatti [555]. Gatto [2260]. Gatzemeier [463]. Gaudart [1245]. Gauge [2919]. Gaukroger [737, 1196, 1020, 2430]. Gauld [1744]. Gauss [2183, 1029]. Gautier [435, 2361]. Gavroglu [488, 1214, 2533]. Gawlick [3049]. Gay [1963]. Gayle [792]. Gebäude [654]. Geber [1465]. Gebiet [2033]. Geburtsstadt [3551]. Gedenkveranstaltung [1120]. Gedichte [1474]. Gee [3574]. geeigneten [174]. Geertz [3470]. Gegenwart [812, 2161]. gehalten [1403]. Gehennical [2169]. Geiger [1585]. Geison [1977, 2317]. Geissler [578]. Geisteswissenschaft [2238]. Gelbart [3346]. geldelijke [2693]. geleerde [2814]. Gelehrten [475]. Gelehrtenschicksal [3003]. Gelehrter [2988]. Gelfand [2762]. geliebter [2457]. Gellner [65]. Gemeinschaft [1830]. Gemistus [3524]. Gemology [3525, 1780]. Gender [3317, 2433, 623, 1326, 478, 1245, 765, 2298, 2513, 2325, 624, 2971, 3037, 3452]. Gendron [2363]. Gene [715, 196, 1674, 590, 1618]. Genealogical [1242]. Genealogy [719, 1324]. Geneeskunde [516, 994, 240, 2335]. Geneeskundigen [2693]. General [769, 1673, 1399, 815, 225, 125, 1041]. générale [1580, 2098, 1176]. Generation [1913, 2307, 3063, 2902]. generatione [727, 2892]. Generationenkonflikt [1932]. generazione [1199]. Genes [1144, 1949]. Genese [2818, 660]. Genesi [2518]. Genesim [1366]. Genesis [629, 1169, 3253, 3440, 1613]. Genethics [238]. Genetic [959, 2781]. genetica [853]. Geneticist [2909]. Genetics [2316, 3355, 3290, 1674, 238, 2534, 86, 2208, 2294, 2494, 958, 1334, 1265, 2015]. genetisch [3375]. Geneva [2344, 60]. Genève [1026]. Genevieve [3339, 1476, 1559, 2031]. génèvois [322]. génèvoise [1201]. Genie [3479, 1762, 1032]. Genius [1213, 2815, 1380, 2112, 2986, 3238, 416, 1581, 3271]. Gennady [2124]. Genocide [1594]. Genome [1144]. Genootschap [994]. Gent [2814]. gente [3445]. Gentle [1750]. Gentleman [3270, 1836]. Gentlemen [3025]. Genuth [3119]. Geo [259]. Geoffrey [2433, 1617, 2616, 3275, 598, 3174]. Geograficas [2813]. Geograph [2988]. Geographic [85]. Geographical [2962, 1453, 3258, 3400]. geographie [1351]. géographique [388]. Geography [627, 140, 703, 1357, 1082, 1154, 537]. geologia [2750, 134]. Geological [3354, 941, 88, 620, 2753, 3353, 265]. geologie [3415, 3144, 1026]. Geologist [2521, 3281, 1836, 2186, 176]. Geologists [1970, 3141, 2788]. Geology [2788, 775, 48, 1613, 2798, 1454, 1741, 3353]. geomagnetici [2963]. Geometers [3115]. geometres [1931]. Geometria [437, 1925]. geometriae [1181]. Geometric [723, 3143]. Geometrical [3242]. Geometrie [1531, 1176, 2437, 2495, 3056, 1809]. Geometrie-Zahlentheorie-Algebra [1531]. géométries [149]. geometrique [2508]. Geometry [719, 2075, 1464, 3029, 268, 3028, 2651, 1064, 1181, 3401, 723]. Geomorphology [813]. Geophysics [2616]. Georg [1932, 3273, 827, 260, 2413, 2253, 1300, 461]. George [202, 2453, 1308, 1371, 437, 2521, 655, 207, 3108, 967, 2084, 1983, 2169, 3175, 1118, 2602, 243, 468, 2655, 120, 1386, 1348, 2245, 1370, 1359, 32, 2276, 1203, 293, 428, 1850, 1833, 3524, 176, 3076, 1261, 456]. Georges [3460, 1657, 331, 1893, 693, 2324, 3354, 1518]. Georgette [1496]. Georgiana [3356]. Georgina [2468]. Geoscience [776]. Gerabek [2754]. Gerald [579, 1513, 1148, 2317, 1452, 3099, 1231, 117, 1977, 1854, 1142]. Geraldine [1133, 3282]. Gerard [601, 1324, 848, 1713, 2619, 1466, 3158, 3422, 1307, 1802, 821]. Gerardi [248]. Gerber [1502]. Gerbert [1808]. Gerd [418, 1360]. Gerda [2178]. Gergonne [3446]. Gerhard [325, 2252, 89, 2988, 90, 1069, 3091]. Gerhards [1360]. Gericke [989]. Gerit [3567]. Gerl [1013]. Germaine [2373]. German [1951, 2386, 3563, 2922, 2779, 3151, 1945, 1755, 759, 1190, 2534, 2364, 3046, 1315, 945, 1198, 1211, 1600, 1952, 2987, 676, 3024, 873, 1315]. Germania [1821]. Germano [564]. Germany [2289, 3151, 258, 3260, 2184, 2024, 957, 2728, 2414, 143, 2204, 3112, 2952, 1857, 1028, 1678, 944, 678]. Germs [1949]. Gernot [2046, 360]. Gerontological [2698]. Gerontology [2697]. Gerrit [1718, 1534, 2808]. Gerson [920, 1000]. Gersonide [1466]. Gersonides [1639]. Gerstner [2992, 2186]. Gert [1304]. Gerta [2353]. Gertrud [3567]. Gesammelte [2620, 1474, 473, 354]. Gesamthochschule [578]. Geschichte [1322, 2464, 517, 3367, 585, 2462, 266, 3118, 2679, 3375, 3256, 1215, 2300, 3327, 1941, 2614, 510, 859, 2923, 196, 2378, 1215, 837, 471, 12, 2884, 1588, 1701, 812]. Geschiedenis [240, 516, 2975, 1535, 2693]. geschiendenis [994]. Gesellschaft [3367, 160, 2462, 1215, 1316, 1588, 1215, 1120]. Geseze [2013]. Gesner [1284]. Gesnerus [520]. Gesprächen [1029]. Gessner [42]. Gessners [42]. Gestalten [3241]. Gesteine [358]. Gesuiti [2000]. gesuitico [2260]. Gesundheitsbeamter [3447]. Get [2797, 284]. Getting [2826]. Getz [1012]. Geuter [1678, 286]. gevolgen [2693]. Geyer [1978]. Geyer-Kordesch [1978]. Gezondheidszorg [1733]. Ghayat [33]. Gheverghese [1348]. Ghiselin [3104, 2908]. Ghislaine [1369]. Ghost [1864]. Giacomoni [1826]. Giambattista [2271]. Giampieri [2032, 2604]. Giampieri-Deutsch [2032, 2604]. Gian [3000]. Gian-Carlo [3000]. Giancarlo [443]. Gianfranco [2347]. Gianni [2241]. Giant [672]. Giants [1820, 839]. Gibson [382, 3269, 2164]. Gideon [3074, 1289]. Gieben [1277]. Gieryn [1517]. Gifford [3380]. gigantes [352]. Gigerenzer [418]. Gil [438, 1917]. Gil-Sotres [438]. Gilain [1484]. Gilbert [1224, 3362]. Gilbertus [1012]. Gilded [1836]. Gill [3367, 2065, 1169, 546]. Gillespie [1407]. Gillian [2719]. Gillies [1873, 3188]. Gillispie [935, 3251, 1613]. Gilly [2342]. Gilman [2493, 3038]. Gilson [2465]. Gimbel [678]. Gindikin [17]. Gingerich [448, 1894, 1139]. Gingras [229, 1133]. Ginneken [1401]. Gino [3054, 454]. Ginsburg [1134]. Ginzberg [876]. Giordano [555, 1560, 2338, 2895, 2585, 1189, 3245]. Giorgio [2587, 913, 521]. Giotto [1464]. Giovan [2744, 1102]. Giovanna [600, 2893]. Giovanni [3104, 2338, 2895, 2063, 2097]. Girard [349]. Girls [3474]. Girolamo [3088, 2811]. Giron [2824]. Gispen [945]. Gispert [1484]. Gist [3574]. Giuliana [1107]. Giuliani [3001]. Giuliano [1169]. Giulio [1737]. Giuseppe [3001, 564, 2063, 1471]. Giusti [521]. Giustina [2155]. Giving [3399]. Gjertsen [720]. Glaciologist [757]. Gladly [803]. Glandular [3026]. Glaser [453]. Glasmacher [1393]. Glass [697, 3456, 86]. Glauser [648]. Gleichschaltung [1142]. Glen [2043]. Glenn [199, 1768, 2540]. Glennan [1863, 1863]. Glenys [3367]. Glick [1261, 58, 2516]. Glisson [2346]. Global [3588]. Globes [346]. Gloeilampenfabrieken [386]. Gloria [1662]. Glossae [1366]. Glover [1924]. Glucksspieltheorie [3535]. Glyndwr [3142]. Go [3218]. God [3249, 2429, 2416, 1866, 534, 3115, 444, 2325, 2889, 2431, 1000, 1983, 2022]. Goddard [3377]. Godefroid [3330]. Godel [2998]. Godelschen [2679]. Gods [1177, 1820, 3363]. Goerke [1930]. Goethe [2238, 1343, 1826, 2752, 364, 2448, 1935]. Goetz [2762]. Gohau [1454, 3415, 3122, 1026]. Goldammer [1283]. Gol’danskii [3090]. Goldberg [2954]. Golden [3463, 2542, 733, 2245, 1458, 1584, 672, 3554, 815]. Goldgar [2351]. Goldhammer [1893, 693, 1035]. Goldie [182]. Goldin [2649]. Goldman [3326, 872, 205, 225]. Goldschmidt [967]. Goldstein [920, 2471, 181, 1095, 2592, 531, 1503]. Goldsworthy [3270]. Golem [2044]. Golinski [1480]. Golland [2199]. Golley [2839, 2127]. Golub [2882]. Golvers [1819]. Golvin [935]. Gomez [3139]. Gone [583]. Gonzalez [1374, 1389]. Gonzalo [1241]. Good [703, 576, 703]. Goode [1203, 1203]. Gooding [417, 881]. Goodman [1910]. Goodrick [1022]. Goodrick-Clarke [1022]. Goodstein [1218]. Goodyear [2453]. Gordon [2521, 2622, 591, 3081, 1133, 225, 3076, 428, 2753, 2802, 1320, 1426, 3180, 2576]. Gorelik [2124]. Gorler [3049]. Gorman [1605]. Goroff [1756]. Gosling [52]. Gospel [218]. Gotffried [463]. Gottfried [2661, 2508, 837, 2089, 3137, 2265, 1531, 30, 1531]. Göttingen [471, 324, 1403]. Gottinger [1660, 1500, 90]. gottliche [1283]. Gottlieb [461, 1566]. Goubert [628]. Goudaroulis [488, 1214]. Gough [2540]. Gouk [34]. Gould [2192, 2613]. Goupil [2275]. Gout [3516]. Govern [1771]. Government [3577, 1649, 968, 176, 1136, 1309]. Governments [1584]. Gowing [3312, 576]. Goyanes [2656]. grabado [132]. Grace [2165]. Gracia [154]. Gradmann [2415]. Gradual [2794]. Graetzer [1502]. Graf [1724]. Grafe [1505]. Grafton [1687, 3409, 714, 923, 1815]. Graham [1708, 3454, 901, 3250, 1864, 2327, 1803]. Grahame [1620, 686]. Grain [3388]. Grains [2496]. Grammar [360, 2153]. Gramsci [298]. Granada [3154, 2517, 1690, 2175, 2090]. Grand [1351, 1361, 775, 1703, 68, 862]. grande [3144]. grandes [1412]. Grandi [2270]. grands [1026]. Grange [2667, 2667]. Granshaw [815]. Grant [1008, 225, 2806, 1950]. Grantham [360]. Granville [185]. Graphs [3224]. Grasshoff [1360]. Grassi [600]. Gratheus [1640]. Grattan [900, 234, 1675, 3251]. Grattan-Guinness [900, 234, 1675, 3251]. Graubard [71]. Grave [2438]. Gravesande [458]. Gravitation [360]. Gravity [831]. Gray [178, 3100]. Grayson [2688]. Graziella [919]. Graznya [65]. Great [601, 2833, 1886, 3165, 1666, 3326, 3142, 1174, 1249, 3004, 3337, 218, 2164, 1458, 2414, 3519, 1710]. Greater [2783, 142, 1607]. Greatest [3529, 131]. grec [1713]. Greci [3332]. greco [2330, 2138, 3427]. greco-romain [3427]. grecs [146, 2330]. Greece [2420, 1628]. Greek [1000, 2969, 426, 1363, 2066, 2327, 2417, 1549, 3330, 610, 1075, 314, 725]. Green [3178, 330, 1833, 3325, 2069, 2985, 2239]. Greene [1627, 661]. Greengrass [2343]. Greenhouse [2836]. Greg [808]. Gregg [1591, 1584]. Gregor [2140, 2909, 2111]. Gregorio [1169]. Gregory [3408, 3317, 2283, 2322, 431, 2327, 686, 1497, 3037, 1203, 3421, 753, 5, 1755]. Greipl [2445]. Grell [1725]. Grelon [2132]. Gren [3149]. Grendler [440]. Grenoble [1043]. Greta [277]. Griechen [145]. Grigorevich [17]. Grimm [2037]. Grinstein [1878]. Grison [2275, 3147]. Grmek [505, 1141, 1023]. Grob [1148]. Groeben [1225, 1747]. Groenevelt [2088]. Groningen [1733]. Grosholz [1192]. Gross [808, 1662, 3484, 3324]. Grosschemie [1588]. Grosseteste [1277, 2809, 310, 2137]. Grossforschung [2923, 875]. Grossforschungseinrichtung [2462]. Grossgerate [2923]. Grossing [1474]. Ground [2284, 197, 2521]. Group [3204, 2794, 3380, 1338, 1051]. Groups [2794, 1031]. Grove [2239]. Groves [225]. Growing [3124]. Growth [2959, 1328, 2999, 1057, 872]. Grubb [3456]. Grubrich [288]. Grubrich-Simitis [288]. Gruder [1766, 1766]. Gruliow [2207, 2207]. Grunbaum [1240]. Gründer [89]. Grundlach [51, 51]. Grundlagen [360, 859, 1240]. Grundlagenforschung [2462]. Grundlagenprobleme [786]. Grundlegung [1927]. Grundlegungsproblematik [842]. Gründungstages [1120]. Guan [1186]. Guard [2227]. Guarnieri [1498]. Gudrun [2411]. Gueny [1466]. Guericke [1924, 1924]. guérit [587]. Guglielmini [2271]. Guicciardini [1101, 932, 1101]. Guidance [595, 1055]. Guide [3020, 300, 2377, 85, 2804, 1378, 393, 1068, 86, 225, 2159, 3518, 3299, 2622, 1427, 2555, 3436, 888]. Guided [2768, 2594, 2498, 2111, 2821]. Guides [1618]. Guido [521, 2155, 2270]. Guigonis [3248]. Guillaume [1026].

Guillaume-Antoine [1026]. Guillaumin [2580]. Guillerme [245]. Guinea [3212]. Guinness [900, 234, 1675, 3251]. Guldentops [3244]. Gundler [1856]. Gundolf [518]. Gunn [1087]. Gunnar [514, 2007, 2390, 1509]. Gunpowder [3041]. Guntau [2103, 1701]. Gunter [3049, 518, 1120]. Gunther [2336, 673, 1120]. Gurney [3270]. Gurtler [429]. Gustav [2984, 2107]. Gusterson [2855]. Gutas [1906, 1550]. guter [3441]. Guthke [632]. Gutting [694]. Guttorm [1280]. Guy [1008, 3244, 1633, 2420, 3248, 3132, 256]. Guyton [2275]. Gynaecology [765, 1248, 2144].

H [627, 2764, 1008, 1395, 1135, 1623, 3252, 571, 1943, 516, 447, 172, 2532, 1774, 2526, 1934, 1878, 1913, 1469, 775, 3270, 194, 474, 1688, 3429, 975, 1411, 2282, 3511, 3151, 1109, 1044, 1853, 1112, 994, 3559, 2959, 1096, 2448, 1297, 3266, 1134, 2337, 996, 1087, 3190, 1506, 3461, 878, 2210, 1535, 1504, 2034, 468, 3290, 2988, 686, 91, 3147, 2101, 280, 3155, 1379, 303, 3513, 2830, 177, 2257, 2466, 756, 205, 2239, 191, 644, 1607, 1458, 3258, 3446, 473, 1478, 1573, 682, 1650, 2022, 597, 1198, 1227, 2029, 2168, 1406, 2292, 658, 590, 2003, 1086, 2116]. H [677, 2814, 2850, 1269]. H.-B [1198]. H.-P [473]. Haar [3461]. Haas [3574]. Haber [3080, 2122, 3568, 15, 1263]. Habermas [489, 489]. Habfast [875]. Habitat [2548]. Habits [1771, 1619]. Hables [3100]. Habsburg [1527]. Hack [813]. Hacker [2213]. Hackett [717]. Hacking [971]. Hadamard [3369]. Hadden [2163]. Haeckel [476, 758]. Hafemeister [225]. Haffenden [1817]. Haffter [520]. Hagelin [1248, 699]. Hagen [1405]. Hagenmeyer [2251]. Hager [2772]. Hahn [273, 1855, 2923]. Hahn-Meitner-Instituts [2923]. Hahns [2665]. Haiken [3476]. Hair [996]. hajj [1718]. Hake [3074]. Hakfoort [2174, 165]. Hakhoe [507]. Hakhoe-ji [507]. hakim [33]. Hal [1502, 966]. Hale [459, 2345, 1763, 2193]. Hales [3170]. Half [3025, 3364, 2687, 3579, 2714, 2925]. Half-Century [3364, 2714]. Halfte [1937]. Hall [140, 1679, 1731, 111, 1281, 1193, 926, 1929, 2461]. Halle [1123]. Halle-Wittenberg [1123]. Haller [1200, 1199, 2114]. Halleux [827, 146, 503]. Halley [839]. hallucinati [254]. Hallyn [1105, 1653]. Hamann [1120]. Hamburg [1941]. Hamburgenses [648]. Hamburger [1501]. Hamburgischen [1941]. Hamesse [1093]. Hamilton [1858]. Hamlet [2136]. Hamlin [1205]. Hammack [142]. Hammond [1578, 3108]. Hammonds [3037]. Han [2408, 1990]. Hand [913, 1996]. Handbook [26, 1175]. Hande [3567]. Handle [1220]. Handlist [3490, 213]. Hands [2663, 2712]. Handschriftensammlung [2253]. Handwriting [3237]. Hane [621, 225]. Hanford [1502, 2540]. Han’guk [507]. Hankins [2644]. Hankinson [883, 1271]. Hanle [682]. Hannah [1228]. Hannaway [110]. Hannes [525]. Hans [517, 3543, 530, 3010, 3118, 471, 489, 1502, 1932, 1946, 324, 163, 453, 1625, 2784, 2705, 2185, 1987, 1398, 357, 1599, 2413, 1303, 673, 2326, 2191]. Hans-Georg [1932, 2413]. Hans-Heinrich [324]. Hans-Joachim [471, 163, 453, 2185, 1599]. Hans-Jorg [3010]. Hans-Liudger [1946].

Hans-Werner [1398]. Hansen [197, 2387]. Hapke [1316]. Haq [2071]. Harald [373]. Haraway [478, 3084, 955]. Harbor [1226]. Harbsmeier [3319]. Harcourt [898]. Hard [1414, 2105]. Harden [963]. Hardetert [2103]. Hardin [1247, 162]. Harding [1900, 1062]. Hardy [1239, 1933]. Harford [3009]. Hargreaves [1109]. Harkányi [1067]. Harlan [3510]. Harley [3267, 577, 1477, 1452, 2145]. Harman [2266, 1487]. Harmke [3464]. Harmond [3108]. Harmonic [2082]. Harmony [2397, 3055, 2472, 1730, 57]. Harold [2088, 3382, 1082, 441, 393, 204, 384]. Harre [206]. Harriet [1604, 2995, 1617]. Harrington [2922, 3328, 3359]. Harris [2032, 167, 2494, 856, 2210, 387, 3171, 1602]. Harrison [3316, 2279]. Harry [1323, 2926, 227, 2536, 2044, 3011, 2323, 385, 1323]. Harth [1384]. Hartlib [2343, 2739, 2739]. Hartman [2678]. Hartmann [1191]. Hartmut [1841, 2413, 2256, 1278, 1531]. Harvard [3252, 3295, 1264, 2878, 3213, 2792]. Harve [3411]. Harvey [138, 571, 2449, 874, 1439, 2507, 3176, 3263, 375, 670, 1999, 688, 2986]. Harwood [1021, 2534]. Harzbergbau [1531]. Hasian [2611]. Haskell [1688]. Haslam [572]. Haslmayr [2342]. Hassan [510]. Hassenfratz [3147]. Hatfield [3064, 937]. Hauke [3367]. Hauptinstrumentes [3253]. Hauser [3535]. Hausmann [1941]. Hauswirtschaftliche [175]. haut [210]. Hauy [2358, 536]. Havana [1695]. Havinga [1134]. Hawaii [3425]. Hawke [274]. Hawkins [3264]. Hawlitschek [3338]. Hawthorne [1407]. hay [1991, 2284, 347]. Hayden [2112]. Hayles [810]. Haynal [2032]. Haynes [2729, 2409]. Hays [3420, 3101, 2783]. Haytham [728, 1809]. Hayum [444]. Hayyân [823]. Hayyan [2071]. Hazard [2121]. Hazardous [3472]. Hazards [3097]. Head [1858]. Headrick [22]. Heal [2982]. Healer [1759]. Healey [690, 1291]. Healing [3151, 1011, 1364, 1693, 1989, 1281, 3291, 3417, 1012, 2113]. Health [2493, 3097, 993, 290, 1934, 947, 1247, 139, 1860, 628, 1027, 1148, 594, 593, 3578, 1499, 470, 1078, 1971, 289, 760, 1732, 1575, 1175, 1404, 1458, 2113, 3474, 2995, 2029, 3576, 2279, 384, 716, 962, 3151, 414, 1281, 762, 676]. Hearn [734]. Heart [3183]. Heat [539, 2418, 2466, 2404, 3384]. Heath [167]. Heaven [1894, 1990]. Heavens [2082, 3244, 2976, 892]. Heaviside [2371]. Heberer [3091]. Hebrew [2808]. Hecht [1531]. Hedin [2036]. Hedley [992]. Heel [1761]. Heerbrant [1649]. Heerding [386]. Heesch [673]. Hegel [656, 2238]. Hegemonikon [428]. Hegemony [2379]. Heiberg [1176, 2242]. Heidegger [870]. Heidelberger [2107]. Heidtmann [954]. Heijmans [2292]. Heike [1214]. Heilbron [841, 1131, 115, 2273]. Heilke [1325]. Heims [1051]. Heinekamp [160]. Heinemann [1766, 1766]. Heinemann-Gruder [1766, 1766]. Heinrich [3367, 550, 259, 2023, 471, 1762, 324, 1375, 2742, 673]. Heinz [345, 2523, 1930, 3327]. Heinz-Peter [2523]. Heirs [1053, 1247]. Heisenberg [2834, 2291, 473, 225, 1617, 3563, 2833]. Heiskanen [1245]. Heitman [1709]. Helaine [1776, 3394]. Held [2739, 3183]. Helden [831, 2814, 2202]. Helen [3317, 1371, 1715, 632, 1552, 1095, 2749, 3037, 685, 2955, 3500, 1076]. Helena [2723, 3143, 1672]. Helene [1484, 1540, 2330, 1172, 3452, 529]. Helfand [1458]. Helffenstein [2453]. Helga [791, 3566, 3375]. Helge [770, 3544, 2831]. Helio [525]. Hella [654]. Hellenistic [2731]. Hellenistische [3049]. Hellenocentrism [1076]. Hellenophilia [1074]. Heller [2769, 2652]. Hellman [2258]. Hellmut [3049]. Helmholtz [2187, 1945, 2365, 938, 937]. Helmholtz-Hering [2187]. Helmut [1322, 2984, 3166, 2965, 2902, 1640, 842, 2033, 586]. Helmuth [1474, 989]. Hematology [3473]. Henderson [1462, 3337, 2907, 1015]. Hendrik [3426]. Hendry [1136]. Henk [2166, 3033]. Henning [1215]. Henri [2237, 535, 1368, 3457, 522, 1756, 2539]. Henri-Jean [522]. Henrich [2447]. Henricus [3244]. Henriette [2457]. Henrika [1490, 2881]. Henriksen [2125, 3477]. Henry [2494, 2797, 3073, 74, 2525, 968, 935, 1908, 2652, 468, 1127, 3377, 560, 3265, 1235, 1890, 225, 1590, 1573, 3351, 2129, 3012, 2388, 1041, 2991, 596, 3116, 2551, 926, 1108, 2186]. Hentschel [1314, 2117, 2606, 1313, 2606]. hepatis [2346]. Heppenheimer [3381, 3587]. Her [573, 1562, 862, 1301, 2793]. Herb [3379]. Herbal [1607]. Herball [2655]. Herbarium [750]. Herbert [3377, 1941, 859, 755, 60, 672, 356, 1587, 1781]. Herbs [3232]. Herbst [3253]. Here [2833, 2693, 2605]. Hereditarian [774]. Heredity [3580, 2694, 3308]. Hereford [2137]. Heresy [3429]. Heribert [442]. heridas [2600]. Hering [2187]. Heritage [1464, 1216, 3510, 3425]. Herken [1584]. Herlihy [2891]. Herman [2299]. Hermann [1751, 851, 654, 247, 2123, 1945, 2365, 2528, 938, 2280, 394, 2270, 1046, 1945]. hermaphrodisme [3427]. Hermeneutical [1669]. Hermes [3285]. hermetico [1811]. Hermione [1125]. Hernandez [2982, 1561, 2158]. Hernschier [3441]. Herodote [1351]. Heron [339]. Herophilus [550]. Herr [2457]. Herries [2753]. Herrlitz [2413]. Herrnstein [3102]. Herrschaft [1946]. Herrschaftsreprasentation [3341]. Hersant [2338, 2895]. Herschel [1491, 2235, 2411, 2480, 3541]. Hersh [3321]. Hershberg [2878]. Hertha [3278]. Hertwig [2024]. Herty [2373]. Hertz [2023]. Hervey [3358]. Herzens [2507]. Herzog [249]. Herzogtum [3150]. Hesiod [3331, 3331]. Hesperies [2743, 834]. Hess [2045, 2410, 1353, 1769]. Hesse [266]. Hessen [1190]. Hetherington [2768, 24, 1621, 3514]. Heuristic [391]. Heuristics [1605]. Heurne [1481]. Heuser [266, 2280, 1572]. Heuss [1225]. Hevly [1328]. Hewitt [2597]. Hewlett [582, 3171]. Hexaemeron [1277]. Hexenkrauter [2081]. Hexenprozess [1474]. Heyd [2662]. Heyden [3061]. Heyman [3329]. Hickel [2620]. Hickman [780]. Hidden [3163, 1534]. Hideo [2536]. hier [2693]. Hierogliphicks [2810]. Hieroglyphicall [2897]. Higby [2322, 1497, 3421, 2283]. High [3566, 704, 192, 965]. High-Temperature [3566]. Higher [3160, 2315]. Highest [3310, 1584]. Highland [2648]. Highlands [941]. Hilary [555, 2047]. Hilbert [1403]. Hilbertprogramm [1500]. Hilde [1812]. Hildebrandt [89, 1322]. Hildegard [2889, 434]. Hildreth [470]. Hilfstein [1373]. Hill [1047, 1134, 483]. Hilla [3179]. Hills [667, 1903]. Hilton [77]. Hilts [1293].

Him [3393]. Himmels [163]. Hinsley [2759]. Hipolito [2272]. hipótesis [340]. Hippocrate [2970]. Hippocrates [1247, 2970, 3240, 1180]. Hippocratic [26, 988]. Hippokrates [2067, 513]. Hippokratischen [549]. Hirnforschung [3375]. Hiroshi [2537]. Hiroshima [2878, 2925, 60, 2384, 225]. Hirsh [681]. Hirt [2214]. Hispanica [523, 2491, 132]. Histoire [587, 516, 1790, 3088, 1412, 1259, 3446, 687, 3415, 150, 2320, 3067, 2567, 1390, 729, 236, 2237, 3020, 1580, 1897, 3144, 3122, 271, 1026]. Histoires [1307]. Historia [525, 1923, 1690, 526, 524, 2158, 2063, 523, 2434, 2027, 1561, 2803, 1261, 507, 321, 134, 2490, 2249, 113]. Historiae [513, 504]. Historiam [523]. Historian [3364, 408]. Historians [3141, 1260]. Historic [1247, 2178, 2520, 3540]. Historical [627, 3104, 486, 559, 1055, 367, 2633, 3573, 2801, 2062, 1457, 866, 1080, 1382, 992, 855, 2273, 1184, 2036, 905, 1978, 3234, 2964, 1980, 1961, 1087, 1621, 3072, 304, 3035, 126, 63, 1624, 808, 303, 1059, 2358, 3513, 3229, 3329, 2379, 1175, 537, 1264, 888, 510, 1977, 491, 1420, 2831, 2231, 586, 1802, 2194, 1703, 3094, 115, 3378, 793, 211, 3036, 506, 679]. Historicism [294]. historie [1042]. historien [1496]. Histories [2886, 2952, 499, 2832]. Historii [512]. Historiographical [2568]. Historiography [1553, 3169, 100, 1138]. Historiques [825, 749]. historischer [1531]. History [1626, 1706, 3346, 1490, 3104, 2684, 2284, 2453, 2764, 411, 298, 1891, 1413, 2678, 902, 818, 2494, 2399, 3419, 2989, 3204, 769, 580, 87, 2839, 1135, 960, 2408, 338, 2725, 2879, 241, 1623, 1795, 1317, 1806, 3299, 2788, 1262, 716, 233, 1434, 2752, 1534, 874, 1597, 1707, 778, 1851, 3479, 2574, 2930, 913, 107, 516, 1457, 3424, 2233, 121, 2115, 2209, 809, 910, 616, 909, 1000, 791, 2529, 3073, 1579, 3007, 11, 3235, 1407, 1790, 2285, 2874, 300, 3206, 2996, 110, 2442, 118, 3370, 2230, 1596, 3381, 2641, 2322, 3232, 106, 1764, 112, 2654, 509, 1997]. History [111, 3010, 3406, 1218, 991, 2798, 2030, 413, 2779, 2125, 220, 1453, 3210, 3413, 235, 1454, 1860, 3318, 1899, 1744, 3211, 3212, 2091, 3224, 3587, 3124, 3383, 2965, 693, 3213, 258, 3517, 2826, 3372, 3214, 2017, 583, 3136, 976, 56, 3215, 2941, 3394, 1223, 386, 1898, 3006, 2063, 2934, 1087, 1224, 507, 2049, 346, 1843, 103, 1935, 3313, 179, 2804, 3033, 1849, 237, 1035, 109, 604, 1089, 1050, 1459, 2144, 1700, 2034, 1452, 2145, 1075, 2886, 2538, 2038, 123, 2873, 1869, 2266, 589, 1226, 3506, 3324, 3386, 2234, 1349, 3032, 598, 3554, 360, 232, 2670, 1432, 2064]. History [1282, 3515, 3041, 2729, 2572, 1, 606, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 620, 2390, 1746, 3513, 1812, 2540, 3387, 141, 508, 1054, 1360, 3476, 722, 1074, 3261, 3040, 1428, 2025, 335, 667, 2245, 2323, 1741, 424, 1370, 1675, 1770, 1327, 963, 1458, 1533, 857, 1865, 3034, 32, 510, 3120, 1901, 2727, 415, 1131, 3046, 274, 225, 3017, 2688, 541, 1590, 1350, 3544, 1903, 2129, 2262, 2488, 2740, 13, 1315, 506, 125, 3323, 3012, 631, 368, 2643, 116, 679, 3490, 3296, 1229, 1508, 1544, 2638, 2685, 1272, 1505, 2127, 3590, 1047, 515, 2802, 114, 630]. History [2003, 1426, 1877, 2544, 2575, 2693, 104, 3221, 381, 1942, 1850, 3465, 3290, 222, 502, 1079, 3199, 3303, 1614, 221, 3205, 2311, 1702, 121, 1972, 110, 111, 3207, 3209, 896, 2627, 1258, 2947, 500, 3219, 3220, 2548, 2681, 1017, 3042, 420, 2321, 911, 3409, 20, 333, 231, 1838, 2851, 2689, 195, 3420, 1174, 2146, 817, 2671, 337, 72, 2215, 3233, 117, 136, 1619, 1803, 1902, 21, 815, 3237, 422, 3095, 850, 225, 369, 119, 1086, 381, 105, 378, 3107, 2880, 1908, 1186, 813, 2555, 832, 120, 505, 2078, 1141, 1907, 234, 1121, 3384, 661, 1046, 649]. Hitler [1211, 1410, 2922, 677, 2461, 3575]. Hittell [3073]. Hiyal [1005]. Hobart [2670]. Hobbes [451, 925, 356, 3497, 2977, 832, 925]. hoch [1637]. hoch- [1637]. Hochgeehrter [2457]. Hochschulalltag [1501]. Hochschuldisziplin [2359]. Hochschule [325, 1856, 2059]. Hochschulwesens [1856]. Hoddeson [3562, 2125, 1865]. Hodgson [1236]. Hodierna [2744]. Hoenen [2253]. Hoeniger [1565]. Hoffer [288, 2032, 2604, 288]. Hoffert [53]. Hoffman [283]. Hoffmann [1481]. Hofmann [1841, 929]. Hogendijk [3242]. Hohenheim [643]. Hohenlohe [1724]. Hoke [1038]. Hold [1774]. Holden [1607]. Hole [323]. Holger [2510, 1658, 1501]. Holism [2922]. Holistic [2471]. Holl [3171, 582]. Hollinger [79]. Hollingworth [1323, 1323]. Holloway [2206, 1174]. Hollywood [1432, 2792]. Holmes [2373, 2913, 3255, 1678, 565, 1977, 2191]. Holocaust [3459]. Holówka [977]. Holton [1854]. Holtzmann [2848]. Holy [2168]. Homann [1181]. Hombach [2018]. hombros [352]. Home [2521, 1502, 1133, 506, 3072, 1085]. Homer [1703]. Homewood [865]. Homicidal [242]. hommage [3415]. homme [1982]. Hommes [2324]. Homosexual [3026]. Homosexualities [3125]. Homosexuality [2701]. Honner [78]. Honor [3310, 3511, 1754, 1263, 32, 428]. Honour [576, 2266]. Hook [1688, 2753]. Hooke [647, 2815, 646]. Hoolihan [2885, 795]. Hooykaas [2946, 2358]. Hope [87, 787, 1778, 2882, 2601]. Hopeful [1784]. Horace [1200, 2178, 2838]. Horace-Benedict [1200, 2178]. Horacio [262]. Horiuchi [2436]. Hormigón [524]. Horn [595]. Hornell [3333]. Horney [2928]. Horoscopes [337]. Horowitz [1513]. Horse [2783]. Horsfield [1392]. Horsman [272]. Horst [530, 654, 259, 861, 51]. Horstmann [3546]. Hortega [1311]. Horticultural [1607]. Hörz [755]. Hosam [2070]. Hoses [3061]. Hoskin [3323, 114]. Hoskovec [3047]. Hospital [3295, 2666, 412, 2296, 815, 2545, 868, 3295, 3473]. Hospitals [2674, 2828, 751, 592, 2981]. Hossam [1649]. Hossein [1805, 1717]. Hossfeld [3091]. Hot [1220]. Hour [1326]. Hourya [1042]. Household [2631]. Households [3113]. Houses [1226]. Houzel [1484]. Howard [2989, 769, 1807, 3481, 1771, 817, 3456, 2061, 2840, 711, 2117, 225, 3077]. Howell [1848, 2545]. Howsam [3018]. Howse [562]. Howson [790]. Hoy [1378]. Hoyningen [1603, 786]. Hoyningen-Huene [1603, 786]. Hristo [81]. HSS [3204]. Hsu [1682]. Huainanzi [1990]. Hubbard [625]. Hubble [2120, 2530, 2235, 2411, 2768]. Huber [1501, 1721]. Hubner [1474]. Huby [1906, 428]. Hud [2877]. Huda [3525]. Hudson [1350]. Huene [1603, 786]. Huertas [183]. Huesca [3053]. Huet [1540]. Hufbauer [1171]. Huff [1798]. Huffman [644]. Hugeniorum [2347]. Huggett [850]. Hugh [1986, 2937, 2020, 2855, 1181]. Hughes [244, 629, 487, 742, 641]. Hugo [435, 3431, 3308]. Hugues [435]. Huguet [2825]. Huib [2515]. Huisman [1733]. Huizenga [1959]. Hulin [657]. Hulin-Jung [657]. Hull [709, 709, 1460, 527]. Hullen [1381]. humain [2658]. Human [1468, 3419, 2549, 121, 1800, 918, 881, 542, 3413, 3420, 3228, 3557, 2737, 907, 237, 952, 3438, 1144, 572, 3371, 2694, 2328, 318, 238, 2784, 2612, 1940, 2243, 483, 2721, 1207, 560]. Humanism [441, 1919]. humanisme [2335]. Humanist [248, 1065, 777]. Humanistic [1149]. Humanists [702, 2002]. humanitären [758]. Humanity [2416, 661]. Humans [1984]. Humboldt [1830, 365, 2185, 1829, 2103, 365]. Humboldts [2185, 845]. Humboldtschen [1303]. Hummon [3176]. Humphreys [1934, 785]. Humphry [1569]. Hund [379]. Hundert [1316]. Hundred [2678, 906, 68, 1702, 2029]. Hunger [247, 2282, 3511]. Hunley [1863]. Hunt [1397, 2974, 2074]. Hunter [647, 2749, 178, 1089, 1378, 3428, 2898, 3526, 2171, 557]. Hunterian [1394]. Hunters [715]. Hurt [2049, 2049]. Husserl [1150, 1064, 688]. Hutcheon [2726, 2726]. Huttig [1274]. Hutton [1108, 813, 1741, 3428]. Huxley [2692, 1396, 468, 1127, 2202, 3025, 2636]. Huygens [257, 2347, 2814, 2900, 2814]. Hwarizmi [2973]. Hybrid [675]. Hybrids [2111]. hydrogenation [1671]. Hygiene [2914, 1414, 595]. Hygienikers [2110]. Hypatia [2419]. hyperbolae [2089]. Hypnosis [1744]. Hypnotism [603, 1744]. hypothese [2567]. Hypotheses [355]. Hysteria [1324, 2800, 2996, 1344].

I-II [3132]. I-III [728]. I-IV [2075]. Iain [2597]. Ian [971, 818, 913, 1208, 3098, 1630, 1901, 3487, 953]. Iannaccone [1365]. Iannone [296]. Iatromechanische [1481]. Ibn [1718, 339, 728, 2877, 1631, 1717, 2070, 1809, 2244, 433, 823, 2071, 3525, 3432, 219, 1276, 2808]. Icarus [1330]. Ice [3164]. Icelandic [1822]. Ich [3441, 2447]. Ichiro [507]. Ida [2284, 658]. Idea [1325, 3466, 1025, 908, 3177, 47, 2103, 208, 1272, 1107]. Ideal [281, 164]. Idealism [726]. Ideals [611, 191]. Ideas [2148, 1708, 2833, 2798, 747, 457, 2938, 3047, 3526, 1549, 752, 1476, 737, 1049, 2064, 1020, 857, 1681, 167, 1349]. idee [1915, 1195]. Ideengeschichte [374]. idées [1982]. Identity [2493, 3206, 3495, 3289]. Ideological [1489]. Ideologies [419]. Ideology [1147, 1353, 693]. Idiocy [2801]. Idioms [2950]. Ido [2371]. Ier [1649, 916]. Ieradiakonou [1001]. If [1866, 2202]. Ignacio [1389]. Ignaz [2108, 1120]. Ignorants [1014]. Igor [2120]. Ihde [1061, 1047]. Ihre [2300]. ihren [266, 2280]. ihrer [325, 1115, 2679, 471]. Ihsanoglu [1005]. II [3276, 378, 2922, 1561, 768, 2581, 2540, 766, 1724, 1361, 2056, 95, 1738, 1585, 1271, 3132, 1416]. III [1176, 549, 728, 2067, 423, 2276, 3296, 2011, 1361]. Ike [3183]. Ikhwan [1717]. Ilana [3291]. Ilario [2063]. Ilich [2835]. Ilie [2439]. Ill [3236, 2524]. Illinois [675]. Illness [1413, 3182, 2493, 1838, 2470, 2339, 2524]. Ills [3579]. Illuminismo [2750]. Illusion [2227, 3280, 538]. Illusions [2548, 2786, 2472, 1773]. Illustrandam [523]. Illustrated [3042, 2725, 1087, 3506, 383, 1248, 2688, 697, 3323, 3518]. Illustrating [1702]. Illustration [1539, 1665]. Illustrations [1750, 2520, 3540, 572]. Illy [2766]. ’ilm [1991]. Ilse [654, 1130, 288, 1669, 3257]. ilusión [261]. Ilustraciones [2803]. Ilustrada [261, 3445, 1389]. Ilyas [533]. Im [2016, 2914, 2018, 286, 325, 1115, 1481, 1660, 851, 3367, 160, 1635, 940, 663, 2382, 2081, 2252, 3003, 1274, 1993, 2110, 266, 3241, 2411, 3278, 2523, 12, 2747, 3150, 2067, 3256, 1215, 2300, 989, 2842, 1120, 2254, 1304, 2583, 1640, 2154, 1813, 2742, 3341, 1941, 461, 1651, 1658, 3060, 1572, 1381, 1501, 2923, 3091, 1142, 1531, 1266]. Image [1081, 2739, 3038, 3185]. imagen [1354]. imagenes [1911]. Imagery [1433, 85]. Images [2399, 623, 2493, 2591, 1645, 85, 3123, 877, 2859, 1458, 225, 3418]. Imagination [2591, 2773, 3522, 2759, 3357, 2758, 1982, 1806, 2104, 1540, 3481, 666, 3164, 2644]. Imaging [2848, 2849]. Imagining [632, 2524, 2904, 1663]. immaginazione [2260]. immagine [443]. Immagini [1116]. Immanuel [3064, 2818]. Immigrant [1949]. Immigrants [3077]. Immortal [2372]. Immunological [1433]. immunologie [1412]. Immunology [63, 1033, 2764]. Impact [3292, 3471, 3300]. Impacts [2542]. Imperative [751, 3578]. Imperfect [2061]. Imperial [3582, 2024, 2012, 944, 1229, 143, 2971, 332]. Imperialism [3495, 1166, 2239, 22, 1167, 3124, 847, 1355]. Impey [2521]. Impolite [2351]. Imponderables [2273]. impone [2000]. Important [699]. impossibile [2526]. Impossible [878, 3143]. Impressions [3572]. Improvement [1482]. Impulse [3267]. Impure [2796, 1511]. Impurity [1205]. Inca [1468]. Incarceration [1037]. Incarnate [3229]. Ince [377]. incertain [3427]. Including [3502, 1656, 1673, 831]. Incorporated [1493]. Incubator [2775]. independant [1671]. Independence [408, 2226]. Indestege [2257]. Index [1795, 3340, 51, 695, 51, 491, 1928, 916]. India [1364, 28, 1309, 1907, 3258, 2279]. Indian [961, 1543, 1309, 508, 1750, 2279, 3223, 508]. Indiana [3210]. Indians [573]. Indias [2158]. Indici [643]. Indien [1631]. Indies [2442]. Indigenous [2813, 332]. Indikatoren [1993]. indische [2973]. Indiscrete [3000]. Individualism [3126]. Individuals [395]. Indivisibles [2896]. Indonesia [130]. industria [2825]. Industrial [580, 396, 3097, 3349, 1954, 96, 628, 2616, 44, 2682, 1089, 45, 1494, 303, 1732, 1265, 1901, 3474, 2480, 2686]. Industrialising [1765]. Industrialization [371, 1493]. Industrie [2633, 2757, 1841, 1535, 2292, 3360]. Industries [249, 2972, 2686]. Industriezeitalters [2281]. Industry [627, 580, 2386, 213, 2285, 2026, 64, 1752, 3483, 3161, 23, 1317, 681, 280, 1136, 3287]. inedit [3460]. inedite [919]. Inediti [2744]. inedits [2010, 3340, 1026, 2508, 161]. Inevitable [2786]. Infant [947, 67, 53]. Infatuation [1151]. Infectious [1933]. Infektionskrankheiten [3327]. Inference [390, 737]. Infertility [2799]. infini [2895, 2004]. Infinite [3167]. Infinitely [2896, 407]. Infinitesimal [159]. Infinitesimals [2896]. Infinity [1213, 2632]. Influence [905, 1112, 854, 2243, 1861, 1511, 1169, 1906]. Influences [1, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 2362, 107, 106]. Influential [2958]. Information [140, 400, 798, 1157, 1522, 1881, 2219, 2558, 2863, 3194, 3197, 3593, 2874, 1299, 2616, 85, 1778, 1599, 1952, 590, 91]. Ingeborg [2032, 2604, 1069]. Ingeniería [921]. ingenieros [262]. Ingenieur [2923, 3088, 1753]. Ingenieurausbildung [325]. Ingenieure [1946]. Ingenieurwissenschaften [325]. Ingenieurwissenschaftler [2033]. ingenios [154]. Ingenious [1038]. Ingenjor [1442]. Ingo [1481]. Ingolstadt [2254]. Ingrao [913]. Inhaltsangaben [2987]. Inherit [1443]. Inheriting [1208]. Inhibition [1579]. inizio [3545]. Inkster [1901]. Inleiding [240]. Inmates [3113]. innere [2107]. innern [2992]. Innig [2457]. Innocenzo [3332]. Innovacion [2825]. Innovating [1136]. Innovation [140, 912, 1704, 434, 23, 3561, 763, 1306, 1399]. Innovator [2005]. inquieto [1194]. Inquiries [1703, 1770]. Inquiry [3472, 1922, 1058, 2961, 1872, 685, 3435, 2568]. Inquisicion [1470]. inquisiti [2259]. Inquisition [3437]. ins [3327]. Insane [335]. Insanity [3236, 2360, 242, 2673]. Inscribing [3589]. insegnamento [124]. Insight [1617]. Insights [317]. insospechada [2516]. Inspirations [3165]. Instauration [2164]. instinct [2680]. institucional [262]. Institut [1941, 1390]. Institute [396, 3367, 650, 300, 1218, 2377, 508, 3490, 466, 3367]. Instituten [3367]. Institution [598, 2019, 1786, 2931, 2171, 2580]. Institutional [2806, 878, 1797]. Institutionalisierung [266]. institutionellen [1313]. Institutions [325, 1286, 2699]. Institutions- [325]. institutos [2027]. Instituts [1215, 2923]. Instrument [713, 843, 1936, 697, 2515, 127]. instrument-maker [2515]. Instrument-Making [843]. Instrumental [1061]. Instrumente [1292]. Instrumentgerichte [127]. instrumentos [1374, 2645]. Instruments [2285, 326, 3422, 697, 1802, 34, 31, 2644]. Insurance [2693]. integrale [2149]. Integrating [289]. Intellect [1092, 13, 3244]. Intellectual [2343, 2527, 848, 2691, 2806, 1045, 2269, 3087, 1187, 3588, 1632, 1831, 3102, 1285, 2430, 13, 1939]. Intellectuals [177]. intellectuele [2335]. intellectuelle [322]. Intelligence [1256, 1851, 1345, 3377, 71, 3188, 596, 3038, 3371]. Intelligent [975]. Intelligenz [2535]. Intensive [2775]. Interaction [2083, 3496, 3485]. Interactions [2319]. Interactive [690]. Interaktion [1115]. Interdisciplinarity [809]. Interdisciplinary [3328, 1948]. interdisziplinarer [1500]. Interdisziplinares [1278]. interes [1561, 134]. intereses [2825]. Interest [2641, 2006, 3584, 2683]. Interests [1262, 1970]. Interface [1061]. Interference [1121]. interioribus [1716]. Interleukin [3291]. Interleukin-2 [3291]. Interlocutor [2055]. intermediaire [2166]. Internal [68]. International [3104, 1317, 2788, 106, 111, 1087, 1742, 507, 2049, 103, 3041, 2995, 515, 2686, 1338]. Internationale [516, 3567, 518, 3552]. internationalen [1120]. internationaler [1842]. Internationales [503]. Internationalis [504]. Internationalism [1676]. Internazionale [521]. Internet [3515]. internis [3531]. Interpretation [2764, 487, 2463, 1206, 690, 2105]. Interpretationen [1314]. Interpretations [1790, 3354, 904, 2800]. Interpreted [1005]. Interpreting [3249, 2429, 2094, 1905]. Interrogator [2055]. Interwar [2868]. Intimate [3113]. intorno [2980]. intra [358]. Intrinsic [3473]. introducción [868]. Introduction [486, 2549, 613, 1176, 735, 167, 1929, 1961, 1459, 1423, 349, 750, 1788, 2426, 1717, 1073, 3217, 3133, 966, 3335, 1367, 452, 3432, 3002, 233, 3422, 3486, 1533, 1064]. Introductions [1386]. Introductory [206, 379, 79]. introdujeron [1911]. Introduzione [443]. Introns [1144]. Intuitie [2101]. inv [2733]. invaliditeit [2693]. inventados [1374]. Inventarium [3248]. Invented [671]. Inventer [1237]. Inventing [987, 3015, 808, 2472, 1055, 2115, 44, 292, 2883, 1260, 47, 3569]. Invention [629, 2233, 1491, 1302, 912, 3186, 1494, 2735, 2263, 538, 869, 2797, 1681]. Inventor [3270, 372, 3285]. Inventory [1795, 1777]. invenzione [973]. Investigation [3473, 1242, 2266, 2722]. Investigations [2969]. Investigative [565]. Invisible [1662, 913, 2263]. Involuntary [1327]. Ioffe [861]. Ionic [2836]. Iorio [1182]. Iovialis [740]. Iowa [1858, 1247]. Ippolito [566]. Ireland [713, 906, 2753, 2736]. Irene [1989]. Irish [253]. Iron [2576, 2240]. Irreverent [3204]. irrt [1762]. Irvine [3042, 1742]. Irving [121]. Isaac [838, 1731, 1112, 249, 360, 741, 1291, 1731, 3530, 558]. Isabelle [2812, 2879, 2247, 1744, 825, 3186, 1728, 1653]. Isaia [1365]. Isenheim [444]. Ishmael [954]. Isidore [2228]. Isis [3017, 3205, 106, 3017]. Islam [1798, 2245, 1909]. Islamic [1451, 1908, 533, 2073, 1452, 640, 1717, 2072, 31]. Islamicate [346]. islamische [251]. Islamischen [510]. Island [2681, 1461, 2239]. Islands [2681, 50, 1392]. Isles [491]. Ismael [2166]. Isolani [2271]. Isolation [3498]. Israel [913, 1947, 2369, 249, 1494]. Issued [835]. Issues [3410, 968, 225]. Istituto [1194]. Istorii [511]. Istoriia [910]. Italia [2000, 2963, 3001, 3268, 3277, 1722, 1498]. Italian [2269, 2142, 613, 744]. italiana [1779, 771]. italiano [645]. italicum [161]. italienischen [1813]. Italy [1182, 1169, 440, 1342, 2086, 2738, 922]. Iter [161]. Itinerari [2340]. itiro [3373, 2537]. Iu [887]. IUTAM [72]. IV [269, 438, 2075]. Ivan [930]. IVF [385]. Ivo [558]. Ivor [900, 234, 1675, 3251]. Ivory [34, 1385]. Ivy [1688]. IX [2242].

J [627, 202, 2546, 3019, 2148, 1484, 2293, 1603, 3249, 713, 281, 858, 3274, 939, 1751, 3079, 906, 3252, 213, 115, 367, 2075, 359, 2392, 1678, 3271, 321, 2574, 633, 3356, 726, 1461, 516, 15, 1007, 1382, 1716, 378, 3007, 781, 2718, 1362, 3354, 3153, 1913, 536, 1800, 3522, 1513, 1922, 336, 2283, 2322, 2273, 2903, 2984, 2975, 2906, 3406, 674, 843, 1557, 2030, 905, 1491, 3429, 1269, 231, 2282, 1837, 1646, 1745, 841, 2501, 3420, 2460, 382, 867, 2924, 710, 2166, 2805, 3517, 380, 793, 1169, 2235, 994, 1808, 849, 1462, 2771, 569, 3322, 2410]. J [3006, 1134, 548, 1011, 1397, 1217, 1840, 3190, 2369, 3176, 3578, 1935, 2804, 3033, 34, 1539, 2671, 2077, 1279, 813, 1302, 1249, 198, 1050, 337, 2144, 2261, 2492, 1127, 1452, 2145, 1906, 3045, 77, 349, 1928, 3180, 383, 292, 3343, 1353, 3147, 1497, 2497, 598, 782, 2280, 808, 360, 1383, 88, 319, 2809, 2676, 1938, 3041, 3363, 2856, 1489, 3450, 2459, 1271, 2499, 2002, 717, 943, 2746, 2045, 1991, 493, 1133, 2117, 2766, 829, 1372, 1495, 766, 3345, 1153, 1987, 1236, 3084, 3541, 1533, 1682, 2451, 2450, 1275, 1636, 531, 1545, 225, 1826, 1016, 955, 1629, 1861]. J [1863, 2129, 2810, 2714, 125, 3421, 2610, 379, 1601, 2685, 2731, 1047, 3576, 2644, 3353, 3055, 252, 3459, 2253, 3574, 1769, 1791, 2928, 794, 1867, 2925, 2693, 1802, 2992, 753, 839, 929, 672, 1777, 2092, 2515, 844, 1090]. J. [3349]. J.-L [349]. J.-P [1484]. Jâbir [823]. Jaan [1673]. Jaap [1401]. Jabir [2071]. Jabr [641]. Jacalyn [1839, 3266, 1141]. Jacek [977]. Jack [582, 3368, 3294, 1209, 3510, 143, 1503, 3171]. Jackson [3543, 572, 250, 1509]. Jacob [654, 320, 1989, 1917, 2270, 3426, 2008, 458, 3315, 1989, 2268]. Jacobaeus [2510]. Jacobi [2984, 2153]. Jacobs [2995]. Jacquart [148, 3189]. Jacqueline [1093, 1746, 1307]. Jacquelyn [2622]. Jacques [281, 3583, 651, 322, 2981, 3536, 2188, 2165, 522, 3329, 146, 2970, 1064, 2116, 706, 3369, 1054]. Jacyna [2176]. Jagdish [2381, 586]. Jahn [654, 2110]. Jahncke [2884]. Jahnke [1303]. Jahr [822]. Jahre [1316, 160, 654, 3586]. Jahren [1855]. Jahrestages [2185]. Jahrhundert [2018, 266, 3150, 2583, 812, 1660, 851, 2081, 1927, 2252, 3003, 3241, 3278, 3327, 2254, 2415, 1658, 348, 3091]. Jahrhunderts [2509, 1638, 830, 730, 663, 2336, 1185, 1932, 1937, 1304, 1634]. jährigen [471]. Jai [2599]. Jaime [2027]. Jakab [869]. Jaki [226, 18]. Jakob [656, 2008, 1393]. Jakub [979]. James [2368, 3094, 3249, 3512, 197, 1132, 655, 1702, 2429, 2115, 3315, 2094, 1460, 2084, 2748, 3326, 683, 3009, 3348, 1953, 48, 3259, 1402, 1761, 3372, 1169, 2017, 1112, 3101, 2650, 1514, 2783, 2778, 1357, 3190, 1849, 3374, 1577, 2130, 468, 2011, 686, 211, 3371, 3333, 1378, 731, 69, 942, 162, 2857, 1598, 1868, 3362, 2379, 343, 3541, 888, 421, 2907, 415, 527, 1487, 2672, 2387, 2844, 1750, 375, 2910, 597, 2388, 1227, 661, 2878, 1086, 1602, 2065, 912, 1938, 2002, 2179, 382, 464, 1741, 757, 2878]. Jami [931]. Jan [1481, 3242, 1480, 3058, 3061, 1486, 1216, 181, 864, 63, 3458, 3433, 3048, 2194]. Jana [1655]. Jandun [729]. Jane [3408, 3339, 1128, 1673, 3005, 304, 323, 1747, 2276, 1318, 2625, 405, 1575]. Janet [1036, 2156, 1748, 3554, 1458, 815]. Janice [282, 1015]. Jannelle [2685]. Janossy [59]. January [2946, 1793, 1573, 1973]. Janus [1380, 516]. Janusz [65]. Janze [3568]. Jaouiche [149]. Japan [465, 3318, 2577, 421, 1345, 507, 330, 2924, 2536, 2538, 1955]. Japanese [995, 996, 2537, 2210, 2226]. japonaises [2436]. Jarcho [566]. Jardin [651, 2090]. Jardine [754, 1058, 109]. Jarmo [2289]. Jarrell [1460]. Jarrett [3297]. Jasanoff [2389, 1330]. Jauregg [1846]. Java [1392, 368]. Javier [261, 2508, 524, 1760]. Jay [2860, 3343, 2938]. Jayawardene [2942]. Jazzar [2808]. Jean [2743, 2509, 3460, 2580, 1649, 791, 1518, 1369, 451, 3298, 1351, 317, 935, 628, 2895, 837, 830, 900, 3538, 460, 925, 2528, 1825, 3542, 3051, 750, 2911, 3458, 1566, 522, 2117, 2539, 2766, 3034, 2333, 729, 1026, 2496, 3505, 545, 210, 320, 3066, 2093, 256]. Jean-Claude [935, 3458, 522, 3034]. Jean-Francois [3460, 3298, 2539]. Jean-Louis [2539, 545]. Jean-Palerne [210]. Jean-Patrice [2333]. Jean-Paul [1649, 837]. Jean-Philippe [3542].

Jean-Yves [2580]. Jed [2569, 2023, 3260, 2117, 173]. Jefferson [567, 2516, 2748, 621, 225]. Jeffrey [2641, 1054, 2775, 944, 1260]. Jeffreys [3405]. Jellison [2200]. Jelved [3543]. Jenkins [372]. Jenkinson [1746]. Jennifer [2724, 2927]. Jensen [1127]. Jeopardy [2633]. Jeremy [2461, 1688]. Jerome [2368, 2910]. Jerrold [1503]. Jerry [489, 3074, 2540]. Jervis [1039]. Jerzy [977]. Jesseph [1386, 1736]. Jessica [828]. Jessie [139]. Jesuit [2653]. Jesuiten [1651]. Jesuitenmissionaren [2265]. Jesuits [1818, 2869]. Jesup [3348]. Jesus [3053]. jeux [2508]. Jevons [1030]. Jewish [2189, 1997, 3038, 3333, 1639, 156, 3077, 3309, 35]. Jews [3038, 2423]. Jeyaraja [732]. Jh [1993]. ji [507]. Jill [66]. Jim [338, 2651]. Jin [1186]. jing [2732]. Jiri [3047]. Jitse [3039]. Jo [3320, 2181, 1380, 2268]. Joachim [740, 1923, 1721, 3531, 837, 471, 1113, 163, 648, 453, 2161, 2185, 143, 1599]. Joan [2453, 1419, 1219, 573, 268, 1506, 1825, 3290, 262, 2113, 1636, 2203, 679]. JoAnne [1851, 3326, 2536]. Joannes [2088]. Joaquin [2090, 1389]. Job [3097, 1506, 2083]. Jobmann [3586]. Jochen [1463]. Jock [2128]. Jodelet [1324]. Jodogne [2331]. Joe [2498]. Joel [1184, 3234, 1709, 3579, 1405, 2360, 1848, 2545, 54, 844, 522]. Joella [257]. Johan [1173, 3426, 3533]. Johann [1371, 3534, 259, 3441, 1375, 2353, 3338, 162]. Johanna [1978]. Johanne [730]. Johannes [342, 1191, 2445, 1474, 1107, 1113, 3239, 42, 998, 1650, 895, 1728, 3055, 354, 42]. Johannis [2249]. John [264, 418, 769, 2189, 2021, 1371, 705, 455, 2700, 1483, 716, 961, 2521, 1537, 78, 1617, 3267, 37, 2114, 1189, 3116, 726, 1702, 3163, 2905, 207, 992, 2958, 1111, 435, 1680, 3107, 1604, 662, 3131, 3078, 1477, 2156, 2748, 3232, 1125, 1699, 1106, 2175, 2998, 1684, 64, 1817, 2088, 1558, 1021, 2689, 2091, 768, 2068, 285, 3557, 924, 1134, 3288, 284, 1957, 733, 1279, 2598, 1959, 1720, 1906, 1519, 572, 3333, 1497, 2497, 974, 3337, 1104, 301, 2572, 1136, 772, 776, 2657, 722, 1619, 1575, 3135, 1392, 2117, 2990, 481, 2164, 1902, 3521, 3541, 2551, 936, 1780]. John [965, 1131, 3017, 1475, 1350, 1995, 1573, 2770, 2844, 1315, 299, 1229, 2685, 1990, 1064, 661, 1422, 373, 678, 2116, 2575, 1046, 104, 2992, 452, 265, 2480, 3207, 223, 1887, 1066, 447, 147, 2337, 780, 1319, 750, 1114, 3393, 2087, 1039, 2601]. Johnson [1662, 199, 936, 2129, 2039, 944]. Johnston [2691, 2651, 2538, 1607]. Joint [3221]. Jole [2062, 638, 1994]. Jolley [2172, 747]. Joly [1729]. Jon [868, 3355, 3337, 177, 3487]. Jonas [579]. Jonathan [2677, 49, 2104, 1001, 1604, 1988, 2737, 878, 2040, 2534, 3263, 966, 1827, 23, 1299]. Jones [1677, 430, 3043, 3245, 751, 277, 2066, 711, 155, 225, 1065, 1760, 1291, 431]. Jonson [733]. Jordan [2309, 386]. Jordanova [623]. Jordanus [1096]. Jordi [2765]. Jorg [3010]. Jorge [1354]. Jørgen [3276]. Jose [1911, 1331, 2700, 2516, 2491, 2600, 2175, 2272, 2490, 3442, 1561, 2158, 2851, 1470, 1354, 3444, 2090, 1311, 2435, 2656, 321, 1024, 132, 752, 526, 1149, 154, 351, 152, 1040, 746, 1311]. Josef [51, 3047, 1941]. Josep [1095, 1917]. Joseph [2754, 2356, 2277, 3528, 1466, 3508, 167, 329, 2423, 2196, 1452, 393, 275, 2704, 3342, 2422, 2972, 1435, 1749, 2451, 225, 1057, 1823, 1573, 682, 3590, 842, 1031, 1137, 3319, 929, 2312, 1348, 3547, 1740, 2598, 3446, 3351]. Joseph-Diez [3446]. Josephson [3089, 3172, 2288]. Joshua [1051, 509, 3124]. Josiah [272]. Jospe [433]. Jost [1724]. Jouanna [2970]. Joule [464]. jour [1124]. Jourdain [1675]. Journal [513, 411, 107, 507, 508, 2262, 116, 1892, 518, 104, 452, 2099, 110, 111, 510, 510, 1298, 121, 118, 260, 114]. Journal. [2991]. Journals [264, 2540, 2990]. Journey [1699, 3365, 2770, 964]. jours [212]. Jovis [1728]. Joy [2449, 1617, 2986, 3263]. Joyce [606]. Joynson [1147]. Jozsef [2766]. Jr [1323, 2634]. Jr. [1998, 74, 2967, 1464, 2998, 2022, 3296, 1497, 2854]. Juan [321, 526, 2334, 239, 2272, 3442, 154, 1197, 2656, 637]. Juanelo [1374]. Jubilee [2452]. Judd [3448, 3357]. Jude [2122]. judios [3053]. judischer [2523]. Judith [3274, 2100, 3163, 1218, 1519, 2841, 1458, 3359]. Judson [2838, 1270]. Judy [3120, 1958]. Juhani [3491]. Juhasz [72]. juin [3538]. juive [642]. Juli [2661]. Julian [622, 1066, 3440, 2197, 2024, 1652, 2202, 2202]. Julie [809, 3087, 3435, 1492]. Julien [2664, 2352]. Juliette [2667]. Julio [2681, 2334, 1633, 2072]. Julius [3447, 266, 476, 2292, 1846, 476]. Jullien [3056]. July [1889, 2634, 3030, 1087, 2956]. Juma [715]. Jumping [1134]. Jun [507]. June [2692, 2313, 3030, 372, 1947, 2985, 1394, 2057, 897]. Jung [2295, 657]. jungen [163]. Jungenschriften [2979]. Jungius [1923, 1113, 648]. Jungmayr [1300]. Jungnickel [3065, 614]. Juni [1531]. Junker [2987, 663]. Junto [3218]. Jupiter [740, 1728]. Jurgen [1474, 2460, 845, 2117, 1865, 1289, 585, 439, 489, 758, 196]. Jurisprudence [1577]. juristischen [3341]. Jussieu [2441]. Just [536]. Justice [225]. Justification [2042]. Justifying [3236]. Justin [1572]. Justus [266, 2280, 1572, 266, 2280, 3155].

Kabbalah [3245, 2660, 156]. Kac [187]. Kafker [395, 2177, 395]. Kagaku [902]. Kagakusi [507]. Kahane [1484, 659]. Kai [2013, 3552]. Kaiser [3367, 1215, 944, 909, 1599, 1215]. Kaiser-Wilhelm- [3367, 1215]. Kaiser-Wilhelm-/Max-Planck-Gesellschaft [3367, 1215]. Kaiser-Wilhelm-Gesellschaft [3367, 1215]. Kaiser-Wilhelm-Instituts [1215]. Kaiserfeld [2036]. Kaiserreich [2914]. Kalahari [3281]. Kalam [2246]. Kalckar [3276]. Kalendarium [2987]. Kalle [513]. Kallinen [2264]. Kalt [1463]. Kalthoff [2609]. Kambartel [463]. Kamerlingh [1214]. Kameshwar [863]. Kamminga [3464, 1173]. Kandall [2927]. Kanigel [1213]. Kannegiesser [385]. Kansas [620, 3303]. Kant [3062, 163, 3064, 1601, 937, 164, 3063, 861, 362, 2094, 2572, 1567]. Kantorovich [1683]. Kants [2818, 2818]. Kanz [2013, 3552]. Kapital [755]. Kapitel [530, 471]. Kapitza [1217, 1217]. Kaplan [1926]. kara [902]. Karamustafa [1452]. Karasszon [424]. Kardel [2510]. Karen [2548, 3543, 2035, 3245, 2282, 3511, 3126, 2928]. Kargon [682]. Kari [2162]. Karin [2755]. Karl [2834, 632, 1171, 358, 2199, 1935, 2291, 2868, 3327, 787, 3485, 1747, 2195, 3112, 2474]. Karl-Heinz [3327]. Karlfried [89]. Karlsbader [1123]. Karnad [3473]. Karpiński [979]. Karrieren [1410]. Karrow [1702, 1998]. Kass [626]. Kass-Simon [626]. Katahiro [2436]. Katalin [1067]. Kate [1458]. Kater [677]. Katerina [1001]. Katherine [810, 3293, 3190, 2200, 3162, 462]. Kathleen [3160, 1387, 1573]. Kathren [2540]. Kathryn [1210]. Katz [2698, 1533]. Kauffman [2536, 2130]. Kaufman [1759]. Kaufmann [2827, 1919]. Kavanagh [686]. Kay [1593, 698]. Kaye [2840]. Kazemi [1215]. Keaney [1906]. Kearns [543]. Keating [1133, 2019, 3093]. Keats [1125]. Keel [229]. Keeney [1391]. Keeping [3098]. Kees [945]. Keiji [2426]. Keil [518]. Keimpe [2417]. Keir [3108]. Keith [2293, 3406, 3536, 1947, 2369, 2786, 3289, 1863, 1318, 416]. Keller [1606, 2608, 2495]. Kelley [1469]. Kellner [1337]. Kelly [1644, 2536, 1981, 482]. Kelvin [846]. Kemp [903, 1010, 2807]. Ken [3443, 2932]. Kendall [3348, 2288]. Ken’ichi [1367]. Kenkyu [507]. Kennard [2201]. Kennedy [2130, 32]. Kenneth [2148, 1664, 1299, 1364, 3099, 1682, 2202, 3319]. Kenney [1133]. Kenney-Wallace [1133]. Kennis [658]. Kenny [1111, 1449]. Kent [753]. Keown [481]. Kepler [2082, 905, 1474, 2083, 1650, 1728, 3055, 354, 1894, 1105, 158, 2258]. Keplero [973]. Kerferd [428]. Kernbauer [475]. Kernforschung [472]. kernphysikalischer [2923]. Kerry [285, 2706]. Kerszberg [671]. Kertz [2646]. Kérygmes [146]. Kesaris [1378]. Kessler [3190]. Ketelsen [2679]. Kettler [2474]. Kevin [1001, 3008, 19, 3572, 3541]. Kevles [2848]. Keweenaw [1495]. Kexueshi [507]. Key [972, 2198]. Keynes [366, 2774]. Keys [2159]. Keywords [1606]. Kh [887]. Khaled [510]. Khazin [2877]. Khazini [1810]. Khwarizmi [1631, 641]. Kicking [2224]. Kidd [3190]. Kidwell [2215, 1544]. Kieckhefer [732]. Kielmeyer [2013, 2013]. Kielmeyer-Bibliographie [2013]. Kilbourne [333, 3032]. Killam [1133]. Killing [2318, 3231]. Kilmister [2198]. Kim [2294]. Kimberly [2037]. Kimura [2208]. kindergeneeskunde [2975]. King [3346, 360, 32, 2276, 2203, 669, 2369, 1909, 31]. Kingdom [3115, 679]. Kings [2968]. Kingsley [2887]. Kipnis [3546, 1121]. Kirdly [59]. Kirk [3379]. Kirke [1323]. Kirschner [3246]. Kirsti [515]. Kirtley [2201]. Kirwan [2275]. Kisin [2120]. Kitab [1005, 2071]. Kitcher [783, 1773, 1948]. Klaas [3279]. Klassifikation [2095]. Klassiker [654]. Klassische [654]. klassischen [360, 755]. Klaus [2018, 1313, 940, 3053, 239, 2153, 3253, 1813, 2117, 2502, 2606, 1314, 225]. Klaus-Dieter [3253]. Klaus-Peter [940, 3053]. Klausner [1281]. Klaver [3353]. Klein [2282, 2460, 2117, 2766, 1927, 809, 3120]. kleinen [875]. Kleinman [2541]. kleinsten [842]. Klemke [79]. Klenke [3367]. Klibansky [1280]. Kliment [570]. Kline [1617, 606, 16, 79]. Klingemann [2842]. Klonk [2819]. Kmita [977]. Knew [1213, 3547]. Knight [3544, 1009, 328, 1569, 1349, 3539]. Knobloch [113]. Knorr [723, 723]. Know [2563, 2044, 2469]. Know-How [2563]. Knowing [462, 299].

Knowledge [819, 2527, 2314, 555, 833, 1407, 1469, 1832, 735, 975, 941, 2964, 1056, 2448, 1062, 3313, 1208, 2598, 2666, 2563, 1585, 1214, 3092, 1757, 685, 2551, 2298, 2047, 1627, 3400, 406, 3185, 808, 3485, 882, 82, 16, 2698, 1961, 294, 3229, 205, 2999, 1083, 3505, 82, 3522, 816, 2796, 753]. Known [554]. Knows [883]. Knox [1673]. Knudsen [515, 3543]. Knudtson [238]. Knut [3467]. Koch [2519, 469, 998]. Koetsier [1420]. Kohl [1371, 2990]. Kohlenforschung [1215]. Kohler [951, 2881]. Kohlstedt [3317, 3037, 1203]. Kolbe [1751, 2280]. Kolbert [1240]. Kollegialverfassung [1941]. Kolloquiums [2185]. Kolumbus [1635]. Komenskiego [1655]. Kommentar [3246, 439, 2067, 822]. kommentierte [394]. Kommentierung [1855]. Kommunikationsmedium [1316]. Kompendium [1638]. komputus [1993]. Konformation [654]. Konig [2757]. Königreich [325]. Konigsberg [1210]. Konrad [2549]. Konrads [2251]. Kontext [2154]. Kontinuitat [1993]. kontra [1123]. Konzepte [2465]. Konzeptionen [2754]. Kopelevich [887]. Kordesch [1978]. Korean [507]. Koreny [350]. Korolev [3009]. Koroth [509]. korrekten [2153]. Korrespondenz [3080]. korrespondiert [2265]. Kosinski [639]. Kosmologie [1185, 297]. Kosmos [1635, 2336, 2411]. Kosmos-Modells [2336]. Kosso [1423]. Kostas [488, 1214, 2533]. Kowalewskaja [2195, 2195]. Kox [2460, 2117, 2766]. Koyre [2872]. Kozhamthadam [2083]. Kozlov [511]. Kozulin [1049]. krachtbegrip [2092]. Kraemer [1184]. Krafft [517, 518]. Kraft [174]. Krafte [2013]. Kraftkonzeptionen [842]. Kragh [770, 3544, 2831]. Kramer [2252, 2780]. Kramers [3461]. Kranakis [2823]. Krankengeschichten [2067]. Krankheit [1813]. Kranzberg [2401]. Krasnick [865]. Krause [1501, 1495]. Krausse [476]. Kraut [1949]. Krayer [1651]. Kraynak [832]. Krebs [3299, 2191]. Kreislaufs [2507]. Krementsov [3560]. Kremer [938]. Kren [1068]. Krieger [1774]. Krige [1046]. Krimsky [1265]. Krisciunas [19]. Krise [1322]. Krishna [2202]. Kristalle [536, 358]. Kristallgeometrie [673]. kristallographischen [536]. Kristie [1857]. Kriterion [428]. Kritik [663, 1813, 1240, 1658]. Kritische [648, 1500]. kritischer [730]. Krivatsy [492]. Kroll [196, 1295]. Kronick [1427]. Krook [2087]. Krop [1368]. Krüger [418]. Krupp [2968]. Kubica [65]. Kudlick [2671]. Kudlien [1011]. Kuehn [258]. Kuei [128]. Kuei-Mei [128]. Kugler [1278]. Kuhl [1951]. Kuhlmann [3531]. Kuhn [3198, 1603, 3030, 2390, 3265, 329, 2742, 1530]. Kuhne [1778, 2506]. Kuhnert [654]. Kuhnke [760]. Kuhns [786]. Kuklick [2881, 1490]. Kultur [837]. kulturelles [3385]. Kulturkampfe [2056]. Kumar [1543, 2669, 1551]. Kümmel [518]. Kunitzsch [905, 2973, 1360]. Kunst [2754]. Kuper [538]. Kurt [1830, 866, 2998, 2471, 3469, 3338, 365, 1029, 515, 348, 1283, 196]. Kurt-R [1830, 365, 1029]. kurz [1013]. Kurzman [225]. Kusch [1242]. Kushner [1807]. Kusukawa [2427]. Kusyar [3432]. Kutchins [3379]. Kwahaksa [507]. Kwartalnik [512]. Kwok [1800]. Kyung [2294]. Kyung-Man [2294].

L [709, 2453, 3419, 2383, 939, 869, 1673, 115, 961, 2493, 1534, 1597, 2622, 2591, 180, 2302, 744, 835, 910, 1000, 3073, 1045, 1680, 2094, 1595, 106, 2273, 1184, 1323, 2906, 1688, 2026, 2963, 1664, 446, 3038, 841, 3254, 185, 2690, 3407, 464, 1096, 2594, 1134, 2080, 145, 268, 1217, 1097, 2317, 470, 3230, 2840, 938, 2512, 226, 1585, 3483, 2038, 18, 2655, 304, 349, 589, 1226, 1197, 2935, 144, 1335, 2489, 2670, 1938, 792, 2830, 2609, 1592, 2540, 2002, 2045, 872, 493, 3393, 667, 1735, 3571, 2728, 395, 2176, 1275, 3359, 3120, 531, 1131, 1977, 225, 1478]. L [369, 1979, 1903, 3550, 1750, 1584, 375, 2731, 2753, 1234, 2644, 2292, 1992, 1426, 3459, 1416, 1188, 1756, 596]. Laass [755]. Lab [2711, 1969]. Labban [3432]. L’Abbé [460]. Labor [2993, 593, 75]. Laboratories [1696, 581]. Laboratory [1967, 1611, 3370, 1514, 1947, 323, 1494, 1868, 759, 2711, 1131, 2855, 1253, 3223, 2039, 1046, 2683, 3171, 3178, 1226]. Labour [3161, 2652]. Lacaille [1298]. Lacan [1054, 1744]. Lacarra [3053]. Lachterman [719]. Lacki [3458]. Lacombe [535]. Lactic [1135]. Lacy [3240]. Ladous [2618]. Lady [1262]. Laennec [3266]. LaFollette [2937, 877, 1511]. Lafuente [321, 1735, 563]. lagartos [2027]. Lagneau [1307]. Lagrange [2354, 2906, 842, 2092]. Lai [191]. Laidler [2293, 3406]. Laing [3468]. Laird [2248]. Lajos [59]. Lakato [1420]. Lakatos [2390, 3486]. Lakes [2990]. Lam [1712]. Lamandé [653]. Lamar [1126]. Lamarck [49]. Lambert [259, 3508]. L’ambigua [1643]. Lambright [2388]. Lamens [3068]. Lamens-van [3068]. Lamp [3140]. Lampert [1775]. L’anatomie [423]. Land [3334, 3163, 573, 2203, 201, 1447]. Landau [570, 2160, 2675, 907]. Landau-Tylkina [570]. Landers [2091]. landlichen [3150]. Landmarks [2621, 2215]. Landriani [2980, 2980]. Lands [3448]. Landsat [969]. Landscape [3274, 2819]. Landscapes [18]. Landwirtschaftlichen [266]. Lane [1041]. Lanfranco [771]. Lang [1552, 2323]. langage [2188, 1412]. Lange [2830]. Langham [953]. Langholm [3247]. Langley [2387]. Langlois [1571]. Langstaff [1839]. Language [2344, 218, 1669, 2113, 1527, 3319, 1006, 2845, 3469, 303]. Languages [3249, 2429, 2899]. langue [3189]. Lankester [2451]. Lankford [3107]. Lanouette [1581]. l’Antiquité [821]. Lanz [1040]. Lapidaire [146]. lapidaires [146]. Lapidge [2582]. Laplace [3405, 3251, 3535]. Laplanche [1571]. Lapland [893]. Laporte [2182]. l’arc [3412]. l’arc-en-ciel [3412]. Lärdomshistoriska [514]. LARG [1338]. Large [244, 1328]. Large-Scale [1328]. Largeault [2528]. Larissa [2081]. L’arithmétique [2165]. Larkin [1039]. L’armonia [2526]. Larrabee [3006]. Larry [2027, 972, 1479, 1491, 1502, 780, 3176, 2676]. Lars [3071]. Larsen [3163, 1415, 3044, 1987]. Larson [2546, 2094, 674, 3363]. Larvor [3486]. Lasby [3183]. Laser [1219]. Laslett [3037]. Last [3547, 632, 2563, 3294]. L’astrolabe [1814]. L’astronomia [1917]. Late [2177, 3160, 1765, 2654, 2571, 2096, 1003, 548, 595, 179, 448, 764, 2025, 2819, 483, 946, 2971, 611, 3400, 2398, 3399]. Late-Enlightenment [2096]. Lateinamerika [1701]. Lateinisch [2502]. Lateinisch-deutsch [2502]. lateinische [2973, 1360]. lateinischen [2332]. Later [2689, 1636, 314]. Latham [39, 1275]. Latin [39, 2075, 1913, 1326, 2426, 2499, 1367, 641, 33, 3330, 452]. latine [727, 914, 718]. Latinoamericana [526]. Latinoamericano [525]. latins [210]. Latinus [1716, 2331]. L’atomo [2518]. Latour [1796, 2642, 974]. Lattice [3223]. Lattis [2650]. L’aube [2960]. Laudan [972, 3425]. Laugier [2539]. Launch [3176]. Launching [2921]. Laura [2978, 2525, 2078, 2025, 2117, 1046, 613, 1342]. Laurance [3283]. Laureate [1958]. Laureates [1133]. Laurel [3348, 2152, 1301]. Laurence [2348, 3043, 1775, 590]. Laurens [935, 3433]. Laurent [2596, 3272, 2441, 3556, 41]. Laures [2738]. Laurie [3562, 3166, 2640]. Laurikainen [189]. Laurinda [2897, 2339]. Lautman [3583]. L’autorite [2573]. Lavater [2361]. LaVerne [760]. Lavoiser [3146]. Lavoisier [3255, 3146, 1744, 3538, 2446, 659, 2274, 41, 2399, 99, 98, 2596, 1801, 2485]. Lavoisieriana [2596]. Law [856, 3079, 2479, 1382, 2389, 2738, 3288, 3008, 1577, 1050, 1618, 2345, 481, 452, 974]. Lawn [2076]. Lawren [225]. Lawrence [3255, 1576, 1617, 3563, 15, 1685, 1184, 3259, 1834, 2412, 867, 2607, 565, 598, 808, 3229, 2190, 171, 1131, 2688, 2191, 2758, 2666, 1131]. Lawrie [3236]. Laws [2083, 3308, 2595, 880]. Lay [1712, 3380]. Laycock [2603]. Layne [2045]. Lazare [2098]. Lazarsfeld [3583]. Lazarus [3273]. Leader [2995]. Leads [385]. Leah [2026]. Lear [2685]. Learned [1557, 3176, 1211]. Learning [1626, 1673, 2801, 1193, 10, 440, 1275, 2351, 542, 2982]. Leary [778, 2566, 3135]. L’eau [1004]. Leavey [1064]. Lebedev [1212]. Lebedeva [1212]. Leben [851, 2123, 1300, 3441, 365, 267]. Lebensbild [1762]. Lebensbilder [89]. Lebrun [3528]. Lecas [3556]. lecteurs [2592]. Lectures [1308, 2416, 1409, 1130, 1394, 2780, 3033]. Lecuyer [3583]. Lederer [2612]. Lederman [1866]. Lee [2453, 2283, 3083, 3585, 1239, 2130, 2541, 188, 2844, 188, 2844]. Leere [41]. Leeuwarden [2515]. Leeuwenhoek [835, 2512, 2511]. Leeuwenhoek-Commission [835]. Leewenhorst [160]. Left [2315]. Legacies [1820]. Legacy [789, 2032, 1617, 3370, 2283, 1216, 1502, 3557, 3288, 3247, 2670, 3462]. legado [2334]. Legal [481]. Legee [1496]. Legem [2000]. Legend [1773]. Legends [3449]. legge [1737, 853]. Legible [2361]. Legislators [1209]. Lehrdichtung [1640]. Lehre [2503, 442]. Leibarztes [3060]. Leibniz [161, 1656, 160, 2508, 2173, 2085, 837, 450, 747, 1379, 2265, 1110, 1531, 2661, 1257, 2089, 3137, 2397, 2348, 2172, 450, 1647, 3138, 559, 1874, 359, 2660, 836, 2265]. Leibniz-Gesellschaft [160]. leibnizien [2348]. Leibowitz [509]. Leibrock [2081]. Leibrock-Plehn [2081]. Leichty [248]. Leiden [3140, 3398, 3068]. Leidy [3547]. Leigh [664]. Leighow [3292]. Leila [3377]. Leiss [1088]. Leisure [2939]. Leitner [3567]. Lejeune [914]. Leland [2039]. Lelorrain [212]. Lemaître [3460, 3460, 2769]. Lemay [1715]. Lemieux [1134]. Lemoine [184]. Lemon [3326]. Leneaux [1008]. Lenn [1910]. Lennon [1820, 1820]. Lennox [2065]. Lenoir [843, 3589]. L’enseignement [212]. Lenses [219]. Leo [1581, 2150, 2207, 246]. Leon [1866, 1744, 1033, 1390, 3576, 3245, 2340, 525]. Leon-Jones [3245]. Leona [646]. Leonard [3547, 304, 2621]. Leonello [2756]. Leonhard [2903, 3534, 842]. Leonhardi [1738]. Leoniceno [1556]. Leplin [3297]. ler [3538]. Lerner [1594, 2432]. Lesley [2066]. Leslie [2343, 3019, 3079, 2392, 3018, 225, 1227, 794, 1105, 1954]. L’esperienza [2893]. L’espressione [443]. Less [1799]. Lessico [643]. Lessons [67, 3215, 3176, 1343]. Lester [1239, 2451]. L’estime [2508]. Let [37]. Leta [1323]. Lethal [1707]. Letter [253, 1974, 2141, 895, 2139]. Letteratura [3332]. Lettere [1101, 2756]. Letters [1626, 2229, 3501, 571, 835, 3031, 2402, 615, 2380, 2635, 3226, 3259, 3254, 102, 1698, 806, 9, 899, 1794, 1217, 2957, 3456, 2512, 313, 410, 1938, 1168, 708, 943, 2058, 566, 166, 1749, 224, 2716, 1487, 2484, 1346, 2351, 2211, 1615, 1795, 2835, 939, 1945, 938, 1378]. Lettie [3061]. Lettinck [2244]. lettre [2326]. Lettres [1808, 161]. Leurquin [2581, 2581]. Leuvense [361]. LeVay [2701]. Levels [1584]. levelse [59]. Leven [3327]. Levenstein [138]. Leverington [2529]. Levi [920]. Levin [2385]. Levina [2124]. Levine [1603, 828, 2129]. Leviton [3073]. Levy [1044, 1853]. Lewin [2531]. Lewis [130, 3339, 2032, 3131, 1476, 2955, 686, 1670, 2456, 596, 3414, 332, 287, 596, 2617]. Lexicography [3430]. Lexikon [345]. Lexique [3189]. lexiques [727]. L’expedition [935]. Leyden [452]. Leys [1326, 779]. L’harmonie [2348]. L’héritage [545]. l’homme [2743, 1875]. l’Huillier [1369]. Liadain [1908]. Liao [1186]. Liba [1548]. Liber [435, 2583, 730, 3431, 1634, 727]. Liberal [3326, 3365]. liberales [1993]. Liberalism [2474, 1509]. liberalismo [752]. Liberation [483]. libertinage [2166]. Libertine [1562]. Liberty [445, 3231]. Libraries [1607, 1689]. Library [2596, 1066, 1702, 492, 2739, 1247, 1688, 2178, 2377, 1566, 301, 1607, 1815, 1248, 646, 2159, 1750, 1338, 3490, 698, 2621, 699, 599]. Library-Anthropology [1338]. Libres [2320]. Libri [2586, 1918, 1556]. Libro [2103]. Libros [2491, 154, 1470]. Licences [493]. Licht [165, 3131]. Lichtenthaeler [549, 2067]. Licoppe [2514]. Liebenau [23]. Lieberknecht [2503]. Lieberman [3576, 3478]. Liebig [266, 2280, 1337, 3155, 1572]. Liebigs [1572]. Lieburg [516, 994, 2975, 2693]. Lies [864]. Liesegang [654]. Lieutenant [1041]. Lieutenant-General [1041]. lieux [388, 2970]. Life [57, 3104, 1213, 3339, 281, 2531, 2069, 2314, 568, 3573, 3271, 1617, 3160, 3356, 2209, 2381, 3148, 1731, 876, 2205, 2608, 2510, 2455, 1596, 2084, 3086, 433, 3270, 2998, 1833, 1899, 3536, 1936, 1003, 2835, 2689, 693, 2690, 2372, 3083, 2771, 3266, 1217, 2537, 1506, 1593, 3090, 3167, 1959, 1632, 1906, 1417, 91, 2724, 469, 950, 1497, 323, 505, 1225, 2830, 3450, 2772, 7, 3059, 3342, 3468, 2112, 964, 3345, 3251, 2243, 2706, 2126, 13, 2995, 3012, 597, 1098, 2202, 1272, 1227, 2307, 282, 1041, 3462, 698, 1503, 3574, 1578, 741, 1301, 2639, 418]. Life [997, 1740, 2406, 1839, 3374, 3015, 1831, 434, 1862, 2720, 1975, 2316, 2619, 684, 3450, 379, 2041]. Lifetime [2837, 3285]. Lifetimes [192]. Lift [74]. Lift-Off [74]. Lifting [3482]. Light [539, 990, 3406, 36, 1929, 3008, 371, 747, 2466, 964, 2713, 2770, 2174, 173, 1306, 3035, 1121, 734]. Lightfoot [750]. Lightman [990, 3069]. Liliane [1649]. Lilienthal [3365]. Lilla [2885]. Lilli [2535]. Lillian [3562, 2125, 1865]. Lilly [2344, 597]. Lily [1593, 698]. Lima [1654, 793]. Lima-de-Faria [793]. Limbrick [554]. Limited [2607, 56]. limites [2247]. Limits [3326, 3294, 3058, 2771, 1289, 2882]. Limoges [118]. Limon [722]. Lin [507, 2615]. Lincecum [3074, 3074]. Lincoln [2683]. Lind [1188]. Linda [2726, 2767, 304, 2045, 679, 1601, 2685]. Lindberg [1100, 3335, 1797]. Linde [1572, 1572]. Lindeboom [240]. Lindee [2384]. Lindemann [3151]. Linden [1557, 2810]. Lindley [1334]. Lindqvist [2036]. Lindroth [320]. Lindsay [1564, 1270, 815]. Lindy [3286]. Line [1434, 1696, 3159]. Lines [3356]. l’infini [2437]. L’influsso [645]. lingua [3332]. Linguist [777]. Linguistics [392]. Link [1359, 368]. Linnaeus [2094, 2440]. Linne [1930, 1659]. Linnean [232]. l’Inquisition [3340]. Linus [2772]. L’inventario [1471]. L’invention [1753, 2031]. Lipson [1221]. Lire [2333]. Lisa [2659, 2777, 1219, 484, 1482]. Lise [1459, 2531, 1855]. Liselotte [1225]. List [803, 3393]. Liste [90]. Lister [1477, 3482]. Lists [87]. Literacy [440]. Literary [2104, 1872, 808]. Literate [1295]. Literatur [2920, 2013, 2920]. Literature [810, 885, 1817, 85, 805, 2902, 83, 722, 2362, 3344, 1669, 2810, 1892, 3353, 2409, 542]. Litten [2123]. litteraires [2330]. litterature [3298]. littereria [1923]. Liudger [1946]. Live [3274, 2792]. Lived [2605]. Liver [2346]. Lives [2833, 2169, 3456, 756, 3162, 1848, 225, 3238, 1062, 2349, 3314, 760]. Livesey [2077, 1279]. Living [1324, 2094, 2917, 3215, 2278, 3112, 3170, 3510, 2243, 2724, 1838, 2784]. Livingstone [1453, 1455]. Livre [2581]. Livres [2242, 3132, 1361, 3340, 1176]. Lloyd [621, 225, 631, 2116, 1606, 2969, 1385, 999]. Llull [524]. Load [959]. Lob [40]. Localisations [2188]. Locality [2481]. Localization [1876]. Localizing [721]. Locations [83]. Lochlainn [2919]. Lock [686]. Locke [452, 1515]. Locqueneux [3067]. lode [613]. Loeb [281, 2116, 199, 1768]. Loeffel [357]. Logan [2996]. Logic [2289, 489, 2151, 2153, 1683, 737, 1006, 344, 3319]. logica [2526, 3129]. Logical [2289, 2998]. Logics [1675]. logique [1063, 3446, 1486]. logiques [1063]. logischen [2153]. Logos [1549]. Logsdon [2685]. Lohff [1122]. Lohmar [1150]. Lohn [12]. Lohrmann [2150]. Lois [1457, 1899, 1009, 3191]. Lola [2765]. Lomax [2828]. Londa [446, 2513]. London [2091, 1087, 301, 2533, 2088, 2666, 232, 467]. Long [2989, 1169, 407, 815, 68]. Longer [839]. Longino [3037, 3317, 685]. Longitude [3252, 1019]. Look [3204, 1530]. Lopez [1911, 2516, 2491, 2490, 3442, 1561, 2158, 1354, 1311, 2435, 3139, 1024, 132, 351, 152, 829]. L’ophtalmologie [2330]. L’Optique [914]. Lord [846, 1106]. Lords [2316]. Loren [1864, 1803, 954, 954]. Lorenz [418, 2549]. L’organisation [657]. L’origine [1713]. Loris [2890]. Lorne [3108]. lornetki [3455]. L’oro [1915]. Lorraine [418]. Lorscher [1638]. Lost [3502, 724, 1755, 1090]. Lothar [654]. Lotze [2526]. Louden [3042]. Loudon [1742]. Louis [2237, 2302, 2275, 2442, 2102, 2317, 1089, 3438, 686, 355, 2539, 1861, 2159, 842, 545, 2457, 672, 1043]. Louis-Bernard [2275]. Louise [1878, 2199, 3386, 546, 2733, 2065]. Loup [2350]. lourds [339]. Louvain [3460]. Louvain-la-Neuve [3460]. Loux [1269]. Love [3472, 2642, 1612, 1578, 2047]. Lovers [1626]. Lovesickness [826]. Low [488, 744, 697, 713]. Lowe [766]. Lowry [1549]. Lowy [3291]. LSD [2566]. Lubbock [3025]. Lubrication [2053]. Luc [2554, 3427, 2895, 1178, 1237]. Luca [2905, 2750]. Lucas [1236]. Luccarelli [2617]. Luciano [2063]. Lucidator [919]. Lucier [3259]. Lucille [854]. Luck [2965]. Ludmilla [623]. Ludwig [1130, 1501, 2457, 812, 755, 3352, 2689, 755]. Ludy [1323]. Luft [2281]. Luftfard [1442]. Lugt [2983]. Luigi [1101, 1920, 1779, 643, 2901]. Luis [1911, 868, 752, 1828, 2600, 2158, 2340, 2249, 746, 637]. Luiz [2027]. Luke [438]. Lull [1914, 552]. Lulofs [1913]. lumière [3340]. lumières [460]. l’un [2895]. Luna [1738]. Lunar [583, 2152]. Lunbeck [2298]. Lundquist [3333]. lunety [3455]. luoghi [1471]. Lupus [1811]. Luqa [1718, 339]. Lurcat [2573]. Lust [301]. Luther [254, 602]. Lutwidge [2456]. Lutz [1721]. Lux [745]. Luyendijk [516, 996, 504]. Luyendijk-Elshout [516, 996, 504]. Luz [1911, 2491, 2982, 2158]. Lychnos [514]. Lydia [141, 922]. Lyell [2136]. Lynch [2552]. Lynda [110, 3505]. Lyndsay [684]. Lynette [3428]. Lynn [2653, 3553, 2470, 3176, 883, 2364, 369]. Lynne [2706]. Lynwood [1089]. Lyon [1315, 14]. Lyons [1508]. Lyra [2419]. Lysenko [2207].

M [1706, 3039, 1336, 1911, 1024, 2764, 1484, 201, 734, 2384, 2189, 1594, 3520, 1306, 846, 2464, 2075, 1210, 233, 2449, 874, 1617, 482, 516, 2828, 172, 2491, 1007, 3349, 287, 2532, 1105, 582, 2205, 429, 1913, 2406, 480, 70, 3078, 967, 669, 2156, 3232, 2975, 2723, 2619, 1491, 1688, 2759, 3442, 3429, 1667, 975, 1119, 2963, 1561, 1723, 3166, 1833, 1838, 3068, 2851, 504, 195, 3124, 185, 2166, 1502, 1418, 994, 1130, 484, 1655, 2650, 1748, 569, 1223, 593, 996, 1097, 3033, 2124, 442, 3365, 3013, 1535, 3438, 1700, 2040, 1906, 3239, 349, 2266, 1197]. M [3333, 2935, 3343, 3455, 3117, 1739, 2101, 3011, 1386, 1831, 3485, 2722, 355, 1820, 1554, 1747, 117, 3539, 3548, 1354, 1732, 1949, 3047, 2362, 1428, 1735, 829, 1311, 2435, 2373, 3143, 182, 1275, 1736, 3359, 3263, 2640, 225, 1487, 1778, 2844, 3012, 3296, 2685, 1973, 1760, 2760, 2730, 114, 3353, 1320, 3055, 1992, 2253, 384, 3171, 2041, 2693, 737, 596, 1777]. M.D [2622, 1825]. mâchoires [820]. ma [1737]. Maarten [2253]. Maatschappij [2693]. Mabahiss [961]. macchina [2241]. macchine [2098, 2155]. Maccubbin [43]. MacDougal [612]. Macfarlane [1417]. Mach [804]. Machian [622]. Machiavelli [1647]. Machiavellian [1372]. Machine [997, 869, 3014, 308, 2234, 1494, 2775, 3496, 2352, 2874, 2664]. Machinery [1708, 445]. Machines [2080, 3081, 1544, 975, 1984, 2105, 419]. Macht [472]. Mack [1707, 969]. Mackay [879]. MacKenzie [3358, 1055]. MacLachlan [2872, 415]. MacLeod [1133, 332, 131, 44]. Maclure [264, 264]. Macmillan [1146]. Macoun [265]. MacPherson [225]. Macrakis [1857]. Macrobius [1274]. Macromolecular [3569]. Macromolecules [279]. Mad [528, 1324, 3358, 2360, 3581]. Mad-Doctoring [3358]. Mad-Doctors [2360]. Madame [3346, 862]. Madders [3008]. Maddin [25]. Made [2384, 2301, 3006, 1134, 2767, 1852, 902, 225, 484, 1538]. Madison [896, 597, 2748, 1047]. Madmen [3238]. Madness [856, 1324, 1299, 1208, 2360, 1037, 3468, 335, 1743, 2673, 572, 2470]. Madrid [897]. Maehle [1658]. Maestlin [2308]. maestri [1267]. Maffioli [744, 2167]. Magazine [515, 2520, 3540]. Magda [1846]. Magdeburg [1924]. Magee [3161]. Magellan [1450]. Maghout [510]. magia [1643]. Magic [1553, 926, 733, 1192, 2887, 156, 732, 1850, 732, 3433, 732]. Magicians [3245]. Magier [1283]. Magierin [1283]. magiques [2165]. Magna [3248]. Magner [1457, 1899, 3191]. Magnetic [2778, 3417, 2817]. Magnetism [2438, 603, 1534]. magnétisme [1043]. Magnitude [583]. Magnolias [2673]. Magnus [1715, 2502, 1642]. Maienschein [1318, 1128, 323]. Maier [2033]. Maija [2264]. Maila [1045]. Maimonides [1184]. Main [3008]. Mainauer [2154]. Mainz [1651]. Maisel [35]. Maistre [3528]. Maistrov [910]. maitres [2820]. Majestad [1374]. Major [3274, 1990]. Make [3096, 703]. Maker [97, 2515, 1936]. Makers [2285, 1752, 3456, 2838]. Making [1135, 396, 2100, 74, 881, 843, 3454, 3087, 372, 3152, 28, 1971, 2109, 772, 2459, 1495, 2451, 225, 2688, 2513, 3544, 1979, 2714, 697, 3435, 2878, 2520, 3540, 3400, 1050, 1085, 3182, 3379, 3313, 2278, 877, 2296, 3012, 1802, 60]. Makinosuke [2537]. mal [2019]. Maladies [1564, 1412]. Malady [1118, 3516]. Malaria [2224]. Malariologist [3269]. Malcolm [882, 3414, 3163, 1477, 1146, 3006, 225]. Male [1754]. Malebranche [1476, 450, 747, 1476]. Malenstein [3068]. Males [2544]. Malet [2896, 3523]. Mali [1823]. Malin [2729]. Malinowski [65]. malle [2350, 2843]. Maloney [436]. Malphrus [2684]. Maltese [1737]. Malthus [1170]. Malusa [2063]. Malvezzi [2271]. Mamie [3099]. Mammoths [1940]. mammouth [2147]. Man [1213, 1581, 3249, 2294, 3547, 3577, 2429, 479, 3009, 1169, 583, 1853, 285, 2916, 3152, 2440, 1620, 1852, 2664, 766, 2986, 368, 2352, 2471, 2352]. Man-Midwifery [3152]. Management [2214, 2616, 1087, 3325, 276, 3326]. Managing [2631]. Mancosu [2588]. Mandarins [2602]. Mandel [2687]. Mandelbaum [49, 1827]. Mandrino [2355]. Manfred [258, 364, 1123, 1215]. Manfredo [828]. Mangle [2550]. Manhattan [3170, 225, 966, 225]. Maniacal [572]. Manias [2453]. maniera [1002]. Manifeste [2342]. Mankynde [1248]. Mann [1087, 518]. Manned [474]. Manner [3130, 1381]. Mannheim [2474]. Manno [2098]. Manual [1461, 431]. Manuals [3401]. Manuel [2645, 1331, 3524, 2090, 1311, 1040, 563]. Manufactures [1038]. Manufacturing [1957, 1407]. Manuscript [2549, 696, 2816, 606, 1612, 347, 599, 698, 2504]. Manuscripta [354]. Manuscripts [2346, 1656, 838, 1816, 1519, 1378, 888, 1689, 1778, 3490, 1291, 1778, 1778]. manuscrits [2508]. Many [386]. Maor [1985]. Map [3414, 1702]. Mapmakers [1998]. Mapmaking [3414]. mappe [435]. Mapping [2785, 1618, 2813, 3400, 3258]. Maps [1434, 1649, 1702, 1998, 85, 2813, 1537, 725, 1999, 3503]. Maquet [2510]. máquinas [154, 1040]. mar [3445]. marasme [3509]. Marat [3066]. Marazzini [3277]. Marc [3271, 1091, 2508, 193, 2847, 1573, 1026, 1877, 1778].

Marc-Auguste [1026]. Marcel [1004, 229, 877, 1511]. Marcella [2793]. Marcello [2391, 1827]. Marcelo [3154, 2517]. Marcet [405]. March [2312, 2625, 707, 1947, 1890, 706, 2057, 1791, 3402]. Marchand [2728]. Marchia [1185]. Marci [1107]. Marcia [38, 2367, 2184, 1352]. Marco [2596, 1284, 1801, 1589, 2756, 2274, 2518, 2270]. Marcus [1107, 817, 2212]. Marelene [1617, 3174]. Marganne [2330]. Margaret [201, 2799, 1934, 885, 2504, 159, 1924, 576, 3121, 2467, 2261, 1383, 1030, 1225, 1117, 1264, 3487, 2268, 2325, 3312, 3532]. Margarita [3386]. Margin [2672]. Marginalia [1169]. Marginality [2844]. Marginalization [3087, 3498]. Margins [2349]. Margit [3568]. Margo [595]. Margolis [1771]. Margot [3477]. Margritt [260]. Marguerite [1454, 1198]. Mari [2285, 2741]. Maria [248, 3475, 132, 1374, 2516, 2491, 2982, 3053, 645, 1284, 973, 2490, 2158, 2271, 1808, 351, 152, 2419, 59, 1199, 2009, 472, 1040, 184, 2491]. Mariafranca [1116]. Mariam [1810]. Marian [3059, 127]. Marianne [1859, 2846]. Mariano [524]. Marie [3339, 939, 1540, 1193, 1466, 1012, 3278, 1412, 2330, 3386, 2331, 2324, 1572, 2326, 212, 270, 2639, 2717]. Marie-Ann [3278]. Marie-Claire [2324]. Marie-Helene [1540, 2330]. Marie-Therese [2331]. Marika [2036]. Marilyn [2655, 2706]. Marin [255]. Marina [2736]. Marine [3121, 323, 2670]. Mariner [2530]. Marinus [153]. Mario [1169, 1102, 3446]. Marion [2625, 1215]. Maritime [2396]. Marius [740]. Marivaux [2361]. Marja [513]. Marjolein [2805]. Marjorie [3423]. Mark [2343, 734, 3451, 2885, 2800, 187, 1960, 36, 2993, 1668, 414, 2466, 957, 2617, 3133, 873, 2126, 1041, 1355, 2279, 276, 573, 2203]. Markel [3077]. Market [2278, 1868]. Marketing [3092]. Markets [3325]. Markley [2899]. Markovich [654]. Markowitz [3099, 1231]. Markowski [1246]. Marks [3011]. Markt [472]. Markteinfuhrung [2994]. Markus [3149]. Marlein [3239]. Marlowe [733]. Maroto [1723]. Marouf [2611]. Marquet [823]. Marriage [1455]. Mars [2643, 161]. Marsh [2799]. Marsha [2449, 1748, 2987]. Marshall [819, 2329, 314]. Marsilio [2501, 2980, 2980]. Marta [1093, 2941, 1194, 1040]. Martha [1324, 1744, 2369, 470, 2211, 1301]. Martians [3175]. Martin [2695, 482, 1774, 3354, 1099, 1477, 1837, 1242, 3582, 2460, 1947, 1980, 2073, 3542, 903, 3483, 522, 2154, 2117, 2766, 3439, 461, 1701, 2103, 225, 1544, 425, 377, 2549, 2653, 1056, 2809, 1652, 3157, 1315]. Martina [2747]. Martinez [239, 1654, 1097, 752]. Martinez-Alier [239]. Martinus [41]. Martius [1337]. Martzloff [3034]. Marum [41, 41]. Marvin [3576]. Marwood [350]. Marx [2033, 3175, 246]. Marxism [3512]. Mary [486, 1661, 2170, 3269, 574, 333, 826, 3151, 2181, 2049, 1510, 3045, 871, 3032, 2939, 151, 2784, 546, 1202, 857, 1714, 2065, 1939, 1078, 1971]. Maryinez [1508]. Masani [584]. Masao [995, 2536, 3318]. Masar [2426]. Maschinen [174]. Masculinity [2846, 1754]. Mask [2225]. Maskelyne [562]. Mass [3286, 2043]. Mass-Extinction [2043]. Massachusetts [3252]. Massey [110, 1513]. Massimo [157, 2589, 2786, 643, 1925, 2901]. Massonie [3542]. Master [734, 1143, 198]. Masterminded [3009]. Masterpiece [3163]. Masters [3358, 3461]. Mastery [1536, 2564, 1919]. Matalon [1307]. matematica [3129, 2270]. matematiche [2260, 1779, 1722, 521]. Materia [822]. Material [1295, 2418, 489, 21]. materiales [2158]. Materiality [3589]. Materials [3419, 3480, 2522]. Materiaux [3189]. Maternal [1742]. Mathe [3348]. Mathematica [2983, 113, 449]. Mathematical [1891, 2570, 2289, 1461, 3529, 2163, 2199, 2282, 900, 2034, 2196, 2588, 1516, 1114, 1030, 2456, 2907, 164, 268]. Mathematician [1922, 1833, 2431, 3369]. Mathematicians [1928, 3071, 17]. mathematicorum [3241]. Mathematics [233, 1461, 767, 2732, 3027, 900, 548, 2588, 3115, 2497, 234, 1533, 3071, 1420, 515, 3321, 2194, 516, 1485, 3033, 1348, 1736, 3034, 16, 2034, 28, 1770, 347, 3071]. Mathematik [530, 3118, 1500, 2585, 859, 1150, 989, 2254, 2238, 1651, 1303]. Mathematiker [1029]. mathematique [1484, 1486, 2004]. mathematiques [2436, 3189, 1711, 150, 931, 1711]. Mathematisations [2102]. Mathematisch [463]. Mathematisch-physikalische [463]. Mathematische [2979, 2238, 1531, 3137]. mathematischen [3338, 1150]. Mathematischer [1531, 2661]. mathematisches [1403]. Mathematism [2289]. Mathematization [257]. Mathematizations [2896]. Mather [2201, 2770]. Mathews [2740]. Matilde [184]. Matkin [1749, 1749]. Matossian [333, 3032]. Matrix [1132]. Matsui [2537]. Matter [2, 93, 215, 306, 399, 3, 4, 94, 216, 307, 402, 495, 608, 700, 797, 403, 404, 496, 609, 701, 800, 890, 983, 1071, 1156, 801, 891, 984, 1072, 1158, 1160, 1251, 1340, 1430, 1521, 1161, 1252, 1341, 1431, 1523, 1525, 1609, 1692, 1783, 1880, 1526, 1610, 1694, 1785, 1882, 1884, 1965, 2051, 2134, 2218, 1885, 1966, 2052, 2135, 2220, 2222, 2305, 2394, 2477, 2557, 2223, 2306, 2395, 2478, 2559, 2561, 2628, 2708, 2790, 2862, 2562, 2629, 2709, 2791, 2864, 2866, 2948, 3022, 3110, 3193, 2867, 2949, 3023]. Matter [3111, 3195, 3200, 3201, 3304, 3396, 3493, 3592, 3202, 3203, 3305, 3397, 3494, 3594, 1801, 2181, 1624, 2657, 2486, 27]. Matters [53, 3082]. Matteucci [2271]. Matthew [148, 2345, 263, 3258, 3285, 1406, 3465]. Matthews [1091, 2492]. Matthias [3378, 654, 463, 656, 3256, 1393]. Mattson [2778]. Maturing [1574]. Matvei [2124]. Mauel [519]. Maulitz [68, 1141]. Maupertuis [1197, 842, 893, 2307]. Maura [253]. Maureen [363]. Maurice [298, 1017, 948, 1176, 1356, 2446, 2357, 3436, 1624, 1458]. Maurizio [1925]. Maushart [3278]. Max [3367, 3562, 654, 2110, 1221, 684, 908, 2258, 2116, 3391]. Max-Planck-Gesellschaft [3367, 1215]. maximos [2741]. Maxwell [1306, 1487, 2404, 2821]. Maxwellians [1397]. May [707, 1793, 3164, 409, 1394, 898, 1760]. May-June [1394]. Maya [2352, 2648]. Mayaud [3340]. Mayenn [2291]. Mayer [1664]. Maynard [284, 3013]. Mayntz [244]. Mayo [1778, 2999]. Mayr [1169, 2917, 445]. Maze [1865]. Mazlish [1984]. Mazumdar [2764]. Mazur [3472]. Maz’ya [3369]. Mazzei [1660]. Mazzini [3332]. Mazzolini [940]. Mazzone [2270]. McAllester [1065]. McAllister [2857]. McBride [289]. McC [2961]. McCandless [2673]. McCay [1510]. McClellan [2011]. McClelland [3156, 1211]. McClung [2844]. McCluskey [3243]. McCombie [2655]. McConchie [3430]. McConnell [1936]. McCormmach [3065]. McCoy [912]. McCulloch [2128]. McCurdy [3481]. McDermott [3295]. McDonald [3553]. McFarland [2383]. McGaw [2100]. McGee [2729]. McGilvray [913]. McGregor [3265]. McGuinness [2199]. McGuire [1597]. McIntosh [2990]. McKay [225]. McKenzie [3163]. McKie [789]. McKirahan [1547]. McKnight [553]. McLaren [1143]. McLaughlin [1289]. McLean [2597]. McMorris [784]. McMullin [1425]. McNeil [818, 363]. McShane [3300]. McVaugh [2250, 3248]. Meade [2125, 1355]. Meaning [420, 1579, 881, 2840, 3035, 1489, 3231, 3401]. Meanings [2720, 1233, 1636]. Means [2152]. Measure [3453, 2734, 419]. Measurement [814, 1214, 689]. Measures [814]. Measuring [2106, 2287, 3377]. Mebane [733]. mecanica [3352, 3440]. mecanicien [1657]. Mecanique [2906, 1296, 2354, 2900, 3002]. mécaniques [339]. Mecca [1718]. meccanica [2347]. mechanica [2092, 1688]. Mechanicae [2903]. Mechanical [1438, 917, 2261, 2563, 285]. Mechanick [713]. Mechanics [690, 2463, 1686, 2906, 900, 1421, 2379, 531, 2404, 1289, 487, 1216, 586, 1756, 1173]. Mechanik [2984, 360, 842]. mechanische [2507, 2092]. Mechanisierung [2507]. Mechanism [3399, 1207, 1861]. Mechanisms [3326, 3573]. Mechanismus [2977]. Mechanist [7]. Mechanistic [2659]. mechanistische [361]. mechanistischen [3060]. Mechthild [3241, 1916]. Meckel [947]. Medal [501]. Medard [2237]. medecine [3132, 3189, 229, 1023, 1412, 549, 516]. Media [1147, 2130, 3004, 3515]. Mediation [1527]. Medica [504, 2491, 438, 730, 1821, 822]. Medical [2054, 1718, 52, 3043, 1576, 2392, 2848, 633, 2035, 579, 2982, 1886, 353, 2030, 2412, 2602, 1945, 1839, 2470, 996, 2849, 1825, 1577, 1050, 2492, 3495, 785, 572, 2278, 2847, 263, 1376, 1746, 1607, 2414, 936, 3359, 2733, 1629, 3012, 679, 3048, 3430, 3191, 794, 2693, 1926, 2070, 110, 2114, 2674, 2505, 1482, 1848, 2648]. médicale [820]. medicaux [2149, 210, 1994]. Medicin [2979]. Medicina [1690, 523, 638, 1828, 351, 3332, 1354, 1267, 1912, 1654, 1002]. Medicinae [513, 523, 504]. medicinalis [1638, 731]. medicinarum [1637]. Medicinas [1911]. Medicine [30, 1564, 2180, 3019, 1232, 2695, 3249, 856, 3042, 3079, 623, 148, 3464, 3577, 2114, 516, 2429, 993, 2777, 110, 1387, 129, 3078, 1460, 550, 2398, 2654, 509, 1247, 1125, 735, 2504, 1209, 988, 2602, 3151, 2778, 3394, 2423, 996, 1364, 2184, 1824, 1286, 2015, 2935, 2278, 263, 469, 2974, 1566, 1759, 289, 2014, 508, 1907, 1933, 1175, 1190, 3344, 467, 332, 1458, 2366, 2176, 2603, 1848, 2074, 3428, 1636, 3026, 444, 2882, 375, 3490, 3576, 1355, 2279, 611, 2374, 699, 2076, 1126, 438, 1457, 110, 1688, 508, 1229, 2725, 241, 68, 3267, 1457, 1776, 1310]. Medicine [110, 492, 3294, 67, 334, 1035, 1459, 2339, 886, 2362, 424, 3521, 26, 1248, 2029, 1992, 480, 540, 1896, 2250, 1249, 1565, 1725, 397, 315, 1086]. Medicines [3421]. medico [1556, 2893, 183]. medicos [1354, 638, 2435]. Medics [493]. Medieval [30, 29, 723, 3243, 2654, 2504, 905, 344, 1552, 1012, 2738, 1645, 3336, 1096, 2423, 1921, 2072, 2426, 2153, 1976, 619, 1818, 1551, 639, 1370, 32, 1281, 1995, 1778, 1367, 310, 2053, 3401, 1641, 1644, 735, 551, 2073, 2152, 1010]. médiévale [2247, 1093, 1633, 727]. Medii [1637, 2253]. Medina [3475]. Medio [251]. medioevo [2890]. medisch [1733, 761]. Meditationes [3527]. Mediterraneo [2645]. Medizin [1481, 2754, 736, 2110, 1813, 517, 1705, 1266, 2614]. Medizinhistorisches [518]. Medizinische [2382]. Medizinischen [1660, 3367]. medizinisches [1638]. Medizinstudenten [1660]. Medizinzum [2647]. Meek [2706]. Meer [3039]. Meeting [3303, 1614, 221, 2311, 1972, 896, 2627, 1258, 2947, 500]. Meets OncoMouse^{T M} [3084]. Megaterio [2516, 3442]. Megenberg [1996]. Mehlman [1054]. Mehra [586, 2381]. Mehrling [3584]. Mehrtens [859]. Mei [128]. Meiji [465]. Meinel [1841, 1923]. meinem [267, 365]. Meites [62]. Meitner [2531, 1855, 2923]. Meja [2474]. Mel [1133]. Melancholy [3153]. Melanchthon [2427]. Mélanges [642]. Meld [3063]. Melhado [1739, 384, 98]. Meli [1656]. Meliteniote [2581]. Melnick [362]. Melodie [1947]. Meloni [2963]. Melosi [1947]. Melville [2996]. Melvin [2401]. Members [2746]. Memoir [2678, 3281, 967, 1991, 379, 3459, 2852]. memoire [3020, 2188]. Memoires [2485]. Memoirs [3467, 3564, 2486, 225, 202, 2981]. Memoranda [2542]. memoria [1195]. memorie [2756]. Memories [2995]. Memory [248, 2480, 2399, 3324, 144, 2025, 2143]. Men [2301, 1441, 1697, 2939, 419, 1940]. Menace [1949, 2226]. Menachem [1488, 1305]. Menagé [1037]. Menato [1284]. Mendel [2909, 278, 2111, 2140]. Mendeleev [2479]. Mendelian [2294, 1334, 774, 3308]. mendeliana [853]. Mendelsohn [118]. Mendenhall [465, 465]. Méndez [752]. Menendez [3444]. Menger [2199, 2868]. Menlo [2369]. Menninger [867]. Menso [2973, 1778, 518]. Mental [2801, 2115, 1605, 855, 2779, 3029, 3379, 1148, 2470, 595, 1648, 1708, 3579]. Mentalities [999]. Mentally [2699]. mente [1107]. Menzies [2972]. Merce [1261]. Mercedes [2825, 2090]. Merchant [624]. Merchants [2163]. Mercier [2227, 3524, 2425]. mercuriales [3444]. Mergler [1872]. Meridiankreises [3253]. Meriwether [2955]. Merleau [522]. Merleau-Ponty [522]. Merricks [2767]. Merriley [1139, 3311]. Merrill [2778, 369]. Mersenne [255, 10]. Mersits [1046]. Merton [928, 387, 2940]. Merzbach [767]. Mesa [2090]. Mesmer [3417]. Message [2826, 3439]. messager [1728, 1653]. Messenger [831]. Messing [3256]. Messingerzeugung [3256]. Mesue [2503]. Mesure [2820, 1931]. Meta [724, 1722]. Meta-Cosmic [724]. Metafisica [2526]. metafori [860]. metafysica [2092]. Metal [3480]. Metals [25, 3502]. Metaphor [1851, 1033]. Metaphors [2608, 3024, 2118, 778, 543]. Metaphysical [616, 1287, 782, 1379]. Metaphysics [709, 559, 1269, 3222, 3064, 1531, 2085, 2497]. Metaphysik [851, 2585]. Metaphysische [394]. Metascience [506]. Metatheoretical [3177]. Metaxas [2776]. Metchnikoff [1033]. Meteorite [2668]. Meteorites [3261]. Meteorology [3094, 942]. Meteors [3343]. Meterology [193]. Metheora [343]. Method [3255, 3061, 1011, 1499, 1192, 1693, 2497, 2267, 311, 1271, 1617, 905, 36, 1602, 3188]. methode [1874, 1916, 2363]. méthodes [931, 2354]. Methodological [488, 1335]. Methodology [605, 1335, 3391, 2941]. Methods [241, 833, 610, 1075, 3040, 1717, 1756]. Methodus [39, 160]. Metiers [2132]. Metraux [2420]. Metropolis [2091, 816]. Metropolitan [3163, 3209]. metropolizacion [2255]. Mettrie [1387, 2664, 2352]. Metzger [1172, 529]. Mexico [2982, 2648, 3483]. Meyenn [2834]. Meyer [2032, 2604, 345, 1069, 2116, 779, 1235]. Meyer-Palmedo [2032, 2604, 1069]. Meyers [3410]. Meyerstein [1237]. Mgr. [3460]. Miasmas [1732]. Micale [2800]. Mich [2733, 3278]. Michael [2726, 3104, 2552, 2343, 3049, 2647, 2962, 1018, 870, 736, 568, 2521, 37, 180, 1424, 3562, 1485, 1536, 1321, 1567, 1605, 2582, 647, 1913, 2662, 2250, 2553, 2898, 1269, 1745, 2501, 2412, 710, 2805, 2235, 3336, 2281, 2769, 2977, 2916, 2144, 2034, 1906, 1378, 2107, 3233, 1938, 2663, 3376, 3387, 419, 1852, 397, 3248, 537, 1944, 2990, 2603, 3541, 543, 2037, 1545, 2908, 966, 3526, 3076, 135, 490, 3323, 2171, 116, 1544, 1179, 3471, 3250, 472, 3459, 2005, 557, 677, 2204, 2378, 1172, 1278, 1532]. Michalski [1811]. Michel [1308, 2463, 1296, 2508, 935, 2432, 2004, 2290, 1178, 237, 3002, 522, 3412, 2028, 2762, 3465, 694, 485]. Michel-Pierre [2432]. Michela [552, 1915]. Michele [3073, 2275, 1003, 1502, 1473]. Micheli [2063, 2241]. Michigan [1848, 2733, 2733]. Michio [2426, 3432]. Microbes [1441]. Microfiche [1795]. Microfilm [1291, 1378]. Microscope [601, 2745, 2263]. Microscopical [601]. Microscopy [3059]. Microtome [3399]. Middelfort [1371]. Middle [148, 726, 2806, 1726, 1280, 826, 1012, 2868, 1632, 3123, 1719, 2152, 1812, 2807, 1636, 3335, 732, 2138, 807]. Midgette [1052]. Midst [2699]. Midwest [336, 3218]. Midwife [3346, 1301]. Midwifery [3152]. Mieczyslaw [1246]. Mieleitner [358]. Mieth [3053]. Migne [1571]. Migration [1034]. Mihai [692]. Mikael [2105]. Mike [2112, 3264]. Mikel [3445]. Mikhailovna [1810]. Mikhajlov [3534]. Mikroelektronik [472]. Mikulas [1706, 1727, 169, 420]. Milan [3104]. milanese [2355]. Milchstrasse [2411]. Mildner [1660]. Mildner-Mazzei [1660]. Miles [2192]. Milestones [3127, 1415, 63]. milieu [322, 1307]. milieux [2166]. militares [262]. Military [1954, 1644, 968, 2851, 3559, 2422, 968]. Military-Industrial-Academic [1954]. Milky [2119, 1853]. Millar [3487, 3487]. Millard [198, 763]. Millennium [1299]. Miller [2622, 2965, 2665, 1950, 3408, 2622, 3348, 3582, 225, 1426]. Millikan [1218]. Mills [3336, 589]. Millstones [1415]. Milos [3341]. Milton [332, 2657]. minas [3444]. Mind [1662, 2287, 2115, 1536, 1579, 1771, 446, 3024, 664, 2187, 2959, 534, 1514, 3469, 2344, 3485, 1456, 543, 665, 2128, 2374, 2480, 1684, 3438, 3532, 1948, 2286]. Minded [2699]. Minden [2035]. Minds [905, 2328, 3377, 2786, 1979]. Mineheads [3179]. Miner [795]. Minerale [358]. Minerva [262, 2439, 3439]. mines [3147, 3336, 827]. mineur [827]. miniature [530]. Minick [1673]. minime [255]. minimis [2183]. Mining [1495, 2939]. Ministerium [1572]. Ministers [1649]. Minneapolis [2311]. Mino [3434].

Minor [1816, 1373]. Minton [596]. mirabilis [3129]. miracle [3509]. Miracles [309]. Miranda [2069]. mirandis [2587]. Mirella [3457]. Mirhady [1906]. Miriam [1878, 2524]. Mirko [1023, 505, 1141]. Mirowski [539, 2121]. Mirror [1557]. Mirrors [219]. Misa [2459]. Misbehavior [1619]. Miscellanea [2983]. Misconduct [1511]. Miscreations [2904]. Misia [907]. Misreading [1536]. Missile [1055, 2129]. Missiles [2422]. Missing [368]. Mission [1670, 1499]. Missions [1345, 583]. Mistakes [2786]. Misurare [1365]. Misuses [3449]. Mitcham [968, 2060]. Mitchell [3230, 302]. Mitglied [1830]. Mithraic [544]. Mitman [1591]. mito [3181, 3545, 2656]. Mitscherlich [1398]. Mitteilungen [1941]. Mittelalter [1635, 1274, 1993, 1993]. mittelalterliche [2503, 1360]. mittelalterlichen [2647, 345]. Mittelalters [2332, 2424, 2070]. Mixed [2061]. Mobile [272, 1589]. Möbius [3112]. Mocchi [1107]. modalita [3054]. Mode [616]. Model [3562, 3264]. modèle [1201]. modeles [1237, 1042]. Modeling [3399, 3016]. Modell [1500, 1304]. modelli [973]. Modells [2336]. Models [1605, 1404, 2068, 2063]. moderado [752]. Modern [1413, 2086, 2054, 3249, 2468, 1325, 1324, 3043, 2700, 1132, 928, 3116, 1649, 2429, 2806, 579, 616, 1080, 1536, 3522, 775, 2163, 1730, 413, 1491, 2926, 3318, 446, 3087, 193, 1148, 1169, 1005, 2959, 2042, 1921, 1087, 1224, 3461, 1775, 2840, 1621, 3045, 2538, 1319, 1798, 3015, 3119, 3328, 1386, 1971, 478, 445, 1282, 1376, 3507, 69, 1647, 556, 2710, 2459, 136, 1141, 2002, 1754, 774, 1083, 692, 2475, 2451, 1260, 2263, 2298, 1895, 2513, 1289, 2696, 1211, 954, 732, 3114, 2194, 703, 1796, 2235, 1154]. moderna [122, 811, 1471]. moderne [1753, 2188, 451, 2031, 859]. modernen [2647, 2920, 1292]. Modernism [3168]. Modernist [484]. Modernity [870, 3356, 3156, 832, 719, 553, 2142]. Modernization [1947, 944]. Moderns [2997]. Modest Witness [3084]. moet [2693]. Moewus [864]. Mogens [3163]. Mohammad [3242, 533]. Moheit [1829]. Mohr [654, 1577]. moitie [2812, 1390]. Moked [457]. Molds [333]. moléculaire [1412]. Molecular [1306, 1221, 864, 1593, 2208, 379, 3465, 2781]. Molecularizing [3464]. Molecules [2404, 698]. Molero [2090]. Molland [437]. Møller [515]. Molloy [3356]. Molnar [1321]. Molyneux [2805]. Moment [2876]. Moments [1886, 865]. Monad [2916]. Monadology [1110]. Monaghan [2111]. Monarchs [1649]. Monastery [1364]. Monastic [5]. Moncada [262]. monde [2743, 718, 1004, 2004, 2960]. mondes [2895]. mondiale [771]. mondo [1471, 1722]. Monetary [3584]. Money [579, 3027, 3584, 384]. Mönnich [736]. Monoclonal [3093]. Monograph [2262]. Monopoly [293]. Monsters [2904, 1784]. Monstres [1259]. Monstrous [1726, 1540]. Montag [686]. Montanari [3139]. Montellano [993]. Montes [2090]. Montgomery [2489, 1979]. Monti [1199]. Montiel [1828]. Montmartre [3147]. Monuments [935]. Monzani [522]. Moon [1956, 1015, 3009]. Moonlight [2673]. Moore [1169, 3019, 2392, 2030, 661, 794, 2282, 950]. Mora [1371]. Moral [3222, 296, 356, 2014]. Morality [3087]. Morals [2328]. Moran [726, 1191, 1286, 1190]. Morange [3465]. Morawski [66]. Moray [2521]. morborum [2587]. morboso [3444]. Mordechai [741]. Moreau [842].

Morelli [1169]. Morfologia [1826]. Morgan [2841, 2116, 2841, 857]. Morin [3190]. Moritz [1190]. Mormino [2347]. Mormon [1542]. Moros [1811]. Morphological [1489]. Morphology [467, 2364, 2171]. Morpurgo [1183]. Morrell [3368]. Morris [213, 16, 396, 781, 2577]. Morris-Suzuki [2577]. Morrison [697, 713]. Morrison-Low [697, 713]. Morse [849]. Mortality [947, 1742]. morte [1134]. Mortlake [2631]. Morton [3019, 2392, 2276, 1710, 794]. Morveau [2275]. Moscow [1217]. Moscucci [765]. Moses [1184]. Moshe [1428]. Mosimann [2154]. Mossing [3256]. Mossing-Werkh [3256]. Most [2958, 1743]. Mothers [3428, 2233]. motibus [1102]. Motion [2695, 622, 3050, 1192, 27, 1289, 2065]. Motions [3499]. Motivation [879]. moto [1737]. Motoo [2208]. mots [388]. Mott [3565, 1627]. Motu [2271, 1386, 1473]. motus [2347]. Motz [621, 225]. Moulin [1412]. Mouloud [1063]. Mount [696]. Mountain [963]. Mountains [2989]. mouvement [1296, 2900]. Movement [3380, 595, 2296, 225, 596, 844, 1499, 2295]. Mover [1789]. Movers [3244]. Mowery [872]. Moyen [3189, 2247, 210, 825]. Moyenne [1307]. Moyer [1617, 3351]. Mozley [3558]. MRC [679]. MRI [2778]. MS [566, 1473]. Mucchielli [3272, 3556]. Mucha [65]. Mueller [2902]. Mueller-Sievers [2902]. Mugnai [1556]. Muhammad [1631]. Muir [3282]. Mujer [1690]. mujeres [1690]. Mukherjee [508]. MULAPIN [247]. mulierum [1715]. Mulkay [1424, 3376]. Mullan [594]. Muller [2884, 3551, 2413, 654, 1481, 2988, 648, 1113]. Muller-Benedict [2413]. Muller-Jahncke [2884]. Mulligan [303]. Mulliken [379]. Multhauf [3061]. Multicultural [2410]. Multiple [1244, 2143, 1145]. Multitude [2034]. Mumford [2617, 1411]. Munchen [1941]. mundi [435]. Mundializacion [1735]. mundo [1911, 2803, 2741]. Mundus [740]. Mundy [2813]. Munevar [1241]. Munk [642]. Munster [1351]. Munsters [2336]. Murad [203]. Murchison [2521, 847, 2521]. Murders [856]. Murdoch [1111, 190]. Murphy [2939, 1126, 1680]. Murr [1559, 1559]. Murray [961, 1404, 2856, 2850]. Murray/Mabahiss [961]. Murs [1369]. Musa [1005, 2877, 1631]. Musallam [510]. Musee [1814, 1129]. Museo [3104]. muséologie [1129]. Museum [2453, 87, 3537, 3163, 1432, 1941, 21, 1458, 1590, 697, 844, 3158, 1202]. Museums [2548, 2086, 179, 1941]. Musgrave [1961]. Music [2082, 1553]. Muslim [2216]. Musschenbroek [3140]. Must [2693]. Mutation [612, 653]. Mutations [3308]. Mutazili [2246]. Mutis [2272, 2517]. Mutual [2693]. Mutuality [3497]. My [803, 182, 2995, 1134]. Myers [2367, 1352, 808]. Myesiers [1914]. Myllyntaus [1765]. Myrdal [1509]. Mysteries [544, 1710]. Mystery [3346, 734, 1534, 1622, 2887, 2668]. Mystic [1285]. Mysticism [2692, 1861]. Myth [1177, 991, 2959, 3485, 1823, 2069, 367, 479, 631]. mythe [1356]. Myths [3449].

N [1008, 1457, 810, 1212, 15, 2906, 1453, 1953, 1899, 3420, 1148, 1169, 386, 297, 1957, 2207, 952, 2655, 1415, 144, 162, 2499, 1309, 2543, 1083, 1669, 2469, 682, 1315, 697, 3191, 1455, 1137, 822]. N.S [825]. NACA [583]. nach [2973, 851, 549, 712, 1150, 1122, 1766, 2535, 2253, 755]. Nachdenken [1381]. Nachlass [2123]. Nachlasse [1337]. Nachum [3117]. Naddaf [1713]. Naddaff [237]. Nadia [237]. Nadler [1730, 1476, 737]. Nagasaki [2925]. Nagatomo [980]. Nagel [2008, 2300]. Nahum [1121]. Naissance [1671, 3272, 1026, 1296, 451, 718, 900]. naïve [1982]. Najm [3525]. Nakayama [1955]. Naked [2848]. Named [1248]. Names [3491, 2071]. Naming [1370, 497, 3469]. Nancy [2723, 735, 1872, 2470, 1687, 962, 383, 1326, 689, 2765, 1799, 1052]. Nantes [653]. Naomi [3454, 1507]. Naoyuki [2208]. Napier [1104]. Naples [3115]. Napoleonic [935]. napoléonienne [270]. napoletano [2260]. Napoli [925]. Nara [3181]. Narayne [59]. Narratio [1728]. Narrative [3073, 123, 3104, 2859, 792]. Narratives [2996, 907]. NASA [583, 3176, 304, 682, 1863, 2387, 2388, 304]. Nash [626, 1114, 381]. Nasim [2216]. Nasir [1991]. Nasr [1805, 1717]. Natalie [970, 2349]. Nath [2599, 2413]. Nathan [2321, 129, 2193, 1204, 414, 872]. Nation [1326, 2459, 47]. National [2684, 2214, 1951, 1666, 3370, 492, 2536, 1727, 3173, 1415, 2670, 2132, 1695, 21, 3095, 873, 2606, 581, 2039, 2683, 3171, 169, 676, 1967, 1735, 2851]. Nationale [2633, 3538]. Nationalism [2113, 1676]. Nationalismus [3552]. Nationalized [1234]. Nationalsozialismus [471, 578, 286, 2059]. nationalsozialistischen [2923]. Nations [1551]. Native [3414, 573, 383, 1495, 1294]. Natur [1946, 2107, 1283, 297, 1403]. natura [3054, 2526, 1183, 1826, 1471, 1643]. naturae [2095]. Natural [2453, 2548, 2989, 2681, 2427, 1548, 2549, 951, 447, 3050, 1000, 417, 1396, 2442, 1832, 2967, 2525, 2441, 2571, 1479, 2264, 96, 905, 2507, 688, 2631, 3213, 3517, 489, 2017, 1130, 976, 1514, 507, 1239, 3313, 179, 2055, 3316, 2886, 77, 2006, 2345, 1432, 2735, 1755, 355, 2002, 3229, 1370, 1652, 1067, 2319, 1590, 369, 1203, 119, 2022, 937, 1942, 265, 1561, 732, 3512, 1627]. Naturale [3104, 2443]. naturales [1828]. naturaleza [2157, 1095, 2103]. naturalis [2249]. Naturalism [454, 2702]. Naturalist [2170, 2455, 3548, 155, 265, 1483, 2407]. Naturalists [3108, 2987, 2367, 1352]. naturaliter [358]. Naturalium [3244, 727]. Nature [539, 2086, 3249, 3284, 1806, 78, 2429, 1536, 309, 2094, 316, 3326, 1871, 2163, 2441, 3210, 167, 1591, 3511, 1112, 1921, 3580, 952, 3403, 1423, 3016, 349, 3035, 3051, 478, 1432, 1665, 1592, 2772, 434, 7, 3225, 560, 2819, 1717, 1202, 1404, 2564, 1919, 1435, 3095, 43, 2174, 1714, 3264, 624, 2758, 3404, 452, 381, 1905, 2348, 574, 970, 2898, 426, 924, 1538, 2886, 257, 2665, 3265, 1799, 955, 1546, 1761, 2064, 1755, 759, 722, 689, 1944, 166, 131, 2513, 1493]. Nature-Nurture [3580]. Naturelle [1201, 1026, 1790, 1390, 2567]. Natures [2071, 784]. Naturfilosofi [1130]. Naturforscher [12, 2013, 1930, 1634]. Naturforschern [2987]. Naturforschung [3060]. Naturgeschichte [163]. Naturhistorischen [1941]. Naturlehre [2154]. naturlichen [1996].

Naturmagie [1283]. Naturphilosophie [1185, 656]. naturphilosophischen [2356]. Naturvorstellungen [1946]. Naturwissenschaft [656, 2095, 1813, 1283]. Naturwissenschaften [517, 2620, 341, 3552, 530, 324]. Naturwissenschaftler [2747, 2988, 1766]. naturwissenschaftlicher [2661, 2535, 1531]. Natuur [165]. natuurkunde [2292]. natuuronderzoek [2101]. natuurwetenschap [361, 1297]. Natuurwetenschappen [516, 994]. Nauchnaia [1212]. nauchnyck [1336]. Nauchnye [887]. Nauki [512]. nautique [146]. Navajo [3577]. Naval [874, 3463, 3445, 3345]. Navarro [2490, 829, 1354, 3444]. nave [2811]. Navegacion [2645]. navegar [152]. Navy [874, 1597]. Nazi [1857, 1951, 1678, 3563]. Nazionale [327]. Nazis [3013]. Nazism [676]. Neal [428]. Neale [3331]. Near [2931, 1987, 32]. Nebel [2411]. Nebulae [2530]. Nebulous [3405]. nec [1134]. Necah [581]. Necessity [2261, 3106, 1152]. Nederland [2680, 2975, 1535, 658]. Need [1805]. Needell [583]. Needham [2312, 329, 2422, 2972, 1133, 3319, 1706]. Néel [1043]. negentiende [228, 658]. Negro [3585]. Neher [35]. nei [638]. Neighbors [703]. Neighbouring [1392]. Neil [2841]. nel [2890, 2590, 645, 2355, 1737, 2586, 2270, 251, 1643, 2259, 2590]. Nélia [1129]. Nelkin [590]. nella [3054, 3545, 853, 2277, 564, 1183, 1199, 1925, 1471, 1821, 1722]. nell’astronomia [1102]. Nellen [2166]. nell’Encyclopédie [1116]. Nelson [2609, 883]. Nemet [598]. Nemore [1096]. Nenuzhnoe [1014]. Nepomuck [3441]. Nerves [1036, 2996]. Nervous [1760]. n’est [2028]. Net [2454]. Netherlands [835, 2547, 2452]. Nets [3372]. Netton [1630]. Networked [54]. Networks [3372]. neu [1304, 3017, 453]. Neu-eröffnete [453]. Neubeginn [712]. neuchey [1014]. neue [297, 998]. neuen [2465]. Neuentdeckung [3273]. Neuer [2067]. neues [2994]. Neufeld [3459]. Neugebauer [1462, 409]. Neuman [304]. Neumann [1319]. Neun [530]. Neural [3372, 752]. Neurasthenia [52]. Neurath [1332, 2765, 789]. Neurociencias [752]. Neurologist [1578]. Neurology [2762]. Neuron [1320]. Neuroscience [1898, 3324]. Neurosciences [2573]. Neuroses [288]. Neuse [3365]. Neustrukturierung [3586]. Neutral [2208]. Neutrons [3085]. Neuve [3460]. Neuzeit [1292, 1923]. neuzeitlichen [1927]. neuzeitlicher [297]. Neve [2412]. Never [1796]. Nevil [562]. Nevill [3565]. Neville [1310, 784, 2414]. Newberry [1702, 1566, 599]. Newborn [2775]. Newcastle [3497]. Newcomb [1617]. Newman [2860, 2169, 1465, 3126]. Newsholme [3078]. Newton [1656, 851, 37, 838, 1731, 905, 1112, 1929, 2266, 2267, 3530, 1704, 1380, 3143, 741, 1291, 839, 2092, 838, 1731, 1789, 1257, 360, 1787, 558, 2350, 1478, 739, 2065, 1657, 38, 2594, 1290, 2268, 2487]. Newtonian [2899, 2269, 1479, 742, 932, 2227]. Newtonianism [2268]. Next [3255]. Niall [2937]. Niccoli [922]. Niccolo [932]. Niche [2527]. Nicholas [754, 447, 1373, 3358, 2172, 1816, 1476, 1058, 3517, 3136, 569, 2080, 109, 747, 2972, 1110, 1022, 569, 1563]. Nichols [225]. Nicholsen [489]. nicht [549, 3278]. Nicolaas [576]. Nicolas [2897, 1032, 1165, 921, 2897].

Nicolas-Claude [1165]. Nicolaus [2506, 358, 3246]. Nicole [561, 900, 273, 657, 2892]. Nicoletta [3268]. Nicolini [3268]. Nicolo [1556]. Nicolosi [450]. niedergeschrieben [2447]. Niederlande [160, 2509]. Niedersen [654]. Niels [57, 78, 2510, 860, 2377, 1804, 2816, 2118, 949, 1303, 3276, 190, 1216, 1408, 3564]. Nielsen [3467]. niemandem [2978]. Nier [1947, 2369]. Nietzsche [1775, 1775]. Nieuwentijt [1297]. Nieves [2740]. Nigel [879, 1458]. Night [371, 1411]. Nightingale [182]. nihil [554]. Niko [2680]. Nikola [3271]. Nikolai [3560]. Nikolas [2883]. Nikolaus [2853, 3534, 1916]. Nimmen [304]. Nina [3346]. Nineteenth [3361, 2180, 1576, 2829, 3267, 3160, 2104, 602, 2578, 2996, 1205, 2156, 574, 2525, 2779, 1491, 1034, 3158, 941, 2602, 1945, 181, 1839, 3260, 3557, 179, 1499, 1208, 1577, 1249, 3072, 371, 2109, 1665, 1755, 2105, 1037, 760, 2823, 2025, 2819, 2113, 3359, 2986, 1391, 1211, 1206, 946, 2279, 1492, 1532, 462, 173, 1942, 3070, 1786, 3399, 217, 3113, 3112, 611, 3114]. Nineteenth- [3260]. Nineteenth-Century [1576, 2829, 3267, 3160, 2104, 2996, 1205, 2156, 574, 2525, 1491, 941, 1945, 1839, 3557, 1499, 1208, 1577, 3072, 2109, 1665, 1037, 760, 2823, 2113, 3359, 2986, 1391, 1206, 2279, 1942, 1786, 3399, 217, 3113, 3112, 611, 3114]. Nineveh [3163]. nisi [1134]. Nissen [1987]. Nixon [1768]. NMR [2778]. No [446, 583, 2536, 3085, 3216, 902, 2510, 1699]. NOAA [2670]. Noah [1645, 3333, 631]. Noam [3086]. Nobel [1958, 1676, 1777, 2618]. Nobelpreistrager [2122]. Nobis [442]. Noble [126, 1799, 1273]. Noel [1324, 3520, 3340, 1744, 1819, 1063]. Noëlle [270]. noi [910]. Nola [295]. Noll [2295, 2699]. Nollet [460]. Nomanul [2071]. Nomanul-Haq [2071]. nombre [2911]. Nombreux [161]. Nomina [1637]. Nomograms [3224]. Non [367, 1776, 3394, 1348, 2326, 149]. Non-Darwinian [367]. non-euclidiennes [149]. Non-European [1348]. Non-Western [1776, 3394]. Nonconservation [188]. nonnullis [2587]. Nonstandard [2196]. Noordwijkerhout [160]. Nora [1395, 1245]. Norbert [1681, 2465, 584]. Nordic [2062]. Nordon [1004]. Noreña [637]. Norimbergae [2901]. Normal [331]. Norman [3479, 111, 1688, 1957, 3176, 839, 833, 1688, 2974, 2074]. Normative [937]. Norrell [858]. Norri [3491]. Norris [1673]. Norriss [2768, 24, 1621]. North [2468, 948, 3326, 905, 3348, 53, 1541, 766, 1750, 245, 2802, 337, 2136, 252, 2753]. Northern [3448, 2738, 1508]. Northington [2296]. Northside [3099]. Norton [846, 2574, 2117, 2488, 2638]. Norway [3013, 3044]. Norwood [1538]. nos [212]. Nose [3467]. Nostradamus [2079, 2079]. Nostrand [3092]. Notable [2706, 2177, 2943]. Note [2872, 2140]. Notebooks [2676]. Notes [2233, 465, 452, 3335, 666]. Nothing [554, 299]. noticias [2158, 2272]. Notion [2898, 2911]. Notker [1185]. Nott [272]. Notturno [3485, 3485]. nouveau [900]. ’nouvelle [2348, 2906, 349]. nouvelliste [2166]. Nova [160]. Novant’anni [3001]. Novel [3297]. November [1614, 3252, 2626, 2634, 3312, 2401, 896, 2627, 2185, 2947, 898, 312, 3402]. novembre [3460]. Novick [126, 188]. Novikov [2120]. Novum [2164].

Nowotny [3566]. NS [1142, 1222]. NS-Biologie [1222]. NS-Pharmazie [1142]. NTM [517]. nucléaire [981]. Nuclear [1055, 1617, 3558, 2205, 3364, 3166, 1502, 2607, 2213, 225, 873, 2686, 2878, 949, 1329, 1867, 225, 2855]. Nucleus [56]. Nueva [2175, 2158, 3154, 2444]. Nuevo [2517, 1911, 1654]. Nuevos [2158]. Nugayev [1336]. nuit [1124]. Nuland [334]. Number [2501, 1096, 1030, 617, 1985]. Numbers [2406, 1764, 1192, 2609, 3143, 375, 1892, 2138, 1595, 106, 3550, 1622, 186, 901]. Numerical [610]. numerorum [1369]. nuncio [1728]. nuncius [831, 1653, 521]. Nunez [3154, 2517]. Nunn [3521]. nuova [860, 454]. Nuove [3277]. Nuovi [2259]. Nuovo [3001]. Nuptial [2501]. Nuremberg [34]. Nurse [3292]. Nursery [2775]. Nurses [3191]. Nursing [70, 751, 3554, 3357]. Nurture [3580, 53]. Nutrition [2148, 993]. Nuts [274]. Nutton [110, 2412, 540]. Nye [3269, 2370, 3424, 2181, 1754, 1233]. Nyhart [2364]. Nylon [3285].

O [76, 1362, 509, 1418, 1462, 725, 260, 2407, 2029, 253]. O.P [2077]. Oak [2039]. obecnych [1655]. Oberdiek [2784]. Objectivity [2406, 1773, 126, 685, 3435]. Objects [3214, 3187]. Objekte [2095]. oblio [1195]. obnoxiae [2183]. O’Boyle [1641]. O’Brian [1740]. Obrist [1094]. Obscurity [2371]. Observation [881, 991, 387, 2643, 1728]. observationibus [39]. Observations [2068, 3074, 572]. observationum [2183]. observatis [1728]. Observatories [600, 3173]. observatorio [563]. Observatory [2684, 2931, 20, 906, 1418]. Observed [1406]. Obsession [862]. Obstetrics [2144, 1248]. Obtained [803]. Ocaña [523]. Occasionalism [1730]. Occident [3509]. occidental [931]. Occult [733, 1190, 355]. occulta [1918]. occulte [1101]. Occupational [1231, 2776]. Ocean [961]. Oceanography [589, 2396]. O’Connell [2991]. O’Connor [3512]. Octavo [1666]. October [3303, 221, 1792, 101, 2311, 1889, 1972, 2542, 2954, 896, 312, 706, 500, 1791]. od [1655]. Odd [3247]. O’Dowd [2144]. Odyssey [60, 959, 2196]. Oefening [3068]. Oelschlaeger [908]. Oestmann [2336]. oeuvre [687, 2079, 2188]. Oeuvres [2338, 2895, 3527, 2338, 3146, 2895]. Oeuvres. [1476]. Off [74]. Office [874, 2641, 872, 225]. Officers [3219]. Official [968]. officiels [2182]. Offray [2664, 2352]. oggi [925]. Ogilvie [2706]. O’Grady [2793]. Ogren [2455]. Oh [1680]. Oil [2386, 2053, 3483]. oiseau [1124]. Oivind [3044]. Oka [3373]. okkulten [1640]. Ökologie [812]. Olaf [1626, 656, 515]. Olaus [2007]. Olav [3178, 1563]. Olbers [2016]. Old [1536, 2698, 2011, 1384, 2990, 945, 375]. Oldroyd [2798, 941]. Oldstone [3233]. Ole [3543, 1725, 515]. Olesko [1210]. Olga [392]. Olive [2053]. Oliver [2371]. Olivia [1276]. Olivier [1764, 2495]. Olmi [1471]. Olson [3343, 1734, 1080]. Olympian [1475]. Olympus [2728]. om [2428]. Omar [262]. O’Meara [726, 548]. Omens [531]. Omnia [3534, 438]. omnium [730]. omocentriche [1102]. Oncology [1087]. onder [2292]. Onderlinge [2693]. onderwijs [3068]. One [1324, 68, 3009, 3470, 2711, 3285, 2029, 1530, 1169]. Onlookers [970]. Only [1693]. Onnes [1214]. ons [2693]. Ontology [1800, 1335, 1669]. ontstaan [2680]. Oosterzee [3159]. Open [2719]. Opening [3085]. Opera [1920, 438, 2726, 3534]. Operation [1679, 2040]. Opere [919]. operis [438]. Ophthalmology [2885]. Ophuijsen [3239]. Opiate [1244, 590]. Opinion [987]. Opinions [3426]. opkomst [761]. Oppenheim [1036]. Oppenheimer [2625, 2125, 672, 1747]. Oppenraaij [1913]. Opposed [1922]. Opsomer [916]. Opstall [996]. Optica [165]. Optical [2594, 903, 173]. Opticks [1929]. Optics [905, 728, 2564, 3133, 2151, 2174]. Optimism [2214]. Opting [2386]. optique [821]. Opvattingen [165]. Oracles [225]. O’Raifeartaigh [2919]. Oral [3372, 2540, 1087]. Orange [3096]. Orbit [8]. Orbital [379]. Orbs [1976]. Order [2348, 1177, 1536, 855, 2595, 3314, 2055, 3227, 945, 3404, 819, 2554, 2021]. Order/Mental [855]. Ordered [70]. Ordering [1895]. Orderly [810]. Orders [583]. Ordinary [2552, 309, 1699, 2088]. Ordnen [2095]. Ordóñez [524]. Oregon [2668]. Orel [2909]. Oresme [2892]. Oresmes [3246]. Orfila [183]. Orford [2668]. Organic [997, 1751, 1134, 177, 168, 2025]. Organisationen [2013]. organischen [2013]. Organism [2309, 2471]. organismische [297]. Organization [3206, 2785, 1057, 406]. Organizational [1162]. Organizations [1211]. Organizing [2685]. Organon [512]. organum [2164]. Oriani [2355]. Oriental [3140, 301]. orientierten [3375]. Orígen [868, 1272]. origenes [3475, 1024]. Origin [2385, 3405, 3166, 1541, 1141, 1669, 3046, 3187, 1064]. Original [1442]. Origination [2794]. origine [1982]. origines [1201, 1004, 2188]. origini [2526, 2589, 2241, 1194]. Origins [627, 1008, 2913, 1077, 544, 1708, 3116, 3356, 866, 3448, 159, 1221, 446, 3380, 3024, 1400, 2959, 133, 2063, 1752, 864, 553, 1319, 2997, 1033, 1405, 2328, 619, 3377, 2239, 1969, 2775, 2295, 1203, 3335, 1179, 2374, 1898, 2849]. Orleans [1972]. Ormiston [792]. Ormsby [1184]. Ornella [765]. Ornithology [3451]. Ornstein [2292]. Oro [1723]. oroscopo [1101]. Orosz [844]. Orphique [146]. Orsted [3543, 3543]. Ortega [1735]. Ortelius [1998]. Orthodoxy [454]. Orthopaedics [1087]. Ortiz [993, 1690]. Osborn [1590]. Osborne [1944]. Oscar [287, 2024]. Osietzki [472]. Osiris [107]. Oskar [3375]. Osler [2261, 1383, 3165]. Osterbrock [2931, 1763]. Österreich [475]. österreichisch [755]. österreichisch-deutscher [755]. österreichischen [475]. Ostuni [2155]. Ostwald [3080, 772, 174]. Ostwalds [654]. Oswald [3027]. Oswaldo [2027, 3181]. Oswaldus [3531]. Ot [910]. Otaola [2915]. Otero [3446]. Other [632, 2463, 3014, 2273, 2017, 2804, 962, 686, 151, 2664, 2164, 566, 953, 1531, 1850]. Otherwise [1784]. Otherwyse [1248]. Othmar [229]. Otis [2025]. Ott [3162]. Ottavia [922]. Otte [343]. ottica [2744]. Otto [789, 3447, 1317, 1332, 1200, 1924, 1762, 995, 2780, 1134, 1415, 445, 756, 1855, 1651, 409, 62, 2765, 3570]. Ottocento [3545]. Our [3299, 2376, 990, 2786, 3510, 877, 13, 125, 2882, 2699, 2693, 1696, 1143, 3288, 782]. Ourselves [2883, 2296]. Ousia [1269]. Outer [1604, 2130]. Outline [164]. Outram [652]. outros [2027]. ouverte [657]. Ouvrage [3415]. Ovary [3416]. Ove [1248, 699]. Ovelton [886]. Overfield [1574]. Overmier [1519]. Overseas [1670]. Overview [288, 3426, 3384]. Owen [1894, 2649, 448, 1139, 1394, 764, 1254]. Own [1143, 1983, 877, 13, 2711]. Owsei [1180]. Oxford [87, 29, 3368, 139, 3136, 3136]. Ozhigova [887].

P [244, 1484, 1999, 629, 2075, 2833, 570, 1702, 887, 1212, 3242, 2532, 1000, 3108, 1460, 2355, 1871, 1983, 2963, 64, 724, 584, 2660, 1808, 817, 2197, 1462, 1217, 346, 2061, 3176, 3056, 2357, 2420, 832, 1928, 2266, 1197, 912, 1378, 2935, 2676, 493, 2775, 1202, 1236, 166, 2720, 2244, 2733, 2907, 473, 2688, 255, 1487, 1064, 2852, 1170, 1777]. P. [2532]. Paarlberg [137]. Pablo [296]. Pabst [2332]. Pace [1722]. Pacey [911]. Pacific [3348, 1663, 571, 131, 180]. Pacifico [2645]. Packard [593, 1957]. Padian [3572]. Padova [2259]. Pagans [1180]. Pagden [2159]. Pagli [3129]. Paicheler [2031]. Paier [2843]. Pain [1087]. Painter [3349, 444]. Painting [3160, 3398, 2976]. Pairs [1702]. Pais [2490, 2640, 1408, 2605]. Pala [1288]. Palace [2785]. Palaestra [3054]. Palas [262]. paleontologia [2516]. Paleontologist [283, 1406, 3571, 3572]. Paleontology [1590]. Palerne [210]. Paley [96]. Palla [1198]. Palm [744, 835, 2512]. Palmarini [2786]. Palmedo [2032, 2604, 1069]. Palombi [3000]. Pamela [248, 969, 1906, 350, 428, 2168, 1952]. Pamphlets [602, 885, 2456]. Pancaldi [1169]. Panchen [1592]. Pandemic [1141]. Pandit [1335]. Pandora [3293]. pane [3268]. Pangenes [3308]. Panos [672]. Pantin [1653, 2812, 1728]. Panza [2270]. Paola [1642, 1826, 1643, 254]. Paolantonio [3277]. Paoli [564]. Paolo [3350, 521, 2588, 3129, 2117, 122, 1195]. Paoloni [2756]. Papal [1467]. Pape [2103]. Paper [2140, 3161, 2274]. Papers [3104, 3148, 2739, 768, 2460, 380, 372, 1947, 2369, 1378, 144, 1379, 1214, 2117, 2766, 225, 1487, 1573, 953, 1227, 1291, 2520, 3540, 1777, 2208]. Papp [1793]. Papper [2362]. Pappus [430]. Papyri [2733]. papyrus [2149, 2330]. para [1374, 154]. Paracelsians [1376]. Paracelsica [2161]. Paracelso [443]. Paracelsus [643, 2161, 1022, 2894, 2256, 2894, 3134]. Paracelsus-Forschung [2256]. Paradigmenwechsel [1932]. Paradigms [776, 1771]. Paradis [468]. Paradise [3262, 3425, 486, 3466, 917, 1942]. paradossi [860]. Paradox [2912, 546]. Paradoxes [2118]. Paragranum [643]. parallèles [149]. Parallels [2795]. Paralysis [1413]. Parameters [2068]. Paranoid [2700]. Paranormal [1769]. Parascandola [1497]. parcours [3147]. Pardo [1911, 2158, 1470]. Parel [1372]. Parents [939, 1945]. Parerga [51, 2161]. Pares [3444]. Parezo [383]. Parfit [139]. Pargeter [572]. Pargetter [1152]. Paris [3429, 2981, 2671, 2028, 2928, 2012, 184]. Pariser [42]. Parisian [927, 762]. Parity [188]. Park [2369, 878, 133, 3035]. Parke [1886]. Parker [1477]. Parkes [1586]. Parkettierungen [673]. Parkinson [1379]. Parks [3095]. Parlors [217]. Parmentier [2508]. Paroxysms [2267]. Parshall [2282, 3511, 2160]. Parsimony [390]. Parsons [3015]. Part [2509, 1708, 3367, 2834, 1913, 329, 2655, 1928, 3137, 2422, 2972, 2970, 739, 354, 3319, 1531]. Parthey [3367]. Participant [2952]. Participation [1329]. Particle [3562, 1866, 3081]. Particles [188, 1059, 457]. Particular [3224, 153, 3130]. Partners [951]. Parts [1280, 3244, 988, 2655, 3491, 2164, 2244, 2839, 386, 2127]. Party [55]. Pasachoff [3454, 3343]. Pascal [1804, 1659, 2431, 2165, 3122, 357, 1390, 3366, 833]. Paschal [140]. Paschali [253]. Pascual [2309]. Pasini [3138]. Paskauskas [1677]. pasos [2915]. passato [2893, 1195]. passé [3509]. Passion [3451, 3356, 2773]. passioni [3139]. Passions [1617, 2267, 3398]. Past [2958, 1460, 333, 3508, 390, 3044, 1989, 720, 274, 1302, 872]. Pasteur [1412, 659, 2317, 2763, 3561]. Pasteur-Pouchet [2763]. Pasteurization [974]. Pastrana [2444]. Patai [3333, 1997]. Patent [44]. Pater [458]. Path [2060]. Pathological [2471, 331]. Pathology [3344]. Paths [2961]. Patient [2012, 2545, 936]. Patients [1036, 1028, 2250, 2278, 1661]. Patillon [1178]. Patrice [2275, 3146, 2333]. Patricia [2032, 2604, 1483, 1262, 1388, 2438, 626, 1293, 1820, 2444, 1458, 2366, 1948, 2452, 946, 2817]. Patrician [3516]. Patrick [3555, 170, 20, 435, 1740, 556, 2991]. Patrizia [1498]. Patron [230]. Patronage [927, 2563, 1165, 3497, 2020, 1286, 745]. Patrons [2603]. Patsy [2186]. Patten [218]. Patterson [69]. Pattin [729]. Pattison [185, 185]. Paty [2290]. Paul [3089, 2214, 2368, 84, 3527, 2678, 1603, 1395, 1656, 2973, 2150, 3563, 1649, 602, 80, 1889, 3148, 2510, 3364, 1241, 905, 2125, 3259, 2851, 317, 2318, 3066, 3511, 837, 849, 1947, 2369, 440, 3172, 785, 2215, 2024, 3506, 1378, 3186, 1494, 2609, 1845, 1360, 2543, 186, 2288, 2190, 2439, 1987, 26, 1573, 682, 599, 1814, 2910, 1229, 2929, 596, 676, 2119, 2926, 3580, 2694, 1542, 1671, 3583]. Paula [2086, 1133]. Pauli [2834, 2291, 189]. Pauline [2764]. Pauling [772, 2772]. Pauly [281]. Pauper [1763]. Pavlov [2630]. Pavlovich [654]. Pawel [1816]. Payer [1636]. pays [149]. Peace [582, 225, 60, 2995]. Peacock [1348]. Peale [3537]. Pearce [2722]. Pearl [2565]. Peat [3167]. Pecham [1720]. Peckhaus [1500]. Pedersen [515, 1626, 515]. Pedro [3053, 2158, 438]. Peenemunde [2852]. Peer [2687, 717, 98]. Peerless [717]. pegaséen [2338]. Peggy [2215, 3074, 1836, 1544]. Peierls [2832]. Peirce [848, 1975, 2760]. Peiresc [1165]. Pelaez [2157]. Pelagonius [2499]. Pelerin [2248, 2248]. Peloux [1037]. Peltzer [2535]. Pencil [541]. Pendulum [3306, 3320, 1019]. Penelope [34]. Penetralia [2347]. Penn [135]. Penny [3159]. Penrose [373]. pensée [2888, 451]. pensiero [1560]. Pension [803]. Peocock [1672]. People [2801, 3448, 922]. Pepe [1920, 1779]. Peptide [1223]. Peptides [1223]. Pera [2391, 1827]. Perception [2805, 904, 2819, 3133, 2262, 937, 382, 2572]. Perceptions [2470]. Percorsi [645]. Percussion [831]. Percy [1045]. Pereira [552, 1915]. perepiska [1212]. Perez [746, 1389]. Perfection [3453]. Perfectionis [1465]. perfil [321]. Performance [3161, 1871]. Perga [1113]. Pergamum [1002]. Perilous [2339]. Perils [2967]. Period [2673, 1275]. periode [2975, 2754]. Periodic [2479]. Periodicals [498, 1427]. Peripatetic [172]. Periphery [2036]. Perleth [3378]. Permutations [168]. Pernick [2695]. Perrin [99, 100, 1781]. Perrone [1918]. Perrot [522]. Perry [3584]. Person [631]. Personal [528, 480, 2196, 3393, 225, 2404]. Personality [2143]. Personhood [2883]. Personlichkeit [2843]. Perspectiva [3335, 1634]. Perspectivae [1634]. Perspective [1087, 1757, 1669, 522, 2062, 1419, 855, 3234, 1980, 283, 1545, 125, 2194]. Perspectives [3104, 3414, 2801, 1382, 2871, 2967, 317, 229, 136, 776, 7, 742, 332, 679, 992, 1264, 2760, 1184]. Persuading [1436]. Persuasion [2298]. pertinentes [2903]. Peruzzi [2811]. pesca [3442]. Peset [2600]. Pesticides [3288]. Pestre [1046, 1043]. Peter [140, 3049, 3262, 960, 2032, 2604, 869, 2019, 213, 367, 3140, 940, 2574, 767, 575, 781, 2718, 2582, 3053, 2996, 492, 2441, 3128, 1221, 167, 2785, 380, 1169, 2160, 3156, 1217, 2523, 1840, 1328, 1302, 3004, 3316, 1423, 288, 2665, 126, 790, 3093, 808, 2360, 59, 1282, 1725, 230, 1059, 3450, 2887, 3170, 238, 774, 1133, 2164, 1987, 2103, 2451, 2450, 2727, 225, 1586, 1289, 2673, 1206, 954, 2852, 2256, 518, 196, 1291, 2428]. Peterfreund [542]. Peters [3361]. Petite [1390]. Petra [3367, 1762, 1300, 3570]. Petrarca [1813]. Petreius [895]. Petrilli [848]. Petrillo [2927]. Petrochemicals [280]. Petroleum [3483]. Petroni [973]. Petrose [1097]. Petroski [2797, 541]. Petrovich [2124]. Petruccioli [860, 2118]. Petrus [1366, 1995]. Pettenkofer [2110]. Pflanzensynonymik [2070]. Pflaum [862]. Ph.D. [308]. Phaenomena [2731]. Phanomen [3385]. Phänomenologie [1150]. phänomenologischen [1150]. Pharaonique [2149, 1711]. Phares [2333]. Pharmaceutical [64, 1012, 1174, 1458]. Pharmaceuticals [1989]. Pharmacie [516]. Pharmacology [1497]. pharmacopée [916]. Pharmacopeia [2283]. Pharmacy [1191, 2322, 1497, 1458, 1458]. Pharmazeut [3447]. pharmazeutischen [3360]. Pharmazie [736, 2110, 2095, 2359, 1142]. phénomènes [2247]. Phenomenology [688, 1239]. Philadelphia [3153, 231, 289, 1458, 1689]. Philanthropy [579, 949]. Philhellenism [2728]. Philip [539, 2913, 281, 2427, 783, 1034, 968, 1773, 386, 1908, 2063, 2665, 2121, 1675, 1327, 1749, 131]. Philipp [2123, 2144, 2123]. Philippe [2511, 3572]. Phillip [3240, 742, 1394]. Phillips [1617, 2068, 3074, 1904, 1262]. Philoponu [147]. Philoponus [342, 915, 1092, 342]. Philosoph [1634]. philosophe [651, 2290, 2093]. Philosopher [1305, 1285, 655, 1639]. Philosopher-Scientist [1639]. Philosophers [2006, 2002]. philosophes [823]. philosophi [1640]. philosophia [1918, 3527]. philosophiae [2253, 3527]. Philosophic [2176]. Philosophica [2063]. Philosophical [789, 1548, 978, 1111, 1922, 1184, 189, 1182, 489, 2964, 2365, 1621, 572, 407, 2358, 1562, 2014, 2705, 692, 86, 1799, 2760, 698, 1797, 1602, 2520, 3540, 1240, 1703, 742, 3250]. Philosophically [2704]. philosophico [2272]. Philosophie [2509, 3049, 2920, 830, 1808, 925, 2977, 297, 1500, 1478, 2238, 2818, 687, 642, 1093, 549, 823, 1559, 3458, 2667, 3446, 3049, 1500, 3060, 2911, 2667, 3103]. Philosophies [3466, 206, 3588, 1647]. philosophique [3157, 3555, 2290, 3002]. philosophische [463, 2107, 1240, 2018]. philosophischen [360, 1531]. philosophorum [1721]. Philosophy [84, 486, 1061, 1603, 2527, 2427, 690, 605, 78, 650, 726, 447, 80, 1485, 207, 1536, 2170, 1000, 300, 2510, 1387, 190, 972, 1151, 1241, 1280, 2571, 1377, 3528, 1479, 2264, 1730, 2860, 905, 1119, 3210, 167, 879, 2689, 426, 2631, 3222, 1130, 2181, 3557, 2063, 548, 1408, 109, 2261, 1423, 686, 3016, 2902, 505, 1282, 355, 1647, 1059, 2887, 720, 2083, 1190, 2765, 1236, 1736, 2263, 510, 2727, 2262, 3398, 1549, 1315, 1420, 506, 79, 1628, 3533, 737, 2194, 3401, 613, 973, 1513, 1817, 2507, 258, 1908, 2151, 980, 2327, 2345, 2735, 457, 3387, 2060]. Philosophy [1652, 1238, 1873, 1685, 1451, 1772, 686, 2588, 1719, 1206, 1962]. Photographic [1458]. Photographs [85, 1458]. Photography [3348, 2676, 3227, 1491, 3275]. Phrenology [3453, 491]. phusis [1713]. Phylogenetic [288]. Physica [2583]. Physical [1008, 2293, 2678, 328, 115, 1617, 616, 3370, 2273, 158, 2880, 133, 3230, 785, 3498, 2014, 1139, 1545, 2246, 2943, 279, 772]. physicam [2903]. Physician [1563, 272, 1209, 1568, 155, 2414, 1126, 3165, 3328, 156]. Physician-Legislators [1209]. Physicians [1388, 2871, 231, 3013, 2420, 3044, 566]. physicien [1043, 460]. Physicist [99, 1833, 2431, 18, 60, 1777, 1315, 1617, 1217, 2537, 61, 2486]. Physicists [2833, 1774, 17, 192]. Physick [1926]. Physicke [3430]. Physics [539, 488, 3276, 1795, 1132, 3050, 3562, 1774, 2787, 1878, 1552, 1358, 1804, 2498, 3130, 1408, 2124, 3461, 1270, 2267, 1290, 3498, 98, 3081, 2458, 2244, 2325, 2194, 900, 1095, 1133, 2288, 672, 2531, 1210, 78, 978, 1287, 1685, 190, 2036, 2569, 621, 2042, 2118, 278, 2244, 1865, 965, 2640, 949, 1210, 207, 2377, 3489, 1133, 225, 2606, 3070]. Physik [851, 2920, 3246, 439, 3278, 2378, 755, 861, 163]. Physikalisch [3367, 466, 2535]. Physikalisch-Technische [466, 2535]. Physikalische [1316, 3546, 463]. physikalistischen [656]. Physikerin [2978]. Physikotheologie [163]. Physiognomical [2361]. Physiologia [2571]. Physiological [2374]. Physiologie [2979, 1122, 1496, 1023]. physiologique [2888]. physiologiques [3075]. physiologischen [266, 3360]. physiologiste [1496]. Physiology [1528, 1611, 3307, 2014, 2243, 2630, 698]. Physique [1201, 1546, 745, 1026, 3460, 2290]. Physis [521]. Pi [649]. Piaget [791, 2375, 2375]. Piaskowski [3455]. Piattelli [2786]. Piattelli-Palmarini [2786]. Piazzi [172]. Picard [256, 2539]. Picatrix [33]. Pickering [1870, 2550, 1939]. Pickover [3238]. Picolet [256]. Picon [1753]. Pictet [1026]. Pictoral [2965]. Pictorial [1837, 2215, 1226, 1458]. Picture [1575, 1458]. Pictures [2695, 1433, 1081, 1458]. Picturing [2493, 3348, 2964, 3180, 1665]. Pieces [2456]. Pierluigi [2277]. Piero [1183, 2421, 1555]. Pierre [2363, 687, 1008, 2743, 2509, 2010, 535, 3583, 3340, 1546, 628, 2895, 830, 2079, 671, 2432, 3538, 1897, 18, 2911, 2331, 2960, 2078, 522, 1636, 2727, 1315, 653, 842, 312, 1315, 3298, 1496, 3251]. Pierre-Noel [3340]. Pierre-Simon [3251]. Pieter [1481, 1777]. Pietro [49, 2443, 111, 919, 46]. Pietsch [2442]. Pig [3212]. Pighetti [645]. Pignon [3433]. Pijper [2310]. Pilcher [2458]. Pile [2953]. Pilgrimage [1424].

Pilgrims [1718]. Pill [1087]. Piltdown [953, 953]. Pinch [417, 2044]. Pindl [1914]. Pindl-Buchel [1914]. Pine [318]. Pineiro [1723]. Piñero [1911, 2516, 2491, 2490, 2158, 1354, 1311, 3442, 1561, 351, 829, 2435, 1024, 132, 152]. Pingree [431, 247, 920, 2071, 3431, 33]. Pinkus [3176]. Pinna [3104, 3104]. Pinto [3416]. Pinto-Correia [3416]. Pio [1311]. Pioneer [1617, 2455, 1323, 738, 1598, 219, 3174, 757, 596, 1039, 373, 3548, 1227]. Pioneering [2367, 1352]. Pioneers [2778, 1957, 2286]. Piore [876]. Piotr [2940]. Pippard [2640]. Pisa [1716, 1094]. Pitirim [2691]. Pitt [1041, 1041]. Pittenger [1668]. Piyo [2341]. Place [1362, 344, 915, 1134, 782, 722, 2296, 282, 1496, 915]. Placebo [3328]. Plagiarism [1511]. Plague [1707, 2829, 2250, 593, 2613, 3581, 2653]. Plagues [3233, 594]. Plain [2802]. Plains [3548, 1446]. Plakins [3237]. Planck [3367, 1215, 3367]. plane [1176]. Planet [2967, 1044, 2643, 2932, 3459]. Planetarium [2712]. Planetary [3007, 1450, 2245, 1137, 739]. planetas [340]. Planetis [2162]. Planets [3520, 904, 1976]. Planning [2617]. Plans [1253]. Plant [759, 2111, 350, 2352]. plantarum [2901, 1359]. plantas [2158, 2803]. Platearius [2074]. Plates [935]. Plato [2501, 2238, 2194]. Platon [1178]. platonicus [3330]. Platonismo [1107]. Play [1871, 692]. Pleasure [2284, 843]. Plehn [2081]. plein [1124]. Plethon [3524]. Pliny [905, 1714]. Plotinus [429]. Plotkin [194]. Plow [1446]. Plunder [678]. Plurality [1109]. plus. [2028]. Pluto [1044]. Plutonium [2540]. Pneuma [2418, 1383]. Pnina [2723]. poder [183]. Poeckern [453]. poema [1811]. poesie [2812]. Poetic [1105]. Poetics [3168]. Poetry [2657, 2362, 665]. Poets [484]. Pogrebyskaia [1212]. Pohl [3447]. Poincaré [1756, 3457, 2985]. Point [622]. Poirier [3538, 2297]. Poisons [333]. Polanco [718]. Poland [441]. polemica [3139]. polémicos [351]. Polemics [893]. Policies [3008, 225]. Policy [1583, 2781, 1148, 2542, 595, 1136, 2541, 1767, 1138, 1584, 2686, 1416, 717]. Policymakers [1330]. Polio [1507, 2613]. Polish [65]. Polite [2433, 2598]. Politic [2782]. politica [771, 2340, 261]. Political [2597, 2386, 1536, 3437, 2748, 67, 1174, 593, 805, 2299, 1869, 1784, 652, 225, 1511]. Politics [1232, 870, 455, 2700, 3310, 2930, 579, 287, 3558, 2641, 2269, 2781, 1209, 3309, 576, 594, 2410, 3504, 2337, 3580, 2767, 962, 277, 198, 1516, 625, 3376, 1231, 2657, 2543, 680, 1401, 186, 3135, 2288, 467, 2617, 2796, 1235, 3100, 2469, 3026, 590, 2003, 3435, 1529, 3280, 377, 676, 3090, 3032, 2765, 225, 682, 3503, 2541]. Politik [1115, 1215, 2523]. politique [2010, 2667, 925, 3458, 3556]. Politisch [940]. Politisch-biologische [940]. politischen [3341]. politischer [1531]. Polity [1408]. Polkinghorne [965]. Pollution [1834, 3234, 3325]. polowy [1655]. polskich [1655, 3455]. Polycarp [2353]. Polyhedra [3128]. Polymath [3269]. Polymer [3569]. Polyp [166]. Pomar [1561]. Pomerans [943]. Pont [3458]. pontificale [2618]. Ponts [1753]. Ponty [522]. Poole [3507]. Poor [1661]. Popkin [1469, 1112, 3533]. Popp [578]. Popper [3485, 787]. Poppi [2259]. Popular [3306, 2365, 376, 3355, 3119, 263, 1432, 2787]. Popularization [1824].

Population [960, 1676, 2208]. por [134, 1561]. Porphyry [547]. Port [2668]. Portela [829]. Portents [1726]. Porter [418, 241, 2449, 3516, 420, 1027, 2412, 1748, 1727, 572, 815, 3263, 169, 2725, 105, 1796, 2642, 2406, 1896, 3270, 2096, 1118, 1027, 1824, 335, 43, 3046, 2048]. Portia [912]. Portrait [3090, 1488]. portraits [320, 833, 591]. Positionastronomie [2979]. Positions [3126]. Positivism [804, 668, 2473]. Posselt [476]. Possessing [2086]. Possible [1134, 3006]. Post [111, 2671, 2475, 2472]. Post-Modern [2475]. Post-Revolutionary [2671]. Post-Traumatic [2472]. Postcards [3572]. Posterior [2649, 1988]. Postmodern [3100]. Postmortem [3359]. Postwar [1696, 1345, 2536, 2541, 1955, 678, 581]. Pot [3426]. Potential [3167]. Potter [2760, 1720, 1732, 26]. Pouchet [2763]. Poultry [1936]. Pound [3168]. pour [2633, 3189, 3088, 931, 1312, 729]. poussière [2496]. pouvoir [900]. Powell [46]. Power [1661, 1537, 3424, 3124, 3392, 2200, 1687, 1089, 2563, 787, 764, 3099, 556, 1083, 2883, 225, 873, 2298, 2047, 1234, 590, 2971, 384, 1329, 2718, 316, 1569, 3247, 1089, 2968, 276, 1447, 3186, 667, 142, 1903]. Powerful [3328]. Powering [2388]. Powers [355]. Pownall [1324]. Pox [3337]. Practica [731, 2074, 1181, 1374, 731]. Practical [1243, 1181]. Practice [2552, 2180, 2962, 1210, 2565, 1680, 2020, 220, 2569, 3158, 67, 2631, 1616, 1087, 1574, 1078, 2588, 1870, 2550, 865, 1087, 809, 735, 542, 263]. Practices [3464, 2704, 141, 1153, 2817, 2847]. Practique [2248]. Practitioners [2250, 2666, 3191, 2693]. praecognitionibus [1103]. praecognitis [1103]. praestigiis [1371]. Pragmatic [780]. praise [613]. Prakrit [3490]. pratique [2514, 820, 2290]. Praxis [1481, 3535]. Pre [2339, 3514, 1732, 1607, 599]. Pre-1500 [599]. Pre-1830 [1607]. Pre-Enlightenment [2339]. Pre-Industrial [1732]. Pre-Telescopic [3514]. Preacher [1388]. Preacher-Physicians [1388]. Preadamite [1455]. Précédeé [349]. précis [931]. Precision [2383, 2285, 843, 2068, 256, 2405]. Preclassical [1289, 1627]. Predecessor [679]. Preestablished [1730]. Preformation [3416]. préhistoire [149, 3067, 2495]. Prehistorian [1620]. Prehistoric [1837]. Prehistory [908, 1940]. Preindustrial [330, 1173]. Prejudice [1594]. Preliminary [838, 2420]. Prelli [808]. Prelog [1134]. Prelude [2237]. premices [3144]. premiere [925, 1390, 1023]. premières [2495]. Premiers [2495, 451]. Premodern [1187]. Premontré [731]. Present [1087, 1904, 2799, 2958, 2529, 3008, 2694, 3044, 619, 720, 425, 2802, 3509]. Presented [531, 1802]. Preserve [620, 1578]. Preserving [3095]. Presidency [3183]. President [2225]. Presidente [2516]. Presidential [1584]. Presocratics [635]. presqu’île [981]. Press [376]. Pressman [3294]. presso [2260]. Preston [948, 140]. Preussen [1304]. Preventing [968]. Prevention [947]. Preventive [1933, 2279]. Price [2034]. Priesner [3256]. Priest [3328]. Priestley [3342, 2277]. prijelomima [930]. prima [3527, 1471]. primaire [3298]. Primar [1393]. Primar- [1393]. Primary [3354, 1781, 2606, 1269]. Primate [2544, 478]. Primer [1409, 3475]. primera [352]. primeras [2158]. primeros [2915]. Primers [217]. Primitive [538].

Prince [1763]. Princeton [1573, 986]. principe [834, 2895]. Principes [2354, 451]. Principia [1656, 838, 2487, 3527]. Principien [1130]. Principle [534, 1121]. Principles [451, 1087, 1114, 1087, 1547]. Pringle [2766]. Print [1099, 2160, 347, 1511]. Printed [492, 600, 1566]. Printing [3127]. Prinzip [842]. Prior [1001]. Priority [1656, 448]. Prisoners [2785]. prisposoblenii [910]. Private [1038, 2317, 95, 2276, 53]. Privilege [61]. Priya [1907]. Prize [1777]. Prizes [3199]. pro [3208]. Probabilistic [1060]. probabilite [2820]. probabilités [2508, 3157]. Probability [418, 2194]. Probable [1382]. Probably [2833]. Problem [3242, 1744, 2805, 1192, 2584, 3485, 166, 3570, 3529, 2985]. problema [2270]. probleme [1237]. problemes [3062]. Problemgeschichte [2018]. Problems [1673, 1162, 426, 2964, 1075, 723, 241]. Proceedings [1656, 3252, 650, 3511, 1011, 1087]. Process [709, 869, 1399, 1584]. Processes [2578]. Processing [211]. Procter [1497]. Proctor [1083, 2469]. prodromus [358]. Production [718, 2578, 1625, 2951, 3286]. Productivity [3161]. proefondervindelijke [994]. Prof [3003, 1941]. profesion [3445]. professeurs [2132, 212]. Profession [1851, 2982, 181, 3044, 945, 126]. Professional [1978, 1497, 2373, 3220, 263]. Professionalisierung [286]. Professionalism [1052, 198]. Professionalization [1678, 1208, 2761, 1211]. Professionals [3025]. Professionen [2131]. Professions [1263, 75, 1211]. Professionselite [3586]. Professor [2540, 2457]. Profile [3367]. Profiles [2997]. Profit [843]. Prognosis [1641]. Prognostication [920]. Prognosticon [549]. Program [225, 2685, 2641, 781, 3382]. progrès [1063]. Progress [2912, 1844, 2916, 1302, 1025, 3011, 22, 3426, 872, 2003, 3280]. Progressive [2673]. Project [3563, 3383, 2130, 966, 2003, 1144, 3170, 225]. Projections [1702]. Projects [3052]. Prolegomena [3064]. Proliferation [3558]. Prologus [1366]. Prometheus [415]. Promising [858]. Promoting [1193]. Promotionen [1660]. Prompt [2924]. Prooemium [2510]. Proofs [1547]. Propagation [2952]. Prophecy [2078, 840, 556, 922]. Prophet [3575]. Prophets [3245]. Proportion [2591]. propos [2320]. propriis [39]. Prose [2996]. Prospect [318]. Prosper [1053]. prosteishikh [910]. protagonistes [1026]. Protagoras [1549]. Protect [3288]. Protection [1529]. Protein [2148]. Proteins [3010]. Protestant [177, 668]. Protestantism [3316]. Protestants [2114]. Proteus [3225]. Protopharmacology [1153]. protsessa [1336]. Proust [1418]. province [1043]. Prozac [3235]. Prozess [1115]. Prusse [2248]. Ps [2421]. Pseudo [1715, 2503, 1465]. Pseudo-Albertus [1715]. Pseudo-Geber [1465]. Pseudo-Mesue [2503]. Pseudosciences [1566]. psichiatria [1498]. Psicología [525]. Psychedelic [2566]. Psychiatric [3379, 181, 3579, 1208, 3046, 2298]. Psychiatrie [2827, 2019, 1937]. Psychiatry [855, 3235, 3024, 3098, 2128, 3046, 865, 665]. Psychical [603]. Psychoanalyse [1240]. Psychoanalysis [2700, 2193, 3582, 1054, 499, 1240, 2700]. Psychoanalyst [2928]. Psychoanalytic [1744]. Psychocultural [1807]. Psychogenesis [791].

Psychological [1662, 1708, 866, 2055, 3417, 1550, 596, 3047]. Psychologie [2018, 286, 1322, 2031]. psychologiegeschichtliche [1322]. Psychologies [462]. Psychologist [3293]. Psychology [1673, 1597, 1678, 2287, 2794, 1049, 382, 2780, 2547, 3469, 2299, 3390, 2807, 1401, 2883, 1758, 2286, 596, 416, 779, 778, 66, 1323, 2965, 1149, 51, 1010, 543, 1009, 2374, 2575, 1597]. Psychopathology [2779]. Psychophysicus [51]. Psychophysik [51]. Psychosomatic [1413]. Psychosurgery [3294]. Psychotherapy [3380]. Pt [2273]. Ptolemaic [2650]. ptolemaico [2741]. Ptolemaus [1360]. Ptolemee [1361, 914]. Ptolemy [1894, 1548, 2068, 1360, 914, 428, 1015, 1548, 3133]. Ptolomeo [340]. Pubblica [1920]. Public [3408, 716, 655, 1480, 1934, 3153, 947, 2406, 2641, 1832, 1479, 576, 2783, 594, 1968, 217, 962, 1499, 470, 1078, 1971, 760, 1732, 3584, 877, 2279, 1329, 3538, 3576, 3537, 2276]. pública [1024]. Published [513, 1518, 1293, 301, 1291]. Publishing [895, 1511]. Puerto [2681, 261]. Pui [3068]. Puig [1095, 432]. Pulkkinen [2289]. Pullman [3413]. Pulse [1019]. Pulte [2984, 842]. Pumpelly [1836]. Pupils [2666]. Pure [3064, 670, 670]. Pureté [823]. Purgantien [2503]. Purgantien-Lehre [2503]. Puritanism [928]. Purity [1083]. Purks [43]. Purple [249]. Purposes [968]. Purrington [3070]. Pursell [2234]. Pursuit [528, 2383, 575, 2406, 2042, 268, 2710, 872]. Pycior [2723, 3143]. Pyenson [3302, 130, 179, 1670, 2601]. Pygmies [2203]. Pyne [3160]. Pythagora [2325]. Pythagoras [548]. Pythagorean [2887].

Q [2276, 1764, 2676]. q-Numbers [1764]. Qadir [1451]. QED [2301]. Qing [2615]. Quadrant [3561]. quadratura [2089]. Quadripartitum [1369]. quadrivium [1993]. Quaestio [2076]. Quaestionen [3246]. Quaestiones [2892]. Quality [2068, 1504]. quanta [1847]. Quantentheorie [2920]. quanti [3277]. Quantification [2734, 2492]. Quantifying [841]. quantique [3002]. quantita [2270]. Quantitative [3255, 2273]. Quantities [2859, 2896]. Quantity [1504]. Quantum [2309, 3276, 487, 690, 78, 2463, 1764, 1216, 2769, 1950, 586, 1686, 3168, 1421, 2379]. Quarantine [3077]. Quark [3081]. Quarrels [2315]. Quarta [3534]. Quarterly [111]. que [1911]. Quebec [2019, 229]. quebrada [261]. Queen [1562]. Queer [2701]. Quellen [1185, 2251]. Quellen-Texte-Wirkungsgeschichte [2251]. Quelleninventar [3367]. Quest [3502, 2700, 1673, 3252, 207, 3320, 3529, 664, 1087, 1504, 2492, 3290, 1263, 873, 2932]. Question [1866, 126]. questione [1920]. Questions [3528, 1058, 1095, 2077, 1279, 3292, 1720]. Questions/Women [3292]. Quhi [2877, 1809]. qui [587]. Quiet [1751]. Quimica [2444]. quina [2272]. Quine [2303, 883]. Quine-Carnap [2303]. Quinn [2639]. Quinologia [2272]. quintae [730]. Quipu [526]. Quiroga [2776]. quirurgica [3445]. Quod [554]. Quorundam [3244]. Quotable [2944]. Qurra [2877]. Qusta [1718, 339, 1718].

R [3089, 1336, 1484, 39, 197, 2962, 1395, 3537, 105, 1795, 487, 1830, 1617, 482, 420, 1864, 2828, 993, 3482, 2532, 207, 2108, 3315, 2680, 2526, 378, 76, 1604, 3410, 3269, 669, 1469, 918, 2250, 2969, 2085, 1218, 2798, 2504, 331, 1106, 2759, 3348, 2169, 1684, 3128, 584, 941, 3536, 1761, 2187, 2448, 3266, 1134, 1038, 595, 836, 2049, 1957, 1208, 920, 3172, 1959, 1452, 2420, 3072, 3483, 1085, 387, 1192, 1928, 2997, 2015, 519, 503, 2927, 1719, 1759, 3485, 1379, 606, 2358, 1271, 742, 1677, 3287, 3510, 3248, 1465, 3261, 3468, 493, 22, 1803, 2288]. R [1404, 1741, 155, 1467, 365, 2859, 1327, 1236, 3541, 888, 1275, 421, 2243, 531, 3263, 1029, 225, 966, 3433, 1475, 3282, 485, 2387, 1315, 506, 631, 1584, 1544, 2202, 2731, 477, 143, 3430, 1207, 661, 2730, 1416, 1503, 1188, 276, 3171, 1802, 1318, 71, 3185, 999]. R&D [95]. R. [1936]. rôle [653]. Raalte [3239]. Rånby [279]. Rabbi [249]. Rabbinic [1989]. Rabbis [3245]. Rabdology [1104]. Rabies [3581]. Rabl [939]. Raby [3262]. Race [2546, 1325, 2189, 1594, 2853, 1143, 3585, 1220, 3495, 1326, 478, 3099, 225, 3231, 2610, 676, 968, 2129, 225, 1956, 474, 1541]. Rachel [598, 3425, 2211, 1510]. Racial [2189, 1900, 272, 3309, 3222, 3028, 1509]. Racism [1951, 1594, 3359, 2696, 1325]. Radar [3007]. Radder [489, 2705]. Radeletde [2438]. Radiation [2213, 1529, 1145, 2925]. radiazioni [3277]. Radical [467]. Radio [2684, 3385, 1586]. Radio- [3385]. Radioactivity [3174]. radioterapia [3475]. Radium [3474]. Radner [490, 490]. Rae [1699, 1902]. Rafael [183]. Rafailovich [1878]. Raffaella [1101]. Rafter [273]. Ragep [1991]. Rahman [1810]. Railroad [246]. Railroads [2992]. Rails [2783]. Rainbow [1752]. Rainer [40, 578]. Raines [2826]. Rainger [1590, 1318]. Raising [3027]. raisons [2004]. Raj [2669]. Ralf [3060]. Ralph [250, 85, 1133]. Raman [3223, 964]. Ramanujan [1213, 2380]. Ramon [3442, 1760, 525, 1914]. Ramona [237]. Ramsey [3131, 263]. Ramunni [3088]. Rand [2162, 779, 387]. Randall [291, 593, 3588]. Randolph [591]. Random [2794]. Randomness [3322]. Rank [2780]. Rankin [2380]. Ranking [2958]. Ranks [3293, 2366]. Ransil [379]. Raoul [1760]. Raphael [654, 433, 1997, 3333, 792, 2475, 1836]. rapides [931]. Rappaport [3141]. Rapport [719, 1728, 931]. rapports [2166]. Raquel [2157]. Rare [1248, 699]. rares [2010]. Rarest [2442]. rari [2744]. Rasch [1215]. Rashed [522, 510, 749, 150, 1809]. Rashomon [3472]. Raspadori [1267]. Rasse [196]. Rassenhygiene [196]. Rate [956]. Rather [939]. Rational [2348, 683, 3528, 2584, 1384, 3286]. rationalen [842]. Rationalism [3522, 2390]. rationalisme [2348]. Rationalist [1893]. Rationalite [2363, 1729]. Rationality [693, 732, 3485, 204]. Rationalizing [2847]. Rationis [3054]. Rattansi [2341]. Rattner [3346]. Ratzel [2988]. Raufer [3325]. Raven [1087]. Rawcliffe [2654]. Ray [2451, 1750, 2685, 2286, 1238]. Raylor [2343]. Raymond [2386, 37, 1280, 1053, 1134, 3524, 2425, 552, 2729, 1814, 1600, 2565]. Rayner [1617, 3174, 3174]. Rayner-Canham [1617, 3174, 3174]. rayon [2496]. razionale [645]. razionalita [3457]. Razon [1811]. Re [3050, 2155]. Re-Examination [3050]. Reaction [1656, 3115, 1198, 1329]. Reactions [1377]. Read [2140, 2153]. Reader [598]. Readers [1995]. Reading [2337, 1279, 1253, 2951, 406, 534, 1423, 1202]. Readings [734, 79, 1022, 1065].

Reagan [3079]. Real [2384, 912, 2517, 1227, 164, 71]. Realism [1061, 1576, 295, 3297, 691]. Realist [1216]. réalité [687]. Realities [3293, 191]. Reality [2734, 1058, 2572, 1243]. Realization [489]. Really [3321]. Reappraisal [1674]. Reappraisals [1977, 636, 1100]. Reappraising [3239]. Rear [2227]. Reardon [1402]. Reardon-Anderson [1402]. Reason [130, 1241, 1744, 840, 2786, 3063, 694, 3058, 3064, 2637]. Reasonable [1382, 2662]. Reasoning [790]. Reassessing [499]. Reassessment [2681]. Réaumur [166]. Rebecca [3549, 2826, 2207]. Rebel [3314, 2366]. Rebellion [1475]. Rebels [3293]. Reborn [692, 1815]. Received [2898]. Recently [2139]. Reception [1888, 2479, 58, 3306, 407, 2244]. Receptor [1244]. Rechenberg [3166, 473, 586]. Rechenbuch [348]. Rechenmaschinen [1544]. Recherches [749, 150, 1711, 3020]. Rechnen [2973, 1013]. Recht [2281]. Récit [1043]. Reckendorf [1505]. Reckenfelderbaumer [2614]. Recluse [2055]. Recollections [3212, 1134]. Reconsideration [143]. reconstitution [545]. Reconstructing [1603, 390]. Reconstruction [1336, 2536, 1489, 3450]. Record [1321, 1458, 3430, 626]. Recording [2026]. Recordings [2461]. Records [3072, 506, 1746, 144, 2676]. recues [1982]. Red [996]. Red-Hair [996]. Redaktion [1996]. Redeeming [3362]. Redirecting [949]. Rediscovering [1332]. Rediscovery [1684]. Redlich [3575]. Redman [605]. Redner [227]. Redondi [111]. Redshaw [3074]. reduccion [1374]. Reduced [617]. Reduction [1192]. Redux [3499]. Reed [3482, 382, 3106, 2373]. Reeling [329]. réelle [1042]. Reenchanted [2922]. Rees [3250]. Reese [372]. Reevaluation [1198]. Reeves [2976]. Reference [3299]. References [605]. Refiguring [2608]. Reflection [2267]. Reflections [1617, 978, 1953, 1315]. Reflects [3364]. Réflexions [1023]. Reflexive [882, 789]. Reflexivity [82]. Reform [947, 96, 3011, 2137, 3507, 1652, 467, 3474, 577, 1303]. reforma [1374]. Reformation [2343, 2894, 1725]. reforme [2508]. Refraction [36]. Refrigeration [2105]. regard [821, 2511]. Regards [2099]. Regelen [2153]. Regeneration [1504, 1760]. Regime [1718, 2747, 751, 2011, 1384]. Regimen [2251, 3154]. Reginald [272]. Regine [3080, 2581]. Regiomontanus [1013, 1098]. Regiomontanus-Bianchini [1013]. Region [2617]. Regional [2783, 721, 2544]. Regions [664]. Regis [2155, 2618]. Register [2987, 1133]. Regius [2658]. Regna [777]. regola [3129]. Regulation [2212, 481, 1867]. Regulatory [2781]. Rehabilitation [1823]. Rehbock [1749, 131]. Reich [3003, 2300, 1501, 2755, 2382, 2842, 1142]. Reichsanstalt [466, 2535]. Reid [3148, 1394]. Reihe [2013]. Reihentheorie [2008]. Reijer [2946]. Reill [2665]. Reilly [1327]. Reingold [1204, 3117]. Reinhard [1762, 2195]. Reinhold [1015]. Reino [2517]. Reinterpretation [1257]. Reinterpreting [367]. Reinventing [2846]. Reinvention [955]. Reisel [3060]. Reiser [1709]. Reisinger [3413]. Reiss [3522]. Reith [519]. Réjane [423]. Rekonstruktii [1336]. Rekonstruktion [786]. relacion [2516]. Relaciones [2813, 2158]. Related [3019, 2303, 253, 1462, 34, 387, 2456, 225]. Relating [3148, 953]. relation [687, 317]. Relations [996]. Relationship [1669, 3564]. Relative [622]. Relativism [972, 1602, 295]. Relativistic [671]. relativistica [2347]. relatività [3277]. Relativitatstheorie [1313, 1314, 2755]. relativité [2900]. Relativity [2952, 769, 58, 1236]. Relativization [6]. Relay [910]. Relevance [420]. Relevant [1585]. releynykh [910]. Reliefs [3163]. religieuses [820]. religieux [255]. Religion [2368, 455, 992, 926, 2269, 2345, 1542, 3362, 3556, 2022, 2910, 46, 732, 1283, 1532, 2667, 447, 3363, 2083, 1455, 2860, 2702, 1275]. Religious [2806, 319, 1799, 3353, 1797]. Reluctant [230]. Remak [2523, 143]. Remane [654, 530]. Remarkable [1680]. Remembered [3463, 2202]. Remembrances [591, 1953]. Reminiscences [3565, 2199]. Remote [85]. Removing [1428]. renacentista [351]. Renacimiento [2434, 921, 152]. Renaissance [902, 1371, 555, 985, 1553, 1469, 2156, 735, 1817, 1182, 2160, 2063, 440, 2737, 553, 3051, 3337, 922, 639, 1815, 3052, 636, 1778, 2159, 739, 736, 1099, 2337, 1565, 2505, 828, 644, 350, 1919, 2324, 1283, 733, 441, 2002]. Renard [2442]. Renate [244, 517, 639]. Renato [940]. Renatus [2238]. Rencontres [1093]. rendez [3088]. rendez-vous [3088]. Rene [2743, 3527, 1673, 1111, 1201, 1192, 1475, 536, 739]. René-Just [536]. Renee [1003, 3190]. Renn [2460, 2117, 1289]. Renneberg [1762]. Repair [3573]. Reparations [678]. Replacement [1177]. Reply [99, 100]. Report [1795, 968, 1969]. Reporting [3400]. Repositories [606]. Representation [2899, 574, 3410]. Representations [1324, 1837, 2665, 639, 2409, 246, 820, 141]. Representative [2152]. Representing [893, 892]. Reprints [28]. Reproduction [2142, 3231, 718, 3376]. Reproductive [3014]. Republic [3537, 2745, 1209, 2113, 2351, 498, 568, 2501, 1600]. Republicain [3147]. Republican [2950]. Republik [2914, 1322]. République [161]. rerum [1379, 730, 3531]. Rescher [1110]. Research [1967, 2931, 1857, 580, 2549, 107, 2565, 2385, 781, 3027, 3566, 2869, 64, 2282, 3158, 968, 2785, 664, 3214, 2701, 2055, 805, 1585, 2555, 603, 3376, 2566, 2541, 1309, 397, 2823, 1133, 2140, 421, 2387, 299, 679, 1031, 2544, 2231, 590, 874, 866, 1328, 2038, 1068, 1876, 1511, 979, 1585, 1977, 1600]. Researches [1392]. Resistance [3102]. Resolute [756]. resolution [3062]. Resonance [2778]. Resort [3294]. Resource [3336, 2159, 1338, 397]. Resources [888, 300]. respirant [2888]. Response [3153, 3420, 1582, 2236, 98]. Responsibility [2225]. Restauration [3552]. Restitution [1113]. Restivo [1770]. Restless [2967, 1859, 2815]. Reston [2084]. Restoration [1295, 455, 1249, 646, 2810]. Restored [917]. Resultate [2256]. Resurrection [2931, 3052]. Retardation [2115]. Retarded [2699]. Rethinking [2610]. Retreat [1594, 865, 665]. Retrospective [2714]. Return [733]. Returns [1516]. Reuben [3321, 3087]. Rev [2521]. Reveal [1750]. Revealed [1293, 1942]. Revealing [3090]. Revelle [2841]. Reverby [70]. Review [248, 392, 57, 1626, 627, 528, 1706, 3346, 2726, 3089, 2251, 1490, 3361, 3104, 2684, 2931, 513, 3039, 1981, 140, 244, 161, 1336, 539, 971, 2363, 202, 2214, 2896, 1911, 3181, 25, 130, 709, 819, 2329, 2546, 2914, 3178, 3408, 2284, 2453, 2548, 2179, 2368, 525, 868, 1024, 2552, 1308, 30, 250, 84, 719, 1564, 2764, 3094, 601, 486, 2343, 236, 2180, 3019, 138, 298, 1147, 1232, 2148, 2695, 3317, 1894, 2433, 562, 740, 1563, 3445, 1043, 1213, 1484, 2018, 559, 687, 1008, 1874, 1891, 2348, 2659, 2743, 3339, 3527, 2509, 1413, 2086, 3475, 2661, 286, 1322]. Review [1857, 1951, 2059, 2843, 789, 3054, 2293, 2678, 1177, 1468, 3334, 201, 264, 1128, 2692, 1581, 2206, 2461, 2639, 715, 418, 1061, 2010, 2597, 3418, 1288, 2000, 3502, 1583, 3352, 1982, 488, 1316, 2237, 1603, 2570, 3249, 734, 2294, 2527, 3049, 2913, 1859, 3284, 330, 902, 818, 39, 2346, 2468, 327, 2647, 291, 1325, 2384, 2120, 3262, 2494, 3555, 132, 713, 3096, 1718, 1661, 2979, 1999, 2645, 2596, 3419, 281, 997, 2989, 3512, 882, 3255, 3545, 3460, 769, 2189, 1923, 2681, 3182, 124, 1313, 2531, 1017, 2383, 858, 2797, 197, 580, 2611, 2962, 3414, 1018, 2812]. Review [87, 1395, 2839, 3581, 2370, 1671, 82, 2289, 1135, 2427, 52, 752, 856, 1036, 1324, 1594, 1690, 960, 1904, 2069, 325, 848, 2021, 3042, 554, 1295, 1371, 2408, 870, 2032, 2604, 2799, 3274, 3537, 338, 430, 544, 1548, 1713, 1986, 2386, 3549, 50, 939, 869, 3520, 622, 1656, 1638, 345, 2725, 170, 587, 2879, 3447, 455, 105, 241, 754, 1028, 1115, 1481, 1660, 2177, 2353, 2754, 3043, 629, 1055, 1623, 1751, 3079, 906, 1662, 2240, 2677, 2853, 3272, 1331, 2434, 2314, 396, 567, 770, 2019, 2301, 2700, 2944, 3276, 736, 1606, 1673, 1708, 568, 328, 1048, 1795]. Review [3252, 487, 690, 362, 1306, 213, 1317, 2846, 1806, 3299, 115, 302, 846, 1483, 2788, 2464, 2549, 49, 367, 623, 1262, 3466, 2973, 3503, 716, 961, 437, 2075, 2150, 571, 2521, 851, 2691, 68, 2493, 1715, 148, 58, 517, 1210, 3367, 605, 2580, 16, 160, 233, 359, 1738, 1830, 3464, 3573, 388, 852, 1081, 1537, 1943, 2449, 2636, 3451, 2356, 2752, 3543, 2915, 1084, 1294, 1734, 1828, 1534, 1576, 2392, 78, 1132, 874, 1440, 1597, 1678, 3271, 1707, 2587, 2653, 2829, 3516, 1152, 778, 1851, 2287, 2801, 2834, 2920, 3563, 2848, 2824, 3267]. Review [1679, 2622, 2833, 3547, 3350, 3577, 3160, 3246, 482, 3479, 1374, 650, 1873, 321, 526, 1543, 3281, 2376, 535, 3140, 981, 3000, 2591, 1635, 2436, 3097, 2303, 2912, 37, 718, 1388, 2062, 2114, 2768, 940, 3415, 180, 663, 2574, 2930, 633, 2100, 913, 3472, 555, 1066, 1189, 951, 2382, 107, 352, 420, 838, 928, 2516, 2942, 3116, 3356, 1807, 3463, 726, 1424, 1461, 1796, 3477, 570, 1910, 2334, 2302, 2321, 240, 516, 744, 835, 927, 1457, 66, 655, 866, 577, 1765, 1864, 3424, 2674, 2828, 3453, 447, 1649, 1702, 1998, 2773, 3240, 3583, 2104, 2429]. Review [1996, 3378, 911, 2233, 121, 2115, 2209, 978, 1332, 810, 970, 2081, 3295, 2035, 993, 2806, 2891, 1927, 2099, 3020, 579, 809, 2252, 3482, 959, 172, 602, 632, 887, 910, 1212, 3242, 1776, 2027, 2157, 2491, 2600, 2982, 15, 3050, 2554, 2381, 2463, 2919, 3562, 3014, 2884, 29, 342, 1007, 1366, 2216, 3349, 3163, 80, 287, 2532, 270, 1080, 1382, 1540, 2405, 3409, 767, 1419, 1485, 1975, 2890, 2905, 207, 290, 909, 992, 1438, 1536, 2170, 2885, 2899, 2958, 1200, 3369, 3481, 1000, 200, 2108, 791, 2260, 2937, 2529, 3082, 2416, 149, 530, 3553, 575, 3073, 417, 833, 1111]. Review [1193, 1287, 2514, 3315, 1321, 2800, 748, 1045, 1774, 1866, 2680, 3003, 1685, 3080, 443, 1641, 1716, 3189, 208, 694, 2658, 1396, 1049, 1579, 1451, 2744, 2578, 2443, 2526, 2590, 391, 3148, 20, 378, 2579, 3007, 2909, 435, 1567, 3558, 1731, 3473, 585, 990, 1956, 2385, 3405, 2646, 1105, 1373, 2082, 2506, 781, 1480, 2275, 2750, 1705, 11, 1605, 2718, 3297, 1051, 2787, 2751, 187, 723, 1680, 1934, 465, 3108, 76, 1060, 582, 876, 2205, 2888, 855, 3235, 1407, 2380, 1230, 1191, 1553, 3448, 253, 1274, 1993, 2247, 2582, 3243, 547, 1001, 1091, 1362, 1547, 2462, 2094, 3165]. Review [2285, 3107, 2608, 2642, 2874, 3053, 239, 2922, 647, 3236, 3354, 1171, 1604, 1878, 2119, 134, 561, 300, 475, 654, 696, 885, 948, 3169, 3488, 3153, 520, 2510, 429, 2996, 1913, 3358, 3373, 2782, 1657, 662, 820, 24, 947, 1201, 1310, 2110, 2777, 3131, 1740, 2455, 110, 1205, 74, 316, 1954, 2406, 536, 673, 480, 2732, 1644, 853, 958, 1334, 3320, 651, 1387, 1518, 1666, 2442, 3410, 118, 3269, 3416, 129, 1800, 70, 3364, 190, 1176, 2242, 370, 3078, 3370, 209, 492, 714, 926, 1099, 1427, 2230, 2739, 967, 669, 1143, 1460, 1596, 2757, 1469, 2445, 322]. Review [1900, 3381, 550, 918, 972, 3159, 1555, 1721, 2825, 2641, 353, 540, 645, 743, 1101, 1284, 1477, 1556, 1726, 1960, 2269, 2355, 2978, 3522, 2156, 266, 2106, 934, 157, 521, 783, 973, 1513, 1771, 1922, 2001, 2084, 2589, 3340, 3437, 261, 1832, 1896, 336, 2454, 2277, 2283, 2322, 1595, 2020, 2748, 2967, 3368, 3232, 574, 3529, 1369, 1464, 3241, 2687, 2745, 691, 1686, 3326, 2662, 106, 1580, 2273, 1439, 775, 881, 1764, 1871, 112, 1296, 2102, 2183, 2354, 2903, 2984, 3534, 3086, 2250, 2654, 2969, 433, 509, 642, 1093, 1184, 1997, 2163]. Review [111, 279, 358, 514, 2036, 2199, 2975, 3531, 2525, 2441, 62, 2837, 3010, 1517, 683, 788, 977, 1151, 1241, 1347, 3270, 1280, 2418, 1323, 1247, 451, 2508, 2571, 2172, 1125, 2906, 479, 1377, 2173, 3406, 3294, 3528, 1479, 2264, 2719, 674, 1983, 1730, 2085, 3440, 3566, 1570, 1218, 1441, 2316, 2193, 2723, 600, 843, 991, 3009, 1988, 2798, 3145, 210, 333, 735, 1557, 2504, 2030, 194, 2619, 2917, 3389, 884, 2553, 189, 2898, 3058, 2860, 3183, 272, 474, 1053, 413, 2098, 1699, 331, 2779, 158, 905, 1106, 1491, 1688, 1816, 1920, 2336, 3088, 2026, 591]. Review [2759, 1034, 2933, 524, 1004, 2175, 2272, 2490, 3442, 1712, 36, 344, 439, 1185, 1552, 3298, 3429, 1351, 2342, 3348, 2125, 2389, 3427, 247, 1466, 1667, 1893, 3118, 975, 1119, 1436, 1569, 1801, 2096, 2169, 2998, 1453, 2963, 1399, 1684, 2569, 1561, 1723, 2158, 2880, 771, 860, 3001, 3508, 1953, 3166, 3454, 3168, 64, 1560, 2338, 3245, 542, 1411, 1817, 724, 1269, 2620, 2781, 3061, 2595, 2926, 3146, 159, 3128, 2698, 2993, 1013, 1103, 1196, 1474, 2007, 1221, 3175, 139, 231, 915, 1182, 2332, 2734, 3244, 463, 584, 1486, 1737, 1833, 2282]. Review [2679, 199, 826, 1012, 3158, 2981, 3413, 48, 235, 941, 1454, 1589, 1837, 3314, 1860, 67, 167, 259, 656, 1664, 2447, 2637, 1733, 2088, 834, 1476, 2507, 2593, 2660, 1838, 3379, 751, 1209, 3068, 1646, 1558, 3062, 3184, 879, 3318, 38, 446, 1021, 1118, 1299, 1899, 1978, 3259, 3536, 2894, 3087, 688, 1242, 1633, 1402, 3038, 968, 2851, 3443, 317, 1058, 1744, 2318, 228, 326, 1219, 1936, 2738, 1591, 1761, 147, 1358, 564, 935, 2178, 3066, 3380, 3582, 339, 825, 917, 1003, 1645, 2155, 2835, 3142, 2689, 1745, 1834, 2091, 2785, 3234, 749, 841, 2093, 2438, 2501]. Review [3511, 426, 549, 1546, 10, 504, 628, 1027, 2412, 2602, 3151, 193, 243, 2827, 3382, 3587, 1187, 823, 195, 3124, 3420, 3383, 1107, 1709, 2897, 3254, 768, 2023, 2460, 2895, 664, 2471, 185, 382, 837, 867, 1148, 1819, 1844, 1924, 2965, 693, 2924, 471, 710, 1109, 2166, 153, 258, 450, 830, 1488, 2187, 2805, 3517, 2826, 2690, 1872, 2079, 1356, 489, 2271, 2377, 3372, 1805, 2581, 227, 3360, 3005, 1216, 380, 793, 1409, 1502, 2964, 2123, 366, 1169, 2017, 2542, 1149, 341, 3489, 583, 671, 863, 1044, 1418, 1450, 1853, 2235, 2372, 2411]. Review [2530, 3083, 3407, 969, 2616, 1112, 3336, 267, 618, 861, 994, 1130, 1945, 2365, 1174, 3559, 1808, 484, 181, 512, 1655, 738, 1005, 3136, 2146, 3101, 329, 1400, 849, 2432, 2650, 2959, 573, 3388, 576, 269, 976, 1748, 1368, 2160, 174, 133, 464, 534, 621, 900, 1753, 1804, 1113, 56, 1762, 2181, 3538, 1146, 1514, 460, 2068, 2783, 1300, 2936, 2335, 2701, 3228, 817, 2042, 2197, 334, 1839, 2778, 2941, 1622, 2495, 3121, 1462, 1932, 2771, 3394, 372, 569, 1096, 3565, 390, 1773, 594, 795, 285, 1056, 1414, 2371, 3260, 3322, 675, 3156, 1946, 2281, 773, 1223, 2192, 1108]. Review [1929, 2594, 3057, 3551, 995, 1947, 2410, 2536, 2577, 2961, 3331, 2769, 386, 1095, 1276, 1908, 364, 2016, 2448, 2780, 3293, 3585, 361, 840, 1297, 1357, 1727, 1980, 684, 2423, 2503, 1898, 2188, 3266, 3557, 2547, 3006, 1365, 2095, 150, 924, 1616, 2063, 2858, 3504, 2934, 1134, 44, 1038, 925, 1961, 2004, 2437, 2977, 2470, 297, 1921, 2337, 89, 324, 1588, 1752, 2080, 2122, 3278, 3568, 593, 996, 145, 438, 548, 595, 728, 1011, 3524, 3579, 3580, 268, 780, 1183, 1397, 511, 1217, 836, 923, 1087, 1559, 2916, 2212, 3288, 2523, 2198, 2519, 1442, 2816, 85, 1768, 2767]. Review [864, 1742, 1412, 2749, 178, 1574, 154, 1224, 2290, 2652, 454, 351, 1363, 1470, 1571, 2586, 432, 2072, 2151, 1082, 163, 2849, 2498, 3130, 814, 2881, 3422, 3275, 3441, 3456, 2755, 507, 2446, 2537, 2836, 3546, 1840, 273, 2049, 2467, 346, 431, 533, 3152, 3190, 3392, 3300, 1239, 284, 3064, 1408, 206, 1097, 3561, 1447, 1843, 1957, 2061, 2200, 2369, 3176, 103, 1364, 2528, 3056, 733, 1687, 441, 725, 1178, 626, 3578, 1935, 2440, 1062, 3486, 1328, 1506, 1593, 2317, 2838, 3313, 179, 2367, 3090, 3167, 2804, 1186, 363, 3033, 12, 2124, 2291, 2320, 2918, 3461, 1775, 440, 2207]. Review [34, 442, 1123, 1835, 2013, 2184, 3075, 2921, 3008, 2668, 3588, 1825, 1421, 2737, 962, 1499, 878, 907, 1849, 2876, 237, 459, 980, 1539, 3273, 277, 1668, 3365, 470, 1035, 1208, 1824, 1897, 2671, 1220, 3374, 553, 831, 2077, 2241, 2651, 1124, 152, 1620, 3247, 2682, 551, 1279, 1463, 2845, 3230, 1173, 35, 920, 904, 3179, 2066, 3013, 952, 1204, 1577, 1659, 1930, 2840, 3403, 2073, 1244, 1102, 1621, 1779, 2811, 3375, 3467, 109, 294, 813, 1302, 1582, 1249, 2598, 2666, 198, 604, 1089, 3542, 1510, 1538, 376, 1050, 1448, 1459, 2210, 3355, 1535, 3469, 1504, 337, 919]. Review [2144, 2261, 3438, 1270, 938, 2492, 2607, 3172, 3411, 2512, 1444, 1700, 3173, 72, 1959, 2043, 2130, 3004, 3085, 1042, 2034, 2196, 3125, 2431, 2747, 468, 1127, 1496, 2357, 226, 903, 1720, 3316, 1144, 1618, 2040, 2213, 3290, 3150, 427, 436, 1632, 1452, 2145, 1906, 2067, 2149, 2327, 2420, 2886, 3239, 1914, 1286, 3256, 2474, 883, 3045, 1446, 1841, 2538, 785, 1585, 1519, 3127, 45, 3072, 2186, 2555, 3483, 2988, 1423, 1772, 3417, 2038, 2299, 3098, 816, 2046, 3139, 14, 2011, 77, 832, 3065, 3436, 3484, 18, 648, 1375, 2813, 572, 2419, 686, 2118, 3002, 123, 2339, 1085]. Review [1417, 3123, 1565, 2655, 304, 387, 2505, 1215, 2300, 3560, 3016, 1630, 211, 257, 349, 871, 989, 1192, 1319, 1869, 1928, 1984, 2089, 2215, 2266, 2347, 2592, 3371, 91, 589, 1226, 2024, 3390, 2165, 128, 931, 1798, 28, 1197, 2340, 2588, 3035, 3180, 2425, 2997, 2006, 2344, 3338, 288, 2786, 886, 912, 2842, 3506, 2902, 2724, 120, 383, 777, 3015, 3333, 2015, 1378, 2935, 2278, 2330, 1415, 2044, 2107, 2665, 3324, 3386, 2236, 828, 2345, 292, 3119, 3343, 1846, 1937, 2234, 3455, 126, 2774, 2847, 263, 1032, 183, 603, 1353, 2985, 3457, 3186, 790, 283, 519]. Review [63, 229, 469, 956, 1033, 2994, 3093, 3289, 3325, 1293, 3233, 3117, 503, 2927, 3147, 1516, 144, 1277, 3514, 3491, 950, 3051, 565, 1172, 1349, 1739, 2584, 1025, 249, 750, 1352, 3074, 1497, 1507, 371, 1335, 2497, 2649, 2911, 3032, 296, 598, 1114, 782, 1237, 1305, 625, 414, 827, 1120, 523, 731, 2249, 2331, 908, 1326, 19, 2694, 2101, 1500, 453, 2280, 3011, 3328, 320, 974, 1624, 2485, 1386, 2267, 808, 666, 53, 3327, 3530, 2599, 3164, 2974, 552, 2426, 3554, 2254, 360, 558, 1290, 1383, 42, 1023, 1963, 88, 176, 232, 393, 659, 787]. Review [1330, 1870, 2489, 3337, 323, 2670, 2841, 1030, 275, 681, 1494, 1704, 2576, 2675, 2109, 478, 505, 1140, 1405, 1665, 2064, 2328, 448, 998, 2360, 51, 1719, 1916, 2417, 2735, 476, 758, 1225, 1304, 2875, 3106, 3458, 59, 319, 1542, 1755, 1831, 3449, 445, 1104, 1282, 1376, 1566, 1725, 764, 1759, 2761, 3099, 3144, 2583, 2809, 3507, 69, 289, 695, 1541, 2822, 280, 2756, 2676, 230, 3485, 2763, 1938, 3515, 2722, 1625, 2663, 3041, 1337, 2939, 3155, 301, 2729, 278, 2153, 792, 1515, 3470, 355, 747, 1379, 1647, 1820, 2572, 2071, 2877, 3564, 3253, 875, 942]. Review [2960, 2142, 3376, 303, 1059, 1214, 606, 2854, 83, 162, 2815, 620, 2201, 643, 730, 1068, 1094, 1380, 1918, 2161, 295, 389, 1683, 2390, 2704, 2857, 1746, 763, 1598, 2358, 2105, 3044, 3513, 2830, 177, 1037, 2609, 3363, 1136, 1868, 2856, 3092, 556, 1019, 1976, 2078, 2152, 2248, 2968, 203, 845, 1231, 1443, 3362, 1554, 1989, 2257, 2424, 2893, 3332, 522, 772, 854, 1757, 2143, 2466, 90, 660, 759, 1489, 1747, 3450, 2887, 2459, 2904, 1592, 1271, 1562, 1640, 1812, 1245, 1674, 2111, 2772, 117, 3539, 619, 712, 318, 784, 2540, 3170, 136, 1141, 1199, 1637, 776, 3037, 2499]. Review [27, 423, 760, 1670, 2132, 356, 457, 711, 1384, 1925, 2002, 2657, 3387, 3229, 2154, 141, 1259, 1754, 2014, 3548, 717, 434, 2541, 1456, 2060, 1354, 847, 3138, 2511, 3059, 3434, 2274, 3129, 742, 3342, 2980, 3377, 1309, 1663, 2938, 2518, 3268, 1915, 943, 1228, 1677, 1845, 151, 638, 922, 1010, 1732, 3412, 3137, 3286, 3036, 2807, 3444, 238, 419, 508, 657, 756, 774, 1054, 1265, 1781, 1852, 1907, 2573, 2669, 3122, 1949, 3287, 532, 1117, 1360, 824, 1729, 3510, 397, 3476, 1813, 3248, 3154, 2486, 3081, 722, 727, 2422, 2972, 2823, 3329, 192, 2473, 2543, 2667, 2375, 412]. Review [1482, 1933, 2746, 720, 640, 2045, 2379, 3102, 2090, 2255, 2444, 2517, 1348, 1465, 1631, 1991, 680, 872, 205, 2083, 2031, 3261, 1401, 1619, 3468, 2742, 2664, 3047, 186, 271, 493, 560, 1307, 1575, 1931, 2121, 2362, 2820, 2999, 81, 2784, 3040, 1002, 2415, 2239, 3584, 1428, 2025, 3393, 1175, 335, 343, 546, 689, 394, 1190, 1818, 537, 957, 2705, 2087, 3135, 667, 862, 22, 1133, 1392, 1314, 1735, 1847, 2117, 1944, 2028, 2539, 2766, 3509, 1472, 2819, 3385, 2775, 1743, 55, 191, 1803, 2288, 3341, 1551, 1717, 2245, 142, 1202, 2112, 2323, 2990, 3571, 829, 1063, 1311]. Review [1568, 2012, 3344, 964, 1404, 1020, 1285, 2265, 2435, 2656, 1741, 2097, 3431, 2615, 644, 424, 685, 1083, 481, 3126, 1767, 3067, 155, 350, 639, 1370, 1467, 1607, 2901, 1652, 2164, 3425, 2195, 1129, 1941, 1067, 1264, 2765, 692, 1372, 2550, 2940, 1040, 1495, 1902, 21, 877, 2373, 234, 467, 766, 932, 1669, 1675, 1770, 2190, 2456, 2883, 3143, 3345, 365, 2185, 2534, 3586, 2070, 746, 916, 1153, 1359, 2502, 113, 347, 1139, 2859, 262, 3439, 332, 461, 566, 2613, 3521, 668, 1862, 2009, 2439, 3251, 1327, 2728, 3277, 1987, 1236, 175, 944, 1398, 1701, 166]. Review [395, 3478, 963, 1449, 1458, 2614, 2617, 2803, 1811, 3227, 1758, 2796, 3084, 3161, 182, 2366, 2458, 3452, 1014, 1810, 1263, 3162, 921, 1642, 1815, 1919, 2176, 2296, 2414, 2603, 40, 3541, 3265, 137, 260, 888, 815, 2612, 2697, 1848, 2475, 2551, 543, 665, 2409, 1533, 3071, 1682, 1749, 3237, 1639, 422, 3095, 3283, 850, 1394, 1829, 1940, 2103, 2364, 2451, 3069, 3556, 1333, 2567, 3103, 936, 2319, 2702, 3133, 2113, 2450, 1110, 2545, 2720, 2966, 3347, 1275, 1809, 2244, 2128, 1736, 1985, 171, 3157, 1909, 1243, 1917, 421, 857, 3177, 1865, 3258, 634, 86]. Review [3532, 1260, 2263, 2391, 3018, 3034, 3570, 32, 510, 2808, 3359, 3052, 156, 26, 146, 822, 2074, 2243, 2733, 3132, 2907, 3120, 2350, 2037, 1235, 1901, 43, 652, 765, 1248, 1799, 2349, 3428, 1281, 1636, 2359, 2983, 2727, 1780, 3141, 357, 531, 2986, 3535, 415, 3552, 3391, 3446, 3474, 1545, 2568, 2908, 3263, 188, 473, 965, 1131, 2640, 3105, 1029, 75, 3046, 2706, 108, 1977, 274, 2295, 2167, 2618, 873, 966, 2855, 3100, 2585, 2776, 3017, 3487, 2469, 3575, 1826, 1389, 1836, 2688, 2298, 444, 1022, 1822, 1895, 2333, 3433]. Review [646, 1057, 1425, 1854, 541, 1052, 1478, 2976, 3526, 122, 255, 811, 1016, 1116, 1195, 1475, 2270, 2430, 2741, 47, 131, 527, 1590, 3076, 1498, 1312, 2605, 369, 1391, 2513, 1711, 2524, 1487, 2404, 1350, 1955, 859, 1150, 1403, 2238, 1121, 1855, 2832, 3567, 2672, 1823, 3238, 2276, 1039, 1995, 636, 1267, 1912, 1994, 3426, 315, 3174, 1238, 1390, 2407, 3282, 3544, 3285, 485, 1194, 1876, 588, 1203, 1573, 1689, 1979, 3351, 529, 135, 293, 682, 821, 914, 955, 1138, 1586, 1629, 1634, 1651, 1724, 1763, 1778, 1861, 1863, 1903, 2129, 2159, 2262, 2324, 2387, 2488]. Review [2770, 2810, 2844, 3187, 3231, 1672, 2126, 2208, 3550, 2740, 490, 599, 637, 729, 1009, 1092, 1658, 1710, 13, 491, 2361, 1549, 246, 1948, 1509, 3292, 1289, 2047, 2452, 2995, 3279, 1315, 2970, 483, 1088, 1814, 3569, 2932, 2174, 1950, 299, 3589, 3060, 1420, 1511, 1958, 506, 119, 125, 697, 1650, 3323, 1750, 2882, 3012, 2171, 2268, 3257, 3421, 631, 2022, 1584, 65, 368, 721, 953, 1338, 2203, 2610, 2696, 945, 375, 379, 3384, 2643, 168, 1122, 745, 2910, 165, 1246, 116, 679, 2351, 3490, 3296, 1069, 1601, 1229, 1508, 2325, 617, 1544, 597, 3291]. Review [2889, 739, 1098, 2638, 2202, 1292, 1273, 2500, 2892, 3335, 79, 786, 880, 1512, 425, 73, 2388, 2685, 428, 635, 1272, 1714, 2731, 1505, 251, 653, 1026, 1198, 1227, 2127, 2496, 2753, 3264, 245, 184, 2191, 2721, 3590, 2029, 563, 842, 2008, 1990, 1047, 1261, 3576, 1367, 31, 1268, 865, 515, 1064, 242, 2673, 1065, 3321, 2533, 2211, 2802, 3518, 282, 477, 624, 1233, 3505, 143, 1234, 1385, 1599, 1681, 1842, 2033, 2465, 3048, 1648, 340, 1361, 3430, 2699, 1471, 1892, 2168, 2831, 1207, 1760, 1827, 1211, 3191, 46, 127, 466, 661, 937, 979, 1031, 1206, 1572, 2131, 2413]. Review [2644, 3423, 2760, 1493, 2736, 2703, 2326, 204, 545, 641, 901, 1179, 1422, 538, 1180, 1628, 2730, 212, 348, 449, 3523, 2246, 114, 256, 458, 1015, 2258, 373, 630, 757, 954, 1041, 1406, 1455, 2929, 3353, 3471, 1320, 41, 3462, 1821, 2352, 2292, 946, 658, 1355, 590, 3357, 33, 732, 1381, 1643, 2003, 3250, 3435, 2279, 3480, 1728, 3055, 592, 3077, 252, 1992, 2162, 930, 1137, 698, 1426, 1877, 2971, 472, 578, 678, 1410, 1501, 1600, 1676, 1766, 1856, 1952, 2535, 2606, 2686, 2852, 2923, 3091, 3459, 1473, 1653, 1722, 2005, 2253, 3330, 2758, 2991]. Review [1416, 1503, 2878, 3280, 3366, 3574, 54, 762, 1086, 1578, 2817, 949, 1962, 3533, 1100, 1188, 581, 2039, 384, 557, 1283, 2256, 2341, 3134, 3404, 2116, 2286, 2374, 2544, 2575, 2762, 2987, 374, 518, 1266, 276, 1170, 2231, 1142, 1222, 196, 677, 2204, 2472, 1769, 2522, 2621, 586, 164, 61, 741, 1046, 1291, 2487, 2683, 3171, 2928, 254, 2259, 1278, 169, 385, 2378, 2232, 699, 794, 1875, 1181, 1797, 1145, 1329, 1867, 2041, 2925, 670, 1602, 1550, 1627, 2693, 2065, 737, 1492, 1532, 1802, 1939, 2048, 2147, 2182, 2520, 2717, 2821, 3540, 104, 354, 755]. Review [1298, 1756, 2457, 2818, 3063, 3149, 60, 1318, 2076, 23, 2315, 2992, 1926, 1126, 1301, 1303, 1858, 452, 933, 462, 753, 1154, 1240, 1393, 71, 173, 381, 3319, 1006, 1905, 1942, 265, 3572, 17, 649, 839, 929, 1531, 2814, 2900, 3070, 1850, 2297, 377, 2194, 3188, 2943, 3185, 2850, 2601, 596, 1703, 672, 1587, 2428, 3525, 416, 2403, 3465, 2648, 3519, 456, 1777, 1654, 2092, 2515, 3432, 812, 676, 761, 844, 2421, 779, 999, 1090]. Reviews [225, 1308]. Revised [2233]. Revision [297]. Revisited [616, 454, 3172, 1167]. Revista [526, 524, 523, 525]. Revistas [2491]. Revival [952]. Revived [548]. revocation [3340]. revolución [752, 352, 829, 2434]. revolutie [361]. Revolution [1751, 367, 2942, 3116, 632, 3315, 2871, 391, 2275, 353, 775, 96, 1196, 3443, 1932, 1727, 44, 2838, 35, 45, 2857, 2002, 2657, 774, 1117, 394, 1404, 461, 652, 3552, 2568, 3526, 2387, 3292, 1530, 3062, 1023, 561, 749, 212, 2570, 3549, 455, 3349, 1464, 2169, 710, 153, 3156, 2774, 3093, 3530, 1290, 776, 3439, 1100, 138, 420, 814, 1282, 2485]. Revolutionary [843, 751, 3024, 2671, 2288, 2012, 1067, 3066]. Revolutionen [1932]. Revolutionized [2950, 1682]. Revolutions [1603, 2288, 624, 1356, 420, 1373, 1512]. Revue [516, 522, 522]. Rewriting [2143]. Rex [620, 2933]. Rey [1374]. reyes [1654]. Reynolds [3513, 2935, 3513]. Rezeption [2920, 663, 3241, 2679, 2894]. Rezeptionsgeschichte [1274]. Reznek [3236]. Rheinberger [3010]. Rheticus [895]. Rhetoric [2611, 1617, 1479, 924, 1968, 3484, 808, 1549, 1436]. Rhoda [3141]. Rhodes [1409, 225]. Rhodesia [1229]. Rhythms [1003]. Ribalka [2352]. Ribera [134]. Ricci [2963, 1560]. Rice [961]. Ricerche [2980, 1722]. ricezione [1821]. Richard [3096, 1594, 1295, 690, 1734, 1678, 2833, 2303, 2381, 1080, 767, 1716, 1731, 2275, 723, 465, 3108, 582, 1407, 3448, 1547, 947, 110, 1800, 1460, 1469, 1922, 1832, 2020, 2687, 2163, 1247, 479, 3528, 1645, 1409, 366, 1112, 2959, 1011, 1087, 1574, 551, 938, 1632, 1630, 956, 1277, 2584, 1114, 681, 2328, 2729, 27, 3036, 3261, 2239, 667, 3478, 3095, 850, 1394, 2295, 225, 1016, 1629, 1903, 721, 2352, 732, 3459, 3533, 2544, 3171, 649, 2624, 1254]. Richards [1232, 1708, 268, 1489, 1952]. Richardson [595, 1104, 1876]. Riche [1808]. Richet [2374]. Richmond [1748, 2449, 1748, 2987]. Richter [753]. Rickards [1458]. Ricklin [2583]. Rico [2681]. Ridder [1812, 1266]. Ridder-Symoens [1812]. Riddle [3232, 2156]. Rider [1795, 841]. Riders [579]. Ridge [2039]. Ridges [1502]. Rieber [1673]. Riehl [2853]. Rienk [1297]. Riet [2331, 727]. riforma [1199]. Rigdon [287]. Right [1134, 284, 1506, 1703]. Righting [626]. Rights [381]. Rikoon [336]. Riley [1126, 2830]. Rima [2936, 1175]. Rimby [3292]. Rime [1097]. Rims [1502]. Rinascimento [2586, 1643]. Ringer [1757, 3391]. Ringvorlesung [578]. Rio [1311]. Riordan [3562]. Riot [1662]. Rip [1138]. Ripa [1037, 212]. Riper [1940]. Risala [1718]. Rise [1528, 928, 66, 579, 3562, 2641, 3522, 1479, 2193, 1399, 671, 2042, 1038, 612, 441, 1593, 2876, 3316, 1798, 414, 1030, 280, 1647, 1265, 412, 1933, 1133, 665, 3351, 739, 73, 590, 54, 949, 586, 2076, 173, 2675]. Risk [3176, 1704, 760]. Risky [3176]. Ristampa [2901]. Rita [2766, 2265, 761]. Ritchey [1763]. Rites [2855]. riti [1643]. Ritter [2992]. Ritvo [854]. Rival [2950]. Rivalries [967]. Rivalry [1952]. Rivals [967]. River [2912, 1834]. Rivers [1041, 2361]. Rivinus [1667]. Rivista [521, 521]. Rivka [3437]. rivoluzione [973, 2001]. Road [225]. Roads [2578]. Rob [2814]. Robbins [2826]. Robert [25, 30, 3019, 1147, 1213, 3418, 3502, 734, 3182, 858, 1673, 3299, 2549, 2392, 1152, 3577, 951, 1702, 1998, 2773, 2209, 207, 2170, 1889, 2899, 1111, 1680, 2380, 2119, 110, 645, 1871, 3086, 683, 2898, 331, 1021, 3536, 2738, 1358, 193, 3124, 2460, 1844, 2690, 1872, 2017, 174, 56, 1514, 3228, 2192, 1108, 1947, 2961, 2780, 1363, 2881, 1097, 2369, 459, 294, 1906, 3098, 832, 304, 349, 3180, 1277, 827, 2576, 448, 1542, 2809, 3507, 2854, 2815, 295, 2248, 1489, 1754, 847, 1781, 2473, 2117, 2766, 1083, 3067, 2940, 21, 3251, 1987, 3265, 3359, 146, 43, 188]. Robert [1131, 108, 2469, 646, 1478, 1876, 682, 1511, 379, 79, 1227, 1234, 2644, 590, 2003, 1100, 2116, 794, 1926, 3070, 672, 310, 1582, 647, 3058, 1664, 773, 2523, 469, 644, 641]. Roberta [3343]. Roberto [978]. Roberts [1066, 1688, 1539, 1134, 177, 3231]. Robertson [1087, 1586]. Roberval [3056]. Robin [1795, 1081, 37, 841, 2422, 3283, 3105, 635, 3435]. Robinet [161]. Robins [1418]. Robinson [773, 2196, 1845, 3319]. Roca [1331, 1917]. Roche [522]. Rochester [965]. Rocke [2280, 1751]. Rockefeller [579, 1593, 3283]. Rocket [73]. Rocks [2201]. Rocky [2989, 963]. Roderick [2521, 847, 1133, 3423, 381, 2956]. Rodger [3094, 942]. Rodis [3339, 1476]. Rodis-Lewis [3339, 1476]. Rodner [3349]. Rodney [198]. Rodnite [1715]. Rodriguez [1917, 523]. Roell [2680]. Roero [2270]. Rogal [1249]. Roger [1649, 1579, 2787, 3488, 353, 2173, 1557, 2125, 2507, 1978, 3536, 3254, 583, 436, 2886, 1585, 128, 3325, 3337, 2841, 2735, 1282, 522, 141, 434, 2074, 2727, 225, 706, 491, 3335, 1531, 651, 2137, 2841, 3335]. Rogers [2186, 1507, 2657]. Roi [651]. rol [994]. Rolando [791]. Roldan [2408]. Role [194, 3130, 3495, 387, 1380, 1309, 1979, 679, 1032]. Roles [988]. Rolf [517, 1705, 89, 1337, 518]. Rolle [2300]. Roller [1890]. ROM [2908, 206]. romain [3427]. romaine [2330]. Romains [427]. Roman [250, 1003, 426, 725, 2499, 2168, 2138, 1714]. Romance [2299, 369, 2362]. Romani [3332]. Romano [2260]. Romantic [3349, 1080, 1739, 3539, 1125]. Romanticism [2159, 754, 169]. Romantik [2754, 2754, 1122]. romantisch [2101]. romantisch-speculatief [2101]. Rome [1467, 1815]. Romero [746]. romischen [989]. Romm [1357]. Ron [1666, 1331, 2851, 1428, 1311]. Ronald [1147, 1617, 1595, 106, 2738, 1872, 1087, 814, 1097, 606, 2609, 318, 2540, 1590, 3550, 375, 1318, 3269]. Roncaglia [3054]. Rond [2093, 1535, 2101]. Roney [1009]. Ronner [2799]. Room [764]. Rooms [2792]. Roots [3295, 2269, 3417, 1348, 1969, 65, 2116, 3525]. Rootselaar [2983]. Rosa [3475, 3176]. Rosalind [666]. Rosalynd [862]. Rosario [3125]. Rosarium [1721]. Rose [2170, 1878, 3563, 2466, 2883, 2047]. Rose-Mary [2170]. Rosell [1331]. Rosemary [592]. Rosen [1373]. Rosenberg [1947, 2369, 1458, 2321, 1516, 872]. Rosenfeld [3372]. Rosenkreuzer [2342]. Rosenmeyer [634]. Rosenthal [2804]. Roser [432]. Roshdi [749, 150, 522, 1809, 510]. Rosholt [304]. Roslynn [2729, 2409]. Rosner [1231, 3099, 1482]. Ross [121, 3269, 3348, 1400]. Rossell [1917]. Rosser [2492]. Rossetti [3139]. Rossi [122, 1195]. Rossiter [2467, 1264]. Rostenberg [646]. Rostock [3551]. Rota [3000]. Rotation [3036]. Rothenberg [1573, 1877]. Rothfield [1576]. Rothman [1838, 3578, 1050, 1502]. Rothra [2455]. Rotterdam [994]. Rouaze [825]. Roudinesco [1054]. Rougle [203]. Rouille [3538]. Roundabout [965]. Rounds [68]. Rouse [2704]. Rousseau [43, 3516]. Rowan [1859]. Rowe [2282, 1403]. Roy [2725, 241, 3516, 420, 1513, 1896, 2096, 1118, 1027, 2412, 1149, 1727, 2668, 1824, 572, 1869, 335, 332, 815, 43, 3046, 131, 65, 169, 2048]. Royal [1626, 601, 835, 927, 1193, 378, 1666, 209, 618, 249, 1458, 3428, 2452, 697, 2171, 745, 1381, 557, 1695, 3497, 1174]. Royale [1649]. Royer [2986]. Rozental [3564]. Rozhanskaya [1810]. Ru [2093]. Rubber [781]. Rubbia [771]. Rubin [3082]. Rubinger [465]. Rubinin [1217]. Rudbeck [2007]. Ruder [650, 930]. Ruderman [156]. Rudi [1902]. Rudolf [939, 345, 1115, 1635, 2303, 940, 2965, 1215, 2832, 2131]. Rudolph [2313, 2256]. Rudwick [3354, 1837]. Ruestow [2745]. Ruetsche [2117]. Ruey [1169]. Ruffini [3350]. Ruggiero [564]. Ruggles [144]. Ruhla [1804]. Ruiz [2272]. Rule [2786]. Ruled [1030]. Rules [2547]. Rumm [1573]. Run [2616]. Rundfunkindustrie [3385]. Runessansu [902]. Runge [654, 175]. Running [1046]. Rupert [1731, 926, 111, 1929]. Rupescissa [730]. Rupke [576]. Rural [480]. Ruse [2916]. Rush [2622]. Rushd [1276]. Rushton [2015]. Russel [1943, 803]. Russell [68, 3163, 207, 3065, 1141, 213, 2416, 2006, 2822, 1260]. Russett [370]. Russia [2677, 3582, 3287, 1803, 2126, 2288, 1404, 957]. Russian [2549, 1217, 2288, 1170, 2231]. Russo [2320]. Rust [297]. Ruth [3302, 2531, 856, 1709, 625, 3171, 779, 201]. Rutherford [2348]. Ruza [1245]. Ryan [3348, 1802]. Rydell [1844].

S [30, 2695, 264, 1603, 2120, 848, 554, 105, 2301, 1673, 2788, 1132, 3516, 2114, 2768, 2942, 570, 744, 632, 3349, 2800, 1731, 2582, 1091, 3165, 3354, 1878, 24, 74, 2085, 331, 905, 1552, 199, 235, 1837, 3068, 2835, 382, 2964, 863, 1418, 1174, 3101, 464, 2783, 386, 1357, 1752, 1087, 1768, 3456, 3190, 1843, 733, 2668, 1621, 436, 2339, 2935, 360, 2109, 448, 3106, 1566, 2956, 3513, 2540, 2624, 1456, 1309, 1228, 508, 727, 2422, 2990, 3439, 2176, 543, 3283, 857, 3034, 32, 43, 2706, 108, 2167, 1549, 2882, 379, 745, 786, 2731, 3590, 1990, 1031, 2292]. S [1503, 1100, 1298, 1154, 649, 2943, 488]. S.J [837, 1819]. S.J. [650]. Sa [1496, 3340]. Sabatier [1671]. Saber [183]. Sabetai [1634]. Sabina [434, 2889]. Sabine [1660, 1855]. Sabra [728, 2151]. Sachiko [2427]. Sachliteratur [2424]. Sachs [2018, 1462, 2498, 278, 248]. Sachs-Hombach [2018]. Sachse [654]. Sachsen [439]. Sacralizing [553]. Sacred [3334, 1922, 1805]. Sadi [174]. Saenger [599]. Saenz [2982]. Safa [1717]. safar [1718]. Safety [2993, 2213, 2993]. Saggi [2756, 1195]. Saggio [973, 1915, 2098]. Sahl [1809, 219]. Sahotra [1674]. Said [1551]. Sail [3345]. Saint [435, 2011, 435]. Saint-Victor [435, 435]. Sal [1770]. Sala [2255]. Salam [191, 191]. Saldaña [321, 321, 526]. Salem [1551, 2496]. Salernitana [1183]. Saliba [2245, 32]. Sallares [1363]. Sally [3317, 302, 3037, 722, 1203]. Salmon [783, 783, 868]. Salomon [3060]. Salon [2307]. Salt [1625]. salubrita [2980]. Salud [868, 1024, 1690]. Salvation [544]. Salvatore [2098, 450]. Salzman [1003]. Samar [3525]. Samfundets [514]. Samh [2877]. Sammlungen [1941]. Sammons [886]. Samso [2334, 1633, 2072]. Samuel [1662, 2891, 2739, 1464, 62, 2924, 3030, 1249, 1263, 1867, 2343, 849, 936]. San [3104, 2947]. sanationum [2587]. Sanatoriumstherapie [374]. Sanches [554]. Sánchez [1331, 1311, 2851, 262]. Sánchez-Ron [2851]. Sandemanian [1532]. Sander [2493, 3038]. Sanders [1975]. Sandia [581]. Sandmann [758]. Sandor [2032, 2604]. Sandra [1900, 1062]. Sandro [860, 2118]. Sane [3098]. Sanford [336, 3380]. Sang [507]. Sang-yong [507]. Sangallo [2080]. Sanger [2540]. sangre [1654]. Sangwan [1543]. sanidad [3445, 261]. sanitaire [1897]. Sanitarians [716]. sanitarias [752]. Sanitatis [2251, 2070]. Sano [2537]. Sanskrit [3490]. Santa [1614]. Santalla [3431]. Santcliment [3523]. Santiago [1760]. Santinello [2063]. Santing [2335]. Santos [2915]. Sant’Uffizio [2445]. sapere [1471]. Sapir [777]. Sapolsky [874]. Sapp [864]. Sara [3119]. Sarah [1108, 2675, 3428]. Sarasohn [2659, 1958]. Sardar [640]. Sarepta [731]. Sareshel [343]. Sarewitz [3280]. Sargent [2170]. Sarjeant [2788]. Sarkar [1674]. Sarmiento [261]. Sarnowsky [439]. Saros [1462]. Sarton [1261, 501]. Sass [3419]. Sassower [792, 2475]. Satellite [969, 85, 387, 1582]. satellites [1728]. satellitibus [1728]. Satpal [1543]. Satzinger [3375]. saubere [2281]. Sauer [2117]. Sauerstoffaktivierung [3570]. Sauglingsfursorge [2914]. Saul [566, 2696]. Saussure [2178, 392, 1200, 1198]. Savage [1490, 3585, 346]. Savage-Smith [346]. Savant [3072, 3460]. savante [1982]. savants [561, 322, 900, 14]. Save [947]. Saverio [1560]. Saving [477]. Savior [226]. Savitt [375, 375]. savoir [320, 2960]. savoirs [1063]. Sawai [2599]. Sawday [2737]. Sawyer [3096]. Saxon [1992]. Saxonicus [1113]. Say [1483]. Sayili [2483]. scala [2095]. Scale [1328]. Scaling [2106]. Scandal [1147]. Scandinavian [513]. Scene [3153]. Scenes [1058, 1837]. Scepticism [1961]. Schaaf [1491, 2676].

Schabas [1030]. Schaffer [108, 417, 647, 1488, 2039]. Schamp [146]. Scharff [2473]. Schatzberg [3480]. Schaxel [476]. Schechner [3119]. Schedule [3176]. Schedvin [580]. Scheffer [1140]. scheikunde [1535]. Schelling [2754, 167, 2356, 656]. Schellings [2754, 2356]. Schemmel [348]. Schepers [761]. schetnykh [910]. Schets [41]. Schiappa [1549]. Schichten [358]. Schicksal [461]. Schiebinger [446, 2513]. Schiffer [2037]. Schimmel [1622]. Schivelbusch [371]. Schlagel [2959]. Schlagintweit [1337]. Schleiden [654, 656, 1393, 1393]. Schleifer [1872]. Schleiner [2505]. Schleissner [2504]. Schleucher [365]. Schling [266]. Schling-Brodersen [266]. Schlor [740]. Schloss [1724]. Schlossberger [266]. Schlumberger [2616]. Schlupmann [239]. Schlüter [851]. Schmaus [2527]. Schmeidler [442]. Schmidt [349, 3467, 1375, 2162]. Schmidt-Biggemann [1375]. Schmidt-Nielsen [3467]. Schmiedebach [2523]. Schmitt [636, 970]. Schmitz [1680]. Schmode [2747]. Schneewind [766]. Schneider [1960, 1185, 1932, 1504, 558]. Schobinger [2509, 830]. Schofield [3342]. Schola [1183]. Scholar [85]. Scholars [2002]. Scholarship [923, 714]. Scholastic [1009, 2076]. Scholasticism [3247]. Scholastische [2153, 439]. Scholz [3281, 536, 1841]. Schoner [2254]. School [1218, 360, 301, 704, 2288, 3163, 1630, 360]. Schooling [217, 440]. Schools [2055, 1977, 392, 10, 596]. Schriften [2356, 2920, 654, 463, 3137, 929, 1531, 1393, 2843, 549, 1463]. Schriftenreihe [517]. Schrist [2973]. Schroder [1142]. Schrödinger [2463, 59, 586, 950]. Schroeer [225]. Schubring [1304]. Schuhflicker [3441]. Schul [453]. Schule [536]. Schull [2925]. Schulmann [2460, 2117, 2766]. Schultze [654]. Schumacher [1337]. Schumpeter [3177]. Schutt [1398]. Schwann [654]. Schwartz [2253]. Schwartzberg [1452]. Schwartzman [1230]. Schwarz [2252]. Schweber [2301]. Schweizer [2923]. Schwerionenforschung [2462]. Schwinger [2301]. Sci [696]. Sci-Tech [696]. Science [3089, 513, 819, 2329, 2546, 2453, 222, 502, 1079, 3199, 3303, 1437, 3317, 687, 3334, 418, 1061, 3249, 2527, 902, 2384, 1614, 3204, 2189, 1077, 3182, 580, 170, 455, 1751, 221, 2549, 623, 605, 3310, 1617, 650, 1873, 2062, 951, 107, 420, 928, 3205, 1424, 516, 2311, 2565, 577, 985, 447, 2104, 2429, 2806, 632, 1776, 15, 3050, 2381, 287, 616, 1080, 2937, 575, 748, 1451, 1605, 2718, 2751, 1230, 807, 2094, 2922, 1171, 300, 1972, 3169, 3206, 2510, 1205, 1954, 2406, 1334, 3410, 370, 2230, 669, 1900, 2641, 1342, 2269, 2106, 1513, 1832, 2020, 2748, 1464]. Science [408, 106, 2273, 1517, 1280, 3406, 3207, 1479, 2719, 843, 991, 2869, 2917, 1688, 3209, 897, 2389, 1786, 220, 1436, 1569, 3168, 542, 1221, 3158, 167, 2637, 1242, 2851, 3211, 3212, 1187, 3124, 3309, 3213, 153, 2690, 3222, 489, 227, 3005, 3214, 1169, 2542, 3559, 484, 2056, 3215, 2701, 3394, 1773, 2536, 1616, 3504, 1961, 923, 3288, 2767, 2652, 454, 2151, 507, 2049, 3064, 206, 3561, 1968, 103, 626, 1343, 1062, 1328, 2317, 179, 217, 3090, 1186, 363, 878, 1279, 2073, 1244, 604, 3355, 2868, 1959, 2043, 3004, 903, 1144, 3290, 3113, 1075, 1286, 3072, 2555, 1772]. Science [816, 2011, 77, 123, 2873, 1085, 3016, 1226, 1798, 886, 1415, 1693, 1516, 1739, 3032, 3011, 100, 1624, 808, 1330, 1704, 2735, 3458, 1755, 1566, 695, 230, 3485, 896, 792, 2142, 3376, 1, 401, 799, 1159, 1524, 1883, 2221, 2560, 2865, 3196, 2390, 2704, 3363, 772, 2710, 1245, 2657, 717, 1166, 1309, 2627, 419, 774, 722, 2667, 2045, 1074, 2083, 1258, 957, 1695, 2705, 1944, 964, 1083, 3126, 1767, 2765, 877, 1669, 2950, 2009, 3496, 944, 2796, 2458, 3452, 1642, 1919, 2176, 2612, 1940, 3069, 2702, 2720, 1275, 2481, 171, 421, 3177, 1165, 32, 156, 2947, 765, 3428]. Science [1636, 415, 2706, 2167, 3017, 1253, 2976, 47, 131, 1823, 2276, 3174, 1203, 1163, 1979, 682, 1861, 3398, 500, 3219, 3220, 1948, 1289, 2452, 1315, 2714, 3569, 3589, 1511, 2882, 2022, 1584, 168, 745, 3291, 2721, 2307, 515, 624, 3505, 1892, 2168, 314, 1455, 590, 732, 2003, 1426, 1877, 1676, 894, 1416, 3280, 611, 1578, 949, 3533, 557, 1797, 1532, 3114, 406, 3185, 672, 2019, 718, 2099, 1201, 973, 1296, 1546, 925, 1571, 2573, 3556, 1312, 513, 1147, 1543, 1491, 542, 3210, 1945, 808, 872, 510, 506, 375, 2202, 1703, 3104, 2684, 3039, 1981, 709, 2552]. Science [84, 2692, 1603, 105, 567, 2752, 3097, 913, 3116, 810, 579, 80, 909, 791, 1547, 972, 934, 2969, 509, 2525, 1151, 884, 2553, 2007, 879, 446, 688, 1745, 1805, 2964, 3407, 1400, 976, 3121, 2771, 995, 2448, 2858, 1082, 814, 2446, 1239, 1935, 2207, 2921, 2066, 109, 3172, 1700, 3085, 1127, 226, 3316, 1423, 3484, 3560, 2044, 3343, 3186, 3051, 1349, 1335, 1305, 478, 1225, 1282, 1059, 295, 2857, 3362, 784, 742, 508, 640, 3135, 55, 2323, 2550, 2940, 3251, 3227, 2697, 1682, 1333, 2391, 2727, 2986, 3105, 1057, 1425, 1854, 1052, 2513, 506, 1273, 79]. Science [635, 1179, 3574, 104, 2076, 265, 377, 2601, 416, 1874, 709, 3408, 2284, 2368, 2180, 2433, 1583, 2150, 874, 1152, 1543, 1776, 2216, 1080, 992, 200, 1045, 208, 1480, 876, 300, 1460, 972, 2748, 3368, 788, 1347, 413, 2389, 2926, 3184, 3318, 1633, 2616, 618, 2365, 329, 576, 2410, 2536, 1980, 1921, 1574, 2881, 3313, 2184, 980, 277, 2682, 1204, 3403, 3125, 2357, 2038, 128, 3074, 1870, 319, 1542, 1515, 3539, 318, 3229, 1309, 2669, 2422, 2972, 2746, 720, 2819, 1803, 2288, 1551, 685, 1264, 3347, 1901, 108, 1854, 1955, 1958, 2910, 1229, 46, 1628, 1355, 678, 1769]. Science [2520, 3540, 2315, 3319]. science-monde [718]. science.’ [2348]. Sciences [1626, 3104, 130, 2453, 1891, 1734, 1617, 516, 835, 3073, 1579, 2273, 1247, 3536, 3213, 489, 618, 900, 3557, 507, 831, 1279, 2357, 816, 2266, 2143, 117, 1670, 2746, 1695, 1167, 1067, 1139, 888, 2319, 510, 531, 2706, 3590, 406, 2943, 236, 3020, 3088, 3075, 1496, 657, 3122, 1026, 1268, 1875, 2182, 121, 110, 522, 3019, 754, 328, 115, 2574, 927, 121, 3553, 417, 3148, 1567, 112, 2723, 235, 1899, 1058, 1709, 693, 489, 1514, 3331, 1506, 2077, 785, 91, 503, 505, 854, 2319, 3391, 2618, 2047, 3257, 1031, 2320, 2667, 3446, 1396, 749, 229]. Scientia [1460, 527]. Scientiarum [507]. Scientiarumque [523]. Scientific [248, 528, 2931, 1336, 1891, 2086, 1857, 498, 1603, 2570, 2294, 3262, 713, 882, 2681, 2309, 580, 1594, 770, 1262, 3543, 1617, 2834, 2942, 3116, 3306, 970, 1212, 3050, 2958, 2108, 575, 3315, 2871, 694, 2275, 3297, 1230, 1604, 885, 783, 1771, 2967, 1464, 691, 775, 881, 977, 843, 474, 905, 1786, 1436, 2169, 1196, 968, 1744, 326, 1936, 2178, 2785, 2023, 664, 1872, 227, 2959, 1345, 1887, 2941, 3293, 1727, 684, 3461, 35, 1621, 1959, 3072, 1970, 871, 1192, 1693, 249, 808, 1971, 3530, 1290, 2489, 1376, 3037, 356, 847, 1133, 2288, 2397, 685, 692, 2534]. Scientific [1848, 2396, 2720, 1799, 2568, 646, 3526, 3076, 1487, 1876, 1689, 2770, 3589, 1511, 697, 2171, 953, 895, 2533, 1031, 3471, 1100, 1797, 1550, 1802, 2520, 3540, 3188, 1309, 2552, 1427, 783, 2569, 1056, 3260, 3422, 790, 1683, 2551, 1057, 2696, 1952, 276, 416]. scientifica [124, 3457, 2001]. Scientification [2105]. scientifiche [919]. scientifico [645]. Scientifics [1749]. scientifique [2514, 3189, 2166, 1063, 1312, 653, 3062, 2099]. Scientifiques [825, 718]. Scientism [1151]. Scientist [1617, 465, 3270, 2175, 3066, 1945, 3090, 2201, 2409, 379, 1503, 1532, 1617, 738, 3467, 1639, 3351, 847]. scientista [3545]. Scientists [180, 951, 744, 927, 2958, 3488, 1983, 2619, 64, 3559, 684, 2467, 3190, 1968, 686, 1353, 1133, 1067, 225, 1052, 1391, 3238, 1584, 3487, 2048, 3121, 2607, 3018]. Scienza [454, 2000, 2260, 2590, 2277, 771, 564, 1107, 2763, 122, 521, 3332]. Scienze [521, 1194, 2526, 1101, 2097, 251, 124]. scienziato [2097]. Scipionis [1274]. scitur [554]. scolastica [3054, 638]. Scope [732]. Scopes [3363]. Scot [1913]. Scotland [697, 697]. Scott [976, 1224, 2489, 1979, 1272]. Scottish [1119, 757, 258, 1746]. Scottus [726]. Scramble [715]. Screening [1437, 2777]. Scriba [929]. Scribal [2563]. Scribano [645]. Scribner [2634]. Scriptures [1922]. Scritti [2756, 1473, 2744]. Scull [3358, 855, 1743]. Sculptors [2420]. scuola [1925]. Scurvy [2179]. SDI [1584]. Se [1712]. Sea [3284, 1461, 3142, 3121, 3345, 1682, 1749, 961]. Seaborg [199, 1768, 2540]. Seafaring [3333]. Seaman [562]. Seamans [2854, 2854]. Search [16, 1256, 1310, 3234, 534, 2198, 1835, 952, 1674, 2928]. Searle [1684]. Seas [2442]. Season [1255]. Seasons [2452]. Seattle [500]. Sebastien [1351]. Sebestik [1486]. sec. [1912]. Seccion [1690]. Seck [1474]. secoli [2963, 638]. secolo [2890, 2590, 1183, 2270, 1498, 3054, 1737]. Second [167, 3559, 1087, 657, 2612, 1530]. Second Millenium.Femaleman [3084]. seconda [1737, 1722]. secondaire [2182, 3298]. Secondary [1781]. seconde [2812]. Secord [48, 2017]. Secret [2461, 2474, 3238, 3509]. secretis [1715]. Secretos [2158]. Secrets [1715, 3508, 1921, 1094, 2795, 2889]. Section [3209]. Sector [1038]. Secular [1983, 553]. Secularisering [1297]. secundus [1634]. Securing [670]. Security [843, 894]. Sedley [1906]. Seductions [3357]. See [3007, 3266]. Seeds [1417, 715]. Seeff [1656]. Seeing [2262, 2525]. Seeliger [374]. Segal [817, 2061]. Segala [1589]. Segel [441]. segno [2590]. Segre [1018]. Sei [1195]. Seicento [645]. Seidel [1131]. Seifer [3271]. seiki [902]. Seils [3149]. Sein [736, 3091]. seine [2994]. seiner [3080, 3551, 2016, 163, 758, 1941, 3149]. Seismology [2869]. seit [1599]. Seite [2107]. Seitz [3479, 2853]. seizieme [2812]. Seki [2436]. Sektion [3367]. Sekundarliteratur [1393]. Selbstdenker [3149]. Selbsterfahrung [2827]. Selbstorganisation [654]. Selbstverständnisses [297]. Selected [2768, 2322, 1893, 2542, 1214, 191, 225, 2208, 2202, 2685, 3543, 1675]. Selection [1132, 1672, 2567]. Selections [1460, 3064, 1617]. Selective [1877]. Self [703, 2904, 1758, 2928, 2902, 2065]. Self-Definition [703]. Self-Generation [2902]. Self-Motion [2065]. Self-Understanding [2928]. Selin [1776, 3394]. Sell [1414]. Sellars [3095]. Sellers [3097]. Selles [2645, 563]. Selling [2130]. selon [3157]. seltener [1762]. Semaan [1551]. Semantic [691, 1601]. Semantics [1308]. Semantische [1637]. semantischen [2153]. Semesters [1134]. Seminar [1210, 3221]. Seminars [2463]. Semmelweis [2108]. Semsel [606]. Semyon [17]. Sen [1309]. Sendivogius [2663]. Senecan [634]. Senior [1519]. sens [729]. Sense [3182, 258, 1005, 1961, 282, 753]. Senses [1896]. Sensibility [3344]. Sensing [85]. Sentences [1279]. Sentiment [3353]. Sentinel [1860]. Sepper [2591, 2594]. Septem [1375]. September [3302, 2313, 3311, 101, 2625, 1120, 2956, 1973]. Sequard [1745]. Sequel [27]. Seraphic [1612]. Serbat [3132]. Serena [395]. Sergeant [2087]. Sergei [2225]. Series [3534, 304, 3120, 2262]. Serjeant [1275]. Serrai [1284]. Service [3295, 1497, 2670, 1909, 594]. Services [2674, 1778]. Servos [772]. Servus [1277]. Ses [2166, 3298, 2592, 3144, 1259, 729]. Set [486, 1675]. settecentesco [2097]. Settecento [1194]. Setting [1978, 1739]. Seurat [903]. Seven [2286]. Seventeenth [2896, 1904, 744, 927, 748, 2578, 1191, 492, 1427, 2230, 353, 2662, 2085, 600, 2088, 3136, 2577, 2588, 2006, 2344, 695, 1562, 3387, 3059, 2087, 1285, 2349, 3526, 745, 2341, 407, 1165, 3398, 3400]. Seventeenth-Century [2896, 927, 492, 2230, 2085, 2088, 3136, 2006, 2344, 1562, 3387, 2087, 1285, 2349, 3526, 745, 3398]. Seventh [2736]. Seventh-Century [2736]. severnoi [887]. Sewers [3326]. Sex [446, 284, 2038, 1445, 652, 1799, 1636, 2546]. sexe [3427]. Sexes [2846]. Sexton [1417]. Sexual [2314, 2782, 1619, 171, 1672, 43, 623, 370]. Sexuality [43, 148]. Seymour [2856, 1554]. Seyyed [1805, 1717]. Sezgin [510]. sfere [1102]. Shackelford [2062]. Shadow [1838, 2446, 2927]. Shadows [1581, 1491, 3348]. shakaishi [902]. Shakespeare [1565, 733]. Shakir [1005]. Shamans [2968]. Shaner [980]. Shanks [2937]. Shape [1786, 2228, 1202, 893]. Shapes [1469, 2469, 2882]. Shapin [3229]. Shaping [2468, 2745, 805, 1497, 1140, 3044, 580, 3032]. Shapiro [3328, 1382, 2266, 2267, 350]. Shaposhnikova [3369]. Sharma [2599, 1907]. Sharon [3230, 1759, 192]. Sharov [2120]. Sharp [185]. Sharples [1906, 2243]. Sharratt [2005]. Shattered [1036]. Shatzmiller [1466, 2423]. Shaw [2021]. Shea [1371, 420, 1192]. Shearer [2706, 2943, 3287, 2706, 2943]. Sheehan [1450, 2372, 904, 2643]. Sheets [3302, 179, 2601]. Sheets-Pyenson [3302, 179, 2601]. Sheffield [2739]. Sheila [2389, 1838, 598, 1330, 1281]. Shelby [1665]. Sheldon [3124, 2210, 1265]. Shelford [1687, 1227]. Shell [3063]. Shem [433]. Shepherd [1320]. Sheridan [974]. Sheriff [2912]. Sherman [3057, 2337]. Sherrard [1908, 1908]. Sherwin [334]. Shield [225, 3296]. Shierry [489]. Shifting [775]. Shigenori [980]. Shigeru [1955]. Shionoya [3177]. Ship [1682]. Shipbuilder [1645]. Shock [1687]. Shocked [3471]. Shocks [2953]. Shoenberg [1217]. Shop [1494]. Shore [3123]. Short [451, 72, 1959, 1803, 1086, 26]. Shorter [1943, 2974, 1413, 290, 3235].

Shorthand [452]. Shortland [37, 397]. Should [2044]. Shoulders [2163, 839]. Shreve [282]. Shteir [2751]. Shuchat [17]. Shukri [1006]. Shumaker [732]. Shuman [3176]. Shumway [485]. Shuttle [74, 3176]. Shuttleworth [722]. Siberian [3172]. sich [2978]. Sicile [914]. Sick [2096]. Sickness [1027, 1508, 592]. Sickroom [2524]. Sida [1412]. Sidereal [831]. sidereo [1728]. Sidereus [831, 1653]. Sidney [1428, 2566, 1973]. Siècle [2189, 3026, 535, 460, 1809, 1390, 2070, 856, 2812, 322, 1093, 2432, 3075, 1023, 1729, 1931, 2820, 2539, 916, 2333, 256, 2900]. siècles [2247, 1026, 1296]. Siege [1404]. Siegel [1306]. Siegen [578]. Sieges [2422]. Siegfried [2462, 459]. Sieglinde [2503]. Siegwart [259]. Siekevitz [968]. Siemens [1842]. Sierra [2824]. Sievers [2902]. Siger [1368]. Sigerist [3012]. Siglo [1723, 3475, 1828, 1354, 3154, 2444, 2435, 262]. siglos [1470, 2803]. Sigmund [2032, 2604, 1321, 288, 943, 1677]. Sign [3140, 2845]. Signal [2826]. signaturis [3531, 2901]. Significance [1080, 497]. Signs [1576, 2845, 1726]. Sigrid [2914]. Sigrist [1201]. Sihha [2070]. Sijzi [3242, 3242]. Silard [1581]. Silbergleit [2835]. Silberstein [943]. Silence [2323]. Silent [1949]. Silicon [3419, 3479]. Silicosis [1231]. Silk [2578]. Silliman [568, 498]. Silurian [48]. Silvan [2301]. Silver [3407]. Silverberg [3452]. Silverman [2644]. Silvia [3245, 2270]. Silvio [567, 1019]. Sime [2531]. Simek [1635]. Simians [955]. Simitis [288]. Simmons [2958, 1905]. Simon [740, 1617, 417, 1230, 647, 3529, 3440, 1488, 2701, 626, 816, 1010, 2807, 108, 2333, 2778, 821, 3251, 2366]. Simone [2331]. Simonton [416]. Simplest [910]. simplicium [1637]. Simplicius [1362, 915]. Simpson [697, 2821, 2455]. Simulating [1605]. Sin [3373, 2537]. Sin-itiro [3373, 2537]. Sina [1717]. Sinaceur [1042]. Sinan [2877]. since [2214, 2695, 1171, 3326, 2779, 1502, 814, 1585, 3162, 3018, 188, 3590, 1416, 23]. Sinding [1412]. sine [2089]. Singh [2599, 3529]. Singular [1697]. Sink [3234]. Sinkankas [1780]. Sinsheimer [2690]. Sir [1617, 3270, 1417, 847, 3541, 3078, 2345, 393, 1291]. Siraisi [735, 1687]. sistem [910]. sistema [1199]. sistemas [2741]. Sister [1449]. Sisters [3428]. Site [1502]. Sites [3334]. Sitter [671]. Situating [3186, 2481]. situation [3538]. sive [438, 3248]. Sivin [129]. Six [2809, 225, 2247]. Sixteenth [2653, 1998, 600, 35, 448, 156, 3430, 2341, 1888]. Sixteenth-Century [448, 156, 3430, 1888]. Skeptical [3533, 2403]. Skepticism [3533]. Skies [3381, 1586]. Skills [633]. Skin [1028]. Skinner [1862]. Sklar [1685, 2199]. Skull [1188]. Sky [1450, 3343, 2729, 3519, 649]. Sky-watching [3519]. Skyscraper [2675]. Skywatchers [2968]. Slack [2723]. Sleep [3417, 2362]. Sleeping [1508]. Sleigh [836]. Slept [2833]. Sloan [3390, 1394]. Slotten [2020]. slowackich [1655]. slozhnykh [910]. Small [2896, 2396, 2828]. Smart [782, 3038]. Smedley [1541]. smeny [1336]. Smil [2146, 3518]. Smith [846, 1483, 852, 110, 1800, 111, 3336, 1748, 346, 176, 1052, 87, 852, 1943, 180, 2104, 1579, 2787, 3488, 3153, 1518, 2106, 36, 269, 284, 1663, 3102, 3133, 225, 3187, 1584, 2168, 1416]. Smithsonian [2215, 1786, 3214, 1667, 3261]. Smocovitis [3389]. Smoking [1619].

Smolenov [81]. Smolin [3083]. Smoller [2078]. Smyth [968, 3393, 2736, 172]. Snakestone [2054]. Snelders [516, 994, 1535, 2101]. Snyder [1702, 590, 3230]. So-Called [1924]. Soap [1414]. Sober [2662, 3390, 390, 1962]. sobre [1024, 1690, 352, 2158, 1197, 1040, 2741]. Sobriety [2872]. Social [1490, 539, 709, 2552, 2597, 902, 1324, 1262, 1734, 1707, 3424, 3553, 11, 2230, 1596, 975, 1399, 1838, 1591, 2785, 1187, 3124, 3309, 489, 969, 1174, 1400, 1056, 3331, 595, 1843, 1849, 2876, 2055, 952, 2840, 2868, 785, 2024, 1415, 2234, 263, 808, 1971, 3554, 2109, 2201, 1309, 1255, 2746, 872, 335, 685, 2940, 1758, 3452, 815, 2319, 2450, 3177, 3391, 1057, 1425, 3544, 1509, 506, 499, 1508, 1233, 590, 2139, 1703, 3272, 110, 683, 855, 884, 1244, 2142, 389, 3264]. Socialism [873, 2606, 1951]. Socialist [276, 1617]. Socialists [1668]. Sociedad [524, 3154]. Societät [1120]. Societatis [513]. Societe [1484, 2633, 1201, 1026, 3088]. Societies [3124, 1746, 2216, 1452, 2145, 332]. Society [513, 2453, 601, 2433, 3182, 856, 1294, 3477, 1457, 15, 575, 1193, 378, 300, 111, 2096, 2734, 1012, 3024, 1187, 1174, 1980, 507, 666, 232, 3515, 1265, 3287, 680, 3047, 1743, 142, 1770, 1458, 1248, 1955, 3220, 2452, 2171, 945, 116, 538, 3353, 699, 222, 502, 1079, 3199, 3303, 1614, 221, 3205, 2311, 1972, 3207, 3209, 896, 2627, 1258, 3497, 2947, 500, 3219, 3220, 2468, 239, 1517, 2423, 507, 774, 86, 3428, 1381, 698, 3533, 557, 2453, 209]. Socio [3290]. Socio-history [3290]. Sociobiological [468]. Sociobiology [2840]. sociohistoriques [229]. Sociological [1424, 220, 1770, 2551, 979]. sociologie [3272, 3583]. Sociology [2527, 882, 82, 1055, 11, 3588, 2844, 878, 1424]. Sociotechnical [2703]. Socrate [146]. Soddy [2767]. Soderqvist [3169]. Soemmerring [364]. Soffler [2818]. Sofja [2195]. Soft [1414]. Soggiest [3204]. Soil [291]. sol [1654, 1738]. Solar [2385, 3405, 1171, 1255, 2932, 3384]. Sold [1853]. soleil [2496]. Solid [407, 1865]. Solid-State [1865]. solido [358]. solidum [358]. Solingen [2686]. Solitary [1743]. Sollenberger [1906]. Solomon [642, 590, 3435]. Soloway [956]. Solution [1327]. Solutions [1594, 1162]. Solve [3529]. Solvency [3222]. Solving [3242]. Somatic [2494, 2524]. Some [992, 972, 591, 3215, 3164, 481, 2319, 1315, 1134]. Sommerfeldschule [2378]. Somnium [1274]. son [3415, 1466, 2539, 729]. sondern [549]. Song [507, 1186]. Sonnemann [517]. Sonntag [1200]. Sons [3478]. Soothoff [1111]. Sophia [433]. Sophie [571, 2428]. Sophisms [2153]. Sorabji [2328, 27]. Soraya [3464]. Sorcerer [3426]. Sorell [1151, 1647]. Sorensen [1513, 2454, 3386]. Sorokin [2691]. Sorrell [434]. Sort [1503]. Sorters [596]. Sorting [1243]. Sotres [438]. Souffrin [1546]. Soul [2584, 747, 2143, 2418]. Souleymane [1124]. souper [2895]. Source [2329, 1997, 1950]. Sourcebook [1878, 3191]. Sources [1008, 3255, 1795, 2504, 1906, 91, 203, 2243, 698, 423, 3239, 393, 2606, 3330, 2654, 606]. South [3142, 593, 1452, 2696, 2673, 2699, 1934, 2224, 375, 2546, 1934, 375]. Southeast [2145]. Southern [2729, 1052, 1586]. Southerner [272]. Southgate [2087]. Southwest [2759, 3483, 383]. Soviet [2206, 1583, 2853, 1864, 2385, 3009, 3582, 3090, 2124, 2207, 1582, 55, 1803, 1404, 276, 1435, 577].

Soviets [1956]. sowie [3246, 3080]. Sowing [3156]. Sowjetische [1766]. sowjetischen [861]. Sowjetunion [1766]. Soyfer [2207]. sozialen [3256]. sozialer [2914]. Sozialgeschichte [1266, 2359]. Soziologie [2842]. Soziologische [2131]. Space [1230, 74, 3382, 3587, 583, 3176, 2921, 3008, 2130, 387, 2417, 203, 27, 1238, 682, 1138, 2129, 2685, 2246, 3459, 164, 1138, 362, 3481, 1620, 1137]. Spaceflight [2387]. Spaces [3170]. Spacetime [1236]. Spadafora [1025]. Spain [2813, 2366, 1888, 2072]. Spallanzani [1116]. Spanier [249]. Spannungsfeld [2523]. Spanungsfeld [1215]. Spariosu [692]. Sparks [2953]. spaten [1932]. Spatial [937]. spatiale [535]. Spatmittelalter [1283, 1640]. spatmittelalterlicher [1637]. Spatmittelalters [1721]. Special [2828, 1236, 888, 2076]. Specialization [299]. Specialty [2934]. specie [1107]. Species [2549, 1669, 2764]. Specificity [2764, 3093]. Spectacles [2804]. Spectacular [3284]. Speculatie [2515]. speculatief [2101]. speculativa [437]. Speculative [2894]. speculis [1367]. Speculum [1642, 3244, 2253]. Speech [1099, 1673]. Speiser [2438]. Spekulant [3149]. spekulativen [656]. Spence [1952]. Spencer [1781, 3556, 1865, 225, 953, 2880, 953, 3555, 1667]. sperimentale [2763]. sperimentalismo [1288]. Sperm [3416]. Spezialisten [1766]. speziellen [1314]. Spheres [2323]. Spherical [2731]. Spiders [1477]. Spiegel [2747, 1266]. Spin [3373]. Spinning [329]. Spinoza [645, 2085, 2593, 3434]. Spirit [3267, 602, 2283, 542, 841, 198, 2105, 1052, 3063]. Spiritism [1353]. Spirits [1353]. Spiritualism [764]. Spitz [280]. Splendid [3450, 1406]. Sponer [3278]. Sponsored [3104]. Spontaneous [686]. Spot [1428]. Spotless [3096]. Spotted [963]. Sprache [837, 859, 1381]. Spring [1226]. Spufford [3164]. Spuren [3551]. Sputnik [1582, 2387]. Sputniks [1138]. Spy [225]. Squinting [498]. St [1181]. St. [434]. Staat [1115, 1588, 1856, 1123]. Staatsbibliothek [348, 1337]. Staatsdienst [2747]. Staatskrachten [658]. Staatssammlung [1941]. Stability [2264, 3006, 2913]. Stabilizing [3439]. Stachel [768, 769]. Stacia [2921]. Staden [550]. Stadsbelang [1733]. Stadt [3150]. Stafford [847, 2009]. Stafleu [153]. Stairways [3519]. Staiti [849]. Stalheim [2029]. Stalin [2225, 3287, 1435, 2206, 2853]. Stalinist [3560]. Stallion [225]. Stamhuis [658]. Stand [3013]. Standard [3562]. Standards [3260]. Standing [839]. standsbesef [1733]. Stanford [1954]. Stanislao [2756]. Stanley [1709, 1898, 226, 18, 572, 1030, 1607, 732, 2233, 2954, 672]. Stanton [1557, 2810, 1133, 1689]. Star [1360, 664]. Stark [489]. Starkey [2169]. Starr [186]. Stars [1177, 905, 1976, 2078, 1272, 254, 3519, 905, 2344]. Start [1858]. Starts [1324, 71]. State [906, 1696, 3078, 1591, 2184, 230, 3287, 1652, 1865, 1052, 135, 2279, 73]. States [1594, 3079, 1597, 2283, 968, 594, 595, 1089, 2555, 3098, 2015, 3011, 1868, 2414, 2688, 1138, 3292, 1877, 3306, 2869, 2056, 1968, 1163, 2115, 2927, 2541, 1327]. Statesman [567, 2202]. Statics [1008]. Station [3382, 1858, 1601].

Stationary [667]. statistica [3352]. Statistical [2119, 531, 2404, 3120]. Statistics [3308, 2121]. Statistiek [658]. statistique [270]. statistischer [3535]. Status [2746]. Staudinger [3569]. Staum [3439]. Stayer [38]. Steacie [669]. Steadman [50]. Stealing [1511]. Steam [667]. Stearn [232]. Steel [2240, 3244, 2459]. Stefan [1951, 3246, 3564]. Stefani [1559]. Stefano [2590, 1912, 2892]. Stefanov [81]. Steffens [2447]. Stefoff [3549]. Steinberg [1493]. Steiner [2238]. Steinman [484]. Steinmetz [3049, 1617]. Steinthal [3273]. Steinway [3478]. Stella [1014]. Steller [260]. Stemming [199]. Sten [320]. Stenehjem [1502]. Stengers [2879, 1744, 3186]. Steno [358, 2510]. Stensen [2510, 2816]. Stenzel [2026]. Stepan [1326]. Stephan [1337]. Stephane [2820]. Stephanie [396, 3163]. Stephen [1583, 3419, 2383, 2930, 2909, 2385, 3243, 2698, 1196, 878, 553, 2651, 72, 2927, 2153, 3513, 680, 55, 1020, 3347, 3034, 2430, 483, 737, 71]. Stephenson [2082, 158, 2448, 3036]. Stereo [2026]. Stereochemistry [7]. Sterilization [1327]. Sterling [3108]. Stern [2785]. Sternkatalog [1360]. Sternkunst [40]. Steroids [1134]. Stetter [1323]. Steve [3408, 82, 208, 1051, 3184, 1772, 389]. Steven [50, 2674, 579, 1866, 1730, 1476, 2187, 2771, 3365, 2077, 1279, 3229, 872, 1385, 2699, 737]. Stevens [1673, 2441, 1778, 592]. Steward [3426]. Stewart [581, 2374, 2183, 1479, 1119, 1276, 776]. Stibbe [3061]. Stichweh [1115, 2131]. stiefkind [1646]. Stig [3386]. Stigler [202]. Still [121]. Stiller [755]. Stillings [748]. Stillman [738, 831, 649, 1792]. Stimson [101]. Stine [2641]. Stockel [2914]. Stocking [120, 1850]. Stoczkowski [1982]. Stoeger [207]. Stoff [966]. Stoffel [3460, 3298]. Stoic [1383, 634]. Stokes [684, 2386, 1053, 3561]. Stolberg [2281]. Stoll [1638]. Stolte [1245]. Stolte-Heiskanen [1245]. Stoltzenberg [2122]. Stone [3419, 2035, 3348, 144]. Stones [3525]. Storia [3104, 327, 521, 771, 3268, 811, 1737, 564, 1779, 1195, 1498, 2890, 122]. storica [2518]. Storicità [124]. storico [2443, 1473]. storico-critica [1473]. Stories [2569]. Stork [2695]. Story [3163, 3373, 621, 2783, 3228, 2778, 594, 2856, 2540, 1985, 965, 2770, 3421, 631, 385, 3482]. Strabon [1351]. Strafe [12]. stran [887]. Strands [2690]. Strange [3163, 2905, 2112, 3238]. Strangelove [3477, 2409]. Stranger [201, 573]. Strangers [412, 1050]. Strassburg [3360]. Strassburger [2336]. Strata [1242]. Strategies [1334, 1342, 1876, 721]. Stratosphere [474]. Straws [2738]. Street [590]. Streets [1933]. streghe [1643]. Streits [859]. Streitwieser [2837]. Stress [2472]. Strikes [1078, 1971]. Stringfellow [373]. Stross [291]. Stroud [1483, 2322, 3421]. Stroup [927]. Structural [3329]. Structure [734, 622, 487, 1105, 153, 808, 1290, 2940, 2358]. Structures [1063, 905, 462]. Struggle [1170]. Struik [2714]. Struktur [1932, 1215, 81]. Strukturen [2411, 654]. strumenti [2963]. Struve [756]. Stuart [1954, 2030, 542, 3379, 3057]. Stubbe [525]. Stubborn [291]. Stückelberger [341]. Students [1482, 1110]. Studi [2000, 1498]. Studia [2185]. Studie [1322]. Studien [2754, 1996, 439, 1808, 1300, 3256, 2046, 730, 3341, 1724, 2583, 2843]. Studienbeginn [2185]. Studienplan [1651]. Studies [248, 513, 2896, 2552, 2343, 2408, 115, 1162, 1617, 1332, 1000, 417, 723, 2273, 884, 2851, 2091, 2941, 1134, 2151, 507, 273, 1097, 1704, 301, 2705, 350, 32, 529, 506, 2721, 3250, 1676, 2925, 1703, 513, 118, 905, 647, 1184, 1242, 1108, 808, 1119, 109, 2327, 1639]. Studium [1134]. Study [2148, 997, 2289, 622, 846, 3563, 2768, 726, 391, 435, 336, 2860, 167, 1402, 1358, 2959, 3121, 684, 1087, 1742, 1775, 436, 2236, 1293, 249, 1344, 3342, 2087, 1717, 1669, 3265, 3497, 682, 2159, 1549, 1289, 2731, 1206, 3330, 2231, 1532, 559, 2594, 2498, 2420, 319, 2111, 1465, 3177, 2821]. Stuhlhofer [12]. Stump [2718, 344]. Sturdy [2746]. Sturlese [2890]. Stuttgart [325]. Style [1612, 1031, 1204]. Styles [1611, 833, 1891, 2534]. Sua [3208]. suan [2732]. Suárez [184]. subactis [2000]. Subalternation [2077, 1279]. Subculture [1528]. Subject [866]. Subjected [2612]. Subjektivitat [2977]. Subjekts [859]. Sublime [2370]. Subscription [3220]. Subscriptions [493]. Subsection [3204]. Substance [3419, 2418, 2085, 546, 2927]. Subversions [1384]. Subversive [1065]. successo [771]. Successors [2177]. Such [3406]. Suche [851, 1122]. Suchodolski [512]. Sudden [67]. Suddenly [2129]. Sudhoffs [518]. Sue [2269, 182]. Suède [320]. Suffering [2384]. Suffrage [2995]. Suffrages [2603]. Sugar [627]. Sugarcane [2972]. Sugars [1134]. Suicide [1807]. sulfamides [587]. sulla [1589, 2756, 2421]. sulle [2098, 1915, 2097]. Sull’ottima [1002]. Sultan [2073]. ’Sulva [28]. ’Sulva-sutra [28]. ’Sulva-sutras [28]. Sum [2839, 2127]. Summa [1465, 1379, 3523]. Summaries [2987]. Summer [2613, 3363]. Sun [3334, 1171, 1015]. Sundials [3320, 34, 1385, 34]. Sunspots [8, 3499]. Suntrup [345]. Supek [930]. super [1366, 2892]. Superbomb [672]. Supercomputer [2856]. Superconductivity [3566]. Superior [3038, 2053]. superiore [2270]. Supermen [2856]. Superstition [2315]. superstructa [39]. Suplemento [2272]. Suppe [691]. Suppl [2273, 1941]. Supplement [2788, 1748, 1869, 1531]. Supplements [255]. Support [2847]. sur/Studies [529]. Surgeon [3345]. Surgery [3476, 3295]. Surgical [3295, 1680, 2362, 1327]. Surrounding [893]. Survey [2681, 1280, 620, 397, 1778, 2753, 265, 1163, 2020, 206, 1624, 1154]. Surveying [1461]. Surveyor [3401]. Surviving [3393, 1857]. Survivors [2384, 2613]. Susan [2639, 3284, 2384, 848, 287, 70, 1517, 1491, 2781, 968, 664, 179, 1244, 2666, 292, 88, 2722, 2612, 246, 3292, 3063, 2601, 3302]. Susana [3139]. Susanna [1562]. Susanne [1001]. Suspicion [2700]. sutra [28]. sutras [28]. Sutton [2433, 116]. Suzanne [1294, 824, 2728, 47, 2297]. Suzuki [238, 2577]. Svante [3080, 475, 2036, 475]. Swann [64]. Swartz [384]. Swastika [1857]. Swearing [1619]. Sweatshop [872]. Sweden [618, 1562, 3071]. Swedish [2036, 618, 1248, 699]. Sweet [38, 1134]. Swerdlow [3520]. Swetz [1461]. Swift [1299]. Swijtink [418]. Swimming [2310]. Swiss [768, 2460, 2117, 1697]. Sword [225]. Sybil [1662]. Sydenham [39, 39]. Sydney [852, 269, 1748]. Syed [2071]. Sylva [524]. Sylves [225].

Sylvester [2759, 2282]. Sylvia [1559, 53, 1711]. Sylwan [1366]. Symbiogenesis [2231]. symbolic [2480]. Symbolism [2817]. symbolistes [824]. Symbols [2795, 2720, 3143]. Symmetrie [536]. Symmetries [188]. Symmetry [880]. Symoens [1812]. Sympathetic [2817]. Symposion [160]. Symposium [1656, 3252, 650, 1087, 188, 583, 1011, 3421]. Symptoms [2779]. Synchrotron [875]. Syndrome [67]. Synergy [2837]. synonymariis [1637]. Synthèse [522]. Synthesis [3389, 2858, 311, 1673, 3016]. Synthesizing [3010]. Synthetic [781, 1752]. Syphilis [376, 2297]. System [3267, 2385, 3405, 1005, 2095, 44, 1038, 75, 2441, 969, 3296, 1760, 412]. Systematic [3104, 3426]. Systematics [2441]. Systematiker [1930]. Systematologie [259]. système [2960, 859]. Systems [910, 11, 85, 3436, 387, 2795, 850, 2522, 244, 2932]. Szasz [1710]. Szilard [1581]. Szollosi [3568]. Szollosi-Janze [3568]. Sztompka [2940]. Szydlo [2663].

T [3104, 84, 3019, 2659, 2032, 2604, 213, 2392, 1132, 3463, 172, 15, 781, 1680, 1191, 1604, 3131, 3381, 3326, 1323, 199, 1021, 3587, 2542, 1655, 1223, 3266, 1768, 2849, 3230, 468, 1452, 1286, 2655, 2266, 288, 232, 51, 69, 2540, 2379, 493, 1190, 2908, 188, 225, 1863, 697, 1584, 164, 1791, 794, 1627, 1090]. T. [2933]. Table [138]. Tables [1361, 255]. tabulis [2089]. Tachoire [2237]. Tacitus [1474]. Tacitus-Ubersetzung [1474]. Tacuini [2070]. tadbir [1718]. Tadeusz [512]. Tadhkira [1991]. Tafuri [828]. Tagebuch [42]. Tagliaferri [2355]. Takacs [3466]. Takahashi [1367]. Takahata [2208]. Takakazu [2436]. Takako [2537]. Take [1774, 3259]. Takebe [2436]. Takes [2364]. Takeshi [3373]. Taking [1619, 2037]. Talbot [1491, 2676]. Tale [1299, 1759, 854, 1009, 1301]. Tales [3324, 17]. Talking [1144, 60, 3372]. Tallis [1458]. Tamara [1513, 3237, 2037]. Tambiah [732]. Taming [971]. Tan [2815]. Tancredi [590]. Tandy [3331]. Tangled [3479]. Tannery [3527]. Tansey [2935]. Tapia [921]. Taquet [3572]. Taqwim [2070]. Tar [1761]. Taraba [748]. tardive [1093]. tardo [645]. Targeting [2309]. Targets [2854]. Tariol [2272]. Tarr [54, 3234]. Tascon [1389]. Tassinari [2421]. Taste [3398]. Tatarewicz [682]. Tate [1411]. Taterewicz [1137]. Taton [739]. Tatyana [3369]. Taub [1548]. Tauber [1033, 2858]. Taubes [1959]. Tayler [378]. Taylor [2672, 3462]. Tazi [237]. Tchudi [3347]. te [3068, 994, 2515]. Teacher [2201]. Teaching [3158, 3213, 2308, 2076, 3403]. Tecce [521]. Tech [696]. Technical [244, 885, 1427, 2125, 3260, 2856, 1309, 3426, 3480, 680]. Technics [2522]. techniek [994, 516]. Technik [517, 3253, 3253, 1705]. Techniker [1766]. Technikgeschichte [519]. Techniki [512]. Technikwissenschaften [2465, 2757]. technique [212, 516]. Techniques [2847, 1625, 522, 1987, 2712, 2324, 1356]. Technische [466, 2535, 2814, 2646, 1856]. Technischen [325, 3367, 2646, 1856]. technischer [2661, 1531]. technisches [3385]. Technologe [3447]. Technological [2370, 2386, 629, 1081, 11, 2205, 336, 968, 2577, 3561, 3578, 2784, 246, 3040]. Technologie [3256]. Technologies [3006, 3014, 3384]. Technology [140, 997, 870, 1543, 2100, 2321, 516, 1765, 1864, 910, 1776, 200, 3558, 1460, 2641, 111, 1479, 2860, 1491, 2389, 2993, 2851, 1645, 1187, 3214, 3559, 484, 329, 3394, 2410, 2536, 1980, 2652, 2049, 2061, 2200, 1968, 3176, 103, 1186, 2682, 1286, 123, 128, 3289, 3032, 666, 1963, 1432, 1566, 1164, 3515, 3513, 2060, 3286, 419, 2422, 2972, 872, 22, 2775, 1767, 3161, 1901, 274, 225, 1955, 682, 483, 1088, 1584, 1233, 1892, 1137, 1877, 678, 1600, 2686, 3280, 2621, 2683, 1492, 23, 300, 111, 3513, 872, 3089, 1061, 818, 2233, 2642, 3169, 1644, 111, 3406, 1218]. Technology [2438, 243, 3101, 2961, 780, 1173, 387, 2234, 296, 1704, 3041, 792, 3092, 2045, 3040, 2705, 415, 1903, 2488, 515, 911, 11, 111, 3336, 817, 681, 3513, 2545, 425, 2971, 54]. Technoscience [3084, 2475]. Technoscientific [1605]. tecnica [1331, 2255, 1389]. tecnicas [3442, 524]. Tecnología [526]. tecnologica [2825]. tedesca [3054, 2890]. Teeter [1380, 2268]. tegen [2693]. Teich [1706, 420, 1727, 169]. Teichmann [1865, 585]. Teilchen [875]. Teitelman [590]. Teixeira [2027]. Tekhelet [249]. tekhniki [910, 511]. Telegraph [3515]. Telegraphy [1494]. téléonomies [1063]. Telephone [1843]. Telescope [3456, 682, 1586]. Telescopes [1763]. Telescopic [904, 3514]. Telkès [212]. Telle [1721, 3531, 3531, 2161]. Teller [672, 225]. Telling [1081, 3057]. Teltschik [1588]. temas [351]. Temkin [1180]. Temmer [967]. Temperate [2802]. Temperature [488, 3566]. Templeton [2416]. tempo [2340]. Temporal [125]. temps [549, 2573, 2028, 3411]. Temptations [2190]. Ten [2609]. Ten-Volume [2609]. Tenon [2981]. Tension [1647]. Tensorkalkuls [2755, 1587]. Tentacles [22]. tenu [3460]. Teologia [2340, 645]. teoria [1589]. teorica [2756]. teorii [1336]. ter [3461]. terapeutica [2158]. Teratogeny [1784]. Tercentenary [1656, 1108]. tercio [3475]. Terence [459]. Teresa [1690, 2271, 1199, 1355]. Teresi [1866]. Terman [596]. Terminology [2499, 3430]. Terms [1629]. Terra [2750, 376, 3123, 1589, 2097, 1648]. Terrada [1911, 2491, 2491, 2158]. Terradas [1331, 1331]. Terrae [2271]. Terre [535, 3418, 2247, 1026]. territorio [2825]. Terry [684, 3513]. tertius [727, 1634]. Tesla [3271]. Tessa [2577]. Test [3010, 2607, 385]. Testament [1536]. Tester [338]. testimonianze [743]. Testing [1851, 2213, 225, 596, 596, 3377, 225]. testo [1102]. Tests [2040]. Teulón [523]. Teun [1420]. Texas [3074]. Texaurus [2155]. Texera [1084]. Text [39, 1638, 2739, 339, 923, 2248, 2808, 3433, 428, 452, 3248]. Textbook [2308]. Texte [2251, 3460, 2506, 1185, 3273, 360, 1176, 2242, 259]. Textes [2010, 2182, 161, 2508, 210, 730]. Textile [329]. Textos [1911]. Texts [905, 1687, 808, 2243, 3589, 1065, 2768, 3354, 1462, 2152]. Textual [723]. Teyler [3158]. TH [1856]. Thackray [107]. Thagard [80, 1512]. Thatcher [1301]. Theater [970, 3051]. Theerman [1656, 3511, 1573]. Their [420, 1998, 2806, 2250, 2619, 905, 988, 317, 836, 346, 349, 1746, 27, 3261, 689, 537, 3174, 1544, 635, 1211, 1381, 611]. Themes [903, 1383, 3040, 1873]. Theo [2658, 1377]. Theodericus [2335]. Theodor [1317, 654, 995, 1134, 1914, 2107, 1225]. Theodore [418, 2406, 2442, 2581, 731, 2540, 1861, 3012, 1493, 3073]. Theodosius [2126]. Theologica [1474]. Theological [1755]. théologique [1571]. Theology [650, 207, 748, 96, 724, 1719, 668, 1642, 1000, 1112, 436, 1114, 1279]. Theon [1361]. Theophrast [643]. Theophrastus [1359, 1550, 1906, 3239, 2243]. theoreme [2592]. Theoretical [2181, 2124]. theoretische [663]. theoretischen [2378]. Theoria [2183]. Theorie [1322, 1481, 439, 259, 163, 3341, 2818, 925, 149, 2508, 660]. Theorien [81]. theoriengeschichtlichen [1313]. Theories [1336, 622, 49, 691, 775, 2569, 153, 2235, 2197, 2724, 2911, 2267, 747, 27, 2245, 1545, 3426, 1009, 2831, 937, 2918, 1517, 235, 710, 153]. Theorist [272]. Theorizing [1784]. Theory [486, 734, 3520, 1306, 809, 76, 1547, 2837, 977, 2418, 36, 220, 542, 2894, 147, 1961, 1087, 2198, 2836, 612, 3588, 310, 1078, 349, 1869, 2358, 1675, 1236, 3133, 2208, 379, 1198, 2703, 2246, 1455, 586, 164, 173, 1090, 1866, 2952, 1008, 2597, 1662, 2919, 1764, 133, 1096, 1033, 3449, 1489, 1422, 1132, 1334, 317]. Therapeutic [3011, 1271, 1153]. Therapie [1813]. Therapieprinzip [2994]. Therapy [2776]. Theresa [595]. Therese [2331, 212]. thermochimie [2237]. thermodynamique [2237, 3067]. These [2474]. Thesis [882, 928]. Theunissen [516, 127, 368]. Thibaut [28]. Thierre [2182]. Thierry [820, 3542, 2149, 3157]. Thijssen [3429]. Thing [2797]. Things [2100, 3010, 2266, 2071]. Thinker [2431]. Thinking [1891, 1673, 1062, 88, 2591, 1605, 2798, 1221, 2060]. Third [1209, 2197]. Thirteenth [917, 103, 2584, 2632]. Thirties [2124]. Thirty [1087, 2776, 314, 188]. Thirty-Five [314]. Thom [144]. Thomae [39]. Thomas [244, 1177, 1316, 1603, 39, 1325, 629, 437, 58, 1617, 663, 2516, 1332, 1007, 465, 3169, 1310, 958, 1922, 2748, 2036, 1517, 451, 3348, 2096, 2507, 3151, 2955, 1947, 2977, 1087, 3456, 2369, 468, 436, 1914, 832, 572, 686, 469, 731, 2583, 1820, 2459, 2772, 619, 1392, 2765, 1919, 2414, 2603, 634, 1822, 3282, 13, 786, 1261, 477, 2644, 2987, 2821, 1393, 2850, 3030, 3227, 3497, 2731, 1483, 1947, 925, 567, 3148, 1127, 356, 2500]. Thomasset [148, 3189]. Thompson [809, 88]. Thomson [1399, 554, 2664]. Thoreau [2525, 3265, 2758, 2991]. Thoren [707, 1106]. Thornton [3237]. Thought [392, 2797, 362, 3160, 1007, 833, 2748, 433, 189, 1591, 1187, 1514, 1357, 1668, 3247, 35, 1906, 832, 3371, 950, 787, 319, 3584, 1612, 2534, 668, 2450, 2243, 1861, 2126, 2361, 1714, 3264, 1990, 1513, 1169, 1170, 298, 3339, 2611, 1731, 1800, 2571, 3413, 3536, 2959, 952, 2997, 1383, 2417, 1380, 636, 2500, 630, 1170, 1513]. Thoughts [3525, 3000, 3050, 2815]. Thousand [1702, 911]. Thousand-Year [911]. Threat [2289, 3264]. Three [2805, 1279, 349, 2985, 1697, 2087, 2349, 1990, 2793, 3519]. Three-Body [2985]. Threshing [336]. Thrice [1557]. Thrice-Famous [1557]. Throm [2994]. Thrower [839]. Thulesius [3178, 1563]. Thumb [1088]. Thurston [1986]. Tian [1712, 2918]. Tibbetts [1452]. Tidal [310]. Tide [199, 2310]. Tieck [940]. Tierversuchs [1658]. Tifashi [3525]. Tihon [3524, 1361].

Tijdschrift [516]. Tiles [2784, 486]. Till [2382]. Tilman [2117]. Tim [1324, 2106]. Time [362, 1707, 3320, 1837, 3259, 2318, 1003, 3057, 1620, 1417, 257, 3164, 1019, 203, 532, 722, 2550, 3120, 1895, 125, 2500, 1503, 164, 1184, 1134, 1362, 13, 254, 15, 3320, 1097, 3507, 225, 1238, 125]. Timee [1178]. Timee/Critias [1178]. Timekeeping [5]. Times [2799, 3271, 2833, 2205, 3270, 1936, 1450, 1087, 533, 1408, 3167, 1775, 1959, 3012, 741, 2297, 2150, 1768, 35, 2474, 3333]. Timiryazev [570]. Timmermans [3062]. Timo [1765]. Timothy [2343, 1680, 3522, 3589]. Tinbergen [2680]. Tinctoria [1428]. Tinfoil [2026]. Tinkering [1683]. Tinkler [813]. ’Tis [43]. Tissot [1568]. Titani [2428]. Titchener [779]. Title [1442]. Titus [225]. Titze [2413]. Tobies [517]. Toby [1798, 2762]. Today [2938, 1672]. Todd [1549, 375, 3339, 2904]. Todes [1170]. Tögel [81]. Together [253, 2426, 620, 428, 452]. Token [1245]. Tolan [1995]. Tollot [1026]. Tolomio [2063]. Tom [3361, 1151, 1647]. Toman [3101, 2783]. Tomas [2158, 1911, 1470]. Tombaugh [1044]. Tomes [2470]. tomis [1375]. Tomlinson [1553, 1539]. Tommaso [449, 736, 2432]. Tomonaga [3373, 2537, 2301]. Tomorrow [2129]. Tony [2974, 2613, 2074]. Too [1257, 595, 3218, 1220]. Tool [1649]. Tools [2941, 1506, 2784]. Topham [3263]. Topographie [998]. Tora [1625]. Torah [433]. Tore [841, 618, 2440, 1739]. Toretti [978]. Torlais [460]. torno [1374]. Torres [2519]. Torrey [2455]. Torricelli [407]. Torrini [1925]. Torsten [2013, 3552]. Tort [3555]. Torti [566]. Torture [3225]. Toscana [3545, 743]. Tosi [743]. tot [240, 2975, 228, 1535, 127, 2092]. Totalitarian [3089]. Totman [330]. Toumey [1983, 2720]. Tour [3085]. tournant [1296]. Tours [5]. tous [1312]. Tov [433]. Town [1825, 3325]. Toxic [3288]. Tra [1556, 3139, 3129, 2260]. Trabant [845]. Traces [3508]. Tracing [2777, 1712]. Tract [451, 253]. Tractatio [1103]. Tractatus [438, 1103]. Tractrix [516]. Trade [713, 2096, 646, 3358]. Tradition [1891, 723, 1453, 2412, 733, 1687, 931, 758, 1376, 434, 1360, 135, 1601, 3533, 2621, 1797, 631, 1196, 2722, 2887, 711, 421, 1549, 65]. Traditional [336, 1452, 2145, 1647, 129]. Traditionen [394, 845]. Traditions [1128, 905, 923, 1539, 1755, 721, 1367, 918]. tradizione [3434]. traducteurs [1093]. Traduction [2149, 727, 914]. Traductions [1093]. Traduits [1176, 2242]. Tragedy [2207, 483]. Trails [2783]. Training [3359]. traite [2070, 451]. Traktat [2502]. Tranquillity [1383]. Transcendence [408]. Transcendentalism [1784]. Transcontinental [47]. Transcripts [452, 1679]. Transfer [1765, 22]. Transference [288]. Transfigured [722]. Transformation [138, 2427, 2891, 3358, 3241, 843, 3087, 3151, 2577, 1330, 681, 1562, 668, 293, 2047, 1126]. Transformations [574, 538]. Transformed [1050]. Transforming [3368, 1759, 872, 2545, 1128]. Transistor [3385]. Transistorisierung [3385]. Transit [3101, 2783]. Transition [876, 3325]. transito [1828, 3154]. Translated [2906]. Translation [39, 2075, 1715, 1716, 1913, 1012, 3331, 3436, 1293, 2426, 2248, 1465, 3133, 3335, 428, 2731, 1198, 1367, 641, 1298, 452, 3114, 768]. Translations [3354, 349, 2808]. Transmission [1089]. Transmutation [3502]. Transvaal [1229]. Tras [2517]. Trash [273]. tratados [432]. Trattati [2590]. Trattato [2421]. Traumatic [2472]. Traunecker [935]. Travaux [825, 2436]. Travel [151]. Traveler [1298]. Travelers [1949]. Travellers [3262]. Travels [750]. travers [1026]. Travis [1752]. Traweek [192]. Treasure [2178]. Treated [3528]. Treatise [3242, 1096, 2731, 3438]. Treatises [349, 732, 2951, 2974]. Treatment [3380, 3579, 2808]. Treaty [2607]. Trebizond [2425]. Trecento [1915]. Trees [2226]. Trefil [534]. tremblements [2247]. Trembley [166, 322]. Trembly [612]. Trennert [3577]. Trent [1688, 2115]. Trepl [812]. tres [1918]. Tresor [3538]. Tresorerie [3538]. Treue [89]. Trevisani [1821]. Trevor [417, 773, 2044]. Trial [3504, 3363, 1958, 649, 674]. Trials [2088]. Trib [2297]. Tribiblos [2581, 2581]. Trigger [630]. trigonometria [2089]. trigonométrie [931]. Trigonometrisch [1013]. Trigonometrisch-astronomisches [1013]. Trinity [225]. Triomphe [3418]. Tristan [2338]. Triumphans [822]. Trocadéro [1129]. Trocchio [853]. Troels [2510]. Trohler [1660, 518]. Trois [1026]. Trolle [3163]. Trollinger [2609]. Tropical [1034, 1035, 2239, 1229]. Tropp [1852]. Trotter [572]. Trouble [3105]. Troubled [1768]. Trousers [2325]. Troyer [1761]. Trudy [2547]. True [1020, 2770]. Truesdell [2008, 112]. Truly [1693]. Truman [2924]. Trust [2406]. Trusted [2724]. Trusting [3400]. Truth [2230, 864, 428, 787]. Truths [1926]. Trypanosomiasis [3378]. Tu [1625]. Tuana [1799]. Tub [1299]. Tube [385, 3010]. Tuberculosis [2468, 1596, 1838, 593, 2538, 2109, 3162]. Tuberkulöse [374]. Tübingen [266]. Tucci [2355, 2963]. Tuchman [2184]. Tucson [3261]. Tudor [1999, 155]. Tufte [2859]. Tummers [2075]. Tunik [1095]. Tunnels [2852]. Turbulent [3381]. Turing [3370]. Turku [2264]. Turm [1313]. Turn [2289, 3307, 2055, 142]. Turn-of-the-Century [2055]. Turnbull [684]. Turner [3349, 878, 2655, 155, 601, 3158, 326, 2187, 3422, 103, 878, 1831, 3426, 1802, 843]. Turney [3355]. Turns [3085]. Turriano [1374]. Tusi [1991]. tussen [2292, 2092]. Tutzke [517]. Tweedale [598]. Tweedle [2655]. Twelfth [2398, 2137]. Twelfth-Century [2137]. Twentieth [2289, 1795, 623, 2848, 2833, 1873, 970, 2608, 2036, 64, 2851, 2602, 3489, 2197, 2771, 3260, 3579, 3190, 1506, 2918, 1504, 2997, 2935, 956, 3289, 2609, 1231, 2025, 1139, 963, 2545, 2844, 483, 592, 381, 2869, 1078, 805, 2640]. Twenty [1702, 583, 2577, 533, 2641]. Twenty-First [2577, 533]. Twenty-Five [583]. Twenty-Two [1702]. Twists [3085]. Two [906, 1702, 1696, 1293, 3470, 854, 65, 2831, 1530, 1702, 831, 2214]. Tyacke [3136]. Tycho [1106, 739, 1253, 2428]. Tyler [1666]. Tylkina [570]. typ [348]. Type [574]. Typhoid [376, 1078, 1971]. Typology [1243]. Tyranny [3013]. Tyson [1483].

U [110, 1134, 2415, 894, 930]. u.a [2834, 2291]. U.S [3463, 2020, 2826, 2542, 3173, 387, 176, 606, 717, 2685]. U.S. [1163].

Überblick [3118]. Überlieferungsgeschichtliche [1996]. Ubersetzung [1474, 1638, 1360]. Übersetzungen [2506]. Udo [1937, 730, 3134]. Uebel [1332, 2765]. Ugo [2000]. Uhr [2336]. Ulansey [544]. Ulfried [286, 1678]. Ulisse [743]. Ullmann [3367, 3145]. Ulm [3338]. Ulrich [1638, 1660, 1625, 348, 1766, 518, 1393, 1301]. Ulrike [3566, 1829, 1046]. Ulserius [2335]. Ultimate [3234]. ultramar [1354]. umanista [1556, 2893]. Umfeld [3256]. Umkreis [1381]. Umweltkonflikte [2281]. Un’analisi [1473]. Unauthorized [43]. Uncertain [2774, 3185]. Uncertainty [1617]. Undergraduate [3213]. Underground [666, 3471]. Understand [3228, 2704]. Understanding [978, 2785, 1673, 207, 2928, 2736]. Undertaking [756]. Underworld [1164]. Underworlds [43]. Unger [1645]. ungluckliche [3441]. Unguru [1634]. Unification [3391, 676]. Unified [76, 2197]. Unifying [3389]. Union [2206, 1803, 2458, 1583, 1864, 3582, 600]. United [1594, 3079, 1597, 2115, 2283, 968, 594, 595, 1089, 2555, 3098, 2015, 2927, 3011, 1868, 2541, 1327, 2414, 2688, 1138, 3292, 679, 1877, 3306, 2869, 2056, 1968, 1163]. Unity [57, 1279, 356, 429, 546]. univers [2895, 2004, 1237]. Universal [2343, 3369, 905, 3143]. Universality [534]. Universe [1548, 1438, 1866, 235, 671, 2235, 2771, 782, 1852, 3265, 156, 2736, 2659, 1177, 2120, 990, 3228, 2769, 2198, 2770, 2831, 252]. Universidad [1690, 2090]. Universität [2646, 2254, 1941, 1651, 2131, 1501, 1115, 471, 1123, 578, 1572]. Universität-Gesamthochschule-Siegen [578]. Universitäten [3360, 2413]. Universitäts [2359]. Universitäts- [2359]. Universitätsreformen [1304]. université [320]. universiteit [2292]. Universities [29, 1585, 1812, 2364]. University [87, 3252, 1191, 948, 2739, 1247, 3429, 3087, 3136, 301, 1812, 619, 3047, 1264, 566, 1047, 1696, 3210, 2308, 2711, 3368]. Unknown [2685]. Unpublished [1656, 2463, 2244]. Unraveling [3508]. Unreason [865]. Unregulated [202]. Unrolling [257]. Unschuld [3567, 518]. Unseen [3007]. Unselfish [3390]. unsere [3567]. Unternehmer [1842]. Untersuchung [842, 2818]. Untersuchungen [1185, 998, 1637]. Unto [3390]. Unvollstandigkeitssatze [2679]. Urania [2428]. Uraniborg [1106]. Uranium [2461]. Urbach [2164, 790]. Urban [2754, 3326, 2738, 1003, 3234, 3325, 53, 2466, 245]. urbano [2825]. Urmson [1362]. Urry [1849]. Urs [42]. Ursula [3447, 1927, 266, 518]. Urteil [12]. Urvoy [1276]. usages [1307]. Use [25, 3299, 2924, 2964, 3336, 2701, 1778, 2076, 3414, 346]. Used [3061, 1717]. Useful [2598, 1926]. Useless [1014]. Uses [2209, 417, 1196]. Using [1618]. uso [1374, 432]. USSR [1404]. Uta [767]. Ute [1410]. utilitaristische [394]. Utility [681]. Utrecht [2292]. Utter [2924]. Uwe [654, 3091].

V [2433, 2691, 3451, 2303, 3295, 1212, 207, 2084, 1184, 189, 1454, 2178, 329, 372, 1947, 386, 1928, 91, 2935, 1172, 1918, 2111, 2002, 493, 2288, 829, 964, 2185, 2458, 1026, 1198, 2029, 3055, 586, 2242]. V-IX [2242]. Vaccari [2097]. Vaclav [2146, 3518]. Vagliente [1008, 2906]. Vaines [3538]. Valdecilla [868]. Vale [3282, 3282]. Valencia [2490]. valenciana [1354]. valencianas [3442]. valencianos [1354, 2435]. Valentin [3003]. Valentina [2124]. Valerie [588]. Valery [2207]. Vallance [1090]. Valles [351]. Vallon [2585]. valore [2526]. Valsiner [1673]. Value [2361, 1083]. Value-Free [1083]. Values [2405, 685, 3347, 2619, 238, 2784]. Valverde [746]. Vance [2609]. Vandeveer [2112]. Vanessa [2296]. Vannevar [3366, 987]. Vanpaemel [361]. Vardalas [1957]. Vardanian [1014]. Variance [1335]. Varieties [1552]. Varisco [2073]. Vassiliki [3389]. Vater [861]. Vatican [1815, 2618]. Vaughan [3176]. Vaulézard [349]. Vaux [339]. Vavilov [2225]. Vay [1087]. Veblen [3027]. Vec [3341]. Veenstra [3433]. Veer [1673]. vegetabilibus [2502]. vegetal [2517]. vegetales [1911]. Vehicles [3101]. veintiun [154]. Velody [121]. Velthuys [1777]. Velthuys-Bechthold [1777]. Venable [948]. Venetischer [2742]. veneto [2097]. Venezuela [1084]. Venice [828]. Venkataraman [964]. Venus [3476]. Vera [3082, 1538]. Verabschiedung [1941]. verbale [2326]. Verbeek [2658, 1377]. Verbiest [1819]. Verbindung [1808, 3535, 1927]. Verbrechen [2382]. verdad [1197]. Verdet [2743]. Verdrängte [471]. verfasst [549]. Vergangenheit [2161]. Vergangenheitsbewaltigung [3586]. vergessen [3278]. Vergleich [1481]. Verhaltens [2464]. Verhaltensforschung [2464]. Verhältnis [656]. Verhaltnisse [2013, 2013]. Verknüpfung [325]. Verkunder [2342]. Verlet [2350]. verlorenen [1113]. verloskundig [3068]. Vermeulen [2408]. Vermij [1297]. Vern [2038]. Vernadsky [2288]. Verne [1688]. Vernet [2334]. Vernon [3125, 1127, 2610]. Vernuft [2515]. vernuftig [2814]. Veronica [1245]. Veronique [2093]. verschiedenen [2013]. Verschuur [1534]. Verschuuren [2619]. Version [33, 914]. Versions [1384]. Versuch [2353, 3273]. Versuchsstationen [266]. versus [1656, 278, 1074, 3586, 2481, 483]. Vertebrate [1254, 1590]. Verteidigung [1658]. Vertinsky [946]. Vertreibung [1410]. Verwendung [1587]. Very [208, 2770]. Verzeichnis [2013, 1393]. Verzekering [2693]. Verzekering-Maatschappij [2693]. Vesalio [2656]. Vesalius [2257]. Vescovini [919]. Vestiges [2017]. Veterinary [2499, 424, 2029]. VI [2502]. viaggio [2355]. viajes [2158]. Viaticum [826]. Vicente [1723]. Vickers [3257]. Vico [1823, 454]. Victor [1008, 1617, 435, 2906, 1106, 2490, 2835, 2124, 1140, 766, 1533, 1227, 61, 1181, 707, 435]. Victoria [963]. Victorian [3262, 1036, 170, 2828, 575, 370, 1832, 574, 48, 268, 468, 1831, 764, 2822, 847, 665, 1940, 47, 369, 2524, 373, 3357, 1164, 1255, 302, 3456, 3069]. Victorians [1302]. Vidal [752, 2375, 1654]. vie [834, 2079, 1032, 2508, 3122, 236]. Vienna [1332, 2199, 1228, 1601, 2868]. Vienne [3583]. vier [3246]. Vierfarbenforschung [673]. Vierhaus [1215]. Vierteljahresschrift [518]. Viète [349]. View [528, 622, 990, 1922, 489, 1216, 3265, 2500, 839, 1421, 1620]. Viewed [1047]. Viewing [969]. Views [970, 904, 1636]. Vigevano [2155]. VII [1358]. VIII [2501]. Viktor [1852, 2715]. Vila [3344]. Viladrich [1261]. Vilain [2900]. Vilhelm [193]. villa [3444]. Villanova [438]. Vincent [2879, 388, 1414, 3056, 1897, 1720, 225, 2760]. Vincent-Pierre [1897].

Vincenzio [2590]. Vincenzo [2269, 521]. Vind [1640]. Vine [2926]. Vinikas [1414]. Viola [1994]. Viollet [734]. Viollet-le-Duc [734]. Virchow [939]. Virchows [940]. Virendra [2599]. Virgin [2681]. Virginia [2453, 166, 2776]. Virology [1505]. Virtues [1315]. viruelas [3154]. Viruses [3233]. Vision [2007, 2187, 1593, 2804, 3324, 2500, 3517, 728, 1117, 444]. Visionary [1745, 434]. Visions [623, 180, 268, 478, 2226, 3120, 869, 2665]. Visual [2777, 3029, 2009, 3133, 2262, 1492, 892, 2859]. Visuality [3126]. Visualization [3348]. Vita [3208, 2443]. Vital [997, 1893, 1576]. vitale [2507]. Vitalii [3090]. Vitalist [7]. Vitaly [2120]. Vitamania [2936]. Vitamin [1232]. Vitamins [2936]. Vitezslav [2909]. Vitrac [1176, 2242, 549]. vivente [1826]. Vives [637]. Vivian [110, 540, 2412, 886]. Vizgin [2197]. Vizionar [930]. Vladimir [3369, 2197, 1134]. Vlastos [2327]. Voegelin [1325]. Vogel [348, 1625]. Vogt [3375, 3458]. Voice [1617, 3175, 2489]. Void [915, 2246, 672]. voie [657]. Voigt [324]. Vol [3104, 2273, 1941]. volgens [2515]. Volk [2554]. Volker [2894, 2474, 1500, 2413, 354]. Volkerpsychologie [3273]. Volksheilstätten [374]. volmaking [994]. volontarismo [2340]. Volpi [3545]. Volta [1827]. Voltaic [2953]. Voltaire [1478]. Volti [1902]. Volume [819, 2329, 2368, 400, 798, 1157, 1522, 1881, 2219, 2558, 2863, 3194, 3197, 3593, 2661, 3249, 3049, 2979, 2032, 2604, 622, 3276, 1673, 852, 2449, 2834, 835, 2429, 602, 2491, 1080, 1111, 1716, 378, 1176, 2242, 1580, 3534, 2975, 1280, 1816, 3118, 2338, 1817, 3146, 1474, 1476, 2438, 768, 2460, 2895, 830, 1169, 1808, 3136, 329, 269, 1748, 372, 2192, 1947, 386, 2063, 2080, 438, 431, 2369, 2291, 1089, 2512, 1452, 2145, 2935, 2974, 3144, 1938, 643, 2609, 1812, 423, 3137, 1360, 3248, 2422, 2972, 2117, 2766, 766, 2456, 1359, 32, 2074, 2243, 3132, 3263]. Volume [1131, 255, 1487, 1573, 2970, 2910, 3490, 2685, 2008, 1892, 212, 3250, 1877, 2991, 586, 1046, 1939, 2182, 2520, 3540, 354, 3319, 1531]. Voluntarism [2283]. Vom [2646, 2755, 812, 851, 1927, 1215, 2185, 1283, 2647, 1660, 3118]. vondsten [2814]. voor [516]. Voprosy [511]. vor [1635, 2978, 1013, 549, 3535, 1278]. Vorgeschichte [325, 163]. vorindustriellen [3256]. Vorlesung [1651]. Vorlesungen [1403, 812, 2984]. Vormarz [1572]. Vorsey [2159]. Vortices [1306]. Vortrage [2620, 2185, 1120]. Vortragsreihe [324]. Vos [996]. vous [3088]. Voyage [260]. Voyager [1450, 3574]. Voyages [3142, 2665, 1663]. Voyaging [2875]. Vrat [1907]. Vrettos [2524]. Vries [3308]. Vucinich [1170]. vue [1026]. Vulgar [713]. vulgarisation [1312, 1571]. Vulgarly [2898]. vychislitel [910]. Vygotsky [1673, 1049]. Vyverberg [560].

W [161, 2214, 1325, 1295, 50, 736, 1673, 571, 160, 1617, 2303, 1998, 2115, 2252, 3349, 3448, 2285, 885, 1954, 2442, 3410, 1800, 669, 1896, 2454, 2183, 2163, 2173, 2553, 3058, 2125, 2998, 1399, 1953, 2926, 2734, 1860, 688, 1761, 1645, 1844, 380, 1169, 1174, 2160, 569, 285, 3331, 2063, 2934, 1217, 1087, 2198, 2467, 725, 1539, 1906, 2327, 3072, 572, 1085, 211, 120, 2234, 2774, 598, 232, 2670, 2675, 1379, 2572, 2857, 1868, 2257, 772, 1309, 1663, 1264, 1862, 260, 2697, 3283, 1110, 1243, 2243, 473, 1131, 225, 1826, 1478, 3174, 1573, 682, 506, 2388, 3430]. W [2292, 2544, 1550, 1802, 452, 265, 839, 1531, 1850, 3455]. waarheid [1368]. Wachstum [2413]. wachtwoord [2693]. Wack [826]. Wad [200]. Wade [3517]. Waerden [145]. Waffentechnik [2252]. Wagener [1008]. Wagenitz [90]. Wagner [1846, 2240, 348]. Wagner-Jauregg [1846]. wahrend [1941]. Wahrscheinlichkeitsrechnung [3535]. Wailoo [3289]. Waiser [265]. Wake [1018, 1663]. Walcott [3571]. Waldby [2782]. Walewska [184]. Wali [863]. Walker [2924, 873, 1355, 1867]. Walking [3282]. Wall [590, 275]. Wallace [1943, 3159, 1133, 3285, 1434, 803, 959, 1103, 3382, 3016, 1674, 1818]. Wallis [1928, 493, 1928]. Wallops [3382]. Walls [2525]. Walsh [253]. Walter [1996, 3003, 2646, 2445, 2026, 2933, 3331, 1588, 950, 1237, 319, 1989, 943, 3535, 1509, 1950, 2022, 1599, 2923, 384, 1045]. Walters [1478]. Wanda [2799]. Wang [1767]. War [2214, 582, 1954, 1502, 2542, 3559, 2056, 2651, 1585, 218, 2540, 95, 2543, 182, 3100, 425, 1416, 2297, 3498, 894, 3559, 277, 3004, 3081, 2612, 2855]. Warburg [1762]. Warburgs [3570]. Ward [1825, 2366, 3291]. Wardy [1358]. ware [3441]. Warfare [2210, 588]. Wargo [3288]. Warm [1463]. Warner [3267]. Warren [2527, 235, 2196, 686, 2685, 3547]. Warren-Findley [2685]. Wars [962, 2469, 1594, 2325]. Warsh [865]. Warshaw [3576]. Wartime [1952]. Was [3079, 1929, 225, 1134, 2447]. waschen [3567]. Waschkies [163]. Washburn [3388]. Washington [85, 888, 500, 1258]. Watanabe [3318, 995]. Watchers [1411]. watching [3519]. Watchword [2693]. Water [2872, 1205, 3061, 628, 245, 2663]. Waterfield [635]. Waters [2167, 2202, 1493]. Watson [285, 3374, 2352, 1388, 1066, 1226, 2116]. Wattenberg [2016]. Watts [3124]. Wave [140, 586, 173]. Waves [1059, 2023]. Way [2570, 2119, 1853]. Wayman [20]. Wayne [732]. Wealth [1806, 592]. Weapon [2056]. Weapons [2607, 2213, 2855, 60, 3364]. Wear [1978, 241, 1457, 353, 2412]. Weart [1865, 225]. Weather [193]. Weaver [1293, 621, 225]. Weaving [1293]. Webb [2388, 1983, 1758]. Weber [489, 3391, 2587, 461, 2519]. Weberian [2105]. Webster [2956, 3423, 636, 3423]. Wechselverhaltnisses [3253]. Wechselwirkung [1316]. Wechsler [2885, 1458]. Wedekind [461]. Wedekinds [461]. Weeks [2894, 1285]. Weg [2818]. Wege [755]. Wegeler [471]. Wegener [2852]. Wehefritz [3003]. Weick [197]. Weidner [1941]. Weierstrass [2195]. Weighing [2273]. Weights [814]. Weigl [1292]. Weikersheim [1724]. Weil [1466, 2008, 2008]. Weil-Gueny [1466]. Weimarer [2914, 1322]. Weimer [2885]. Weinberg [1866, 2205]. Weindling [2024, 676]. Weiner [1419, 2981, 2012, 1404]. Weingart [196]. Weintraub [1869]. Weir [1133]. Weird [1959]. Weisgall [2040]. Weiss [2923, 1046, 2923]. Weisser [518]. Weisskopf [1617, 61]. Weisz [2602]. weite [2033]. Weland [1113]. Welch [2026]. Well [137]. Well-Fed [137]. Wellborn [957]. Wellcome [3490, 2935]. Wellens [1649]. Wellens-De [1649]. Weller [2063]. Welling [1300]. Wellman [1387]. Wellmann [1766]. Wells [1761]. Welt [740, 2742]. Weltauffassung [2107]. Weltbild [1635, 1278]. Weltburger [3003].

Welteislehre [2300]. Weltkarte [1278]. Weltkrieg [2535]. Welzijn [458]. Wende [1927]. Wendell [2721]. Wenjao [507]. Wenzel [364]. Werck [453]. Werck-Schul [453]. Werdegang [461]. Were [3141, 225]. wereld [2680]. Werk [2894, 2123, 1841, 1300]. Werke [2979, 2356, 1474, 2438, 473, 2008, 354]. Werke/Collected [473]. Werkes [1113]. Werkh [3256]. Werkkonkordanz [1069]. Werner [2754, 1617, 1762, 1398, 473, 1381, 518, 3570, 1842]. Wertheim [2325]. Wesencraft [2655]. Wesley [783, 1778]. Wess [2276]. Wessels [1601]. West [2386, 2578, 3513, 2990, 3095, 2891, 15, 3232, 3318, 1798, 3347, 1710]. Western [338, 2150, 1702, 2773, 1776, 1080, 714, 2734, 2412, 3420, 3394, 995, 2961, 1752, 814, 1700, 903, 1632, 2328, 434, 419, 332, 2409, 131, 599, 1273, 2621, 1797, 1077, 807, 3114, 3042, 301, 3423]. Westfall [2624, 1731, 2125, 649]. Westman [448, 1100]. Wet [3554, 2663]. Wetenschap [228, 127, 2515, 458, 2101, 2292]. wetenschappelijke [361]. wetenschapsbeoefening [228]. Wetenschapsgeschiedenis [127]. Wetenschapsrevolutie [2814]. Weyer [1371, 1724]. Weyl [2528]. Wheatley [579]. Wheaton [1795]. Wheeler [119]. Where [583, 2578, 3159, 864]. Wherein [3528]. Whether [1922]. Whewell [1832, 1488, 1305, 2907]. Which [1922, 2663]. Whigs [2176]. Whistle [1958]. Whistle-Blower [1958]. White [1491, 542, 1545, 3577, 593, 273]. Whitehaven [1825]. Whitehead [297, 766]. Whiteside [2266]. Whitfield [231]. Whitney [155, 917, 2802]. Whitrow [1846, 125]. Whittet [493]. Who [1213, 2301, 3547, 1853, 1852, 883]. Whose [1062, 1062]. Wider [3265]. Widmaier [2265]. Wiebe [2703]. Wiederkehr [1120]. wieku [1655]. Wieland [1223]. Wiener [584, 1681]. Wiesing [2754]. Wife [938]. Wijngaard [2846]. wijsbegeerte [994, 458]. Wiktor [1982]. Wilbur [723]. Wilcomb [3388]. Wilcox [2759]. Wilda [933]. Wildberg [342, 915, 147]. Wilderness [2548, 908, 2802]. Wildlife [477]. Wiley [76]. Wilfried [1842]. Wilhelm [2661, 1481, 2754, 3367, 2356, 3080, 2922, 3531, 2508, 167, 837, 174, 89, 1841, 1375, 1215, 2089, 3137, 756, 2265, 2185, 260, 1531, 2016]. Wilkins [319]. Wilkinson [1459]. Wilks [2500]. Will [803, 3064, 2261]. Willem [1637, 2702, 458]. Willey [591, 3261]. Willhoft [2935]. William [2368, 2180, 264, 1859, 1623, 3252, 2788, 571, 420, 3356, 3295, 3349, 207, 3473, 1956, 110, 118, 1832, 1491, 2169, 1817, 1103, 2542, 583, 1450, 2372, 1921, 2337, 1087, 2749, 2804, 1825, 904, 1504, 2043, 1906, 2538, 572, 3123, 2655, 387, 3016, 1192, 1319, 2997, 2344, 1497, 176, 1030, 1104, 3155, 303, 83, 2609, 2424, 1465, 186, 1818, 644, 1458, 1243, 2907, 225, 1876, 1689, 1092, 1088, 1650, 1584, 2643, 2910, 3574, 1578, 2925, 2601, 1581, 96, 218, 655, 1439, 2507, 1488, 1305, 155, 2672, 1406]. Williams [482, 1045, 2285, 3142, 3254, 773, 468, 666, 787, 2014, 537, 225, 966, 2610, 1544]. Williamson [2685]. Wills [1144]. Wilma [1370]. Wilson [1564, 559, 37, 1726, 905, 628, 3228, 3152, 3390, 1333, 2263, 3034, 2407, 739]. Wilton [2354, 2175]. Wiltshire [1575, 936]. Winau [89, 518, 1705]. Wind [734, 2738, 1903, 2852, 1443]. Windmill [1903]. Wine [2872, 2926].

Winfried [2505]. Wing [3468]. Wingless [3482]. Wings [2773, 3480, 3480]. Winkler [2041]. Winks [3283]. Winnard [1895]. Winner [1777]. Winning [2029]. Winsor [1202]. Winston [3515]. Winterbourne [164]. Winters [1338]. wir [3567]. Wires [2792]. Wirklichkeit [1292]. Wirkung [842, 1841]. Wirkungsgeschichte [2251]. Wirtschaft [1856]. Wisconsin [896, 1047]. Wisdom [356]. Wise [846]. Wiskunde [516]. Wismuttherapie [2614]. Wissens [3241]. Wissenschaft [2754, 2757, 1808, 297, 2523, 12, 1841, 1640, 712, 2131, 472, 578]. Wissenschaften [851, 3367, 1993, 654, 3338, 1701, 1215, 510, 3367, 1830]. wissenschaftlich [2107]. wissenschaftlich-philosophische [2107]. wissenschaftliche [837]. wissenschaftlichen [3241, 259, 1120, 394]. Wissenschaftlicher [2834, 1932, 81, 2291]. Wissenschaftlichkeit [1122]. Wissenschaftsbeziehungen [3552, 755]. Wissenschaftsentwicklung [2046]. Wissenschaftsgeschichte [81, 3586, 518, 517, 2154, 518, 530]. Wissenschaftshistorischer [2356]. Wissenschaftskritik [2620]. Wissenschaftsphilosophie [786]. Wissenschaftsphilosophische [1393]. Wissenschaftspolitik [1304]. wissenschaftstheoretische [2614]. Wissenschaftstheorie [81, 2614]. Wissink [1007]. Witches [1371]. Witelo [1634, 1634]. Witelonis [1634]. within [1490, 2372, 3293, 3035]. without [1773, 1273, 1170]. Witkop [267]. Witness [151]. Witnesses [2935]. Witnessing [2360]. Wittenberg [1123]. Wittich [448]. Wives [3113]. Wizard [3271]. Wizards [2856]. Wohl [2773]. Wojcik [3058]. Wojtkowiak [1671]. Woker [3567]. Woldeck [1113]. Woldemar [3049]. Wolf [2884, 1308, 2262, 2374]. Wolf-Devine [2262]. Wolf-Dieter [2884]. Wolfe [1323, 3176, 3336]. Wolfenbuttel [3150]. Wolfgang [2834, 748, 2757, 189, 2965, 89, 3441, 2291, 371, 2415, 2901, 1724, 578, 374, 755, 456, 2834]. Wolfschmidt [2411]. Wollen [3567]. Wollner [2540]. wollte [3441]. Woman [1028, 639, 765, 1799, 946]. Womanhood [370]. Womans [1248]. Women [1981, 3079, 1262, 3553, 2751, 1604, 1310, 446, 3005, 3489, 2200, 1062, 2367, 1538, 2339, 383, 3386, 2927, 1352, 625, 764, 2939, 1037, 1245, 1384, 2711, 3474, 2706, 3174, 955, 3292, 1273, 3505, 946, 2943, 1528, 1564, 3317, 1715, 1878, 2467, 626, 2066, 1352, 1963, 1037, 1245, 1175, 2349, 3428, 2706, 2232]. Wonders [2548, 3404]. Wood [1537, 954, 3480, 3148]. Woodforde [3259]. Woodring [574]. Woods [323]. Woodward [1452, 2145]. Wooldridge [2287, 3463]. Woolgar [82, 208]. Woolhouse [2173, 2085]. Word [3249, 1295, 2429, 497, 2962]. Words [1081, 1099]. Work [1967, 2962, 3050, 287, 1666, 2175, 2998, 2993, 2835, 153, 2372, 2371, 1506, 2847, 750, 3342, 3286, 680, 3468, 766, 1861, 1289, 1315, 2202, 1227, 282, 1041, 2303, 1833, 2440, 1098]. Workers [680]. Working [1641, 3564, 3292, 2540]. Works [1395, 3276, 1673, 605, 3543, 835, 2532, 1816, 3331, 431, 2043, 3393, 1607, 617, 1373, 1518, 473]. Workshop [3140, 1947]. Workshops [697]. World [2214, 2293, 3284, 2962, 1594, 3274, 3043, 1483, 911, 1007, 1105, 2156, 3529, 1034, 897, 915, 710, 3559, 2146, 1223, 2410, 2767, 1585, 3436, 263, 3032, 1030, 478, 2540, 151, 192, 2543, 335, 1190, 2705, 1551, 21, 497, 2612, 2396, 936, 2244, 2263, 188, 225, 2159, 2022, 2203, 617, 1273, 1180, 3330, 1416, 254, 2138, 2056, 95, 544, 2376, 342, 1774, 1451, 1983, 2917, 1837, 1363, 551, 2261, 1720, 19, 1989, 862, 137, 3055, 385, 3185, 1844]. Worlds [632, 3159, 2525, 1687, 65, 1109, 1450]. Worldview [1541]. Worldwide [1584]. Wormald [2003]. Worstelende [228]. Worster [1806, 2064, 136]. Worth [633, 3153, 3345, 1153, 2850]. Worthen [1177]. Wosk [1492]. Wounded [946]. Wozniak [2116]. Wrestling [3225]. Wright [869, 2801, 3500, 2781, 968, 2995, 2730]. Wrighting [882]. Wrigley [45]. Writing [1099, 542, 3005, 2337, 151, 1987, 2758, 808, 722, 3130, 1515, 721]. Writings [1111, 2759, 1893, 1012, 768, 2460, 2594, 2367, 1906, 2117, 2766, 2243, 2889, 1943, 2017, 2664, 1550]. Written [2962, 428]. Wroclaw [731]. Wrong [1703]. Wujastyk [3490]. Wuketits [2464]. Wundt [3273]. wurde [549]. Württemberg [325]. Wurzburger [2754]. Wurzen [3535]. Wussing [517, 530, 3118]. Wyatt [91]. Wybrow [1536, 1536]. Wyndham [2192]. Wynes [479]. Wytze [2515].

X [3025]. Xavier [718, 1917]. Xe [916, 1809]. Xia [1186]. XIe [2070]. XII [835, 1810, 1183]. XIII [1634, 2890, 638]. XIII-XV [638]. XIV [355]. XIVe [1093]. XIX [3154, 1354, 1828]. XIXe [322, 3075, 2820, 1390]. XV [638]. XVe [2333]. XVI [1470, 2510, 1354]. XVII [3054, 2590, 1655, 2733, 255, 1470, 2435]. XVIIe [322, 1296, 2432, 1023, 1729, 1026, 256, 2900]. XVIII [1828, 1737, 3154, 2444, 262]. XVIIIe [535, 1296, 1931, 1026]. XX [3244, 354, 3475]. XX-XXIII [3244]. XXIII [3244]. xylographisches [348].

Ya [2835, 2124, 1852]. Yabe [507]. Yakov [2835]. Yakovlevich [2715]. Yamamoto [2426]. Yanjiu [507]. Yankee [2714]. Yankees [1038]. Yannick [1037, 212]. Yano [2426, 3432]. Yans [827]. Yapp [1370]. Yaron [1330]. Yasu [902, 3569]. Yasuo [980]. Yates [3370, 2422]. Yavetz [2371]. Year [3255, 2681, 911, 2425, 2452]. Yearbook [516]. Years [906, 68, 1702, 959, 2641, 2026, 2125, 3211, 768, 2460, 583, 1087, 2369, 360, 323, 2117, 2766, 3218, 188, 2776, 2753, 2029, 314, 601, 2678, 199, 3452, 1573, 2039]. Yeats [3168]. Yehuda [350, 108]. Yell [1825]. Yellow [3153, 795, 1035, 1934, 1934]. Yemeni [2073]. Yeo [1832]. Yeomans [991]. Yerkes [2931]. Yes [3085]. Yeu [1767]. Yeu-Farn [1767]. Yih [128, 3050]. Yochelson [3571]. Yoder [257]. Yolanda [1084]. Yolton [2572]. Yong [1712, 507]. Yorgos [488, 1214]. York [3163, 3326, 816, 2675, 142, 13, 3576, 3077, 672, 3209, 704, 60, 3583]. Yosemite [3282]. Yoshida [1625, 2536]. Yoshikawa [2536]. Young [568, 1275, 375, 2236, 2472, 670]. Younger [2080]. Yours [1786]. Youssef [1667]. Youthful [1475]. Youtie [2733]. Ysidro [2121]. Yu [2918]. Yuan [1186]. Yuasa [980]. Yuichi [3177]. Yusuf [3525]. Yves [2580, 2338, 823, 2895, 925, 229, 1133, 3459].

Z [87, 1686, 2569, 1269, 2023, 3260, 173]. Zabusky [2921]. Zacharias [1503]. Zachary [3366]. Zahar [391]. Zahl [1808]. Zahlenbedeutungen [345]. Zahlentheorie [1531]. Zahlstein [3118]. Zaire [1508]. Zambelli [1642, 1643, 254]. Zamorensis [2249]. Zande [3533]. Zanella [1704]. Zangger [1697]. Zanier [443, 2578]. Zanine [1861]. zapadnoi [887]. Zarka [925]. Zbigniew [2663]. zbiorach [3455]. Zebrowski [2967]. Zeeberg [2428]. Zeeland [228]. Zeeman [1777]. Zehn [812]. Zeit [3375, 1841, 394, 1122, 2818, 3149, 1808, 2016, 758]. Zeitalter [2110, 461]. Zeitalters [2046]. Zeitbegriffs [2818]. Zeitgenossen [1314]. Zeitschrift [1316, 518, 510]. Zellenlehre [654]. Zeller [47]. Zemon [2349]. Zenderland [3377]. Zeno [418]. Zeremonialwissenschaft [3341]. Zerner [1484]. Zerubavel [1648]. zestiende [228]. Zev [1290]. zeventiende [1297]. Zhabotinsky [654]. Zhou [2732]. Ziauddin [640]. Ziegler [59, 2238]. ziemiach [1655]. Ziggelaar [2816]. Zij [431]. zijn [2693]. Ziman [299]. Zimmerman [870]. Zimmermann [2894]. Zinner [442, 1098]. Ziporyn [376]. Ziran [507]. Znajomosc [1655]. znanosti [930]. Zola [2361]. Zolo [789]. Zonabend [981]. Zoological [1665, 1392]. zoologico [3442]. Zoologischen [1463, 1941]. Zoology [3112, 1202]. Zott [3080]. Zousmer [50]. Zuckerman [1604]. Zugleich [266, 1640]. Zuidervaart [2515]. Zum [656, 1660, 1927, 654, 439, 3118, 3256, 712, 1637, 2415]. Zuñiga [1888]. Zupko [2738, 814]. Zur [1115, 2679, 297, 374, 2843, 1923, 1481, 736, 517, 3367, 2920, 3246, 1996, 2646, 1274, 654, 536, 266, 439, 1185, 2620, 259, 656, 549, 3360, 1808, 174, 1932, 2755, 3256, 2046, 1120, 998, 730, 1640, 81, 394, 3341, 1941, 2070, 3570, 822, 2359, 842, 3091, 1283, 2818, 1587, 812, 654, 1322, 585, 2615]. Zürich [536]. Zusammenarbeit [1500, 3552]. zusammengestellt [2123]. Zweckbronner [325]. Zwei [2583]. Zweiter [549, 1996]. zwischen [2914, 1316, 851, 1993, 2757, 2095, 1500, 3253, 2195, 2185, 3552, 2033, 2465, 2747]. zwolf [2067]. Zysk [1364].

References

Neu:1990:CBH

[1] John Neu. Current bibliography of the history of science and its cultural influences 1990. Isis, 81(??):1–268, ???? 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

      stable/234436.

Anonymous:1990:BMa

[2] Anonymous. Back matter. Isis, 81(??):??, ???? 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

      /www.jstor.org/stable/234437.

Anonymous:1990:FMa

[3] Anonymous. Front matter. Isis, 81(??):i–xvi, ???? 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234435.

Anonymous:1990:FMb

[4] Anonymous. Front matter. Isis, 81(1):1–7, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234079.

McCluskey:1990:GTM

[5] Stephen C. McCluskey. Gregory of tours, monastic timekeeping, and Early Christian attitudes to astronomy. Isis, 81(1):8–22, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234080.

Meli:1990:RCF

[6] Domenico Bertoloni Meli. The relativization of centrifugal force. Isis, 81(1):23–43, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234081.

Palladino:1990:SNL

[7] Paolo Palladino. Stereochemistry and the nature of life: Mechanist, vitalist, and evolutionary perspectives. Isis, 81(1):44–67, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234082.

Hutchison:1990:SGO

[8] Keith Hutchison. Sunspots, Galileo, and the orbit of the Earth. Isis, 81 (1):68–74, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234083. See

[3499].

Hughes:1990:LE

[9] Thomas P. Hughes and John Henry. Letters to the Editor. Isis, 81(1): 75–76, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234084.

Hannaway:1990:BRB

[10] Owen Hannaway. Book review: Mersenne and the Learning of the Schools. Isis, 81(1):77–80, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234085.

Douglas:1990:BRB

[11] Susan J. Douglas. Book review: The Social Construction of Technological Systems: New Directions in the Sociology and History of Technology. Isis, 81(1):80–83, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234086.

Kragh:1990:BRB

[12] Helge Kragh. Book review: Lohn und Strafe in der Wissenschaft: Naturforscher im Urteil der Geschichte by Franz Stuhlhofer. Isis, 81(1):84–85, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234087.

Sokal:1990:BRB

[13] Michael M. Sokal. Book review: New York Intellect: A History of Intellectual Life in New York City, from 1750 to the Beginning of Our Own Time by Thomas Bender. Isis, 81(1):85, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234088.

Lux:1990:BRB

[14] David S. Lux. Book review: Lyon: Cité de savants. Isis, 81(1):85– 86, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234089.

Cullen:1990:BRB

[15] Christopher Cullen. Book review: Time, Science, and Society in China and the West by J. T. Fraser; N. Lawrence; F. C. Haber. Isis, 81(1): 86–87, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234090.

Calinger:1990:BRBa

[16] Ronald L. Calinger. Book review: Mathematics and the Search for Knowledge by Morris Kline. Isis, 81(1):87–88, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234091.

Yoder:1990:BRB

[17] Joella Yoder. Book review: Tales of Physicists and Mathematicians by Semyon Grigorevich Gindikin; Alan Shuchat. Isis, 81(1):88–89, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234092.

Lyon:1990:BRB

[18] John Lyon. Book review: The Physicist as Artist: The Landscapes of Pierre Duhem by Stanley L. Jaki. Isis, 81(1):89–90, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234093.

Meadows:1990:BRB

[19] A. J. Meadows. Book review: Astronomical Centers of the World by Kevin Krisciunas. Isis, 81(1):90–91, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234094.

Dick:1990:BRBa

[20] Steven J. Dick. Book review: Dunsink Observatory, 1785–1985: A Bicentennial History by Patrick A. Wayman. Isis, 81(1):91–92, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234095.

Rhees:1990:BRB

[21] David J. Rhees. Book review: A Material World: An Exhibition at the National Museum of American History by Robert Friedel. Isis, 81(1): 92–93, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234096.

Pyenson:1990:BRB

[22] Lewis Pyenson. Book review: The Tentacles of Progress: Technology Transfer in the Age of Imperialism, 1850–1940 by Daniel R. Headrick. Isis, 81(1):93, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234097.

Wise:1990:BRB

[23] George Wise. Book review: The Challenge of New Technology: Innovation in British Industry since 1850 by Jonathan Liebenau. Isis, 81(1): 93–94, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234098.

Evans:1990:BRB

[24] James Evans. Book review: Ancient Astronomy and Civilization by Norriss S. Hetherington. Isis, 81(1):94–95, March 1990. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234099.

Allen:1990:BRBa

[25] Ralph Allen. Book review: The Beginning of the Use of Metals and Alloys by Robert Maddin. Isis, 81(1):95–96, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234100.

Scarborough:1990:BRBa

[26] John Scarborough. Book review: Short Handbook of Hippocratic Medicine by Paul Potter. Isis, 81(1):96–97, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234101.

Osborne:1990:BRB

[27] Catherine Osborne. Book review: Matter, Space, and Motion: Theories in Antiquity and Their Sequel by Richard Sorabji. Isis, 81(1):97–98, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234102.

Majumdar:1990:BRB

[28] Pradip Kumar Majumdar. Book review: Mathematics in the Making in Ancient India: Reprints of “On the ’Sulva-sutras” and “Baudhyayana ’Sulva-sutra” by G. Thibaut; Debiprasad Chattopadhyaya. Isis, 81(1): 98–99, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234103.

Dales:1990:BRBa

[29] Richard C. Dales. Book review: The Medieval English Universities: Oxford and Cambridge to c. 1500 by Alan B. Cobban. Isis, 81(1):99, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234104.

Amundsen:1990:BRBa

[30] Darrel W. Amundsen. Book review: Doctors and Medicine in Medieval England, 1340–1530 by Robert S. Gottfried. Isis, 81(1):100–101, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234105.

Thibodeau:1990:BRB

[31] Sharon Gibbs Thibodeau. Book review: Islamic Astronomical Instruments by David A. King. Isis, 81(1):101–102, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234106.

Savage-Smith:1990:BRB

[32] Emilie Savage-Smith. Book review: From Deferent to Equant: A Volume of Studies in the History of Science in the Ancient and Medieval Near East in Honor of E. S. Kennedy by David A. King; George Saliba. Isis, 81(1):102–103, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234107.

Vickers:1990:BRB

[33] Brian Vickers. Book review: Picatrix: The Latin Version of the Ghayat al-h.akim by David Pingree. Isis, 81(1):103–104, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234108.

Kremer:1990:BRB

[34] Richard Kremer. Book review: Sundials and Related Instruments by David J. Bryden; The Ivory Sundials of Nuremberg, 1500–1700 by Penelope Gouk. Isis, 81(1):104–105, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234109.

Langermann:1990:BRB

[35] Y. Tzvi Langermann. Book review: Jewish Thought and the Scientific Revolution of the Sixteenth Century: David Gans (1541–1613) and His Times by André Neher; David Maisel. Isis, 81(1):105–107, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234110.

Goddu:1990:BRBa

[36] André Goddu. Book review: Descartes’s Theory of Light and Refraction: A Discourse on Method by A. Mark Smith. Isis, 81(1):107–108, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234111.

Christianson:1990:BRB

[37] Gale E. Christianson. Book review: Let Newton Be! by John Fauvel; Raymond Flood; Michael Shortland; Robin Wilson. Isis, 81(1):109, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234112.

Guerrini:1990:BRB

[38] Anita Guerrini. Book review: Newton’s Dream by Marcia Sweet Stayer. Isis, 81(1):109–110, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234113.

Bates:1990:BRB

[39] Don G. Bates. Book review: Thomae Sydenham Methodus curandi febres, propriis observationibus superstructa: The Latin Text of the 1666 and 1668 Editions with English Translation from R. G. Latham (1848) by Thomas Sydenham. Isis, 81(1):110–111, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234114.

Ross:1990:BRB

[40] Joseph T. Ross. Book review: Das Lob der Sternkunst: Astronomie in der deutschen Aufklärung by Rainer Baasner. Isis, 81(1):111–112, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234115.

Vanpaemel:1990:BRB

[41] G. H. W. Vanpaemel. Book review: Schets der Leere van Lavoisier door Martinus van Marum by Antoine-Laurent Lavoisier; Martinus van Marum. Isis, 81(1):113, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234116.

Michler:1990:BRB

[42] Markwart Michler. Book review: Johannes Gessners Pariser Tagebuch, 1727 by Johannes Gessner; Urs Boschung. Isis, 81(1):113–114, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234117.

Schiebinger:1990:BRB

[43] Londa Schiebinger. Book review: ’Tis Nature’s Fault: Unauthorized Sexuality during the Enlightenment by Robert Purks Maccubbin; Sexual Underworlds of the Enlightenment by G. S. Rousseau; Roy Porter. Isis, 81(1):114–115, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234118.

Jeremy:1990:BRB

[44] David J. Jeremy. Book review: Inventing the Industrial Revolution: The English Patent System, 1660–1800 by Christine MacLeod. Isis, 81(1):

115–116, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234119.

Lubar:1990:BRB

[45] Steven Lubar. Book review: Continuity, Chance and Change: The Character of the Industrial Revolution in England by E. A. Wrigley. Isis, 81 (1):116–117, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234120.

Turner:1990:BRBa

[46] Frank M. Turner. Book review: Science and Religion: Baden Powell and the Anglican Debate, 1800–1860 by Pietro Corsi. Isis, 81(1):117– 118, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234121.

Sheets-Pyenson:1990:BRBa

[47] Susan Sheets-Pyenson. Book review: Inventing Canada: Early Victorian Science and the Idea of a Transcontinental Nation by Suzanne Zeller. Isis, 81(1):118–119, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234122.

Greene:1990:BRBa

[48] Mott T. Greene. Book review: Controversy in Victorian Geology: The Cambrian-Silurian Dispute by James A. Secord. Isis, 81(1):119–120, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234123.

Burlingame:1990:BRBa

[49] Leslie J. Burlingame. Book review: The Age of Lamarck: Evolutionary Theories in France, 1790–1830 by Pietro Corsi; Jonathan Mandelbaum. Isis, 81(1):120–122, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234124.

Bowman:1990:BRB

[50] Robert I. Bowman. Book review: Galápagos: Discovery on Darwin’s Islands by David W. Steadman; Steven Zousmer. Isis, 81(1):122–123, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234125.

Molella:1990:BRB

[51] Arthur P. Molella. Book review: G. T. Fechner and Psychology by Josef Brožek; Horst Grundlach; Index Psychophysicus: Bio- und bibliographischer Index zu Fechners Elementen der Psychophysik und den Parerga by Horst Grundlach. Isis, 81(1):124, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234126.

Blustein:1990:BRB

[52] Bonnie Ellen Blustein. Book review: Before Freud: Neurasthenia and the American Medical Community, 1870–1910 by F. G. Gosling. Isis, 81(1):124–125, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234127.

Melosh:1990:BRB

[53] Barbara Melosh. Book review: Private Matters: American Attitudes toward Childbearing and Infant Nurture in the Urban North, 1800– 1860 by Sylvia D. Hoffert. Isis, 81(1):125–127, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234128.

Warner:1990:BRB

[54] Sam Bass Warner, Jr. Book review: Technology and the Rise of the Networked City in Europe and America by Joel A. Tarr; Gabriel Dupuy. Isis, 81(1):127–128, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234129.

Rabkin:1990:BRBa

[55] Yakov M. Rabkin. Book review: The Communist Party and Soviet Science by Stephen Fortescue. Isis, 81(1):128–129, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234130.

Holl:1990:BRB

[56] Jack M. Holl. Book review: Nucleus: The History of Atomic Energy Canada, Limited by Robert Bothwell. Isis, 81(1):129–130, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234131.

Aaserud:1990:BRB

[57] Finn Aaserud. Book review: Harmony and Unity: The Life of Niels Bohr by Niels Blædel. Isis, 81(1):130–131, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234132.

Cahan:1990:BRB

[58] David Cahan. Book review: The Comparative Reception of Relativity by Thomas F. Glick. Isis, 81(1):131–132, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234133.

Moore:1990:BRBa

[59] Walter Moore. Book review: Janossy Lajos (1912–1978) es Erwin Schrödinger (1887–1961) levelse by Peter Kirdly; Maria Narayne Ziegler. Isis, 81(1):132–133, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234134.

Winkler:1990:BRB

[60] Allan M. Winkler. Book review: Making Weapons, Talking Peace: A Physicist’s Odyssey from Hiroshima to Geneva by Herbert F. York. Isis, 81(1):133–134, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234135.

Westfall:1990:BRB

[61] Catherine Westfall. Book review: The Privilege of Being a Physicist by Victor F. Weisskopf. Isis, 81(1):134, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234136.

Fruton:1990:BRB

[62] Joseph S. Fruton. Book review: Otto Folin: America’s First Clinical Biochemist by Samuel Meites. Isis, 81(1):134–135, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234137.

Mazumdar:1990:BRBa

[63] Pauline M. H. Mazumdar. Book review: Milestones in Immunology: A Historical Exploration by Debra Jan Bibel. Isis, 81(1):135–136, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234138.

Gortler:1990:BRB

[64] Leon Gortler. Book review: Academic Scientists and the Pharmaceutical Industry: Cooperative Research in Twentieth–Century America by John P. Swann. Isis, 81(1):136–137, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234139.

Stocking:1990:BRBa

[65] George W. Stocking, Jr. Book review: Malinowski between Two Worlds: The Polish Roots of an Anthropological Tradition by Roy Ellen; Ernest Gellner; Graznya Kubica; Janusz Mucha. Isis, 81(1):137–138, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234140.

Coon:1990:BRB

[66] Deborah J. Coon. Book review: The Rise of Experimentation in American Psychology by Jill G. Morawski. Isis, 81(1):138–140, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234141.

Gregory:1990:BRBa

[67] Geoff Gregory. Book review: The “Discovery” of Sudden Infant Death Syndrome: Lessons in the Practice of Political Medicine by Abraham B. Bergman. Isis, 81(1):140, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234142.

Bynum:1990:BRB

[68] W. F. Bynum. Book review: Grand Rounds: One Hundred Years of Internal Medicine by Russell C. Maulitz; Diana E. Long. Isis, 81(1): 141, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234143.

Morman:1990:BRBa

[69] Edward T. Morman. Book review: The Dread Disease: Cancer and Modern American Culture by James T. Patterson. Isis, 81(1):141–142, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234144.

Farr:1990:BRB

[70] Gail E. Farr. Book review: Ordered to Care: The Dilemma of American Nursing, 1850–1945 by Susan M. Reverby. Isis, 81(1):142–144, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234145.

Woolgar:1990:BRB

[71] Steve Woolgar. Book review: The Artificial Intelligence Debate: False Starts, Real Foundations by Stephen R. Graubard. Isis, 81(1):144–145, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234146.

Levinson:1990:BRB

[72] Mark Levinson. Book review: IUTAM: A Short History by Stephen Juhasz. Isis, 81(1):145–146, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234147.

Tatarewicz:1990:BRB

[73] Joseph N. Tatarewicz. Book review: The Final Frontier: The Rise and Fall of the American Rocket State by Dale Carter. Isis, 81(1):146–147, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234148.

Ezell:1990:BRB

[74] Edward C. Ezell. Book review: Before Lift-Off: The Making of a Space Shuttle Crew by Henry S. F. Cooper, Jr. Isis, 81(1):147–148, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234149.

Scull:1990:BRB

[75] Andrew Scull. Book review: The System of Professions: An Essay on the Division of Expert Labor by Andrew Abbott. Isis, 81(1):148–149, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234150.

Dupre:1990:BRB

[76] John Dupré. Book review: Evolution as Entropy: Toward a Unified Theory of Biology by Daniel R. Brooks; E. O. Wiley. Isis, 81(1):149– 150, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234151.

Lynch:1990:BRB

[77] Michael Lynch. Book review: Contemporary Science and Natural Explanation: Commonsense Conceptions of Causality by Denis J. Hilton. Isis, 81(1):150–151, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234152.

Cao:1990:BRB

[78] Tian Yu Cao. Book review: The Description of Nature: Niels Bohr and the Philosophy of Quantum Physics by John Honner. Isis, 81(1):151– 152, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234153.

Tang:1990:BRB

[79] Paul C. L. Tang. Book review: Introductory Readings in the Philosophy of Science by E. D. Klemke; Robert Hollinger; A. David Kline. Isis, 81(1):152–153, March 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234154.

Darden:1990:BRB

[80] Lindley Darden. Book review: Computational Philosophy of Science by Paul Thagard. Isis, 81(1):153–154, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234155.

Poser:1990:BRB

[81] Hans Poser. Book review: Struktur und Dynamik wissenschaftlicher Theorien: Beiträge zur Wissenschaftsgeschichte und Wissenschaftstheorie aus der bulgarischen Forschung by Christfried Tögel; Hristo Smolenov; Anguel Stefanov. Isis, 81(1):154–155, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234156.

Bloor:1990:BRB

[82] David Bloor. Book review: Knowledge and Reflexivity: New Frontiers in the Sociology of Knowledge by Steve Woolgar. Isis, 81(1):155–156, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234157.

Neville:1990:BRBa

[83] Roy G. Neville. Book review: Chemical Literature, 1700–1860: A Bibliography with Annotations, Detailed Descriptions, Comparisons and Locations by William A. Cole. Isis, 81(1):156–157, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234158.

Amundson:1990:BRB

[84] Ron Amundson. Book review: Dictionary of Concepts in the Philosophy of Science by Paul T. Durbin. Isis, 81(1):157–158, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234159.

Karrow:1990:BRB

[85] Robert W. Karrow, Jr. Book review: Scholar’s Guide to Washington, D.C., for Cartography and Remote Sensing Imagery (Maps, Charts, Aerial Photographs, Satellite Images, Cartographic Literature, and Geographic Information Systems) by Ralph E. Ehrenberg. Isis, 81(1):158– 159, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234160.

Sapp:1990:BRB

[86] Jan Sapp. Book review: A Guide to the Genetics Collection of the American Philosophical Society by Bentley Glass. Isis, 81(1):159, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234161.

Blaisdell:1990:BRB

[87] Muriel Blaisdell. Book review: A History of the Hope Entomological Collections in the University Museum, Oxford, with Lists of Archives and Collections by Audrey Z. Smith. Isis, 81(1):159–160, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234162.

Millbrooke:1990:BRBa

[88] Anne Millbrooke. Book review: A Chronology of Geological Thinking from Antiquity to 1899 by Susan J. Thompson. Isis, 81(1):160–161, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234163.

Johnson:1990:BRBa

[89] Jeffrey A. Johnson. Book review: Berlinische Lebensbilder by Wilhelm Treue; Gerhard Hildebrandt; Rolf Winau; Wolfgang Treue; Karlfried Gründer. Isis, 81(1):161–163, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234164.

Nyhart:1990:BRB

[90] Lynn K. Nyhart. Book review: Göttinger Biologen 1737–1945: Eine biographisch-bibliographische Liste by Gerhard Wagenitz. Isis, 81(1): 163, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234165.

Maienschein:1990:BRB

[91] Jane Maienschein. Book review: Information Sources in the Life Sciences by H. V. Wyatt. Isis, 81(1):164, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234166.

Anonymous:1990:Ca

[92] Anonymous. Collections. Isis, 81(1):164–173, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234167.

Anonymous:1990:BMb

[93] Anonymous. Back matter. Isis, 81(1):174–180, March 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234168.

Anonymous:1990:FMc

[94] Anonymous. Front matter. Isis, 81(2):181–187, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233682.

Owens:1990:FAA

[95] Larry Owens. MIT and the Federal “angel”: Academic R&D and federal– private cooperation before World War II. Isis, 81(2):188–213, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233683.

Gillespie:1990:DDI

[96] Neal C. Gillespie. Divine design and the industrial revolution: William Paley’s abortive reform of natural theology. Isis, 81(2):214–229, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233684.

Biagioli:1990:GEM

[97] Mario Biagioli. Galileo the emblem maker. Isis, 81(2):230–258, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233685.

Perrin:1990:CPP

[98] C. E. Perrin. Chemistry as peer of physics: A response to Donovan and Melhado on Lavoisier. Isis, 81(2):259–270, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233686. See [99, 100].

Donovan:1990:LCE

[99] Arthur Donovan. Lavoisier as chemist and experimental physicist: A reply to Perrin. Isis, 81(2):270–272, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233687. See [98].

Melhado:1990:HSR

[100] Evan M. Melhado. On the historiography of science: A reply to Perrin. Isis, 81(2):273–276, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233688.

See [99, 98].

Cohen:1990:EDS

[101] I. Bernard Cohen. Eloge: Dorothy Stimson, 10 October 1890–19 September 1988. Isis, 81(2):277–278, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233689.

Hentschel:1990:LE

[102] Klaus Hentschel. Letters to the Editor. Isis, 81(2):279–280, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233690.

Knight:1990:BRB

[103] David Knight. Book review: Annals of Science: An International Review of the History of Science and Technology from the Thirteenth Century by G. L’E. Turner. Isis, 81(2):281–282, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233691.

Wilson:1990:BRBa

[104] David B. Wilson. Book review: British Journal for the History of Science by John Brooke. Isis, 81(2):282–285, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233692.

Broman:1990:BRBa

[105] Thomas Broman. Book review: History of Science by R. S. Porter. Isis, 81(2):285–287, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233693.

Fox:1990:BRB

[106] Robert Fox. Book review: Isis: An International Review Devoted to the History of Science and Its Cultural Influences by Ronald L. Numbers. Isis, 81(2):287–288, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233694.

Cohen:1990:BRBa

[107] I. Bernard Cohen. Book review: Osiris: A Research Journal Devoted to the History of Science and Its Cultural Influences by Arnold Thackray. Isis, 81(2):288–289, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233695.

Secord:1990:BRB

[108] James A. Secord. Book review: Science in Context by Robert S. Cohen; Yehuda Elkana; Simon Schaffer; Gad Freudenthal. Isis, 81(2):289–290, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233696.

Laudan:1990:BRBa

[109] Rachel Laudan. Book review: Studies in History and Philosophy of Science by Nicholas Jardine. Isis, 81(2):290–292, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233697.

Eyler:1990:BRB

[110] John M. Eyler. Book review: Bulletin of the History of Medicine by Caroline Hannaway; Journal of the History of Medicine and Allied Sciences by Robert U. Massey; Medical History by William F. Bynum; Vivian Nutton; Social History of Medicine by Anne Digby; Richard Smith; Lynda Bryder. Isis, 81(2):292–293, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233698.

Friedel:1990:BRB

[111] Robert Friedel. Book review: Technology and Culture: The International Quarterly of the Society for the History of Technology by Robert C. Post; History of Technology by A. Rupert Hall; Norman A. F. Smith; History and Technology: An International Journal by Pietro Redondi. Isis, 81(2): 294–296, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233699.

Fraser:1990:BRB

[112] Craig G. Fraser. Book review: Archive for History of Exact Sciences by Clifford Truesdell. Isis, 81(2):296–297, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233700.

Rider:1990:BRBa

[113] Robin E. Rider. Book review: Historia Mathematica by Eberhard Knobloch. Isis, 81(2):297–298, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233701.

VanHelden:1990:BRBa

[114] Albert Van Helden. Book review: Journal for the History of Astronomy by M. A. Hoskin. Isis, 81(2):298–299, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233702.

Burchfield:1990:BRBa

[115] Joe D. Burchfield. Book review: Historical Studies in the Physical and Biological Sciences by J. L. Heilbron. Isis, 81(2):299–300, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233703.

Sturchio:1990:BRB

[116] Jeffrey L. Sturchio. Book review: Ambix. Journal of the Society for the History of Alchemy and Chemistry by Michael A. Sutton. Isis, 81(2): 300–302, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233704.

Oldroyd:1990:BRB

[117] David Oldroyd. Book review: Earth Sciences History by Gerald M. Friedman. Isis, 81(2):302–303, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233705.

Farley:1990:BRB

[118] John Farley. Book review: Journal of the History of Biology by Everett Mendelsohn; Studies in History of Biology by William Coleman; Camille Limoges. Isis, 81(2):303–304, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233706.

Stephens:1990:BRBa

[119] Lester D. Stephens. Book review: Archives of Natural History by Alwyne Wheeler. Isis, 81(2):304–305, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233707.

Mark:1990:BRBa

[120] Joan Mark. Book review: History of Anthropology by George W. Stocking. Isis, 81(2):305–306, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233708.

Cravens:1990:BRB

[121] Hamilton Cravens. Book review: Journal of the History of the Behavioral Sciences by Barbara Ross; History of the Human Sciences by Arthur Still; Irving Velody. Isis, 81(2):306–308, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233709.

Shea:1990:BRBa

[122] William R. Shea. Book review: Storia della scienza moderna e contemporanea by Paolo Rossi. Isis, 81(2):309, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233710.

MacLachlan:1990:BRB

[123] James MacLachlan. Book review: The History of Science and Technology: A Narrative Chronology. Isis, 81(2):310–311, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233711.

Bevilacqua:1990:BRB

[124] Fabio Bevilacqua. Book review: Storicità e attualità della cultura scientifica e insegnamento delle scienze by C.I.D.I. di Firenze. Isis, 81(2): 311–312, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233712.

Stephens:1990:BRBb

[125] Carlene E. Stephens. Book review: Time in History: The Evolution of Our General Awareness of Time and Temporal Perspective by G. J. Whitrow. Isis, 81(2):312–313, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233713.

Matthews:1990:BRB

[126] Fred Matthews. Book review: That Noble Dream: The “Objectivity Question” and the American Historical Profession by Peter Novick. Isis, 81(2):313–314, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233714.

Turner:1990:BRBb

[127] G. L’E. Turner. Book review: Instrument in de Wetenschap: Bijdragen tot de Instrumentgerichte Wetenschapsgeschiedenis by Marian Fournier; Bert Theunissen. Isis, 81(2):314–315, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233715.

Major:1990:BRB

[128] John S. Major. Book review: Science and Technology in Chinese Civilization by Cheng-Yih Chen; Roger Cliff; Kuei-Mei Chen. Isis, 81(2): 315–316, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233716.

Farquhar:1990:BRB

[129] Judith Farquhar. Book review: Traditional Medicine in Contemporary China by Nathan Sivin. Isis, 81(2):316–317, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233717.

Allen:1990:BRBb

[130] Donald S. Allen. Book review: Empire of Reason: Exact Sciences in Indonesia, 1840–1940 by Lewis Pyenson. Isis, 81(2):317–318, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233718.

Sheets-Pyenson:1990:BRBb

[131] Susan Sheets-Pyenson. Book review: Nature in Its Greatest Extent: Western Science in the Pacific by Roy MacLeod; Philip F. Rehbock. Isis, 81(2):318, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233719.

Bedini:1990:BRB

[132] Silvio A. Bedini. Book review: El grabado en la ciencia hispánica by José Maria López Piñero. Isis, 81(2):318–319, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233720.

Hofmann:1990:BRB

[133] James R. Hofmann. Book review: The How and the Why: An Essay on the Origins and Development of Physical Theory by David Park. Isis, 81(2):319–320, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233721.

Elena:1990:BRB

[134] Alberto Elena. Book review: Historia del interés anglosajón por la geología de España by Estanislao Ribera i Faig. Isis, 81(2):320–321, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233722.

Smith:1990:BRBa

[135] David C. Smith. Book review: The College of Agriculture at Penn State: A Tradition of Excellence by Michael Bezilla. Isis, 81(2):321–322, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233723.

Opie:1990:BRB

[136] John Opie. Book review: The Ends of the Earth: Perspectives on Modern Environmental History by Donald Worster. Isis, 81(2):322–323, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233724.

Rossiter:1990:BRB

[137] Margaret W. Rossiter. Book review: Toward a Well-Fed World by Don Paarlberg. Isis, 81(2):323, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233725.

Apple:1990:BRBa

[138] Rima D. Apple. Book review: Revolution at the Table: The Transformation of the American Diet by Harvey A. Levenstein. Isis, 81(2):323–324, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233726.

Granshaw:1990:BRBa

[139] Lindsay Granshaw. Book review: The Health of a City: Oxford, 1790– 1974 by Jessie Parfit. Isis, 81(2):324–325, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233727.

Aitken:1990:BRBa

[140] Hugh G. J. Aitken. Book review: The Carrier Wave: New Information Technology and the Geography of Innovation, 1846–2003 by Peter Hall; Paschal Preston. Isis, 81(2):325–327, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233728.

Outram:1990:BRB

[141] Dorinda Outram. Book review: Cultural History: Between Practices and Representations by Roger Chartier; Lydia G. Cochrane. Isis, 81(2):327– 328, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233729.

Rainger:1990:BRB

[142] Ronald Rainger. Book review: Power and Society: Greater New York at the Turn of the Century by David C. Hammack. Isis, 81(2):328–329, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233730.

Todd:1990:BRB

[143] Edmund N. Todd. Book review: Another Germany: A Reconsideration of the Imperial Era by Jack R. Dukes; Joachim Remak. Isis, 81(2):329– 330, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233731.

McCluskey:1990:BRB

[144] Stephen McCluskey. Book review: Records in Stone: Papers in Memory of Alexander Thom by C. L. N. Ruggles. Isis, 81(2):330–331, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233732.

Jones:1990:BRB

[145] Alexander Jones. Book review: Die Astronomie der Griechen: Eine Einführung by B. L. van der Waerden. Isis, 81(2):331–332, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233733.

Scarborough:1990:BRBb

[146] John Scarborough. Book review: Les lapidaires grecs: Lapidaire Orphique; Kérygmes lapidaires d’Orphée; Socrate et Denys; Lapidaire nautique; Damigéron–Evax by Robert Halleux; Jacques Schamp. Isis, 81(2): 332–334, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233734.

Hahm:1990:BRB

[147] David E. Hahm. Book review: John Philoponus’ Criticism of Aristotle’s Theory of Aether by Christian Wildberg. Isis, 81(2):334–335, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233735.

Cadden:1990:BRB

[148] Joan Cadden. Book review: Sexuality and Medicine in the Middle Ages by Danielle Jacquart; Claude Thomasset; Matthew Adamson. Isis, 81(2): 335–336, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233736.

DeYoung:1990:BRB

[149] Gregg De Young. Book review: La théorie des parallèles en pays d’Islam: Contribution à la préhistoire des géométries non-euclidiennes by K. Jaouiche. Isis, 81(2):336–337, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233737.

Jaouiche:1990:BRB

[150] K. Jaouiche. Book review: Entre arithmétique et algèbre: Recherches sur l’histoire des mathématiques arabes by Roshdi Rashed. Isis, 81(2): 337–338, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233738.

Park:1990:BRB

[151] Katharine Park. Book review: The Witness and the Other World: Exotic European Travel Writing, 400–1600 by Mary B. Campbell. Isis, 81(2): 338–339, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233739.

Lamb:1990:BRB

[152] Ursula Lamb. Book review: El arte de navegar en la España del Renacimiento by José Maria López Piñero. Isis, 81(2):339–340, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233740.

Hatfield:1990:BRBa

[153] Gary Hatfield. Book review: Theories at Work: On the Structure and Functioning of Theories in Science, in Particular during the Copernican Revolution by Marinus Dirk Stafleu. Isis, 81(2):340–341, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233741.

Keller:1990:BRB

[154] A. G. Keller. Book review: Un autor aragonés para “Los veintiun libros de los ingenios y de las máquinas.” by Juan A. Frago Gracia; José García–Diego. Isis, 81(2):341–342, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233742.

Reeds:1990:BRBa

[155] Karen Reeds. Book review: William Turner: Tudor Naturalist, Physician, and Divine by Whitney R. D. Jones. Isis, 81(2):342–343, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233743.

Sawyer:1990:BRB

[156] Ronald C. Sawyer. Book review: Kabbalah, Magic, and Science: The Cultural Universe of a Sixteenth-Century Jewish Physician by David B. Ruderman. Isis, 81(2):343–344, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233744.

Finocchiaro:1990:BRB

[157] Maurice A. Finocchiaro. Book review: Carteggio by Benedetto Castelli; Massimo Bucciantini. Isis, 81(2):344–345, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233745.

Gingerich:1990:BRB

[158] Owen Gingerich. Book review: Kepler’s Physical Astronomy by Bruce Stephenson. Isis, 81(2):345–346, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233746.

Grabiner:1990:BRB

[159] Judith V. Grabiner. Book review: The Origins of the Infinitesimal Calculus by Margaret E. Baron. Isis, 81(2):346–347, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233747.

Calinger:1990:BRBb

[160] Ronald Calinger. Book review: 300 Jahre “Nova methodus” von G. W. Leibniz (1684–1984): Symposion der Leibniz-Gesellschaft im Congresscentrum “Leewenhorst” in Noordwijkerhout, Niederlande, 28. bis 30. August 1984 by Albert Heinekamp. Isis, 81(2):347–349, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233748.

Aiton:1990:BRB

[161] E. J. Aiton. Book review: G. W. Leibniz: Iter italicum (mars 1689– mars 1690): La dynamique de la République des lettres: Nombreux textes inédits by André Robinet. Isis, 81(2):349, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233749.

Neville:1990:BRBb

[162] Roy G. Neville. Book review: Johann Christoph Ettner: Eine beschreibende Bibliographie by James N. Hardin. Isis, 81(2):349–350, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233750.

Kerszberg:1990:BRB

[163] Pierre Kerszberg. Book review: Physik und Physikotheologie des jungen Kant: Die Vorgeschichte seiner Allgemeinen Naturgeschichte und Theorie des Himmels by Hans-Joachim Waschkies. Isis, 81(2):350–351, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233751.

West:1990:BRB

[164] Grant West. Book review: The Ideal and the Real: An Outline of Kant’s Theory of Space, Time, and Mathematical Construction by A. T. Winterbourne. Isis, 81(2):351–352, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233752.

Struik:1990:BRB

[165] D. J. Struik. Book review: Optica in de eeuw van Euler: Opvattingen over de Natuur van het Licht, 1700–1795 by Casper Hakfoort. Isis, 81(2): 352–354, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233753.

Roe:1990:BRBa

[166] Shirley A. Roe. Book review: Nature’s Enigma: The Problem of the Polyp in the Letters of Bonnet, Trembley, and Réaumur by Virginia P. Dawson. Isis, 81(2):354–355, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233754.

Gregory:1990:BRBb

[167] Frederick Gregory. Book review: Ideas for a Philosophy of Nature as Introduction to the Study of This Science, 1797: Second Edition, 1803 by Friedrich Wilhelm Joseph von Schelling; Errol E. Harris; Peter Heath. Isis, 81(2):355–356, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233755.

Strickland:1990:BRB

[168] Stuart Walker Strickland. Book review: Approaches to Organic Form: Permutations in Science and Culture by Frederick Burwick. Isis, 81(2): 356–357, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233756.

Wetzels:1990:BRB

[169] Walter D. Wetzels. Book review: Romanticism in National Context by Roy Porter; Mikulas Teich. Isis, 81(2):357–358, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233757.

Brock:1990:BRB

[170] W. H. Brock. Book review: Energy and Entropy: Science and Culture in Victorian Britain by Patrick Brantlinger. Isis, 81(2):358–359, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233758.

Sahli:1990:BRB

[171] Nancy Sahli. Book review: Consuming Desire: Sexual Science and the Emergence of a Culture of Abundance, 1871–1914 by Lawrence Birken. Isis, 81(2):359–360, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233759.

Crowe:1990:BRB

[172] Michael J. Crowe. Book review: The Peripatetic Astronomer: Charles Piazzi Smyth by H. A. Brück; M. T. Brück. Isis, 81(2):360–362, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233760.

Worrall:1990:BRB

[173] John Worrall. Book review: The Rise of the Wave Theory of Light: Optical Theory and Experiment in the Early Nineteenth Century by Jed Z. Buchwald. Isis, 81(2):362–363, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233761.

Hoffmann:1990:BRB

[174] Dieter Hoffmann. Book review: Betrachtungen über die bewegende Kraft des Feuers und die zur Entwicklung dieser Kraft geeigneten Maschinen by Sadi Carnot; Robert Fox; Wilhelm Ostwald. Isis, 81(2):363–364, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233762.

Rocke:1990:BRB

[175] A. J. Rocke. Book review: Hauswirtschaftliche Briefe: Erstes bis drittes Dutzend by Friedlieb Ferdinand Runge. Isis, 81(2):364, June 1990. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233763.

Millbrooke:1990:BRBb

[176] Anne Millbrooke. Book review: George William Featherstonhaugh: The First U.S. Government Geologist by Edmund Berkeley; Dorothy Smith Berkeley. Isis, 81(2):364–365, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233764.

Noll:1990:BRB

[177] Mark A. Noll. Book review: Darwin and the Divine in America: Protestant Intellectuals and Organic Evolution, 1859–1900 by Jon H. Roberts. Isis, 81(2):365–366, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233765.

Keeney:1990:BRB

[178] Elizabeth Barnaby Keeney. Book review: Asa Gray: American Botanist, Friend of Darwin by A. Hunter Dupree. Isis, 81(2):366–368, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233766.

Kohlstedt:1990:BRB

[179] Sally Gregory Kohlstedt. Book review: Cathedrals of Science: The Development of Colonial Natural History Museums during the Late Nineteenth Century by Susan Sheets-Pyenson. Isis, 81(2):368–369, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233767.

Cittadino:1990:BRB

[180] Eugene Cittadino. Book review: Pacific Visions: California Scientists and the Environment, 1850–1915 by Michael L. Smith. Isis, 81(2):369– 370, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233768.

Hildreth:1990:BRB

[181] Martha Hildreth. Book review: Console and Classify: The French Psychiatric Profession in the Nineteenth Century by Jan Goldstein. Isis, 81(2):370–371, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233769.

Rosenberg:1990:BRB

[182] Charles E. Rosenberg. Book review: “I Have Done My Duty”: Florence Nightingale in the Crimean War, 1854–56 by Sue M. Goldie. Isis, 81 (2):371, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233770.

Mauskopf:1990:BRB

[183] Seymour H. Mauskopf. Book review: Orfila: Saber y poder médico by Rafael Huertas. Isis, 81(2):371–372, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233771.

TePaske:1990:BRB

[184] John Jay TePaske. Book review: Beauperthuy: De Cumaná a la Academia de Ciencias de Paris by Walewska Lemoine; María Matilde Suárez. Isis, 81(2):372–373, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233772.

Harris:1990:BRBa

[185] Ruth R. Harris. Book review: Granville Sharp Pattison: Anatomist and Antagonist, 1791–1851 by F. L. M. Pattison. Isis, 81(2):373–374, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233773.

Porter:1990:BRBa

[186] Theodore M. Porter. Book review: The Politics of Numbers by William Alonso; Paul Starr. Isis, 81(2):374–375, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233774.

Drucker:1990:BRB

[187] Thomas Drucker. Book review: Enigmas of Chance: An Autobiography by Mark Kac. Isis, 81(2):375–376, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233775.

Schweber:1990:BRB

[188] Silvan S. Schweber. Book review: Symmetries, Asymmetries, and the World of Particles by T. D. Lee; Thirty Years since Parity Nonconservation: A Symposium for T. D. Lee by Robert Novick. Isis, 81(2): 376–377, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233776.

Gieser:1990:BRB

[189] Suzanne Gieser. Book review: Beyond the Atom: The Philosophical Thought of Wolfgang Pauli by K. V. Laurikainen. Isis, 81(2):377–378, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233777.

Faye:1990:BRB

[190] Jan Faye. Book review: Niels Bohr’s Philosophy of Physics by Dugald Murdoch. Isis, 81(2):378–379, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233778.

Rabkin:1990:BRBb

[191] Yakov M. Rabkin. Book review: Ideals and Realities: Selected Essays of Abdus Salam by Abdus Salam; C. H. Lai. Isis, 81(2):380–381, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233779.

Pickering:1990:BRB

[192] Andy Pickering. Book review: Beamtimes and Lifetimes: The World of High Energy Physicists by Sharon Traweek. Isis, 81(2):381–382, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233780.

Hansen:1990:BRBa

[193] James R. Hansen. Book review: Appropriating the Weather: Vilhelm Bjerknes and the Construction of a Modern Meterology by Robert Marc Friedman. Isis, 81(2):382–383, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233781.

Ghiselin:1990:BRB

[194] Michael T. Ghiselin. Book review: The Role of Behavior in Evolution by H. C. Plotkin. Isis, 81(2):383, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233782.

Harden:1990:BRB

[195] Victoria A. Harden. Book review: AIDS: The Burdens of History by Elizabeth Fee; Daniel M. Fox. Isis, 81(2):383–384, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233783.

Weiss:1990:BRB

[196] Sheila Faith Weiss. Book review: Rasse, Blut und Gene: Geschichte der Eugenik und Rassenhygiene in Deutschland by Peter Weingart; Jürgen Kroll; Kurt Bayertz. Isis, 81(2):384–386, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233784.

Bilstein:1990:BRB

[197] Roger E. Bilstein. Book review: From the Ground Up: The Autobiography of an Aeronautical Engineer by Fred E. Weick; James R. Hansen. Isis, 81(2):386–387, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233785.

Layton:1990:BRB

[198] Edwin T. Layton, Jr. Book review: The Master Spirit of the Age: Canadian Engineers and the Politics of Professionalism by J. Rodney Millard. Isis, 81(2):387, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233786.

Greb:1990:BRB

[199] G. Allen Greb. Book review: Stemming the Tide: Arms Control in the Johnson Years by Glenn T. Seaborg; Benjamin S. Loeb. Isis, 81(2):388– 389, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233787.

DeGregori:1990:BRB

[200] Thomas R. DeGregori. Book review: Science, Technology, and Development by Atul Wad. Isis, 81(2):389–390, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233788.

Avrith:1990:BRB

[201] Gale Avrith. Book review: Ruth Benedict: Stranger in This Land by Margaret M. Caffrey. Isis, 81(2):390–391, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233789.

Alchon:1990:BRB

[202] Guy Alchon. Book review: Memoirs of an Unregulated Economist by George J. Stigler. Isis, 81(2):392–393, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233790.

Norton:1990:BRB

[203] John D. Norton. Book review: Conceptions of Space and Time: Sources, Evolution, Directions by Murad D. Akhundov; Charles Rougle. Isis, 81 (2):393–394, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233791.

Ungar:1990:BRB

[204] A. M. Ungar. Book review: Rationality by Harold I. Brown. Isis, 81 (2):394, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233792.

Pitt:1990:BRB

[205] Joseph C. Pitt. Book review: Empirical Knowledge by Alan H. Goldman. Isis, 81(2):395, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233793.

Kleiner:1990:BRB

[206] Scott A. Kleiner. Book review: The Philosophies of Science: An Introductory Survey by Rom Harre. Isis, 81(2):395–396, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233794.

Davis:1990:BRBa

[207] Edward B. Davis. Book review: Physics, Philosophy, and Theology: A Common Quest for Understanding by Robert John Russell; William R. Stoeger; George V. Coyne. Isis, 81(2):396–397, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233795.

Dennis:1990:BRB

[208] Michael Aaron Dennis. Book review: Science: The Very Idea by Steve Woolgar. Isis, 81(2):397–398, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233796.

Feingold:1990:BRB

[209] Mordechai Feingold. Book review: The Archives of the Royal Society. Isis, 81(2):398–399, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233797.

Getz:1990:BRBa

[210] Faye Marie Getz. Book review: Bibliographie des textes médicaux latins: Antiquité et haut moyen âge by Centre Jean-Palerne. Isis, 81(2):399– 400, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233798.

Mahoney:1990:BRBa

[211] Michael S. Mahoney. Book review: Historical Dictionary of Data Processing by James W. Cortada. Isis, 81(2):400–401, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233799.

Vampola:1990:BRB

[212] David Vampola. Book review: Les professeurs du Collège de France: Dictionnaire biographique 1901–1939 by Christophe Charle; Eva Telkès; L’enseignement technique de la Révolution à nos jours. Volume I: 1789– 1926 by Thérèse Charmasson; Anne-Marie Lelorrain; Yannick Ripa. Isis, 81(2):401–402, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233800.

Bud:1990:BRB

[213] Robert Bud. Book review: Archives of the British Chemical Industry, 1750–1914: A Handlist by Peter J. T. Morris; Colin A. Russell. Isis, 81(2):402–403, June 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233801.

Anonymous:1990:Cb

[214] Anonymous. Collections. Isis, 81(2):404–408, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233802.

Anonymous:1990:BMc

[215] Anonymous. Back matter. Isis, 81(2):409–416, June 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233803.

Anonymous:1990:FMd

[216] Anonymous. Front matter. Isis, 81(3):417–423, September 1990. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233420.

Kohlstedt:1990:PPP

[217] Sally Gregory Kohlstedt. Parlors, primers, and public schooling: Education for science in Nineteenth-Century America. Isis, 81(3):424–445, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233421.

Mitman:1990:EGW

[218] Gregg Mitman. Evolution as gospel: William patten, the language of democracy, and the great war. Isis, 81(3):446–463, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233422.

Rashed:1990:PAI

[219] Roshdi Rashed. A pioneer in anaclastics: Ibn Sahl on burning mirrors and lenses. Isis, 81(3):464–491, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233423.

Golinski:1990:TPP

[220] Jan Golinski. The theory of practice and the practice of theory: Sociological approaches in the history of science. Isis, 81(3):492–505, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233424.

Brush:1990:AMH

[221] Stephen G. Brush, Michael M. Sokal, and Albert Moyer. Annual meeting of the History of Science Society, Gainesville, Florida, 26–29 October 1989. Isis, 81(3):506–512, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233425.

Anonymous:1990:CHS

[222] Anonymous. Contributors to the History of Science Society. Isis, 81 (3):512–513, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233426.

Greene:1990:EJG

[223] John Greene. Eloge: John Garrett Burke (12 August 1917–21 February 1989). Isis, 81(3):514–516, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233427.

Schulmann:1990:LE

[224] Robert Schulmann and John Stachel. Letters to the Editor. Isis, 81 (3):517–518, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233428.

Seidel:1990:BRB

[225] Robert Seidel. Book reviews: Books on the bomb. Atomic Bomb Scientists: Memoirs, 1939-1945 by Joseph J. Ermenc. The End of the World That Was: Six Lives in the Atomic Age by Peter Goldman. Manhattan: The Army and the Atomic Bomb by Vincent C. Jones. Day of the Bomb: Countdown to Hiroshima by Dan Kurzman. The General and the Bomb: A Biography of General Leslie R. Groves, Director of the Manhattan Project by William Lawren. Time Bomb: Fermi, Heisenberg, and the Race for the Atomic Bomb by Malcolm C. MacPherson. The Making of the Atomic Age by Alwyn McKay. The Road to Trinity: A Personal Account of How America’s Nuclear Policies Were Made by K. D. Nichols. The Making of the Atomic Bomb by Richard Rhodes. Stallion Gate by Martin Cruz Smith. The Atomic Scientists: A Biographical History by Henry A. Boorse. Lloyd Motz, and Jefferson Hane Weaver. Forging the Atomic Shield: Excerpts from the Office Diary of Gordon E. Dean by Gordon E. Dean and Roger M. Anders. The Nuclear Oracles: A Political History of the General Advisory Committee of the Atomic Energy Commission, 1947-1977 by Richard T. Sylves. Better a Shield Than a Sword by Edward Teller. Klaus Fuchs, Atom Spy by Robert Chadwell Williams. Justice Downwind: America’s Atomic Testing Program in the 1950s by Howard Ball. The Atomic Papers: A Citizen’s Guide to Selected Books and Articles on the Bomb, the Arms Race, Nuclear Power, the Peace Movement, and Related Issues by Grant Burns. Physics, Technology and the Nuclear Arms Race by D. W. Hafemeister and D. Schroeer. Under the Cloud: The Decades of Nuclear Testing by Richard L. Miller. Bombs in the Backyard: Atomic Testing and American Politics by A. Costandina Titus. Nuclear Fear: A History of Images by Spencer R. Weart. Isis, 81 (3):519–537, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233429.

Lindberg:1990:BRB

[226] David C. Lindberg. Book review: The Savior of Science by Stanley L. Jaki. Isis, 81(3):538–539, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233430.

Heinz:1990:BRB

[227] Lisa C. Heinz. Book review: The Ends of Science: An Essay in Scientific Authority by Harry Redner. Isis, 81(3):539–540, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233431.

Hackmann:1990:BRBa

[228] Willem D. Hackmann. Book review: Worstelende Wetenschap: Aspecten van wetenschapsbeoefening in Zeeland van de zestiende tot in de negentiende eeuw. Isis, 81(3):540–541, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233432.

Mazumdar:1990:BRBb

[229] Pauline M. H. Mazumdar. Book review: Sciences et médecine au Québec: Perspectives sociohistoriques by Marcel Fournier; Yves Gingras; Othmar Keel. Isis, 81(3):541–542, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233433.

Moseley:1990:BRB

[230] Russell Moseley. Book review: The Reluctant Patron: Science and the State in Britain, 1850–1920 by Peter Alter; Angela Davies. Isis, 81(3): 542–543, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233434.

Granshaw:1990:BRBb

[231] Lindsay Granshaw. Book review: The College of Physicians of Philadelphia: A Bicentennial History by Whitfield J. Bell. Isis, 81(3):543–544, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233435.

Miller:1990:BRBa

[232] David Philip Miller. Book review: A Bicentenary History of the Linnean Society of London by A. T. Gage; W. T. Stearn. Isis, 81(3):544–545, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233436.

Calinger:1990:BRBc

[233] Ronald Calinger. Book review: The History of Mathematics: An Introduction by David M. Burton. Isis, 81(3):545–546, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233437.

Richards:1990:BRB

[234] Joan L. Richards. Book review: History in Mathematics Education by Ivor Grattan-Guinness. Isis, 81(3):546–547, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233438.

Greene:1990:BRBb

[235] Mott T. Greene. Book review: Theories of the Earth and Universe: A History of Dogma in the Earth Sciences by S. Warren Carey. Isis, 81 (3):547–549, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233439.

Appel:1990:BRB

[236] Toby A. Appel. Book review: Histoire du concept d’espèce dans les sciences de la vie. Isis, 81(3):549, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233440.

Kuriyama:1990:BRB

[237] Shigehisa Kuriyama. Book review: Fragments for a History of the Human Body by Michel Feher; Ramona Naddaff; Nadia Tazi. Isis, 81(3):550–551, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233441.

Paul:1990:BRB

[238] Diane Paul. Book review: Genethics: The Clash between the New Genetics and Human Values by David Suzuki; Peter Knudtson. Isis, 81 (3):551–552, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233442.

Egerton:1990:BRB

[239] Frank N. Egerton. Book review: Ecological Economics: Energy, Environment, and Society by Juan Martinez-Alier; Klaus Schlupmann. Isis, 81(3):552, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233443.

Cook:1990:BRB

[240] Harold J. Cook. Book review: Inleiding tot de Geschiedenis der Geneeskunde by G. A. Lindeboom. Isis, 81(3):553–554, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233444.

Broman:1990:BRBb

[241] Thomas Broman. Book review: Problems and Methods in the History of Medicine by Roy Porter; Andrew Wear. Isis, 81(3):554–555, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233445.

Tighe:1990:BRB

[242] Janet A. Tighe. Book review: Homicidal Insanity, 1800–1985 by Janet Colaizzi. Isis, 81(3):555–556, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233446.

Hansen:1990:BRBb

[243] James R. Hansen. Book review: The Evolution of Technology by George Basalla. Isis, 81(3):556–557, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233447.

Aitken:1990:BRBb

[244] Hugh G. J. Aitken. Book review: The Development of Large Technical Systems by Renate Mayntz; Thomas P. Hughes. Isis, 81(3):557–558, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233448.

TeBrake:1990:BRB

[245] William H. TeBrake. Book review: The Age of Water: The Urban Environment in the North of France, A.D. 300–1800 by André E. Guillerme. Isis, 81(3):558–559, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233449.

Sollors:1990:BRB

[246] Werner Sollors. Book review: The Railroad in American Art: Representations of Technological Change by Susan Danly; Leo Marx. Isis, 81 (3):559–561, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233450.

Goldstein:1990:BRB

[247] Bernard R. Goldstein. Book review: MULAPIN: An Astronomical Compendium in Cuneiform by Hermann Hunger; David Pingree. Isis, 81(3): 561–562, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233451.

Aaboe:1990:BRB

[248] Asger Aaboe. Book review: A Scientific Humanist: Studies in Memory of Abraham Sachs by Erle Leichty; Maria deJ. Ellis; Pamela Gerardi. Isis, 81(3):562–563, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233452.

McGovern:1990:BRB

[249] Patrick E. McGovern. Book review: The Royal Purple and the Biblical Blue (Argaman and Tekhelet): The Study of Chief Rabbi Dr. Isaac Herzog on the Dye Industries in Ancient Israel and Recent Scientific Contributions by Ehud Spanier. Isis, 81(3):563–565, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233453.

Amundsen:1990:BRBb

[250] Darrel W. Amundsen. Book review: Doctors and Diseases in the Roman Empire by Ralph Jackson. Isis, 81(3):565–566, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233454.

Taylor:1990:BRB

[251] Richard C. Taylor. Book review: La diffusione delle scienze islamische nel Medio Evo europeo. Isis, 81(3):566–567, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233455.

Voigts:1990:BRB

[252] Linda Ehrsam Voigts. Book review: Chaucer’s Universe by J. D. North. Isis, 81(3):567–569, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233456.

Eastwood:1990:BRB

[253] Bruce Eastwood. Book review: Cummian’s Letter De controversia Paschali, Together with a Related Irish Computistical Tract by Maura Walsh; Dáibhí Ó Cróinín. Isis, 81(3):569, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233457.

Westman:1990:BRB

[254] Robert S. Westman. Book review: “Astrologi hallucinati”: Stars and the End of the World in Luther’s Time by Paola Zambelli. Isis, 81(3): 569–571, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233458.

Shea:1990:BRBb

[255] William R. Shea. Book review: Correspondance du P. Marin Mersenne, religieux minime. Volume XVII: Supplements, tables et bibliographie by P. Marin Mersenne; Armand Beaulieu. Isis, 81(3):571–572, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233459.

VanHelden:1990:BRBb

[256] Albert Van Helden. Book review: Jean Picard et les débuts de l’astronomie de précision au XVIIe siècle by Guy Picolet. Isis, 81(3): 572–573, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233460.

Mahoney:1990:BRBb

[257] Michael S. Mahoney. Book review: Unrolling Time: Christiaan Huygens and the Mathematization of Nature by Joella G. Yoder. Isis, 81(3):573– 574, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233461.

Hatfield:1990:BRBb

[258] Gary Hatfield. Book review: Scottish Common Sense in Germany, 1768– 1800: A Contribution to the History of Critical Philosophy by Manfred Kuehn. Isis, 81(3):574–575, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233462.

Gregory:1990:BRBc

[259] Frederick Gregory. Book review: Texte zur Systematologie und zur Theorie der wissenschaftlichen Erkenntnis by Johann Heinrich Lambert; Geo Siegwart; Horst D. Brandt. Isis, 81(3):575–576, September 1990. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233463.

Rothenberg:1990:BRB

[260] Marc Rothenberg. Book review: Journal of a Voyage with Bering, 1741– 1742 by Georg Wilhelm Steller; O. W. Frost; Margritt A. Engel. Isis, 81(3):576, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233464.

Fireman:1990:BRB

[261] Janet R. Fireman. Book review: La ilusión quebrada: Botánica, sanidad y política científica en la España Ilustrada by Francisco Javier Puerto Sarmiento. Isis, 81(3):576–577, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233465.

Ringrose:1990:BRB

[262] David R. Ringrose. Book review: De Palas a Minerva: La formación científica y la estructura institucional de los ingenieros militares en el siglo XVIII by Horacio Capel; Joan Eugeni Sánchez; Omar Moncada. Isis, 81(3):577–578, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233466.

Maulitz:1990:BRB

[263] Russell C. Maulitz. Book review: Professional and Popular Medicine in France, 1770–1830: The Social World of Medical Practice by Matthew Ramsey. Isis, 81(3):579–580, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233467.

Baatz:1990:BRB

[264] Simon Baatz. Book review: The European Journals of William Maclure by William Maclure; John S. Doskey. Isis, 81(3):580–581, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233468.

Yochelson:1990:BRB

[265] Ellis L. Yochelson. Book review: The Field Naturalist: John Macoun, the Geological Survey, and Natural Science by W. A. Waiser. Isis, 81 (3):581–582, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233469.

Finlay:1990:BRB

[266] Mark R. Finlay. Book review: Justus von Liebig und Julius Eugen Schlossberger in ihren Briefen von 1844–1860: Zugleich ein Beitrag zur Geschichte physiologischen Chemie in Tübingen by Justus von Liebig; Julius Eugen Schlossberger; Fritz Hesse; Emil Heuser; Entwicklung und Institutionalisierung der Agriculturchemie im 19. Jahrhundert: Liebig und die Landwirtschaftlichen Versuchsstationen by Ursula Schling-Brodersen. Isis, 81(3):582–584, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233470.

Hiebert:1990:BRB

[267] Erwin N. Hiebert. Book review: Aus meinem Leben by Emil Fischer; Bernhard Witkop. Isis, 81(3):584–585, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233471.

Jongsma:1990:BRB

[268] Calvin Jongsma. Book review: Mathematical Visions: The Pursuit of Geometry in Victorian England by Joan L. Richards. Isis, 81(3):585–586, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233472.

Hodge:1990:BRB

[269] M. J. S. Hodge. Book review: The Correspondence of Charles Darwin. Volume IV: 1847–1850 by Charles Darwin; Frederick Burkhardt; Sydney Smith. Isis, 81(3):586–588, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233473.

Daston:1990:BRB

[270] Lorraine Daston. Book review: Déchiffrer la France: La statistique départementale à l’époque napoléonienne by Marie–Noëlle Bourguet. Isis, 81(3):588–589, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233474.

Porter:1990:BRBb

[271] Theodore M. Porter. Book review: Histoire du calcul économique en France by François Etner. Isis, 81(3):589–590, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233475.

Gillespie:1990:BRB

[272] Neal C. Gillespie. Book review: Josiah Nott of Mobile: Southerner, Physician, and Racial Theorist by Reginald Horsman. Isis, 81(3):590– 591, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233476.

Kimmelman:1990:BRB

[273] Barbara A. Kimmelman. Book review: White Trash: The Eugenic Family Studies, 1877–1919 by Nicole Hahn Rafter. Isis, 81(3):591–592, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233477.

Segal:1990:BRB

[274] Howard P. Segal. Book review: Nuts and Bolts of the Past: A History of American Technology, 1776–1860 by David Freeman Hawke. Isis, 81 (3):592–593, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233478.

Misa:1990:BRB

[275] Thomas J. Misa. Book review: Andrew Carnegie by Joseph Frazier Wall. Isis, 81(3):593–594, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233479.

Weiner:1990:BRB

[276] Douglas R. Weiner. Book review: Scientific Management, Socialist Discipline, and Soviet Power by Mark R. Beissinger. Isis, 81(3):594–595, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233480.

Kuznick:1990:BRB

[277] Peter J. Kuznick. Book review: Science, Politics, and the Cold War by Greta Jones. Isis, 81(3):596–597, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233481.

Murdoch:1990:BRB

[278] Dugald Murdoch. Book review: Einstein versus Bohr: The Continuing Controversies in Physics by Mendel Sachs. Isis, 81(3):597, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233482.

Friedman:1990:BRBa

[279] Robert Marc Friedman. Book review: Physical Chemistry of Colloids and Macromolecules by Bengt Rånby. Isis, 81(3):597–598, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233483.

Morris:1990:BRB

[280] Peter J. T. Morris. Book review: Petrochemicals: The Rise of an Industry by Peter H. Spitz. Isis, 81(3):599–600, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233484.

Benson:1990:BRB

[281] Keith R. Benson. Book review: Controlling Life: Jacques Loeb and the Engineering Ideal in Biology by Philip J. Pauly. Isis, 81(3):600–601, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233485.

Tobey:1990:BRB

[282] Ronald C. Tobey. Book review: A Sense of Place: The Life and Work of Forrest Shreve by Janice Emily Bowers. Isis, 81(3):601–602, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233486.

Mayr:1990:BRB

[283] Ernst Mayr. Book review: Arguments on Evolution: A Paleontologist’s Perspective by Antoni Hoffman. Isis, 81(3):602, September 1990. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233487.

Kitcher:1990:BRB

[284] Philip Kitcher. Book review: Did Darwin Get It Right? Essays on Games, Sex and Evolution by John Maynard Smith. Isis, 81(3):603–604, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233488.

Hunt:1990:BRB

[285] Elizabeth Hunt. Book review: Mechanical Man: John Broadus Watson and the Beginnings of Behaviorism by Kerry W. Buckley. Isis, 81(3): 604–605, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233489.

Ash:1990:BRB

[286] Mitchell G. Ash. Book review: Die Professionalisierung der deutschen Psychologie im Nationalsozialismus by Ulfried Geuter. Isis, 81(3):605– 606, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233490.

Darnell:1990:BRB

[287] Regna Darnell. Book review: The Culture Facade: Art, Science, and Politics in the Work of Oscar Lewis by Susan M. Rigdon. Isis, 81(3): 606–607, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233491.

Manier:1990:BRB

[288] Edward Manier. Book review: A Phylogenetic Fantasy: Overview of the Transference Neuroses by Sigmund Freud; Ilse Grubrich-Simitis; Axel Hoffer; Peter T. Hoffer. Isis, 81(3):607–608, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233492.

Morman:1990:BRBb

[289] Edward T. Morman. Book review: Integrating the City of Medicine: Blacks in Philadelphia Health Care, 1910–1965 by David McBride. Isis, 81(3):608–609, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233493.

Davis:1990:BRBb

[290] Audrey B. Davis. Book review: The Health Century by Edward Shorter. Isis, 81(3):609–610, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233494.

Bays:1990:BRB

[291] Daniel H. Bays. Book review: The Stubborn Earth: American Agriculturalists on Chinese Soil, 1898–1937 by Randall E. Stross. Isis, 81(3): 610–611, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233495.

Marvin:1990:BRB

[292] Carolyn Marvin. Book review: Inventing American Broadcasting, 1899– 1922 by Susan J. Douglas. Isis, 81(3):611–613, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233496.

Smith:1990:BRBb

[293] John Kenly Smith, Jr. Book review: From Monopoly to Competition: The Transformation of Alcoa, 1888–1986 by George David Smith. Isis, 81(3):613, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233497.

Laudan:1990:BRBb

[294] Rachel Laudan. Book review: Historicism and Knowledge by Robert D’Amico. Isis, 81(3):613–614, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233498.

Nickles:1990:BRBa

[295] Thomas Nickles. Book review: Relativism and Realism in Science by Robert Nola. Isis, 81(3):614–615, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233499.

McMahon:1990:BRB

[296] Michal McMahon. Book review: Contemporary Moral Controversies in Technology by A. Pablo Iannone. Isis, 81(3):615–616, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233500.

John:1990:BRB

[297] Peter C. John. Book review: Die organismische Kosmologie von Alfred N. Whitehead: Zur Revision des Selbstverständnisses neuzeitlicher Philosophie und Wissenschaft durch eine neue Philosophie der Natur by Alois Rust. Isis, 81(3):616–617, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233501.

Apple:1990:BRBb

[298] Michael W. Apple. Book review: Gramsci and the History of Dialectical Thought by Maurice A. Finocchiaro. Isis, 81(3):617–618, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233502.

Stehr:1990:BRB

[299] Nico Stehr. Book review: Knowing Everything about Nothing: Specialization and Change in Research Careers by John Ziman. Isis, 81(3): 618–619, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233503.

Elliott:1990:BRB

[300] Clark A. Elliott. Book review: Science in Society: An Annotated Guide to Resources by Institute for the History and Philosophy of Science and Technology. Isis, 81(3):619, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233504.

Mungello:1990:BRB

[301] D. E. Mungello. Book review: Western Books on China Published up to 1850 in the Library of the School of Oriental and African Studies, University of London: A Descriptive Catalogue by John Lust. Isis, 81 (3):619–620, September 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233505.

Burchfield:1990:BRBb

[302] Joe D. Burchfield. Book review: Victorian Britain: An Encyclopedia by Sally Mitchell. Isis, 81(3):620–621, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233506.

Nelson:1990:BRB

[303] Daniel Nelson. Book review: A Historical Dictionary of American Industrial Language by William H. Mulligan. Isis, 81(3):621–622, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233507.

Mack:1990:BRBa

[304] Pamela E. Mack. Book review: NASA Historical Data Book. (The NASA Historical Series.) by Jane Van Nimmen; Leonard C. Bruno; Robert L. Rosholt; Linda Neuman Ezell. Isis, 81(3):622, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233508.

Anonymous:1990:Cc

[305] Anonymous. Collections. Isis, 81(3):623–625, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233509.

Anonymous:1990:BMd

[306] Anonymous. Back matter. Isis, 81(3):626–630, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233510.

Anonymous:1990:FMe

[307] Anonymous. Front matter. Isis, 81(4):631–637, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233808.

Kohler:1990:PDM

[308] Robert E. Kohler. The Ph.D. machine: Building on the collegiate base. Isis, 81(4):638–662, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233809.

Dear:1990:MEO

[309] Peter Dear. Miracles, experiments, and the ordinary course of nature. Isis, 81(4):663–683, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233810.

Laird:1990:RGA

[310] Edgar S. Laird. Robert Grosseteste, Albumasar, and Medieval tidal theory. Isis, 81(4):684–694, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233811.

Naylor:1990:GMA

[311] Ronald H. Naylor. Galileo’s method of analysis and synthesis. Isis, 81 (4):695–707, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233812.

Shea:1990:EPC

[312] William R. Shea. Eloge: Pierre Costabel, 24 October 1912–20 November 1989. Isis, 81(4):708–709, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233813.

Mackay:1990:LE

[313] Alan Mackay, Maurice Crosland, Jack Neufeld, and Walter A. Mc-Dougall. Letters to the Editor. Isis, 81(4):710–712, December 1990.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233814.

Vallance:1990:MCG

[314] J. T. Vallance. Marshall Clagett’s Greek science in antiquity: Thirty-five years later. Isis, 81(4):713–721, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233815.

Sivin:1990:BRB

[315] Nathan Sivin. Book review: Medicine in China. Isis, 81(4):722–731, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233816.

Ezrahi:1990:BRB

[316] Yaron Ezrahi. Book review: The Nature of Power by Barry Barnes. Isis, 81(4):732–733, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233817.

Hacking:1990:BRB

[317] Ian Hacking. Book review: Theory and Experiment: Recent Insights and New Perspectives on Their Relation by Diderik Batens; Jean Paul van Bendegem. Isis, 81(4):733, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233818.

Olson:1990:BRB

[318] Richard Olson. Book review: Science and the Human Prospect by Ronald C. Pine. Isis, 81(4):733–734, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233819.

Moore:1990:BRBb

[319] James Moore. Book review: Science and Religious Thought: A Darwinism Case Study by Walter J. Wilkins. Isis, 81(4):734–735, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233820.

Melhado:1990:BRB

[320] Evan M. Melhado. Book review: Les chemins du savoir en Suède de la foundation de l’université d’Upsal à Jacob Berzelius: Etudes et portraits by Sten Lindroth; Jean–François Battail. Isis, 81(4):735–736, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233821.

Chambers:1990:BRB

[321] David Wade Chambers. Book review: Historia de las ciencias by Antonio Lafuente; Juan J. Saldaña; El perfil de la ciencia en América by Juan José Saldaña. Isis, 81(4):736–737, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233822.

Feldman:1990:BRB

[322] Theodore S. Feldman. Book review: Les savants génèvois dans l’Europe intellectuelle: Du XVIIe au milieu du XIXe siècle by Jacques Trembley. Isis, 81(4):737–738, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233823.

Mills:1990:BRB

[323] Eric Mills. Book review: 100 Years Exploring Life 1888–1988: The Marine Biological Laboratory at Woods Hole by Jane Maienschein. Isis, 81(4):738–739, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233824.

Johnson:1990:BRBb

[324] Jeffrey A. Johnson. Book review: Naturwissenschaften in Göttingen: Eine Vortragsreihe by Hans-Heinrich Voigt. Isis, 81(4):739–740, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233825.

Bolenz:1990:BRB

[325] Eckhard Bolenz. Book review: Ingenieurausbildung im Königreich Württemberg: Vorgeschichte, Einrichtung und Ausbau der Technischen Hochschule Stuttgart und ihrer Ingenieurwissenschaften bis 1900– Eine Verknüpfung von Institutions- und Disziplingeschichte by Gerhard Zweckbronner. Isis, 81(4):740–741, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233826.

Hackmann:1990:BRBb

[326] Willem D. Hackmann. Book review: Early Scientific Instruments: Europe, 1400–1800 by Anthony Turner. Isis, 81(4):742–743, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233827.

Battimelli:1990:BRB

[327] Giovanni Battimelli. Book review: Atti dell’VIII Congresso Nazionale di Storia della Fisica by Fabio Bevilacqua. Isis, 81(4):743–744, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233828.

Brush:1990:BRB

[328] Stephen G. Brush. Book review: A Companion to the Physical Sciences by David Knight. Isis, 81(4):744, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233829.

Hilts:1990:BRB

[329] Patricia Hilts. Book review: Science and Civilisation in China. Volume V: Chemistry and Chemical Technology. Part 9: Textile Technology: Spinning and Reeling by Joseph Needham; Dieter Kuhn. Isis, 81 (4):744–745, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233830.

Bartholomew:1990:BRB

[330] James R. Bartholomew. Book review: The Green Archipelago: Forestry in Preindustrial Japan by Conrad Totman. Isis, 81(4):745–746, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233831.

Gilman:1990:BRB

[331] Sander Gilman. Book review: The Normal and the Pathological by Georges Canguilhem; Carolyn R. Fawcett; Robert S. Cohen. Isis, 81 (4):746–748, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233832.

Risse:1990:BRB

[332] Guenter B. Risse. Book review: Imperial Medicine and Indigenous Societies by David Arnold; Disease, Medicine, and Empire: Perspectives on Western Medicine and the Experience of European Expansion by Roy MacLeod; Milton Lewis. Isis, 81(4):748–749, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233833.

Getz:1990:BRBb

[333] Faye M. Getz. Book review: Poisons of the Past: Molds, Epidemics, and History by Mary Kilbourne Matossian. Isis, 81(4):749–750, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233834.

Howell:1990:BRB

[334] Joel D. Howell. Book review: Doctors: The Biography of Medicine by Sherwin B. Nuland. Isis, 81(4):750–751, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233835.

Pressman:1990:BRB

[335] Jack D. Pressman. Book review: A Social History of Madness: The World through the Eyes of the Insane by Roy Porter. Isis, 81(4):751– 752, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233836.

Fitzgerald:1990:BRB

[336] Deborah Fitzgerald. Book review: Threshing in the Midwest, 1820–1940: A Study of Traditional Culture and Technological Change by J. Sanford Rikoon. Isis, 81(4):752–753, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233837.

Lemay:1990:BRB

[337] Richard Lemay. Book review: Horoscopes and History by J. D. North. Isis, 81(4):753–754, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233838.

Bowden:1990:BRB

[338] Mary Ellen Bowden. Book review: A History of Western Astrology by Jim Tester. Isis, 81(4):754–756, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233839.

Hall:1990:BRB

[339] Bert S. Hall. Book review: Les mécaniques, ou l’élévateur des corps lourds. The Arabic Text of Qust.a Ibn Luqa by Heron of Alexandria; B. Carra de Vaux. Isis, 81(4):756–757, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233840.

Toomer:1990:BRB

[340] G. J. Toomer. Book review: Las hipótesis de los planetas by Claudio Ptolomeo. Isis, 81(4):757–758, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233841.

Hershbell:1990:BRB

[341] Jackson P. Hershbell. Book review: Einführung in die antiken Naturwissenschaften by Alfred Stückelberger. Isis, 81(4):758–759, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233842.

Dales:1990:BRBb

[342] Richard C. Dales. Book review: Philoponus against Aristotle on the Eternity of the World by Johannes Philoponus; Christian Wildberg. Isis, 81(4):759, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233843.

Preus:1990:BRB

[343] Anthony Preus. Book review: Alfred of Sareshel’s Commentary on the Metheora of Aristotle by James K. Otte. Isis, 81(4):759–760, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233844.

Goddu:1990:BRBb

[344] André Goddu. Book review: Dialectic and Its Place in the Development of Medieval Logic by Eleonore Stump. Isis, 81(4):760–761, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233845.

Brey:1990:BRB

[345] Gerhard Brey. Book review: Lexikon der mittelalterlichen Zahlenbedeutungen by Heinz Meyer; Rudolf Suntrup. Isis, 81(4):761–762, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233846.

King:1990:BRB

[346] David A. King. Book review: Islamicate Celestial Globes: Their History, Construction, and Use by Emilie Savage-Smith; Andrea P. A. Belloli. Isis, 81(4):762–764, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233847.

Rider:1990:BRBb

[347] Robin E. Rider. Book review: Mathematics from Manuscript to Print, 1300–1600 by Cynthia Hay. Isis, 81(4):764, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233848.

VanEgmond:1990:BRB

[348] Warren Van Egmond. Book review: Das Bamberger Blockbuch: Inc. typ. 1c 144 der Staatsbibliothek Bamberg: Ein xylographisches Rechenbuch aus dem 15. Jahrhundert by Kurt Vogel; Bernhard Schemmel; Das Bamberger Rechenbuch von 1483 by Ulrich Wagner. Isis, 81(4):764–765, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233849.

Mahoney:1990:BRBc

[349] Michael S. Mahoney. Book review: The Early Theory of Equations: On Their Nature and Constitution: Translations of Three Treatises by Viète, Girard, and De Beaune by Robert Schmidt; Ellen Black; La nouvelle algèbre de M. Viète: Précédeé de Introduction en l’art analytique by J.-L. de Vaulézard. Isis, 81(4):765–766, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233850.

Reeds:1990:BRBb

[350] Karen Reeds. Book review: Albrecht Dürer and the Animal and Plant Studies of the Renaissance by Fritz Koreny; Pamela Marwood; Yehuda Shapiro. Isis, 81(4):766–768, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233851.

Kelter:1990:BRB

[351] Irving A. Kelter. Book review: Los temas polémicos de la medicina renacentista: Las Controversias (1556) de Francisco Valles by José Maria López Piñero; Francisco Calero. Isis, 81(4):768–769, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233852.

Cohen:1990:BRBb

[352] I. Bernard Cohen. Book review: A hombros de gigantes: Estudios sobre la primera revolución científica by Alberto Elena. Isis, 81(4):769–770, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233853.

Findlen:1990:BRB

[353] Paula Findlen. Book review: The Medical Revolution of the Seventeenth Century by Roger French; Andrew Wear. Isis, 81(4):770–771, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233854.

Wilson:1990:BRBb

[354] Curtis Wilson. Book review: Gesammelte Werke. Volume XX, Part 1: Manuscripta Astronomica (I) by Johannes Kepler; Volker Bialas; Friederike Boockmann. Isis, 81(4):771–772, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233855.

Nadler:1990:BRB

[355] Steven Nadler. Book review: Occult Powers and Hypotheses: Cartesian Natural Philosophy under Louis XIV by Desmond M. Clarke. Isis, 81 (4):772–773, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233856.

Osler:1990:BRB

[356] Margaret J. Osler. Book review: Thomas Hobbes: The Unity of Scientific and Moral Wisdom by Gary B. Herbert. Isis, 81(4):773–774, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233857.

Schneider:1990:BRB

[357] Ivo Schneider. Book review: Blaise Pascal, 1623–1662 by Hans Loeffel. Isis, 81(4):774, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233858.

Friedman:1990:BRBb

[358] Gerald M. Friedman. Book review: De solido intra solidum naturaliter contento dissertationis prodromus: Über Gesteine, Minerale und Kristalle, Fossilien, Schichten und Berge by Nicolaus Steno; Karl Mieleitner; Eginhard Fabian. Isis, 81(4):775–776, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233859.

Calinger:1990:BRBd

[359] Ronald Calinger. Book review: Leibniz: A Biography by E. J. Aiton. Isis, 81(4):776–777, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233860.

Meyer:1990:BRB

[360] Eric R. Meyer. Book review: The History of the King’s School, Grantham: 660 Years of a Grammar School by S. J. Branson; Über die Gravitation…: Texte zu den philosophischen Grundlagen der klassischen Mechanik by Isaac Newton; Gernot Böhme. Isis, 81(4):777–778, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233861.

Jacob:1990:BRB

[361] Margaret C. Jacob. Book review: Echo’s van een wetenschappelijke revolutie: De mechanistische natuurwetenschap aan de Leuvense Artesfaculteit (1650–1797) by G. Vanpaemel. Isis, 81(4):778, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233862.

Buchdahl:1990:BRB

[362] Gerd Buchdahl. Book review: Space, Time, and Thought in Kant by Arthur Melnick. Isis, 81(4):779, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233863.

Korshin:1990:BRB

[363] Paul J. Korshin. Book review: Under the Banner of Science: Erasmus Darwin and His Age by Maureen McNeil. Isis, 81(4):780–781, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233864.

Jackson:1990:BRB

[364] Myles W. Jackson. Book review: Goethe und Soemmerring: Briefwechsel, 1784–1828 by Manfred Wenzel. Isis, 81(4):781, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233865.

Richmond:1990:BRB

[365] Marsha L. Richmond. Book review: Aus meinem Leben: Autobiographische Bekanntnisse by Alexander von Humboldt; Kurt-R. Biermann; Alexander von Humboldt by Kurt Schleucher. Isis, 81(4):781–783, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233866.

Herbert:1990:BRB

[366] Sandra Herbert. Book review: Charles Darwin’s Beagle Diary by Richard Darwin Keynes. Isis, 81(4):783–784, December 1990. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233867.

Burlingame:1990:BRBb

[367] Leslie J. Burlingame. Book review: The Non-Darwinian Revolution: Reinterpreting a Historical Myth by Peter J. Bowler. Isis, 81(4):784–785, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233868.

Stocking:1990:BRBb

[368] George W. Stocking, Jr. Book review: Eugène Dubois and the Ape-Man from Java: The History of the First “Missing Link” and Its Discoverer by Bert Theunissen. Isis, 81(4):785–786, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233869.

Shteir:1990:BRB

[369] Ann B. Shteir. Book review: The Romance of Victorian Natural History by Lynn L. Merrill. Isis, 81(4):786–787, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233870.

Fee:1990:BRB

[370] Elizabeth Fee. Book review: Sexual Science: The Victorian Construction of Womanhood by Cynthia Eagle Russett. Isis, 81(4):787–788, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233871.

McKay:1990:BRB

[371] John P. McKay. Book review: Disenchanted Night: The Industrialization of Light in the Nineteenth Century by Wolfgang Schivelbusch; Angela Davies. Isis, 81(4):788–790, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233872.

Hughes:1990:BRB

[372] Thomas P. Hughes. Book review: The Papers of Thomas A. Edison. Volume I: The Making of an Inventor, February 1847–June 1873 by Reese V. Jenkins. Isis, 81(4):790–791, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233873.

VanRiper:1990:BRB

[373] A. Bowdoin Van Riper. Book review: An Ancient Air: A Biography of John Stringfellow of Chard, the Victorian Aeronautical Pioneer by Harald Penrose. Isis, 81(4):792, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233874.

Weindling:1990:BRB

[374] Paul Weindling. Book review: Die “Volksheilstätten–Bewegung” in Deutschland um 1900: Zur Ideengeschichte der Sanatoriumstherapie für Tuberkulöse by Wolfgang Seeliger. Isis, 81(4):792–793, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233875.

Stowe:1990:BRB

[375] Steven Stowe. Book review: Disease and Distinctiveness in the American South by Todd L. Savitt; James Harvey Young; Science and Medicine in the Old South by Ronald L. Numbers; Todd L. Savitt. Isis, 81(4):793– 794, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233876.

Lederer:1990:BRB

[376] Susan E. Lederer. Book review: Disease in the Popular American Press: The Case of Diphtheria, Typhoid Fever, and Syphilis, 1870– 1920 by Terra Ziporyn. Isis, 81(4):794–795, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233877.

Yoxen:1990:BRB

[377] Edward Yoxen. Book review: The Politics of British Science by Martin Ince. Isis, 81(4):795–796, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233878.

Dick:1990:BRBb

[378] Steven J. Dick. Book review: History of the Royal Astronomical Society. Volume II: 1920–1980 by R. J. Tayler. Isis, 81(4):796–797, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233879.

Stranges:1990:BRB

[379] Anthony N. Stranges. Book review: Life of a Scientist: An Autobiographical Account of the Development of Molecular Orbital Theory with an Introductory Memoir by Friedrich Hund by Robert S. Mulliken; Bernard J. Ransil. Isis, 81(4):797, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233880.

Henriksen:1990:BRB

[380] Paul W. Henriksen. Book review: The Collected Papers of Peter J. W. Debye by Peter J. W. Debye. Isis, 81(4):798, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233881.

Worster:1990:BRB

[381] Donald Worster. Book review: Ecology in the Twentieth Century: A History by Anna Bramwell; The Rights of Nature: A History of Environmental Ethics by Roderick Frazier Nash. Isis, 81(4):798–800, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233882.

Harris:1990:BRBb

[382] Benjamin Harris. Book review: James J. Gibson and the Psychology of Perception by Edward S. Reed. Isis, 81(4):800–801, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233883.

Mark:1990:BRBb

[383] Joan Mark. Book review: Daughters of the Desert: Women Anthropologists and the Native American Southwest, 1880–1980: An Illustrated Catalogue by Barbara Babcock; Nancy J. Parezo. Isis, 81(4):801–802, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233884.

Webster:1990:BRB

[384] Charles Webster. Book review: Money, Power, and Health Care by Evan M. Melhado; Walter Feinberg; Harold M. Swartz. Isis, 81(4):802–803, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233885.

Whatley:1990:BRB

[385] Mariamne H. Whatley. Book review: Conception in the Test Tube: The IVF Story: How Australia Leads the World by Harry Kannegiesser. Isis, 81(4):803–804, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233886.

Israel:1990:BRB

[386] Paul B. Israel. Book review: The History of N. V. Philips’ Gloeilampenfabrieken. Volume I: A Company of Many Parts by A. Heerding; Derek S. Jordan. Isis, 81(4):804–805, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233887.

Mack:1990:BRBb

[387] Pamela E. Mack. Book review: RAND’s Role in the Evolution of Balloon and Satellite Observation Systems and Related U.S. Space Technology by Merton E. Davies; William R. Harris. Isis, 81(4):805, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233888.

Camerini:1990:BRB

[388] Jane R. Camerini. Book review: Des mots et des lieux: La dynamique du discours géographique by Vincent Berdoulay. Isis, 81(4):805–806, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233889.

Nickles:1990:BRBb

[389] Thomas Nickles. Book review: Social Epistemology by Steve Fuller. Isis, 81(4):806–808, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233890.

Hull:1990:BRB

[390] David L. Hull. Book review: Reconstructing the Past: Parsimony, Evolution, and Inference by Elliott Sober. Isis, 81(4):808–809, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233891.

DiSalle:1990:BRB

[391] Robert DiSalle. Book review: Einstein’s Revolution: A Study in Heuristic by Elie Zahar. Isis, 81(4):809–810, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233892.

Aarsleff:1990:BRB

[392] Hans Aarsleff. Book review: Schools of Thought: The Development of Linguistics from Bopp to Saussure by Olga Amsterdamska. Isis, 81(4): 810–811, December 1990. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233893.

Miller:1990:BRBb

[393] David Philip Miller. Book review: Sir Joseph Banks (1743–1820): A Guide to Biographical and Bibliographical Sources by Harold B. Carter. Isis, 81(4):811–812, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233894.

Principe:1990:BRB

[394] Lawrence M. Principe. Book review: Metaphysische, experimentelle und utilitaristische Traditionen in der Antimonliteratur zur Zeit der “wissenschaftlichen Revolution” (1520–1820): Eine kommentierte Auswahl-Bibliographie by Hermann Fischer. Isis, 81(4):812–813, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233895.

Roe:1990:BRBb

[395] Shirley A. Roe. Book review: The Encyclopedists as Individuals: A Biographical Dictionary of the Authors of the Encyclopédie by Frank A. Kafker; Serena L. Kafker. Isis, 81(4):813–814, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233896.

Brown:1990:BRB

[396] Jack Brown. Book review: The Franklin Institute and the Making of Industrial America by Stephanie A. Morris. Isis, 81(4):814–815, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233897.

Pernick:1990:BRB

[397] Martin S. Pernick. Book review: Medicine and Film: A Checklist, Survey, and Research Resource by Michael Shortland. Isis, 81(4):815–816, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233898.

Anonymous:1990:Cd

[398] Anonymous. Collections. Isis, 81(4):816–818, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233899.

Anonymous:1990:BMe

[399] Anonymous. Back matter. Isis, 81(4):819–854, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233900.

Anonymous:1990:VI

[400] Anonymous. Volume information. Isis, 81(4):829–853, December 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233807.

Neu:1991:CBH

[401] John Neu. Current bibliography of the history of science and its cultural influences, 1991. Isis, 82(??):1–271, ???? 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234170.

Anonymous:1991:BMa

[402] Anonymous. Back matter. Isis, 82(??):??, ???? 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234171.

Anonymous:1991:FMa

[403] Anonymous. Front matter. Isis, 82(??):i–xvi, ???? 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234169.

Anonymous:1991:FMb

[404] Anonymous. Front matter. Isis, 82(1):1–7, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233511.

Lindee:1991:ACJ

[405] M. Susan Lindee. The American career of Jane Marcet’s conversations on chemistry, 1806–1853. Isis, 82(1):8–23, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233512.

Yeo:1991:RES

[406] Richard Yeo. Reading encyclopedias: Science and the organization of knowledge in British dictionaries of arts and sciences, 1730–1850. Isis, 82(1):24–49, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233513.

Mancosu:1991:TIL

[407] Paolo Mancosu and Ezio Vailati. Torricelli’s infinitely long solid and its philosophical reception in the Seventeenth Century. Isis, 82(1):50– 70, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233514.

Forman:1991:ITH

[408] Paul Forman. Independence, not transcendence, for the historian of science. Isis, 82(1):71–86, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233515.

Pingree:1991:EON

[409] David Pingree. Eloge: Otto Neugebauer, 26 May 1899–19 February 1990. Isis, 82(1):87–88, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233516.

McEvoy:1991:LE

[410] John G. McEvoy, Maurice Crosland, C. Truesdell, Craig Fraser, Gideon Freudenthal, and Gad Freudenthal. Letters to the Editor. Isis, 82(1): 89–90, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233517.

Anonymous:1991:CJH

[411] Anonymous. Correction: Journal of the history of biology. Isis, 82(1): 90, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233518.

Pickstone:1991:BRB

[412] John Pickstone. Book review: The Care of Strangers: The Rise of America’s Hospital System. Isis, 82(1):91–94, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233519.

Gillispie:1991:BRB

[413] Charles C. Gillispie. Book review: Companion to the History of Modern Science. Isis, 82(1):94–98, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233520.

McNeill:1991:BRB

[414] William H. McNeill. Book review: Health and the Rise of Civilization by Mark Nathan Cohen. Isis, 82(1):99–100, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233521.

Schofield:1991:BRB

[415] Robert E. Schofield. Book review: Children of Prometheus: A History of Science and Technology by James MacLachlan. Isis, 82(1):100–101, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233522.

Zuckerman:1991:BRB

[416] Harriet Zuckerman. Book review: Scientific Genius: A Psychology of Science by Dean Keith Simonton. Isis, 82(1):101–102, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233523.

Dear:1991:BRB

[417] Peter Dear. Book review: The Uses of Experiment: Studies in the Natural Sciences by David Gooding; Trevor Pinch; Simon Schaffer. Isis, 82(1): 102–103, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233524.

Baird:1991:BRB

[418] Davis Baird. Book review: The Empire of Chance: How Probability Changed Science and Everyday Life by Gerd Gigerenzer; Zeno Swijtink; Theodore Porter; Lorraine Daston; John Beatty; Lorenz Krüger. Isis, 82(1):103–105, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233525.

Paul:1991:BRBa

[419] John J. Paul. Book review: Machines as the Measure of Men: Science, Technology, and Ideologies of Western Dominance by Michael Adas. Isis, 82(1):105–106, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233526.

Cohen:1991:BRB

[420] I. Bernard Cohen. Book review: Revolution in History by Roy Porter; Mikulás Teich; Revolutions in Science: Their Meaning and Relevance by William R. Shea. Isis, 82(1):106–107, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233527.

Samuels:1991:BRB

[421] Richard J. Samuels. Book review: The Formation of Science in Japan: Building a Research Tradition by James R. Bartholomew. Isis, 82(1): 107–108, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233528.

Rudwick:1991:BRB

[422] Martin Rudwick. Book review: Catastrophic Episodes in Earth History by Claude C. Albritton. Isis, 82(1):108–109, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233529.

Osborne:1991:BRBa

[423] Michael A. Osborne. Book review: Aux sources de la biologie: Volume III: L’anatomie by Réjane Bernier. Isis, 82(1):109, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233530.

Rauch:1991:BRB

[424] Alan Rauch. Book review: A Concise History of Veterinary Medicine by D. Karasszon; E. Farkas. Isis, 82(1):109–110, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233531.

Tango:1991:BRB

[425] Gerard J. Tango. Book review: Technology and War: From 2000 B.C. to the Present by Martin van Creveld. Isis, 82(1):110–111, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233532.

Hankinson:1991:BRBa

[426] R. J. Hankinson. Book review: Cosmic Problems: Essays on Greek and Roman Philosophy of Nature by David Furley. Isis, 82(1):111–112, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233533.

Lipton:1991:BRB

[427] Joshua Lipton. Book review: Le ciel des Romains by A. Le Boeuffle. Isis, 82(1):112–113, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233534.

Taub:1991:BRBa

[428] Liba C. Taub. Book review: The Criterion of Truth: Essays Written in Honor of George Kerferd, together with a Text and Translation (With Annotations) of Ptolemy’s “On the Kriterion and Hegemonikon” by Pamela Huby; Gordon Neal. Isis, 82(1):113–114, March 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233535.

Emilsson:1991:BRB

[429] Eyjolfur Kjalar Emilsson. Book review: Plotinus: The Experience of Unity by Gary M. Gurtler. Isis, 82(1):114–115, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233536.

Bowen:1991:BRBa

[430] Alan C. Bowen. Book review: Book 7 of the Collection by Pappus of Alexandria; Alexander Jones. Isis, 82(1):115–116, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233537.

King:1991:BRBa

[431] David A. King. Book review: The Astronomical Works of Gregory Chioniades. Volume I: Zij al-⊂Ala⊃ i by Gregory Chioniades; David Pingree; An Eleventh-Century Manual of Arabo-Byzantine Astronomy by Alexander Jones. Isis, 82(1):116–118, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233538.

Kennedy:1991:BRB

[432] E. S. Kennedy. Book review: Los tratados de construcción uso de la azafea de Azarquiel by Roser Puig Aguilar. Isis, 82(1):118, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233539.

Freudenthal:1991:BRBa

[433] Gad Freudenthal. Book review: Torah and Sophia: The Life and Thought of Shem Tov ibn Falaquera by Raphael Jospe. Isis, 82(1):118–119, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233540.

Ovitt:1991:BRB

[434] George Ovitt, Jr. Book review: Hildegard of Bingen, 1098–1179: A Visionary Life by Sabina Flanagan; St. Francis of Assisi and Nature: Tradition and Innovation in Western Christian Attitudes toward the Environment by Roger D. Sorrell. Isis, 82(1):119–121, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233541.

Dilke:1991:BRB

[435] O. A. W. Dilke. Book review: La “Descriptio mappe mundi” de Hugues de Saint-Victor by Hugo of Saint-Victor; Patrick Gautier Dalché; John de Foxton’s Liber Cosmographiae (1408): An Edition and Codicological Study by John de Foxton; John B. Friedman. Isis, 82(1):121–123, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233542.

Livesey:1991:BRB

[436] Steven J. Livesey. Book review: Compendium of the Study of Theology by Roger Bacon; Thomas S. Maloney. Isis, 82(1):123–124, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233543.

Busard:1991:BRB

[437] H. L. L. Busard. Book review: Geometria speculativa by Thomas Bradwardine; George Molland. Isis, 82(1):124–125, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233544.

Jones:1991:BRBa

[438] Peter Murray Jones. Book review: Arnaldi de Villanova opera medica omnia. Volume IV: Tractatus de consideracionibus operis medicine sive de flebotomia by Arnald of Villanova; Luke Demaitre; Pedro Gil-Sotres. Isis, 82(1):125, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233545.

Goddu:1991:BRB

[439] André L. Goddu. Book review: Die aristotelisch-scholastische Theorie der Bewegung: Studien zum Kommentar Alberts von Sachsen zur Physik des Aristoteles by Jürgen Sarnowsky. Isis, 82(1):125–127, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233546.

Kraye:1991:BRB

[440] Jill Kraye. Book review: Schooling in Renaissance Italy: Literacy and Learning, 1300–1600 by Paul F. Grendler. Isis, 82(1):127–128, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233547.

Knoll:1991:BRB

[441] Paul W. Knoll. Book review: Renaissance Culture in Poland: The Rise of Humanism, 1470–1543 by Harold B. Segel. Isis, 82(1):128–129, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233548.

Kremer:1991:BRB

[442] Richard L. Kremer. Book review: Entstehung und Ausbreitung der copernicanischen Lehre by Ernst Zinner; Heribert M. Nobis; Felix Schmeidler. Isis, 82(1):129–130, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233549.

Demaitre:1991:BRB

[443] Luke E. Demaitre. Book review: L’espressione e l’immagine: Introduzione a Paracelso by Giancarlo Zanier. Isis, 82(1):130–131, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233550.

Shackelford:1991:BRB

[444] Jole Shackelford. Book review: The Isenheim Altarpiece: God’s Medicine and the Painter’s Vision by Andree Hayum. Isis, 82(1):131, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233551.

Moran:1991:BRB

[445] Bruce T. Moran. Book review: Authority, Liberty and Automatic Machinery in Early Modern Europe by Otto Mayr. Isis, 82(1):132–133, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233552.

Guerrini:1991:BRB

[446] Anita Guerrini. Book review: The Mind Has No Sex? Women in the Origins of Modern Science by Londa L. Schiebinger. Isis, 82(1):133– 134, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233553.

Cormack:1991:BRB

[447] Lesley B. Cormack. Book review: John Dee’s Natural Philosophy: Between Science and Religion by Nicholas H. Clulee. Isis, 82(1):134–135, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233554.

Moesgaard:1991:BRB

[448] Kr. Peder Moesgaard. Book review: The Wittich Connection: Conflict and Priority in Late Sixteenth-Century Cosmology by Owen Gingerich; Robert S. Westman. Isis, 82(1):135–136, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233555.

VanEgmond:1991:BRB

[449] Warren Van Egmond. Book review: Mathematica by Tommaso Campanella; Armando Brissoni. Isis, 82(1):136, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233556.

Hatfield:1991:BRB

[450] Gary Hatfield. Book review: Il dualismo da Cartesio a Leibniz: Cartesio, Cordemoy, La Forge, Malebranche, Leibniz by Salvatore Nicolosi. Isis, 82(1):136–137, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233557.

Gabbey:1991:BRB

[451] Alan Gabbey. Book review: Court traité des premiers principes: Le “Short Tract on First Principles” de 1630–1631: La naissance de Thomas Hobbes à la pensée moderne by Thomas Hobbes; Jean Bernhardt. Isis, 82(1):137–138, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233558.

Wood:1991:BRB

[452] P. B. Wood. Book review: Essays on the Law of Nature: The Latin Text with a Translation, Introduction and Notes, together with Transcripts of Locke’s Shorthand in His Journal for 1676 by John Locke; W. von Leyden. Isis, 82(1):138–139, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233559.

Meinel:1991:BRB

[453] Christoph Meinel. Book review: Neu-eröffnete chymische Artzney- und Werck-Schul by Christoph Glaser; Hans-Joachim Poeckern. Isis, 82(1): 139–140, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233560.

Kelley:1991:BRB

[454] Donald R. Kelley. Book review: Vico Revisited: Orthodoxy, Naturalism, and Science in the Scienza nuova by Gino Bedani. Isis, 82(1):140–141, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233561.

Brockliss:1991:BRB

[455] L. W. B. Brockliss. Book review: Cambridge in the Age of the Enlightenment: Science, Religion, and Politics from the Restoration to the French Revolution by John Gascoigne. Isis, 82(1):141–142, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233562.

dePater:1991:BRB

[456] C. de Pater. Book review: George Berkeley, 1685–1753 by Wolfgang Breidert. Isis, 82(1):142–143, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233563.

Osler:1991:BRBa

[457] Margaret J. Osler. Book review: Particles and Ideas: Bishop Berkeley’s Corpuscularian Philosophy by Gabriel Moked. Isis, 82(1):143–144, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233564.

VanHelden:1991:BRB

[458] Albert Van Helden. Book review: Welzijn, wijsbegeerte, en wetenschap by Willem Jacob’s Gravesande; C. de Pater. Isis, 82(1):144–145, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233565.

Kuriyama:1991:BRB

[459] Shigehisa Kuriyama. Book review: Albinus on Anatomy by Bernard Siegfried Albinus; Robert Beverly Hale; Terence Coyle. Isis, 82(1):145– 146, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233566.

Home:1991:BRB

[460] R. W. Home. Book review: Un physicien au siècle des lumières: L’Abbé Nollet, 1700–1770 by Jean Torlais. Isis, 82(1):146–147, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233567.

Risse:1991:BRBa

[461] Guenter B. Risse. Book review: Georg Christian Gottlieb Wedekind, 1761–1831: Werdegang und Schicksal eines Arztes im Zeitalter der Aufklärung und der Französischen Revolution. Mit einem Anhang: Wedekinds Diätetikvorlesung von 1789/90 by Martin Weber. Isis, 82 (1):147–148, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233568.

Woodward:1991:BRBa

[462] William R. Woodward. Book review: Structures of Knowing: Psychologies of the Nineteenth Century by Katherine Arens. Isis, 82(1):148–149, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233569.

Grattan-Guinness:1991:BRBa

[463] I. Grattan-Guinness. Book review: Mathematisch-physikalische und philosophische Schriften, 1842–1843 by Bernard Bolzano; Gotffried Gabriel; Matthias Gatzemeier; Friedrich Kambartel. Isis, 82(1):149, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233570.

Hofmann:1991:BRBa

[464] James R. Hofmann. Book review: James Joule: A Biography by Donald S. L. Cardwell. Isis, 82(1):149–150, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233571.

Dugan:1991:BRB

[465] Kathleen G. Dugan. Book review: An American Scientist in Early Meiji Japan: The Autobiographical Notes of Thomas C. Mendenhall by Thomas C. Mendenhall; Richard Rubinger. Isis, 82(1):151, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233572.

Turner:1991:BRBa

[466] R. Steven Turner. Book review: An Institute for an Empire: The Physikalisch-Technische Reichsanstalt, 1871–1918 by David Cahan. Isis, 82(1):151–152, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233573.

Richards:1991:BRBa

[467] Evelleen Richards. Book review: The Politics of Evolution: Morphology, Medicine, and Reform in Radical London by Adrian Desmond. Isis, 82 (1):152–153, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233574.

Lightman:1991:BRB

[468] Bernard Lightman. Book review: Evolution and Ethics: T. H. Huxley’s Evolution and Ethics with New Essays on Its Victorian and Sociobiological Context by Thomas Henry Huxley; James Paradis; George C. Williams. Isis, 82(1):154–155, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233575.

Mazumdar:1991:BRB

[469] Pauline M. H. Mazumdar. Book review: Robert Koch: A Life in Medicine and Bacteriology by Thomas D. Brock. Isis, 82(1):155–156, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233576.

LaBerge:1991:BRB

[470] Ann La Berge. Book review: Doctors, Bureaucrats, and Public Health in France, 1888–1902 by Martha L. Hildreth. Isis, 82(1):156–157, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233577.

Harwood:1991:BRB

[471] Jonathan Harwood. Book review: Die Universität Göttingen unter dem Nationalsozialismus: Das Verdrängte Kapitel ihrer 250-jährigen Geschichte by Heinrich Becker; Hans-Joachim Dahms; Cornelia Wegeler. Isis, 82(1):157–158, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233578.

Walker:1991:BRBa

[472] Mark Walker. Book review: Wissenschaft für Macht und Markt: Kernforschung und Mikroelektronik in der Bundesrepublik Deutschland by Michael Eckert; Maria Osietzki. Isis, 82(1):158–159, March 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233579.

Schweber:1991:BRB

[473] S. S. Schweber. Book review: Gesammelte Werke/Collected Works by Werner Heisenberg; W. Blum; H.-P. Dürr; H. Rechenberg. Isis, 82(1): 159–160, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233580.

Gillespie:1991:BRB

[474] Richard Gillespie. Book review: Race to the Stratosphere: Manned Scientific Ballooning in America by David H. DeVorkin. Isis, 82(1):160–161, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233581.

Elliott:1991:BRBa

[475] Eric Elliott. Book review: Svante Arrhenius’ Beziehungen zu österreichischen Gelehrten: Briefe aus Österreich an Svante Arrhenius (1891–1926) by Alois Kernbauer. Isis, 82(1):161–162, March 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233582.

Montgomery:1991:BRBa

[476] William Montgomery. Book review: Julius Schaxel an Ernst Haeckel, 1906–1917 by Julius Schaxel; Erika Krausse; Doris Posselt. Isis, 82(1): 162, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233583.

Tobey:1991:BRBa

[477] Ronald C. Tobey. Book review: Saving America’s Wildlife by Thomas R. Dunlap. Isis, 82(1):162–163, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233584.

Mitman:1991:BRBa

[478] Gregg Mitman. Book review: Primate Visions: Gender, Race, and Nature in the World of Modern Science by Donna Haraway. Isis, 82(1): 163–165, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233585.

Gamble:1991:BRB

[479] Vanessa Northington Gamble. Book review: Charles Richard Drew: The Man and the Myth by Charles E. Wynes. Isis, 82(1):165–166, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233586.

Fan:1991:BRB

[480] Dainian Fan. Book review: Medicine in Rural China: A Personal Account by C. C. Chen; Frederica M. Bunge. Isis, 82(1):166–167, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233587.

Reagan:1991:BRB

[481] Leslie J. Reagan. Book review: Abortion in England, 1900–1967 by Barbara Brookes; Abortion, Doctors, and the Law: Some Aspects of the Legal Regulation of Abortion in England from 1803 to 1982 by John Keown. Isis, 82(1):167–168, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233588.

Ceruzzi:1991:BRB

[482] Paul Ceruzzi. Book review: The Early British Conferences by M. R. Williams; Martin Campbell-Kelly. Isis, 82(1):168–169, March 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233589.

Staudenmaier:1991:BRB

[483] John Staudenmaier. Book review: The Tragedy of Technology: Human Liberation versus Domination in the Late Twentieth Century by Stephen Hill. Isis, 82(1):170–171, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233590.

Hilden:1991:BRB

[484] Patricia J. Hilden. Book review: Made in America: Science, Technology, and American Modernist Poets by Lisa M. Steinman. Isis, 82(1):171– 172, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233591.

Sloan:1991:BRB

[485] Phillip R. Sloan. Book review: Michel Foucault by David R. Shumway. Isis, 82(1):172–173, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233592.

Anellis:1991:BRB

[486] Irving H. Anellis. Book review: The Philosophy of Set Theory: A Historical Introduction to Cantor’s Paradise by Mary Tiles. Isis, 82(1):173–174, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233593.

Bub:1991:BRBa

[487] Jeffrey Bub. Book review: The Structure and Interpretation of Quantum Mechanics by R. I. G. Hughes. Isis, 82(1):174–175, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233594.

Barkan:1991:BRB

[488] Diana Barkan. Book review: Methodological Aspects of the Development of Low Temperature Physics, 1881–1956: Concepts Out of Context(s) by Kostas Gavroglu; Yorgos Goudaroulis. Isis, 82(1):175–177, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233595.

Heelan:1991:BRB

[489] Patrick A. Heelan. Book review: On the Logic of the Social Sciences by Jürgen Habermas; Shierry Weber Nicholsen; Jerry A. Stark; The Material Realization of Science: A Philosophical View on the Experimental Natural Sciences, Developed in Discussion with Habermas by Hans Radder. Isis, 82(1):177–178, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233596.

Sobol:1991:BRBa

[490] Peter Sobol. Book review: Animal Consciousness by Daisie Radner; Michael Radner. Isis, 82(1):178–179, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233597.

Sokal:1991:BRB

[491] Michael M. Sokal. Book review: Phrenology in the British Isles: An Annotated Historical Biobibliography and Index by Roger Cooter. Isis, 82(1):179–180, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233598.

Feingold:1991:BRBa

[492] Mordechai Feingold. Book review: A Catalogue of Seventeenth-Century Printed Books in the National Library of Medicine by Peter Krivatsy. Isis, 82(1):180, March 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233599.

Porter:1991:BRBa

[493] Dorothy E. Porter. Book review: Eighteenth-Century Medics: Subscriptions, Licences, Apprenticeships by P. J. Wallis; R. V. Wallis; J. L. L. Burnby; T. D. Whittet. Isis, 82(1):180–181, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233600.

Anonymous:1991:Ca

[494] Anonymous. Collections. Isis, 82(1):181–184, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233601.

Anonymous:1991:BMb

[495] Anonymous. Back matter. Isis, 82(1):185–190, March 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233602.

Anonymous:1991:FMc

[496] Anonymous. Front matter. Isis, 82(2):191–197, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234818.

Roberts:1991:WWS

[497] Lissa Roberts. A word and the world: The significance of naming the calorimeter. Isis, 82(2):198–222, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234819.

Baatz:1991:SSS

[498] Simon Baatz. “Squinting at Silliman”: Scientific periodicals in the Early American Republic, 1810–1833. Isis, 82(2):223–244, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234820.

Sulloway:1991:RFC

[499] Frank J. Sulloway. Reassessing Freud’s case histories: The social construction of psychoanalysis. Isis, 82(2):245–275, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234821.

Sokal:1991:AMH

[500] Michael M. Sokal and Albert Moyer. Annual meeting of the History of Science Society, Seattle, Washington, 25–28 October 1990. Isis, 82(2): 276–280, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234822.

Forman:1991:SMC

[501] Paul Forman. Sarton Medal citation. Isis, 82(2):281–283, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234823.

Anonymous:1991:CHS

[502] Anonymous. Contributors to the History of Science Society. Isis, 82(2): 283–284, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234824.

McClellan:1991:BRB

[503] James E. McClellan, III. Book review: Archives Internationales d’Histoire des Sciences by R. Halleux. Isis, 82(2):285–286, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234825.

Hannaway:1991:BRBa

[504] Caroline Hannaway. Book review: Clio Medica: Acta Academiae Internationalis Historiae Medicinae by A. M. Luyendijk-Elshout. Isis, 82(2): 286–287, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234826.

Mitman:1991:BRBb

[505] Gregg Mitman. Book review: History and Philosophy of the Life Sciences by Mirko D. Grmek; Bernandino Fantini. Isis, 82(2):288–289, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234827.

Stenhouse:1991:BRB

[506] John Stenhouse. Book review: Historical Records of Australian Science by R. W. Home; Metascience: Biannual Review of the Australasian Association for the History, Philosophy and Social Studies of Science by W. R. Albury. Isis, 82(2):289–290, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234828.

Kim:1991:BRB

[507] Yung Sik Kim. Book review: Han’guk Kwahaksa Hakhoe-ji: Journal of the Korean History of Science Society by Song Sang-yong; Historia Scientiarum: The International Journal of the History of Science Society of Japan by Jun Fujimura; Kagakusi Kenkyu: Journal of History of Science, Japan by Ichiro Yabe; Ziran Kexueshi Yanjiu (Studies in the History of Natural Sciences) by Lin Wenjao. Isis, 82(2):291–293, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234829.

Paul:1991:BRBb

[508] John J. Paul. Book review: Indian Journal of History of Science by S. K. Mukherjee; Indian Journal of History of Medicine; Bulletin of the Institute of History of Medicine. Isis, 82(2):293–294, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234830.

Freudenthal:1991:BRBb

[509] Gad Freudenthal. Book review: Koroth: A Bulletin Devoted to the History of Medicine and Science by Joshua O. Leibowitz. Isis, 82(2):295– 296, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234831.

Savage-Smith:1991:BRB

[510] Emilie Savage-Smith. Book review: Arabic Sciences and Philosophy: A Historical Journal by Basim Musallam; Journal for the History of Arabic Science by Ahmad Y. al-Hassan; Khaled Maghout; Roshdi Rashed; Zeitschrift fur Geschichte der Arabisch-Islamischen Wissenschaften by Fuat Sezgin. Isis, 82(2):296–298, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234832.

Josephson:1991:BRB

[511] Paul R. Josephson. Book review: Voprosy Istorii Estestvoznaniia i Tekhniki by B. I. Kozlov. Isis, 82(2):298–300, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234833.

Hilfstein:1991:BRB

[512] Erna Hilfstein. Book review: Kwartalnik Historii Nauki i Techniki by Tadeusz Bieńkowski; Organon by Bogdan Suchodolski. Isis, 82(2):300– 301, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234834.

Ahonen:1991:BRB

[513] Kathleen Ahonen. Book review: Hippokrates: Annales Societatis Historiae Medicinae Fennicae by Kalle Achté; Science Studies: A Scandinavian Journal Published by the Finnish Society for Science Studies by Marja Alestalo. Isis, 82(2):301–302, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234835.

Friedman:1991:BRB

[514] Robert Marc Friedman. Book review: Lychnos: Lärdomshistoriska Samfundets Årsbok by Gunnar Eriksson. Isis, 82(2):302–304, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234836.

Thoren:1991:BRB

[515] Victor Thoren. Book review: Centaurus: International Magazine of the History of Mathematics, Science, and Technology by Kirsti Andersen; Ole Knudsen; Kurt Møller Pedersen; Olaf Pedersen. Isis, 82(2):304, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234837.

Cook:1991:BRBa

[516] Harold J. Cook. Book review: Janus: Revue Internationale de l’Histoire des Sciences, de la Médecine, de la Pharmacie et de la Technique by H. A. M. Snelders; M. J. van Lieburg; E. M. Bruins; Tijdschrift voor de Geschiedenis der Geneeskunde, Natuurwetenschappen, Wiskunde, en Techniek by A. M. Luyendijk-Elshout; Tractrix: Yearbook for the History of Science, Medicine, Technology, and Mathematics by H. Floris Cohen; Bert Theunissen. Isis, 82(2):304–306, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234838.

Cahan:1991:BRB

[517] David Cahan. Book review: Berichte zur Wissenschaftsgeschichte by Fritz Krafft; NTM: Schriftenreihe für Geschichte der Naturwissenschaften, Technik und Medizin by Rolf Sonnemann; Dietrich Tutzke; Hans Wussing; Renate Tobies. Isis, 82(2):306–309, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234839.

Weindling:1991:BRB

[518] Paul Weindling. Book review: Medizinhistorisches Journal: Internationale Vierteljahresschrift für Wissenschaftsgeschichte by Gunter Mann; Werner F. Kümmel; Ulrich Tröhler; Ursula Weisser; Sudhoffs Archiv: Zeitschrift für Wissenschaftsgeschichte by Peter Dilg; Menso Folkerts; Gundolf Keil; Fritz Krafft; Rolf Winau; Paul Unschuld. Isis, 82(2):309– 311, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234840.

Mayr:1991:BRB

[519] Otto Mayr. Book review: Technikgeschichte by K. Mauel; R. Reith. Isis, 82(2):311–313, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234841.

Emch-Deriaz:1991:BRB

[520] Antoinette Emch-Dériaz. Book review: Gesnerus by Carl Haffter. Isis, 82(2):313–314, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234842.

Finocchiaro:1991:BRBa

[521] Maurice A. Finocchiaro. Book review: Bolletino di Storia delle Scienze Matematiche by Enrico Giusti; Nuncius: Annali di Storia della Scienza by Paolo Galluzzi; Physis: Rivista Internazionale di Storia della Scienza by Vincenzo Cappelletti; Guido Cimino; Rivista de Storia della Scienza by Giorgio Tecce. Isis, 82(2):314–317, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234843.

Nye:1991:BRBa

[522] Mary Jo Nye. Book review: Revue de Synthèse by Dominique Bourel; Eric Brian; Roger Chartier; Joël Cornette; Ernest Coumet; Henri-Jean Martin; Jacques Merleau-Ponty; Pierre Monzani; Jean-Claude Perrot; Roshdi Rashed; Daniel Roche; Revue d’Histoire des Sciences by Michel Blay; Sciences et Techniques en Perspective by Jean Dhombres. Isis, 82(2):317–319, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234844.

McVaugh:1991:BRBa

[523] Michael R. McVaugh. Book review: Asclepio: Revista de Historia de la Medicina y de la Ciencia by Agustín Albarracín Teulón; Dynamis: Acta Hispanica ad Medicinae Scientiarumque Historiam Illustrandam by Esteban Rodríguez Ocaña. Isis, 82(2):319–320, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234845.

Glick:1991:BRB

[524] Thomas F. Glick. Book review: Llull: Revista de la Sociedad Española de Historia de las Ciencias y de las Técnicas by Mariano Hormigón; Sylva Clius: Revista de Historia de la Ciencia by Javier Ordóñez. Isis, 82(2): 320–321, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234846.

Alvarado:1991:BRB

[525] Carlos S. Alvarado. Book review: Revista de Historia de la Psicología by Helio Carpintero; Archivo Latinoamericano de Historia de la Psicología y Ciencias Afines by Hannes Stubbe; Ramón León. Isis, 82(2):321–323, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234847.

Chambers:1991:BRB

[526] David Wade Chambers. Book review: Quipu: Revista Latinoamericana de Historia de las Ciencias y la Tecnología by Juan José Saldaña. Isis, 82(2):323–324, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234848.

Sheets-Pyenson:1991:BRB

[527] Susan Sheets-Pyenson. Book review: Scientia Canadensis by James Hull. Isis, 82(2):324–325, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234849.

Abir-Am:1991:BRB

[528] Pnina G. Abir-Am. Book review: What Mad Pursuit: A Personal View of Scientific Discovery. Isis, 82(2):326–343, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234850.

Smeaton:1991:BRB

[529] William A. Smeaton. Book review: Études sur/Studies on Hélène Metzger by Gad Freudenthal. Isis, 82(2):344–345, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234851.

DeYoung:1991:BRB

[530] Gregg De Young. Book review: Wissenschaftsgeschichte en miniature: Neun Kapitel aus der Entwicklung der Mathematik und der Naturwissenschaften by Hans Wussing; Horst Remane. Isis, 82(2):345, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234852.

Schneider:1991:BRB

[531] Ivo H. Schneider. Book review: From Ancient Omens to Statistical Mechanics: Essays on the Exact Sciences Presented to Asger Aaboe by J. L. Berggren; B. R. Goldstein. Isis, 82(2):345–346, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234853.

Pedersen:1991:BRB

[532] Olaf Pedersen. Book review: Empires of Time: Calendars, Clocks, and Cultures by Anthony F. Aveni. Isis, 82(2):346–348, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234854.

King:1991:BRBb

[533] David A. King. Book review: Astronomy of Islamic Times for the Twenty-First Century by Mohammad Ilyas. Isis, 82(2):348–349, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234855.

Hofmann:1991:BRBb

[534] James R. Hofmann. Book review: Reading the Mind of God: In Search of the Principle of Universality by James Trefil. Isis, 82(2):349–350, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234856.

Chapin:1991:BRB

[535] Seymour L. Chapin. Book review: La figure de la Terre du XVIIIe siècle à l’ère spatiale by Henri Lacombe; Pierre Costabel. Isis, 82(2): 350–351, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234857.

Fabian:1991:BRBa

[536] Eginhard Fabian. Book review: Die Symmetrie der Kristalle: Von René-Just Haüy zur kristallographischen Schule in Zürich by J. J. Burckhardt; Erhard Scholz. Isis, 82(2):351–352, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234858.

Proctor:1991:BRB

[537] James D. Proctor. Book review: Americans and Their Forests: A Historical Geography by Michael Williams. Isis, 82(2):352–353, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234859.

Urry:1991:BRB

[538] James Urry. Book review: The Invention of Primitive Society: Transformations of an Illusion by Adam Kuper. Isis, 82(2):353–354, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234860.

Alborn:1991:BRB

[539] Timothy Alborn. Book review: More Heat than Light: Economics as Social Physics, Physics as Nature’s Economics by Philip Mirowski. Isis, 82(2):354–355, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234861.

Findlen:1991:BRBa

[540] Paula Findlen. Book review: Medicine at the Courts of Europe, 1500– 1837 by Vivian Nutton. Isis, 82(2):355, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234862.

Shapiro:1991:BRB

[541] Henry D. Shapiro. Book review: The Pencil: A History of Design and Circumstance by Henry Petroski. Isis, 82(2):355–356, June 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234863.

Gossin:1991:BRB

[542] Pamela Gossin. Book review: Literature and Science: Theory and Practice by Stuart Peterfreund; Science and the Human Spirit: Contexts for Writing and Learning by Fred D. White. Isis, 82(2):356–358, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234864.

Rousseau:1991:BRBa

[543] G. S. Rousseau. Book review: Metaphors of Mind in Fiction and Psychology by Michael S. Kearns. Isis, 82(2):358–359, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234865.

Bowen:1991:BRBb

[544] Alan C. Bowen. Book review: The Origins of the Mithraic Mysteries: Cosmology and Salvation in the Ancient World by David Ulansey. Isis, 82(2):359–360, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234866.

Unguru:1991:BRBa

[545] Sabetai Unguru. Book review: L’héritage épistémologique d’Eudoxe de Cnide: Un essai de reconstitution by Jean-Louis Gardies. Isis, 82(2): 360–362, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234867.

Preus:1991:BRB

[546] Anthony Preus. Book review: Aristotle on Substance: The Paradox of Unity by Mary Louise Gill. Isis, 82(2):362–363, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234868.

Ebbesen:1991:BRB

[547] Sten Ebbesen. Book review: Aristotle’s Categories and Porphyry by Christos Evangeliou. Isis, 82(2):363–364, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234869.

Jones:1991:BRBb

[548] Alexander Jones. Book review: Pythagoras Revived: Mathematics and Philosophy in Late Antiquity by Dominic J. O’Meara. Isis, 82(2):364– 365, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234870.

Hankinson:1991:BRBb

[549] R. J. Hankinson. Book review: Das Prognosticon wurde nicht vor, sondern nach den Epidemienbüchern III und I verfasst: Zweiter Beitrag zur Chronologie der echten Hippokratischen Schriften by Charles Lichtenthaeler; Medecine et philosophie au temps d’Hippocrate by Bernard Vitrac. Isis, 82(2):365–366, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234871.

Ferngren:1991:BRB

[550] Gary B. Ferngren. Book review: Herophilus: The Art of Medicine in Early Alexandria by Heinrich von Staden. Isis, 82(2):366–367, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234872.

Lang:1991:BRB

[551] Helen S. Lang. Book review: Medieval Discussions of the Eternity of the World by Richard C. Dales. Isis, 82(2):367–368, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234873.

Mertens:1991:BRB

[552] Michèle Mertens. Book review: The Alchemical Corpus Attributed to Raymond Lull by Michela Pereira. Isis, 82(2):368–369, June 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234874.

Laird:1991:BRB

[553] W. R. Laird. Book review: Sacralizing the Secular: The Renaissance Origins of Modernity by Stephen A. McKnight. Isis, 82(2):369–371, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234875.

Bono:1991:BRB

[554] James J. Bono. Book review: That Nothing Is Known (Quod nihil scitur) by Francisco Sanches; Elaine Limbrick; Douglas F. S. Thomson. Isis, 82 (2):371–372, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234876.

Clulee:1991:BRB

[555] Nicholas H. Clulee. Book review: The Renaissance Drama of Knowledge: Giordano Bruno in England by Hilary Gatti. Isis, 82(2):372–373, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234877.

North:1991:BRB

[556] J. D. North. Book review: Prophecy and Power: Astrology in Early Modern England by Patrick Curry. Isis, 82(2):373–374, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234878.

Webster:1991:BRB

[557] Charles Webster. Book review: Establishing the New Science: The Experience of the Early Royal Society by Michael Hunter. Isis, 82(2):375–376, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234879.

Meyer:1991:BRB

[558] Eric R. Meyer. Book review: Isaac Newton by Ivo Schneider. Isis, 82(2): 376–377, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234880.

Ariew:1991:BRBa

[559] Roger Ariew. Book review: Leibniz’s Metaphysics: A Historical and Comparative Study by Catherine Wilson. Isis, 82(2):377, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234881.

Porter:1991:BRBb

[560] Roy Porter. Book review: Human Nature, Cultural Diversity, and the French Enlightenment by Henry Vyverberg. Isis, 82(2):378, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234882.

Elena:1991:BRB

[561] Alberto Elena. Book review: Les savants en Révolution, 1789–1799 by Nicole Dhombres. Isis, 82(2):378–379, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234883.

Apt:1991:BRBa

[562] Adam Apt. Book review: Nevil Maskelyne: The Seaman’s Astronomer by Derek Howse. Isis, 82(2):379–380, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234884.

Ten:1991:BRB

[563] Antonio E. Ten. Book review: El observatorio de Cádiz (1753–1831) by Antonio Lafuente; Manuel Sellés. Isis, 82(2):380–381, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234885.

Hahn:1991:BRB

[564] Roger Hahn. Book review: Ruggiero Giuseppe Boscovich nella scienza e nella storia del ’700 by Germano Paoli. Isis, 82(2):381–382, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234886.

McEvoy:1991:BRB

[565] John G. McEvoy. Book review: Eighteenth-Century Chemistry as an Investigative Enterprise by Frederic Lawrence Holmes. Isis, 82(2):382, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234887.

Risse:1991:BRBb

[566] Guenter B. Risse. Book review: Clinical Consultations and Letters by Ippolito Francesco Albertini, Francesco Torti, and Other Physicians: University of Bologna MS 2089-1 by Saul Jarcho. Isis, 82(2):383, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234888.

Brown:1991:BRBa

[567] Chandos Michael Brown. Book review: Thomas Jefferson: Statesman of Science by Silvio A. Bedini. Isis, 82(2):383–384, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234889.

Bruce:1991:BRB

[568] Robert V. Bruce. Book review: Benjamin Silliman: A Life in the Young Republic by Chandos Michael Brown. Isis, 82(2):384–386, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234890.

Hughes:1991:BRB

[569] Miranda Hughes. Book review: Charles Darwin in Australia by F. W. Nicholas; J. M. Nicholas. Isis, 82(2):386, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234891.

Colp:1991:BRB

[570] Ralph Colp, Jr. Book review: Kliment Timiryazev by S. P. Landau-Tylkina; G. G. Egorov. Isis, 82(2):386–387, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234892.

Butcher:1991:BRB

[571] Barry W. Butcher. Book review: The Contented Botanist: Letters of W. H. Harvey about Australia and the Pacific by William H. Harvey; Sophie C. Ducker. Isis, 82(2):387, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234893.

MacDonald:1991:BRB

[572] Michael MacDonald. Book review: Illustrations of Madness by John Haslam; Roy Porter; Observations on Maniacal Disorders by William Pargeter; Stanley W. Jackson; An Essay, Medical, Philosophical, and Chemical on Drunkenness and Its Effects on the Human Body by Thomas Trotter. Isis, 82(2):387–388, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234894.

Hinsley:1991:BRB

[573] Curtis M. Hinsley. Book review: A Stranger in Her Native Land: Alice Fletcher and the American Indians by Joan Mark. Isis, 82(2):389, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234895.

Fletcher:1991:BRB

[574] Ian Christopher Fletcher. Book review: The Artist as Anthropologist: The Representation of Type and Character in Victorian Art by Mary Cowling; Nature into Art: Cultural Transformations in Nineteenth-Century Britain by Carl Woodring. Isis, 82(2):389–391, June 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234896.

Dean:1991:BRB

[575] Dennis R. Dean. Book review: In Pursuit of a Scientific Culture: Science, Art, and Society in the Victorian Age by Peter Allan Dale. Isis, 82(2):391–393, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234897.

Hoch:1991:BRB

[576] Paul K. Hoch. Book review: Science, Politics and the Public Good: Essays in Honour of Margaret Gowing by Nicolaas A. Rupke. Isis, 82 (2):393–394, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234898.

Coopersmith:1991:BRB

[577] Jonathan Coopersmith. Book review: Soviet Science on the Edge of Reform by Harley D. Balzer. Isis, 82(2):394, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234899.

Walker:1991:BRBb

[578] Mark Walker. Book review: Wissenschaft und Nationalsozialismus: Eine Ringvorlesung an der Universität-Gesamthochschule-Siegen by Rainer Geissler; Wolfgang Popp. Isis, 82(2):394–395, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234900.

Cross:1991:BRB

[579] Stephen J. Cross. Book review: The Circuit Riders: Rockefeller Money and the Rise of Modern Science by Gerald Jonas; The Politics of Philanthropy: Abraham Flexner and Medical Education by Steven C. Wheatley. Isis, 82(2):395–396, June 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234901.

Bindon:1991:BRB

[580] George H. Bindon. Book review: Shaping Science and Industry: A History of Australia’s Council for Scientific and Industrial Research, 1926–49 by C. B. Schedvin. Isis, 82(2):396–398, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234902.

Webb:1991:BRB

[581] George E. Webb. Book review: Sandia National Laboratories: The Postwar Decade by Necah Stewart Furman. Isis, 82(2):398–399, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234903.

Dupree:1991:BRB

[582] A. Hunter Dupree. Book review: Atoms for Peace and War, 1953– 1961: Eisenhower and the Atomic Energy Commission by Richard G. Hewlett; Jack M. Holl. Isis, 82(2):399, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234904.

Hetherington:1991:BRBa

[583] Norriss S. Hetherington. Book review: Orders of Magnitude: A History of the NACA and NASA, 1915–1990 by Roger E. Bilstein; Where No Man Has Gone Before: A History of Apollo Lunar Exploration Missions by William David Compton; The First Twenty-Five Years in Space: A Symposium by Allan A. Needell. Isis, 82(2):400–401, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234905.

Grattan-Guinness:1991:BRBb

[584] I. Grattan-Guinness. Book review: Norbert Wiener, 1894–1964 by P. R. Masani. Isis, 82(2):401–402, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234906.

Doel:1991:BRB

[585] Ronald E. Doel. Book review: Zur Geschichte der Festkörperphysik: Farbzentrenforschung bis 1940 by Jürgen Teichmann. Isis, 82(2):402– 403, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234907.

Wessels:1991:BRB

[586] Linda Wessels. Book review: The Historical Development of Quantum Theory. Volume V: Erwin Schrödinger and the Rise of Wave Mechanics by Jagdish Mehra; Helmut Rechenberg. Isis, 82(2):404–405, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234908.

Brock:1991:BRB

[587] W. H. Brock. Book review: Une chimie qui guérit: Histoire de la découverte des sulfamides by Daniel Bovet. Isis, 82(2):405–406, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234909.

Slotten:1991:BRB

[588] Hugh R. Slotten. Book review: Chemical Warfare, Chemical Disarmament by Valerie Adams. Isis, 82(2):407, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234910.

Maienschein:1991:BRB

[589] Jane Maienschein. Book review: Biological Oceanography: An Early History, 1870–1960 by Eric L. Mills. Isis, 82(2):407–408, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234911.

Vicedo:1991:BRB

[590] Marga Vicedo. Book review: Dangerous Diagnostics: The Social Power of Biological Information by Dorothy Nelkin; Laurence Tancredi; Brainstorming: The Science and Politics of Opiate Research by Solomon H. Snyder; Gene Dreams: Wall Street, Academia, and the Rise of Biotechnology by Robert Teitelman. Isis, 82(2):408–409, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234912.

Givens:1991:BRB

[591] Douglas R. Givens. Book review: Portraits in American Archaeology: Remembrances of Some Distinguished Americanists by Gordon Randolph Willey. Isis, 82(2):409–410, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234913.

Vogel:1991:BRB

[592] Morris J. Vogel. Book review: In Sickness and in Wealth: American Hospitals in the Twentieth Century by Rosemary Stevens. Isis, 82(2): 410–411, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234914.

Johnston:1991:BRB

[593] William Johnston. Book review: White Plague, Black Labor: Tuberculosis and the Political Economy of Health and Disease in South Africa by Randall M. Packard. Isis, 82(2):411–412, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234915.

Humphreys:1991:BRBa

[594] Margaret Humphreys. Book review: Plagues and Politics: The Story of the United States Public Health Service by Fitzhugh Mullan. Isis, 82(2): 412–413, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234916.

Jones:1991:BRBc

[595] Kathleen W. Jones. Book review: Before It’s Too Late: The Child Guidance Movement in the United States, 1922–1945 by Margo Horn; The Century of the Child: The Mental Hygiene Movement and Social Policy in the United States and Canada by Theresa R. Richardson. Isis, 82(2): 413–415, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234917.

Zenderland:1991:BRB

[596] Leila Zenderland. Book review: Schools as Sorters: Lewis M. Terman, Applied Psychology, and the Intelligence Testing Movement, 1890–1930 by Paul Davis Chapman; Lewis M. Terman: Pioneer in Psychological Testing by Henry L. Minton. Isis, 82(2):415–416, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234918.

Swann:1991:BRB

[597] John P. Swann. Book review: Eli Lilly: A Life, 1885–1977 by James H. Madison. Isis, 82(2):416–417, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234919.

McMahon:1991:BRB

[598] A. Michal McMahon. Book review: A History of the Institution of Electrical Engineers, 1871–1971 by W. J. Reader; Rachel Lawrence; Sheila Nemet; Geoffrey Tweedale. Isis, 82(2):417–418, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234920.

Sobol:1991:BRBb

[599] Peter Sobol. Book review: A Catalogue of the Pre-1500 Western Manuscript Books at the Newberry Library by Paul Saenger. Isis, 82 (2):418–419, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234921.

Genuth:1991:BRB

[600] Sara Schechner Genuth. Book review: Union Catalogue of Printed Books of the Fifteenth, Sixteenth, and Seventeenth Centuries in European Astronomical Observatories by Giovanna Grassi. Isis, 82(2):419, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234922.

Andrewes:1991:BRB

[601] William Andrewes. Book review: The Great Age of the Microscope: The Collection of the Royal Microscopical Society through 150 Years by Gerard L’E. Turner. Isis, 82(2):419–420, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234923.

Crowe:1991:BRBa

[602] Michael J. Crowe. Book review: Bibliography of Astronomers: Books and Pamphlets in English by and about Astronomers. Volume I: The Spirit of the Nineteenth Century by Paul Luther. Isis, 82(2):420, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234924.

Mauskopf:1991:BRB

[603] Seymour H. Mauskopf. Book review: Animal Magnetism, Early Hypnotism, and Psychical Research, 1766–1925: An Annotated Bibliography by Adam Crabtree. Isis, 82(2):421, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234925.

Layton:1991:BRB

[604] Edwin T. Layton. Book review: The History of Engineering Science: An Annotated Bibliography by David F. Channell. Isis, 82(2):421–422, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234926.

Caldwell:1991:BRB

[605] Bruce J. Caldwell. Book review: Economic Methodology: A Bibliography with References to Works in the Philosophy of Science, 1860–1988 by Deborah A. Redman. Isis, 82(2):422, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234928.

Neu:1991:BRB

[606] John Neu. Book review: Sources in Electrical History: Archives and Manuscript Collections in U.S. Repositories by Joyce E. Bedi; Ronald R. Kline; Craig Semsel. Isis, 82(2):422, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234927.

Anonymous:1991:Cb

[607] Anonymous. Collections. Isis, 82(2):423–425, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234929.

Anonymous:1991:BMc

[608] Anonymous. Back matter. Isis, 82(2):426–432, June 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234930.

Anonymous:1991:FMd

[609] Anonymous. Front matter. Isis, 82(3):433–439, September 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233224.

Jones:1991:ABM

[610] Alexander Jones. The adaptation of Babylonian methods in Greek numerical astronomy. Isis, 82(3):440–453, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233225.

Warner:1991:IST

[611] John Harley Warner. Ideals of science and their discontents in late Nineteenth-Century American medicine. Isis, 82(3):454–478, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233226.

Kingsland:1991:BBD

[612] Sharon E. Kingsland. The battling botanist: Daniel Trembly MacDougal, mutation theory, and the rise of experimental evolutionary biology in America, 1900–1912. Isis, 82(3):479–509, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233227.

Elena:1991:LDF

[613] Alberto Elena. “In lode della filosofessa di Bologna”: An introduction to Laura Bassi. (Italian) [In praise of philosophy in Bologna …]. Isis, 82 (3):510–518, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233228.

Pyenson:1991:ECJ

[614] Lewis Pyenson. Eloge: Christa Jungnickel, 11 April 1935–12 August 1990. Isis, 82(3):519–520, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233229.

Diamond:1991:LE

[615] Solomon Diamond and John Vallance. Letters to the Editor. Isis, 82(3): 521, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233230.

Daston:1991:HSE

[616] Lorraine Daston. History of science in an elegiac mode: E. A. Burtt’s metaphysical foundations of modern physical science revisited. Isis, 82 (3):522–531, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233231.

Swade:1991:BRW

[617] Doron Swade. Book review: The world reduced to number: The Works of Charles Babbage. Isis, 82(3):532–536, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233232.

Hiebert:1991:BRB

[618] Erwin N. Hiebert. Book review: Science in Sweden: The Royal Swedish Academy of Sciences, 1739–1989 by Tore Frängsmyr. Isis, 82(3):537, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233233.

Olesko:1991:BRBa

[619] Kathryn M. Olesko. Book review: The University and the City: From Medieval Origins to the Present by Thomas Bender. Isis, 82(3):538–539, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233234.

Newell:1991:BRB

[620] Julie R. Newell. Book review: “To Bring Together, Correlate, and Preserve”: A History of the Kansas Geological Survey, 1864–1989 by Rex C. Buchanan. Isis, 82(3):539, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233235.

Hofmann:1991:BRBc

[621] James R. Hofmann. Book review: The Story of Physics by Lloyd Motz; Jefferson Hane Weaver. Isis, 82(3):539–540, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233236.

Brackenridge:1991:BRB

[622] J. Bruce Brackenridge. Book review: Absolute or Relative Motion? A Study from a Machian Point of View of the Discovery and the Structure of Dynamical Theories. Volume I: The Discovery of Dynamics by Julian B. Barbour. Isis, 82(3):540–541, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233237.

Burlingame:1991:BRB

[623] Leslie J. Burlingame. Book review: Sexual Visions: Images of Gender in Science and Medicine between the Eighteenth and Twentieth Centuries by Ludmilla Jordanova. Isis, 82(3):541–542, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233238.

Tobey:1991:BRBb

[624] Ronald Tobey. Book review: Ecological Revolutions: Nature, Gender, and Science in New England by Carolyn Merchant. Isis, 82(3):542–543, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233239.

McNeil:1991:BRB

[625] Maureen McNeil. Book review: The Politics of Women’s Biology by Ruth Hubbard. Isis, 82(3):543–544, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233240.

Koblitz:1991:BRB

[626] Ann Hibner Koblitz. Book review: Women of Science: Righting the Record by G. Kass-Simon; Patricia Farnes; Deborah Nash. Isis, 82(3): 544–545, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233241.

Aay:1991:BRB

[627] Henk Aay. Book review: The Sugar Cane Industry: An Historical Geography from Its Origins to 1914 by J. H. Galloway. Isis, 82(3):545, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233242.

Hannaway:1991:BRBb

[628] Caroline Hannaway. Book review: The Conquest of Water: The Advent of Health in the Industrial Age by Jean–Pierre Goubert; Andrew Wilson. Isis, 82(3):546–547, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233243.

Bromberg:1991:BRB

[629] Joan Lisa Bromberg. Book review: American Genesis: A Century of Invention and Technological Enthusiasm, 1870–1970 by Thomas P. Hughes. Isis, 82(3):547–548, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233244.

VanRiper:1991:BRBa

[630] A. Bowdoin Van Riper. Book review: A History of Archaeological Thought by Bruce G. Trigger. Isis, 82(3):548–549, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233245.

Stiling:1991:BRB

[631] Rodney L. Stiling. Book review: The Flood Myth by Alan Dundes; Noah: The Person and the Story in History and Tradition by Lloyd R. Bailey.

Isis, 82(3):549–550, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233246.

Crowe:1991:BRBb

[632] Michael J. Crowe. Book review: The Last Frontier: Imagining Other Worlds, from the Copernican Revolution to Science Fiction by Karl S. Guthke; Helen Atkins. Isis, 82(3):550–551, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233247.

Clagett:1991:BRB

[633] Marshall Clagett. Book review: The Medical Skills of Ancient Egypt by J. Worth Estes. Isis, 82(3):551–552, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233248.

Santirocco:1991:BRB

[634] Matthew S. Santirocco. Book review: Senecan Drama and Stoic Cosmology by Thomas G. Rosenmeyer. Isis, 82(3):552–553, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233249.

Taub:1991:BRBb

[635] Liba Taub. Book review: Before “Eureka”: The Presocratics and Their Science by Robin Waterfield. Isis, 82(3):553–554, September 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233250.

Siraisi:1991:BRB

[636] Nancy G. Siraisi. Book review: Reappraisals in Renaissance Thought by Charles B. Schmitt; Charles Webster. Isis, 82(3):554–555, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233251.

Sobol:1991:BRBc

[637] Peter Sobol. Book review: Juan Luis Vives and the Emotions by Carlos G. Noreña. Isis, 82(3):555–556, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233252.

Park:1991:BRB

[638] Katharine Park. Book review: Edocere medicos: Medicina scolastica nei secoli XIII-XV by Jole Agrimi; Chiara Crisciani. Isis, 82(3):556–557, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233253.

Reeds:1991:BRB

[639] Karen Reeds. Book review: Not of Woman Born: Representations of Caesarean Birth in Medieval and Renaissance Culture by Renate Blumenfeld-Kosinski. Isis, 82(3):557–558, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233254.

Pinch:1991:BRB

[640] Trevor Pinch. Book review: Explorations in Islamic Science by Ziauddin Sardar. Isis, 82(3):558–559, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233255.

Unguru:1991:BRBb

[641] S. Unguru. Book review: Robert of Chester’s Latin Translation of al-Khwarizmi’s Al-Jabr by Al-Khwarizmi; Barnabas B. Hughes. Isis, 82(3): 559, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233256.

Freudenthal:1991:BRBc

[642] Gad Freudenthal. Book review: Mélanges de philosophie juive et arabe by Solomon Munk. Isis, 82(3):559–560, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233257.

Newman:1991:BRBa

[643] William R. Newman. Book review: Lessico del Paragranum di Theophrast von Hohenheim detto Paracelsus. Volume I: Indici by Massimo Luigi Bianchi. Isis, 82(3):560–561, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233258.

Rattansi:1991:BRB

[644] P. M. Rattansi. Book review: Robert Fludd and the End of the Renaissance by William H. Huffman. Isis, 82(3):561–562, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233259.

Findlen:1991:BRBb

[645] Paula Findlen. Book review: L’influsso scientifico di Robert Boyle nel tardo ’600 italiano by Clelia Pighetti; Da Descartes a Spinoza: Percorsi della teologia razionale nel Seicento by Maria Emanuela Scribano. Isis, 82 (3):563–564, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233260.

Shapin:1991:BRB

[646] Steven Shapin. Book review: The Library of Robert Hooke: The Scientific Book Trade of Restoration England by Leona Rostenberg. Isis, 82(3): 564–565, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233261.

Ehrlich:1991:BRB

[647] Mark Ehrlich. Book review: Robert Hooke: New Studies by Michael Hunter; Simon Schaffer. Isis, 82(3):565–566, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233262.

Maas:1991:BRB

[648] Jorg F. Maas. Book review: Disputationes Hamburgenses: Kritische Edition by Joachim Jungius; Clemens Müller–Glauser. Isis, 82(3):566, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233263.

Yoder:1991:BRB

[649] Joella Yoder. Book review: History of Free Fall: Aristotle to Galileo: With an Epilogue on Pi in the Sky by Stillman Drake; Essays on the Trial of Galileo by Richard S. Westfall. Isis, 82(3):566–567, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233264.

Challey:1991:BRB

[650] James F. Challey. Book review: The Philosophy of Science of Ruder Bosković: Proceedings of the Symposium of the Institute of Philosophy and Theology, S.J. Isis, 82(3):567–568, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233265.

Farber:1991:BRB

[651] Paul Lawrence Farber. Book review: Buffon: Un philosophe au Jardin du Roi by Jacques Roger. Isis, 82(3):568–569, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233266.

Schiebinger:1991:BRBa

[652] Londa Schiebinger. Book review: The Body and the French Revolution: Sex, Class, and Political Culture by Dorinda Outram. Isis, 82(3):569–570, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233267.

Taylor:1991:BRB

[653] Kenneth L. Taylor. Book review: La mutation de l’enseignement scientifique en France (1750–1810) et le rôle des écoles centrales: L’exemple de Nantes by Pierre Lamandé. Isis, 82(3):570, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233268.

Elliott:1991:BRBb

[654] Eric Elliott. Book review: Ein Fundament zum Gebäude der Wissenschaften: Einhundert Jahre Ostwalds Klassiker der exakten Wissenschaften (1889–1989) by Lothar Dunsch; Hella Müller; Selbstorganisation chemischer Strukturen by Friedlieb Ferdinand Runge; Raphael Eduard Liesegang; Boris Pavlovich Belousov; Anatol Markovich Zhabotinsky; Lothar Kuhnert; Uwe Niedersen; Zur Konformation des Cyclohexans by Hermann Sachse; Ernst Mohr; Horst Remane; Klassische Schriften zur Zellenlehre by Matthias Jacob Schleiden; Theodor Schwann; Max Schultze; Ilse Jahn. Isis, 82(3):570–571, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233269.

Coon:1991:BRB

[655] Deborah J. Coon. Book review: William James: Public Philosopher by George Cotkin. Isis, 82(3):571–572, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233270.

Gregory:1991:BRB

[656] Frederick Gregory. Book review: Schelling’s und Hegel’s Verhältnis der Naturwissenschaft: Zum Verhältnis der physikalistischen Naturwissenschaft zur spekulativen Naturphilosophie by Matthias Jakob Schleiden; Olaf Breidbach. Isis, 82(3):572–573, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233271.

Paul:1991:BRBc

[657] H. W. Paul. Book review: L’organisation de l’enseignement des sciences: La voie ouverte par le Second Empire by Nicole Hulin-Jung. Isis, 82(3): 573–574, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233272.

Vermij:1991:BRB

[658] R. H. Vermij. Book review: “Cijfers en Aequaties” en “Kennis der Staatskrachten”: Statistiek in Nederland in de negentiende eeuw by Ida H. Stamhuis. Isis, 82(3):574, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233273.

Miller:1991:BRBa

[659] Jane A. Miller. Book review: Boussingault entre Lavoisier et Pasteur: Biographie cordiale by Ernest Kahane. Isis, 82(3):574–575, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233274.

Nyhart:1991:BRBa

[660] Lynn K. Nyhart. Book review: Genèse de la théorie cellulaire by François Duchesneau. Isis, 82(3):575–576, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233275.

Turner:1991:BRBb

[661] Frank M. Turner. Book review: History, Humanity, and Evolution: Essays for John C. Greene by James R. Moore. Isis, 82(3):576–578, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233276.

Erskine:1991:BRB

[662] Fiona Erskine. Book review: Charles Darwin: A Biography by John Bowlby. Isis, 82(3):578–579, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233277.

Cittadino:1991:BRB

[663] Eugene Cittadino. Book review: Darwinismus und Botanik: Rezeption, Kritik und theoretische Alternativen im Deutschland des 19. Jahrhunderts by Thomas Junker. Isis, 82(3):579–580, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233278.

Harrington:1991:BRB

[664] Anne Harrington. Book review: Regions of the Mind: Brain Research and the Quest for Scientific Certainty by Susan Leigh Star. Isis, 82(3): 580–581, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233279.

Rousseau:1991:BRBb

[665] G. S. Rousseau. Book review: Retreat into the Mind: Victorian Poetry and the Rise of Psychiatry by Ekbert Faas. Isis, 82(3):581–582, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233280.

Melnick:1991:BRB

[666] Jeffrey Melnick. Book review: Notes on the Underground: An Essay on Technology, Society, and the Imagination by Rosalind Williams. Isis, 82 (3):582–583, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233281.

Pursell:1991:BRB

[667] Carroll Pursell. Book review: Power from Steam: A History of the Stationary Steam Engine by Richard L. Hills. Isis, 82(3):583–584, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233282.

Roberts:1991:BRB

[668] Jon H. Roberts. Book review: The Transformation of Theology, 1830– 1890: Positivism and Protestant Thought in Britain and America by Charles D. Cashdollar. Isis, 82(3):584–585, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233283.

Feldberg:1991:BRB

[669] Georgiana Feldberg. Book review: E. W. R. Steacie and Science in Canada by M. Christine King. Isis, 82(3):585–586, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233284.

Whorton:1991:BRB

[670] James Whorton. Book review: Pure Food: Securing the Pure Food and Drugs Act of 1906 by James Harvey Young. Isis, 82(3):586–588, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233285.

Hetherington:1991:BRBb

[671] Norriss S. Hetherington. Book review: The Invented Universe: The Einstein–De Sitter Controversy (1916–17) and the Rise of Relativistic Cosmology by Pierre Kerszberg. Isis, 82(3):588–589, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233286.

Ziegler:1991:BRS

[672] Charles Ziegler. Book review: Stanley A. Blumberg and Louis G. Panos: Edward Teller: Giant of the Golden Age of Physics. Herbert F. York: The Advisors: Oppenheimer, Teller, and the Superbomb. J. Robert Oppenheimer: Atom and Void: Essays on Science and Community. Isis, 82 (3):589–590, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233287.

Fabian:1991:BRBb

[673] Eginhard Fabian. Book review: Heinrich Heesch: Kristallgeometrie, Parkettierungen, Vierfarbenforschung by Hans–Günther Bigalke. Isis, 82 (3):590–591, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233288.

Gatewood:1991:BRB

[674] Willard B. Gatewood, Jr. Book review: Trial and Error: The American Controversy over Creation and Evolution by Edward J. Larson. Isis, 82 (3):591–592, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233289.

Hurt:1991:BRB

[675] R. Douglas Hurt. Book review: The Business of Breeding: Hybrid Corn in Illinois, 1890–1940 by Deborah Fitzgerald. Isis, 82(3):592–593, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233290.

vanHeteren:1991:BRBa

[676] G. van Heteren. Book review: Health, Race, and German Politics between National Unification and Nazism, 1870–1945 by Paul Weindling. Isis, 82 (3):593–595, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233291.

Weiss:1991:BRB

[677] Sheila Faith Weiss. Book review: Doctors under Hitler by Michael H. Kater. Isis, 82(3):595–596, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233292.

Walker:1991:BRBc

[678] Mark Walker. Book review: Science, Technology, and Reparations: Exploitation and Plunder in Postwar Germany by John Gimbel. Isis, 82 (3):596–597, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233293.

Sturdy:1991:BRB

[679] Steve Sturdy. Book review: Historical Perspectives on the Role of the MRC: Essays in the History of the Medical Research Council of the United Kingdom and Its Predecessor, the Medical Research Committee, 1913–1953 by Joan Austoker; Linda Bryder. Isis, 82(3):597–598, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233294.

Pisano:1991:BRB

[680] Dominick A. Pisano. Book review: Technical Workers in an Advanced Society: The Work, Careers, and Politics of French Engineers by Stephen Crawford. Isis, 82(3):598, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233295.

Misa:1991:BRB

[681] Thomas J. Misa. Book review: Technology and Transformation in the American Electric Utility Industry by Richard F. Hirsh. Isis, 82(3):598– 600, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233296.

Smith:1991:BRB

[682] Harlan J. Smith. Book review: The Space Telescope: A Study of NASA, Science, Technology, and Politics by Robert W. Smith; Paul A. Hanle; Robert H. Kargon; Joseph N. Tatarewicz. Isis, 82(3):600–601, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233297.

Fuller:1991:BRBa

[683] Steve Fuller. Book review: The Rational and the Social by James Robert Brown. Isis, 82(3):601–602, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233298.

Jacobs:1991:BRB

[684] Struan Jacobs. Book review: Life among the Scientists: An Anthropological Study of an Australian Scientific Community by Max Charlesworth; Lyndsay Farrall; Terry Stokes; David Turnbull. Isis, 82(3):602, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233299.

Ravetz:1991:BRB

[685] Jerome R. Ravetz. Book review: Science as Social Knowledge: Values and Objectivity in Scientific Inquiry by Helen E. Longino. Isis, 82(3): 602–603, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233300.

MacIntyre:1991:BRB

[686] Alasdair MacIntyre. Book review: Philosophy and the Spontaneous Philosophy of the Scientists and Other Essays by Louis Althusser; Gregory Elliott; Ben Brewster; James H. Kavanagh; Thomas E. Lewis; Grahame Lock; Warren Montag. Isis, 82(3):603–604, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233301.

Ariew:1991:BRBb

[687] Roger Ariew. Book review: Duhem: Science, réalité et apparence: La relation entre philosophie et histoire dans l’oeuvre de Pierre Duhem by Anastasios Brenner. Isis, 82(3):604, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233302.

Gutting:1991:BRB

[688] Gary Gutting. Book review: Husserl’s Phenomenology and the Foundations of Natural Science by Charles W. Harvey. Isis, 82(3):604–605, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233303.

Price:1991:BRB

[689] Huw Price. Book review: Nature’s Capacities and Their Measurement by Nancy Cartwright. Isis, 82(3):605–606, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233304.

Bub:1991:BRBb

[690] Jeffrey Bub. Book review: The Philosophy of Quantum Mechanics: An Interactive Interpretation by Richard Healey. Isis, 82(3):606–607, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233305.

Forge:1991:BRB

[691] John Forge. Book review: The Semantic Conception of Theories and Scientific Realism by Frederick Suppe. Isis, 82(3):607–608, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233306.

Reiss:1991:BRB

[692] Timothy J. Reiss. Book review: Dionysus Reborn: Play and the Aesthetic Dimension in Modern Philosophical and Scientific Discourse by Mihai I. Spariosu. Isis, 82(3):608–609, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233307.

Harvey:1991:BRB

[693] Joy Harvey. Book review: Ideology and Rationality in the History of the Life Sciences by Georges Canguilhem; Arthur Goldhammer. Isis, 82(3):610, September 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233308.

DesChene:1991:BRB

[694] Dennis Des Chene. Book review: Michel Foucault’s Archaeology of Scientific Reason by Gary Gutting. Isis, 82(3):610–611, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233309.

Morman:1991:BRB

[695] Edward T. Morman. Book review: Biographical Index to American Science: The Seventeenth Century to 1920 by Clark A. Elliott. Isis, 82(3): 611, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233310.

Elliott:1991:BRBc

[696] Clark A. Elliott. Book review: Sci-Tech Archives and Manuscript Collections by Ellis Mount. Isis, 82(3):611–612, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233311.

Stephens:1991:BRB

[697] Carlene Stephens. Book review: Brass and Glass: Scientific Instrument Making Workshops in Scotland as Illustrated by Instruments from the Arthur Frank Collection at the Royal Museum of Scotland by T. N. Clarke; A. D. Morrison-Low; A. D. C. Simpson. Isis, 82(3):612–613, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233312.

Wagner:1991:BRB

[698] Stephen C. Wagner. Book review: Molecules, Cells, and Life: An Annotated Bibliography of Manuscript Sources on Physiology, Biochemistry, and Biophysics, 1900–1960, in the Library of the American Philosophical Society by Lily E. Kay. Isis, 82(3):613, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233313.

Whitcomb:1991:BRBa

[699] Dorothy Whitcomb. Book review: Rare and Important Books in the Library of the Swedish Society of Medicine: A Descriptive and Annotated Catalogue by Ove Hagelin. Isis, 82(3):613–614, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233314.

Anonymous:1991:BMd

[700] Anonymous. Back matter. Isis, 82(3):615–620, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233315.

Anonymous:1991:FMe

[701] Anonymous. Front matter. Isis, 82(4):621–627, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233317.

Laird:1991:AAH

[702] W. R. Laird. Archimedes among the humanists. Isis, 82(4):628–638, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233318.

Cormack:1991:GFM

[703] Lesley B. Cormack. “Good fences make good neighbors”: Geography as self-definition in Early Modern England. Isis, 82(4):639–661, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233319.

Pauly:1991:DHS

[704] Philip J. Pauly. The development of high school biology: New York City, 1900–1925. Isis, 82(4):662–688, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233320.

Brackenridge:1991:EEJ

[705] J. Bruce Brackenridge. Eloge: Eric John Aiton, 1920–21 February 1991. Isis, 82(4):689–691, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233321.

Sloan:1991:EJR

[706] Phillip R. Sloan. Eloge: Jacques Roger, 24 October 1920–26 March 1990. Isis, 82(4):691–693, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233322.

Gingerich:1991:EVT

[707] Owen Gingerich. Eloge: Victor E. Thoren, 13 May 1935–9 March 1991. Isis, 82(4):693–694, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233323.

North:1991:LE

[708] J. D. North, Richard Lemay, Louis I. Kuslan, and M. Susan Lindee. Letters to the Editor. Isis, 82(4):695–697, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233324.

Allen:1991:BRB

[709] Garland E. Allen. Book review: Science as a Process: An Evolutionary Account of the Social and Conceptual Development of Science by David L. Hull; The Metaphysics of Evolution by David L. Hull. Isis, 82(4): 698–704, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233325.

Hatch:1991:BRB

[710] Robert A. Hatch. Book review: Theories of the World from Antiquity to the Copernican Revolution by Michael J. Crowe. Isis, 82(4):705, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233326.

Osler:1991:BRBb

[711] Margaret J. Osler. Book review: The Epicurean Tradition by Howard Jones. Isis, 82(4):706, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233327.

Olesko:1991:BRBb

[712] Kathryn M. Olesko. Book review: Wissenschaft in Berlin: Von den Anfängen bis zum Neubeginn nach 1945. Isis, 82(4):706–707, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233328.

Bedini:1991:BRB

[713] Silvio A. Bedini. Book review: “Vulgar and Mechanick”: The Scientific Instrument Trade in Ireland, 1650–1921 by J. E. Burnett; A. D. Morrison-Low. Isis, 82(4):707–708, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233329.

Feingold:1991:BRBb

[714] Mordechai Feingold. Book review: Forgers and Critics: Creativity and Duplicity in Western Scholarship by Anthony Grafton. Isis, 82(4):709, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233330.

Badham:1991:BRB

[715] Richard J. Badham. Book review: The Gene Hunters: Biotechnology and the Scramble for Seeds by Calestous Juma. Isis, 82(4):709–710, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233331.

Burnstein:1991:BRB

[716] Daniel Eli Burnstein. Book review: The Sanitarians: A History of American Public Health by John Duffy. Isis, 82(4):710–711, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233332.

Overmann:1991:BRB

[717] Ronald J. Overmann. Book review: Peerless Science: Peer Review and U.S. Science Policy by Daryl E. Chubin; Edward J. Hackett. Isis, 82 (4):711–712, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233333.

Christianson:1991:BRB

[718] Eric Howard Christianson. Book review: Naissance et développement de la science-monde: Production et reproduction des communautés scientifiques en Europe et en Amérique latine by Xavier Polanco. Isis, 82 (4):712–713, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233334.

Anastaplo:1991:BRB

[719] George M. D. Anastaplo. Book review: The Ethics of Geometry: A Genealogy of Modernity by David Rapport Lachterman. Isis, 82(4):713– 714, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233335.

Pierson:1991:BRB

[720] Stuart Pierson. Book review: Science and Philosophy: Past and Present by Derek Gjertsen. Isis, 82(4):714–715, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233336.

Stocking:1991:BRB

[721] George W. Stocking, Jr. Book review: Localizing Strategies: Regional Traditions of Ethnographic Writing by Richard Fardon. Isis, 82(4):715– 716, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233337.

Peterfreund:1991:BRB

[722] Stuart Peterfreund. Book review: Nature Transfigured: Science and Literature, 1700–1900 by John Christie; Sally Shuttleworth; The Place of Fiction in the Time of Science: A Disciplinary History of American Writing by John Limon. Isis, 82(4):716–718, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233338.

Drucker:1991:BRB

[723] Thomas Drucker. Book review: The Ancient Tradition of Geometric Problems by Wilbur Richard Knorr; Textual Studies in Ancient and Medieval Geometry by Wilbur Richard Knorr. Isis, 82(4):718–720, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233339.

Gottlieb:1991:BRB

[724] Paula Gottlieb. Book review: Cosmic and Meta-Cosmic Theology in Aristotle’s Lost Dialogues by A. P. Bos. Isis, 82(4):720–721, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233340.

Knorr:1991:BRB

[725] Wilbur R. Knorr. Book review: Greek and Roman Maps by O. A. W. Dilke. Isis, 82(4):721–722, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233341.

Colish:1991:BRB

[726] Marcia L. Colish. Book review: The Philosophy of John Scottus Eriugena: A Study of Idealism in the Middle Ages by Dermot Moran; Eriugena by John J. O’Meara. Isis, 82(4):722–724, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233342.

Peterson:1991:BRB

[727] Kristin E. Peterson. Book review: Liber tertius naturalium: De generatione et corruptione: Édition critique de la traduction latine médiévale et lexiques by Avicenna; S. van Riet. Isis, 82(4):724, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233343.

Jones:1991:BRBd

[728] Alexander Jones. Book review: The Optics of Ibn al-Haytham, Books I-III: On Direct Vision by Ibn al-Haytham; A. I. Sabra. Isis, 82(4): 724–726, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233344.

Sobol:1991:BRBd

[729] Peter G. Sobol. Book review: Pour l’histoire du sens agent: La controverse entre Barthélemy de Bruges et Jean de Jandun, ses antécédents et son évolution by Adrien Pattin. Isis, 82(4):726, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233345.

Newman:1991:BRBb

[730] William R. Newman. Book review: Johannes’ de Rupescissa “Liber de consideratione quintae essentiae omnium rerum” deutsch: Studien zur Al-chemia medica des 15. bis 17. Jahrhunderts mit kritischer Edition des Textes by Udo Benzenhöfer. Isis, 82(4):726–727, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233346.

McVaugh:1991:BRBb

[731] Michael R. McVaugh. Book review: Practica medicinalis: A Critical Edition of the “Practica medicinalis” of Thomas of Wroclaw, Premontré Bishop of Sarepta (1297–c. 1378) by Thomas of Wroclaw; Theodore James Antry. Isis, 82(4):727–728, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233347.

Vickers:1991:BRB

[732] Brian Vickers. Book review: Magic in the Middle Ages by Richard Kieckhefer; Natural Magic and Modern Science: Four Treatises, 1590–1657 by Wayne Shumaker; Magic, Science, Religion, and the Scope of Rationality by Stanley Jeyaraja Tambiah. Isis, 82(4):728–731, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233348.

Knoespel:1991:BRB

[733] Kenneth Knoespel. Book review: Renaissance Magic and the Return of the Golden Age: The Occult Tradition and Marlowe, Jonson, and Shakespeare by John S. Mebane. Isis, 82(4):731–732, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233349.

Barnes:1991:BRB

[734] Carl F. Barnes, Jr. Book review: Light, Wind, and Structure: The Mystery of the Master Builders by Robert Mark; The Architectural Theory of Viollet-le-Duc: Readings and Commentary by Eugène–Emmanuel Viollet-le-Duc; M. F. Hearn. Isis, 82(4):732–733, December 1991. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233350.

Getz:1991:BRB

[735] Faye M. Getz. Book review: Medieval and Early Renaissance Medicine: An Introduction to Knowledge and Practice by Nancy G. Siraisi. Isis, 82 (4):733–734, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233351.

Browne:1991:BRB

[736] Alice Browne. Book review: Tommaso Campanella: Sein Beitrag zur Medizin und Pharmazie der Renaissance by Michael W. Mönnich. Isis, 82(4):734–735, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233352.

Williams:1991:BRB

[737] William H. Williams. Book review: Cartesian Logic: An Essay on Descartes’s Conception of Inference by Stephen Gaukroger; Arnauld and the Cartesian Philosophy of Ideas by Steven M. Nadler. Isis, 82(4):735– 737, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233353.

Hill:1991:BRB

[738] David K. Hill. Book review: Galileo: Pioneer Scientist by Stillman Drake. Isis, 82(4):737–738, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233354.

Swerdlow:1991:BRB

[739] N. M. Swerdlow. Book review: Planetary Astronomy from the Renaissance to the Rise of Astrophysics. Part A: Tycho Brahe to Newton by René Taton; Curtis Wilson. Isis, 82(4):738–740, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233355.

Apt:1991:BRBb

[740] Adam J. Apt. Book review: Mundus Iovialis: Die Welt des Jupiter by Simon Marius; Joachim Schlor. Isis, 82(4):740, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233356.

Westfall:1991:BRB

[741] Richard S. Westfall. Book review: Before Newton: The Life and Times of Isaac Barrow by Mordechai Feingold. Isis, 82(4):740–741, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233357.

Palter:1991:BRB

[742] Robert Palter. Book review: Philosophical Perspectives on Newtonian Science by Phillip Bricker; R. I. G. Hughes. Isis, 82(4):741–742, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233358.

Findlen:1991:BRBc

[743] Paula Findlen. Book review: Ulisse Aldrovandi e la Toscana: Carteggio e testimonianze documentarie by Alessandro Tosi. Isis, 82(4):742–743, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233359.

Cook:1991:BRBb

[744] Harold J. Cook. Book review: Italian Scientists in the Low Countries in the Seventeenth and Eighteenth Centuries by C. S. Maffioli; L. C. Palm. Isis, 82(4):743–745, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233360.

Stroup:1991:BRB

[745] Alice Stroup. Book review: Patronage and Royal Science in Seventeenth-Century France: The Academie de Physique in Caen by David S. Lux. Isis, 82(4):745–746, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233361.

Riddle:1991:BRB

[746] John M. Riddle. Book review: Drogas americanas en fuentes de escritores franciscanos y dominicos by José Luis Valverde; José A. Perez Romero. Isis, 82(4):746–747, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233362.

Nadler:1991:BRB

[747] Steven Nadler. Book review: The Light of the Soul: Theories of Ideas in Leibniz, Malebranche, and Descartes by Nicholas Jolley. Isis, 82(4): 747–748, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233363.

Degen:1991:BRB

[748] Peter A. Degen. Book review: The Theology of Electricity: On the Encounter and Explanation of Theology and Science in the Seventeenth and Eighteenth Centuries by Ernst Benz; Dennis Stillings; Wolfgang Taraba. Isis, 82(4):748–749, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233364.

Hankins:1991:BRB

[749] Thomas L. Hankins. Book review: Sciences à l’époque de la Révolution française: Recherches historiques by Roshdi Rashed. Isis, 82(4):749, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233365.

McGrath:1991:BRB

[750] Sylvia W. McGrath. Book review: John Lightfoot, His Work and Travels: With a Biographical Introduction and a Catalogue of the Lightfoot Herbarium by Jean K. Bowden. Isis, 82(4):750, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233366.

Groopman:1991:BRB

[751] Leonard C. Groopman. Book review: The Charitable Imperative: Hospitals and Nursing in Ancien Régime and Revolutionary France by Colin Jones. Isis, 82(4):750–751, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233367.

Blustein:1991:BRB

[752] Bonnie Ellen Blustein. Book review: Neurociencias y revolución científica en España: La circulación neural by Alvaro Martínez Vidal; Francisco Méndez Alvaro (1806–1883) y las ideas sanitarias del liberalismo moderado by José Luis Fresquet Febrer. Isis, 82(4):751–752, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233368.

Woodward:1991:BRBb

[753] William R. Woodward. Book review: Knowledge, Belief, and Aesthetic Sense by J. F. Fries; Frederick Gregory; Kent Richter. Isis, 82(4):752– 753, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233369.

Broman:1991:BRB

[754] Thomas Broman. Book review: Romanticism and the Sciences by Andrew Cunningham; Nicholas Jardine. Isis, 82(4):753–754, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233370.

Wilson:1991:BRB

[755] Andrew Wilson. Book review: Ludwig Boltzmanns Wege nach Berlin: Ein Kapital österreichisch-deutscher Wissenschaftsbeziehungen by Herbert Hörz; Andreas Laass; Ludwig Boltzmann: Altmeister der klassischen Physik by Wolfgang Stiller. Isis, 82(4):754–755, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233371.

Paul:1991:BRBd

[756] Erich Robert Paul. Book review: Resolute and Undertaking Characters: The Lives of Wilhelm and Otto Struve by Alan H. Batten. Isis, 82(4): 755–756, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233372.

VanRiper:1991:BRBb

[757] A. Bowdoin Van Riper. Book review: James David Forbes: Pioneer Scottish Glaciologist by Frank F. Cunningham. Isis, 82(4):756–757, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233373.

Montgomery:1991:BRBb

[758] William Montgomery. Book review: Der Bruch mit der humanitären Tradition: Die Biologisierung der Ethik bei Ernst Haeckel und anderen Darwinisten seiner Zeit by Jürgen Sandmann. Isis, 82(4):757–758, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233374.

Nyhart:1991:BRBb

[759] Lynn K. Nyhart. Book review: Nature as the Laboratory: Darwinian Plant Ecology in the German Empire, 1880–1900 by Eugene Cittadino. Isis, 82(4):758–759, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233375.

Osborne:1991:BRBb

[760] Michael A. Osborne. Book review: Lives at Risk: Public Health in Nineteenth-Century Egypt by LaVerne Kuhnke. Isis, 82(4):759, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233376.

vanHeteren:1991:BRBb

[761] G. van Heteren. Book review: De opkomst van het medisch beroep in Belgie by Rita Schepers. Isis, 82(4):759–760, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233377.

Warner:1991:BRB

[762] John Harley Warner. Book review: Health Care in the Parisian Countryside, 1800–1914 by Evelyn Bernette Ackerman. Isis, 82(4):760–761, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233378.

Nier:1991:BRB

[763] Keith A. Nier. Book review: Edison and the Business of Innovation by Andre Millard. Isis, 82(4):761–762, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233379.

Morantz-Sanchez:1991:BRB

[764] Regina Morantz-Sánchez. Book review: The Darkened Room: Women, Power, and Spiritualism in Late Victorian England by Alex Owen. Isis, 82(4):762–763, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233380.

Schiebinger:1991:BRBb

[765] Londa Schiebinger. Book review: The Science of Woman: Gynaecology and Gender in England, 1800–1929 by Ornella Moscucci. Isis, 82(4): 763–764, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233381.

Richards:1991:BRBb

[766] Joan L. Richards. Book review: Alfred North Whitehead: The Man and His Work. Volume II: 1910–1947 by Victor Lowe; J. B. Schneewind. Isis, 82(4):764–765, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233382.

Dauben:1991:BRB

[767] Joseph W. Dauben. Book review: A Century of Mathematics in America by Peter Duren; Richard A. Askey; Uta C. Merzbach. Isis, 82(4):765– 768, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233383.

Harman:1991:BRA

[768] P. M. Harman. Book review: Albert Einstein and John Stachel: The Collected Papers of Albert Einstein. Volume II: The Swiss Years: Writings, 1900–1909. Albert Einstein and Anna Beck: The Collected Papers of Albert Einstein. Volume II: The Swiss Years: Writings, 1990–1909. English Translation. Isis, 82(4):768–769, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233384.

Bergmann:1991:BRB

[769] Peter G. Bergmann. Book review: Einstein and the History of General Relativity by Don Howard; John Stachel. Isis, 82(4):769, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233385.

Brown:1991:BRBb

[770] Laurie M. Brown. Book review: Dirac: A Scientific Biography by Helge Kragh. Isis, 82(4):769–770, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233386.

Goodstein:1991:BRL

[771] Judith Goodstein. Book review: Lanfranco Belloni: Da Fermi a Rubbia: Storia e politica di un successo mondiale della scienza italiana. Isis, 82 (4):770–771, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233387.

Nye:1991:BRBb

[772] Mary Jo Nye. Book review: Physical Chemistry from Ostwald to Pauling: The Making of a Science in America by John W. Servos. Isis, 82(4):771– 772, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233388.

Ihde:1991:BRB

[773] Aaron J. Ihde. Book review: Robert Robinson, Chemist Extraordinary by Trevor I. Williams. Isis, 82(4):772–773, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233389.

Paul:1991:BRBe

[774] Diane B. Paul. Book review: The Mendelian Revolution: The Emergence of Hereditarian Concepts in Modern Science and Society by Peter Bowler. Isis, 82(4):773–774, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233390.

Frankel:1991:BRB

[775] Simon J. Frankel. Book review: Drifting Continents and Shifting Theories: The Modern Revolution in Geology and Scientific Change by H. E.

Le Grand. Isis, 82(4):774–775, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233391.

Oreskes:1991:BRB

[776] Naomi Oreskes. Book review: Drifting Continents and Colliding Paradigms: Perspectives on the Geoscience Revolution by John A. Stewart. Isis, 82(4):775–776, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233392.

Mark:1991:BRB

[777] Joan Mark. Book review: Edward Sapir: Linguist, Anthropologist, Humanist by Regna Darnell. Isis, 82(4):776, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233393.

Carson:1991:BRB

[778] John Carson. Book review: Metaphors in the History of Psychology by David E. Leary. Isis, 82(4):776–777, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233394.

vonMayrhauser:1991:BRB

[779] Richard T. von Mayrhauser. Book review: Defining American Psychology: The Correspondence between Adolf Meyer and Edward Bradford Titchener by Ruth Leys; Rand B. Evans. Isis, 82(4):777–778, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233395.

Jordan:1991:BRB

[780] John M. Jordan. Book review: John Dewey’s Pragmatic Technology by Larry A. Hickman. Isis, 82(4):778–779, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233396.

Donnelly:1991:BRB

[781] J. F. Donnelly. Book review: The American Synthetic Rubber Research Program by Peter J. T. Morris. Isis, 82(4):779–780, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233397.

McMullin:1991:BRB

[782] Ernan McMullin. Book review: Our Place in the Universe: A Metaphysical Discussion by J. J. C. Smart. Isis, 82(4):780–781, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233398.

Finocchiaro:1991:BRBb

[783] Maurice A. Finocchiaro. Book review: Scientific Explanation by Philip Kitcher; Wesley C. Salmon; Four Decades of Scientific Explanation by Wesley C. Salmon. Isis, 82(4):781–782, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233399.

Olson:1991:BRB

[784] Richard Olson. Book review: The Natures of Science by Neville Mc-Morris. Isis, 82(4):782–783, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233400.

Lowe:1991:BRB

[785] E. J. Lowe. Book review: The Chances of Explanation: Causal Explanation in the Social, Medical, and Physical Sciences by Paul Humphreys. Isis, 82(4):783–784, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233401.

Tang:1991:BRB

[786] Paul C. L. Tang. Book review: Die Wissenschaftsphilosophie Thomas S. Kuhns: Rekonstruktion und Grundlagenprobleme by Paul Hoyningen-Huene. Isis, 82(4):784–785, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233402.

Miller:1991:BRBb

[787] David Miller. Book review: Truth, Hope, and Power: The Thought of Karl Popper by Douglas E. Williams. Isis, 82(4):785–786, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233403.

Fuller:1991:BRBb

[788] Steve Fuller. Book review: Science and Its Fabrication by Alan Chalmers. Isis, 82(4):786–787, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233404.

Ashmore:1991:BRB

[789] Malcolm Ashmore. Book review: Reflexive Epistemology: The Philosophical Legacy of Otto Neurath by Danilo Zolo; David McKie. Isis, 82 (4):787–788, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233405.

Mayo:1991:BRB

[790] Deborah G. Mayo. Book review: Scientific Reasoning: The Bayesian Approach by Colin Howson; Peter Urbach. Isis, 82(4):788–789, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233406.

DeMey:1991:BRB

[791] Marc De Mey. Book review: Psychogenesis and the History of Science by Jean Piaget; Rolando Garcia; Helga Feider. Isis, 82(4):789–791, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233407.

Myers:1991:BRB

[792] Greg Myers. Book review: Narrative Experiments: The Discursive Authority of Science and Technology by Gayle L. Ormiston; Raphael Sassower. Isis, 82(4):791–792, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233408.

Henriksen:1991:BRB

[793] Paul W. Henriksen. Book review: Historical Atlas of Crystallography by J. Lima-de-Faria. Isis, 82(4):792–793, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233409.

Whitcomb:1991:BRBb

[794] Dorothy Whitcomb. Book review: A Bibliography of Medical and Biomedical Biography by Leslie T. Morton; Robert J. Moore. Isis, 82 (4):792–794, December 1991. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233410.

Humphreys:1991:BRBb

[795] Margaret Humphreys. Book review: An Annotated Catalog of the Miner Yellow Fever Collection by Christopher Hoolihan. Isis, 82(4):794, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233411.

Anonymous:1991:Cc

[796] Anonymous. Collections. Isis, 82(4):794–796, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233412.

Anonymous:1991:BMe

[797] Anonymous. Back matter. Isis, 82(4):797–832, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233413.

Anonymous:1991:VI

[798] Anonymous. Volume information. Isis, 82(4):808–831, December 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233316.

Neu:1992:CBH

[799] John Neu. Current bibliography of the history of science and its cultural influences, 1992. Isis, 83(??):1–286, ???? 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235815.

Anonymous:1992:BMa

[800] Anonymous. Back matter. Isis, 83(??):??, ???? 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235816.

Anonymous:1992:FMa

[801] Anonymous. Front matter. Isis, 83(??):i–xvi, ???? 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235814.

Numbers:1992:E

[802] Ronald L. Numbers. Editorial. Isis, 83(1):1, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233990.

Colp:1992:WGD

[803] Ralph Colp, Jr. “I will gladly do my best”: How Charles Darwin obtained a civil list pension for Alfred Russel Wallace. Isis, 83(1):2–26, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233991.

Holton:1992:EMF

[804] Gerald Holton. Ernst Mach and the fortunes of positivism in America. Isis, 83(1):27–60, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233992.

Lederer:1992:PAS

[805] Susan E. Lederer. Political animals: The shaping of biomedical research literature in Twentieth–Century America. Isis, 83(1):61–79, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233993.

Hill:1992:LE

[806] David K. Hill, Ron Naylor, Lissa Roberts, Olga Amsterdamska, and Paul Forman. Letters to the Editor. Isis, 83(1):80–83, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233994.

Eastwood:1992:CWS

[807] Bruce S. Eastwood. On the continuity of Western science from the Middle Ages: A. C. Crombie’s Augustine to Galileo. Isis, 83(1):84–99, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233995.

Melia:1992:BRB

[808] Trevor Melia. Book review: The Literary Structure of Scientific Argument: Historical Studies by Peter Dear; The Rhetoric of Science by Alan G. Gross; Writing Biology: Texts in the Social Construction of Scientific Knowledge by Greg Myers; A Rhetoric of Science: Inventing Scientific Discourse by Lawrence J. Prelli. Isis, 83(1):100–106, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233996.

Cross:1992:BRB

[809] Stephen J. Cross. Book review: Interdisciplinarity: History, Theory, and Practice by Julie Thompson Klein. Isis, 83(1):107, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233997.

Creed:1992:BRBa

[810] Walter Creed. Book review: Chaos Bound: Orderly Disorder in Contemporary Literature and Science by N. Katherine Hayles. Isis, 83(1): 107–108, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233998.

Shea:1992:BRBa

[811] William R. Shea. Book review: Caos e armonia: Storia della fisica moderna e contemporanea by Enrico Bellone. Isis, 83(1):108–109, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233999.

vanDommelen:1992:BRB

[812] Ad van Dommelen. Book review: Geschichte der Ökologie: Vom 17. Jahrhundert bis zur Gegenwart: Zehn Vorlesungen by Ludwig Trepl. Isis, 83(1):109–110, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234000.

Laudan:1992:BRB

[813] Rachel Laudan. Book review: History of Geomorphology: From Hutton to Hack by K. J. Tinkler. Isis, 83(1):110–111, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234001.

Kidwell:1992:BRB

[814] Peggy Aldrich Kidwell. Book review: Revolution in Measurement: Western European Weights and Measures since the Age of Science by Ronald Edward Zupko. Isis, 83(1):111, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234002.

Rothman:1992:BRB

[815] David J. Rothman. Book review: The Hospital in History by Lindsay Granshaw; Roy Porter; The American General Hospital: Communities and Social Contexts by Diana Elizabeth Long; Janet Golden. Isis, 83 (1):111–113, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234003.

Lurie:1992:BRB

[816] Edward Lurie. Book review: Knowledge, Culture, and Science in the Metropolis: The New York Academy of Sciences, 1817–1970 by Simon Baatz. Isis, 83(1):113–114, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234004.

Hounshell:1992:BRB

[817] David A. Hounshell. Book review: Technology in America: A Brief History by Alan I. Marcus; Howard P. Segal. Isis, 83(1):114–115, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234005.

Basalla:1992:BRB

[818] George Basalla. Book review: An Encyclopaedia of the History of Technology by Ian McNeil. Isis, 83(1):115–117, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234006.

Allen:1992:BRB

[819] James P. Allen. Book review: Ancient Egyptian Science. Volume 1: Knowledge and Order by Marshall Clagett. Isis, 83(1):117, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234007.

Estes:1992:BRB

[820] J. Worth Estes. Book review: Dents et mâchoires dans les représentations religieuses et la pratique médicale de l’Égypte ancienne by Thierry Bardinet. Isis, 83(1):117–118, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234008.

Smith:1992:BRBa

[821] A. Mark Smith. Book review: Le regard, l’être et l’apparence dans l’optique de l’Antiquité by Gérard Simon. Isis, 83(1):118–119, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234009.

Scarborough:1992:BRB

[822] John Scarborough. Book review: Dioscurides Triumphans: Ein anonymer arabischer Kommentar (Ende 12. Jahr. n. Chr.) zur Materia medica by Albert Dietrich. Isis, 83(1):119–120, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234010.

Haq:1992:BRB

[823] S. Nomanul Haq. Book review: La philosophie des alchimistes et l’alchimie des philosophes: Jâbir ibn Hayyân et les “Frères de la Pureté” by Yves Marquet. Isis, 83(1):120–121, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234011.

Pereira:1992:BRB

[824] Michela Pereira. Book review: Le code alchimique dévoilé: Distillateurs, alchimistes et symbolistes by Suzanne Colnort-Bodet. Isis, 83(1):121– 122, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234012.

Hall:1992:BRBa

[825] Bert Hall. Book review: Un atelier de distillation du moyen âge. (Bulletin Archeologique du Comite des Travaux Historiques et Scientifiques, N.S., 22) by Isabelle Rouaze. Isis, 83(1):122–123, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234013.

Green:1992:BRBa

[826] Monica H. Green. Book review: Lovesickness in the Middle Ages: The “Viaticum” and Its Commentaries by Mary Frances Wack. Isis, 83(1): 123–124, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234014.

McPeak:1992:BRBa

[827] William J. McPeak. Book review: Bermannus (Le mineur): Un dialogue sur les mines by Georg Agricola; Robert Halleux; Albert Yans. Isis, 83(1):124–125, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234015.

Martin:1992:BRB

[828] John Martin. Book review: Venice and the Renaissance by Manfredo Tafuri; Jessica Levine. Isis, 83(1):125–126, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234016.

Ramirez-Castaneda:1992:BRB

[829] Santiago Ramirez-Castañeda. Book review: La revolución científica by J. M. López Piñero; V. Navarro; E. Portela. Isis, 83(1):126, March 1992.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234017.

Hatfield:1992:BRB

[830] Gary Hatfield. Book review: Die Philosophie des 17. Jahrhunderts. Volume 3: England by Jean–Pierre Schobinger. Isis, 83(1):126–128, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234018.

Laird:1992:BRB

[831] W. R. Laird. Book review: Sidereus nuncius, or the Sidereal Messenger by Galileo Galilei; Albert Van Helden; Two New Sciences: Including Centers of Gravity and Force of Percussion by Galileo Galilei; Stillman Drake. Isis, 83(1):128–130, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234019.

Lynch:1992:BRB

[832] William T. Lynch. Book review: History and Modernity in the Thought of Thomas Hobbes by Robert P. Kraynak. Isis, 83(1):130–131, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234020.

Dear:1992:BRBa

[833] Peter Dear. Book review: Portraits of Thought: Knowledge, Methods, and Styles in Pascal by Buford Norman. Isis, 83(1):131–132, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234021.

Grene:1992:BRB

[834] Marjorie Grene. Book review: Le principe de vie chez Descartes by Annie Bitbol–Hespériès. Isis, 83(1):132, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234022.

Cook:1992:BRBa

[835] Harold J. Cook. Book review: The Collected Letters of Antoni van Leeuwenhoek: The Complete Works of Van Leeuwenhoek, Issued and Annotated under the Auspices of the Leeuwenhoek-Commission of the Royal Netherlands Academy of Arts and Sciences. Volume XII: 1696–1699 by Antoni van Leeuwenhoek; L. C. Palm. Isis, 83(1):132–133, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234023.

Joy:1992:BRBa

[836] Lynn S. Joy. Book review: Leibniz and Arnauld: A Commentary on Their Correspondence by R. C. Sleigh. Isis, 83(1):133–134, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234024.

Harris:1992:BRBa

[837] Steven J. Harris. Book review: Eine wissenschaftliche Akademie für China: Briefe des Chinamissionars Joachim Bouvet S.J. an Gottfried Wilhelm Leibniz und Jean-Paul Bignon über die Erforschung der chinesischen Kultur, Sprache und Geschichte by Claudia von Collani. Isis, 83(1):134–135, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234025.

Cohen:1992:BRBa

[838] I. Bernard Cohen. Book review: The Preliminary Manuscripts for Isaac Newton’s 1687 Principia, 1684–1685 by Isaac Newton. Isis, 83(1):135– 136, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234026.

Yoder:1992:BRBa

[839] Joella G. Yoder. Book review: Standing on the Shoulders of Giants: A Longer View of Newton and Halley by Norman J. W. Thrower. Isis, 83(1):136–137, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234027.

Jacob:1992:BRB

[840] Margaret C. Jacob. Book review: Prophecy and Reason: The Dutch Collegiants in the Early Enlightenment by Andrew C. Fix. Isis, 83(1): 137–138, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234028.

Hankins:1992:BRB

[841] Thomas L. Hankins. Book review: The Quantifying Spirit in the Eighteenth Century by Tore Frängsmyr; J. L. Heilbron; Robin E. Rider. Isis, 83(1):138–140, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234029.

Terrall:1992:BRB

[842] Mary Terrall. Book review: Das Prinzip der kleinsten Wirkung und die Kraftkonzeptionen der rationalen Mechanik: Eine Untersuchung zur Grundlegungsproblematik bei Leonhard Euler, Pierre Louis Moreau de Maupertuis und Joseph Louis Lagrange by Helmut Pulte. Isis, 83(1): 140–141, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234030.

Genuth:1992:BRBa

[843] Sara Schechner Genuth. Book review: Le Citoyen Lenoir: Scientific Instrument Making in Revolutionary France by J. A. Bennett; From Pleasure and Profit to Science and Security: Etienne Lenoir and the Transformation of Precision Instrument-Making in France, 1760–1830 by A. J. Turner. Isis, 83(1):141–142, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234031.

vanKeuren:1992:BRB

[844] David K. van Keuren. Book review: Curators and Culture: The Museum Movement in America, 1740–1870 by Joel J. Orosz. Isis, 83(1):142– 143, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234032.

Norton:1992:BRB

[845] Robert E. Norton. Book review: Traditionen Humboldts by Jurgen Trabant. Isis, 83(1):143–144, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234033.

Burchfield:1992:BRB

[846] Joe D. Burchfield. Book review: Energy and Empire: A Biographical Study of Lord Kelvin by Crosbie Smith; M. Norton Wise. Isis, 83(1): 144–146, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234034.

Page:1992:BRB

[847] Leroy E. Page. Book review: Scientist of Empire: Sir Roderick Murchison, Scientific Exploration, and Victorian Imperialism by Robert A. Stafford. Isis, 83(1):146–147, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234035.

Boler:1992:BRB

[848] John Boler. Book review: Charles S. Peirce: An Intellectual Biography by Gerard Deledalle; Susan Petrilli. Isis, 83(1):147, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234036.

Hindle:1992:BRB

[849] Brooke Hindle. Book review: Samuel F. B. Morse by Paul J. Staiti. Isis, 83(1):147–148, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234037.

Rupke:1992:BRB

[850] Nicolaas Rupke. Book review: Catastrophism: Systems of Earth History by Richard Huggett. Isis, 83(1):148–149, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234038.

Buttner:1992:BRB

[851] Stefan Büttner. Book review: Die Wissenschaften vom Leben zwischen Physik und Metaphysik: Auf der Suche nach dem Newton der Biologie im 19. Jahrhundert by Hermann Schlüter. Isis, 83(1):149–150, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234039.

Camerini:1992:BRBa

[852] Jane R. Camerini. Book review: The Correspondence of Charles Darwin. Volume 5: 1851–1855 by Charles Darwin; Frederick Burkhardt; Sydney Smith; The Correspondence of Charles Darwin. Volume 6: 1856–1857 by Charles Darwin; Frederick Burkhardt; Sydney Smith. Isis, 83(1): 150–151, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234040.

Fantini:1992:BRBa

[853] Bernardino Fantini. Book review: Legge e caso nella genetica mendeliana by Federico Di Trocchio. Isis, 83(1):151, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234041.

Nye:1992:BRB

[854] Robert A. Nye. Book review: Darwin’s Influence on Freud: A Tale of Two Sciences by Lucille B. Ritvo. Isis, 83(1):151–152, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234042.

Dwyer:1992:BRB

[855] Ellen Dwyer. Book review: Social Order/Mental Disorder: Anglo– American Psychiatry in Historical Perspective by Andrew Scull. Isis, 83(1):152–153, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234043.

Blustein:1992:BRBa

[856] Bonnie Ellen Blustein. Book review: Murders and Madness: Medicine, Law, and Society in the Fin de Siècle by Ruth Harris. Isis, 83(1):153– 154, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234044.

Samuels:1992:BRB

[857] Warren J. Samuels. Book review: The History of Econometric Ideas by Mary S. Morgan. Isis, 83(1):154–155, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234045.

Biggs:1992:BRB

[858] Lindy Biggs. Book review: A Promising Field: Engineering at Alabama, 1837–1987 by Robert J. Norrell. Isis, 83(1):155–156, March 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234046.

Sigurdsson:1992:BRBa

[859] Skuli Sigurdsson. Book review: Moderne Sprache Mathematik: Eine Geschichte des Streits um die Grundlagen der Disziplin und des Subjekts formaler Systeme by Herbert Mehrtens. Isis, 83(1):156–158, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234047.

Goodstein:1992:BRB

[860] Judith R. Goodstein. Book review: Atomi metafori paradossi: Niels Bohr e la costruzione di una nuova fisica by Sandro Petruccioli. Isis, 83(1): 158, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234048.

Hiebert:1992:BRBa

[861] Erwin N. Hiebert. Book review: Abram Fedorovic Ioffe: Vater der sowjetischen Physik by Horst Kant. Isis, 83(1):158–159, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234049.

Pycior:1992:BRB

[862] Helena M. Pycior. Book review: Grand Obsession: Madame Curie and Her World by Rosalynd Pflaum. Isis, 83(1):159–160, March 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234050.

Hetherington:1992:BRBa

[863] Norriss S. Hetherington. Book review: Chandra: A Biography of S. Chandrasekhar by Kameshwar C. Wali. Isis, 83(1):160, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234051.

Kay:1992:BRB

[864] Lily E. Kay. Book review: Where the Truth Lies: Franz Moewus and the Origins of Molecular Biology by Jan Sapp. Isis, 83(1):160–161, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234052.

Thielman:1992:BRB

[865] Samuel B. Thielman. Book review: Moments of Unreason: The Practice of Canadian Psychiatry and the Homewood Retreat, 1883–1923 by Cheryl Krasnick Warsh. Isis, 83(1):161–162, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234053.

Coon:1992:BRB

[866] Deborah J. Coon. Book review: Constructing the Subject: Historical Origins of Psychological Research by Kurt Danziger. Isis, 83(1):162– 163, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234054.

Harris:1992:BRBb

[867] Benjamin Harris. Book review: Menninger: The Family and the Clinic by Lawrence J. Friedman. Isis, 83(1):163–164, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234055.

Alvarado:1992:BRBa

[868] Carlos S. Alvarado. Book review: La Casa de Salud Valdecilla: Orígen y antecedentes: La introducción del hospital contemporáneo en España by Fernando Salmón; Luis García Ballester; Jon Arrizabalaga. Isis, 83 (1):164–165, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234056.

Boyer:1992:BRB

[869] Paul Boyer. Book review: Visions of a Flying Machine: The Wright Brothers and the Process of Invention by Peter L. Jakab. Isis, 83(1): 165–166, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234057.

Borgmann:1992:BRB

[870] Albert Borgmann. Book review: Heidegger’s Confrontation with Modernity: Technology, Politics, and Art by Michael E. Zimmerman. Isis, 83(1):166–167, March 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234058.

Mahoney:1992:BRBa

[871] Michael S. Mahoney. Book review: Early Scientific Computing in Britain by Mary Croarken. Isis, 83(1):167–168, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234059.

Pisano:1992:BRB

[872] Dominick A. Pisano. Book review: The Electronic Sweatshop: How Computers Are Transforming the Office of the Future into the Factory of the Past by Barbara Garson; Science, Technology, and Social Progress by Steven L. Goldman; Technology and the Pursuit of Economic Growth by David C. Mowery; Nathan Rosenberg. Isis, 83(1):168–170, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234060.

Seidel:1992:BRBa

[873] Robert W. Seidel. Book review: German National Socialism and the Quest for Nuclear Power, 1939–1949 by Mark Walker. Isis, 83(1):170, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234061.

Capshew:1992:BRB

[874] James H. Capshew. Book review: Science and the Navy: The History of the Office of Naval Research by Harvey M. Sapolsky. Isis, 83(1):170– 171, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234062.

Nebeker:1992:BRBa

[875] Frederik Nebeker. Book review: Grossforschung mit kleinen Teilchen: Das Deutsche Elektronen-Synchrotron DESY 1956–1970 by Claus Habfast. Isis, 83(1):171–172, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234063.

Dupree:1992:BRB

[876] A. Hunter Dupree. Book review: Science and Academic Life in Transition by Emanuel Piore; Eli Ginzberg. Isis, 83(1):172–173, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234064.

Rhees:1992:BRB

[877] David J. Rhees. Book review: Making Science Our Own: Public Images of Science, 1910–1955 by Marcel C. LaFollette. Isis, 83(1):173–174, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234065.

Kuklick:1992:BRBa

[878] Henrika Kuklick. Book review: The Impossible Science: An Institutional Analysis of American Sociology by Stephen Park Turner; Jonathan H. Turner. Isis, 83(1):174–175, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234066.

Grunbaum:1992:BRB

[879] Adolf Grunbaum. Book review: Motivation and Explanation: An Essay on Freud’s Philosophy of Science by Nigel Mackay. Isis, 83(1):175–176, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234067.

Tang:1992:BRB

[880] Paul C. L. Tang. Book review: Laws and Symmetry by Bas C. van Fraassen. Isis, 83(1):176–177, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234068.

Franklin:1992:BRB

[881] Allan Franklin. Book review: Experiment and the Making of Meaning: Human Agency in Scientific Observation and Experiment by David Gooding. Isis, 83(1):177–178, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234069.

Berbrier:1992:BRB

[882] Mitchell Berbrier. Book review: The Reflexive Thesis: Wrighting Sociology of Scientific Knowledge by Malcolm Ashmore. Isis, 83(1):178–179, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234070.

Longino:1992:BRB

[883] Helen E. Longino. Book review: Who Knows: From Quine to a Feminist Empiricism by Lynn Hankinson Nelson. Isis, 83(1):179, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234071.

Gieryn:1992:BRB

[884] Thomas F. Gieryn. Book review: Social Studies of Science by Bernard Barber. Isis, 83(1):179–181, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234072.

Elliott:1992:BRBa

[885] Clark A. Elliott. Book review: Early American Scientific and Technical Literature: An Annotated Bibliography of Books, Pamphlets, and Broadsides by Margaret W. Batschelet. Isis, 83(1):181, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234073.

Manning:1992:BRBa

[886] Kenneth R. Manning. Book review: Blacks in Science and Medicine by Vivian Ovelton Sammons. Isis, 83(1):181–183, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234074.

Crowe:1992:BRBa

[887] Gregory D. Crowe. Book review: Nauchnye akademii stran zapadnoi evropy i severnoi ameriki by Iu. Kh. Kopelevich; E. P. Ozhigova. Isis, 83(1):183, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234075.

Rothenberg:1992:BRB

[888] Marc Rothenberg. Book review: Guide to Historical Resources in the Atmospheric Sciences: Archives, Manuscripts, and Special Collections in the Washington, D.C., Area by James R. Fleming. Isis, 83(1):183– 184, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234076.

Anonymous:1992:Ca

[889] Anonymous. Collections. Isis, 83(1):184–188, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234077.

Anonymous:1992:BMb

[890] Anonymous. Back matter. Isis, 83(1):189–194, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234078.

Anonymous:1992:FMb

[891] Anonymous. Front matter. Isis, 83(1):??, March 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233989.

Winkler:1992:RHG

[892] Mary G. Winkler and Albert Van Helden. Representing the heavens: Galileo and visual astronomy. Isis, 83(2):195–217, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234504.

Terrall:1992:RES

[893] Mary Terrall. Representing the Earth’s shape: The polemics surrounding Maupertuis’s expedition to Lapland. Isis, 83(2):218–237, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234505.

Wang:1992:SSC

[894] Jessica Wang. Science, security, and the Cold War: The case of E. U. Condon. Isis, 83(2):238–269, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234506.

Swerdlow:1992:ASP

[895] N. M. Swerdlow. Annals of scientific publishing: Johannes Petreius’s letter to rheticus. Isis, 83(2):270–274, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234507.

Moyer:1992:AMH

[896] Albert E. Moyer, Richard F. Hirsh, Michael M. Sokal, and Roger Hahn. Annual meeting of the History of Science Society, Madison, Wisconsin, 30 October–3 November 1991. Isis, 83(2):275–282, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234508.

Glick:1992:SDC

[897] Thomas F. Glick. Science, discovery, and the colonial world: Madrid, 25–28 June 1991. Isis, 83(2):282–286, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234509.

Schofield:1992:EHB

[898] Robert E. Schofield. Eloge: Harcourt Brown, 30 May 1900–17 November 1990. Isis, 83(2):286–287, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234510.

Hunter:1992:LE

[899] Andrew Hunter, H. L. L. Busard, P. Masani, and Theodore W. Pietsch. Letters to the Editor. Isis, 83(2):288–290, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234511.

Hofmann:1992:BRB

[900] James R. Hofmann. Book review: Naissance d’un nouveau pouvoir: Sciences et savants en France (1793–1824) by Nicole Dhombres; Jean Dhombres; Convolutions in French Mathematics, 1800–1840: From the Calculus and Mechanics to Mathematical Analysis and Mathematical Physics by Ivor Grattan-Guinness. Isis, 83(2):291–297, June 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234512.

Unguru:1992:BRB

[901] Sabetai Unguru. Book review: Numbers through the Ages by Graham Flegg. Isis, 83(2):298–299, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234513.

Bartholomew:1992:BRB

[902] James R. Bartholomew. Book review: Kagaku no shakaishi: Runessansu kara 20 seiki made. (A Social History of Science: From the Renaissance to the 20th Century) by Yasu Furukawa. Isis, 83(2):299–300, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234514.

Lindberg:1992:BRB

[903] David C. Lindberg. Book review: The Science of Art: Optical Themes in Western Art from Brunelleschi to Seurat by Martin Kemp. Isis, 83(2): 300–301, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234515.

Lankford:1992:BRB

[904] John Lankford. Book review: Planets and Perception: Telescopic Views and Interpretations, 1609–1909 by William Sheehan. Isis, 83(2):301– 302, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234516.

Gingerich:1992:BRBa

[905] Owen Gingerich. Book review: Astronomy and Optics from Pliny to Descartes: Texts, Diagrams, and Conceptual Structures by Bruce S. Eastwood; The Arabs and the Stars: Texts and Traditions on the Fixed Stars, and Their Influence on Medieval Europe by Paul Kunitzsch; Stars, Minds, and Fate: Essays in Ancient and Medieval Cosmology by J. D. North; The Universal Frame: Historical Essays in Astronomy, Natural Philosophy, and Scientific Method by J. D. North; Astronomy from Kepler to Newton: Historical Studies by Curtis Wilson. Isis, 83(2):302–303, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234517.

Brooks:1992:BRB

[906] Randall C. Brooks. Book review: Church, State, and Astronomy in Ireland: Two Hundred Years of Armagh Observatory by J. A. Bennett. Isis, 83(2):303–304, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234518.

Kuklick:1992:BRBb

[907] Henrika Kuklick. Book review: Narratives of Human Evolution by Misia Landau. Isis, 83(2):304, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234519.

Mcgrath:1992:BRB

[908] Sylvia W. Mcgrath. Book review: The Idea of Wilderness: From Prehistory to the Age of Ecology by Max Oelschlaeger. Isis, 83(2):304–305, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234520.

Davis:1992:BRBa

[909] Edward B. Davis. Book review: Creation and the History of Science by Christopher B. Kaiser. Isis, 83(2):305–306, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234521.

Crowe:1992:BRBb

[910] Gregory D. Crowe. Book review: Istoriia vychislitel’ noi tekhniki: Ot prosteishikh schetnykh prisposoblenii do slozhnykh releynykh sistem. (A History of Computer Technology: From the Simplest Counting Devices to Complex Relay Systems) by I. A. Apokin; L. E. Maistrov. Isis, 83(2): 306–307, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234522.

Cowan:1992:BRB

[911] Ruth Schwartz Cowan. Book review: Technology in World Civilization: A Thousand-Year History by Arnold Pacey. Isis, 83(2):307–308, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234523.

Manning:1992:BRBb

[912] Kenneth R. Manning. Book review: The Real McCoy: African-American Invention and Innovation, 1619–1930 by Portia P. James. Isis, 83(2): 308–309, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234524.

Clark:1992:BRB

[913] Charles M. A. Clark. Book review: The Invisible Hand: Economic Equilibrium in the History of Science by Bruna Ingrao; Giorgio Israel; Ian McGilvray. Isis, 83(2):309–310, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234525.

Smith:1992:BRBb

[914] A. Mark Smith. Book review: L’Optique de Claude Ptolémée dans la version latine d’après l’arabe de l’émir Eugène de Sicile: Édition critique et exégétique augmentée d’une traduction française et de compléments by Claudius Ptolemy; Albert Lejeune. Isis, 83(2):310–311, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234526.

Grant:1992:BRB

[915] Edward Grant. Book review: Place, Void, and Eternity. Philoponus: “Corollaries on Place and Void.” by David Furley; Simplicius: “Against Philoponus on the Eternity of the World.” by Christian Wildberg. Isis, 83(2):311–312, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234527.

Riddle:1992:BRBa

[916] John M. Riddle. Book review: Index de la pharmacopée du ier au Xe siècle by Carmélia Opsomer. Isis, 83(2):312, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234528.

Hall:1992:BRBb

[917] Bert Hall. Book review: Paradise Restored: The Mechanical Arts from Antiquity through the Thirteenth Century by Elspeth Whitney. Isis, 83 (2):312–313, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234529.

Ferngren:1992:BRB

[918] Gary B. Ferngren. Book review: The Human Embryo: Aristotle and the Arabic and European Traditions by G. R. Dunstan. Isis, 83(2):313– 314, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234530.

Lemay:1992:BRB

[919] Richard Lemay. Book review: Il “Lucidator dubitabilium astronomiae” di Pietro d’Abano: Opere scientifiche inedite by Graziella Federici Vescovini. Isis, 83(2):314–316, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234531.

Langermann:1992:BRB

[920] Y. Tzvi Langermann. Book review: Levi ben Gerson’s Prognostication for the Conjunction of 1345 by Bernard R. Goldstein; David Pingree. Isis, 83(2):316–317, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234532.

Rosenthal:1992:BRB

[921] Earl E. Rosenthal. Book review: Ingeniería y arquitectura en el Renacimiento español by Nicolás García Tapia. Isis, 83(2):317, June 1992.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234533.

Park:1992:BRBa

[922] Katharine Park. Book review: Prophecy and People in Renaissance Italy by Ottavia Niccoli; Lydia G. Cochrane. Isis, 83(2):318, June 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234534.

Joy:1992:BRBb

[923] Lynn S. Joy. Book review: Defenders of the Text: The Traditions of Scholarship in an Age of Science, 1450–1800 by Anthony Grafton. Isis, 83(2):319, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234535.

Jardine:1992:BRB

[924] Lisa Jardine. Book review: Francis Bacon and the Rhetoric of Nature by John C. Briggs. Isis, 83(2):319–320, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234536.

Jesseph:1992:BRB

[925] Douglas M. Jesseph. Book review: Hobbes oggi by Andrea Napoli; Thomas Hobbes: Philosophie premiere, theorie de la science et politique by Yves Charles Zarka; Jean Bernhardt. Isis, 83(2):320–322, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234537.

Feingold:1992:BRBa

[926] Mordechai Feingold. Book review: Henry More: Magic, Religion, and Experiment by A. Rupert Hall. Isis, 83(2):322–323, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234538.

Cook:1992:BRBb

[927] Harold J. Cook. Book review: A Company of Scientists: Botany, Patronage, and Community at the Seventeenth-Century Parisian Royal Academy of Sciences by Alice Stroup. Isis, 83(2):323–324, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234539.

Cohen:1992:BRBb

[928] H. Floris Cohen. Book review: Puritanism and the Rise of Modern Science: The Merton Thesis by I. Bernard Cohen. Isis, 83(2):324–325, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234540.

Yoder:1992:BRBb

[929] Joella Yoder. Book review: Ausgewählte Schriften by Joseph Ehrenfried Hofmann; Christoph J. Scriba. Isis, 83(2):325–326, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234541.

Vucinich:1992:BRB

[930] Alexander Vucinich. Book review: Ruder Bosković: Vizionar u prijelomima filozofije, znanosti i drustva by Ivan Supek. Isis, 83(2):326– 327, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234542.

Major:1992:BRB

[931] John S. Major. Book review: Les méthodes rapides pour la trigonométrie et le rapport précis du cercle (1774): Tradition chinoise et apport occidental en mathématiques by Catherine Jami. Isis, 83(2):327–328, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234543.

Richards:1992:BRB

[932] Joan Richards. Book review: The Development of Newtonian Calculus in Britain, 1700–1800 by Niccolo Guicciardini. Isis, 83(2):328–329, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234544.

Wood:1992:BRB

[933] Paul Wood. Book review: Diderot’s Dream by Wilda Anderson. Isis, 83 (2):329, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234545.

Finn:1992:BRB

[934] Bernard S. Finn. Book review: Benjamin Franklin’s Science by I. Bernard Cohen. Isis, 83(2):329–330, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234546.

Hahn:1992:BRB

[935] Roger Hahn. Book review: Monuments of Egypt: The Napoleonic Edition: The Complete Archaeological Plates from “La Description de l’Egypte” by Charles Coulston Gillispie; Michel Dewachter; L’expedition d’Egypte, 1798–1801 by Henry Laurens; Charles C. Gillispie; Jean-Claude Golvin; Claude Traunecker. Isis, 83(2):330–332, June 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234547.

Rusnock:1992:BRB

[936] Andrea Rusnock. Book review: Samuel Johnson in the Medical World: The Doctor and the Patient by John Wiltshire. Isis, 83(2):332–333, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234548.

Turner:1992:BRBa

[937] R. Steven Turner. Book review: The Natural and the Normative: Theories of Spatial Perception from Kant to Helmholtz by Gary Hatfield. Isis, 83(2):333, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234549.

Lenoir:1992:BRB

[938] Timothy Lenoir. Book review: Letters of Hermann von Helmholtz to His Wife, 1847–1859 by Richard L. Kremer. Isis, 83(2):333–335, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234550.

Boyd:1992:BRB

[939] Byron Boyd. Book review: Letters to His Parents, 1839–1864 by Rudolf Virchow; Marie Rabl; L. J. Rather. Isis, 83(2):335, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234551.

Churchill:1992:BRB

[940] Frederick B. Churchill. Book review: Politisch-biologische Analogien im Fruhwerk Rudolf Virchows by Renato G. Mazzolini; Klaus-Peter Tieck. Isis, 83(2):335–337, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234552.

Greene:1992:BRB

[941] Mott T. Greene. Book review: The Highlands Controversy: Constructing Geological Knowledge through Fieldwork in Nineteenth-Century Britain by David R. Oldroyd. Isis, 83(2):337–338, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234553.

Nebeker:1992:BRBb

[942] Frederik Nebeker. Book review: Meteorology in America, 1800–1870 by James Rodger Fleming. Isis, 83(2):338, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234554.

Parisi:1992:BRB

[943] Thomas Parisi. Book review: The Letters of Sigmund Freud to Eduard Silberstein, 1871–1881 by Walter Boehlich; Arnold J. Pomerans. Isis, 83(2):338–339, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234555.

Rocke:1992:BRB

[944] Alan J. Rocke. Book review: The Kaiser’s Chemists: Science and Modernization in Imperial Germany by Jeffrey Allan Johnson. Isis, 83(2): 339–340, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234556.

Stokes:1992:BRB

[945] Raymond G. Stokes. Book review: New Profession, Old Order: Engineers and German Society, 1815–1914 by Kees Gispen. Isis, 83(2):340– 341, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234557.

Verbrugge:1992:BRB

[946] Martha H. Verbrugge. Book review: The Eternally Wounded Woman: Women, Doctors, and Exercise in the Late Nineteenth Century by Patricia Vertinsky. Isis, 83(2):341–342, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234558.

Evans:1992:BRB

[947] H. Hughes Evans. Book review: Save the Babies: American Public Health Reform and the Prevention of Infant Mortality, 1850–1929 by Richard A. Meckel. Isis, 83(2):342, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234559.

Elliott:1992:BRBb

[948] J. Eric Elliott. Book review: Francis Preston Venable of the University of North Carolina by Maurice Bursey. Isis, 83(2):342–343, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234560.

Wasserman:1992:BRB

[949] Neil Wasserman. Book review: Redirecting Science: Niels Bohr, Philanthropy, and the Rise of Nuclear Physics by Finn Aaserud. Isis, 83(2): 343–344, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234561.

McCormmach:1992:BRB

[950] Russell McCormmach. Book review: Schrödinger: Life and Thought by Walter Moore. Isis, 83(2):344–345, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234562.

Coben:1992:BRB

[951] Stanley Coben. Book review: Partners in Science: Foundations and Natural Scientists, 1900–1945 by Robert E. Kohler. Isis, 83(2):345– 346, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234563.

Larson:1992:BRB

[952] Edward J. Larson. Book review: In Search of Human Nature: The Decline and Revival of Darwinism in American Social Thought by Carl N. Degler. Isis, 83(2):346–347, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234564.

Stocking:1992:BRB

[953] George W. Stocking, Jr. Book review: Piltdown: A Scientific Forgery by Frank Spencer; Ian Langham; The Piltdown Papers, 1908–1955: The Correspondence and Other Documents Relating to the Piltdown Forgery by Frank Spencer. Isis, 83(2):347–349, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234565.

VanRiper:1992:BRBa

[954] A. Bowdoin Van Riper. Book review: Fox at the Wood’s Edge: A Biography of Loren Eiseley by Gale E. Christianson; Loren Eiseley: A Modern Ishmael by Peter Heidtmann. Isis, 83(2):349–350, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234566.

Smith:1992:BRBc

[955] Roger Smith. Book review: Simians, Cyborgs, and Women: The Reinvention of Nature by Donna J. Haraway. Isis, 83(2):350–351, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234567.

Mazumdar:1992:BRBa

[956] Pauline M. H. Mazumdar. Book review: Demography and Degeneration: Eugenics and the Declining Birth Rate in Twentieth–Century Britain by Richard A. Soloway. Isis, 83(2):351–352, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234568.

Proctor:1992:BRB

[957] Robert N. Proctor. Book review: The Wellborn Science: Eugenics in Germany, France, Brazil, and Russia by Mark B. Adams. Isis, 83(2): 352–353, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234569.

Fantini:1992:BRBb

[958] Bernardino Fantini. Book review: The Emergence of Bacterial Genetics by Thomas D. Brock. Isis, 83(2):353–354, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234570.

Crow:1992:BRB

[959] James F. Crow. Book review: Fifty Years of Genetic Load: An Odyssey by Bruce Wallace. Isis, 83(2):354–355, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234571.

Bocking:1992:BRB

[960] Stephen Bocking. Book review: Elton’s Ecologists: A History of the Bureau of Animal Population by Peter Crowcroft. Isis, 83(2):355–356, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234572.

Burstyn:1992:BRB

[961] Harold L. Burstyn. Book review: Deep-Sea Challenge: The John Murray/Mabahiss Expedition to the Indian Ocean, 1933–34 by A. L. Rice. Isis, 83(2):356–357, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234573.

Kudlick:1992:BRB

[962] Catherine J. Kudlick. Book review: Egypt’s Other Wars: Epidemics and the Politics of Public Health by Nancy Elizabeth Gallagher. Isis, 83(2): 357–358, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234574.

Rogers:1992:BRB

[963] Naomi Rogers. Book review: Rocky Mountain Spotted Fever: History of a Twentieth–Century Disease by Victoria A. Harden. Isis, 83(2):358– 359, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234575.

Randjbar-Daemi:1992:BRB

[964] S. Randjbar-Daemi. Book review: Journey into Light: Life and Science of C. V. Raman by G. Venkataraman. Isis, 83(2):359, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234576.

Schweber:1992:BRBa

[965] Samuel S. Schweber. Book review: Rochester Roundabout: The Story of High Energy Physics by John Polkinghorne. Isis, 83(2):359–361, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234577.

Seidel:1992:BRBb

[966] Robert W. Seidel. Book review: The Manhattan Project: A Documentary Introduction to the Atomic Age by Michael B. Stoff; Jonathan F. Fanton; R. Hal Williams. Isis, 83(2):361, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234578.

Feld:1992:BRB

[967] Bernard T. Feld. Book review: Atomic Rivals: A Candid Memoir of Rivalries among the Allies over the Bomb by Bertrand Goldschmidt; George M. Temmer. Isis, 83(2):361–362, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234579.

Hacker:1992:BRB

[968] Barton C. Hacker. Book review: Ethical Issues Associated with Scientific and Technological Research for the Military by Carl Mitcham; Philip Siekevitz; Atomic Energy for Military Purposes: The Official Report on the Development of the Atomic Bomb under the Auspices of the United States Government, 1940–1945 by Henry DeWolf Smyth; Preventing a Biological Arms Race by Susan Wright. Isis, 83(2):362–364, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234580.

Hevly:1992:BRB

[969] Bruce W. Hevly. Book review: Viewing the Earth: The Social Construction of the Landsat Satellite System by Pamela E. Mack. Isis, 83(2): 364–365, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234581.

Creed:1992:BRBb

[970] Walter Creed. Book review: Actors and Onlookers: Theater and Twentieth–Century Scientific Views of Nature by Natalie Crohn Schmitt. Isis, 83(2):365–366, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234582.

Alborn:1992:BRB

[971] Timothy L. Alborn. Book review: The Taming of Chance by Ian Hacking. Isis, 83(2):366–367, June 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234583.

Feyerabend:1992:BRB

[972] Paul Feyerabend. Book review: Science and Relativism: Some Key Controversies in the Philosophy of Science by Larry Laudan. Isis, 83(2): 367–368, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/234584.

Finocchiaro:1992:BRB

[973] Maurice A. Finocchiaro. Book review: I modelli, l’invenzione e la conferma: Saggio su Keplero, la rivoluzione copernicana e la “New philosophy of science.” by Angelo Maria Petroni. Isis, 83(2):368–369, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234585.

Melhado:1992:BRB

[974] Evan M. Melhado. Book review: The Pasteurization of France by Bruno Latour; Alan Sheridan; John Law. Isis, 83(2):369–371, June 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234586.

Golinski:1992:BRBa

[975] Jan V. Golinski. Book review: Artificial Experts: Social Knowledge and Intelligent Machines by H. M. Collins. Isis, 83(2):371–372, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234587.

Hodge:1992:BRB

[976] M. J. S. Hodge. Book review: Cognitive Foundations of Natural History: Towards an Anthropology of Science by Scott Atran. Isis, 83(2):372– 373, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234588.

Fuller:1992:BRBa

[977] Steve Fuller. Book review: Essays on the Theory of Scientific Cognition by Jerzy Kmita; Jacek Holówka. Isis, 83(2):373, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234589.

Creath:1992:BRB

[978] Richard Creath. Book review: Creative Understanding: Philosophical Reflections on Physics by Roberto Toretti. Isis, 83(2):373–374, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234590.

Turner:1992:BRBb

[979] Stephen P. Turner. Book review: Causality in Sociological Research by Jakub Karpiński. Isis, 83(2):374–375, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234591.

Kuriyama:1992:BRB

[980] Shigehisa Kuriyama. Book review: Science and Comparative Philosophy by David Edward Shaner; Shigenori Nagatomo; Yasuo Yuasa. Isis, 83(2): 375–376, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234592.

Charvolin:1992:BRB

[981] Florian Charvolin. Book review: La presqu’île au nucléaire by Françoise Zonabend. Isis, 83(2):376, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234593.

Anonymous:1992:Cb

[982] Anonymous. Collections. Isis, 83(2):376–379, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234594.

Anonymous:1992:BMc

[983] Anonymous. Back matter. Isis, 83(2):380–386, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234595.

Anonymous:1992:FMc

[984] Anonymous. Front matter. Isis, 83(2):??, June 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234503.

Copenhaver:1992:DSR

[985] Brian P. Copenhaver. Did science have a renaissance? Isis, 83(3):387– 407, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233902.

Livingstone:1992:DCB

[986] David N. Livingstone. Darwinism and Calvinism: The Belfast–Princeton connection. Isis, 83(3):408–428, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233903.

Goldberg:1992:ICO

[987] Stanley Goldberg. Inventing a climate of opinion: Vannevar Bush and the decision to build the bomb. Isis, 83(3):429–452, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233904.

Gundert:1992:PTR

[988] Beate Gundert. Parts and their roles in Hippocratic medicine. Isis, 83 (3):453–465, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233905.

Mahoney:1992:BRBb

[989] Michael S. Mahoney. Book review: Mathematik im Abendland: Von den romischen Feldmessern bis zu Descartes by Helmuth Gericke. Isis, 83(3):466, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233906.

Doel:1992:BRB

[990] Ronald E. Doel. Book review: Ancient Light: Our Changing View of the Universe by Alan Lightman. Isis, 83(3):466–467, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233907.

Genuth:1992:BRBb

[991] Sara Schechner Genuth. Book review: Comets: A Chronological History of Observation, Science, Myth, and Folklore by Donald K. Yeomans. Isis, 83(3):468–469, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233908.

Davis:1992:BRBb

[992] Edward B. Davis. Book review: Science and Religion: Some Historical Perspectives by John Hedley Brooke. Isis, 83(3):469–470, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233909.

Crosby:1992:BRB

[993] Alfred W. Crosby. Book review: Aztec Medicine, Health, and Nutrition by Bernard R. Ortiz de Montellano. Isis, 83(3):470, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233910.

Hiebert:1992:BRBb

[994] Erwin N. Hiebert. Book review: De bevordering en volmaking der proefondervindelijke wijsbegeerte: De rol van het Bataafsch Genootschap te Rotterdam in de geschiendenis van de natuurwetenschappen, geneeskunde en techniek (1769–1988) by M. J. van Lieburg; H. A. M. Snelders. Isis, 83 (3):470–471, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233911.

Inkster:1992:BRB

[995] Ian Inkster. Book review: The Japanese and Western Science by Masao Watanabe; Otto Theodor Benfey. Isis, 83(3):471–472, September 1992.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233912.

Johnston:1992:BRB

[996] William D. Johnston. Book review: Red-Hair Medicine: Dutch-Japanese Medical Relations by H. Beukers; A. M. Luyendijk-Elshout; M. E. van Opstall; F. Vos. Isis, 83(3):472–473, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233913.

Benson:1992:BRB

[997] Keith R. Benson. Book review: The Vital Machine: A Study of Technology and Organic Life by David F. Channell. Isis, 83(3):473–474, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233914.

Moesgaard:1992:BRB

[998] Kristian Peder Moesgaard. Book review: Neue Untersuchungen zur Topographie des babylonischen Fixsternhimmels by Johannes Koch. Isis, 83 (3):474–475, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233915.

vonStaden:1992:BRBa

[999] Heinrich von Staden. Book review: Demystifying Mentalities by G. E. R. Lloyd. Isis, 83(3):475–476, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233916.

DeFilippo:1992:BRB

[1000] Joseph G. DeFilippo. Book review: God and Greek Philosophy: Studies in the Early History of Natural Theology by L. P. Gerson. Isis, 83(3): 476–477, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233917.

Ebbesen:1992:BRBa

[1001] Sten Ebbesen. Book review: On Aristotle’s Prior Analytics 1.7-7 by Alexander of Aphrodisias; Jonathan Barnes; Susanne Bobzien; Kevin Flannery; Katerina Ieradiakonou. Isis, 83(3):477–478, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233918.

Potter:1992:BRB

[1002] Paul Potter. Book review: Sull’ottima maniera d’insegnare; Esortazione alla medicina by Adelmo Barigazzi; Galen of Pergamum. Isis, 83(3):478, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233919.

Hall:1992:BRBc

[1003] Bert S. Hall. Book review: On Roman Time: The Codex-Calendar of 354 and the Rhythms of Urban Life in Late Antiquity by Michele Renee Salzman. Isis, 83(3):478–479, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233920.

Glick:1992:BRB

[1004] Thomas F. Glick. Book review: L’eau conquise: Les origines et le monde antique by Marcel Nordon. Isis, 83(3):479–480, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233921.

Hill:1992:BRB

[1005] Donald R. Hill. Book review: The Book “Kitab al-Hiyal” of Banu Musa bin Shakir: Interpreted in Sense of Modern System and Control Engineering by Atilla Bir; Ekmeleddin Ihsanoglu. Isis, 83(3):480, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233922.

Yaqub:1992:BRB

[1006] Aladdin M. Yaqub. Book review: Aristotelian Logic and the Arabic Language in Alfarabi by Shukri B. Abed. Isis, 83(3):480–481, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233923.

Dales:1992:BRB

[1007] Richard C. Dales. Book review: The Eternity of the World: In the Thought of Thomas Aquinas and His Contemporaries by J. B. M. Wissink. Isis, 83(3):481, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233924.

Ariew:1992:BRB

[1008] Roger Ariew. Book review: The Origins of Statics: The Sources of Physical Theory by Pierre Duhem; Grant F. Leneaux; Victor N. Vagliente; Guy H. Wagener. Isis, 83(3):482, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233925.

Sobol:1992:BRBa

[1009] Peter G. Sobol. Book review: Chaucer’s Knight’s Tale and Theories of Scholastic Psychology by Lois Roney. Isis, 83(3):482–483, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233926.

Park:1992:BRBb

[1010] Katharine Park. Book review: Medieval Psychology by Simon Kemp. Isis, 83(3):483–484, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233927.

Jones:1992:BRB

[1011] Peter Murray Jones. Book review: Galen’s Method of Healing: Proceedings of the 1982 Galen Symposium by Fridolf Kudlien; Richard J. Durling. Isis, 83(3):484–485, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233928.

Green:1992:BRBb

[1012] Monica H. Green. Book review: Healing and Society in Medieval England: A Middle English Translation of the Pharmaceutical Writings of Gilbertus Anglicus by Faye Marie Getz. Isis, 83(3):485–486, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233929.

Grafton:1992:BRBa

[1013] Anthony Grafton. Book review: Trigonometrisch-astronomisches Rechnen kurz vor Copernicus: Der Briefwechsel Regiomontanus-Bianchini by Armin Gerl. Isis, 83(3):486–487, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233930.

Rosenfeld:1992:BRB

[1014] Boris A. Rosenfeld. Book review: Nenuzhnoe dlya neuchey [Useless for Ignorants] by Amirdovlat Amasiaci; Stella A. Vardanian. Isis, 83(3): 487–488, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233931.

VanHelden:1992:BRB

[1015] Albert Van Helden. Book review: On the Distances between Sun, Moon, and Earth according to Ptolemy, Copernicus, and Reinhold by Janice Adrienne Henderson. Isis, 83(3):488, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233932.

Shea:1992:BRBb

[1016] William R. Shea. Book review: Galileo, Bellarmine, and the Bible by Richard J. Blackwell. Isis, 83(3):488–489, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233933.

Biagioli:1992:BRB

[1017] Mario Biagioli. Book review: The Galileo Affair: A Documentary History by Maurice A. Finocchiaro. Isis, 83(3):490, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233934.

Blackwell:1992:BRB

[1018] Richard J. Blackwell. Book review: In the Wake of Galileo by Michael Segre. Isis, 83(3):490–491, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233935.

North:1992:BRB

[1019] J. D. North. Book review: The Pulse of Time: Galileo Galilei, the Determination of Longitude, and the Pendulum Clock by Silvio A. Bedini. Isis, 83(3):491–492, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233936.

Ranea:1992:BRB

[1020] Alberto Guillermo Ranea. Book review: On True and False Ideas by A. Arnauld; Stephen Gaukroger. Isis, 83(3):492–493, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233937.

Guerrini:1992:BRBa

[1021] Anita Guerrini. Book review: The Early Essays and Ethics of Robert Boyle by Robert Boyle; John T. Harwood. Isis, 83(3):493–494, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233938.

Shackelford:1992:BRB

[1022] Jole Shackelford. Book review: Essential Readings by Paracelsus; Nicholas Goodrick-Clarke. Isis, 83(3):494, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233939.

Miles:1992:BRB

[1023] Sara Joan Miles. Book review: La première révolution biologique: Réflexions sur la physiologie et la médecine du XVIIe siècle by Mirko D. Grmek. Isis, 83(3):495, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233940.

Alvarado:1992:BRBb

[1024] Carlos S. Alvarado. Book review: Los orígenes en España de los estudios sobre la salud pública by José M. López Piñero. Isis, 83(3):495–496, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233941.

McGeachie:1992:BRB

[1025] James McGeachie. Book review: The Idea of Progress in Eighteenth-Century Britain by David Spadafora. Isis, 83(3):496–497, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233942.

Taylor:1992:BRB

[1026] Kenneth L. Taylor. Book review: Histoire des sciences de la terre entre 1790 et 1815 vue à travers les documents inédits de la Société de Physique et d’Histoire Naturelle de Genève: Trois grands protagonistes: Marc-Auguste Pictet, Guillaume-Antoine Deluc et Jean Tollot by Albert V. Carozzi; Les sciences de la terre aux XVIIe et XVIIIe siècles: Naissance de la géologie by Gabriel Gohau. Isis, 83(3):497–499, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233943.

Hannaway:1992:BRB

[1027] Caroline Hannaway. Book review: In Sickness and in Health: The British Experience, 1650–1850 by Roy Porter; Dorothy Porter. Isis, 83(3):499– 500, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233944.

Broman:1992:BRBa

[1028] Thomas Broman. Book review: The Woman Beneath the Skin: A Doctor’s Patients in Eighteenth-Century Germany by Barbara Duden; Thomas Dunlap. Isis, 83(3):500–501, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233945.

Scriba:1992:BRB

[1029] Christoph J. Scriba. Book review: Der “Fürst der Mathematiker” in Briefen und Gesprächen by Carl Friedrich Gauss; Kurt-R. Biermann. Isis, 83(3):501, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233946.

Mirowski:1992:BRB

[1030] Philip Mirowski. Book review: A World Ruled by Number: William Stanley Jevons and the Rise of Mathematical Economics by Margaret Schabas. Isis, 83(3):501–502, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233947.

Turner:1992:BRBc

[1031] R. Steven Turner. Book review: Contrasts in Scientific Style: Research Groups in the Chemical and Biochemical Sciences by Joseph S. Fruton. Isis, 83(3):503, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233948.

Maulitz:1992:BRB

[1032] Russell C. Maulitz. Book review: Bichat: La vie fulgurante d’un génie by Nicolas Dobo; André Role. Isis, 83(3):503–504, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233949.

Mazumdar:1992:BRBb

[1033] Pauline M. H. Mazumdar. Book review: Metchnikoff and the Origins of Immunology: From Metaphor to Theory by Alfred I. Tauber; Leon Chernyak. Isis, 83(3):504–505, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233950.

Gladstone:1992:BRB

[1034] Jo Gladstone. Book review: Death by Migration: Europe’s Encounter with the Tropical World in the Nineteenth Century by Philip D. Curtin. Isis, 83(3):505–506, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233951.

LaBerge:1992:BRB

[1035] Ann La Berge. Book review: The History of Yellow Fever: An Essay on the Birth of Tropical Medicine by François Delaporte; Arthur Goldhammer. Isis, 83(3):506–507, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233952.

Blustein:1992:BRBb

[1036] Bonnie Ellen Blustein. Book review: “Shattered Nerves”: Doctors, Patients, and Depression in Victorian England by Janet Oppenheim. Isis, 83(3):507–508, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233953.

Noll:1992:BRB

[1037] Steven Noll. Book review: Women and Madness: The Incarceration of Women in Nineteenth-Century France by Yannick Ripa; Catherine du Peloux Menagé. Isis, 83(3):508–509, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/233954.

Jeremy:1992:BRB

[1038] David J. Jeremy. Book review: Ingenious Yankees: The Rise of the American System of Manufactures in the Private Sector by Donald R. Hoke. Isis, 83(3):509–511, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233955.

Sinclair:1992:BRB

[1039] Bruce Sinclair. Book review: John B. Jervis: An American Engineering Pioneer by F. Daniel Larkin. Isis, 83(3):511, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233956.

Reynolds:1992:BRB

[1040] Terry S. Reynolds. Book review: Ensayo sobre la composición de las máquinas by José María de Lanz; Augustín de Betancourt; Manuel Díaz– Marta. Isis, 83(3):511–512, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233957.

VanRiper:1992:BRBb

[1041] A. Bowdoin Van Riper. Book review: Pitt Rivers: The Life and Archaeological Work of Lieutenant-General Augustus Henry Lane Fox Pitt Rivers, DCL, FRS, FSA by Mark Bowden. Isis, 83(3):512–513, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/233958.

Lewis:1992:BRB

[1042] Albert C. Lewis. Book review: Corps et modèles: Essai sur l’historie de l’algèbre réelle by Hourya Sinaceur. Isis, 83(3):513–514, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233959.

Archibald:1992:BRB

[1043] Thomas Archibald. Book review: Louis Néel, le magnétisme et Grenoble: Récit de la création d’un empire physicien dans la province française, 1940–1965 by Dominique Pestre. Isis, 83(3):514–515, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233960.

Hetherington:1992:BRBb

[1044] Norriss S. Hetherington. Book review: Clyde Tombaugh: Discoverer of Planet Pluto by David H. Levy. Isis, 83(3):515, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233961.

Degen:1992:BRB

[1045] Peter A. Degen. Book review: Science and Cultural Crisis: An Intellectual Biography of Percy Williams Bridgman (1882–1961) by Maila L. Walter. Isis, 83(3):515–516, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233962.

Westfall:1992:BRB

[1046] Catherine Westfall. Book review: History of CERN. Volume 2: Building and Running the Laboratory, 1954–1965 by Armin Hermann; John Krige; Ulrike Mersits; Dominique Pestre; Laura Weiss. Isis, 83(3):516– 517, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/233963.

Thackray:1992:BRB

[1047] Arnold Thackray. Book review: Chemistry, as Viewed from Bascom’s Hill: A History of the Chemistry Department at the University of Wisconsin in Madison by Aaron J. Ihde. Isis, 83(3):517–518, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233964.

Brush:1992:BRB

[1048] Stephen G. Brush. Book review: The Age of the Earth by G. Brent Dalrymple. Isis, 83(3):518, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233965.

Dewsbury:1992:BRB

[1049] Donald A. Dewsbury. Book review: Vygotsky’s Psychology: A Biography of Ideas by Alex Kozulin. Isis, 83(3):518–519, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233966.

Lederer:1992:BRB

[1050] Susan E. Lederer. Book review: Strangers at the Bedside: A History of How Law and Bioethics Transformed Medical Decision Making by David J. Rothman. Isis, 83(3):519–520, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233967.

Downing:1992:BRB

[1051] Kevin J. Downing. Book review: The Cybernetics Group by Steve Joshua Heims. Isis, 83(3):520–521, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233968.

Shapiro:1992:BRB

[1052] Henry D. Shapiro. Book review: To Foster the Spirit of Professionalism: Southern Scientists and State Academies of Science by Nancy Smith Midgette. Isis, 83(3):521, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233969.

Gillingham:1992:BRB

[1053] John Gillingham. Book review: Divide and Prosper: The Heirs of I. G. Farben under Allied Authority, 1945–1951 by Raymond G. Stokes. Isis, 83(3):521–522, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233970.

Paul:1992:BRB

[1054] H. W. Paul. Book review: Jacques Lacan & Co.: A History of Psychoanalysis in France, 1925–1985 by Elisabeth Roudinesco; Jeffrey Mehlman. Isis, 83(3):522–523, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/233971.

Bromberg:1992:BRB

[1055] Joan Lisa Bromberg. Book review: Inventing Accuracy: A Historical Sociology of Nuclear Missile Guidance by Donald MacKenzie. Isis, 83 (3):523–524, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233972.

Hunt:1992:BRB

[1056] Elizabeth Hunt. Book review: Scientific Knowledge in Controversy: The Social Dynamics of the Fluoridation Debate by Brian Martin. Isis, 83 (3):524–525, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233973.

Shapin:1992:BRB

[1057] Steven Shapin. Book review: Scientific Growth: Essays on the Social Organization and Ethos of Science by Joseph Ben-David; Gad Freudenthal. Isis, 83(3):525–526, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233974.

Hacking:1992:BRB

[1058] Ian Hacking. Book review: The Scenes of Inquiry: On the Reality of Questions in the Sciences by Nicholas Jardine. Isis, 83(3):526–527, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233975.

Nersessian:1992:BRB

[1059] Nancy J. Nersessian. Book review: Particles and Waves: Historical Essays in the Philosophy of Science by Peter Achinstein. Isis, 83(3):

527–528, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233976.

Dupre:1992:BRB

[1060] John Dupré. Book review: Probabilistic Causality by Ellery Eells. Isis, 83 (3):528–529, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233977.

Baird:1992:BRB

[1061] Davis Baird. Book review: Instrumental Realism: The Interface between Philosophy of Science and Philosophy of Technology by Don Ihde. Isis, 83 (3):529–530, September 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/233978.

Koertge:1992:BRB

[1062] Noretta Koertge. Book review: Whose Science? Whose Knowledge? Thinking from Women’s Lives by Sandra Harding. Isis, 83(3):530, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233979.

Ramirez:1992:BRB

[1063] Santiago Ramirez. Book review: Les assises logiques et épistémologiques du progrès scientifique: Structures et téléonomies dans une logique des savoirs évolutifs by Noël Mouloud. Isis, 83(3):531, September 1992. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233980.

Tieszen:1992:BRB

[1064] Richard Tieszen. Book review: Edmund Husserl’s Origin of Geometry: An Introduction by Jacques Derrida; John P. Leavey. Isis, 83(3):531–532, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233981.

Tiles:1992:BRB

[1065] Mary Tiles. Book review: Gaston Bachelard, Subversive Humanist: Texts and Readings by Mary McAllester Jones. Isis, 83(3):532–533, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/233982.

Clulee:1992:BRB

[1066] Nicholas H. Clulee. Book review: John Dee’s Library Catalogue by Julian Roberts; Andrew G. Watson. Isis, 83(3):533–534, September 1992.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/233983.

Reingold:1992:BRB

[1067] Nathan Reingold. Book review: The Natural Sciences and American Scientists in the Revolutionary Era: A Bibliography by Katalin Harkányi. Isis, 83(3):534, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

233984.

Newman:1992:BRBa

[1068] William R. Newman. Book review: Alchemy in Europe: A Guide to Research by Claudia Kren. Isis, 83(3):534–536, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233985.

Sulloway:1992:BRB

[1069] Frank J. Sulloway. Book review: Freud-Bibliographie mit Werkkonkordanz by Ingeborg Meyer-Palmedo; Gerhard Fichtner. Isis, 83(3):536, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/233986.

Anonymous:1992:Cc

[1070] Anonymous. Collections. Isis, 83(3):536–540, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233987.

Anonymous:1992:BMd

[1071] Anonymous. Back matter. Isis, 83(3):541–546, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233988.

Anonymous:1992:FMd

[1072] Anonymous. Front matter. Isis, 83(3):??, September 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/233901.

Rochberg:1992:I

[1073] F. Rochberg. Introduction. Isis, 83(4):547–553, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234256.

Pingree:1992:HVH

[1074] David Pingree. Hellenophilia versus the history of science. Isis, 83(4): 554–563, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234257.

Lloyd:1992:MPH

[1075] G. E. R. Lloyd. Methods and problems in the history of ancient science: The Greek case. Isis, 83(4):564–577, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234258.

vonStaden:1992:AEH

[1076] Heinrich von Staden. Affinities and elisions: Helen and hellenocentrism. Isis, 83(4):578–595, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234259.

Bernal:1992:AOW

[1077] Martin Bernal. Animadversions on the origins of Western science. Isis, 83 (4):596–607, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234260.

Leavitt:1992:TMS

[1078] Judith Walzer Leavitt. “Typhoid Mary” strikes back bacteriological theory and practice in Early Twentieth–Century public health. Isis, 83 (4):608–629, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234261.

See [1971].

Anonymous:1992:CHS

[1079] Anonymous. Contributors to the History of Science Society. Isis, 83 (4):630–631, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234262.

Daston:1992:BRB

[1080] Lorraine Daston. Book review: Science Deified and Science Defied: The Historical Significance of Science in Western Culture. Volume 2: From the Early Modern Age through the Early Romantic Era, ca. 1640 to ca. 1820 by Richard Olson. Isis, 83(4):632–633, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234263.

Camerini:1992:BRBb

[1081] Jane R. Camerini. Book review: The Telling Image: The Changing Balance between Pictures and Words in a Technological Age by Duncan Davies; Diana Bathurst; Robin Bathurst. Isis, 83(4):633–634, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234264.

Kenzer:1992:BRB

[1082] Martin S. Kenzer. Book review: The Geography of Science by Harold Dorn. Isis, 83(4):634–635, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234265.

Ravetz:1992:BRB

[1083] Jerome R. Ravetz. Book review: Value-Free Science? Purity and Power in Modern Knowledge by Robert N. Proctor. Isis, 83(4):635–636, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234266.

Canizares:1992:BRB

[1084] Jorge Cañizares. Book review: La exploración botánica en Venezuela, 1754–1950 by Yolanda Texera Arnal. Isis, 83(4):636–637, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234267.

MacLeod:1992:BRB

[1085] Roy MacLeod. Book review: Australian Science in the Making by R. W. Home. Isis, 83(4):637–638, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234268.

Warner:1992:BRB

[1086] John Harley Warner. Book review: Medicine in America: A Short History by James H. Cassedy. Isis, 83(4):638–639, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234269.

Joy:1992:BRBc

[1087] Robert J. T. Joy. Book review: Oral Contraception in Perspective: Thirty Years of Clinical Experience with the Pill by A. D. G. Gunn; An Illustrated History of Contraception: A Concise Account of the Quest for Fertility Control by William H. Robertson; The History of the Management of Pain: From Early Principles to Present Practice by Ronald D. Mann; The Theory and Practice of Oncology: Historical Evolution and Present Principles by Ronald W. Raven; The History of Anesthesia: Proceedings of the Second International Symposium on the History of Anesthesia, London, July 1987 by Richard S. Atkinson; Thomas B. Boulton; The History of Orthopaedics: An Account of the Study and Practice of Orthopaedics from the Earliest Times to the Modern Era by David Le Vay. Isis, 83(4):639–641, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234270.

Staudenmaier:1992:BRB

[1088] John M. Staudenmaier. Book review: Under Technology’s Thumb by William Leiss. Isis, 83(4):641–642, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234271.

Layton:1992:BRB

[1089] Edwin T. Layton. Book review: A History of Industrial Power in the United States, 1780–1930. Volume 3: The Transmission of Power by Louis C. Hunter; Lynwood Bryant. Isis, 83(4):642, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234272.

vonStaden:1992:BRBb

[1090] Heinrich von Staden. Book review: The Lost Theory of Asclepiades of Bithynia by J. T. Vallance. Isis, 83(4):642–643, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234273.

Ebbesen:1992:BRBb

[1091] Sten Ebbesen. Book review: On Aristotle’s Categories by Ammonius; S. Marc Cohen; Gareth B. Matthews. Isis, 83(4):643–644, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234274.

Sobol:1992:BRBb

[1092] Peter G. Sobol. Book review: On Aristotle on the Intellect (de Anima 3.4-8) by Philoponus; William Charlton; Fernand Bossier. Isis, 83(4): 644–645, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234275.

Freudenthal:1992:BRB

[1093] Gad Freudenthal. Book review: Rencontres de cultures dans la philosophie médiévale: Traductions et traducteurs de l’antiquité tardive au XIVe siècle by Jacqueline Hamesse; Marta Fattori. Isis, 83(4):645–646, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/234276.

Newman:1992:BRBb

[1094] William R. Newman. Book review: The Book of the Secrets of Alchemy by Constantine of Pisa; Barbara Obrist. Isis, 83(4):646–647, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234277.

Ivry:1992:BRB

[1095] Alfred L. Ivry. Book review: Epítome de fisica (Filosofía de la naturaleza) by Averroes; Josep Puig; Averroes’ Questions in Physics by Helen Tunik Goldstein. Isis, 83(4):647–649, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234278.

Hughes:1992:BRB

[1096] Barnabas Hughes. Book review: Jordanus de Nemore, De elementis arithmetice artis: A Medieval Treatise on Number Theory by H. L. L. Busard. Isis, 83(4):649–650, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234279.

Kleinhenz:1992:BRB

[1097] Christopher Kleinhenz. Book review: Time and the Crystal: Studies in Dante’s “Rime Petrose.” by Robert M. Durling; Ronald L. Martinez. Isis, 83(4):650, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234280.

Swerdlow:1992:BRB

[1098] N. M. Swerdlow. Book review: Regiomontanus: His Life and Work by Ernst Zinner; Ezra Brown. Isis, 83(4):650–652, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234281.

Feingold:1992:BRBb

[1099] Mordechai Feingold. Book review: Authorizing Words: Speech, Writing, and Print in the English Renaissance by Martin Elsky. Isis, 83(4):652, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234282.

Wear:1992:BRB

[1100] Andrew Wear. Book review: Reappraisals of the Scientific Revolution by David C. Lindberg; Robert S. Westman. Isis, 83(4):652–654, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234283.

Findlen:1992:BRB

[1101] Paula Findlen. Book review: I Guicciardini e le scienze occulte: L’oroscopo di Francesco Guicciardini: Lettere di alchimia, astrologia e cabala a Luigi Guicciardini by Raffaella Castagnola. Isis, 83(4):654–655, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234284.

Lattis:1992:BRB

[1102] James M. Lattis. Book review: Le sfere omocentriche di Giovan Battista Amico nell’astronomia del Cinquecento: Con il testo del “De motibus corporum coelestium… ” by Mario Di Bono. Isis, 83(4):655–656, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234285.

Grafton:1992:BRBb

[1103] Anthony Grafton. Book review: Tractatus de praecognitionibus et praecognitis and Tractatio de demonstratione by Galileo Galilei; William F. Edwards; William A. Wallace. Isis, 83(4):656–657, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234286.

Moran:1992:BRB

[1104] Bruce T. Moran. Book review: Rabdology by John Napier; William Frank Richardson. Isis, 83(4):657, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234287.

Donahue:1992:BRB

[1105] William H. Donahue. Book review: The Poetic Structure of the World: Copernicus and Kepler by Fernand Hallyn; Donald M. Leslie. Isis, 83 (4):657–658, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234288.

Gingerich:1992:BRBb

[1106] Owen Gingerich. Book review: The Lord of Uraniborg: A Biography of Tycho Brahe by Victor E. Thoren; John R. Christianson. Isis, 83 (4):658–660, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234289.

Harkness:1992:BRB

[1107] Deborah E. Harkness. Book review: Idea, mente, specie: Platonismo e scienza in Johannes Marcus Marci (1595–1667) by Giuliana Mocchi. Isis, 83(4):660, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234290.

Iliffe:1992:BRB

[1108] Rob Iliffe. Book review: Henry More (1614–1687): Tercentenary Studies by Sarah Hutton; Robert Crocker. Isis, 83(4):660–661, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234291.

Hatch:1992:BRB

[1109] Robert A. Hatch. Book review: Conversations on the Plurality of Worlds by Bernard le Bovier de Fontenelle; H. A. Hargreaves. Isis, 83(4):661– 662, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234292.

Rutherford:1992:BRB

[1110] Donald Rutherford. Book review: G. W. Leibniz’s Monadology: An Edition for Students by Nicholas Rescher. Isis, 83(4):662–663, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234293.

Dear:1992:BRBb

[1111] Peter Dear. Book review: The Philosophical Writings of Descartes. Volume 3: The Correspondence by Rene Descartes; John Cottingham; Robert Soothoff; Dugald Murdoch; Anthony Kenny. Isis, 83(4):663–664, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234294.

Heyd:1992:BRB

[1112] Michael Heyd. Book review: Essays on the Context, Nature, and Influence of Isaac Newton’s Theology by James E. Force; Richard H. Popkin. Isis, 83(4):664–665, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

234295.

Hogendijk:1992:BRB

[1113] Jan P. Hogendijk. Book review: “Apollonius Saxonicus”: Die Restitution eines verlorenen Werkes des Apollonius von Perga durch Joachim Jungius, Woldeck Weland und Johannes Müller by Bernd Elsner. Isis, 83 (4):665–666, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234296.

McMullen:1992:BRB

[1114] Tom McMullen. Book review: John Craige’s Mathematical Principles of Christian Theology by Richard Nash. Isis, 83(4):666, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234297.

Broman:1992:BRBb

[1115] Thomas Broman. Book review: Der frühmoderne Staat und die europäische Universität: Zur Interaktion von Politik und Erziehungssystem im Prozess ihrer Ausdifferenzierung by Rudolf Stichweh. Isis, 83 (4):667–668, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234298.

Shea:1992:BRBc

[1116] William R. Shea. Book review: Immagini di Descartes nell’Encyclopédie by Mariafranca Spallanzani. Isis, 83(4):668, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234299.

Pedersen:1992:BRB

[1117] Kurt Moller Pedersen. Book review: Berkeley’s Revolution in Vision by Margaret Atherton. Isis, 83(4):668–669, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234300.

Guerrini:1992:BRBb

[1118] Anita Guerrini. Book review: The English Malady (1733) by George Cheyne; Roy Porter. Isis, 83(4):669–670, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234301.

Golinski:1992:BRBb

[1119] Jan Golinski. Book review: Studies in the Philosophy of the Scottish Enlightenment by M. A. Stewart. Isis, 83(4):670–671, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234302.

McPeak:1992:BRBb

[1120] William J. McPeak. Book review: Über Ignaz von Born und die Societät der Bergbaukunde: Vorträge einer Gedenkveranstaltung zur 200. Wiederkehr des Gründungstages im September 1786 der ältesten internationalen wissenschaftlichen Gesellschaft by Günter B. Fettweis; Günther Hamann. Isis, 83(4):671, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234303.

Silliman:1992:BRB

[1121] Robert H. Silliman. Book review: History of the Principle of Interference of Light by Nahum Kipnis. Isis, 83(4):671–672, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234304.

Strickland:1992:BRB

[1122] Stuart Walker Strickland. Book review: Die Suche nach der Wissenschaftlichkeit der Physiologie in der Zeit der Romantik by Brigitte Lohff. Isis, 83(4):672–673, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234305.

Kremer:1992:BRB

[1123] Richard L. Kremer. Book review: Staat kontra Universität: Die Universität Halle-Wittenberg und die Karlsbader Beschlüsse, 1819–1848 by Manfred Brümmer. Isis, 83(4):673–674, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234306.

Laita:1992:BRB

[1124] Luis M. Laita. Book review: Boole, 1815–1864: L’oiseau de nuit en plein jour by Souleymane Bachir Diagne. Isis, 83(4):674–675, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234307.

Gall:1992:BRB

[1125] Gilbert J. Gall. Book review: Romantic Medicine and John Keats by Hermione De Almeida. Isis, 83(4):675–676, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234308.

Wolfe:1992:BRB

[1126] Thomas J. Wolfe. Book review: Enter the Physician: The Transformation of Domestic Medicine, 1760–1860 by Lamar Riley Murphy. Isis, 83 (4):676–677, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234309.

Lightman:1992:BRB

[1127] Bernard Lightman. Book review: Thomas Henry Huxley: Communicating for Science by J. Vernon Jensen. Isis, 83(4):677–678, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234310.

Baatz:1992:BRB

[1128] Simon Baatz. Book review: Transforming Traditions in American Biology, 1880–1915 by Jane Maienschein. Isis, 83(4):678–679, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234311.

Reichenbach:1992:BRB

[1129] Herman Reichenbach. Book review: Le Musée d’Ethnographie du Trocadéro (1878–1908): Anthropologie et muséologie en France by Nélia Dias. Isis, 83(4):679–680, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234312.

Hiebert:1992:BRBc

[1130] Erwin N. Hiebert. Book review: Principien der Naturfilosofi: Lectures on Natural Philosophy, 1903–1906 by Ludwig Boltzmann; Ilse M. Fasol-Boltzmann. Isis, 83(4):680–681, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234313.

Schweber:1992:BRBb

[1131] S. S. Schweber. Book review: Lawrence and His Laboratory: A History of the Lawrence Berkeley Laboratory. Volume 1 by John L. Heilbron; Robert W. Seidel. Isis, 83(4):681–682, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234314.

Cao:1992:BRB

[1132] Tian Yu Cao. Book review: Theory Construction and Selection in Modern Physics: The S Matrix by James T. Cushing. Isis, 83(4):682–684, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234315.

Pyenson:1992:BRB

[1133] Lewis Pyenson. Book review: Physics in Australia to 1945: Bibliography and Biographical Register by Roderick Weir Home; Paula J. Needham; Physics and the Rise of Scientific Research in Canada by Yves Gingras; Peter Keating; In Celebration of Canadian Scientists: A Decade of Killam Laureates by Geraldine A. Kenney-Wallace; Mel G. MacLeod; Ralph Gordon Stanton. Isis, 83(4):684–685, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234316.

Jensen:1992:BRB

[1134] William B. Jensen. Book review: Some Recollections of Gap Jumping by Derek H. R. Barton; From Design to Discovery by Donald J. Cram; Steroids Made It Possible by Carl Djerassi; From Cologne to Chapel Hill by Ernest L. Eliel; Enjoying Organic Chemistry by Egbert Havinga; Explorations with Sugars: How Sweet It Was by Raymond U. Lemieux; My 132 Semesters of Chemistry Studies: Studium chymiae nec nisi cum morte finitur by Vladimir Prelog; Otto Theodor Benfey; David Ginsburg; The Right Place at the Right Time by John D. Roberts. Isis, 83 (4):685–687, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234317.

Blukis:1992:BRB

[1135] Uldis Blukis. Book review: A History of Lactic Acid Making: A Chapter in the History of Biotechnology by H. Benninga. Isis, 83(4):687–688, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234318.

Norberg:1992:BRB

[1136] Arthur Norberg. Book review: Innovating for Failure: Government Policy and the Early British Computing Industry by John Hendry. Isis, 83 (4):688–690, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234319.

Waff:1992:BRB

[1137] Craig B. Waff. Book review: Space Technology and Planetary Astronomy by Joseph N. Taterewicz. Isis, 83(4):690–691, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234320.

Smith:1992:BRBd

[1138] Robert W. Smith. Book review: The Sputniks Crisis and Early United States Space Policy: A Critique of the Historiography of Space by Rip Bulkeley. Isis, 83(4):691–692, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234321.

Rider:1992:BRB

[1139] Robin E. Rider. Book review: The Biological Sciences in the Twentieth Century by Merriley Borell; The Physical Sciences in the Twentieth Century by Owen Gingerich. Isis, 83(4):692–693, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234322.

Mitman:1992:BRB

[1140] Gregg Mitman. Book review: The Shaping of Environmentalism in America by Victor B. Scheffer. Isis, 83(4):693, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234323.

Oppenheimer:1992:BRB

[1141] Gerald M. Oppenheimer. Book review: History of AIDS: Emergence and Origin of a Modern Pandemic by Mirko D. Grmek; Russell C. Maulitz; Jacalyn Duffin. Isis, 83(4):693–694, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234324.

Weingart:1992:BRB

[1142] Peter Weingart. Book review: NS-Pharmazie: Gleichschaltung des deutschen Apothekenwesens im Dritten Reich by Gerald Schroder. Isis, 83(4):695, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234325.

Feldberg:1992:BRB

[1143] Gina D. Feldberg. Book review: Our Own Master Race: Eugenics in Canada, 1885–1945 by Angus McLaren. Isis, 83(4):695–696, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234326.

Lindee:1992:BRB

[1144] M. Susan Lindee. Book review: Exons, Introns, and Talking Genes: The Science behind the Human Genome Project by Christopher Wills. Isis, 83 (4):696–697, December 1992. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/234327.

Whittemore:1992:BRB

[1145] Gilbert Whittemore. Book review: Multiple Exposures: Chronicles of the Radiation Age by Catherine Caufield. Isis, 83(4):697–698, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234328.

Holt:1992:BRB

[1146] Robert R. Holt. Book review: Freud Evaluated: The Completed Arc. by Malcolm Macmillan. Isis, 83(4):698, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234329.

Apple:1992:BRB

[1147] Michael W. Apple. Book review: The Burt Affair by Robert B. Joynson; Science, Ideology, and the Media: The Cyril Burt Scandal by Ronald Fletcher. Isis, 83(4):698–700, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/234330.

Harris:1992:BRBc

[1148] Benjamin Harris. Book review: From Asylum to Community: Mental Health Policy in Modern America by Gerald N. Grob. Isis, 83(4):701– 702, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234331.

Herman:1992:BRB

[1149] Ellen Herman. Book review: The Founders of Humanistic Psychology by Roy José DeCarvalho. Isis, 83(4):702, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/234332.

Sigurdsson:1992:BRBb

[1150] Skuli Sigurdsson. Book review: Phänomenologie der Mathematik: Elemente einer phänomenologischen Aufklärung der mathematischen Erkenntnis nach Husserl by Dieter Lohmar. Isis, 83(4):702–704, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/234333.

Fuller:1992:BRBb

[1151] Steve Fuller. Book review: Scientism: Philosophy and the Infatuation with Science by Tom Sorell. Isis, 83(4):704, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234334.

Carroll:1992:BRB

[1152] John W. Carroll. Book review: Science and Necessity by John Bigelow; Robert Pargetter. Isis, 83(4):704–705, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/234335.

Riddle:1992:BRBb

[1153] John M. Riddle. Book review: Dictionary of Protopharmacology: Therapeutic Practices, 1700–1850 by J. Worth Estes. Isis, 83(4):705–706, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/234336.

Woodward:1992:BRB

[1154] David Woodward. Book review: Modern Geography: An Encyclopedic Survey by Gary S. Dunbar. Isis, 83(4):706, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234337.

Anonymous:1992:Cd

[1155] Anonymous. Collections. Isis, 83(4):706–709, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234338.

Anonymous:1992:BMe

[1156] Anonymous. Back matter. Isis, 83(4):710–719, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234339.

Anonymous:1992:VI

[1157] Anonymous. Volume information. Isis, 83(4):720–740, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/234254.

Anonymous:1992:FMe

[1158] Anonymous. Front matter. Isis, 83(4):??, December 1992. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/234255.

Neu:1993:CBH

[1159] John Neu. Current bibliography of the history of science and its cultural influences, 1993. Isis, 84(??):1–297, ???? 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235818.

Anonymous:1993:BMa

[1160] Anonymous. Back matter. Isis, 84(??):??, ???? 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235819.

Anonymous:1993:FMa

[1161] Anonymous. Front matter. Isis, 84(??):i–xvi, ???? 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235817.

Cain:1993:CPC

[1162] Joseph Allen Cain. Common problems and cooperative solutions: Organizational activity in evolutionary studies, 1936–1947. Isis, 84(1):1–25, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235551.

Slotten:1993:DSU

[1163] Hugh Richard Slotten. The dilemmas of science in the United States: Alexander Dallas Bache and the U.S. Coast Survey. Isis, 84(1):26–49, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235552.

Morus:1993:CUE

[1164] Iwan Rhys Morus. Currents from the underworld: Electricity and the technology of display in Early Victorian England. Isis, 84(1):50–69, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235553.

Sarasohn:1993:NCF

[1165] Lisa T. Sarasohn. Nicolas-claude fabri de peiresc and the patronage of the new science in the Seventeenth Century. Isis, 84(1):70–90, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235554.

Palladino:1993:SI

[1166] Paolo Palladino and Michael Worboys. Science and imperialism. Isis, 84(1):91–102, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235555.

Pyenson:1993:CIE

[1167] Lewis Pyenson. Cultural imperialism and exact sciences revisited. Isis, 84(1):103–108, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235556.

Murad:1993:LE

[1168] Hazim Murad, Gerald Holton, Garry J. Tee, and Lamar Riley Murphy. Letters to the Editor. Isis, 84(1):109–112, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235557.

Herbert:1993:BRB

[1169] Sandra Herbert. Book review: Darwin by Adrian Desmond; James Moore; Charles Darwin: The Man and His Influence by Peter J. Bowler; One Long Argument: Charles Darwin and the Genesis of Modern Evolutionary Thought by Ernst Mayr; Darwin in Italy: Science across Cultural Frontiers by Giuliano Pancaldi; Ruey Brodine Morelli; Charles Darwin’s Marginalia. Volume 1 by Charles Darwin; Mario A. Di Gregorio; N. W. Gill. Isis, 84(1):113–124, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235558.

Weiner:1993:BRB

[1170] Douglas R. Weiner. Book review: Darwin without Malthus: The Struggle for Existence in Russian Evolutionary Thought by Daniel P. Todes; Darwin in Russian Thought by Alexander Vucinich. Isis, 84(1):124–127, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235559.

Eisberg:1993:BRB

[1171] Joann Eisberg. Book review: Exploring the Sun: Solar Science since Galileo by Karl Hufbauer. Isis, 84(1):128, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235560.

McEvoy:1993:BRBa

[1172] John G. McEvoy. Book review: Chemistry by Helene Metzger; Colette V. Michael. Isis, 84(1):128–129, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235561.

Langdon:1993:BRB

[1173] John Langdon. Book review: Mechanics of Preindustrial Technology by Brian Cotterell; Johan Kamminga. Isis, 84(1):129–130, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235562.

Higby:1993:BRB

[1174] Gregory J. Higby. Book review: Royal Pharmaceutical Society of Great Britain, 1841–1991: A Political and Social History by S. W. F. Holloway. Isis, 84(1):130–131, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235563.

Prescott:1993:BRB

[1175] Heather Munro Prescott. Book review: Women, Health, and Medicine in America: A Historical Handbook by Rima D. Apple. Isis, 84(1):131– 132, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235564.

Federspiel:1993:BRB

[1176] Michel Federspiel. Book review: Les éléments. Traduits du texte de Heiberg. Volume 1: Introduction générale by Euclide d’Alexandrie; Maurice Caveing; Livres I–III: Géometrie plane by Bernard Vitrac. Isis, 84 (1):132–133, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235565.

Aveni:1993:BRBa

[1177] Anthony F. Aveni. Book review: The Myth of Replacement: Stars, Gods, and Order in the Universe by Thomas D. Worthen. Isis, 84(1):133, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235566.

Knorr:1993:BRB

[1178] Wilbur R. Knorr. Book review: Timee/Critias by Platon; Luc Brisson; Michel Patillon. Isis, 84(1):134–135, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235567.

Upton:1993:BRB

[1179] Thomas V. Upton. Book review: The Origins of Aristotelian Science by Michael Ferejohn. Isis, 84(1):135–136, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235568.

Vallance:1993:BRB

[1180] J. T. Vallance. Book review: Hippocrates in a World of Pagans and Christians by Owsei Temkin. Isis, 84(1):136–137, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235569.

Whitney:1993:BRB

[1181] Elspeth Whitney. Book review: Practical Geometry (Practica geometriae) by Hugh of St. Victor; Frederick A. Homann. Isis, 84(1):137, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235570.

Grant:1993:BRB

[1182] Edward Grant. Book review: The Aristotelians of Renaissance Italy: A Philosophical Exposition by Dominick A. Iorio. Isis, 84(1):137–138, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235571.

Jordan:1993:BRBa

[1183] Mark D. Jordan. Book review: Filosofia della natura nella Schola Salernitana del secolo XII by Piero Morpurgo. Isis, 84(1):138–139, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235572.

Freudenthal:1993:BRB

[1184] Gad Freudenthal. Book review: Perspectives on Maimonides: Philosophical and Historical Studies by Joel L. Kraemer; Lawrence V. Berman; Moses Maimonides and His Time by Eric L. Ormsby. Isis, 84(1):139– 141, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235573.

Goddu:1993:BRB

[1185] Andre Goddu. Book review: Die Kosmologie des Franciscus de Marchia: Texte, Quellen und Untersuchungen zur Naturphilosophie des 14. Jahrhunderts by Notker Schneider. Isis, 84(1):141, March 1993. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235574.

Konings:1993:BRB

[1186] Patricia Konings. Book review: History of Science and Technology in the Song, Liao, Xia, Jin, and Yuan Dynasties by Guan Chengxue. Isis, 84(1):141–142, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235575.

Hanson:1993:BRB

[1187] Marta Hanson. Book review: Chinese Thought, Society, and Science: The Intellectual and Social Background of Science and Technology in Premodern China by Derk Bodde. Isis, 84(1):142–143, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235576.

Wear:1993:BRB

[1188] Andrew Wear. Book review: On Fracture of the Skull or Cranium by Berengario da Carpi; L. R. Lind. Isis, 84(1):143–144, March 1993. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235577.

Clulee:1993:BRB

[1189] Nicholas H. Clulee. Book review: Giordano Bruno and the Embassy Affair by John Bossy. Isis, 84(1):144–145, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235578.

Principe:1993:BRB

[1190] Lawrence M. Principe. Book review: The Alchemical World of the German Court: Occult Philosophy and Chemical Medicine in the Circle of Moritz of Hessen (1572–1632) by Bruce T. Moran. Isis, 84(1):145, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235579.

Eamon:1993:BRB

[1191] William Eamon. Book review: Chemical Pharmacy Enters the University: Johannes Hartmann and the Didactic Care of Chymiatria in the Early Seventeenth Century by Bruce T. Moran. Isis, 84(1):145–146, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235580.

Mahoney:1993:BRBa

[1192] Michael S. Mahoney. Book review: Cartesian Method and the Problem of Reduction by Emily R. Grosholz; The Magic of Numbers and Motion: The Scientific Career of Rene Descartes by William R. Shea. Isis, 84 (1):146–148, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235581.

Dear:1993:BRBa

[1193] Peter Dear. Book review: Promoting Experimental Learning: Experiment and the Royal Society, 1660–1727 by Marie Boas Hall. Isis, 84(1):148– 149, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235582.

Sloan:1993:BRB

[1194] Jan Butin Sloan. Book review: Settecento inquieto: Alle origini dell’ Istituto delle Scienze di Bologna by Marta Cavazza. Isis, 84(1):149– 150, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235583.

Shea:1993:BRBa

[1195] William R. Shea. Book review: Il passato, la memoria, l’oblio: Sei saggi di storia delle idee by Paolo Rossi. Isis, 84(1):150–151, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235584.

Grafton:1993:BRBa

[1196] Anthony Grafton. Book review: The Uses of Antiquity: The Scientific Revolution and the Classical Tradition by Stephen Gaukroger. Isis, 84 (1):151–152, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235585.

Malet:1993:BRB

[1197] Antoni Malet. Book review: Apariencia y verdad: Estudio sobre la filosofia de P. L. M. de Maupertuis (1698–1759) by Juan Arana Canedo-Arguelles. Isis, 84(1):152, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235586.

Taylor:1993:BRBa

[1198] Kenneth L. Taylor. Book review: Pallas’ Theory of the Earth in German (1778): Translation and Reevaluation; Reaction by a Contemporary: H.- B. de Saussure by Albert V. Carozzi; Marguerite Carozzi. Isis, 84(1):

152–153, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235587.

Oppenheimer:1993:BRB

[1199] Jane Oppenheimer. Book review: Congettura ed esperienza nella fisiologia di Haller: La riforma dell’ anatomia e il sistema della generazione by Maria Teresa Monti. Isis, 84(1):153, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235588.

Dawson:1993:BRB

[1200] Virginia P. Dawson. Book review: The Correspondence between Albrecht von Haller and Horace-Benedict de Saussure by Otto Sonntag. Isis, 84 (1):153–154, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235589.

Evans:1993:BRBa

[1201] James Evans. Book review: Les origines de la Société de Physique et d’Histoire Naturelle (1790–1822): La science génèvoise face au modèle français by Rene Sigrist. Isis, 84(1):154–155, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235590.

Rainger:1993:BRB

[1202] Ronald Rainger. Book review: Reading the Shape of Nature: Comparative Zoology at the Agassiz Museum by Mary P. Winsor. Isis, 84(1): 155–156, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235591.

Slotten:1993:BRB

[1203] Hugh R. Slotten. Book review: The Origins of Natural Science in America: The Essays of George Brown Goode by George Brown Goode; Sally Gregory Kohlstedt. Isis, 84(1):156, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235592.

Larson:1993:BRB

[1204] Edward J. Larson. Book review: Science, American Style by Nathan Reingold. Isis, 84(1):157, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235593.

Eyler:1993:BRB

[1205] John M. Eyler. Book review: A Science of Impurity: Water Analysis in Nineteenth-Century Britain by Christopher Hamlin. Isis, 84(1):157– 158, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235594.

Turner:1993:BRB

[1206] Frank M. Turner. Book review: Philosophy and Biblical Interpretation: A Study in Nineteenth-Century Conflict by Peter Addinall. Isis, 84(1): 158–159, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235595.

Trumpler:1993:BRB

[1207] Maria Trumpler. Book review: The Mechanism of Human Facial Expression by G.-B. Duchenne de Boulogne; R. Andrew Cuthbertson. Isis, 84(1):159–160, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235596.

LaBerge:1993:BRB

[1208] Ann F. La Berge. Book review: Inheriting Madness: Professionalization and Psychiatric Knowledge in Nineteenth-Century France by Ian R. Dowbiggin. Isis, 84(1):160–161, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235597.

Groopman:1993:BRB

[1209] Leonard Groopman. Book review: The Physician-Legislators of France: Medicine and Politics in the Early Third Republic, 1870–1914 by Jack D. Ellis. Isis, 84(1):161–162, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235598.

Cahan:1993:BRB

[1210] David Cahan. Book review: Physics as a Calling: Discipline and Practice in the Konigsberg Seminar for Physics by Kathryn M. Olesko. Isis, 84 (1):162–163, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235599.

Tuchman:1993:BRB

[1211] Arleen Tuchman. Book review: The German Experience of Professionalization: Modern Learned Professions and Their Organizations from the Early Nineteenth Century to the Hitler Era by Charles E. McClelland. Isis, 84(1):163–164, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235600.

Crowe:1993:BRB

[1212] Gregory D. Crowe. Book review: Nauchnaia perepiska P. N. Lebedeva [The Scientific Correspondence of P. N. Lebedev] by P. N. Lebedev; V. A. Fabrikant; E. I. Pogrebyskaia. Isis, 84(1):164–165, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235601.

Archibald:1993:BRBa

[1213] Tom Archibald. Book review: The Man Who Knew Infinity: A Life of the Genius Ramanujan by Robert Kanigel. Isis, 84(1):165–166, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235602.

Nersessian:1993:BRB

[1214] Nancy J. Nersessian. Book review: Through Measurement to Knowledge: The Selected Papers of Heike Kamerlingh Onnes, 1853–1926 by Heike Kamerlingh Onnes; Kostas Gavroglu; Yorgos Goudaroulis. Isis, 84(1): 166–167, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235603.

Macrakis:1993:BRB

[1215] Kristie Macrakis. Book review: Forschung im Spanungsfeld von Politik und Gesellschaft: Geschichte und Struktur der Kaiser-Wilhelm-/Max-Planck-Gesellschaft by Rudolf Vierhaus; Bernhard vom Brocke; Chronik der Kaiser-Wilhelm-Gesellschaft zu Forderung der Wissenschaften by Eckart Henning; Marion Kazemi; Geschichte des Kaiser-Wilhelm-Instituts fur Kohlenforschung, 1913–1943 by Manfred Rasch. Isis, 84 (1):167–168, March 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235604.

Hendry:1993:BRB

[1216] John Hendry. Book review: Niels Bohr: His Heritage and Legacy: An Anti-Realist View of Quantum Mechanics by Jan Faye. Isis, 84(1):169, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235605.

Josephson:1993:BRB

[1217] Paul Josephson. Book review: Kapitza in Cambridge and Moscow: Life and Letters of a Russian Physicist by Peter L. Kapitza; J. W. Boag; P. E. Rubinin; D. Shoenberg. Isis, 84(1):169–170, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235606.

Geiger:1993:BRB

[1218] Roger Geiger. Book review: Millikan’s School: A History of the California Institute of Technology by Judith R. Goodstein. Isis, 84(1):170–171, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235607.

Hackmann:1993:BRB

[1219] Willem Hackmann. Book review: The Laser in America, 1950–1970 by Joan Lisa Bromberg. Isis, 84(1):171–172, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235608.

LaFollette:1993:BRB

[1220] Marcel C. LaFollette. Book review: Too Hot to Handle: The Race for Cold Fusion by Frank Close. Isis, 84(1):172–173, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235609.

Grandy:1993:BRB

[1221] David Grandy. Book review: Thinking about Science: Max Delbruck and the Origins of Molecular Biology by Ernst Peter Fischer; Carol Lipson. Isis, 84(1):173–174, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235610.

Weingart:1993:BRB

[1222] Peter Weingart. Book review: NS-Biologie by Anne Baumer. Isis, 84(1): 174, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235611.

Ihde:1993:BRB

[1223] Aaron J. Ihde. Book review: The World of Peptides: A Brief History of Peptide Chemistry by T. Wieland; M. Bodanszky. Isis, 84(1):174– 175, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235612.

Keller:1993:BRB

[1224] Evelyn Fox Keller. Book review: A Conceptual History of Modern Embryology by Scott Gilbert. Isis, 84(1):175–176, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235613.

Montgomery:1993:BRBa

[1225] William Montgomery. Book review: Anton Dohrn: A Life for Science by Theodor Heuss; Christiane Groeben; Liselotte Dieckmann; Margaret Boveri. Isis, 84(1):176–177, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235614.

Maienschein:1993:BRB

[1226] Jane Maienschein. Book review: Houses for Science: A Pictorial History of Cold Spring Harbor Laboratory by Elizabeth L. Watson. Isis, 84(1): 177, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235615.

Taylor:1993:BRBb

[1227] Peter J. Taylor. Book review: Pioneer Ecologist: The Life and Work of Victor Ernest Shelford, 1877–1968 by Robert A. Croker; Foundations of Ecology: Classic Papers with Commentaries by Leslie A. Real; James H. Brown. Isis, 84(1):177–179, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235616.

Parisi:1993:BRB

[1228] Thomas Parisi. Book review: Freud, Dora, and Vienna 1900 by Hannah S. Decker. Isis, 84(1):179, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235617.

Sutphen:1993:BRBa

[1229] Molly Sutphen. Book review: Science and Empire: East Coast Fever in Rhodesia and the Transvaal by Paul F. Cranefield; Bilharzia: A History of Imperial Tropical Medicine by John Farley. Isis, 84(1):179–181, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235618.

Eakin:1993:BRB

[1230] Marshall C. Eakin. Book review: A Space for Science: The Development of the Scientific Community in Brazil by Simon Schwartzman. Isis, 84(1):

181, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235619.

Nugent:1993:BRB

[1231] Angela Nugent. Book review: Deadly Dust: Silicosis and the Politics of Occupational Disease in Twentieth–Century America by David Rosner; Gerald Markowitz. Isis, 84(1):181–182, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235620.

Apple:1993:BRB

[1232] Rima D. Apple. Book review: Vitamin C and Cancer: Medicine or Politics? by Evelleen Richards. Isis, 84(1):182–183, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235621.

Tobey:1993:BRB

[1233] Ronald C. Tobey. Book review: Electrifying America: Social Meanings of a New Technology, 1880–1940 by David E. Nye. Isis, 84(1):183–184, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235622.

Todd:1993:BRBa

[1234] Edmund N. Todd. Book review: Alternating Currents: Nationalized Power in France, 1946–1970 by Robert L. Frost. Isis, 84(1):184–185, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235623.

Schatzberg:1993:BRB

[1235] Eric Schatzberg. Book review: Airshipmen, Businessmen, and Politics, 1890–1940 by Henry Cord Meyer. Isis, 84(1):185–186, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235624.

Roche:1993:BRB

[1236] John J. Roche. Book review: Spacetime and Electromagnetism: An Essay on the Philosophy of the Special Theory of Relativity by J. R. Lucas; P. E. Hodgson. Isis, 84(1):186–187, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235625.

McMullin:1993:BRB

[1237] Ernan McMullin. Book review: Inventer l’univers: Le probleme de la connaissance et les modeles cosmologiques by Luc Brisson; E. Walter Meyerstein. Isis, 84(1):187–188, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235626.

Sklar:1993:BRB

[1238] Lawrence Sklar. Book review: Time, Space, and Philosophy by Christopher Ray. Isis, 84(1):189, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235627.

Kisiel:1993:BRB

[1239] Theodore Kisiel. Book review: Phenomenology of Natural Science by Lee Hardy; Lester Embree. Isis, 84(1):189–190, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235628.

Woodward:1993:BRBa

[1240] William R. Woodward. Book review: Die Grundlagen der Psychoanalyse: Eine philosophische Kritik. The Foundations of Psychoanalysis: A Philosophical Critique by Adolf Grunbaum; Christa Kolbert. Isis, 84(1): 190–191, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235629.

Fuller:1993:BRBa

[1241] Steve Fuller. Book review: Beyond Reason: Essays on the Philosophy of Paul Feyerabend by Gonzalo Munevar. Isis, 84(1):191–193, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235630.

Gutting:1993:BRB

[1242] Gary Gutting. Book review: Foucault’s Strata and Fields: An Investigation into Archaeological and Genealogical Science Studies by Martin Kusch. Isis, 84(1):193, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235631.

Salmon:1993:BRB

[1243] Merrilee H. Salmon. Book review: Archaeological Typology and Practical Reality: A Dialectical Approach to Artifact Classification and Sorting by William Y. Adams; Ernest W. Adams. Isis, 84(1):193–194, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235632.

Latour:1993:BRB

[1244] Bruno Latour. Book review: Social Control and Multiple Discovery in Science: The Opiate Receptor Case by Susan E. Cozzens. Isis, 84(1):194– 195, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235633.

Ogilvie:1993:BRB

[1245] Marilyn Bailey Ogilvie. Book review: Women in Science: Token Women or Gender Equality? by Veronica Stolte-Heiskanen; Feride Acar; Nora Ananieva; Dorothea Gaudart; Ruza Furst-Dilic. Isis, 84(1):195–196, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235634.

Stuhlhofer:1993:BRB

[1246] Franz Stuhlhofer. Book review: Astronomica et astrologica Cracoviensia ante annum 1550 by Mieczyslaw Markowski. Isis, 84(1):196–197, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235635.

Fye:1993:BRB

[1247] W. Bruce Fye. Book review: Heirs of Hippocrates: The Development of Medicine in a Catalogue of Historic Books in the Hardin Library for the Health Sciences, The University of Iowa by Richard Eimas. Isis, 84(1): 197, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235636.

Schiebinger:1993:BRB

[1248] Londa Schiebinger. Book review: The Byrth of Mankynde, Otherwyse Named The Womans Booke: Embryology, Obstetrics, Gynaecology through Four Centuries: An Illustrated and Annotated Catalogue of Rare Books in the Library of the Swedish Society of Medicine by Ove Hagelin. Isis, 84(1):197–198, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235637.

Lawrence:1993:BRB

[1249] Susan C. Lawrence. Book review: Medicine in Great Britain from the Restoration to the Nineteenth Century, 1660–1800: An Annotated Bibliography by Samuel J. Rogal. Isis, 84(1):198, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235638.

Anonymous:1993:Ca

[1250] Anonymous. Collections. Isis, 84(1):198–205, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235639.

Anonymous:1993:BMb

[1251] Anonymous. Back matter. Isis, 84(1):206–210, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235640.

Anonymous:1993:FMb

[1252] Anonymous. Front matter. Isis, 84(1):??, March 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235550.

Shackelford:1993:TBL

[1253] Jole Shackelford. Tycho Brahe, laboratory design, and the aim of science: Reading plans in context. Isis, 84(2):211–230, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236232.

Rupke:1993:ROV

[1254] Nicolaas A. Rupke. Richard Owen’s vertebrate archetype. Isis, 84(2): 231–251, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236233.

Pang:1993:SES

[1255] Alex Soojung-Kim Pang. The social event of the season: Solar eclipse expeditions and Victorian culture. Isis, 84(2):252–277, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236234.

Carson:1993:AAA

[1256] John Carson. Army alpha, Army brass, and the search for Army intelligence. Isis, 84(2):278–309, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236235.

Feingold:1993:NLB

[1257] Mordechai Feingold. Newton, Leibniz, and Barrow too: An attempt at a reinterpretation. Isis, 84(2):310–338, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236236.

Porter:1993:AMH

[1258] Theodore M. Porter, Karl Huibauer, Michael M. Sokal, Joan Richards, and Marshall Clagett. Annual meeting of the History of Science Society Washington, D.C., 27–30 December 1992. Isis, 84(2):339–346, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236237.

Outram:1993:BRB

[1259] Dorinda Outram. Book review: Monstres: Histoire du corps et de ses defauts. Isis, 84(2):347–352, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236238.

Sargent:1993:BRB

[1260] Steven D. Sargent. Book review: Inventing the Flat Earth: Columbus and Modern Historians by Jeffrey Burton Russell. Isis, 84(2):353, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236239.

Thackray:1993:BRB

[1261] Arnold Thackray. Book review: George Sarton i la historia de la ciencia a Espanya by Thomas F. Glick; Merce Viladrich. Isis, 84(2):353–354, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236240.

Burlingame:1993:BRB

[1262] Leslie J. Burlingame. Book review: The Scientific Lady: A Social History of Women’s Scientific Interests, 1520–1918 by Patricia Phillips. Isis, 84 (2):354–355, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236241.

Rosenkrantz:1993:BRB

[1263] Barbara Gutmann Rosenkrantz. Book review: The Quest for Authority and Honor in the American Professions, 1750–1900 by Samuel Haber. Isis, 84(2):355–356, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236242.

Reingold:1993:BRB

[1264] Nathan Reingold. Book review: Science at Harvard University: Historical Perspectives by Clark A. Elliott; Margaret W. Rossiter. Isis, 84(2): 356–357, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236243.

Paul:1993:BRB

[1265] Diane Paul. Book review: Biotechnics and Society: The Rise of Industrial Genetics by Sheldon Krimsky. Isis, 84(2):357–358, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236244.

Weindling:1993:BRB

[1266] Paul Weindling. Book review: Im Spiegel der Arznei: Sozialgeschichte der Medizin by Paul Ridder. Isis, 84(2):358, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236245.

Siraisi:1993:BRB

[1267] Nancy G. Siraisi. Book review: I maestri di medicina ed arti dell’Universita di Ferrara, 1391–1950 by Francesco Raspadori. Isis, 84 (2):358–359, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236246.

Thibodeau:1993:BRB

[1268] Kenneth Thibodeau. Book review: Les sciences en Alsace, 1538–1988. Isis, 84(2):359–360, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236247.

Gottlieb:1993:BRB

[1269] Paula Gottlieb. Book review: Primary Ousia: An Essay on Aristotle’s Metaphysics Z and H by Michael J. Loux. Isis, 84(2):360–361, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236248.

Lennox:1993:BRB

[1270] James G. Lennox. Book review: Aristotle’s Physics: A Collection of Essays by Lindsay Judson. Isis, 84(2):361–363, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236249.

ONeill:1993:BRB

[1271] Ynez Viole O’Neill. Book review: On the Therapeutic Method: Books I and II by Galen; R. J. Hankinson. Isis, 84(2):363–364, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236250.

Taub:1993:BRB

[1272] Liba Taub. Book review: Origen and the Life of the Stars: A History of an Idea by Alan Scott. Isis, 84(2):364, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236251.

Sylla:1993:BRB

[1273] Edith Sylla. Book review: A World without Women: The Christian Clerical Culture of Western Science by David F. Noble. Isis, 84(2):364– 366, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236252.

Eastwood:1993:BRB

[1274] Bruce S. Eastwood. Book review: Macrobius im Mittelalter: Ein Beitrag zur Rezeptionsgeschichte der Commentarii in Somnium Scipionis by Albrecht Huttig. Isis, 84(2):366–367, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236253.

Sabra:1993:BRB

[1275] A. I. Sabra. Book review: Religion, Learning, and Science in the^{c} Abbasid Period by M. J. L. Young; J. D. Latham; R. B. Serjeant. Isis, 84(2):367– 368, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236254.

Ivry:1993:BRB

[1276] Alfred L. Ivry. Book review: Ibn Rushd (Averroes) by Dominique Urvoy; Olivia Stewart. Isis, 84(2):368, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236255.

McCluskey:1993:BRB

[1277] Stephen C. McCluskey. Book review: Hexaemeron by Robert Grosseteste; Richard C. Dales; Servus Gieben. Isis, 84(2):368–369, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236256.

Westrem:1993:BRB

[1278] Scott D. Westrem. Book review: Ein Weltbild vor Columbus: Die Ebstorfer Weltkarte: Interdisziplinares Colloquium 1988 by Hartmut Kugler; Eckhard Michael. Isis, 84(2):369–370, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236257.

Lang:1993:BRBa

[1279] Helen S. Lang. Book review: Theology and Science in the Fourteenth Century: Three Questions on the Unity and Subalternation of the Sciences from John of Reading’s Commentary on the Sentences by Steven J. Livesey. Isis, 84(2):370–371, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236258.

Funkenstein:1993:BRB

[1280] Amos Funkenstein. Book review: Contemporary Philosophy: A New Survey. Volume 6: Philosophy and Science in the Middle Ages: Parts 1 and 2 by Guttorm Floistad; Raymond Klibansky. Isis, 84(2):371– 372, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236259.

Schleissner:1993:BRB

[1281] Margaret Schleissner. Book review: Health, Disease, and Healing in Medieval Culture by Sheila Campbell; Bert Hall; David Klausner. Isis, 84(2):372, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236260.

Moran:1993:BRBa

[1282] Bruce T. Moran. Book review: Revolution and Continuity: Essays in the History and Philosophy of Early Modern Science by Peter Barker; Roger Ariew. Isis, 84(2):372–373, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236261.

Webster:1993:BRB

[1283] Charles Webster. Book review: Der gottliche Magier und die Magierin Natur: Religion, Naturmagie und die Anfange der Naturwissenschaft vom Spatmittelalter bis zur Renaissance by Kurt Goldammer. Isis, 84 (2):373–374, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236262.

Findlen:1993:BRBa

[1284] Paula Findlen. Book review: Conrad Gesner by Alfredo Serrai; Maria Cochetti; Marco Menato. Isis, 84(2):374–375, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236263.

Ranea:1993:BRB

[1285] Alberto Guillermo Ranea. Book review: Boehme: An Intellectual Biography of the Seventeenth-Century Philosopher and Mystic by Andrew Weeks. Isis, 84(2):375–376, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236264.

Long:1993:BRB

[1286] Pamela O. Long. Book review: Patronage and Institutions: Science, Technology, and Medicine at the European Court, 1500–1750 by Bruce T. Moran. Isis, 84(2):376–377, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236265.

Dear:1993:BRBb

[1287] Peter Dear. Book review: Descartes’s Metaphysical Physics by Daniel Garber. Isis, 84(2):377–378, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236266.

Baldwin:1993:BRB

[1288] Martha Baldwin. Book review: Descartes e lo sperimentalismo francese (1600–1650) by Alberto Pala. Isis, 84(2):378–379, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236267.

Souffrin:1993:BRB

[1289] Pierre Souffrin. Book review: Exploring the Limits of Preclassical Mechanics: A Study of Conceptual Development in Early Modern Science: Free Fall and Compound Motion in the Work of Descartes, Galileo, and Beeckman by Peter Damerow; Gideon Freudenthal; Peter McLaughlin; Jurgen Renn. Isis, 84(2):379–380, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236268.

Meyer:1993:BRBa

[1290] Eric R. Meyer. Book review: Newton’s Physics and the Conceptual Structure of the Scientific Revolution by Zev Bechler. Isis, 84(2):380–381, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236269.

Westfall:1993:BRB

[1291] Richard S. Westfall. Book review: Sir Isaac Newton: A Catalogue of Manuscripts and Papers Collected and Published on Microfilm by Chadwyck-Healey by Peter Jones. Isis, 84(2):381, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236270.

Swijtink:1993:BRB

[1292] Zeno G. Swijtink. Book review: Instrumente der Neuzeit: Die Entdeckung der modernen Wirklichkeit by Engelhard Weigl. Isis, 84(2):381–382, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236271.

Mazzaoui:1993:BRB

[1293] Maureen F. Mazzaoui. Book review: The Weaver’s Art Revealed: Facsimile, Translation, and Study of the First Two Published Books on Weaving by Patricia Hilts. Isis, 84(2):382–383, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236272.

Canizares:1993:BRB

[1294] Jorge Canizares. Book review: Native Society and Disease in Colonial Ecuador by Suzanne Austin Alchon. Isis, 84(2):383–384, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236273.

Bono:1993:BRBa

[1295] James J. Bono. Book review: The Material Word: Literate Culture in the Restoration and Early Eighteenth Century by Richard W. F. Kroll. Isis, 84(2):384–386, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236274.

Fraser:1993:BRB

[1296] Craig Fraser. Book review: La naissance de la mecanique analytique: La science du mouvement au tournant des XVIIe et XVIIIe siècles by Michel Blay. Isis, 84(2):386–387, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236275.

Jacob:1993:BRBa

[1297] Margaret C. Jacob. Book review: Secularisering en natuurwetenschap in de zeventiende en achttiende eeuw: Bernard Nieuwentijt by Rienk H. Vermij. Isis, 84(2):387–388, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236276.

Wilson:1993:BRB

[1298] Curtis Wilson. Book review: Lacaille: Astronomer, Traveler: With a New Translation of His Journal by David S. Evans. Isis, 84(2):388– 389, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236277.

Guerrini:1993:BRB

[1299] Anita Guerrini. Book review: Jonathan Swift and the Millennium of Madness: The Information Age in Swift’s a Tale of a Tub by Kenneth Craven. Isis, 84(2):389, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236278.

Hoppe:1993:BRB

[1300] Brigitte Hoppe. Book review: Georg von Welling (1655–1727): Studien zu Leben und Werk by Petra Jungmayr. Isis, 84(2):390, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236279.

Wolfe:1993:BRB

[1301] Thomas J. Wolfe. Book review: A Midwife’s Tale: The Life of Martha Ballard, Based on Her Diary, 1785–1812 by Laurel Thatcher Ulrich. Isis, 84(2):390–391, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236280.

Laudan:1993:BRB

[1302] Rachel Laudan. Book review: The Invention of Progress: The Victorians and the Past by Peter J. Bowler. Isis, 84(2):391–392, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236281.

Wolff:1993:BRB

[1303] Stefan L. Wolff. Book review: Mathematik und Bildung in der Humboldtschen Reform by Hans Niels Jahnke. Isis, 84(2):392–393, June 1993.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236282.

Montgomery:1993:BRBb

[1304] William Montgomery. Book review: “Einsamkeit und Freiheit” neu besichtigt: Universitätsreformen und Disziplinenbildung in Preussen als Modell fur Wissenschaftspolitik im Europa des 19. Jahrhunderts by Gert Schubring. Isis, 84(2):393–394, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236283.

McNally:1993:BRB

[1305] Donald H. McNally. Book review: William Whewell, Philosopher of Science by Menachem Fisch. Isis, 84(2):394–395, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236284.

Buchwald:1993:BRB

[1306] Jed Z. Buchwald. Book review: Innovation in Maxwell’s Electromagnetic Theory: Molecular Vortices, Displacement Current, and Light by Daniel M. Siegel. Isis, 84(2):395–396, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236285.

Porter:1993:BRB

[1307] Theodore M. Porter. Book review: Moyenne, milieu, centre: Histoires et usages by Jacqueline Feldman; Gerard Lagneau; Benjamin Matalon. Isis, 84(2):396–397, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236286.

Amsterdamska:1993:BRB

[1308] Olga Amsterdamska. Book review: The Beginnings of Semantics: Essays, Lectures, and Reviews by Michel Breal; George Wolf. Isis, 84(2): 397–398, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236287.

Pang:1993:BRB

[1309] Alex Soojung-Kim Pang. Book review: Science and Social Science Research in British India, 1780–1880: The Role of Anglo-Indian Associations and Government by Edward W. Ellsworth; Scientific and Technical Education in India, 1781–1900 by S. N. Sen. Isis, 84(2):398–399, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236288.

Evans:1993:BRBb

[1310] Hughes Evans. Book review: To the Ends of the Earth: Women’s Search for Education in Medicine by Thomas Neville Bonner. Isis, 84(2):399– 400, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236289.

Ramirez:1993:BRB

[1311] Santiago Ramirez and Ana Barahona. Book review: Pio del Rio Hortega by Jose M. Lopez Piñero; Jose Echegaray by Jose Manuel Sánchez Ron. Isis, 84(2):401–402, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236290.

Shortland:1993:BRB

[1312] Michael Shortland. Book review: La science pour tous: Sur la vulgarisation scientifique en France de 1850 a 1914 by Bruno Beguet. Isis, 84(2): 402–403, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236291.

Beyler:1993:BRB

[1313] Richard H. Beyler. Book review: Der Einstein-Turm: Erwin F. Freundlich und die Relativitatstheorie: Ansatze zu einer “dichten Beschreibung” von institutionellen, biographischen und theoriengeschichtlichen Aspekten by Klaus Hentschel. Isis, 84(2):403–404, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236292.

Pyenson:1993:BRK

[1314] Lewis Pyenson. Book review: Klaus Hentschel: Interpretationen und Fehlinterpretationen der speziellen und der allgemeinen Relativitatstheorie durch Zeitgenossen Albert Einsteins. Isis, 84(2):404–405, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236293.

Stamp:1993:BRB

[1315] David Stamp. Book review: German Science: Some Reflections on German Science: German Science and German Virtues by Pierre Duhem; John Lyon; Pierre Duhem: Philosophy and History in the Work of a Believing Physicist by R. N. D. Martin. Isis, 84(2):405–406, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236294.

Barkan:1993:BRB

[1316] Diana Barkan. Book review: Die Zeitschrift fur Physikalische Chemie: Hundert Jahre Wechselwirkung zwischen Fachwissenschaft, Kommunikationsmedium und Gesellschaft by Thomas Hapke. Isis, 84(2):406–407, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236295.

Bud:1993:BRB

[1317] Robert Bud. Book review: A History of the International Chemical Industry by Fred Aftalion; Otto Theodor Benfey. Isis, 84(2):407–408, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236296.

Winsor:1993:BRB

[1318] Mary P. Winsor. Book review: The Expansion of American Biology by Keith R. Benson; Jane Maienschein; Ronald Rainger. Isis, 84(2): 408, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236297.

Mahoney:1993:BRBb

[1319] Michael S. Mahoney. Book review: John von Neumann and the Origins of Modern Computing by William Aspray. Isis, 84(2):408–409, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236298.

Vanast:1993:BRB

[1320] Walter J. Vanast. Book review: Foundations of the Neuron Doctrine by Gordon M. Shepherd. Isis, 84(2):409–410, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236299.

Decker:1993:BRB

[1321] Hannah S. Decker. Book review: The Diary of Sigmund Freud, 1929– 1939: A Record of the Final Decade by Sigmund Freud; Michael Molnar. Isis, 84(2):410–411, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236300.

Ash:1993:BRB

[1322] Mitchell G. Ash. Book review: Zur Bedeutung des Begriffs der Alltagspsychologie in Theorie und Geschichte der Psychologie: Eine psychologiegeschichtliche Studie anhand der Krise der Psychologie in der Weimarer Republik by Helmut Hildebrandt. Isis, 84(2):411–412, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236301.

Furumoto:1993:BRB

[1323] Laurel Furumoto. Book review: Harry Kirke Wolfe: Pioneer in Psychology by Ludy T. Benjamin, Jr.; Leta Stetter Hollingworth: A Biography by Harry L. Hollingworth. Isis, 84(2):412–413, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236302.

Blustein:1993:BRB

[1324] Bonnie Ellen Blustein. Book review: Fits and Starts: A Genealogy of Hysteria in Modern France by Martha Noel Evans; Madness and Social Representations: Living with the Mad in One French Community by Denise Jodelet; Gerard Duveen; Tim Pownall. Isis, 84(2):413–415, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236303.

Beardsley:1993:BRB

[1325] E. H. Beardsley. Book review: Voegelin on the Idea of Race: An Analysis of Modern European Racism by Thomas W. Heilke. Isis, 84(2):415– 416, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236304.

Meade:1993:BRB

[1326] Teresa Meade. Book review: “The Hour of Eugenics”: Race, Gender, and Nation in Latin America by Nancy Leys Stepan. Isis, 84(2):416, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236305.

Robertson:1993:BRB

[1327] John A. Robertson. Book review: The Surgical Solution: A History of Involuntary Sterilization in the United States by Philip R. Reilly. Isis, 84 (2):417, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236306.

Kohler:1993:BRBa

[1328] Robert E. Kohler. Book review: Big Science: The Growth of Large-Scale Research by Peter Galison; Bruce Hevly. Isis, 84(2):417–419, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236307.

Whittemore:1993:BRB

[1329] Gilbert Whittemore. Book review: Chain Reaction: Expert Debate and Public Participation in American Commercial Nuclear Power, 1945– 1975 by Brian Balogh. Isis, 84(2):419–420, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236308.

Miller:1993:BRB

[1330] David Philip Miller. Book review: The Descent of Icarus: Science and the Transformation of Contemporary Democracy by Yaron Ezrahi; The Fifth Branch: Science Advisers as Policymakers by Sheila Jasanoff. Isis, 84(2):420–422, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236309.

Brotons:1993:BRB

[1331] Victor Navarro Brotons. Book review: Esteban Terradas by Antoni Roca Rosell; Esteban Terradas (1883–1950): Ciencia y tecnica en la España contemporanea by Antoni Roca Rosell; Jose Manuel Sánchez Ron. Isis, 84(2):422–423, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236310.

Creath:1993:BRB

[1332] Richard Creath. Book review: Rediscovering the Forgotten Vienna Circle: Austrian Studies on Otto Neurath and the Vienna Circle by Thomas E. Uebel. Isis, 84(2):423–424, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236311.

Ruse:1993:BRB

[1333] Michael Ruse. Book review: Empiricism and Darwin’s Science by Fred Wilson. Isis, 84(2):424, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236312.

Fantini:1993:BRB

[1334] Bernardino Fantini. Book review: Theory Change in Science: Strategies from Mendelian Genetics by Lindley Darden. Isis, 84(2):424–425, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236313.

McKinney:1993:BRB

[1335] William J. McKinney. Book review: Methodological Variance: Essays in Epistemological Ontology and the Methodology of Science by G. L. Pandit.

Isis, 84(2):425–426, June 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236314.

Akhundov:1993:BRB

[1336] Murad D. Akhundov. Book review: Rekonstruktii a protsessa smeny fundamental’nykh nauchnyck teorii (Reconstruction of the Change of Fundamental Scientific Theories) by R. M. Nugayev. Isis, 84(2):426–427, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236315.

Munday:1993:BRB

[1337] Pat Munday. Book review: Die Nachlasse von Martius, Liebig und den Brudern Schlagintweit in der Bayerischen Staatsbibliothek by Anne Buchler; Rolf Schumacher; Stephan Kellner. Isis, 84(2):427–428, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236316.

Stocking:1993:BRB

[1338] George W. Stocking, Jr. Book review: International Dictionary of Anthropologists by Christopher Winters; Library-Anthropology Resource Group (LARG). Isis, 84(2):428–429, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236317.

Anonymous:1993:Cb

[1339] Anonymous. Collections. Isis, 84(2):429–434, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236318.

Anonymous:1993:BMc

[1340] Anonymous. Back matter. Isis, 84(2):435–440, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236319.

Anonymous:1993:FMc

[1341] Anonymous. Front matter. Isis, 84(2):??, June 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236231.

Findlen:1993:SCE

[1342] Paula Findlen. Science as a career in Enlightenment Italy: The strategies of Laura Bassi. Isis, 84(3):441–469, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235642.

Koerner:1993:GBL

[1343] Lisbet Koerner. Goethe’s botany: Lessons of a feminine science. Isis, 84 (3):470–495, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235643.

Micale:1993:DHS

[1344] Mark S. Micale. On the “disappearance” of hysteria: A study in the clinical deconstruction of a diagnosis. Isis, 84(3):496–526, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235644.

Home:1993:PSI

[1345] R. W. Home and Morris F. Low. Postwar scientific intelligence missions to Japan. Isis, 84(3):527–537, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235645.

Simms:1993:LE

[1346] D. L. Simms, Martin Bernal, Yves Gingras, and Lewis Pyenson. Letters to the Editor. Isis, 84(3):538–541, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235646.

Fuller:1993:BRBb

[1347] Steve Fuller. Book review: Coping with Science. Isis, 84(3):542–547, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235647.

Pingree:1993:BRBa

[1348] David Pingree. Book review: The Crest of the Peacock: Non-European Roots of Mathematics by George Gheverghese Joseph. Isis, 84(3):548– 549, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235648.

McEvoy:1993:BRBb

[1349] John G. McEvoy. Book review: Ideas in Chemistry: A History of the Science by David Knight. Isis, 84(3):549, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235649.

Siegfried:1993:BRB

[1350] Robert Siegfried. Book review: The History of Chemistry by John Hudson. Isis, 84(3):549–550, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235650.

Godlewska:1993:BRB

[1351] Anne Godlewska. Book review: Determinisme et geographie: Herodote, Strabon, Albert le Grand et Sebastien Munster by Jean Bergevin. Isis, 84 (3):550–551, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235651.

McGrath:1993:BRB

[1352] Sylvia W. McGrath. Book review: Women in the Field: America’s Pioneering Women Naturalists by Marcia Myers Bonta. Isis, 84(3):551– 552, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235652.

Mauskopf:1993:BRB

[1353] Seymour H. Mauskopf. Book review: Spirits and Scientists: Ideology, Spiritism, and Brazilian Culture by David J. Hess. Isis, 84(3):552–553, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235653.

Padilla:1993:BRB

[1354] Amado M. Padilla. Book review: Clasicos medicos valencianos del siglo XVI by Jose M. Lopez Piñero; La imagen de ultramar en la medicina valenciana del siglo XIX by Jorge Navarro. Isis, 84(3):553–554, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235654.

Vessuri:1993:BRB

[1355] Hebe Vessuri. Book review: Science, Medicine, and Cultural Imperialism by Teresa Meade; Mark Walker. Isis, 84(3):554–555, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235655.

Hechi:1993:BRB

[1356] Gabrielle Hechi. Book review: Le cheval de Cesar: Ou le mythe des revolutions techniques by Maurice Daumas. Isis, 84(3):555–556, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235656.

Jacob:1993:BRBb

[1357] Christian Jacob. Book review: The Edges of the Earth in Ancient Thought: Geography, Exploration, and Fiction by James S. Romm. Isis, 84(3):556, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235657.

Hahm:1993:BRB

[1358] David E. Hahm. Book review: The Chain of Change: A Study of Aristotle’s Physics VII by Robert Wardy. Isis, 84(3):556–557, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235658.

Riddle:1993:BRB

[1359] John M. Riddle. Book review: De causis plantarum. Volume 2 by Theophrastus; Benedict Einarson; George K. K. Link; De causis plantarum. Volume 3. Isis, 84(3):557–558, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235659.

Pedersen:1993:BRB

[1360] Olaf Pedersen. Book review: The History of Ptolemy’s Star Catalogue by Gerd Grasshoff; Der Sternkatalog des Almagest: Die arabischmittelalterliche Tradition. Volume 2: Die lateinische Übersetzung Gerhards von Cremona by Claudius Ptolemaus; Paul Kunitzsch. Isis, 84 (3):558–560, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235660.

Toomer:1993:BRB

[1361] G. J. Toomer. Book review: Le “Grand Commentaire” de Theon d’Alexandrie aux Tables Faciles de Ptolemee: Livres II et III by Theon of Alexandria; Anne Tihon. Isis, 84(3):560, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235661.

Ebbesen:1993:BRB

[1362] Sten Ebbesen. Book review: Corollaries on Place and Time by Simplicius; J. O. Urmson. Isis, 84(3):560–561, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235662.

Kelter:1993:BRBa

[1363] Irving Kelter. Book review: The Ecology of the Ancient Greek World by Robert Sallares. Isis, 84(3):561–562, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235663.

Knipe:1993:BRB

[1364] David M. Knipe. Book review: Asceticism and Healing in Ancient India: Medicine in the Buddhist Monastery by Kenneth G. Zysk. Isis, 84(3): 562–563, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235664.

Jami:1993:BRB

[1365] Catherine Jami. Book review: Misurare il cielo: L’antica astronomia cinese by Isaia Iannaccone. Isis, 84(3):563–564, September 1993. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235665.

Dales:1993:BRB

[1366] Richard C. Dales. Book review: Glossae super Genesim: Prologus et capitula 1–3 by Petrus Cantor; Agneta Sylwan. Isis, 84(3):564, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235666.

Theisen:1993:BRB

[1367] Wilfred Theisen. Book review: The Medieval Latin Traditions of Euclid’s Catoptrica: A Critical Edition of De speculis with an Introduction, English Translation, and Commentary by Ken’ichi Takahashi. Isis, 84(3): 565, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235667.

Hoenen:1993:BRB

[1368] Maarten J. F. M. Hoenen. Book review: Siger van Brabant: De dubbele waarheid by Henri Krop. Isis, 84(3):565–566, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235668.

Folkerts:1993:BRBa

[1369] Menso Folkerts. Book review: Le Quadripartitum numerorum de Jean de Murs by Ghislaine l’Huillier. Isis, 84(3):566–567, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235669.

Reeds:1993:BRBa

[1370] Karen Reeds. Book review: The Naming of the Beasts: Natural History in the Medieval Bestiary by Wilma George; Brunsdon Yapp. Isis, 84 (3):567–568, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235670.

Bono:1993:BRBb

[1371] James J. Bono. Book review: Witches, Devils, and Doctors in the Renaissance: Johann Weyer, De praestigiis daemonum by Johann Weyer; George Mora; Benjamin Kohl; Erik Middelfort; Helen Bacon; John Shea. Isis, 84(3):568–569, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235671.

Reiss:1993:BRB

[1372] Timothy J. Reiss. Book review: The Machiavellian Cosmos by Anthony J. Parel. Isis, 84(3):569–570, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235672.

Donahue:1993:BRB

[1373] William H. Donahue. Book review: On the Revolutions by Nicholas Copernicus; Edward Rosen; Minor Works by Nicholas Copernicus; Edward Rosen; Erna Hilfstein. Isis, 84(3):570–572, September 1993. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235673.

Chabas:1993:BRB

[1374] Jose Chabas. Book review: Breve discurso a su Majestad el Rey Catolico en torno a la reduccion del ano y reforma del calendario: Con la explicacion de los instrumentos inventados para ensenar su uso en la practica by Juanelo Turriano; Jose Maria Gonzalez Aboin. Isis, 84(3):572–573, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235674.

Maas:1993:BRB

[1375] Jorg F. Maas. Book review: Encyclopaedia [Septem tomis distincta] by Johann Heinrich Alsted; Wilhelm Schmidt-Biggemann. Isis, 84(3): 573–575, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235675.

Moran:1993:BRBb

[1376] Bruce T. Moran. Book review: The French Paracelsians: The Chemical Challenge to Medical and Scientific Tradition in Early Modern France by Allen G. Debus. Isis, 84(3):575–576, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235676.

Garber:1993:BRB

[1377] Daniel Garber. Book review: Descartes and the Dutch: Early Reactions to Cartesian Philosophy, 1637–1650 by Theo Verbeek. Isis, 84(3):576– 577, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235677.

Markley:1993:BRB

[1378] Robert Markley. Book review: Letters and Papers of Robert Boyle: A Guide to the Manuscripts and Microfilm by Michael Hunter; Paul Kesaris; James P. Hoy. Isis, 84(3):577, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235678.

Nadler:1993:BRB

[1379] Steven Nadler. Book review: De summa rerum: Metaphysical Papers, 1675–1676 by G. W. Leibniz; G. H. R. Parkinson. Isis, 84(3):577–578, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235679.

Newman:1993:BRB

[1380] William Newman. Book review: The Janus Faces of Genius: The Role of Alchemy in Newton’s Thought by Betty Jo Teeter Dobbs. Isis, 84 (3):578–579, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235680.

Vickers:1993:BRB

[1381] Brian Vickers. Book review: “Their Manner of Discourse”: Nachdenken über Sprache im Umkreis der Royal Society by Werner Hullen. Isis, 84 (3):579–580, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235681.

Daston:1993:BRB

[1382] Lorraine Daston. Book review: “Beyond Reasonable Doubt” and “Probable Cause”: Historical Perspectives on the Anglo–American Law of Evidence by Barbara J. Shapiro. Isis, 84(3):580–581, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235682.

Meyer:1993:BRBb

[1383] Eric R. Meyer. Book review: Atoms, Pneuma, and Tranquillity: Epicurean and Stoic Themes in European Thought by Margaret J. Osler. Isis, 84(3):581–582, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235683.

Osler:1993:BRB

[1384] Margaret J. Osler. Book review: Cartesian Women: Versions and Subversions of Rational Discourse in the Old Regime by Erica Harth. Isis, 84 (3):582–583, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235684.

Todd:1993:BRBb

[1385] W. David Todd and Peggy Aldrich Kidwell. Book review: Ivory Diptych Sundials, 1570–1750 by Steven A. Lloyd. Isis, 84(3):583–584, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235685.

Meli:1993:BRB

[1386] D. Bertoloni Meli. Book review: De motu and The Analyst: A Modern Edition, with Introductions and Commentary by George Berkeley; Douglas M. Jesseph. Isis, 84(3):584–585, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235686.

Farber:1993:BRBa

[1387] Paul Lawrence Farber. Book review: La Mettrie: Medicine, Philosophy, and Enlightenment by Kathleen Wellman. Isis, 84(3):585, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235687.

Christianson:1993:BRB

[1388] Eric Howard Christianson. Book review: The Angelical Conjunction: The Preacher-Physicians of Colonial New England by Patricia Ann Watson. Isis, 84(3):585–586, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235688.

Servin-Massieu:1993:BRB

[1389] Manuel Servin-Massieu. Book review: Ciencia, tecnica y estado en la España ilustrada by Joaquin Fernandez Perez; Ignacio Gonzalez Tascon.

Isis, 84(3):586–587, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235689.

Slack:1993:BRB

[1390] Nancy Slack. Book review: Petite histoire naturelle de la premiere moitie du XIXe siècle: Leon Dufour, correspondant de l’Institut (1780–1865) by Pascal Duris; Elvire Diaz. Isis, 84(3):587–588, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235690.

Shteir:1993:BRB

[1391] Ann B. Shteir. Book review: The Botanizers: Amateur Scientists in Nineteenth-Century America by Elizabeth B. Keeney. Isis, 84(3):588– 589, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235691.

Pyenson:1993:BRBb

[1392] Lewis Pyenson. Book review: Zoological Researches in Java, and the Neighbouring Islands by Thomas Horsfield; John Bastin. Isis, 84(3): 589–590, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235692.

Woodward:1993:BRBb

[1393] William R. Woodward. Book review: Fries-Apelt-Schleiden: Verzeichnis der Primar- und Sekundarliteratur 1798–1988 by Thomas Glasmacher; Wissenschaftsphilosophische Schriften by Matthias Jakob Schleiden; Ulrich Charpa. Isis, 84(3):590–591, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235693.

Rupke:1993:BRB

[1394] Nicolaas A. Rupke. Book review: The Hunterian Lectures in Comparative Anatomy, May-June, 1837 by Richard Owen; Phillip Reid Sloan. Isis, 84 (3):591–592, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235694.

Blaisdell:1993:BRB

[1395] Muriel L. Blaisdell. Book review: The Works of Charles Darwin by Charles Darwin; Paul H. Barrett; R. B. Freeman; Nora Barlow. Isis, 84 (3):592–594, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235695.

Desmond:1993:BRB

[1396] Adrian Desmond. Book review: Darwin, Huxley, and the Natural Sciences. Isis, 84(3):594–595, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235696.

Jordan:1993:BRBb

[1397] D. W. Jordan. Book review: The Maxwellians by Bruce J. Hunt. Isis, 84 (3):595–596, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235697.

Rocke:1993:BRB

[1398] A. J. Rocke. Book review: Eilhard Mitscherlich: Baumeister am Fundament der Chemie by Hans-Werner Schutt. Isis, 84(3):596–597, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235698.

Gooday:1993:BRB

[1399] Graeme J. N. Gooday. Book review: Innovation as a Social Process: Elihu Thomson and the Rise of General Electric, 1870–1900 by W. Bernard Carlson. Isis, 84(3):597–598, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235699.

Hilts:1993:BRB

[1400] Victor L. Hilts. Book review: The Origins of American Social Science by Dorothy Ross. Isis, 84(3):598–599, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235700.

Pols:1993:BRB

[1401] Johannes C. Pols. Book review: Crowds, Psychology, and Politics, 1871– 1899 by Jaap van Ginneken. Isis, 84(3):599, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235701.

Haas:1993:BRB

[1402] William J. Haas. Book review: The Study of Change: Chemistry in China, 1840–1949 by James Reardon-Anderson. Isis, 84(3):599–600, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235702.

Sigurdsson:1993:BRB

[1403] Skuli Sigurdsson. Book review: Natur und mathematisches Erkennen: Vorlesungen, gehalten 1919–1920 in Gottingen by David Hilbert; David E. Rowe. Isis, 84(3):600–602, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235703.

Randolph:1993:BRB

[1404] Robert H. Randolph. Book review: Ecocide in the USSR: Health and Nature under Siege by Murray Feshbach; Alfred Friendly; Models of Nature: Ecology, Conservation, and Cultural Revolution in Soviet Russia by Douglas R. Weiner. Isis, 84(3):602–604, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235704.

Mitman:1993:BRB

[1405] Gregg Mitman. Book review: An Entangled Bank: The Origins of Ecosystem Ecology by Joel B. Hagen. Isis, 84(3):604–605, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235705.

VanRiper:1993:BRBa

[1406] A. Bowdoin Van Riper. Book review: William Diller Matthew, Paleontologist: The Splendid Drama Observed by Edwin H. Colbert. Isis, 84(3):605, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235706.

Dyer:1993:BRB

[1407] Davis Dyer. Book review: Manufacturing Knowledge: A History of the Hawthorne Experiments by Richard Gillespie. Isis, 84(3):606, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235707.

Klein:1993:BRB

[1408] Martin J. Klein. Book review: Niels Bohr’s Times: In Physics, Philosophy, and Polity by Abraham Pais. Isis, 84(3):606–607, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235708.

Henriksen:1993:BRBa

[1409] Paul W. Henriksen. Book review: The Los Alamos Primer: The First Lectures on How to Build an Atomic Bomb by Richard Rhodes. Isis, 84 (3):607–608, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235709.

Walker:1993:BRBa

[1410] Mark Walker. Book review: Biologen unter Hitler: Vertreibung, Karrieren, Forschung by Ute Deichmann. Isis, 84(3):608–609, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235710.

Gossin:1993:BRB

[1411] Pamela Gossin. Book review: Mumford, Tate, Eiseley: Watchers in the Night by Gale H. Carrithers. Isis, 84(3):609–610, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235711.

Keating:1993:BRB

[1412] Peter Keating. Book review: Le dernier langage de la médécine: Histoire de l’immunologie de Pasteur au Sida by Anne Marie Moulin; Le clinicien et le chercheur: Des grandes maladies de carence à la médécine moléculaire, 1880–1980 by Christiane Sinding. Isis, 84(3):610–611, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235712.

Armstrong:1993:BRB

[1413] David Armstrong. Book review: From Paralysis to Fatigue: A History of Psychosomatic Illness in the Modern Era by Edward Shorter. Isis, 84 (3):611–612, September 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235713.

Hunt:1993:BRB

[1414] Elizabeth E. Hunt. Book review: Soft Soap, Hard Sell: American Hygiene in an Age of Advertising by Vincent Vinikas. Isis, 84(3):612–613, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235714.

Marrett:1993:BRB

[1415] Cora Bagley Marrett. Book review: Milestones and Millstones: Social Science at the National Science Foundation, 1945–1991 by Otto N. Larsen. Isis, 84(3):613–614, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235715.

Wang:1993:BRBa

[1416] Zuoyue Wang. Book review: American Science Policy since World War II by Bruce L. R. Smith. Isis, 84(3):614–615, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235716.

MacLeod:1993:BRB

[1417] Roy MacLeod. Book review: The Seeds of Time: The Life of Sir Macfarlane Burnet by Christopher Sexton. Isis, 84(3):615–617, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235717.

Hetherington:1993:BRBa

[1418] Norriss S. Hetherington. Book review: The Creation of the Anglo-Australian Observatory by S. C. B. Gascoigne; K. M. Proust; M. O. Robins. Isis, 84(3):617–618, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235718.

Dauben:1993:BRBa

[1419] Joseph W. Dauben. Book review: Frege in Perspective by Joan Weiner. Isis, 84(3):618–619, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235719.

Steiner:1993:BRB

[1420] Mark Steiner. Book review: Lakatos’ Philosophy of Mathematics: A Historical Approach by Teun Koetsier. Isis, 84(3):619–620, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235720.

Kronz:1993:BRB

[1421] Frederick M. Kronz. Book review: Quantum Mechanics: An Empiricist View by Bas C. van Fraassen. Isis, 84(3):620–621, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235721.

Urbach:1993:BRB

[1422] Peter Urbach. Book review: Bayes or Bust? A Critical Examination of Bayesian Confirmation Theory by John Earman. Isis, 84(3):621, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235722.

Ludwig:1993:BRB

[1423] Jan Keith Ludwig. Book review: Reading the Book of Nature: An Introduction to the Philosophy of Science by Peter Kosso. Isis, 84(3):621–622, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235723.

Collins:1993:BRBa

[1424] H. M. Collins. Book review: Sociology of Science: A Sociological Pilgrimage by Michael Mulkay. Isis, 84(3):622–623, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235724.

Shapin:1993:BRB

[1425] Steven Shapin. Book review: The Social Dimensions of Science by Ernan McMullin. Isis, 84(3):623–624, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235725.

Wagner:1993:BRB

[1426] Stephen C. Wagner. Book review: The History of Science: An Annotated Bibliography by Gordon L. Miller. Isis, 84(3):624–625, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235726.

Feingold:1993:BRB

[1427] Mordechai Feingold. Book review: Scientific and Technical Periodicals of the Seventeenth and Eighteenth Centuries: A Guide by David A. Kronick. Isis, 84(3):625, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235727.

Pratt:1993:BRB

[1428] Herbert T. Pratt. Book review: Bibliotheca Tinctoria: Annotated Catalog of the Sidney M. Edelstein Collection in the History of Bleaching, Dyeing, Finishing, and Spot Removing by Moshe Ron. Isis, 84(3):625– 626, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235728.

Anonymous:1993:Cc

[1429] Anonymous. Collections. Isis, 84(3):626–630, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235729.

Anonymous:1993:BMd

[1430] Anonymous. Back matter. Isis, 84(3):631–636, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235730.

Anonymous:1993:FMd

[1431] Anonymous. Front matter. Isis, 84(3):??, September 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235641.

Mitman:1993:CNH

[1432] Gregg Mitman. Cinematic nature: Hollywood technology, popular culture, and the American museum of natural history. Isis, 84(4):637–661, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235103.

Cambrosio:1993:EBP

[1433] Alberto Cambrosio, Daniel Jacobi, and Peter Keating. Ehrlich’s “beautiful pictures” and the controversial beginnings of immunological imagery. Isis, 84(4):662–699, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235104.

Camerini:1993:EBM

[1434] Jane R. Camerini. Evolution, biogeography, and maps: An early history of Wallace’s line. Isis, 84(4):700–727, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235105.

Rossianov:1993:ENJ

[1435] Kirill O. Rossianov. Editing nature: Joseph Stalin and the “new” Soviet biology. Isis, 84(4):728–745, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235106.

Golinski:1993:BRB

[1436] Jan Golinski. Book review: Persuading Science: The Art of Scientific Rhetoric. Isis, 84(4):746–749, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235107.

Apple:1993:SS

[1437] Rima D. Apple and Michael W. Apple. Screening science. Isis, 84(4): 750–754, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235108.

Davis:1993:BRB

[1438] Edward B. Davis. Book review: “The Mechanical Universe”; “The Mechanical Universe and beyond”. Isis, 84(4):755–756, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235109.

Frank:1993:BRB

[1439] Robert G. Frank, Jr. Book review: William Harvey and the Circulation of the Blood. Isis, 84(4):756–758, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235110.

Capshew:1993:BRB

[1440] James H. Capshew. Book review: Frankenstein. Isis, 84(4):758–761, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235111.

Geison:1993:BRB

[1441] Gerald L. Geison. Book review: “Microbes and Men”. Isis, 84(4):761– 762, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235112.

Kaiserfeld:1993:BRB

[1442] Thomas Kaiserfeld. Book review: The Flight of the Eagle. (Original Title: Ingenjor Andrees Luftfard). Isis, 84(4):762–763, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235113.

Numbers:1993:BRB

[1443] Ronald L. Numbers. Book review: Inherit the Wind. Isis, 84(4):763–764, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235114.

Levere:1993:BRB

[1444] Trevor H. Levere. Book review: Coppermine. Isis, 84(4):764–766, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235115.

Pernick:1993:SEF

[1445] Martin S. Pernick. Sex education films. Isis, 84(4):766–768, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235116.

Lovely:1993:BRB

[1446] Robert Lovely. Book review: The Plow That Broke the Plains. Isis, 84 (4):768–769, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235117.

Kline:1993:BRB

[1447] Ronald Kline. Book review: Power and the Land. Isis, 84(4):769–771, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235118.

Lederer:1993:BRBa

[1448] Susan E. Lederer. Book review: Arrowsmith. Isis, 84(4):771–772, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235119.

Rogers:1993:BRBa

[1449] Naomi Rogers. Book review: Sister Kenny. Isis, 84(4):772–774, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235120.

Hetherington:1993:BRBb

[1450] Norriss S. Hetherington. Book review: Worlds in the Sky: Planetary Discovery from Earliest Times through Voyager and Magellan by William Sheehan. Isis, 84(4):775, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235121.

Dhanani:1993:BRB

[1451] Alnoor Dhanani. Book review: Philosophy and Science in the Islamic World by C. A. Qadir. Isis, 84(4):775–777, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235122.

Livingstone:1993:BRB

[1452] David N. Livingstone. Book review: The History of Cartography. Volume 2, Book 1: Cartography in the Traditional Islamic and South Asian Societies by J. B. Harley; David Woodward; Joseph E. Schwartzberg; Gerald R. Tibbetts; Ahmet T. Karamustafa. Isis, 84(4):777–778, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235123.

Good:1993:BRB

[1453] Gregory A. Good. Book review: The Geographical Tradition: Episodes in the History of a Contested Enterprise by David N. Livingstone. Isis, 84 (4):778–779, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235124.

Greene:1993:BRB

[1454] Mott T. Greene. Book review: A History of Geology by Gabriel Gohau; Albert V. Carozzi; Marguerite Carozzi. Isis, 84(4):779–780, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235125.

VanRiper:1993:BRBb

[1455] A. Bowdoin Van Riper. Book review: The Preadamite Theory and the Marriage of Science and Religion by David N. Livingstone. Isis, 84(4): 780–781, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235126.

Pacey:1993:BRB

[1456] Arnold Pacey. Book review: Engineering and the Mind’s Eye by Eugene S. Ferguson. Isis, 84(4):781, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235127.

Cook:1993:BRB

[1457] Harold J. Cook. Book review: A History of Medicine by Lois N. Magner; Medicine in Society: Historical Essays by Andrew Wear. Isis, 84(4):781– 783, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235128.

Rogers:1993:BRBb

[1458] Naomi Rogers. Book review: Pictures of Health: A Photographic History of Health Care in Philadelphia, 1860–1945 by Janet Golden; Charles E. Rosenberg; The Picture of Health: Images of Medicine and Pharmacy from the William H. Helfand Collection by William H. Helfand; Patricia Eckert Boyer; Judith Wechsler; Maurice Rickards; Pharmacy History: A Pictorial Record: Photographs from the Museum of the Royal Pharmaceutical Society of Great Britain by Nigel Tallis; Kate Arnold-Forster. Isis, 84(4):783–785, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235129.

Lederer:1993:BRBb

[1459] Susan E. Lederer. Book review: Animals and Disease: An Introduction to the History of Comparative Medicine by Lise Wilkinson. Isis, 84(4): 785–786, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235130.

Feldberg:1993:BRB

[1460] Georgina Feldberg. Book review: Science, Technology, and Medicine in Canada’s Past: Selections from Scientia Canadensis by Richard A. Jarrell; James P. Hull. Isis, 84(4):786, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235131.

Collins:1993:BRBb

[1461] Randall Collins. Book review: The Sea Island Mathematical Manual: Surveying and Mathematics in Ancient China by Frank J. Swetz. Isis, 84(4):786–787, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235132.

Huber:1993:BRB

[1462] Peter J. Huber. Book review: Saros Cycle Dates and Related Babylonian Astronomical Texts by A. Aaboe; J. P. Britton; J. A. Henderson; O. Neugebauer; A. J. Sachs. Isis, 84(4):787–788, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235133.

Lang:1993:BRBb

[1463] Helen S. Lang. Book review: Warm, Kalt, Flussig und Fest bei Aristoteles: Die Elementarqualitaten in den Zoologischen Schriften by Jochen Althoff. Isis, 84(4):788, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235134.

Folkerts:1993:BRBb

[1464] Menso Folkerts. Book review: The Heritage of Giotto’s Geometry: Art and Science on the Eve of the Scientific Revolution by Samuel Y. Edgerton, Jr. Isis, 84(4):788–789, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235135.

Pingree:1993:BRBb

[1465] David Pingree. Book review: The Summa Perfectionis of Pseudo-Geber: A Critical Edition, Translation, and Study by William R. Newman. Isis, 84(4):789–790, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235136.

Goldstein:1993:BRB

[1466] Bernard R. Goldstein. Book review: La bibliothèque de Gersonide: D’apres son catalogue autographe by Gerard E. Weil; Frederic Chartrain; Anne-Marie Weil-Gueny; Joseph Shatzmiller. Isis, 84(4):790–791, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235137.

Reeds:1993:BRBb

[1467] Karen Reeds. Book review: Gardens and Gardening in Papal Rome by David R. Coffin. Isis, 84(4):791–792, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235138.

Aveni:1993:BRBb

[1468] Anthony F. Aveni. Book review: Inca Cosmology and the Human Body by Constance Classen. Isis, 84(4):792–793, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235139.

Feldhay:1993:BRB

[1469] Rivka Feldhay. Book review: The Shapes of Knowledge from the Renaissance to the Enlightenment by Donald R. Kelley; Richard H. Popkin. Isis, 84(4):793–795, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235140.

Kelter:1993:BRBb

[1470] Irving A. Kelter. Book review: Ciencia y censura: La Inquisicion Española y los libros cientificos en los siglos XVI y XVII by Jose Pardo Tomas. Isis, 84(4):795, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235141.

Tribby:1993:BRB

[1471] Jay Tribby. Book review: L’inventario del mondo: Catalogazione della natura e luoghi del sapere nella prima eta moderna by Giuseppe Olmi.

Isis, 84(4):795–796, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235142.

Pyle:1993:BRB

[1472] Cynthia M. Pyle. Book review: Durer’s Animals by Colin Eisler. Isis, 84 (4):796–797, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235143.

Wallace:1993:BRB

[1473] William A. Wallace. Book review: Gli scritti De motu antiquiora di Galileo Galilei: Il Ms. Gal. 71: Un’analisi storico-critica by Michele Camerota. Isis, 84(4):797, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235144.

Grafton:1993:BRBb

[1474] Anthony Grafton. Book review: Gesammelte Werke. Volume 12: Theologica, Hexenprozess, Tacitus-Ubersetzung, Gedichte by Johannes Kepler; Jurgen Hubner; Helmuth Grossing; Friederike Boockmann; Friedrich Seck. Isis, 84(4):798–799, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235145.

Shea:1993:BRBb

[1475] William R. Shea. Book review: The Olympian Dreams and Youthful Rebellion of Rene Descartes by John R. Cole. Isis, 84(4):799–800, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235146.

Grene:1993:BRB

[1476] Marjorie Grene. Book review: Oeuvres. Volume 2 by Nicholas Malebranche; Genevieve Rodis-Lewis; Malebranche and Ideas by Steven Nadler. Isis, 84(4):800–801, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235147.

Findlen:1993:BRBb

[1477] Paula Findlen. Book review: English Spiders (1678) by Martin Lister; Malcolm Davies; Basil Harley; John Parker. Isis, 84(4):801, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235148.

Shapiro:1993:BRB

[1478] Alan E. Shapiro. Book review: Elements de la Philosophie de Newton by Voltaire; Robert L. Walters; W. H. Barber. Isis, 84(4):802, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235149.

Gascoigne:1993:BRB

[1479] John Gascoigne. Book review: The Rise of Public Science: Rhetoric, Technology, and Natural Philosophy in Newtonian Britain, 1660–1750 by Larry Stewart. Isis, 84(4):802–803, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235150.

Donovan:1993:BRB

[1480] Arthur Donovan. Book review: Science as Public Culture: Chemistry and Enlightenment in Britain, 1760–1820 by Jan Golinski. Isis, 84(4): 803–804, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235151.

Broman:1993:BRB

[1481] Thomas Broman. Book review: Iatromechanische Theorie und arztliche Praxis im Vergleich zur galenistischen Medizin (Friedrich Hoffmann– Pieter van Foreest, Jan van Heurne) by Ingo Wilhelm Muller. Isis, 84 (4):804–805, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235152.

Pickstone:1993:BRB

[1482] John Pickstone. Book review: Medical Education in the Age of Improvement: Edinburgh Students and Apprentices, 1760–1826 by Lisa Rosner. Isis, 84(4):805, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235153.

Burchsted:1993:BRB

[1483] Frederic F. Burchsted. Book review: The Correspondence of John Bartram, 1734–1777 by John Bartram; Edmund Berkeley; Dorothy Smith Berkeley; Thomas Say: New World Naturalist by Patricia Tyson Stroud. Isis, 84(4):805–807, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235154.

Archibald:1993:BRBb

[1484] Thomas Archibald. Book review: La France mathematique: La Societe Mathematique de France (1872–1914) by Helene Gispert; R. Bkouche; C. Gilain; C. Houzel; J.-P. Kahane; M. Zerner. Isis, 84(4):807–808, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235155.

Dauben:1993:BRBb

[1485] Joseph W. Dauben. Book review: Frege: Philosophy of Mathematics by Michael Dummett. Isis, 84(4):808–809, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235156.

Grattan-Guinness:1993:BRB

[1486] I. Grattan-Guinness. Book review: Logique et mathematique chez Bernard Bolzano by Jan Sebestik. Isis, 84(4):809–810, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235157.

Siegel:1993:BRB

[1487] Daniel M. Siegel. Book review: The Scientific Letters and Papers of James Clerk Maxwell. Volume 1: 1846–1862 by James Clerk Maxwell; P. M. Harman. Isis, 84(4):810, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235158.

Hatfield:1993:BRB

[1488] Gary Hatfield. Book review: William Whewell: A Composite Portrait by Menachem Fisch; Simon Schaffer. Isis, 84(4):811, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235159.

Nyhart:1993:BRB

[1489] Lynn K. Nyhart. Book review: The Meaning of Evolution: The Morphological Construction and Ideological Reconstruction of Darwin’s Theory by Robert J. Richards. Isis, 84(4):811–812, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235160.

Adas:1993:BRB

[1490] Michael Adas. Book review: The Savage within: The Social History of British Anthropology, 1885–1945 by Henrika Kuklick. Isis, 84(4):812– 813, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235161.

Gingerich:1993:BRB

[1491] Owen Gingerich. Book review: Out of the Shadows: Herschel, Talbot, and the Invention of Photography by Larry J. Schaaf; The Daguerreotype: Nineteenth-Century Technology and Modern Science by M. Susan Barger; William B. White. Isis, 84(4):814, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235162.

Williams:1993:BRB

[1492] Rosalind Williams. Book review: Breaking Frame: Technology and the Visual Arts in the Nineteenth Century by Julie Wosk. Isis, 84(4):814– 815, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235163.

Tweedale:1993:BRB

[1493] Geoffrey Tweedale. Book review: Nature Incorporated: Industrialization and the Waters of New England by Theodore Steinberg. Isis, 84(4): 815–816, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235164.

Misa:1993:BRB

[1494] Thomas J. Misa. Book review: From Machine Shop to Industrial Laboratory: Telegraphy and the Changing Context of American Invention, 1830–1920 by Paul Israel. Isis, 84(4):816–818, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235165.

Reynolds:1993:BRB

[1495] Terry S. Reynolds. Book review: The Making of a Mining District: Keweenaw Native Copper, 1500–1870 by David J. Krause. Isis, 84(4): 818, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235166.

Limoges:1993:BRB

[1496] Camille Limoges. Book review: Pierre Flourens, 1794–1867, physiologiste et historien des sciences: Sa place dans l’evolution de la physiologie experimentale by Georgette Legee. Isis, 84(4):818–819, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235167.

McGraw:1993:BRBa

[1497] Donald J. McGraw. Book review: In Service to American Pharmacy: The Professional Life of William Procter, Jr. by Gregory J. Higby; The Development of American Pharmacology: John J. Abel and the Shaping of a Discipline by John Parascandola. Isis, 84(4):819–820, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235168.

Shorter:1993:BRB

[1498] Edward Shorter. Book review: La storia della psichiatria: Un secolo di studi in Italia by Patrizia Guarnieri. Isis, 84(4):820–821, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235169.

Kudlick:1993:BRB

[1499] Catherine J. Kudlick. Book review: Mission and Method: The Early Nineteenth-Century French Public Health Movement by Ann F. La Berge. Isis, 84(4):821–822, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235170.

Mehrtens:1993:BRB

[1500] Herbert Mehrtens. Book review: Hilbertprogramm und kritische Philosophie: Das Gottinger Modell interdisziplinarer Zusammenarbeit zwischen Mathematik und Philosophie by Volker Peckhaus. Isis, 84(4):823, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235171.

Walker:1993:BRBb

[1501] Mark Walker. Book review: Hochschulalltag im “Dritten Reich”: Die Hamburger Universität, 1933–1945 by Eckart Krause; Ludwig Huber; Holger Fischer. Isis, 84(4):823–824, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235172.

Henriksen:1993:BRBb

[1502] Paul W. Henriksen. Book review: On the Home Front: The Cold War Legacy of the Hanford Nuclear Site by Michele Stenehjem Gerber; The Atomic Bomb: An Annotated Bibliography by Hans G. Graetzer; Larry M. Browning; On Rims and Ridges: The Los Alamos Area since 1880 by Hal K. Rothman. Isis, 84(4):824–825, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235173.

Wang:1993:BRBb

[1503] Zuoyue Wang. Book review: A Different Sort of Time: The Life of Jerrold R. Zacharias, Scientist, Engineer, Educator by Jack S. Goldstein. Isis, 84(4):825–826, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235174.

Lemaine:1993:BRB

[1504] Gerard Lemaine. Book review: Quality and Quantity: The Quest for Biological Regeneration in Twentieth–Century France by William H. Schneider. Isis, 84(4):826–827, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235175.

Tauber:1993:BRB

[1505] Alfred I. Tauber. Book review: A History of Experimental Virology by Alfred Grafe; Elvira Reckendorf. Isis, 84(4):827–828, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235176.

Kohler:1993:BRBb

[1506] Robert E. Kohler. Book review: The Right Tools for the Job: At Work in Twentieth–Century Life Sciences by Adele E. Clarke; Joan H. Fujimura. Isis, 84(4):828–829, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235177.

McGraw:1993:BRBb

[1507] Donald J. McGraw. Book review: Dirt and Disease: Polio before FDR by Naomi Rogers. Isis, 84(4):829–830, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235178.

Sutphen:1993:BRBb

[1508] Molly Sutphen. Book review: The Colonial Disease: A Social History of Sleeping Sickness in Northern Zaire, 1900–1940 by Maryinez Lyons. Isis, 84(4):830–831, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235179.

Solovey:1993:BRB

[1509] Mark Solovey. Book review: Gunnar Myrdal and America’s Conscience: Social Engineering and Racial Liberalism, 1938–1987 by Walter A. Jackson. Isis, 84(4):831–832, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235180.

Lear:1993:BRB

[1510] Linda J. Lear. Book review: Rachel Carson by Mary A. McCay. Isis, 84 (4):832–833, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235181.

Steneck:1993:BRB

[1511] Nicholas H. Steneck. Book review: Impure Science: Fraud, Compromise, and Political Influence in Scientific Research by Robert Bell; Stealing into Print: Fraud, Plagiarism, and Misconduct in Scientific Publishing by Marcel C. LaFollette. Isis, 84(4):833–834, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235182.

Tang:1993:BRB

[1512] Paul C. L. Tang. Book review: Conceptual Revolutions by Paul Thagard. Isis, 84(4):834–835, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235183.

Finocchiaro:1993:BRB

[1513] Maurice A. Finocchiaro. Book review: Thought Experiments in Science and Philosophy by Tamara Horowitz; Gerald J. Massey; Thought Experiments by Roy A. Sorensen. Isis, 84(4):835–836, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235184.

Holton:1993:BRB

[1514] Gerald Holton. Book review: The Laboratory of the Mind: Thought Experiments in the Natural Sciences by James Robert Brown. Isis, 84 (4):836–838, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235185.

Myers:1993:BRB

[1515] Greg Myers. Book review: Science as Writing by David Locke. Isis, 84(4):838, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235186.

McCloskey:1993:BRB

[1516] Donald N. McCloskey. Book review: Economics: Mathematical Politics or Science of Diminishing Returns? by Alexander Rosenberg. Isis, 84 (4):838–839, December 1993. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235187.

Fujimura:1993:BRB

[1517] Joan H. Fujimura. Book review: Theories of Science in Society by Susan E. Cozzens; Thomas F. Gieryn. Isis, 84(4):839–841, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235188.

Farber:1993:BRBb

[1518] Paul Lawrence Farber. Book review: Georges Cuvier: An Annotated Bibliography of His Published Works by Jean Chandler Smith. Isis, 84 (4):841, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235189.

Lowood:1993:BRB

[1519] Henry Lowood. Book review: Books and Manuscripts of the Bakken by Judith A. Overmier; John Edward Senior. Isis, 84(4):841–842, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235190.

Anonymous:1993:Cd

[1520] Anonymous. Collections. Isis, 84(4):842–843, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235191.

Anonymous:1993:BMe

[1521] Anonymous. Back matter. Isis, 84(4):844–874, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235192.

Anonymous:1993:VI

[1522] Anonymous. Volume information. Isis, 84(4):856–873, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235101.

Anonymous:1993:FMe

[1523] Anonymous. Front matter. Isis, 84(4):??, December 1993. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235102.

Neu:1994:CBH

[1524] John Neu. Current bibliography of the history of science and its cultural influences, 1994. Isis, 85(??):1–300, ???? 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235821.

Anonymous:1994:BMa

[1525] Anonymous. Back matter. Isis, 85(??):??, ???? 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235822.

Anonymous:1994:FMa

[1526] Anonymous. Front matter. Isis, 85(??):i–xvi, ???? 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235820.

Smith:1994:ALM

[1527] Pamela H. Smith. Alchemy as a language of mediation at the Habsburg Court. Isis, 85(1):1–25, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235894.

Appel:1994:PAW

[1528] Toby A. Appel. Physiology in American Women’s colleges: The rise and decline of a female subculture. Isis, 85(1):26–56, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235895.

Walker:1994:AEC

[1529] J. Samuel Walker. The Atomic Energy Commission and the politics of radiation protection, 1967–1971. Isis, 85(1):57–78, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235896.

Westman:1994:TCO

[1530] Robert S. Westman. Two cultures or one?: A second look at Kuhn’s the Copernican revolution. Isis, 85(1):79–115, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235897.

Yoder:1994:BRB

[1531] Joella Yoder. Book review: Allgemeiner, politischer und historischer Briefwechsel, by Gottfried Wilhelm Leibniz; Mathematische Schriften. Volume 1: 1672–1676, Geometrie-Zahlentheorie-Algebra (Part 1) by Gottfried Wilhelm Leibniz; Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. Volume 3: 1680–Juni 1683 by Gottfried Wilhelm Leibniz; Discourse on Metaphysics and Other Essays by G. W. Leibniz; Daniel Garber; Roger Ariew; Gottfried Wilhelm Leibniz im philosophischen Diskurs über Geometrie und Erfahrung by Hartmut Hecht. Supplement: Harzbergbau, 1692–1696. Isis, 85(1):116–119, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235898.

Williams:1994:BRBa

[1532] L. Pearce Williams. Book review: Michael Faraday: Sandemanian and Scientist: A Study of Science and Religion in the Nineteenth Century. Isis, 85(1):120–124, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235899.

Rowe:1994:BRB

[1533] David E. Rowe. Book review: A History of Mathematics: An Introduction by Victor J. Katz. Isis, 85(1):125, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235900.

Cantor:1994:BRBa

[1534] Geoffrey Cantor. Book review: Hidden Attraction: The History and Mystery of Magnetism by Gerrit L. Verschuur. Isis, 85(1):125–126, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235901.

Leegwater:1994:BRB

[1535] Arie Leegwater. Book review: De geschiedenis van de scheikunde in Nederland: Van alchemie tot chemie en chemische industrie rond 1900 by H. A. M. Snelders. Isis, 85(1):126–127, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235902.

Davis:1994:BRB

[1536] Edward B. Davis. Book review: Creation, Nature, and Political Order in the Philosophy of Michael Foster (1903–1959): The Classic Mind Articles and Others, with Modern Critical Essays by Cameron Wybrow; The Bible, Baconianism, and Mastery over Nature: The Old Testament and Its Modern Misreading by Cameron Wybrow. Isis, 85(1):127–129, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235903.

Camerini:1994:BRB

[1537] Jane R. Camerini. Book review: The Power of Maps by Denis Wood; John Fels. Isis, 85(1):129–130, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235904.

Lear:1994:BRB

[1538] Linda J. Lear. Book review: Made from This Earth: American Women and Nature by Vera Norwood. Isis, 85(1):130–131, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235905.

Kuriyama:1994:BRB

[1539] Shigehisa Kuriyama. Book review: The Fabric of the Body: European Traditions of Anatomical Illustration by K. B. Roberts; J. D. W. Tomlinson. Isis, 85(1):131–132, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235906.

Daston:1994:BRB

[1540] Lorraine Daston. Book review: Monstrous Imagination by Marie-Helene Huet. Isis, 85(1):132, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235907.

Morman:1994:BRB

[1541] Edward T. Morman. Book review: Race in North America: Origin and Evolution of a Worldview by Audrey Smedley. Isis, 85(1):132–133, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235908.

Moore:1994:BRBa

[1542] James Moore. Book review: Science, Religion, and Mormon Cosmology by Erich Robert Paul. Isis, 85(1):133–134, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235909.

Chambers:1994:BRB

[1543] David Wade Chambers. Book review: Science and Empire: Essays in Indian Context (1700–1947) by Deepak Kumar; Science, Technology, and Colonisation: An Indian Experience, 1757–1857 by Satpal Sangwan. Isis, 85(1):134–136, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235910.

Swade:1994:BRB

[1544] Doron Swade. Book review: The Calculating Machines (Die Rechenmaschinen): Their History and Development by Ernst Martin; Peggy Aldrich Kidwell; Michael R. Williams. Isis, 85(1):136–137, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235911.

Schulman:1994:BRB

[1545] Adam L. Schulman. Book review: The Continuous and the Discrete: Ancient Physical Theories from a Contemporary Perspective by Michael J. White. Isis, 85(1):137–138, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235912.

Hankinson:1994:BRB

[1546] R. J. Hankinson. Book review: La Physique d’Aristote et les conditions d’une science de la nature by François de Gandt; Pierre Souffrin. Isis, 85(1):138–139, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235913.

Ebbesen:1994:BRB

[1547] Sten Ebbesen. Book review: Principles and Proofs: Aristotle’s Theory of Demonstrative Science by Richard D. McKirahan. Isis, 85(1):139– 140, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235914.

Bowen:1994:BRBa

[1548] Alan C. Bowen. Book review: Ptolemy’s Universe: The Natural Philosophical and Ethical Foundations of Ptolemy’s Astronomy by Liba Chaia Taub. Isis, 85(1):140–141, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235915.

Sollenberger:1994:BRB

[1549] Michael G. Sollenberger. Book review: The Archaeology of Economic Ideas: The Classical Greek Tradition by S. Todd Lowry; Protagoras and Logos: A Study in Greek Philosophy and Rhetoric by Edward Schiappa. Isis, 85(1):141–145, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235916.

Wians:1994:BRBa

[1550] William Wians. Book review: His Psychological, Doxographical, and Scientific Writings by Theophrastus; William W. Fortenbaugh; Dimitri Gutas. Isis, 85(1):145, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235917.

Ragep:1994:BRBa

[1551] F. Jamil Ragep. Book review: Science in the Medieval World: “Book of the Categories of Nations” by Said al-Andalusi; Semaan I. Salem; Alok Kumar. Isis, 85(1):145–146, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235918.

Goddu:1994:BRB

[1552] Andre Goddu. Book review: Aristotle’s Physics and Its Medieval Varieties by Helen S. Lang. Isis, 85(1):146–147, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235919.

Eardley:1994:BRB

[1553] Steve Eardley. Book review: Music in Renaissance Magic: Toward a Historiography of Others by Gary Tomlinson. Isis, 85(1):147–148, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235920.

Nutton:1994:BRB

[1554] Vivian Nutton. Book review: Bartholomaeus Anglicus and His Encyclopedia by M. C. Seymour. Isis, 85(1):148–149, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235921.

Field:1994:BRB

[1555] J. V. Field. Book review: Piero della Francesca by Carlo Bertelli; Edward Farelly. Isis, 85(1):149–150, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235922.

Findlen:1994:BRBa

[1556] Paula Findlen. Book review: La biblioteca di Nicolo Leoniceno: Tra Aristotele e Galeno: Cultura e libri di un medico umanista by Daniela Mugnai Carrara. Isis, 85(1):150–151, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235923.

Getz:1994:BRB

[1557] Faye M. Getz. Book review: The Mirror of Alchimy: Composed by the Thrice-Famous and Learned Fryer, Roger Bachon by Roger Bacon; Stanton J. Linden. Isis, 85(1):151, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235924.

Grosholz:1994:BRB

[1558] Emily R. Grosholz. Book review: The Cambridge Companion to Descartes by John Cottingham. Isis, 85(1):151–152, March 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235925.

Joy:1994:BRB

[1559] Lynn S. Joy. Book review: Abrégé de la philosophie de Gassendi by François Bernier; Sylvia Murr; Genevieve Stefani; Bernier et les gassendistes by Sylvia Murr. Isis, 85(1):152–154, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235926.

Gosselin:1994:BRB

[1560] Edward A. Gosselin. Book review: La fortuna del pensiero di Giordano Bruno, 1600–1750 by Saverio Ricci. Isis, 85(1):154–155, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235927.

Goodman:1994:BRBa

[1561] David Goodman. Book review: El codice Pomar (ca. 1590), el interes de Felipe II por la historia natural y la expedicion Hernandez a America by Jose M. Lopez Piñero. Isis, 85(1):155, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235928.

ONeill:1994:BRB

[1562] Eileen O’Neill. Book review: Queen Christina of Sweden and Her Circle: The Transformation of a Seventeenth-Century Philosophical Libertine by Susanna Akerman. Isis, 85(1):155–156, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235929.

Apt:1994:BRB

[1563] Adam Jared Apt. Book review: Nicholas Culpeper: English Physician and Astrologer by Olav Thulesius. Isis, 85(1):156–157, March 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235930.

Anderson:1994:BRB

[1564] Nancy Anderson. Book review: Women and Medicine in the French Enlightenment: The Debate over Maladies des Femmes by Lindsay Wilson. Isis, 85(1):157–158, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235931.

Macgill:1994:BRB

[1565] Elizabeth Rawson Macgill. Book review: Medicine and Shakespeare in the English Renaissance by F. David Hoeniger. Isis, 85(1):158, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235932.

Moran:1994:BRBa

[1566] Bruce T. Moran. Book review: A Checklist of the Newberry Library’s Printed Books in Science, Medicine, Technology, and the Pseudosciences ca. 1460–1750 by Jean S. Gottlieb. Isis, 85(1):158–159, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235933.

Disalle:1994:BRB

[1567] Robert Disalle. Book review: Kant and the Exact Sciences by Michael Friedman. Isis, 85(1):159–160, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235934.

Ramsey:1994:BRB

[1568] Matthew Ramsey. Book review: Tissot: Physician of the Enlightenment by Antoinette Emch-Deriaz. Isis, 85(1):160–161, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235935.

Golinski:1994:BRBa

[1569] Jan Golinski. Book review: Humphry Davy: Science and Power by David Knight. Isis, 85(1):161, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235936.

Gayon:1994:BRB

[1570] Jean Gayon. Book review: Darwin, darwinisme, evolutionnisme by Daniel Becquemont. Isis, 85(1):161–162, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235937.

Kelter:1994:BRB

[1571] Irving A. Kelter. Book review: La science catholique: L’“Encyclopédie théologique” de Migne (1844–1873) entre apologétique et vulgarisation by Claude Langlois; François Laplanche. Isis, 85(1):162–163, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235938.

Turner:1994:BRB

[1572] R. Steven Turner. Book review: Universität und Ministerium im Vormarz: Justus Liebigs Briefwechsel mit Justin von Linde by Justus Liebig; Justin von Linde; Eva-Marie Felschow; Emil Heuser. Isis, 85(1):163–164, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235939.

Slotten:1994:BRBa

[1573] Hugh Richard Slotten. Book review: The Papers of Joseph Henry. Volume 6: January 1844–December 1846: The Princeton Years by Joseph Henry; Marc Rothenberg; Kathleen W. Dorman; John C. Rumm; Paul H. Theerman. Isis, 85(1):164–165, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235940.

Keeney:1994:BRB

[1574] Elizabeth B. Keeney. Book review: Science with Practice: Charles E. Bessey and the Maturing of American Botany by Richard A. Overfield. Isis, 85(1):165, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235941.

Porter:1994:BRB

[1575] Roy Porter. Book review: Jane Austen and the Body: “The Picture of Health” by John Wiltshire. Isis, 85(1):165–166, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235942.

Cantor:1994:BRBb

[1576] David Cantor. Book review: Vital Signs: Medical Realism in Nineteenth-Century Fiction by Lawrence Rothfield. Isis, 85(1):166–167, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235943.

Larson:1994:BRBa

[1577] Edward J. Larson. Book review: Doctors and the Law: Medical Jurisprudence in Nineteenth-Century America by James C. Mohr. Isis, 85(1):167–168, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235944.

Warner:1994:BRB

[1578] John Harley Warner. Book review: Preserve Your Love for Science: Life of William A. Hammond, American Neurologist by Bonnie Ellen Blustein. Isis, 85(1):168–169, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235945.

Dewsbury:1994:BRB

[1579] Donald A. Dewsbury. Book review: Inhibition: History and Meaning in the Sciences of Mind and Brain by Roger Smith. Isis, 85(1):169– 170, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235946.

Fox:1994:BRB

[1580] Robert Fox. Book review: Histoire générale de l’électricité en France. Volume 1: Espoirs et conquêtes, 1881–1918 by François Caron; Fabienne Cardot. Isis, 85(1):170–172, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235947.

Badash:1994:BRW

[1581] Lawrence Badash. Book review: William Lanouette and Bela Silard: Genius in the Shadows: A Biography of Leo Szilard, the Man behind the Bomb. Isis, 85(1):172–173, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235948.

Launius:1994:BRR

[1582] Roger D. Launius. Book review: Robert A. Divine: The Sputnik Challenge: Eisenhower’s Response to the Soviet Satellite. Isis, 85(1):173–174, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235949.

Balzer:1994:BRB

[1583] Harley Balzer. Book review: Science Policy in the Soviet Union by Stephen Fortescue. Isis, 85(1):174, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235950.

Stine:1994:BRB

[1584] Jeffrey K. Stine. Book review: Worldwide Science and Technology Advice to the Highest Levels of Governments by William T. Golden; Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI by Gregg Herken; The Advisers: Scientists in the Policy Process by Bruce L. R. Smith. Isis, 85(1):174–177, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235951.

Lowen:1994:BRB

[1585] Rebecca S. Lowen. Book review: Research and Relevant Knowledge: American Research Universities since World War II by Roger L. Geiger. Isis, 85(1):177, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235952.

Smith:1994:BRBa

[1586] Robert W. Smith. Book review: Beyond Southern Skies: Radio Astronomy and the Parkes Telescope by Peter Robertson. Isis, 85(1):177–178, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235953.

Ziegler:1994:BRB

[1587] Renatus Ziegler. Book review: Die Entstehung des Tensorkalkuls: Von den Anfangen in der Elastizitatstheorie bis zur Verwendung in der Baustatik by Dieter Herbert. Isis, 85(1):178–179, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235954.

Johnson:1994:BRBa

[1588] Jeffrey A. Johnson. Book review: Geschichte der deutschen Grosschemie: Entwicklung und Einfluss in Staat und Gesellschaft by Walter Teltschik. Isis, 85(1):179, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235955.

Greene:1994:BRBa

[1589] Mott T. Greene. Book review: La favola della terra mobile: La controversia sulla teoria della deriva dei continenti by Marco Segala. Isis, 85(1):179–180, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235956.

Sheets-Pyenson:1994:BRB

[1590] Susan Sheets-Pyenson. Book review: An Agenda for Antiquity: Henry Fairfield Osborn and Vertebrate Paleontology at the American Museum of Natural History, 1890–1935 by Ronald Rainger. Isis, 85(1):180–181, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235957.

Hagen:1994:BRBa

[1591] Joel B. Hagen. Book review: The State of Nature: Ecology, Community, and American Social Thought, 1900–1950 by Gregg Mitman. Isis, 85(1): 181–182, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235958.

OHara:1994:BRB

[1592] Robert J. O’Hara. Book review: Classification, Evolution, and the Nature of Biology by Alec L. Panchen. Isis, 85(1):182–183, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235959.

Kohler:1994:BRB

[1593] Robert E. Kohler. Book review: The Molecular Vision of Life: Caltech, the Rockefeller Foundation, and the Rise of the New Biology by Lily E. Kay. Isis, 85(1):183–184, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235960.

Blustein:1994:BRBa

[1594] Bonnie Ellen Blustein. Book review: The Retreat of Scientific Racism: Changing Concepts of Race in Britain and the United States between the World Wars by Elazar Barkan; Final Solutions: Biology, Prejudice, and Genocide by Richard M. Lerner. Isis, 85(1):184–186, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235961.

Fleming:1994:BRB

[1595] Donald Fleming. Book review: The Creationists by Ronald L. Numbers. Isis, 85(1):186–187, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235962.

Feldberg:1994:BRB

[1596] Georgina Feldberg. Book review: Bargaining for Life: A Social History of Tuberculosis, 1876–1938 by Barbara Bates. Isis, 85(1):187–188, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235963.

Capshew:1994:BRBa

[1597] James H. Capshew. Book review: Psychology Aweigh! A History of Clinical Psychology in the United States Navy, 1900–1988 by Frederick L. McGuire. Isis, 85(1):188–189, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235964.

Nier:1994:BRB

[1598] Keith A. Nier. Book review: Alexanderson: Pioneer in American Electrical Engineering by James E. Brittain. Isis, 85(1):189–190, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235965.

Todd:1994:BRBa

[1599] Edmund N. Todd. Book review: Energiewirtschaft, Automatisierung, Information seit 1914 by Hans-Joachim Braun; Walter Kaiser. Isis, 85 (1):190–191, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235966.

Walker:1994:BRBa

[1600] Mark Walker. Book review: Research and Technology in the Former German Democratic Republic by Raymond Bentley. Isis, 85(1):191, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235967.

Suppe:1994:BRB

[1601] Frederick Suppe. Book review: The Semantic Tradition from Kant to Carnap: To the Vienna Station by J. Alberto Coffa; Linda Wessels. Isis, 85(1):191–192, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235968.

Whyte:1994:BRB

[1602] J. T. Whyte. Book review: Against Relativism: A Philosophical Defense of Method by James F. Harris. Isis, 85(1):192–193, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235969.

Barker:1994:BRB

[1603] Peter Barker. Book review: Reconstructing Scientific Revolutions: Thomas S. Kuhn’s Philosophy of Science by Paul Hoyningen-Huene; Alexander J. Levine. Isis, 85(1):193–195, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235970.

Eisberg:1994:BRBa

[1604] Joann Eisberg. Book review: The Outer Circle: Women in the Scientific Community by Harriet Zuckerman; Jonathan R. Cole; John T. Bruer. Isis, 85(1):195, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235971.

Downes:1994:BRB

[1605] Stephen Downes. Book review: Simulating Science: Heuristics, Mental Models, and Technoscientific Thinking by Michael E. Gorman. Isis, 85 (1):195–197, March 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235972.

Browne:1994:BRB

[1606] Janet Browne. Book review: Keywords in Evolutionary Biology by Evelyn Fox Keller; Elisabeth A. Lloyd. Isis, 85(1):197–198, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235973.

Reeds:1994:BRB

[1607] Karen Reeds. Book review: The Cleveland Herbal, Botanical, and Horticultural Collections: A Descriptive Bibliography of Pre-1830 Works from the Libraries of the Holden Arboretum, the Cleveland Medical Library Association, and the Garden Center of Greater Cleveland by Stanley H. Johnston. Isis, 85(1):198, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235974.

Anonymous:1994:Ca

[1608] Anonymous. Collections. Isis, 85(1):198–200, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235975.

Anonymous:1994:BMb

[1609] Anonymous. Back matter. Isis, 85(1):201–206, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235976.

Anonymous:1994:FMb

[1610] Anonymous. Front matter. Isis, 85(1):i–viii, March 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235893.

Cueto:1994:LSA

[1611] Marcos Cueto. Laboratory styles in Argentine physiology. Isis, 85(2): 228–246, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236489.

Principe:1994:STE

[1612] Lawrence M. Principe. Style and thought of the early Boyle: Discovery of the 1648 manuscript of seraphic love. Isis, 85(2):247–260, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236490.

Rupke:1994:CCG

[1613] Nicolaas A. Rupke. C. c. Gillispie’s Genesis and geology. Isis, 85(2):261– 270, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236491.

Benson:1994:AMH

[1614] Keith R. Benson and Daniel J. Kevles. Annual meeting of the History of Science Society: Santa Fe, 11–14 November 1993. Isis, 85(2):271– 277, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236492.

Weir:1994:LE

[1615] Gary E. Weir. Letters to the Editor. Isis, 85(2):278, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236493.

Jardine:1994:BRB

[1616] Nicholas Jardine. Book review: Galileo, Courtier: The Practice of Science in the Culture of Absolutism. Isis, 85(2):279–283, June 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236494.

Carson:1994:RBS

[1617] Cathryn Carson and Silvan S. Schweber. Recent biographical studies in the physical sciences: David C. Cassidy: Uncertainty: The Life and Science of Werner Heisenberg. Ronald R. Kline: Steinmetz: Engineer and Socialist. Albert E. Moyer: A Scientist’s Voice in American Culture: Simon Newcomb and the Rhetoric of Scientific Method. Marelene F. Rayner-Canham and Geoffrey W. Rayner-Canham: Harriet Brooks: Pioneer Nuclear Scientist. John M. Thomas and David Phillips: Selections and Reflections: The Legacy of Sir Lawrence Bragg. Victor Weisskopf: The Joy of Insight: Passions of a Physicist. Isis, 85(2):284–292, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236495.

Lindee:1994:BRB

[1618] M. Susan Lindee. Book review: Gene Mapping: Using Law and Ethics as Guides. Isis, 85(2):293–296, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236496.

Pols:1994:BRB

[1619] Johannes C. Pols. Book review: Bad Habits: Drinking, Smoking, Taking Drugs, Gambling, Sexual Misbehavior, and Swearing in American History by John C. Burnham. Isis, 85(2):297, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236497.

Landau:1994:BRB

[1620] Misia Landau. Book review: Space, Time, and Man: A Prehistorian’s View by Grahame Clark. Isis, 85(2):297–299, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236498.

Lattis:1994:BRBa

[1621] James Lattis. Book review: Encyclopedia of Cosmology: Historical, Philosophical, and Scientific Foundations of Modern Cosmology by Norriss S. Hetherington. Isis, 85(2):299–300, June 1994. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236499.

Hoyrup:1994:BRB

[1622] Jens Hoyrup. Book review: The Mystery of Numbers by Annemarie Schimmel. Isis, 85(2):300–301, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236500.

Brooke:1994:BRBa

[1623] John Hedley Brooke. Book review: The Fontana History of Chemistry by William H. Brock. Isis, 85(2):301–302, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236501.

Melhado:1994:BRB

[1624] Evan M. Melhado. Book review: The Science of Matter: A Historical Survey by Maurice Crosland. Isis, 85(2):302–303, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236502.

Multhauf:1994:BRB

[1625] R. P. Multhauf. Book review: Salt Production Techniques in Ancient China: The Aobo Tu by Yoshida Tora; Hans Ulrich Vogel. Isis, 85(2): 303, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236503.

Aaserud:1994:BRB

[1626] Finn Aaserud. Book review: Lovers of Learning: A History of the Royal Danish Academy of Sciences and Letters, 1742–1992 by Olaf Pedersen. Isis, 85(2):303–304, June 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236504.

Wians:1994:BRBb

[1627] William Wians. Book review: Natural Knowledge in Preclassical Antiquity by Mott T. Greene. Isis, 85(2):304–305, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236505.

Vallance:1994:BRB

[1628] J. T. Vallance. Book review: Science and Philosophy in Classical Greece by Alan C. Bowen. Isis, 85(2):305–306, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236506.

Smith:1994:BRBb

[1629] Wesley D. Smith. Book review: A Dictionary of Medical Terms in Galen by Richard J. Durling. Isis, 85(2):306–307, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236507.

Mahdi:1994:BRB

[1630] Muhsin Mahdi. Book review: Al-Farabi and His School by Ian Richard Netton. Isis, 85(2):307, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236508.

Pingree:1994:BRB

[1631] David Pingree. Book review: Le calcul Indien (algorismus) by Muhammad Ibn Musa Al-Khwarizmi; Andre Allard. Isis, 85(2):307–308, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236509.

Livesey:1994:BRB

[1632] Steven J. Livesey. Book review: The Intellectual Life of Western Europe in the Middle Ages by Richard C. Dales. Isis, 85(2):308–309, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236510.

Gutwirth:1994:BRB

[1633] Eleazar Gutwirth. Book review: Science médiévale d’Espagne et d’alentour by Guy Beaujouan; Las ciencias de los antiguos en al-Andalus by Julio Samso. Isis, 85(2):309–310, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236511.

Smith:1994:BRBc

[1634] A. Mark Smith. Book review: Witelo: Ein Philosoph und Naturforscher des XIII. Jahrhunderts by Clemens Baeumker; Witelonis Perspectivae: Liber secundus et liber tertius. Books 2 and 3 of Witelo’s Perspectiva by Witelo; Sabetai Unguru. Isis, 85(2):310–311, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236512.

Chekin:1994:BRB

[1635] Leonid S. Chekin. Book review: Erde und Kosmos im Mittelalter: Das Weltbild vor Kolumbus by Rudolf Simek. Isis, 85(2):311–312, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236513.

Schleissner:1994:BRB

[1636] Margaret Schleissner. Book review: Meanings of Sex Difference in the Middle Ages: Medicine, Science, and Culture by Joan Cadden; The Bridling of Desire: Views of Sex in the Later Middle Ages by Pierre J. Payer. Isis, 85(2):312–313, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236514.

Opsomer-Halleux:1994:BRB

[1637] C. Opsomer-Halleux. Book review: Nomina simplicium medicinarum ex synonymariis Medii Aevi collecta: Semantische Untersuchungen zum Fachwortschatz hoch- und spatmittelalterlicher Drogenkunde by Willem F. Daems. Isis, 85(2):313–314, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236515.

Brevart:1994:BRB

[1638] Francis B. Brevart. Book review: Das “Lorscher Arzneibuch”: Ein medizinisches Kompendium des 8. Jahrhunderts (Codex Bambergensis medicinalis 1): Text, Übersetzung und Fachglossar by Ulrich Stoll. Isis, 85(2): 314–315, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236516.

Ruderman:1994:BRB

[1639] David B. Ruderman. Book review: Studies on Gersonides: A Fourteenth-Century Jewish Philosopher-Scientist by Gad Freudenthal. Isis, 85(2): 315, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236517.

Obrist:1994:BRB

[1640] Barbara Obrist. Book review: Die alchemistische Lehrdichtung des Gratheus filius philosophi in Cod. Vind. 2372: Zugleich ein Beitrag zur okkulten Wissenschaft im Spatmittelalter by Helmut Birkhan. Isis, 85(2): 315–317, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236518.

Demaitre:1994:BRBa

[1641] Luke E. Demaitre. Book review: Medieval Prognosis and Astrology: A Working Edition of the Aggregationes de crisi et creticis diebus by Cornelius O’Boyle. Isis, 85(2):317, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236519.

Rosinska:1994:BRBa

[1642] Grazyna Rosinska. Book review: The Speculum astronomiae and Its Enigma: Astrology, Theology, and Science in Albertus Magnus and His Contemporaries by Paola Zambelli. Isis, 85(2):317–318, June 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236520.

Vickers:1994:BRB

[1643] Brian Vickers. Book review: L’ambigua natura della magia: Filosofi, streghe, riti nel Rinascimento by Paola Zambelli. Isis, 85(2):318–320, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236521.

Fanning:1994:BRB

[1644] Steven Fanning. Book review: Medieval Military Technology by Kelly DeVries. Isis, 85(2):320, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236522.

Hall:1994:BRB

[1645] Bert Hall. Book review: The Art of Medieval Technology: Images of Noah the Shipbuilder by Richard W. Unger. Isis, 85(2):320–321, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236523.

Grootendorst:1994:BRB

[1646] A. W. Grootendorst. Book review: Clio’s stiefkind by E. J. Dijksterhuis; K. Van Berkel. Isis, 85(2):321–322, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236524.

Nadler:1994:BRBa

[1647] Steven Nadler. Book review: The Rise of Modern Philosophy: The Tension between the New and Traditional Philosophies from Machiavelli to Leibniz by Tom Sorell. Isis, 85(2):322–323, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236525.

Tolmacheva:1994:BRB

[1648] Marina Tolmacheva. Book review: Terra Cognita: The Mental Discovery of America by Eviatar Zerubavel. Isis, 85(2):323–324, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236526.

Cormack:1994:BRBa

[1649] Lesley B. Cormack. Book review: Monarchs, Ministers, and Maps: The Emergence of Cartography as a Tool of Government in Early Modern Europe by David Buisseret; Cartes des Ameriques: Dans les collections de la BibliothÈQue Royale Albert Ier by Hossam Elkhadem; Jean-Paul Heerbrant; Liliane Wellens-De Donder; Roger Calcoen. Isis, 85(2):324– 325, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236527.

Stephenson:1994:BRB

[1650] Bruce Stephenson. Book review: New Astronomy by Johannes Kepler; William H. Donahue. Isis, 85(2):326–327, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236528.

Smith:1994:BRBd

[1651] Pamela H. Smith. Book review: Mathematik im Studienplan der Jesuiten: Die Vorlesung von Otto Cattenius an der Universität Mainz (1610/11) by Albert Krayer. Isis, 85(2):327–328, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236529.

Rees:1994:BRB

[1652] Graham Rees. Book review: Francis Bacon, the State, and the Reform of Natural Philosophy by Julian Martin. Isis, 85(2):328–329, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236530.

Wallace:1994:BRBa

[1653] William A. Wallace. Book review: Sidereus nuncius: Le messager céleste by Galileo Galilei; Isabelle Pantin; Le messager des étoiles by Galileo Galilei; Fernand Hallyn. Isis, 85(2):329–330, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236531.

deRomo:1994:BRB

[1654] Ana Cecilia Rodriguez de Romo. Book review: El nuevo sol de la medicina en la ciudad de los reyes: Federico Bottoni y la Evidencia de la circulacion de la sangre (Lima, 1723) by Alvar Martinez Vidal. Isis, 85 (2):330, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236532.

Hilfstein:1994:BRB

[1655] Erna Hilfstein. Book review: Znajomosc dziel Jana Amosa Komenskiego na ziemiach czeskich, slowackich i polskich od polowy XVII wieku do czasow obecnych by M. Beckova; T. Bienkowski; D. Capkova. Isis, 85(2): 330–331, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236533.

Brackenridge:1994:BRB

[1656] J. Bruce Brackenridge. Book review: Equivalence and Priority: Newton versus Leibniz: Including Leibniz’s Unpublished Manuscripts on the Principia by Domenico Bertoloni Meli; Action and Reaction: Proceedings of a Symposium to Commemorate the Tercentenary of Newton’s Principia by Paul Theerman; Adele F. Seeff. Isis, 85(2):332–334, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236534.

Erlichson:1994:BRB

[1657] Herman Erlichson. Book review: Newton mecanicien du cosmos by Georges Barthelemy. Isis, 85(2):334–335, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236535.

Sobol:1994:BRBa

[1658] Peter Sobol. Book review: Kritik und Verteidigung des Tierversuchs: Die Anfange der Diskussion im 17. und 18. Jahrhundert by Andreas-Holger Maehle. Isis, 85(2):335–336, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236536.

Larson:1994:BRBb

[1659] James Larson. Book review: Linne et la France (1780–1850) by Pascal Duris. Isis, 85(2):336–337, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236537.

Broman:1994:BRB

[1660] Thomas Broman. Book review: Vom Medizinstudenten zum Doktor: Die Gottinger Medizinischen Promotionen im 18. Jahrhundert by Ulrich Trohler; Sabine Mildner-Mazzei. Isis, 85(2):337–338, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236538.

Beier:1994:BRB

[1661] A. L. Beier. Book review: Patients, Power, and the Poor in Eighteenth-Century Bristol by Mary E. Fissell. Isis, 85(2):338–339, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236539.

Brooks:1994:BRB

[1662] G. P. Brooks. Book review: This Invisible Riot of the Mind: Samuel Johnson’s Psychological Theory by Gloria Sybil Gross. Isis, 85(2):339– 340, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236540.

Pang:1994:BRB

[1663] Alex Soojung-Kim Pang. Book review: Imagining the Pacific: In the Wake of the Cook Voyages by Bernard W. Smith. Isis, 85(2):340–341, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236541.

Gregory:1994:BRB

[1664] Frederick Gregory. Book review: Robert Mayer and the Conservation of Energy by Kenneth L. Caneva. Isis, 85(2):341–342, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236542.

Mitman:1994:BRB

[1665] Gregg Mitman. Book review: Picturing Nature: American Nineteenth-Century Zoological Illustration by Ann Shelby Blum. Isis, 85(2):342– 344, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236543.

Farber:1994:BRB

[1666] Paul Lawrence Farber. Book review: Audubon’s Great National Work: The Royal Octavo Edition of The Birds of America by Ron Tyler. Isis, 85 (2):344, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236544.

Goldstein:1994:BRB

[1667] Daniel Goldstein. Book review: Spencer Baird of the Smithsonian by E. F. Rivinus; E. M. Youssef. Isis, 85(2):344–345, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236545.

Kuznick:1994:BRB

[1668] Peter J. Kuznick. Book review: American Socialists and Evolutionary Thought, 1870–1920 by Mark Pittenger. Isis, 85(2):345–346, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236546.

Richards:1994:BRBa

[1669] Robert J. Richards. Book review: The Language of Science: A Study of the Relationship between Literature and Science in the Perspective of a Hermeneutical Ontology, with a Case Study of Darwin’s The Origin of Species by Ilse N. Bulhof. Isis, 85(2):346–347, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236547.

Osborne:1994:BRB

[1670] Michael A. Osborne. Book review: Civilizing Mission: Exact Sciences and French Overseas Expansion, 1830–1940 by Lewis Pyenson. Isis, 85 (2):347–348, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236548.

Blondel-Megrelis:1994:BRB

[1671] Marika Blondel-Megrelis. Book review: Paul Sabatier, un chimiste independant (1854–1941): Naissance de l’hydrogenation catalytique by Bruno Wojtkowiak. Isis, 85(2):348–349, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236549.

Smocovitis:1994:BRB

[1672] Vassiliki Betty Smocovitis. Book review: The Ant and the Peocock: Altruism and Sexual Selection from Darwin to Today by Helena Cronin. Isis, 85(2):349–351, June 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236550.

Brozek:1994:BRBa

[1673] Josef Brožek. Book review: The Collected Works of L. S. Vygotsky. Volume 1: Problems of General Psychology. Including the Volume Thinking and Speech by L. S. Vygotsky; Robert W. Rieber; Aaron S. Carton; Norris Minick; The Collected Works of L. S. Vygotsky. Volume 2: The Fundamentals of Defectology (Abnormal Psychology and Learning Disabilities) by L. S. Vygotsky; Robert W. Rieber; Aaron S. Carton; Jane E. Knox; Carol B. Stevens; Understanding Vygotsky: A Quest for Synthesis by Rene van der Veer; Jaan Valsiner. Isis, 85(2):351–353, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236551.

Olby:1994:BRB

[1674] Robert Olby. Book review: The Founders of Evolutionary Genetics: A Centenary Reappraisal by Sahotra Sarkar; The Search for the Gene by Bruce Wallace. Isis, 85(2):353–354, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236552.

Richards:1994:BRBb

[1675] Joan L. Richards. Book review: Selected Essays on the History of Set Theory and Logics (1906–1918) by Philip E. B. Jourdain; Ivor Grattan-Guinness. Isis, 85(2):354–355, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236553.

Walker:1994:BRBb

[1676] Mark Walker. Book review: Nationalism and Internationalism in Science, 1880–1939: Four Studies of the Nobel Population by Elisabeth Crawford. Isis, 85(2):355–356, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236554.

Parisi:1994:BRBa

[1677] Thomas Parisi. Book review: The Complete Correspondence of Sigmund Freud and Ernest Jones, 1908–1939 by Sigmund Freud; Ernest Jones; R. Andrew Paskauskas. Isis, 85(2):356–357, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236555.

Capshew:1994:BRBb

[1678] James H. Capshew. Book review: The Professionalization of Psychology in Nazi Germany by Ulfried Geuter; Richard J. Holmes. Isis, 85(2):357– 358, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236556.

Cassidy:1994:BRB

[1679] David Cassidy. Book review: Operation Epsilon: The Farm Hall Transcripts. Isis, 85(2):358–359, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236557.

Duffin:1994:BRB

[1680] Jacalyn Duffin. Book review: The Remarkable Surgical Practice of John Benjamin Murphy by Robert L. Schmitz; Timothy T. Oh. Isis, 85(2): 359, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236558.

Todd:1994:BRBb

[1681] Edmund N. Todd. Book review: Invention: The Care and Feeding of Ideas by Norbert Wiener. Isis, 85(2):359–360, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236559.

Rozwadowski:1994:BRBa

[1682] Helen Rozwadowski. Book review: Challenger at Sea: A Ship That Revolutionized Earth Science by Kenneth J. Hsu. Isis, 85(2):360–361, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236560.

Nickles:1994:BRB

[1683] Thomas Nickles. Book review: Scientific Discovery: Logic and Tinkering by Aharon Kantorovich. Isis, 85(2):361–362, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236561.

Gooding:1994:BRB

[1684] David C. Gooding. Book review: The Rediscovery of the Mind by John R. Searle. Isis, 85(2):362–363, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236562.

Deltete:1994:BRB

[1685] Robert J. Deltete. Book review: Philosophy of Physics by Lawrence Sklar. Isis, 85(2):363–364, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236563.

Forge:1994:BRB

[1686] John Forge. Book review: Quantum Mechanics and Experience by David Z. Albert. Isis, 85(2):364–365, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236564.

Knoespel:1994:BRB

[1687] Kenneth J. Knoespel. Book review: New Worlds, Ancient Texts: The Power of Tradition and the Shock of Discovery by Anthony Grafton; April Shelford; Nancy Siraisi. Isis, 85(2):365–366, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236565.

Gingerich:1994:BRBa

[1688] Owen Gingerich. Book review: Bibliotheca Mechanica by Verne L. Roberts; Ivy Trent; The Haskell F. Norman Library of Science and Medicine by Diana H. Hook; Jeremy M. Norman. Isis, 85(2):366–367, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236566.

Slotten:1994:BRBb

[1689] Hugh R. Slotten. Book review: American Scientific Exploration, 1803– 1860: Manuscripts in Four Philadelphia Libraries by William Stanton. Isis, 85(2):367, June 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236567.

Blustein:1994:BRBb

[1690] Bonnie Ellen Blustein. Book review: Mujer, salud y ciencia (1900– 1991): Fondos bibliograficos sobre mujeres en la Seccion de Historia de la Medicina y de la Ciencia de la Biblioteca de la Universidad de Granada by Teresa Ortiz. Isis, 85(2):367–368, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236568.

Anonymous:1994:Cb

[1691] Anonymous. Collections. Isis, 85(2):368–371, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236569.

Anonymous:1994:BMc

[1692] Anonymous. Back matter. Isis, 85(2):372–376, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236570.

Martin:1994:OTS

[1693] Steven C. Martin. “The only truly scientific method of healing”: Chiropractic and American science, 1895–1990. Isis, 85(2):viii + 207–227, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236488.

Anonymous:1994:FMc

[1694] Anonymous. Front matter. Isis, 85(2):i–vii, June 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236487.

Pruna:1994:NSC

[1695] Pedro M. Pruna. National science in a colonial context: The Royal Academy of Sciences of Havana, 1861–1898. Isis, 85(3):412–426, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235461.

Dennis:1994:OFL

[1696] Michael Aaron Dennis. “Our first line of defense”: Two University laboratories in the postwar American state. Isis, 85(3):427–455, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235462.

Medicus:1994:FAT

[1697] Heinrich A. Medicus. The friendship among three singular men: Einstein and his Swiss friends Besso and Zangger. Isis, 85(3):456–478, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235463.

Hilfstein:1994:LE

[1698] Erna Hilfstein and William H. Donahue. Letters to the Editor. Isis, 85 (3):479–481, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235464.

Gillmor:1994:BRB

[1699] C. Stewart Gillmor. Book review: No Ordinary Journey: John Rae, Arctic Explorer, 1813–1893. Isis, 85(3):482–485, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235465.

Levere:1994:BRB

[1700] Trevor H. Levere. Book review: A History of Western Science by Anthony M. Alioto. Isis, 85(3):486–487, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235466.

Rodriguez:1994:BRB

[1701] Julia Emilia Rodriguez. Book review: Geschichte der Wissenschaften in Lateinamerika by Martin Guntau. Isis, 85(3):487–488, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235467.

Cormack:1994:BRBb

[1702] Lesley B. Cormack. Book review: Flattening the Earth: Two Thousand Years of Map Projections by John P. Snyder; Two by Two: Twenty-Two Pairs of Maps from the Newberry Library Illustrating Five Hundred Years of Western Cartographic History by James Akerman; Robert Karrow; David Buisseret. Isis, 85(3):488–489, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235468.

Zhu:1994:BRB

[1703] Yuelin Zhu. Book review: Experiment, Right or Wrong by Allan Franklin; Experimental Inquiries: Historical, Philosophical, and Social Studies of Experimentation in Science by Homer E. Le Grand. Isis, 85(3):489–491, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235469.

Misa:1994:BRB

[1704] Thomas J. Misa. Book review: Discovery, Innovation, and Risk: Case Studies in Science and Technology by Newton Copp; Andrew Zanella. Isis, 85(3):491–493, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235470.

Dorries:1994:BRB

[1705] Matthias Dorries. Book review: Technik und Medizin by Rolf Winau. Isis, 85(3):493, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235471.

Abir-Am:1994:BRB

[1706] Pnina Abir-Am. Book review: A Documentary History of Biochemistry 1770–1940 by Mikulas Teich; Dorothy M. Needham. Isis, 85(3):493–495, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235472.

Carmichael:1994:BRB

[1707] Ann G. Carmichael. Book review: In Time of Plague: The History and Social Consequences of Lethal Epidemic Disease by Arien Mack. Isis, 85 (3):495–496, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235473.

Brozek:1994:BRBb

[1708] Josef Brožek. Book review: Mental Machinery: The Origins and Consequences of Psychological Ideas. Part 1: 1600–1850 by Graham Richards. Isis, 85(3):496–497, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235474.

Harkness:1994:BRB

[1709] Jon M. Harkness. Book review: The Ethical Dimensions of the Biological Sciences by Ruth Ellen Bulger; Elizabeth Heitman; Stanley Joel Reiser. Isis, 85(3):497–498, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235475.

Sobol:1994:BRBb

[1710] Peter G. Sobol. Book review: Great Mysteries of the West by Ferenc Morton Szasz. Isis, 85(3):498, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235476.

Shute:1994:BRB

[1711] Charles Shute. Book review: Mathematiques egyptiennes: Recherches sur les connaissances mathematiques de l’Egypte pharaonique by Sylvia Couchoud. Isis, 85(3):498–499, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235477.

Glicksman:1994:BRB

[1712] Maurice Glicksman. Book review: Fleeting Footsteps: Tracing the Conception of Arithmetic and Algebra in Ancient China by Lam Lay Yong; Ang Tian Se. Isis, 85(3):499–500, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235478.

Bowen:1994:BRBb

[1713] Alan C. Bowen. Book review: L’origine et l’évolution du concept grec de phusis by Gerard Naddaf. Isis, 85(3):500–501, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235479.

Taub:1994:BRB

[1714] Liba Taub. Book review: Roman Nature: The Thought of Pliny the Elder by Mary Beagon. Isis, 85(3):501–502, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235480.

Cabre:1994:BRB

[1715] Montserrat Cabre. Book review: Women’s Secrets: A Translation of Pseudo-Albertus Magnus’s De secretis mulierum with Commentaries by Helen Rodnite Lemay. Isis, 85(3):502–503, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235481.

Demaitre:1994:BRBb

[1716] Luke Demaitre. Book review: Galenus Latinus. Volume 2: Burgundio of Pisa’s Translation of Galen’s ΠEP IT ΩNΠEΠONΘΩNT OΠΩN, “De interioribus” by Galen; Richard J. Durling. Isis, 85(3):503–504, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235482.

Ragep:1994:BRBb

[1717] F. Jamil Ragep. Book review: An Introduction to Islamic Cosmological Doctrines: Conceptions of Nature and Methods Used for Its Study by the Ikhwan al-Safa, al-Biruni, and Ibn Sina by Seyyed Hossein Nasr. Isis, 85 (3):504–505, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235483.

Behrens-Abouseif:1994:BRB

[1718] Doris Behrens-Abouseif. Book review: Qusta Ibn Luqa’s Medical Regime for the Pilgrims to Mecca: The Risala fi tadbir safar al-hajj by Qusta Ibn Luqa; Gerrit Bos. Isis, 85(3):505, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235484.

Molland:1994:BRB

[1719] George Molland. Book review: Philosophy and Theology in the Middle Ages by G. R. Evans. Isis, 85(3):505–506, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235485.

Lindberg:1994:BRB

[1720] David C. Lindberg. Book review: Questions concerning the Eternity of the World by John Pecham; Vincent G. Potter. Isis, 85(3):506–507, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235486.

Figala:1994:BRB

[1721] Karin Figala and Ulrich Petzold. Book review: Rosarium philosophorum: Ein alchemisches Florilegium des Spatmittelalters by Joachim Telle; Lutz Claren; Joachim Huber. Isis, 85(3):507, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235487.

Wallace:1994:BRBb

[1722] William A. Wallace. Book review: Le matematiche e il mondo: Ricerche su un dibattito in Italia nella seconda meta del Cinquecento by Anna De Pace. Isis, 85(3):507–508, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235488.

Goodman:1994:BRBb

[1723] David Goodman. Book review: Aspectos de la ciencia aplicada en la España del Siglo de Oro by M. I. Vicente Maroto; M. Esteban Pineiro. Isis, 85(3):508–509, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235489.

Smith:1994:BRBe

[1724] Pamela H. Smith. Book review: Graf Wolfgang II von Hohenlohe und die Alchemie: Alchemistische Studien in Schloss Weikersheim, 1587– 1610 by Jost Weyer. Isis, 85(3):509–510, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235490.

Moran:1994:BRBb

[1725] Bruce T. Moran. Book review: Medicine and the Reformation by Ole Peter Grell; Andrew Cunningham. Isis, 85(3):510–511, September 1994.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235491.

Findlen:1994:BRBb

[1726] Paula Findlen. Book review: Signs and Portents: Monstrous Births from the Middle Ages to the Enlightenment by Dudley Wilson. Isis, 85(3):511, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235492.

Jacob:1994:BRB

[1727] Margaret C. Jacob. Book review: The Scientific Revolution in National Context by Roy Porter; Mikulas Teich. Isis, 85(3):511–513, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235493.

Voelkel:1994:BRB

[1728] James R. Voelkel. Book review: Dissertatio cum nuncio sidereo; Narratio de observatis Jovis satellitibus: Discussion avec le messager celeste; Rapport sur l’observation des satellites de Jupiter by Johannes Kepler; Isabelle Pantin. Isis, 85(3):513, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235494.

Pereira:1994:BRB

[1729] Michela Pereira. Book review: La rationalite de l’alchimie au XVIIe siècle by Bernard Joly. Isis, 85(3):513–514, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235495.

Gaukroger:1994:BRB

[1730] Stephen Gaukroger. Book review: Causation in Early Modern Philosophy: Cartesianism, Occasionalism, and Preestablished Harmony by Steven Nadler. Isis, 85(3):514–515, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235496.

Dobbs:1994:BRB

[1731] B. J. T. Dobbs. Book review: Isaac Newton: Adventurer in Thought by A. Rupert Hall; The Life of Isaac Newton by Richard S. Westfall. Isis, 85 (3):515–517, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235497.

Park:1994:BRB

[1732] Katharine Park. Book review: Miasmas and Disease: Public Health and the Environment in the Pre-Industrial Age by Carlo M. Cipolla; Elizabeth Potter. Isis, 85(3):517, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235498.

Grell:1994:BRB

[1733] O. P. Grell. Book review: Stadsbelang en standsbesef: Gezondheidszorg en medisch beroep in Groningen 1500–1730 by Frank Huisman. Isis, 85 (3):517–518, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235499.

Canizares:1994:BRBa

[1734] Jorge Canizares. Book review: The Emergence of the Social Sciences, 1642–1792 by Richard Olson. Isis, 85(3):518–519, September 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235500.

Pyenson:1994:BRBa

[1735] Lewis Pyenson and Jean-François Gauvin. Book review: Mundializacion de la ciencia y cultura national by A. Lafuente; A. Elena; M. L. Ortega. Isis, 85(3):519–520, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235501.

Sageng:1994:BRB

[1736] Erik Sageng. Book review: Berkeley’s Philosophy of Mathematics by Douglas M. Jesseph. Isis, 85(3):520–521, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235502.

Grattan-Guinness:1994:BRBa

[1737] I. Grattan-Guinness. Book review: La storia di “F = ma”: Le seconda legge del moto nel XVIII secolo by Giulio Maltese. Isis, 85(3):521, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235503.

Calinger:1994:BRBa

[1738] Ronald Calinger. Book review: Sol et Luna II by Leonhardi Euleri; Charles Blanc. Isis, 85(3):521–522, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235504.

McEvoy:1994:BRB

[1739] John G. McEvoy. Book review: Enlightenment Science in the Romantic Era: The Chemistry of Berzelius and Its Cultural Setting by Evan M. Melhado; Tore Frangsmyr. Isis, 85(3):522–523, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235505.

Ewan:1994:BRB

[1740] Joseph Ewan. Book review: Joseph Banks: A Life by Patrick O’Brian. Isis, 85(3):523–524, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235506.

Rappaport:1994:BRB

[1741] Rhoda Rappaport. Book review: James Hutton and the History of Geology by Dennis R. Dean. Isis, 85(3):524–525, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235507.

Kearns:1994:BRB

[1742] Gerry Kearns. Book review: Death in Childbirth: An International Study of Maternal Care and Maternal Mortality, 1800–1950 by Irvine Loudon. Isis, 85(3):525–526, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235508.

Rabin:1994:BRB

[1743] Dana Rabin. Book review: The Most Solitary of Afflictions: Madness and Society in Britain, 1700–1900 by Andrew Scull. Isis, 85(3):526–527, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235509.

Hacking:1994:BRB

[1744] Ian Hacking. Book review: A History of Hypnotism by Alan Gauld; A Critique of Psychoanalytic Reason: Hypnosis as a Scientific Problem from Lavoisier to Lacan by Leon Chertok; Isabelle Stengers; Martha Noel Evans. Isis, 85(3):527–528, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235510.

Hamilton:1994:BRB

[1745] David Hamilton. Book review: Brown-Sequard: A Visionary of Science by Michael J. Aminoff. Isis, 85(3):528–529, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235511.

Nicolson:1994:BRB

[1746] Malcolm Nicolson. Book review: Scottish Medical Societies, 1731–1939: Their History and Records by Jacqueline Jenkinson. Isis, 85(3):529–530, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235512.

Nyhart:1994:BRB

[1747] Lynn K. Nyhart. Book review: Correspondence by Karl Ernst von Baer; Anton Dohrn; Christiane Groeben; Jane M. Oppenheimer. Isis, 85(3): 530, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235513.

Hodge:1994:BRB

[1748] M. J. S. Hodge. Book review: The Correspondence of Charles Darwin. Volume 7: 1858–1859. Supplement to the Correspondence, 1821–1857 by Charles Darwin; Frederick Burkhardt; Sydney Smith; Janet Browne; Marsha Richmond; The Correspondence of Charles Darwin. Volume 8: 1860 by Charles Darwin; Frederick Burkhardt; Duncan M. Porter; Janet Browne; Marsha Richmond. Isis, 85(3):530–531, September 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235514.

Rozwadowski:1994:BRBb

[1749] Helen Rozwadowski. Book review: At Sea with the Scientifics: The Challenger Letters of Joseph Matkin by Joseph Matkin; Philip F. Rehbock. Isis, 85(3):531–532, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235515.

Stevens:1994:BRB

[1750] P. F. Stevens. Book review: The European Discovery of the Indian Flora by Ray Desmond; Gentle Conquest: The Botanical Discovery of North America with Illustrations from the Library of Congress by James L. Reveal. Isis, 85(3):532–534, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235516.

Brooke:1994:BRBb

[1751] John Hedley Brooke. Book review: The Quiet Revolution: Hermann Kolbe and the Science of Organic Chemistry by Alan J. Rocke. Isis, 85 (3):534–535, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235517.

Johnson:1994:BRBb

[1752] Jeffrey A. Johnson. Book review: The Rainbow Makers: The Origins of the Synthetic Dyestuffs Industry in Western Europe by Anthony S. Travis. Isis, 85(3):535–536, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235518.

Hofmann:1994:BRBa

[1753] James R. Hofmann. Book review: L’invention de l’ingenieur moderne: L’Ecole des Ponts et Chaussees, 1747–1851 by Antoine Picon. Isis, 85 (3):536–537, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235519.

Outram:1994:BRB

[1754] Dorinda Outram. Book review: Masculinity and Male Codes of Honor in Modern France by Robert A. Nye. Isis, 85(3):537–538, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235520.

Moore:1994:BRBb

[1755] James Moore. Book review: Nature Lost? Natural Science and the German Theological Traditions of the Nineteenth Century by Frederick Gregory. Isis, 85(3):538–539, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235521.

Wilson:1994:BRB

[1756] Curtis Wilson. Book review: New Methods of Celestial Mechanics by Henri Poincaré; Daniel L. Goroff. Isis, 85(3):539–540, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235522.

Nye:1994:BRB

[1757] Robert Nye. Book review: Fields of Knowledge: French Academic Culture in Comparative Perspective, 1890–1920 by Fritz Ringer. Isis, 85 (3):540–541, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235523.

Rosario:1994:BRB

[1758] Vernon A. Rosario, II. Book review: The Self between: From Freud to the New Social Psychology of France by Eugene Webb. Isis, 85(3):541–542, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235524.

Morantz-Sanchez:1994:BRB

[1759] Regina Morantz-Sánchez. Book review: The Healer’s Tale: Transforming Medicine and Culture by Sharon R. Kaufman. Isis, 85(3):542–543, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235525.

Trumpler:1994:BRBa

[1760] Maria Trumpler. Book review: Cajal’s Degeneration and Regeneration of the Nervous System by Santiago Ramon y Cajal; Raoul M. May; Javier DeFelipe; Edward G. Jones. Isis, 85(3):543–544, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235526.

Hagen:1994:BRBb

[1761] Joel B. Hagen. Book review: Nature’s Champion: B. W. Wells, Tar Heel Ecologist by James R. Troyer. Isis, 85(3):544, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235527.

Holmes:1994:BRB

[1762] Frederic L. Holmes. Book review: Ein Genie irrt seltener…: Otto Heinrich Warburg: Ein Lebensbild in Dokumenten by Petra Werner; Reinhard Renneberg. Isis, 85(3):544–546, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235528.

Smith:1994:BRBf

[1763] Robert W. Smith. Book review: Pauper and Prince: Ritchey, Hale, and Big American Telescopes by Donald E. Osterbrock. Isis, 85(3):546, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235529.

Franklin:1994:BRBa

[1764] Allan Franklin. Book review: From c-Numbers to q-Numbers: The Classical Analogy in the History of Quantum Theory by Olivier Darrigol. Isis, 85(3):546–547, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235530.

Coopersmith:1994:BRBa

[1765] Jonathan Coopersmith. Book review: Electrifying Finland: The Transfer of a New Technology into a Late Industrialising Economy by Timo Myllyntaus. Isis, 85(3):547–548, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235531.

Walker:1994:BRBc

[1766] Mark Walker. Book review: Die Spezialisten: Deutsche Naturwissenschaftler und Techniker in der Sowjetunion nach 1945 by Ulrich Albrecht; Andreas Heinemann-Gruder; Arend Wellmann; 0Die Sowjetische Atombombe by Andreas Heinemann-Gruder. Isis, 85(3):548–549, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235532.

Reardon-Anderson:1994:BRB

[1767] James Reardon-Anderson. Book review: China’s Science and Technology Policy: 1949–1989 by Wang Yeu-Farn. Isis, 85(3):549–550, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235533.

Kauffman:1994:BRB

[1768] George B. Kauffman and Laurie M. Kauffman. Book review: The Atomic Energy Commission under Nixon: Adjusting to Troubled Times by Glenn T. Seaborg; Benjamin S. Loeb. Isis, 85(3):550–551, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235534.

Weldon:1994:BRB

[1769] Stephen Weldon. Book review: Science in the New Age: The Paranormal, Its Defenders and Debunkers, and American Culture by David J. Hess. Isis, 85(3):551–552, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235535.

Richards:1994:BRBc

[1770] Joan Richards. Book review: Mathematics in Society and History: Sociological Inquiries by Sal Restivo. Isis, 85(3):552–553, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235536.

Finocchiaro:1994:BRB

[1771] Maurice A. Finocchiaro. Book review: Paradigms and Barriers: How Habits of Mind Govern Scientific Beliefs by Howard Margolis. Isis, 85 (3):553–554, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235537.

Ludwig:1994:BRB

[1772] Jan Keith Ludwig. Book review: Philosophy of Science and Its Discontents by Steve Fuller. Isis, 85(3):554, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235538.

Hull:1994:BRB

[1773] David L. Hull. Book review: The Advancement of Science: Science without Legend, Objectivity without Illusions by Philip Kitcher. Isis, 85 (3):554–555, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235539.

Degen:1994:BRBa

[1774] Peter A. Degen. Book review: Doing Physics: How Physicists Take Hold of the World by Martin H. Krieger. Isis, 85(3):555–556, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235540.

Krasnoff:1994:BRB

[1775] Larry Krasnoff. Book review: Nietzsche and Modern Times: A Study of Bacon, Descartes, and Nietzsche by Laurence Lampert. Isis, 85(3): 556–557, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235541.

Cueto:1994:BRB

[1776] Marcos Cueto. Book review: Science across Cultures: An Annotated Bibliography of Books on Non-Western Science, Technology, and Medicine by Helaine Selin. Isis, 85(3):557, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235542.

dePater:1994:BRB

[1777] C. de Pater. Book review: Inventory of the Papers of Pieter Zeeman (1865–1943), Physicist and Nobel Prize Winner: c. 1877–1946 by P. J. M. Velthuys-Bechthold. Isis, 85(3):557–558, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235543.

Smith:1994:BRBg

[1778] Shawn M. Smith and Steven J. Livesey. Book review: The Use of Computers in Cataloging Medieval and Renaissance Manuscripts by Menso Folkerts; Andreas Kuhne; MARC Cataloging for Medieval Manuscripts by Hope Mayo; Bibliographic Access to Medieval and Renaissance Manuscripts: A Survey of Computerized Data Bases and Information Services by Wesley M. Stevens. Isis, 85(3):558–559, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235544.

Lattis:1994:BRBb

[1779] James M. Lattis. Book review: Bibliografia italiana di storia delle matematiche, 1961–1990 by Francesco Barbieri; Luigi Pepe. Isis, 85 (3):559–560, September 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235545.

Schneer:1994:BRB

[1780] Cecil J. Schneer. Book review: Gemology: An Annotated Bibliography by John Sinkankas. Isis, 85(3):560–561, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235546.

Paul:1994:BRBa

[1781] Diane B. Paul. Book review: Herbert Spencer: A Primary and Secondary Bibliography by Robert G. Perrin. Isis, 85(3):561–562, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235547.

Anonymous:1994:Cc

[1782] Anonymous. Collections. Isis, 85(3):562–565, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235548.

Anonymous:1994:BMd

[1783] Anonymous. Back matter. Isis, 85(3):566–572, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235549.

Richards:1994:PAM

[1784] Evelleen Richards. A political anatomy of monsters, hopeful and otherwise: Teratogeny, transcendentalism, and evolutionary theorizing. Isis, 85(3):viii + 377–411, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235460.

Anonymous:1994:FMd

[1785] Anonymous. Front matter. Isis, 85(3):i–vii, September 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235459.

Goldstein:1994:YSS

[1786] Daniel Goldstein. “Yours for science”: The Smithsonian Institution’s correspondents and the shape of scientific community in Nineteenth-Century America. Isis, 85(4):573–599, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235279.

Shapiro:1994:ACN

[1787] Alan E. Shapiro. Artists’ colors and Newton’s colors. Isis, 85(4):600–630, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235280.

McVaugh:1994:I

[1788] Michael R. McVaugh. Introduction. Isis, 85(4):631–632, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235281.

Dobbs:1994:NFC

[1789] B. J. T. Dobbs. Newton as final cause and first mover. Isis, 85(4): 633–643, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235282.

Eddy:1994:BHN

[1790] John H. Eddy, Jr. Buffon’s histoire naturelle: History? A critique of recent interpretations. Isis, 85(4):644–661, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235283.

Westfall:1994:EBJ

[1791] Richard S. Westfall. Eloge: B. J. T. Dobbs, 19 October 1930–29 March 1994. Isis, 85(4):662–663, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235284.

Buchwald:1994:ESD

[1792] Jed Z. Buchwald and Noel M. Swerdlow. Eloge: Stillman Drake, 24 December 1910–6 October 1993. Isis, 85(4):663–666, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235285.

Loncarica:1994:EDP

[1793] Alfredo G. Kohn Loncarica. Eloge: Desiderio Papp, 21 May 1895–31 January 1993. Isis, 85(4):666–667, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235286.

Joseph:1994:LE

[1794] George Gheverghese Joseph, David Pingree, S. I. Salem, and F. Jamil Ragep. Letters to the Editor. Isis, 85(4):668–670, December 1994. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235287.

Brush:1994:BRB

[1795] Stephen G. Brush. Book review: Inventory of Sources for History of Twentieth–Century Physics: Report and Microfiche Index to 700,000 Letters by Bruce R. Wheaton; Robin E. Rider. Isis, 85(4):671–672, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235288.

Collins:1994:BRB

[1796] H. M. Collins. Book review: We Have Never Been Modern by Bruno Latour; Catherine Porter. Isis, 85(4):672–674, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235289.

Whitney:1994:BRB

[1797] Elspeth Whitney. Book review: The Beginnings of Western Science: The European Scientific Tradition in Philosophical, Religious, and Institutional Context, 600 B.C. to A.D. 1450 by David C. Lindberg. Isis, 85 (4):674–675, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235290.

Major:1994:BRB

[1798] John S. Major. Book review: The Rise of Early Modern Science: Islam, China, and the West by Toby E. Huff. Isis, 85(4):675–676, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235291.

Schiebinger:1994:BRB

[1799] Londa Schiebinger. Book review: The Less Noble Sex: Scientific, Religious, and Philosophical Conceptions of Woman’s Nature by Nancy Tuana. Isis, 85(4):676–677, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235292.

Farquhar:1994:BRB

[1800] Judith Farquhar. Book review: Cosmology, Ontology, and Human Efficacy: Essays in Chinese Thought by Richard J. Smith; D. W. Y. Kwok. Isis, 85(4):677–678, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235293.

Golinski:1994:BRBb

[1801] Jan Golinski. Book review: The Enlightenment of Matter: The Definition of Chemistry from Agricola to Lavoisier by Marco Beretta. Isis, 85(4): 678–679, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235294.

Williams:1994:BRBb

[1802] Mari E. W. Williams. Book review: Making Instruments Count: Essays on Historical Scientific Instruments Presented to Gerard L’Estrange Turner by R. G. W. Anderson; J. A. Bennett; W. F. Ryan. Isis, 85(4): 679–680, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235295.

Rabkin:1994:BRB

[1803] Yakov M. Rabkin. Book review: Science in Russia and the Soviet Union: A Short History by Loren R. Graham. Isis, 85(4):680, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235296.

Hofmann:1994:BRBb

[1804] James R. Hofmann. Book review: The Physics of Chance: From Blaise Pascal to Niels Bohr by Charles Ruhla; G. Barton. Isis, 85(4):680–681, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235297.

Heinen:1994:BRB

[1805] Anton Heinen. Book review: The Need for a Sacred Science by Seyyed Hossein Nasr. Isis, 85(4):681–682, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235298.

Bunting:1994:BRB

[1806] Robert R. Bunting. Book review: The Wealth of Nature: Environmental History and the Ecological Imagination by Donald Worster. Isis, 85(4): 682–683, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235299.

Colaizzi:1994:BRB

[1807] Janet Colaizzi. Book review: American Suicide: A Psychocultural Exploration by Howard I. Kushner. Isis, 85(4):683, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235300.

Hildebrandt:1994:BRB

[1808] M. M. Hildebrandt. Book review: Correspondance. Volume 1: Lettres 1 à 129 by Gerbert d’Aurillac; P. Riche; J. P. Callu; Correspondance. Volume 2: Lettres 130 a 220 by Gerbert d’Aurillac; P. Riche; J. P. Callu; Zeit, Zahl und Bild: Studien zur Verbindung von Philosophie und Wissenschaft bei Abbo von Fleury by Eva-Maria Engelen. Isis, 85(4):683–685, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235301.

Sabra:1994:BRB

[1809] A. I. Sabra. Book review: Geometrie et dioptrique au Xe siècle: Ibn Sahl, al-Quhi, Ibn al-Haytham by Roshdi Rashed. Isis, 85(4):685–686, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235302.

Rosenfeld:1994:BRB

[1810] Boris Rosenfeld. Book review: Abu’l-Fath Abd al-Rahman al-Khazini (XII Century) by Mariam Mikhailovna Rozhanskaya. Isis, 85(4):686– 688, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235303.

Ronca:1994:BRB

[1811] Italo Ronca. Book review: La Razon de Lupus de Moros: Un poema hermetico by Andre Michalski. Isis, 85(4):688–689, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235304.

Ocker:1994:BRB

[1812] Christopher Ocker. Book review: A History of the University in Europe. Volume 1: Universities in the Middle Ages by Hilde De Ridder-Symoens. Isis, 85(4):689–690, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235305.

Pesenti:1994:BRB

[1813] Tiziana Pesenti. Book review: Arzt, Krankheit und Therapie bei Petrarca: Die Kritik an Medizin und Naturwissenschaft im italienischen Fruhhumanismus by Klaus Bergdolt. Isis, 85(4):690–691, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235306.

Stautz:1994:BRB

[1814] Burkhard Stautz. Book review: L’astrolabe: Les astrolabes du Musee Paul Dupuy by Raymond D’Hollander. Isis, 85(4):691, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235307.

Rosinska:1994:BRBb

[1815] Grazyna Rosinska. Book review: Rome Reborn: The Vatican Library and Renaissance Culture by Anthony Grafton. Isis, 85(4):691–692, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235308.

Gingerich:1994:BRBb

[1816] Owen Gingerich. Book review: Complete Works. Volume 4: The Manuscripts of Nicholas Copernicus’ Minor Works: Facsimiles by Nicholas Copernicus; Pawel Czartoryski. Isis, 85(4):692, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235309.

Gossin:1994:BRB

[1817] Pamela Gossin. Book review: Essays on Renaissance Literature. Volume 1: Donne and the New Philosophy by William Empson; John Haffenden.

Isis, 85(4):692–693, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235310.

Principe:1994:BRB

[1818] Lawrence M. Principe. Book review: Galileo, the Jesuits, and the Medieval Aristotle by William A. Wallace. Isis, 85(4):694, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235311.

Harris:1994:BRBa

[1819] Steven J. Harris. Book review: The “Astronomia Europaea” of Ferdinand Verbiest, S.J. (Dillingen, 1687) by Ferdinand Verbiest; Noel Golvers. Isis, 85(4):694–695, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235312.

Nadler:1994:BRBb

[1820] Steven Nadler. Book review: The Cartesian Empiricism of François Bayle by Thomas M. Lennon; Patricia Ann Easton; The Battle of the Gods and Giants: The Legacies of Descartes and Gassendi, 1655–1715 by Thomas M. Lennon. Isis, 85(4):695–696, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235313.

Verbeek:1994:BRB

[1821] Theo Verbeek. Book review: Descartes in Germania: La ricezione del cartesianesimo nella Facolta filosofica e medica di Duisburg (1652–1703) by Francesco Trevisani. Isis, 85(4):696–697, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235314.

Shackelford:1994:BRB

[1822] Jole Shackelford. Book review: Experiments on Birefringent Icelandic Crystal by Erasmus Bartholin; Thomas Archibald. Isis, 85(4):697, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235315.

Simonsuuri:1994:BRB

[1823] Kirsti K. Simonsuuri. Book review: The Rehabilitation of Myth: Vico’s “New Science.” by Joseph Mali. Isis, 85(4):697–698, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235316.

LaBerge:1994:BRB

[1824] Ann La Berge. Book review: The Popularization of Medicine, 1650– 1850 by Roy Porter. Isis, 85(4):698–699, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235317.

Kronick:1994:BRB

[1825] David A. Kronick. Book review: The Medical Casebook of William Brownrigg, M.D., F.R.S. (1712–1800) of the Town of Whitehaven in Cumberland by William Brownrigg; Jean E. Ward; Joan Yell. Isis, 85 (4):699–700, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235318.

Sepper:1994:BRB

[1826] Dennis L. Sepper. Book review: Le forme e il vivente: Morfologia e filosofia della natura in J. W. Goethe by Paola Giacomoni. Isis, 85(4): 700–701, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235319.

Trumpler:1994:BRBb

[1827] Maria Trumpler. Book review: The Ambiguous Frog: The Galvani-Volta Controversy on Animal Electricity by Marcello Pera; Jonathan Mandelbaum. Isis, 85(4):701–702, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235320.

Canizares:1994:BRBb

[1828] Jorge Canizares. Book review: La corona de las ciencias naturales: La medicina en el transito del siglo XVIII al XIX by Elvira Arquiola; Luis Montiel. Isis, 85(4):702–703, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235321.

Rupke:1994:BRB

[1829] Nicolaas A. Rupke. Book review: Briefe aus Amerika, 1799–1804 by Alexander von Humboldt; Ulrike Moheit. Isis, 85(4):703–704, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235322.

Calinger:1994:BRBb

[1830] Ronald Calinger. Book review: Begluckende Ermunterung durch die akademische Gemeinschaft: Alexander von Humboldt als Mitglied der Berliner Akademie der Wissenschaften by Kurt-R. Biermann. Isis, 85 (4):704, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235323.

Moore:1994:BRBc

[1831] James Moore. Book review: Contesting Cultural Authority: Essays in Victorian Intellectual Life by Frank M. Turner. Isis, 85(4):704–706, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235324.

Fisch:1994:BRB

[1832] Menachem Fisch. Book review: Defining Science: William Whewell, Natural Knowledge, and Public Debate in Early Victorian Britain by Richard Yeo. Isis, 85(4):706–707, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235325.

Grattan-Guinness:1994:BRBb

[1833] I. Grattan-Guinness. Book review: George Green: Mathematician and Physicist, 1793–1841: The Background to His Life and Work by D. M. Cannell. Isis, 85(4):707, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235326.

Hamlin:1994:BRB

[1834] Christopher Hamlin. Book review: The British Experience with River Pollution, 1865–1876 by Lawrence E. Breeze. Isis, 85(4):707–708, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235327.

Kremer:1994:BRB

[1835] Richard L. Kremer. Book review: The Search for Christian Doppler by Alec Eden. Isis, 85(4):708–709, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235328.

Servos:1994:BRB

[1836] John W. Servos. Book review: Raphael Pumpelly: Gentleman Geologist of the Gilded Age by Peggy Champlin. Isis, 85(4):709, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235329.

Greene:1994:BRBb

[1837] Mott T. Greene. Book review: Scenes from Deep Time: Early Pictorial Representations of the Prehistoric World by Martin J. S. Rudwick. Isis, 85(4):709–711, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235330.

Grob:1994:BRB

[1838] Gerald N. Grob. Book review: Living in the Shadow of Death: Tuberculosis and the Social Experience of Illness in American History by Sheila M. Rothman. Isis, 85(4):711–712, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235331.

Howell:1994:BRB

[1839] Joel D. Howell. Book review: Langstaff: A Nineteenth-Century Medical Life by Jacalyn Duffin. Isis, 85(4):712–713, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235332.

Kimler:1994:BRB

[1840] William Kimler. Book review: Darwinism by Peter J. Bowler. Isis, 85(4):713, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235333.

Low:1994:BRB

[1841] Reinhard Low. Book review: Die Allianz von Wissenschaft und Industrie: August Wilhelm Hofmann (1818–1892): Zeit, Werk, Wirkung by Christoph Meinel; Hartmut Scholz. Isis, 85(4):713–714, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235334.

Todd:1994:BRBc

[1842] Edmund N. Todd. Book review: Werner von Siemens: Erfinder und internationaler Unternehmer by Wilfried Feldenkirchen. Isis, 85(4):714– 715, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235335.

Kline:1994:BRB

[1843] Ronald R. Kline. Book review: America Calling: A Social History of the Telephone to 1940 by Claude S. Fischer. Isis, 85(4):715–716, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235336.

Harris:1994:BRBb

[1844] Neil Harris. Book review: World of Fairs: The Century-of-Progress Expositions by Robert W. Rydell. Isis, 85(4):716–718, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235337.

Parisi:1994:BRBb

[1845] Thomas Parisi. Book review: Freud and His Critics by Paul Robinson. Isis, 85(4):718–719, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235338.

Marx:1994:BRB

[1846] Otto M. Marx. Book review: Julius Wagner-Jauregg (1857–1940) by Magda Whitrow. Isis, 85(4):719, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235339.

Pyenson:1994:BRBb

[1847] Lewis Pyenson. Book review: Einstein 1905: De l’ether aux quanta by Françoise Balibar. Isis, 85(4):719–720, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235340.

Rothstein:1994:BRB

[1848] William G. Rothstein. Book review: Medical Lives and Scientific Medicine at Michigan, 1891–1969 by Joel D. Howell. Isis, 85(4):720– 721, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235341.

Kuklick:1994:BRB

[1849] Henrika Kuklick. Book review: Before Social Anthropology: Essays on the History of British Anthropology by James Urry. Isis, 85(4):721–722, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235342.

Young:1994:BRB

[1850] Michael W. Young. Book review: The Ethnographer’s Magic and Other Essays in the History of Anthropology by George W. Stocking. Isis, 85 (4):722, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235343.

Carson:1994:BRBb

[1851] John Carson. Book review: The Definition of a Profession: The Authority of Metaphor in the History of Intelligence Testing, 1890–1930 by JoAnne Brown. Isis, 85(4):722–723, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235344.

Paul:1994:BRBb

[1852] E. Robert Paul. Book review: Alexander A. Friedmann: The Man Who Made the Universe Expand by Eduard A. Tropp; Viktor Ya. Frenkel; Artur D. Chernin; Alexander Dron; Michael Burov. Isis, 85(4):723–724, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235345.

Hetherington:1994:BRB

[1853] Norriss S. Hetherington. Book review: The Man Who Sold the Milky Way: A Biography of Bart Bok by David H. Levy. Isis, 85(4):724–725, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235346.

Shapin:1994:BRB

[1854] Steven Shapin. Book review: Science and Anti-Science by Gerald Holton. Isis, 85(4):725–726, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235347.

Sime:1994:BRL

[1855] Ruth Lewin Sime. Book review: Lise Meitner an Otto Hahn: Briefe aus den Jahren 1912 bis 1924: Edition und Kommentierung by Sabine Ernst. Isis, 85(4):726, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235348.

Walker:1994:BRBd

[1856] Mark Walker. Book review: Technische Bildung, Hochschule, Staat und Wirtschaft: Entwicklungslinien des Technischen Hochschulwesens 1914– 1930: Das Beispiel der TH Braunschweig by Bettina Gundler. Isis, 85 (4):726–727, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235349.

Ash:1994:BRB

[1857] Mitchell G. Ash. Book review: Surviving the Swastika: Scientific Research in Nazi Germany by Kristie Macrakis. Isis, 85(4):727–729, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235350.

Wollons:1994:BRB

[1858] Roberta Wollons. Book review: Before Head Start: The Iowa Station and America’s Children by Hamilton Cravens. Isis, 85(4):729–730, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235351.

Barrow:1994:BRB

[1859] Mark V. Barrow, Jr. Book review: Restless Energy: A Biography of William Rowan, 1891–1957 by Marianne G. Ainley. Isis, 85(4):730–731, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235352.

Greenough:1994:BRB

[1860] Paul Greenough. Book review: Sentinel for Health: A History of the Centers for Disease Control by Elizabeth W. Etheridge. Isis, 85(4):731– 732, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235353.

Smith:1994:BRBh

[1861] Richard Candida Smith. Book review: Mechanism and Mysticism: The Influence of Science on the Thought and Work of Theodore Dreiser by Louis J. Zanine. Isis, 85(4):732–733, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235354.

Roberts:1994:BRB

[1862] Jon H. Roberts. Book review: B. F. Skinner: A Life by Daniel W. Bjork. Isis, 85(4):733–734, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235355.

Smith:1994:BRBi

[1863] Robert W. Smith. Book review: The Birth of NASA: The Diary of T. Keith Glennan by T. Keith Glennan; J. D. Hunley. Isis, 85(4):734–735, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235356.

Coopersmith:1994:BRBb

[1864] Jonathan Coopersmith. Book review: The Ghost of the Executed Engineer: Technology and the Fall of the Soviet Union by Loren R. Graham.

Isis, 85(4):735, December 1994. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235357.

Sanchez-Ron:1994:BRB

[1865] Jose M. Sánchez-Ron. Book review: Out of the Crystal Maze: Chapters from the History of Solid-State Physics by Lillian Hoddeson; Ernest Braun; Jurgen Teichmann; Spencer Weart. Isis, 85(4):735–736, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235358.

Degen:1994:BRBb

[1866] Peter A. Degen. Book review: The God Particle: If the Universe Is the Answer, What Is the Question? by Leon Lederman; Dick Teresi; Dreams of a Final Theory by Steven Weinberg. Isis, 85(4):736–738, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235359.

Whittemore:1994:BRB

[1867] Gilbert Whittemore. Book review: Containing the Atom: Nuclear Regulation in a Changing Environment, 1963–1971 by J. Samuel Walker. Isis, 85(4):738–739, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235360.

Norberg:1994:BRB

[1868] Arthur L. Norberg. Book review: The Computer in the United States: From Laboratory to Market, 1930–1960 by James W. Cortada. Isis, 85 (4):739, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235361.

Mahoney:1994:BRB

[1869] Michael S. Mahoney. Book review: Toward a History of Game Theory. (Annual Supplement to History of Political Economy, 24.) by E. Roy Weintraub. Isis, 85(4):739–740, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235362.

Miller:1994:BRB

[1870] David Philip Miller. Book review: Science as Practice and Culture by Andrew Pickering. Isis, 85(4):740–742, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235363.

Franklin:1994:BRBb

[1871] Allan Franklin. Book review: The Play of Nature: Experimentation as Performance by Robert P. Crease. Isis, 85(4):742–743, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235364.

Hayles:1994:BRB

[1872] N. Katherine Hayles. Book review: Culture and Cognition: The Boundaries of Literary and Scientific Inquiry by Ronald Schleifer; Robert Con Davis; Nancy Mergler. Isis, 85(4):743–744, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235365.

Chalmers:1994:BRB

[1873] Alan Chalmers. Book review: Philosophy of Science in the Twentieth Century: Four Central Themes by Donald Gillies. Isis, 85(4):744, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235366.

Ariew:1994:BRB

[1874] Roger Ariew. Book review: Leibniz et la methode de la science by François Duchesneau. Isis, 85(4):744–745, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235367.

White:1994:BRB

[1875] Owen White. Book review: Des sciences contre l’homme by Claude Blanckaert. Isis, 85(4):745–746, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235368.

Sloan:1994:BRB

[1876] Phillip R. Sloan. Book review: Discovering Complexity: Decomposition and Localization as Strategies in Scientific Research by William Bechtel; Robert C. Richardson. Isis, 85(4):746–747, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235369.

Wagner:1994:BRB

[1877] Stephen C. Wagner. Book review: The History of Science and Technology in the United States: A Critical and Selective Bibliography. Volume 2 by Marc Rothenberg. Isis, 85(4):747, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235370.

Eisberg:1994:BRBb

[1878] Joann Eisberg. Book review: Women in Chemistry and Physics: A Biobibliographic Sourcebook by Louise S. Grinstein; Rose K. Rose; Miriam H. Rafailovich. Isis, 85(4):747–748, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235371.

Anonymous:1994:Cd

[1879] Anonymous. Collections. Isis, 85(4):748–755, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235372.

Anonymous:1994:BMe

[1880] Anonymous. Back matter. Isis, 85(4):756–767, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235373.

Anonymous:1994:VI

[1881] Anonymous. Volume information. Isis, 85(4):768–790, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235277.

Anonymous:1994:FMe

[1882] Anonymous. Front matter. Isis, 85(4):??, December 1994. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235278.

Neu:1995:CBH

[1883] John Neu. Current bibliography of the history of science and its cultural influences, 1995. Isis, 86(??):1–331, ???? 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235891.

Anonymous:1995:BMa

[1884] Anonymous. Back matter. Isis, 86(??):??, ???? 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235892.

Anonymous:1995:FMa

[1885] Anonymous. Front matter. Isis, 86(??):i–xvii, ???? 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235890.

Duffin:1995:GMP

[1886] Jacalyn Duffin and Alison Li. Great moments: Parke, Davis and company and the creation of medical art. Isis, 86(1):1–29, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236407.

Hong:1995:FSE

[1887] Sungook Hong. Forging scientific electrical engineering: John Ambrose Fleming and the Ferranti effect. Isis, 86(1):30–51, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236408.

Brotons:1995:RCS

[1888] Victor Navarro Brotons. The reception of Copernicus in Sixteenth-Century Spain: The case of Diego de Zuñiga. Isis, 86(1):52–78, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236409.

Davis:1995:ERP

[1889] Edward B. Davis. Eloge: E. Robert Paul, 23 July 1943–12 October 1994. Isis, 86(1):79–80, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236410.

Schofield:1995:EDH

[1890] Robert E. Schofield. Eloge: Duane Henry Dubose Roller, 14 March 1920– 22 August 1994. Isis, 86(1):80–81, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236411.

Ariew:1995:BRB

[1891] Roger Ariew. Book review: Styles of Scientific Thinking in the European Tradition: The History of Argument and Explanation Especially in the Mathematical and Biomedical Sciences and Arts by A. C. Crombie. Isis, 86(1):82–83, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236412.

Tribby:1995:BRB

[1892] Jay Tribby. Book review: Configurations: A Journal of Literature, Science, and Technology. Volume 1, Numbers 1–3, 1993. Isis, 86(1):83–84, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236413.

Goldstein:1995:BRB

[1893] Jan Goldstein. Book review: A Vital Rationalist: Selected Writings from Georges Canguilhem by Georges Canguilhem; François Delaporte; Arthur Goldhammer. Isis, 86(1):84–85, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236414.

Applebaum:1995:BRB

[1894] Wilbur Applebaum. Book review: The Eye of Heaven: Ptolemy, Copernicus, Kepler by Owen Gingerich. Isis, 86(1):85–86, March 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236415.

Shank:1995:BRB

[1895] Michael H. Shank. Book review: The Ordering of Time: From the Ancient Computus to the Modern Computer by Arno Borst; Andrew Winnard. Isis, 86(1):86–87, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236416.

Fissell:1995:BRB

[1896] Mary E. Fissell. Book review: Medicine and the Five Senses by W. F. Bynum; Roy Porter. Isis, 86(1):87–88, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236417.

LaBerge:1995:BRB

[1897] Ann F. La Berge. Book review: Histoire du droit sanitaire en France by Vincent-Pierre Comiti. Isis, 86(1):88, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236418.

Jacyna:1995:BRBa

[1898] L. S. Jacyna. Book review: Origins of Neuroscience: A History of Explorations into Brain Function by Stanley Finger. Isis, 86(1):88–89, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236419.

Guerrini:1995:BRBa

[1899] Anita Guerrini. Book review: A History of the Life Sciences by Lois N. Magner. Isis, 86(1):89, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236420.

Fergerson:1995:BRB

[1900] Gerard Fergerson. Book review: The “Racial” Economy of Science: Toward a Democratic Future by Sandra Harding. Isis, 86(1):89–91, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236421.

Schatzberg:1995:BRB

[1901] Eric Schatzberg. Book review: Science and Technology in History: An Approach to Industrial Development by Ian Inkster. Isis, 86(1):91–92, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236422.

Reynolds:1995:BRB

[1902] Terry S. Reynolds. Book review: The Engineer in History by John Rae; Rudi Volti. Isis, 86(1):92–93, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236423.

Smith:1995:BRBa

[1903] Norman A. F. Smith. Book review: Power from Wind: A History of Windmill Technology by Richard L. Hills. Isis, 86(1):93, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236424.

Bocking:1995:BRB

[1904] Stephen Bocking. Book review: The Environmentalists: A Biographical Dictionary from the Seventeenth Century to the Present by Alan Axelrod; Charles Phillips. Isis, 86(1):93–94, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236425.

Yearley:1995:BRB

[1905] Steven Yearley. Book review: Interpreting Nature: Cultural Constructions of the Environment by I. G. Simmons. Isis, 86(1):94–95, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236426.

Lloyd:1995:BRBa

[1906] G. E. R. Lloyd. Book review: Theophrastus of Eresus: Sources for His Life, Writings, Thought, and Influence by William W. Fortenbaugh; Pamela M. Huby; Robert W. Sharples; Dimitri Gutas; Andrew D. Barker; John J. Keaney; David C. Mirhady; David Sedley; Michael G. Sollenberger. Isis, 86(1):95–96, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236427.

Paul:1995:BRB

[1907] John J. Paul. Book review: History of Medicine in India: From Antiquity to 1000 A.D. by Priya Vrat Sharma. Isis, 86(1):96–97, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236428.

Ivry:1995:BRB

[1908] Alfred L. Ivry. Book review: History of Islamic Philosophy by Henry Corbin; Liadain Sherrard; Philip Sherrard. Isis, 86(1):97, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236429.

Saliba:1995:BRB

[1909] George Saliba. Book review: Astronomy in the Service of Islam by David A. King. Isis, 86(1):97–98, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236430.

Conrad:1995:BRB

[1910] Lawrence I. Conrad. Book review: Avicenna by Lenn E. Goodman. Isis, 86(1):98–99, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236431.

Alfonso-Goldfarb:1995:BRB

[1911] Ana Maria Alfonso-Goldfarb. Book review: Medicinas, drogas y alimentos vegetales del nuevo mundo: Textos e imagenes españolas que los introdujeron en Europa by Jose M. Lopez Piñero; Jose Luis Fresquet Febrer; Maria Luz Lopez Terrada; Jose Pardo Tomas. Isis, 86(1):99– 100, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236432.

Siraisi:1995:BRBa

[1912] Nancy G. Siraisi. Book review: I codici di Bernardo Campagna: Filosofia e medicina alla fine del sec. by Stefano Caroti. Isis, 86(1):100–101, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236433.

Endress:1995:BRB

[1913] Gerhard Endress. Book review: De animalibus: Michael Scot’s Arabic-Latin Translation. Part 3, Books 15–19: Generation of Animals by Aristotle; Aafke M. I. van Oppenraaij; H. J. Drossaart Lulofs. Isis, 86(1): 101–102, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236434.

Lohr:1995:BRB

[1914] Charles H. Lohr. Book review: Ramon Lull und die Erkenntnislehre Thomas Le Myesiers by Theodor Pindl-Buchel. Isis, 86(1):102–103, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236435.

Papathanassiou:1995:BRB

[1915] Maria Papathanassiou. Book review: L’oro dei filosofi: Saggio sulle idee di un alchimista del Trecento by Michela Pereira. Isis, 86(1):103, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236436.

Molland:1995:BRB

[1916] George Molland. Book review: Nikolaus von Amiens: Ars fidei catholicae–Ein Beispielwerk axiomatischer Methode by Mechthild Dreyer. Isis, 86(1):103–104, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236437.

Samso:1995:BRB

[1917] Julio Samso. Book review: L’astronomia de Jacob ben David Bonjorn by Josep Chabas i Bergon; Antoni Roca i Rossell; Xavier Rodriguez i Gil. Isis, 86(1):104–105, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236438.

Newman:1995:BRBa

[1918] William Newman. Book review: De occulta philosophia: Libri tres by Cornelius Agrippa; V. Perrone Compagni. Isis, 86(1):105–106, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236439.

Rosinska:1995:BRB

[1919] Grazyna Rosinska. Book review: The Mastery of Nature: Aspects of Art, Science, and Humanism in the Renaissance by Thomas DaCosta Kaufmann. Isis, 86(1):106, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236440.

Gingerich:1995:BRB

[1920] Owen Gingerich. Book review: Copernico e la questione copernicana: Opera della Pubblica Biblioteca di Ferrara by Luigi Pepe. Isis, 86(1): 106–107, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236441.

Johns:1995:BRB

[1921] Adrian Johns. Book review: Science and the Secrets of Nature: Books of Secrets in Medieval and Early Modern Culture by William Eamon. Isis, 86(1):107–108, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236442.

Finocchiaro:1995:BRBa

[1922] Maurice A. Finocchiaro. Book review: A Defense of Galileo, the Mathematician from Florence: Which Is an Inquiry as to Whether the Philosophical View Advocated by Galileo Is in Agreement with, or Is Opposed to, the Sacred Scriptures by Thomas Campanella; Richard J. Blackwell. Isis, 86(1):108–109, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236443.

Bernstein:1995:BRB

[1923] Howard R. Bernstein. Book review: Die Bibliothek des Joachim Jungius: Ein Beitrag zur Historia littereria der fruhen Neuzeit by Christoph Meinel. Isis, 86(1):109–110, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236444.

Harris:1995:BRB

[1924] Steven J. Harris. Book review: The New (So-Called) Magdeburg Experiments of Otto von Guericke by Otto von Guericke; Margaret Glover Foley Ames. Isis, 86(1):110–111, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236445.

Osler:1995:BRBa

[1925] Margaret J. Osler. Book review: Geometria e atomismo nella scuola galileiana by Massimo Bucciantini; Maurizio Torrini. Isis, 86(1):111, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236446.

Wojcik:1995:BRB

[1926] Jan Wojcik. Book review: “Divulging of Useful Truths in Physick”: The Medical Agenda of Robert Boyle by Barbara Beigun Kaplan. Isis, 86(1): 111–112, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236447.

Crosland:1995:BRBa

[1927] Maurice Crosland. Book review: Verbindung und Affinitat: Die Grundlegung der neuzeitlichen Chemie an der Wende vom 17. zum 18. Jahrhundert by Ursula Klein. Isis, 86(1):112–113, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236448.

Mahoney:1995:BRBa

[1928] Michael S. Mahoney. Book review: Index of British Mathematicians. Part 3: 1701–1800 by R. V. Wallis; P. J. Wallis. Isis, 86(1):113, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236449.

Iliffe:1995:BRB

[1929] Rob Iliffe. Book review: All Was Light: An Introduction to Newton’s Opticks by A. Rupert Hall. Isis, 86(1):113–114, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236450.

Larson:1995:BRB

[1930] James Larson. Book review: Carl von Linne, 1707–1778: Arzt– Naturforscher–Systematiker by Heinz Goerke. Isis, 86(1):114–115, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236451.

Porter:1995:BRBa

[1931] Theodore M. Porter. Book review: La mesure de l’Etat: Administrateurs et geometres au XVIIIe siècle by Eric Brian. Isis, 86(1):115–116, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236452.

Hufbauer:1995:BRB

[1932] Karl Hufbauer. Book review: Paradigmenwechsel und Generationenkonflikt: Eine Fallstudie zur Struktur wissenschaftlicher Revolutionen: Die Revolution der Chemie des spaten 18. Jahrhunderts by Hans-Georg Schneider. Isis, 86(1):116–117, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236453.

Pickstone:1995:BRB

[1933] John V. Pickstone. Book review: The Epidemic Streets: Infectious Disease and the Rise of Preventive Medicine, 1856–1900 by Anne Hardy. Isis, 86(1):117, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236454.

Duffy:1995:BRB

[1934] John Duffy. Book review: Yellow Fever and Public Health in the New South by John H. Ellis; Yellow Fever and the South by Margaret Humphreys. Isis, 86(1):117–118, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236455.

Koerner:1995:BRB

[1935] Lisbet Koerner. Book review: Goethe’s History of Science by Karl J. Fink. Isis, 86(1):118–119, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236456.

Hackmann:1995:BRB

[1936] Wm. Kent Hackmann. Book review: R. B. Bate of the Poultry, 1782– 1847: The Life and Times of a Scientific Instrument Maker by Anita McConnell. Isis, 86(1):119–120, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236457.

Marx:1995:BRB

[1937] Otto M. Marx. Book review: Psychiatrie und Anthropologie in der ersten Halfte des 19. Jahrhunderts by Udo Benzenhofer. Isis, 86(1):120– 121, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236458.

Moyer:1995:BRB

[1938] Albert E. Moyer. Book review: The Correspondence of Michael Faraday. Volume 1: 1811–December 1831, Letters 1–524; The Correspondence of Michael Faraday. Volume 2: 1832–December 1840, Letters 525–1333 by Michael Faraday; Frank A. J. L. James. Isis, 86(1):121–122, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236459.

Williams:1995:BRBa

[1939] Elizabeth A. Williams. Book review: Auguste Comte: An Intellectual Biography. Volume 1 by Mary Pickering. Isis, 86(1):122–123, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236460.

Rupke:1995:BRBa

[1940] Nicolaas A. Rupke. Book review: Men among the Mammoths: Victorian Science and the Discovery of Human Prehistory by A. Bowdoin Van Riper. Isis, 86(1):123–124, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236461.

Reichenbach:1995:BRB

[1941] Herman Reichenbach. Book review: Bilder aus der Geschichte des Zoologischen Museums der Universität Hamburg: Die Zoologischen Sammlungen im Naturhistorischen Museum zu Hamburg wahrend seiner Kollegialverfassung, 1843 bis 1882. (Mitteilungen aus dem Hamburgischen Zoologischen Museum und Institut, Suppl. to Vol. 90.) by Herbert Weidner; Chronik der Zoologischen Staatssammlung Munchen: Festschrift zur Verabschiedung des Direktors der Zoologischen Staatssammlung Munchen, Prof. Dr. Ernst Josef Fittkau, 1976–1992 by E. Diller; A. Hausmann. Isis, 86(1):124–125, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236462.

Yeo:1995:BRB

[1942] Richard Yeo. Book review: Paradise Revealed: Natural History in Nineteenth-Century Australia by Colin Finney. Isis, 86(1):125–126, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236463.

Camerini:1995:BRB

[1943] Jane Camerini. Book review: Alfred Russel Wallace: An Anthology of His Shorter Writings by Alfred Russel Wallace; Charles H. Smith. Isis, 86(1):126–127, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236464.

Pyenson:1995:BRBa

[1944] Lewis Pyenson. Book review: Nature, the Exotic, and the Science of French Colonialism by Michael A. Osborne. Isis, 86(1):127–128, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236465.

Hiebert:1995:BRB

[1945] Erwin N. Hiebert. Book review: Letters of Hermann von Helmholtz to His Parents: The Medical Education of a German Scientist, 1837– 1846 by David Cahan; Hermann von Helmholtz and the Foundations of Nineteenth-Century Science by David Cahan. Isis, 86(1):128–130, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236466.

Husberg:1995:BRB

[1946] Volker Husberg. Book review: Herrschaft über die Natur? Naturvorstellungen deutscher Ingenieure, 1871–1914 by Hans-Liudger Dienel. Isis, 86(1):130–131, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236467.

Inkster:1995:BRB

[1947] Ian Inkster. Book review: The Papers of Thomas A. Edison. Volume 2: From Workshop to Laboratory, June 1873–March 1876 by Thomas A. Edison; Robert A. Rosenberg; Paul B. Israel; Keith A. Nier; Melodie Andrews; Thomas A. Edison and the Modernization of America by Martin V. Melosi. Isis, 86(1):131–132, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236468.

Solms:1995:BRB

[1948] Mark Solms. Book review: Freud’s Dream: A Complete Interdisciplinary Science of Mind by Patricia Kitcher. Isis, 86(1):132–133, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236469.

Pauly:1995:BRB

[1949] Philip J. Pauly. Book review: Silent Travelers: Germs, Genes, and the “Immigrant Menace.” by Alan M. Kraut. Isis, 86(1):133–135, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236470.

Stehle:1995:BRB

[1950] Philip Stehle. Book review: Early Quantum Electrodynamics: A Source Book by Arthur I. Miller; Walter Grant. Isis, 86(1):135–136, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236471.

Ash:1995:BRBa

[1951] Mitchell G. Ash. Book review: The Nazi Connection: Eugenics, American Racism, and German National Socialism by Stefan Kuhl. Isis, 86 (1):136–137, March 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236472.

Walker:1995:BRB

[1952] Mark Walker. Book review: Scientific Information in Wartime: The Allied-German Rivalry, 1939–1945 by Pamela Spence Richards. Isis, 86 (1):137, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236473.

Gorelik:1995:BRB

[1953] Gennady Gorelik. Book review: G. I. Budker: Reflections and Remembrances by Boris N. Breizman; James W. Van Dam. Isis, 86(1):137–139, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236474.

Ezrahi:1995:BRB

[1954] Yaron Ezrahi. Book review: The Cold War and American Science: The Military-Industrial-Academic Complex at MIT and Stanford by Stuart W. Leslie. Isis, 86(1):139–140, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236475.

Sigurdson:1995:BRB

[1955] Jon Sigurdson. Book review: Science, Technology, and Society in Postwar Japan by Shigeru Nakayama. Isis, 86(1):140–141, March 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236476.

Doel:1995:BRB

[1956] Ronald E. Doel. Book review: Race to the Moon: America’s Duel with the Soviets by William B. Breuer. Isis, 86(1):141–143, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236477.

Kline:1995:BRBa

[1957] Ronald Kline. Book review: Ferranti-Packard: Pioneers in Canadian Electrical Manufacturing by Norman R. Ball; John N. Vardalas. Isis, 86 (1):143, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236478.

Steneck:1995:BRB

[1958] Nicholas Steneck. Book review: Science on Trial: The Whistle-Blower, the Accused, and the Nobel Laureate by Judy Sarasohn. Isis, 86(1):143– 144, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236479.

Lewenstein:1995:BRBa

[1959] Bruce V. Lewenstein. Book review: Cold Fusion: The Scientific Fiasco of the Century by John R. Huizenga; Bad Science: The Short Life and Weird Times of Cold Fusion by Gary Taubes. Isis, 86(1):144–146, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236480.

Findlen:1995:BRB

[1960] Paula Findlen. Book review: Culture and Enchantment by Mark A. Schneider. Isis, 86(1):146–147, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236481.

Jesseph:1995:BRBa

[1961] Douglas M. Jesseph. Book review: Common Sense, Science, and Scepticism: A Historical Introduction to the Theory of Knowledge by Alan Musgrave. Isis, 86(1):147, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236482.

Waters:1995:BRB

[1962] C. Kenneth Waters. Book review: Philosophy of Biology by Elliott Sober. Isis, 86(1):147–148, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236483.

Miles:1995:BRB

[1963] Sara Joan Miles. Book review: Women and Technology: An Annotated Bibliography by Cynthia Gay Bindocci. Isis, 86(1):148–149, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236484.

Anonymous:1995:Ca

[1964] Anonymous. Collections. Isis, 86(1):149–161, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236485.

Anonymous:1995:BMb

[1965] Anonymous. Back matter. Isis, 86(1):162–166, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236486.

Anonymous:1995:FMb

[1966] Anonymous. Front matter. Isis, 86(1):??, March 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236406.

Acker:1995:ALW

[1967] Caroline Jean Acker. Addiction and the laboratory: The work of the National Research Council’s Committee on Drug Addiction, 1928–1939. Isis, 86(2):167–193, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236321.

Kline:1995:CTA

[1968] Ronald Kline. Construing “technology” as “applied science”: Public rhetoric of scientists and engineers in the United States, 1880–1945. Isis, 86(2):194–221, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236322.

Price:1995:RDC

[1969] Matt Price. Roots of dissent: The Chicago Met Lab and the origins of the Franck Report. Isis, 86(2):222–244, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236323.

Lucier:1995:CIS

[1970] Paul Lucier. Commercial interests and scientific disinterestedness: Consulting geologists in antebellum America. Isis, 86(2):245–267, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236324.

Mendelsohn:1995:TMS

[1971] J. Andrew Mendelsohn. “Typhoid Mary” strikes again: The social and the scientific in the making of modern public health. Isis, 86(2):268– 277, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236325. See [1078].

Elliott:1995:AMH

[1972] Clark A. Elliott, Richard L. Kremer, Keith R. Benson, and Nicholas H. Clulee. Annual meeting of the History of Science Society New Orleans, 13–16 October 1994. Isis, 86(2):278–285, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236326.

Thackray:1995:ESM

[1973] Arnold Thackray. Eloge: Sidney M. Edelstein, 22 January 1912–18 September 1994. Isis, 86(2):285–288, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236327.

Eddy:1995:LE

[1974] John H. Eddy, Jr. Letter to the Editor. Isis, 86(2):289, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236328.

Dauben:1995:BRB

[1975] Joseph W. Dauben. Book review: Charles Sanders Peirce: A Life. Isis, 86(2):290–299, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236329.

North:1995:BRBa

[1976] J. D. North. Book review: Planets, Stars, and Orbs: The Medieval Cosmos, 1200–1687. Isis, 86(2):300–301, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236330.

Secord:1995:BRB

[1977] James A. Secord. Book review: Research Schools: Historical Reappraisals by Gerald L. Geison; Frederic L. Holmes. Isis, 86(2):301–302, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236331.

Guerrini:1995:BRBb

[1978] Anita Guerrini. Book review: Doctors and Ethics: The Earlier Historical Setting of Professional Ethics by Andrew Wear; Johanna Geyer-Kordesch; Roger French. Isis, 86(2):302–303, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236332.

Slotten:1995:BRB

[1979] Hugh Richard Slotten. Book review: Minds for the Making: The Role of Science in American Education, 1750–1990 by Scott L. Montgomery. Isis, 86(2):303, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236333.

Jacob:1995:BRB

[1980] Margaret C. Jacob. Book review: Science, Technology, and Society: A Historical Perspective by Martin Fichman. Isis, 86(2):303–304, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236334.

Ainley:1995:BRB

[1981] Marianne Gosztonyi Ainley. Book review: On the Edge of Discovery: Australian Women in Science by Farley Kelly. Isis, 86(2):304–305, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236335.

Barberis:1995:BRB

[1982] Daniela Barberis. Book review: Anthropologie naïve, anthropologie savante: De l’origine de l’homme, de l’imagination et des idées reçues by Wiktor Stoczkowski. Isis, 86(2):305, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236336.

Gatewood:1995:BRB

[1983] Willard B. Gatewood. Book review: God’s Own Scientists: Creationists in a Secular World by Christopher P. Toumey; The Evolution Controversy in America by George E. Webb. Isis, 86(2):305–307, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236337.

Mahoney:1995:BRBb

[1984] Michael S. Mahoney. Book review: The Fourth Discontinuity: The Co-Evolution of Humans and Machines by Bruce Mazlish. Isis, 86(2):307– 308, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236338.

Sageng:1995:BRB

[1985] Erik Sageng. Book review: e: The Story of a Number by Eli Maor. Isis, 86(2):308, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236339.

Bowen:1995:BRB

[1986] Alan C. Bowen. Book review: Early Astronomy by Hugh Thurston. Isis, 86(2):309, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236340.

Rochberg:1995:BRB

[1987] F. Rochberg. Book review: Archaic Bookkeeping: Early Writing and Techniques of Economic Administration in the Ancient Near East by Hans J. Nissen; Peter Damerow; Robert K. Englund; Paul Larsen. Isis, 86(2):309–310, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236341.

Gerson:1995:BRB

[1988] Lloyd P. Gerson. Book review: Posterior Analytics by Aristotle; Jonathan Barnes. Isis, 86(2):310–311, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236342.

Nutton:1995:BRB

[1989] Vivian Nutton. Book review: The Healing Past: Pharmaceuticals in the Biblical and Rabbinic World by Irene Jacob; Walter Jacob. Isis, 86(2): 311–312, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236343.

Tessenow:1995:BRB

[1990] Hermann Tessenow. Book review: Heaven and Earth in Early Han Thought: Chapters Three, Four, and Five of the Huainanzi by John S. Major. Isis, 86(2):312–313, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236344.

Pingree:1995:BRB

[1991] David Pingree. Book review: Nasir al-Din al-Tusi’s Memoir on Astronomy (al Tadhkira fi ’ilm al-hay a) by F. J. Ragep. Isis, 86(2):313–314, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236345.

Voigts:1995:BRBa

[1992] Linda Ehrsam Voigts. Book review: Anglo-Saxon Medicine by M. L. Cameron. Isis, 86(2):314–315, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236346.

Eastwood:1995:BRB

[1993] Bruce S. Eastwood. Book review: Die Artes liberales im fruhen Mittelalter (5.-9. Jh.): Das quadrivium und der komputus als Indikatoren fur Kontinuitat und Erneuerung der exakten Wissenschaften zwischen Antike und Mittelalter by Brigitte Englisch. Isis, 86(2):315–316, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236347.

Siraisi:1995:BRBb

[1994] Nancy G. Siraisi. Book review: Les consilia medicaux by Jole Agrimi; Chiara Crisciani; Caroline Viola. Isis, 86(2):316, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236348.

Sinex:1995:BRB

[1995] M. A. Sinex. Book review: Petrus Alfonsi and His Medieval Readers by John Tolan. Isis, 86(2):316–317, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236349.

Courtenay:1995:BRB

[1996] William J. Courtenay. Book review: Megenberg aus zweiter Hand: Überlieferungsgeschichtliche Studien zur Redaktion B des Buchs von den naturlichen Dingen by Walter Buckl. Isis, 86(2):317–318, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236350.

Freudenthal:1995:BRBa

[1997] Gad Freudenthal. Book review: The Jewish Alchemists: A History and Source Book by Raphael Patai. Isis, 86(2):318–319, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236351.

Cormack:1995:BRB

[1998] Lesley B. Cormack. Book review: Mapmakers of the Sixteenth Century and Their Maps: Biobibliographies of the Cartographers of Abraham Ortelius, 1570 by Robert W. Karrow, Jr. Isis, 86(2):319, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236352.

Bennett:1995:BRB

[1999] Jim Bennett. Book review: Maps in Tudor England by P. D. A. Harvey. Isis, 86(2):319–321, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236353.

Baldwin:1995:BRB

[2000] Martha Baldwin. Book review: Legem impone subactis: Studi su filosofia e scienza dei Gesuiti in Italia, 1540–1632 by Ugo Baldini. Isis, 86(2): 321–322, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236354.

Finocchiaro:1995:BRBb

[2001] Maurice A. Finocchiaro. Book review: Forme di esperienza e rivoluzione scientifica by Gabriele Baroncini. Isis, 86(2):322–323, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236355.

Osler:1995:BRBb

[2002] Margaret J. Osler. Book review: Renaissance and Revolution: Humanists, Scholars, Craftsmen, and Natural Philosophers in Early Modern Europe by J. V. Field; Frank A. J. L. James. Isis, 86(2):323–324, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236356.

Vickers:1995:BRB

[2003] Brian Vickers. Book review: Francis Bacon and the Project of Progress by Robert K. Faulkner; Francis Bacon: History, Politics, and Science, 1561–1626 by B. H. G. Wormald. Isis, 86(2):324–325, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236357.

Jesseph:1995:BRBb

[2004] Douglas M. Jesseph. Book review: Les raisons de l’infini: Du monde clos á l’univers mathématique by Michel Blay. Isis, 86(2):325–326, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236358.

Wallace:1995:BRB

[2005] William A. Wallace. Book review: Galileo: Decisive Innovator by Michael Sharratt. Isis, 86(2):326–327, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236359.

Mandelbrote:1995:BRB

[2006] Scott Mandelbrote. Book review: The “Arabick” Interest of the Natural Philosophers in Seventeenth-Century England by G. A. Russell. Isis, 86 (2):327–328, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236360.

Grafton:1995:BRB

[2007] Anthony Grafton. Book review: The Atlantic Vision: Olaus Rudbeck and Baroque Science by Gunnar Eriksson. Isis, 86(2):328, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236361.

Terrall:1995:BRB

[2008] Mary Terrall. Book review: Der Briefwechsel von Jacob Bernoulli by Jacob Bernoulli; Andre Weil; Clifford Truesdell; Fritz Nagel; Die Werke von Jakob Bernoulli. Volume 4: Reihentheorie by Jakob Bernoulli; Andre Weil. Isis, 86(2):329–330, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236362.

Roberts:1995:BRB

[2009] Lissa Roberts. Book review: Artful Science: Enlightenment Entertainment and the Eclipse of Visual Education by Barbara Maria Stafford. Isis, 86(2):330, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236363.

Baker:1995:BRB

[2010] Keith Michael Baker. Book review: Arthmetique politique: Textes rares ou inedits (1767–1789) by Condorcet; Bernard Bru; Pierre Crepel. Isis, 86(2):330–331, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236364.

Lux:1995:BRB

[2011] David S. Lux. Book review: Colonialism and Science: Saint Domingue in the Old Regime by James E. McClellan III. Isis, 86(2):331–332, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236365.

Ramsey:1995:BRB

[2012] Matthew Ramsey. Book review: The Citizen-Patient in Revolutionary and Imperial Paris by Dora B. Weiner. Isis, 86(2):332–333, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236366.

Kremer:1995:BRBa

[2013] Richard L. Kremer. Book review: Über die Verhaltnisse der organischen Krafte unter einander in der Reihe der verschiedenen Organisationen, die Geseze und Folgen dieser Verhaltnisse by Carl Friedrich Kielmeyer; Kielmeyer-Bibliographie: Verzeichnis der Literatur von und über den Naturforscher Carl Friedrich Kielmeyer (1765–1844) by Kai Torsten Kanz. Isis, 86(2):333–334, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236367.

Outram:1995:BRB

[2014] Dorinda Outram. Book review: The Physical and the Moral: Anthropology, Physiology, and Philosophical Medicine in France, 1750–1850 by Elizabeth A. Williams. Isis, 86(2):334–335, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236368.

Markel:1995:BRB

[2015] Howard Markel. Book review: Genetics and Medicine in the United States, 1800–1922 by Alan R. Rushton. Isis, 86(2):335, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236369.

Jackson:1995:BRB

[2016] Myles W. Jackson. Book review: Wilhelm Olbers: Im Briefwechsel mit Astronomen seiner Zeit by Diedrich Wattenberg. Isis, 86(2):335–336, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236370.

Herbert:1995:BRB

[2017] Sandra Herbert. Book review: Vestiges of the Natural History of Creation: And Other Evolutionary Writings by Robert Chambers; James A. Secord. Isis, 86(2):336–337, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236371.

Arens:1995:BRB

[2018] Katherine Arens. Book review: Philosophische Psychologie im 19. Jahrhundert: Entstehung und Problemgeschichte by Klaus Sachs-Hombach. Isis, 86(2):337, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236372.

Brown:1995:BRB

[2019] T. E. Brown. Book review: La science du mal: L’institution de la psychiatrie au Quebec, 1800–1914 by Peter Keating. Isis, 86(2):337–338, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236373.

Fleming:1995:BRB

[2020] James Rodger Fleming. Book review: Patronage, Practice, and the Culture of American Science: Alexander Dallas Bache and the U.S. Coast Survey by Hugh Richard Slotten. Isis, 86(2):338–339, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236374.

Bonner:1995:BRB

[2021] Thomas Neville Bonner. Book review: Order out of Chaos: John Shaw Billings and America’s Coming of Age by Carleton B. Chapman. Isis, 86(2):339–340, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236375.

Stiling:1995:BRB

[2022] Rodney L. Stiling. Book review: God and the Natural World: Religion and Science in Antebellum America by Walter H. Conser, Jr. Isis, 86 (2):340, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236376.

Harman:1995:BRB

[2023] P. M. Harman. Book review: The Creation of Scientific Effects: Heinrich Hertz and Electric Waves by Jed Z. Buchwald. Isis, 86(2):341–342, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236377.

Maienschein:1995:BRB

[2024] Jane Maienschein. Book review: Darwinism and Social Darwinism in Imperial Germany: The Contribution of the Cell Biologist Oscar Hertwig (1849–1922) by Paul Julian Weindling. Isis, 86(2):342, June 1995.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236378.

Pratt:1995:BRB

[2025] Dale J. Pratt. Book review: Organic Memory: History and the Body in the Late Nineteenth and Early Twentieth Centuries by Laura Otis. Isis, 86(2):342–343, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236379.

Gitelman:1995:BRB

[2026] Lisa Gitelman. Book review: From Tinfoil to Stereo: The Acoustic Years of the Recording Industry, 1877–1929 by Walter L. Welch; Leah Brodbeck Stenzel Burt. Isis, 86(2):343–344, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236380.

Cueto:1995:BRBa

[2027] Marcos Cueto. Book review: Cobras, lagartos y outros bichos, uma historia comparada dos institutos Oswaldo Cruz e Butantan by Jaime Larry Benchimol; Luiz Antonio Teixeira. Isis, 86(2):344, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236381.

Pyenson:1995:BRBb

[2028] Lewis Pyenson. Book review: Einstein a Paris: Le temps n’est plus. … by Michel Biezunski. Isis, 86(2):344–345, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236382.

Teigen:1995:BRB

[2029] Philip M. Teigen. Book review: The Winning of Animal Health: One Hundred Years of Veterinary Medicine by O. H. V. Stalheim. Isis, 86 (2):345, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236383.

Gevitz:1995:BRB

[2030] Norman Gevitz. Book review: Chiropractic in America: The History of a Medical Alternative by J. Stuart Moore. Isis, 86(2):345–346, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236384.

Plas:1995:BRB

[2031] Regine Plas. Book review: L’invention de la psychologie moderne by Genevieve Paicheler. Isis, 86(2):346–347, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236385.

Borossa:1995:BRB

[2032] Julia Borossa. Book review: The Legacy of Sandor Ferenczi by Lewis Aron; Adrienne Harris; The Correspondence of Sigmund Freud and Sandor Ferenczi. Volume 1: 1908–1914 by Sigmund Freud; Sandor Ferenczi; Eva Brabant; Ernst Falzeder; Patricia Giampieri-Deutsch; Andre Haynal; Ingeborg Meyer-Palmedo; Peter T. Hoffer. Isis, 86(2):347–348, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236386.

Todd:1995:BRB

[2033] Edmund N. Todd. Book review: Erwin Marx (1893–1980): Ingenieurwissenschaftler in Braunschweig, und die Forschung und Entwicklung auf dem Gebiet der elektrischen Energieubertragung auf weite Entfernungen zwischen 1918 und 1950 by Helmut Maier. Isis, 86(2):348–349, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236387.

Lewis:1995:BRBa

[2034] Albert C. Lewis. Book review: Mathematics for the Multitude? A History of the Mathematical Association by Michael H. Price. Isis, 86(2):349– 350, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236388.

Croizier:1995:BRB

[2035] Ralph Croizier. Book review: Bamboo Stone: The Evolution of a Chinese Medical Elite by Karen Minden. Isis, 86(2):350, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236389.

Friedman:1995:BRBa

[2036] Robert Marc Friedman. Book review: Center on the Periphery: Historical Aspects of Twentieth–Century Swedish Physics by Svante Lindqvist; Marika Hedin; Thomas Kaiserfeld. Isis, 86(2):350–351, June 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236390.

Scharff:1995:BRB

[2037] Virginia Scharff. Book review: Taking Charge: The Electric Automobile in America by Michael Brian Schiffer; Tamara C. Butts; Kimberly K. Grimm. Isis, 86(2):351–352, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236391.

Lunbeck:1995:BRB

[2038] Elizabeth Lunbeck. Book review: Science in the Bedroom: A History of Sex Research by Vern L. Bullough. Isis, 86(2):352–353, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236392.

Webb:1995:BRB

[2039] George E. Webb. Book review: Oak Ridge National Laboratory: The First Fifty Years by Leland Johnson; Daniel Schaffer. Isis, 86(2):353, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236393.

Lindee:1995:BRBa

[2040] Susan Lindee. Book review: Operation Crossroads: The Atomic Tests at Bikini Atoll by Jonathan M. Weisgall. Isis, 86(2):353–355, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236394.

Whittemore:1995:BRB

[2041] Gilbert Whittemore. Book review: Life under a Cloud: American Anxiety about the Atom by Allan M. Winkler. Isis, 86(2):355, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236395.

Hovis:1995:BRBa

[2042] R. Corby Hovis. Book review: The Rise and Fall of the Fifth Force: Discovery, Pursuit, and Justification in Modern Physics by Allan Franklin. Isis, 86(2):355–356, June 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236396.

Lewenstein:1995:BRBb

[2043] Bruce V. Lewenstein. Book review: Mass-Extinction Debates: How Science Works in a Crisis by William Glen. Isis, 86(2):356–357, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236397.

Marsden:1995:BRB

[2044] Ben Marsden. Book review: The Golem: What Everyone Should Know about Science by Harry Collins; Trevor Pinch. Isis, 86(2):357–358, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236398.

Pinch:1995:BRB

[2045] Trevor Pinch. Book review: The Anthropology of Science and Technology by David J. Hess; Linda L. Layne. Isis, 86(2):358, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236399.

Luthy:1995:BRB

[2046] Christoph Luthy. Book review: Am Ende des Baconschen Zeitalters: Studien zur Wissenschaftsentwicklung by Gernot Bohme. Isis, 86(2): 358–359, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236400.

Spallone:1995:BRB

[2047] Pat Spallone. Book review: Love, Power, and Knowledge: Towards a Feminist Transformation of the Sciences by Hilary Rose. Isis, 86(2): 359–360, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236401.

Williams:1995:BRBb

[2048] L. Pearce Williams. Book review: The Biographical Dictionary of Scientists by Roy Porter. Isis, 86(2):360–361, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236402.

Kimmelman:1995:BRB

[2049] Barbara A. Kimmelman. Book review: The History of Agricultural Science and Technology: An International Annotated Bibliography by R. Douglas Hurt; Mary Ellen Hurt. Isis, 86(2):361, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236403.

Anonymous:1995:Cb

[2050] Anonymous. Collections. Isis, 86(2):361–367, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236404.

Anonymous:1995:BMc

[2051] Anonymous. Back matter. Isis, 86(2):368–372, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236405.

Anonymous:1995:FMc

[2052] Anonymous. Front matter. Isis, 86(2):??, June 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236320.

Muendel:1995:FLM

[2053] John Muendel. Friction and lubrication in Medieval Europe: The emergence of olive oil as a superior agent. Isis, 86(3):373–393, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235019.

Baldwin:1995:SEE

[2054] Martha Baldwin. The Snakestone Experiments: An early modern medical debate. Isis, 86(3):394–418, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235020.

Kusch:1995:RII

[2055] Martin Kusch. Recluse, interlocutor, interrogator: Natural and social order in turn-of-the-century psychological research schools. Isis, 86(3): 419–439, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235021.

Hollinger:1995:SWK

[2056] David A. Hollinger. Science as a weapon in Kulturkampfe in the United States during and after World War II. Isis, 86(3):440–454, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235022.

Stigler:1995:ECE

[2057] Stephen M. Stigler. Eloge: Churchill Eisenhart, 11 March 1910–25 June 1994. Isis, 86(3):455–456, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235023.

Pepe:1995:LE

[2058] Luigi Pepe and James M. Lattis. Letters to the Editor. Isis, 86(3):457, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235024.

Ash:1995:BRBb

[2059] Mitchell G. Ash. Book review: Hochschule und Nationalsozialismus. Isis, 86(3):458–462, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235025.

Pacey:1995:BRB

[2060] Arnold Pacey. Book review: Thinking through Technology: The Path between Engineering and Philosophy by Carl Mitcham. Isis, 86(3):463, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235026.

Kline:1995:BRBb

[2061] Ronald Kline. Book review: Future Imperfect: The Mixed Blessings of Technology in America by Howard P. Segal. Isis, 86(3):463–464, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235027.

Christianson:1995:BRBa

[2062] J. R. Christianson. Book review: Nordic Science in Historical Perspective by Jole Shackelford. Isis, 86(3):464–465, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235028.

Jardine:1995:BRB

[2063] Nicholas Jardine. Book review: Models of the History of Philosophy. Volume 1: From Its Origins in the Renaissance to the “Historia Philosophica.” by Francesco Bottin; Luciano Malusa; Giuseppe Micheli; Giovanni Santinello; Ilario Tolomio; Constance W. Blackwell; Philip Weller. Isis, 86(3):465–466, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235029.

Mitman:1995:BRB

[2064] Gregg Mitman. Book review: Nature’s Economy: A History of Ecological Ideas by Donald Worster. Isis, 86(3):466–467, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235030.

Wildberg:1995:BRB

[2065] Christian Wildberg. Book review: Self-Motion: From Aristotle to Newton by Mary Louise Gill; James G. Lennox. Isis, 86(3):467–468, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235031.

Laqueur:1995:BRB

[2066] Thomas W. Laqueur. Book review: Women’s Bodies in Classical Greek Science by Lesley Dean-Jones. Isis, 86(3):468–469, September 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235032.

Lloyd:1995:BRBb

[2067] G. E. R. Lloyd. Book review: Neuer Kommentar zu den ersten zwolf Krankengeschichten im III Epidemienbuch des Hippokrates by Charles Lichtenthaeler. Isis, 86(3):469, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235033.

Hon:1995:BRB

[2068] Giora Hon. Book review: Models and Precision: The Quality of Ptolemy’s Observations and Parameters by John Phillips Britton. Isis, 86 (3):469–470, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235034.

Bodson:1995:BRB

[2069] Liliane Bodson. Book review: Animals in Celtic Life and Myth by Miranda Green. Isis, 86(3):470–472, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235035.

Richter-Bernburg:1995:BRB

[2070] Lutz Richter-Bernburg. Book review: Die Dioskurides-Erklarung des Ibn al-Baitar: Ein Beitrag zur arabischen Pflanzensynonymik des Mittelalters by Albert Dietrich; Le Taqwim al-Sihha (Tacuini Sanitatis) d’Ibn Butlan: Un traite medical du XIe siècle by Hosam Elkhadem. Isis, 86 (3):472–475, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235036.

Nasr:1995:BRB

[2071] Seyyed Hossein Nasr. Book review: Names, Natures, and Things: The Alchemist Jabir ibn Hayyan and His Kitab al-ahjar by Syed Nomanul-Haq; David E. Pingree. Isis, 86(3):475, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235037.

Kennedy:1995:BRBa

[2072] E. S. Kennedy. Book review: Islamic Astronomy and Medieval Spain by Julio Samso. Isis, 86(3):475–476, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235038.

Latham:1995:BRB

[2073] J. Derek Latham. Book review: Medieval Agriculture and Islamic Science: The Almanac of a Yemeni Sultan by Daniel Martin Varisco. Isis, 86 (3):476–477, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235039.

Scarborough:1995:BRB

[2074] John Scarborough. Book review: Anglo-Norman Medicine. Volume 1: Roger Frugard’s Chirurgia and the Practica brevis of Platearius by Tony Hunt. Isis, 86(3):477–478, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235040.

Busard:1995:BRBa

[2075] H. L. L. Busard. Book review: The Latin Translation of Anaritius’ Commentary on Euclid’s Elements of Geometry, Books I-IV by Anaritius; P. M. J. E. Tummers. Isis, 86(3):478, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235041.

Wippel:1995:BRB

[2076] John F. Wippel. Book review: The Rise and Decline of the Scholastic “Quaestio disputata”: With Special Emphasis on Its Use in the Teaching of Medicine and Science by Brian Lawn. Isis, 86(3):478–479, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235042.

Laird:1995:BRB

[2077] W. R. Laird. Book review: Antonius de Carlensis, O.P.: Four Questions on the Subalternation of the Sciences by Antonius de Carlensis; Steven J. Livesey. Isis, 86(3):479–480, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235043.

North:1995:BRBb

[2078] J. D. North. Book review: History, Prophecy, and the Stars: The Christian Astrology of Pierre d’Ailly, 1350–1420 by Laura Ackerman Smoller. Isis, 86(3):480–481, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235044.

Heath:1995:BRB

[2079] Michael Heath. Book review: Nostradamus astrophile: Les astres et l’astrologie dans la vie et l’oeuvre de Nostradamus by Pierre Brind’Amour. Isis, 86(3):481–482, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235045.

Johnson:1995:BRBa

[2080] Eugene J. Johnson. Book review: The Architectural Drawings of Antonio da Sangallo the Younger and His Circle. Volume 1: Fortifications, Machines, and Festival Architecture by Christoph L. Frommel; Nicholas Adams. Isis, 86(3):482–483, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235046.

Crellin:1995:BRB

[2081] J. K. Crellin. Book review: Hexenkrauter oder Arznei: Die Abtreibungsmittel im 16. und 17. Jahrhundert by Larissa Leibrock-Plehn. Isis, 86(3):483–484, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235047.

Donahue:1995:BRB

[2082] William H. Donahue. Book review: The Music of the Heavens: Kepler’s Harmonic Astronomy by Bruce Stephenson. Isis, 86(3):484–485, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235048.

Pitt:1995:BRB

[2083] Joseph C. Pitt. Book review: The Discovery of Kepler’s Laws: The Interaction of Science, Philosophy, and Religion by Job Kozhamthadam. Isis, 86(3):485–486, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235049.

Finocchiaro:1995:BRBc

[2084] Maurice A. Finocchiaro. Book review: Galileo: For Copernicanism and for the Church by Annibale Fantoli; George V. Coyne; Galileo: A Life by James Reston. Isis, 86(3):486–488, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235050.

Gaukroger:1995:BRB

[2085] Stephen Gaukroger. Book review: Descartes, Spinoza, Leibniz: The Concept of Substance in Seventeenth-Century Metaphysics by R. S. Woolhouse. Isis, 86(3):488, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235051.

Arnold:1995:BRB

[2086] Ken Arnold. Book review: Possessing Nature: Museums, Collecting, and Scientific Culture in Early Modern Italy by Paula Findlen. Isis, 86 (3):488–489, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235052.

Pumfrey:1995:BRB

[2087] Stephen Pumfrey. Book review: John Sergeant and His Circle: A Study of Three Seventeenth-Century English Aristotelians by Dorothea Krook; Beverley C. Southgate. Isis, 86(3):489–490, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235053.

Grell:1995:BRB

[2088] Ole Peter Grell. Book review: Trials of an Ordinary Doctor: Joannes Groenevelt in Seventeenth-Century London by Harold John Cook. Isis, 86 (3):490–491, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235054.

Mahoney:1995:BRBc

[2089] Michael S. Mahoney. Book review: De quadratura arithmetica circuli ellipseos et hyperbolae cujus corollarium est trigonometria sine tabulis by Gottfried Wilhelm Leibniz. Isis, 86(3):491, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235055.

Pinero:1995:BRB

[2090] J. M. Lopez Piñero. Book review: El Jardin Botanico de la Universidad de Granada by Mercedes Fernandez-Carrion; Jose Manuel Garcia Montes; Joaquin Molero Mesa. Isis, 86(3):491–492, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235056.

Hamlin:1995:BRB

[2091] Christopher Hamlin. Book review: Death and the Metropolis: Studies in the Demographic History of London, 1670–1830 by John Landers. Isis, 86(3):492, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235057.

vanBerkel:1995:BRB

[2092] Klaas van Berkel. Book review: Het mechanische van de mechanica: Het krachtbegrip tussen metafysica en mechanica van Newton tot Lagrange by J. Christiaan Boudri. Isis, 86(3):493, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235058.

Hankins:1995:BRB

[2093] Thomas L. Hankins. Book review: Jean Le Rond d’Alembert philosophe by Veronique Le Ru. Isis, 86(3):494, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235059.

Eddy:1995:BRB

[2094] John H. Eddy, Jr. Book review: Interpreting Nature: The Science of Living Form from Linnaeus to Kant by James L. Larson. Isis, 86(3): 494–495, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235060.

Janal:1995:BRB

[2095] Martin J. Janal. Book review: Klassifikation–System–“scala naturae”: Das Ordnen der Objekte in Naturwissenschaft und Pharmazie zwischen 1700 und 1850 by Annette Diekmann. Isis, 86(3):495–496, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235061.

Golinski:1995:BRBa

[2096] Jan Golinski. Book review: Doctor of Society: Thomas Beddoes and the Sick Trade in Late-Enlightenment England by Roy Porter. Isis, 86 (3):496–497, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235062.

Rappaport:1995:BRB

[2097] Rhoda Rappaport. Book review: Giovanni Arduino (1714–1795): Il contributo di uno scienziato veneto al dibattito settecentesco sulle scienze della Terra by Ezio Vaccari. Isis, 86(3):497–498, September 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235063.

Gillispie:1995:BRB

[2098] Charles C. Gillispie. Book review: Saggio sulle macchine in generale by Lazare Carnot; Antonino Drago; Salvatore D. Manno. Isis, 86(3):498, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235064.

Crosland:1995:BRBb

[2099] Maurice Crosland. Book review: Regards sur la science: Le journal scientifique. Isis, 86(3):498–499, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235065.

Clancey:1995:BRB

[2100] Gregory Clancey. Book review: Early American Technology: Making and Doing Things from the Colonial Era to 1850 by Judith A. McGaw. Isis, 86(3):499–500, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235066.

Mehos:1995:BRB

[2101] Donna C. Mehos. Book review: Wetenschap en Intuitie: Het Duitse romantisch-speculatief natuuronderzoek rond 1800 by H. A. M. Snelders. Isis, 86(3):500–501, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235067.

Fraser:1995:BRBa

[2102] Craig Fraser. Book review: Mathematisations: Augustin-Louis Cauchy et l’Ecole Française by Amy Dahan Dalmedico. Isis, 86(3):501–502, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235068.

Rupke:1995:BRBb

[2103] Nicolaas A. Rupke. Book review: Alejandro de Humboldt: La naturaleza, idea y aventura: Libro de la exposicion by Martin Guntau; Peter Hardetert; Martin Pape. Isis, 86(3):502–503, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235069.

Cosslett:1995:BRB

[2104] Tess Cosslett. Book review: Fact and Feeling: Baconian Science and the Nineteenth-Century Literary Imagination by Jonathan Smith. Isis, 86 (3):503–504, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235070.

Nijhuis:1995:BRB

[2105] Ton Nijhuis. Book review: Machines Are Frozen Spirit: The Scientification of Refrigeration and Brewing in the Nineteenth Century — a Weberian Interpretation by Mikael Hard. Isis, 86(3):504, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235071.

Finlayson:1995:BRB

[2106] Alan Christopher Finlayson. Book review: Scaling Fisheries: The Science of Measuring the Effects of Fishing, 1855–1955 by Tim D. Smith. Isis, 86(3):504–505, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235072.

Marshall:1995:BRB

[2107] Marilyn Marshall. Book review: Die innere Seite der Natur: Gustav Theodor Fechners wissenschaftlich-philosophische Weltauffassung by Michael Heidelberger. Isis, 86(3):505–506, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235073.

DeLacy:1995:BRB

[2108] Margaret DeLacy. Book review: Childbed Fever: A Scientific Biography of Ignaz Semmelweis by K. Codell Carter; Barbara R. Carter. Isis, 86 (3):506–507, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235074.

Mitchell:1995:BRB

[2109] Allan Mitchell. Book review: The Making of a Social Disease: Tuberculosis in Nineteenth-Century France by David S. Barnes. Isis, 86(3): 507–508, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235075.

Evans:1995:BRB

[2110] Richard J. Evans. Book review: Die Cholera in Medizin und Pharmazie im Zeitalter des Hygienikers Max von Pettenkofer by Ellen Jahn. Isis, 86(3):508, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235076.

Olby:1995:BRB

[2111] Robert Olby. Book review: Gregor Mendel’s Experiments on Plant Hybrids: A Guided Study by Alain F. Corcos; Floyd V. Monaghan. Isis, 86 (3):508–509, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235077.

Rainger:1995:BRB

[2112] Ronald Rainger. Book review: Strange Genius: The Life of Ferdinand Vandeveer Hayden by Mike Foster. Isis, 86(3):509–510, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235078.

Russett:1995:BRB

[2113] Cynthia Russett. Book review: Healing the Republic: The Language of Health and the Culture of Nationalism in Nineteenth-Century America by Joan Burbick. Isis, 86(3):510–511, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235079.

Christianson:1995:BRBb

[2114] Eric Howard Christianson. Book review: Medical Protestants: The Eclectics in American Medicine, 1825–1939 by John S. Haller. Isis, 86(3): 511–512, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235080.

Cravens:1995:BRB

[2115] Hamilton Cravens. Book review: Inventing the Feeble Mind: A History of Mental Retardation in the United States by James W. Trent. Isis, 86 (3):512–513, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235081.

Weidman:1995:BRB

[2116] Nadine Weidman. Book review: The Roots of Behaviourism by Robert H. Wozniak; Conwy Lloyd Morgan; Jacques Loeb; Max Meyer; John B. Watson. Isis, 86(3):513–514, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235082.

Pyenson:1995:BRA

[2117] Lewis Pyenson. Book review: Albert Einstein, Martin J. Klein, A. J. Kox, Jurgen Renn, Jed Buchwald, Jean Eisenstaedt, Don Howard, John Norton, Tilman Sauer: The Collected Papers of Albert Einstein. Volume 3: The Swiss Years: Writings, 1909–1911. Albert Einstein and Anna Beck: The Collected Papers of Albert Einstein. Volume 3: The Swiss Years: Writings, 1909–1911. Albert Einstein, Martin J. Klein, A. J. Kox, Robert Schulmann, Paolo Brenni, Klaus Hentschel, Jurgen Renn, and Laura Ruetsche:The Collected Papers of Albert Einstein. Volume 5: The Swiss Years: Correspondence, 1902–1914. Isis, 86(3):514–515, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235083.

MacKinnon:1995:BRB

[2118] Edward MacKinnon. Book review: Atoms, Metaphors, and Paradoxes: Niels Bohr and the Construction of a New Physics by Sandro Petruccioli. Isis, 86(3):515, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235084.

Eisberg:1995:BRB

[2119] Joann Eisberg. Book review: The Milky Way Galaxy and Statistical Cosmology, 1890–1924 by Erich Robert Paul. Isis, 86(3):515–516, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235085.

Beaver:1995:BRB

[2120] Donald DeB. Beaver. Book review: Edwin Hubble: The Discoverer of the Big Bang Universe by Alexander S. Sharov; Igor D. Novikov; Vitaly Kisin. Isis, 86(3):516–517, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235086.

Porter:1995:BRBb

[2121] Theodore M. Porter. Book review: Edgeworth on Chance, Economic Hazard, and Statistics by Francis Ysidro Edgeworth; Philip Mirowski. Isis, 86(3):517–518, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235087.

Johnson:1995:BRBb

[2122] Jeffrey A. Johnson. Book review: Fritz Haber: Chemiker, Nobelpreistrager, Deutscher, Jude by Dietrich Stoltzenberg. Isis, 86(3):518–519, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235088.

Hentschel:1995:BRB

[2123] Klaus Hentschel. Book review: Philipp Fauth–Leben und Werk: Aus dem autobiographischen Nachlass zusammengestellt von Hermann Fauth by Philipp Fauth; Freddy Litten. Isis, 86(3):519–520, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235089.

Kragh:1995:BRB

[2124] Helge Kragh. Book review: Matvei Petrovich Bronstein and Soviet Theoretical Physics in the Thirties by Gennady E. Gorelik; Victor Ya. Frenkel; Valentina M. Levina. Isis, 86(3):520, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235090.

Goldberg:1995:BRB

[2125] Stanley Goldberg. Book review: Critical Assembly: A Technical History of Los Alamos during the Oppenheimer Years, 1943–1945 by Lillian Hoddeson; Paul W. Henriksen; Roger A. Meade; Catherine Westfall. Isis, 86 (3):520–522, September 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235091.

Smocovitis:1995:BRBa

[2126] Vassiliki Betty Smocovitis. Book review: The Evolution of Theodosius Dobzhansky: Essays on His Life and Thought in Russia and America by Mark B. Adams. Isis, 86(3):522–523, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235092.

Taylor:1995:BRB

[2127] Peter J. Taylor. Book review: A History of the Ecosystem Concept in Ecology: More than the Sum of the Parts by Frank Benjamin Golley. Isis, 86(3):523–524, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235093.

Sadowsky:1995:BRB

[2128] Jonathan Sadowsky. Book review: Colonial Psychiatry and the “African Mind.” by Jock McCulloch. Isis, 86(3):524–525, September 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235094.

Smith:1995:BRBb

[2129] Robert W. Smith. Book review: Suddenly, Tomorrow Came…: A History of the Johnson Space Center by Henry C. Dethloff; The Missile and Space Race by Alan J. Levine. Isis, 86(3):525–526, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235095.

Lewenstein:1995:BRBc

[2130] Bruce V. Lewenstein. Book review: Selling Outer Space: The Kennedy Administration, the Media, and Funding for Project Apollo, 1961–1963 by James Lee Kauffman. Isis, 86(3):526, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235096.

Turner:1995:BRB

[2131] R. Steven Turner. Book review: Wissenschaft, Universität, Professionen: Soziologische Analysen by Rudolf Stichweh. Isis, 86(3):526–527, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235097.

Osborne:1995:BRB

[2132] Michael A. Osborne. Book review: Les professeurs du Conservatoire National des Arts et Metiers: Dictionnaire biographique, 1794–1955 by Claudine Fontanon; Andre Grelon. Isis, 86(3):527–528, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235098.

Anonymous:1995:Cc

[2133] Anonymous. Collections. Isis, 86(3):528–534, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235099.

Anonymous:1995:BMd

[2134] Anonymous. Back matter. Isis, 86(3):535–540, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235100.

Anonymous:1995:FMd

[2135] Anonymous. Front matter. Isis, 86(3):??, September 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235018.

Silliman:1995:HAC

[2136] Robert H. Silliman. The Hamlet Affair: Charles Lyell and the North Americans. Isis, 86(4):541–561, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235376.

Moreton:1995:BGR

[2137] Jennifer Moreton. Before Grosseteste: Roger of Hereford and calendar reform in Eleventh- and Twelfth-Century England. Isis, 86(4):562–586, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235377.

Williams:1995:FNG

[2138] Burma P. Williams and Richard S. Williams. Finger numbers in the Greco–Roman world and the Early Middle Ages. Isis, 86(4):587–608, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235378.

Weikart:1995:RDD

[2139] Richard Weikart. A recently discovered Darwin letter on social Darwinism. Isis, 86(4):609–611, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235379.

Rheinberger:1995:WDC

[2140] Hans-Jorg Rheinberger. When did Carl Correns read Gregor Mendel’s paper? A research note. Isis, 86(4):612–616, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235380.

Leavitt:1995:LE

[2141] Judith Walzer Leavitt. Letter to the Editor. Isis, 86(4):617–618, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235381.

Nelkin:1995:BRB

[2142] Dorothy Nelkin. Book review: Social Bodies: Science, Reproduction, and Italian Modernity. Isis, 86(4):619–621, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235382.

Nye:1995:BRB

[2143] Robert A. Nye. Book review: Rewriting the Soul: Multiple Personality and the Sciences of Memory. Isis, 86(4):622–624, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235383.

Lemay:1995:BRB

[2144] Helen Rodnite Lemay. Book review: The History of Obstetrics and Gynaecology by Michael J. O’Dowd; Elliot E. Philipp. Isis, 86(4):624–625, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235384.

Livingstone:1995:BRB

[2145] David N. Livingstone. Book review: The History of Cartography. Volume 2, Book 2: Cartography in the Traditional East and Southeast Asian Societies by J. B. Harley; David Woodward. Isis, 86(4):625–626, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235385.

Hills:1995:BRB

[2146] Richard L. Hills. Book review: Energy in World History by Vaclav Smil. Isis, 86(4):626–627, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235386.

Williams:1995:BRBc

[2147] Elizabeth A. Williams. Book review: Le destin du mammouth by Claudine Cohen. Isis, 86(4):627–628, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235387.

Apple:1995:BRB

[2148] Rima D. Apple. Book review: Protein and Energy: A Study of Changing Ideas in Nutrition by Kenneth J. Carpenter. Isis, 86(4):628, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235388.

Lloyd:1995:BRBc

[2149] Geoffrey Lloyd. Book review: Les papyrus medicaux de l’Egypte Pharaonique: Traduction integrale et commentaire by Thierry Bardinet.

Isis, 86(4):628–629, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235389.

Busard:1995:BRBb

[2150] H. L. L. Busard. Book review: Science in Western and Eastern Civilization in Carolingian Times by Paul Leo Butzer; Dietrich Lohrmann. Isis, 86(4):629–630, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235390.

Kennedy:1995:BRBb

[2151] E. S. Kennedy. Book review: Optics, Astronomy, and Logic: Studies in Arabic Science and Philosophy by A. I. Sabra. Isis, 86(4):630–631, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235391.

North:1995:BRBc

[2152] J. D. North. Book review: Medieval Lunar Astrology: A Collection of Representative Middle English Texts by Laurel Means. Isis, 86(4):631– 632, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235392.

Murdoch:1995:BRB

[2153] John E. Murdoch. Book review: Argumentationstheorie: Scholastische Forschungen zu den logischen und semantischen Regelen korrekten Folgerns by Klaus Jacobi; Sophisms in Medieval Logic and Grammar by Stephen Read. Isis, 86(4):632–633, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235393.

Otte:1995:BRB

[2154] James K. Otte. Book review: Die “Mainauer Naturlehre” im Kontext der Wissenschaftsgeschichte by Martin Mosimann. Isis, 86(4):633–634, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235394.

Hall:1995:BRB

[2155] A. Rupert Hall. Book review: Le macchine del Re: Il Texaurus Regis Francie di Guido da Vigevano by Giustina Ostuni. Isis, 86(4):634–635, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235395.

Fine:1995:BRB

[2156] Eve Fine. Book review: Contraception and Abortion in Nineteenth-Century America by Janet Farrell Brodie; Contraception and Abortion from the Ancient World to the Renaissance by John M. Riddle. Isis, 86 (4):635–636, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235396.

Cueto:1995:BRBb

[2157] Marcos Cueto. Book review: La conquista de la naturaleza americana by Raquel Alvarez Pelaez. Isis, 86(4):636–637, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235397.

Goodman:1995:BRB

[2158] David Goodman. Book review: La experiencia americana y la terapeutica en los Secretos de chirurgia (1567) de Pedro Arias de Benavides by Jose Luis Fresquet Febrer; Nuevos materiales y noticias sobre la Historia de las plantas de Nueva España, de Francisco Hernandez by Jose Maria Lopez Piñero; Jose Pardo Tomas; Las primeras noticias sobre plantas americanas en las relaciones de viajes y cronicas de Indias (1493–1553) by Jose Pardo Tomas; Maria Luz Lopez Terrada. Isis, 86(4):637–638, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235398.

Smith:1995:BRBc

[2159] Pamela H. Smith. Book review: European Encounters with the New World: From Renaissance to Romanticism by Anthony Pagden; Keys to the Encounter: A Library of Congress Resource Guide for the Study of the Age of Discovery by Louis De Vorsey. Isis, 86(4):638–640, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235399.

Hoeniger:1995:BRB

[2160] F. David Hoeniger. Book review: The Renaissance Print: 1470–1550 by David Landau; Peter W. Parshall. Isis, 86(4):640–641, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235400.

Newman:1995:BRBb

[2161] William R. Newman. Book review: Parerga Paracelsica: Paracelsus in Vergangenheit und Gegenwart by Joachim Telle. Isis, 86(4):641–642, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235401.

Voigts:1995:BRBb

[2162] Linda Ehrsam Voigts. Book review: The Authorship of The Equatorie of the Planetis by Kari Anne Rand Schmidt. Isis, 86(4):642, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235402.

Freudenthal:1995:BRBb

[2163] Gideon Freudenthal. Book review: On the Shoulders of Merchants: Exchange and the Mathematical Conception of Nature in Early Modern Europe by Richard W. Hadden. Isis, 86(4):642–643, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235403.

Rees:1995:BRB

[2164] Graham Rees. Book review: Novum organum: With Other Parts of the Great Instauration by Francis Bacon; Peter Urbach; John Gibson. Isis, 86 (4):643–644, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235404.

Maistrellis:1995:BRB

[2165] Nicholas Maistrellis. Book review: L’arithmétique de la Grace: Pascal et les carrés magiques by Jacques Darriulat. Isis, 86(4):644, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235405.

Hatch:1995:BRB

[2166] Robert A. Hatch. Book review: Ismael Boulliau (1605–1694): Astronome, epistolier, nouvelliste et intermediaire scientifique: Ses rapports avec les milieux du “libertinage erudit.” by Henk J. M. Nellen. Isis, 86 (4):645–646, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235406.

Segre:1995:BRB

[2167] Michael Segre. Book review: Out of Galileo: The Science of Waters, 1628–1718 by Cesare S. Maffioli. Isis, 86(4):646–647, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235407.

Tribe:1995:BRB

[2168] Keith Tribe. Book review: The Business of Alchemy: Science and Culture in the Holy Roman Empire by Pamela H. Smith. Isis, 86(4):647–648, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235408.

Golinski:1995:BRBb

[2169] Jan Golinski. Book review: Gehennical Fire: The Lives of George Starkey, an American Alchemist in the Scientific Revolution by William R. Newman. Isis, 86(4):648–649, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235409.

Davis:1995:BRB

[2170] Edward B. Davis. Book review: The Diffident Naturalist: Robert Boyle and the Philosophy of Experiment by Rose-Mary Sargent. Isis, 86(4): 649, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235410.

Stewart:1995:BRB

[2171] Ian G. Stewart. Book review: The Royal Society and Its Fellows, 1660– 1700: The Morphology of an Early Scientific Institution by Michael Hunter. Isis, 86(4):649–650, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235411.

Gale:1995:BRB

[2172] George Gale. Book review: The Cambridge Companion to Leibniz by Nicholas Jolley. Isis, 86(4):650–651, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235412.

Garber:1995:BRB

[2173] Daniel Garber. Book review: G. W. Leibniz: Critical Assessments by Roger Woolhouse. Isis, 86(4):651–652, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235413.

Steffens:1995:BRB

[2174] Henry Steffens. Book review: Optics in the Age of Euler: Conceptions of the Nature of Light, 1700–1795 by Casper Hakfoort. Isis, 86(4):652–653, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235414.

Glick:1995:BRB

[2175] Thomas F. Glick. Book review: Francisco Jose de Caldas: A Scientist at Work in Nueva Granada by John Wilton Appel. Isis, 86(4):653, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235415.

Rosner:1995:BRB

[2176] Lisa Rosner. Book review: Philosophic Whigs: Medicine, Science, and Citizenship in Edinburgh, 1789–1848 by L. S. Jacyna. Isis, 86(4):653– 654, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235416.

Broman:1995:BRB

[2177] Thomas Broman. Book review: Notable Encyclopedias of the Late Eighteenth Century: Eleven Successors of the Encyclopedie by Frank A. Kafker. Isis, 86(4):654–655, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235417.

Hahn:1995:BRB

[2178] Roger Hahn. Book review: The Scientific Library of Horace-Benedict de Saussure (1797): Annotated Catalog of an Eighteenth-Century Bibliographic and Historic Treasure by Albert V. Carozzi; Gerda Bouvier. Isis, 86(4):655–656, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235418.

Alsop:1995:BRB

[2179] J. D. Alsop. Book review: James Cook and the Conquest of Scurvy by Francis E. Cuppage. Isis, 86(4):656, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235419.

Appel:1995:BRB

[2180] Toby A. Appel. Book review: Science and the Practice of Medicine in the Nineteenth Century by William F. Bynum. Isis, 86(4):656–658, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235420.

Holmes:1995:BRB

[2181] Frederic L. Holmes. Book review: From Chemical Philosophy to Theoretical Chemistry: Dynamics of Matter and Dynamics of Disciplines, 1800– 1950 by Mary Jo Nye. Isis, 86(4):658–659, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235421.

Williams:1995:BRBd

[2182] L. Pearce Williams. Book review: Les sciences dans l’enseignement secondaire français: Textes officiels. Volume 1: 1789–1914 by Bruno Belhoste; Claudette Balpe; Thierre Laporte. Isis, 86(4):659–660, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235422.

Fraser:1995:BRBb

[2183] Craig Fraser. Book review: Theoria combinationis observationum erroribus minimis obnoxiae by Carl Friedrich Gauss; G. W. Stewart. Isis, 86(4):660–661, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235423.

Kremer:1995:BRBb

[2184] Richard L. Kremer. Book review: Science, Medicine, and the State in Germany: The Case of Baden, 1815–1871 by Arleen Marcia Tuchman. Isis, 86(4):661–662, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235424.

Richmond:1995:BRB

[2185] Marsha Richmond. Book review: Studia Fribergensia: Vortrage des Alexander-von-Humboldt-Kolloquiums in Freiberg vom 8. bis 10. November 1991 aus Anlass des 200. Jahrestages von A. V. Humboldts Studienbeginn an der Bergakademie Freiberg; Briefwechsel zwischen Alexander von Humboldt und Friedrich Wilhelm Bessel by Hans-Joachim Felber; Alexander von Humboldt; Friedrich Wilhelm Bessel. Isis, 86(4):662–663, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235425.

Lucier:1995:BRB

[2186] Paul Lucier. Book review: Henry Darwin Rogers, 1808–1866: American Geologist by Patsy Gerstner. Isis, 86(4):663–664, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235426.

Hatfield:1995:BRB

[2187] Gary Hatfield. Book review: In the Eye’s Mind: Vision and the Helmholtz-Hering Controversy by R. Steven Turner. Isis, 86(4):664–665, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235427.

Jacyna:1995:BRBb

[2188] L. S. Jacyna. Book review: Aux origines du cerveau moderne: Localisations, langage et memoire dans l’oeuvre de Charcot by Jacques Gasser. Isis, 86(4):665–666, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235428.

Berkowitz:1995:BRB

[2189] Michael Berkowitz. Book review: Defenders of the Race: Jewish Doctors and Racial Science in Fin-de-Siècle Europe by John M. Efron. Isis, 86 (4):666–667, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235429.

Richards:1995:BRB

[2190] Robert J. Richards. Book review: The Temptations of Evolutionary Ethics by Paul Lawrence Farber. Isis, 86(4):667–668, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235430.

Teich:1995:BRB

[2191] Mikulas Teich. Book review: Hans Krebs by Frederic Lawrence Holmes. Isis, 86(4):668–670, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235431.

Ihde:1995:BRB

[2192] Aaron J. Ihde. Book review: American Chemists and Chemical Engineers. Volume 2 by Wyndham D. Miles; Robert F. Gould. Isis, 86(4): 670–671, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235432.

Gellner:1995:BRB

[2193] Ernest Gellner. Book review: The Rise and Crisis of Psychoanalysis in America: Freud and the Americans, 1917–1985 by Nathan G. Hale. Isis, 86(4):671, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235433.

Zabell:1995:BRB

[2194] S. L. Zabell. Book review: Creating Modern Probability: Its Mathematics, Physics, and Philosophy in Historical Perspective by Jan von Plato.

Isis, 86(4):671–672, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235434.

Reich:1995:BRB

[2195] Karin Reich. Book review: Briefwechsel zwischen Karl Weierstrass und Sofja Kowalewskaja by Karl Weierstrass; Sofja Kowalewskaja; Reinhard Bolling. Isis, 86(4):672–673, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235435.

Lewis:1995:BRBb

[2196] Albert C. Lewis. Book review: Abraham Robinson: The Creation of Nonstandard Analysis, a Personal and Mathematical Odyssey by Joseph Warren Dauben. Isis, 86(4):673–674, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235436.

Hovis:1995:BRBb

[2197] R. Corby Hovis. Book review: Unified Field Theories in the First Third of the Twentieth Century by Vladimir P. Vizgin; Julian B. Barbour. Isis, 86(4):674–675, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235437.

Kaiser:1995:BRB

[2198] David Kaiser. Book review: Eddington’s Search for a Fundamental Theory: A Key to the Universe by C. W. Kilmister. Isis, 86(4):675, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235438.

Friedman:1995:BRBb

[2199] Michael Friedman. Book review: Reminiscences of the Vienna Circle and the Mathematical Colloquium by Karl Menger; Louise Golland; Brian McGuinness; Abe Sklar. Isis, 86(4):675–676, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235439.

Kline:1995:BRBc

[2200] Ronald Kline. Book review: Entitled to Power: Farm Women and Technology, 1913–1963 by Katherine Jellison. Isis, 86(4):676–677, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235440.

Newell:1995:BRB

[2201] Julie R. Newell. Book review: Cracking Rocks and Defending Democracy: Kirtley Fletcher Mather: Scientist, Teacher, Social Activist, 1888–1978 by Kennard Baker Bork. Isis, 86(4):677–678, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235441.

Swetlitz:1995:BRB

[2202] Marc Swetlitz. Book review: If I Am to Be Remembered: The Life and Work of Julian Huxley with Selected Correspondence by Krishna R. Dronamraju; Julian Huxley: Biologist and Statesman of Science by C. Kenneth Waters; Albert Van Helden. Isis, 86(4):678–679, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235442.

Stocking:1995:BRB

[2203] George W. Stocking, Jr. Book review: The King of the World in the Land of the Pygmies by Joan Mark. Isis, 86(4):679–680, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235443.

Weiss:1995:BRB

[2204] Sheila Faith Weiss. Book review: Death and Deliverance: “Euthanasia” in Germany, 1900–1945 by Michael Burleigh. Isis, 86(4):680–681, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235444.

Dupree:1995:BRB

[2205] A. Hunter Dupree. Book review: The First Nuclear Era: The Life and Times of a Technological Fixer by Alvin M. Weinberg. Isis, 86(4):681– 682, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235445.

Badash:1995:BRB

[2206] Lawrence Badash. Book review: Stalin and the Bomb: The Soviet Union and Atomic Energy, 1939–1956 by David Holloway. Isis, 86(4):682–683, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235446.

Krementsov:1995:BRB

[2207] Nikolai Krementsov. Book review: Lysenko and the Tragedy of Soviet Science by Valery N. Soyfer; Leo Gruliow; Rebecca Gruliow. Isis, 86 (4):683–685, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235447.

Smocovitis:1995:BRBb

[2208] Vassilki Betty Smocovitis. Book review: Population Genetics, Molecular Evolution, and the Neutral Theory: Selected Papers by Motoo Kimura; Naoyuki Takahata. Isis, 86(4):685, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235448.

Creager:1995:BRB

[2209] Angela N. H. Creager. Book review: The Uses of Life: A History of Biotechnology by Robert Bud. Isis, 86(4):685–687, December 1995. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235449.

Lederer:1995:BRB

[2210] Susan E. Lederer. Book review: Factories of Death: Japanese Biological Warfare 1932–45 and the American Cover-Up by Sheldon H. Harris. Isis, 86(4):687, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235450.

Tjossem:1995:BRB

[2211] Sara F. Tjossem. Book review: Always, Rachel: The Letters of Rachel Carson and Dorothy Freeman, 1952–1964 by Rachel Carson; Dorothy Freeman; Martha Freeman. Isis, 86(4):687–688, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235451.

Junod:1995:BRB

[2212] Suzanne White Junod. Book review: Cancer from Beef: DES, Federal Food Regulation, and Consumer Confidence by Alan I. Marcus. Isis, 86 (4):688–689, December 1995. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235452.

Lindee:1995:BRBb

[2213] Susan Lindee. Book review: Elements of Controversy: The Atomic Energy Commission and Radiation Safety in Nuclear Weapons Testing, 1947–1974 by Barton C. Hacker. Isis, 86(4):689–690, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235453.

Alexander:1995:BRB

[2214] Thomas G. Alexander. Book review: A Conspiracy of Optimism: Management of the National Forests since World War Two by Paul W. Hirt. Isis, 86(4):690–691, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235454.

Mahoney:1995:BRBd

[2215] Michael S. Mahoney. Book review: Landmarks in Digital Computing: A Smithsonian Pictorial History by Peggy A. Kidwell; Paul E. Ceruzzi. Isis, 86(4):691–692, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235455.

Dallal:1995:BRB

[2216] Ahmad S. Dallal. Book review: Science and Muslim Societies by Nasim Butt. Isis, 86(4):692, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235456.

Anonymous:1995:Cd

[2217] Anonymous. Collections. Isis, 86(4):692–699, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235457.

Anonymous:1995:BMe

[2218] Anonymous. Back matter. Isis, 86(4):700–736, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235458.

Anonymous:1995:VI

[2219] Anonymous. Volume information. Isis, 86(4):711–735, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235374.

Anonymous:1995:FMe

[2220] Anonymous. Front matter. Isis, 86(4):??, December 1995. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235375.

Neu:1996:CBH

[2221] John Neu. Current bibliography of the history of science and its cultural influences, 1996. Isis, 87(??):1–333, ???? 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235275.

Anonymous:1996:BMa

[2222] Anonymous. Back matter. Isis, 87(??):??, ???? 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235276.

Anonymous:1996:FMa

[2223] Anonymous. Front matter. Isis, 87(??):i–xvi, ???? 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235274.

Humphreys:1996:KDD

[2224] Margaret Humphreys. Kicking a dying dog: DDT and the demise of malaria in the American South, 1942–1950. Isis, 87(1):1–17, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235732.

Kojevnikov:1996:PSA

[2225] Alexei Kojevnikov. President of Stalin’s Academy: The mask and responsibility of Sergei Vavilov. Isis, 87(1):18–50, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235733.

Pauly:1996:BMJ

[2226] Philip J. Pauly. The beauty and menace of the Japanese cherry trees: Conflicting visions of American ecological independence. Isis, 87(1):51– 73, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235734.

Evans:1996:FIA

[2227] James Evans. Fraud and illusion in the anti-Newtonian rear guard: The Coultaud–Mercier Affair and Bertier’s experiments, 1767–1777. Isis, 87 (1):74–107, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235735.

McCready:1996:IAS

[2228] William D. McCready. Isidore, the antipodeans, and the shape of the Earth. Isis, 87(1):108–127, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235736.

Anderson:1996:LE

[2229] Leland Anderson, Sungook Hong, Gennady Gorelik, and Helge Kragh. Letters to the Editor. Isis, 87(1):128–130, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235737.

Feingold:1996:BRB

[2230] Mordechai Feingold. Book review: A Social History of Truth: Civility and Science in Seventeenth-Century England. Isis, 87(1):131–139, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235738.

Weiner:1996:BRB

[2231] Douglas R. Weiner. Book review: Concepts of Symbiogenesis: A Historical and Critical Study of the Research of Russian Botanists. Isis, 87(1):140–141, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235739.

Whelan:1996:BRB

[2232] Mary K. Whelan. Book review: Women in Archaeology by Cheryl Claassen. Isis, 87(1):142–143, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235740.

Cowan:1996:BRB

[2233] Ruth Schwartz Cowan. Book review: Mothers and Daughters of Invention: Notes for a Revised History of Technology by Autumn Stanley. Isis, 87(1):143–144, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235741.

Marx:1996:BRB

[2234] Leo Marx. Book review: The Machine in America: A Social History of Technology by Carroll W. Pursell. Isis, 87(1):144–145, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235742.

Hetherington:1996:BRBa

[2235] Norriss S. Hetherington. Book review: Modern Theories of the Universe from Herschel to Hubble by Michael J. Crowe. Isis, 87(1):145–146, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235743.

Marston:1996:BRB

[2236] V. Paul Marston. Book review: The Biblical Flood: A Case Study of the Church’s Response to Extrabiblical Evidence by Davis A. Young. Isis, 87(1):146–147, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235744.

Barkan:1996:BRB

[2237] Diana Barkan. Book review: Histoire de la thermochimie: Prelude a la thermodynamique chimique by Louis Medard; Henri Tachoire. Isis, 87(1):147–148, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235745.

Sigurdsson:1996:BRB

[2238] Skuli Sigurdsson. Book review: Mathematik und Geisteswissenschaft: Mathematische Einfuhrung in die Philosophie als Geisteswissenschaft in Ankupfung an Plato, Cusanus, Goethe, Hegel und Steiner by Renatus Ziegler. Isis, 87(1):148–149, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235746.

Prasad:1996:BRB

[2239] Archana Prasad. Book review: Green Imperialism: Colonial Expansion, Tropical Island Edens, and the Origins of Environmentalism, 1600–1860 by Richard H. Grove. Isis, 87(1):149, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235747.

Brooks:1996:BRB

[2240] E. Bruce Brooks. Book review: Iron and Steel in Ancient China by Donald B. Wagner. Isis, 87(1):149–150, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235748.

Laird:1996:BRB

[2241] W. R. Laird. Book review: Le origini del concetto di macchina by Gianni Micheli. Isis, 87(1):150–151, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235749.

Federspiel:1996:BRB

[2242] Michel Federspiel. Book review: Les elements (Traduits du texte de Heiberg). Volume 2: Livres V-IX by Euclide d’Alexandrie; Bernard Vitrac. Isis, 87(1):151–152, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235750.

Scarborough:1996:BRB

[2243] John Scarborough. Book review: Theophrastus of Eresus: Sources for His Life, Writings, Thought, and Influence. Commentary Volume 5: Sources on Biology (Human Physiology, Living Creatures, Botany: Texts 328– 435) by R. W. Sharples. Isis, 87(1):152–153, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235751.

Sabra:1996:BRB

[2244] A. I. Sabra. Book review: Aristotle’s Physics and Its Reception in the Arabic World: With an Edition of the Unpublished Parts of Ibn Bajja’s Commentary on the Physics by P. Lettinck. Isis, 87(1):153–154, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235752.

Ragep:1996:BRB

[2245] F. Jamil Ragep. Book review: A History of Arabic Astronomy: Planetary Theories during the Golden Age of Islam by George Saliba. Isis, 87(1): 154–155, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235753.

VanEss:1996:BRB

[2246] Josef Van Ess. Book review: The Physical Theory of Kalam: Atoms, Space, and Void in Basrian Mutazili Cosmology by Alnoor Dhanani. Isis, 87(1):155–156, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235754.

Eastwood:1996:BRB

[2247] Bruce Eastwood. Book review: Eclipses, comètes, autres phénomènes célestes et tremblements de terre au moyen age: Enquête sur six siècles d’historiographie médiévale dans les limites de la Belgique actuelle (600– 1200) by Isabelle Draelants. Isis, 87(1):156–157, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235755.

North:1996:BRB

[2248] J. D. North. Book review: Pelerin de Prusse on the Astrolabe: Text and Translation of His Practique de astralable by Pelerin de Prusse; Edgar Laird; Robert Fischer. Isis, 87(1):157, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235756.

McVaugh:1996:BRBa

[2249] Michael R. McVaugh. Book review: Historia naturalis by Johannis Aegidii Zamorensis; Avelino Dominguez Garcia; Luis Garcia Ballester. Isis, 87(1):158, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235757.

French:1996:BRB

[2250] Roger French. Book review: Medicine before the Plague: Practitioners and Their Patients in the Crown of Aragon, 1285–1345 by Michael R. McVaugh. Isis, 87(1):158–159, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235758.

Adamson:1996:BRB

[2251] Melitta Weiss Adamson. Book review: Das Regimen Sanitatis Konrads von Eichstatt: Quellen-Texte-Wirkungsgeschichte by Christa Hagenmeyer. Isis, 87(1):159–160, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235759.

Crossgrove:1996:BRB

[2252] William Crossgrove. Book review: Berthold Schwarz: Chemie und Waffentechnik im 15. Jahrhundert by Gerhard W. Kramer. Isis, 87(1):160– 161, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235760.

Wallace:1996:BRB

[2253] William A. Wallace. Book review: Speculum philosophiae medii aevi: Die Handschriftensammlung des Dominikaners Georg Schwartz (d. nach 1484) by Maarten J. F. M. Hoenen. Isis, 87(1):161–162, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235761.

Methuen:1996:BRB

[2254] Charlotte Methuen. Book review: Mathematik und Astronomie an der Universität Ingolstadt im 15. und 16. Jahrhundert by Christoph Schoner. Isis, 87(1):162–163, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235762.

Pinero:1996:BRBa

[2255] Jose M. Lopez Piñero. Book review: Ciencia y tecnica en la metropolizacion de America by Jose Sala Catala. Isis, 87(1):163–164, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235763.

Webster:1996:BRBa

[2256] Charles Webster. Book review: Resultate und desiderate der Paracelsus-Forschung by Peter Dilg; Hartmut Rudolph. Isis, 87(1):164–165, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235764.

Nutton:1996:BRBa

[2257] Vivian Nutton. Book review: Andreas Vesalius, a Belgian Census: Contributions towards a New Edition of H. W. Cushing’s Bibliography by Elly Cockx-Indestege. Isis, 87(1):165, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235765.

VanHelden:1996:BRB

[2258] Albert Van Helden. Book review: Kepler by Max Caspar; C. Doris Hellman. Isis, 87(1):165–166, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235766.

Westman:1996:BRB

[2259] Robert S. Westman. Book review: Cremonini e Galilei inquisiti a Padova nel 1604: Nuovi documenti d’archivio by Antonino Poppi. Isis, 87(1): 166–167, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235767.

DePace:1996:BRB

[2260] Anna De Pace. Book review: Tra scienza e immaginazione: Le matematiche presso il Collegio gesuitico napoletano (1552–1670 ca.) by Romano Gatto. Isis, 87(1):167–168, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235768.

Lennon:1996:BRB

[2261] Thomas M. Lennon. Book review: Divine Will and the Mechanical Philosophy: Gassendi and Descartes on Contingency and Necessity in the Created World by Margaret J. Osler. Isis, 87(1):168–169, March 1996.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235769.

Smith:1996:BRBa

[2262] A. Mark Smith. Book review: Descartes on Seeing: Epistemology and Visual Perception. (Journal of the History of Philosophy Monograph Series) by Celia Wolf-Devine. Isis, 87(1):169–170, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235770.

Sargent:1996:BRBa

[2263] Rose-Mary Sargent. Book review: The Invisible World: Early Modern Philosophy and the Invention of the Microscope by Catherine Wilson. Isis, 87(1):170–171, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235771.

Gascoigne:1996:BRB

[2264] John Gascoigne. Book review: Change and Stability: Natural Philosophy at the Academy of Turku (1640–1713) by Maija Kallinen. Isis, 87(1): 171, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235772.

Ranea:1996:BRBa

[2265] Alberto Guillermo Ranea. Book review: Leibniz korrespondiert mit China: Der Briefwechsel mit den Jesuitenmissionaren (1689–1714) by Gottfried Wilhelm Leibniz; Rita Widmaier. Isis, 87(1):172, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235773.

Mahoney:1996:BRBa

[2266] Michael S. Mahoney. Book review: The Investigation of Difficult Things: Essays on Newton and the History of the Exact Sciences in Honour of D. T. Whiteside by P. M. Harman; Alan E. Shapiro. Isis, 87(1):172– 174, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235774.

Meli:1996:BRB

[2267] Domenico Bertoloni Meli. Book review: Fits, Passions, and Paroxysms: Physics, Method, and Chemistry and Newton’s Theories of Colored Bodies and Fits of Easy Reflection by Alan E. Shapiro. Isis, 87(1):174–175, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235775.

Stewart:1996:BRB

[2268] Larry Stewart. Book review: Newton and the Culture of Newtonianism by Betty Jo Teeter Dobbs; Margaret C. Jacob. Isis, 87(1):175, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235776.

Findlen:1996:BRBa

[2269] Paula Findlen. Book review: The Intellectual Roots of the Italian Englightenment: Newtonian Science, Religion, and Politics in the Early Eighteenth Century by Vincenzo Ferrone; Sue Brotherton. Isis, 87(1): 175–176, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235777.

Shea:1996:BRBa

[2270] William R. Shea. Book review: Carteggio (1708–1714) by Guido Grandi; Jacob Hermann; Silvia Mazzone; Clara Silvia Roero; La forma della quantita: Analisi algebrica e analisi superiore: Il problema dell’unita della matematica nel secolo dell’illuminismo by Marco Panza. Isis, 87 (1):176–177, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235778.

Heilbron:1996:BRBa

[2271] J. L. Heilbron. Book review: Carteggio “de Diurno Terrae Motu”: Canterzani, Isolani, Matteucci, Bonfioli Malvezzi, Caldani, Calandrelli, Bonati by Giambattista Guglielmini; Maria Teresa Borgato; Alessandra Fiocca. Isis, 87(1):177, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235779.

Glick:1996:BRBa

[2272] Thomas F. Glick. Book review: El arcano de la quina by Jose Celestino Mutis; Quinologia and Suplemento a la quinologia by Hipolito Ruiz; Noticias de el caphe: Discvrso philosophico by Juan de Tariol. Isis, 87(1): 178, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235780.

Fox:1996:BRB

[2273] Robert Fox. Book review: Weighing Imponderables and Other Quantitative Science around 1800. (Historical Studies in the Physical and Biological Sciences, Suppl. Vol. 24, Pt. 1.) by J. L. Heilbron. Isis, 87(1): 178–179, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235781.

Palmer:1996:BRB

[2274] Louise Y. Palmer. Book review: A New Course in Chemistry: Lavoisier’s First Chemical Paper by Marco Beretta. Isis, 87(1):179–180, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235782.

Donovan:1996:BRB

[2275] Arthur Donovan. Book review: A Scientific Correspondence during the Chemical Revolution: Louis-Bernard Guyton de Morveau and Richard Kirwan, 1782–1802 by Louis-Bernard Guyton de Morveau; Richard Kirwan; Emmanuel Grison; Michele Goupil; Patrice Bret. Isis, 87(1):180– 181, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235783.

Simpson:1996:BRB

[2276] A. D. C. Simpson. Book review: Public and Private Science: The King George III Collection by Alan Q. Morton; Jane A. Wess. Isis, 87(1):181– 182, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235784.

Fitzpatrick:1996:BRB

[2277] Martin Fitzpatrick and E. B. Fryde. Book review: Dogmatismo ed eresia nella scienza: Joseph Priestley by Pierluigi Barrotta. Isis, 87(1):182– 183, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235785.

Marland:1996:BRB

[2278] Hilary Marland. Book review: Making a Medical Living: Doctors and Patients in the English Market for Medicine, 1720–1911 by Anne Digby. Isis, 87(1):183–184, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235786.

Vicziany:1996:BRB

[2279] Marika Vicziany. Book review: Colonizing the Body: State Medicine and Epidemic Disease in Nineteenth-Century India by David Arnold; Public Health in British India: Anglo-Indian Preventive Medicine, 1859–1914 by Mark Harrison. Isis, 87(1):184–186, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235787.

Meinel:1996:BRB

[2280] Christoph Meinel. Book review: Justus von Liebig und Hermann Kolbe in ihren Briefen, 1846–1873 by Justus von Liebig; Hermann Kolbe; Alan J. Rocke; Emil Heuser. Isis, 87(1):186, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235788.

Husberg:1996:BRB

[2281] Volker Husberg. Book review: Ein Recht auf saubere Luft? Umweltkonflikte am Beginn des Industriezeitalters by Michael Stolberg. Isis, 87(1): 186–187, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235789.

Grattan-Guinness:1996:BRB

[2282] I. Grattan-Guinness. Book review: The Emergence of the American Mathematical Research Community, 1876–1900: J. J. Sylvester, Felix Klein, and E. H. Moore by Karen Hunger Parshall; David E. Rowe. Isis, 87(1):187–188, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235790.

Flannery:1996:BRBa

[2283] Michael A. Flannery. Book review: The Spirit of Voluntarism: A Legacy of Commitment and Contribution: The United States Pharmacopeia, 1820–1995 by Lee Anderson; Gregory J. Higby. Isis, 87(1):188–189, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235791.

Allison:1996:BRBa

[2284] Steven W. Allison. Book review: Science in the Pleasure Ground: A History of the Arnold Arboretum by Ida Hay. Isis, 87(1):189–190, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235792.

Edgerton:1996:BRB

[2285] D. E. H. Edgerton. Book review: The Precision Makers: A History of the Instruments Industry in Britain and France, 1870–1939 by Mari E. W. Williams. Isis, 87(1):190, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235793.

Weidman:1996:BRBa

[2286] Nadine Weidman. Book review: Seven Pioneers of Psychology: Behaviour and Mind by Ray Fuller. Isis, 87(1):190–191, March 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235794.

Carson:1996:BRB

[2287] John Carson. Book review: Measuring the Mind: Education and Psychology in England, c. 1860–c. 1990 by Adrian Wooldridge. Isis, 87(1): 191–192, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235795.

Rabkin:1996:BRB

[2288] Yakov M. Rabkin. Book review: Science and Russian Culture in an Age of Revolutions: V. I. Vernadsky and His Scientific School, 1863–1945 by Kendall E. Bailes; Physics and Politics in Revolutionary Russia by Paul R. Josephson. Isis, 87(1):192–194, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235796.

Bloor:1996:BRB

[2289] David Bloor. Book review: The Threat of Logical Mathematism: A Study on the Critique of Mathematical Logic in Germany at the Turn of the Twentieth Century by Jarmo Pulkkinen. Isis, 87(1):194–195, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235797.

Keller:1996:BRB

[2290] Pierre Keller. Book review: Einstein philosophe: La physique comme pratique philosophique by Michel Paty. Isis, 87(1):195–196, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235798.

Kragh:1996:BRBa

[2291] Helge Kragh. Book review: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg u.a. Volume 3: 1940–1949 by Wolfgang Pauli; Karl von Mayenn. Isis, 87(1):196, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235799.

Verbong:1996:BRB

[2292] G. P. J. Verbong. Book review: Wetenschap tussen universiteit en industrie: De experimentele natuurkunde in Utrecht onder W. H. Julius en L. S. Ornstein 1896–1940 by H. G. Heijmans. Isis, 87(1):196–197, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235800.

Assmus:1996:BRB

[2293] Alexi Assmus. Book review: The World of Physical Chemistry by Keith J. Laidler. Isis, 87(1):197–198, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235801.

Barnes:1996:BRBa

[2294] Barry Barnes. Book review: Explaining Scientific Consensus: The Case of Mendelian Genetics by Kyung-Man Kim. Isis, 87(1):198–199, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235802.

Segal:1996:BRB

[2295] Robert A. Segal. Book review: The Jung Cult: Origins of a Charismatic Movement by Richard Noll. Isis, 87(1):199, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235803.

Rosner:1996:BRBa

[2296] David Rosner. Book review: Making a Place for Ourselves: The Black Hospital Movement, 1920–1945 by Vanessa Northington Gamble. Isis, 87(1):199–200, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235804.

Young:1996:BRB

[2297] James Harvey Young. Book review: Chicago’s War on Syphilis, 1937– 1940: The Times, the Trib, and the Clap Doctor by Suzanne Poirier. Isis, 87(1):200–201, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235805.

Severino:1996:BRB

[2298] Sally K. Severino. Book review: The Psychiatric Persuasion: Knowledge, Gender, and Power in Modern America by Elizabeth Lunbeck. Isis, 87 (1):201–202, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235806.

Lunbeck:1996:BRB

[2299] Elizabeth Lunbeck. Book review: The Romance of American Psychology: Political Culture in the Age of Experts by Ellen Herman. Isis, 87(1):202– 203, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235807.

Macrakis:1996:BRB

[2300] Kristie Macrakis. Book review: Die Welteislehre: Ihre Geschichte und ihre Rolle im “Dritten Reich” by Brigitte Nagel. Isis, 87(1):203–204, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235808.

Brown:1996:BRB

[2301] Laurie M. Brown. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger, and Tomonaga by Silvan S. Schweber. Isis, 87(1): 204–205, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235809.

Constant:1996:BRBa

[2302] Ed Constant. Book review: Designing Engineers by Louis L. Bucciarelli. Isis, 87(1):205, March 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235810.

Chiara:1996:BRB

[2303] Maria Luisa Dalla Chiara. Book review: Dear Carnap, Dear Van: The Quine-Carnap Correspondence and Related Work by W. V. Quine; Rudolf Carnap; Richard Creath. Isis, 87(1):205–206, March 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235811.

Anonymous:1996:Ca

[2304] Anonymous. Collections. Isis, 87(1):206–210, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235812.

Anonymous:1996:BMb

[2305] Anonymous. Back matter. Isis, 87(1):211–216, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235813.

Anonymous:1996:FMb

[2306] Anonymous. Front matter. Isis, 87(1):??, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235731.

Terrall:1996:SAB

[2307] Mary Terrall. Salon, academy, and boudoir: Generation and desire in Maupertuis’s science of life. Isis, 87(2):217–229, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236061.

Methuen:1996:MTC

[2308] Charlotte Methuen. Maestlin’s teaching of Copernicus: The evidence of his University textbook and disputations. Isis, 87(2):230–247, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236062.

Beyler:1996:TOS

[2309] Richard H. Beyler. Targeting the organism: The scientific and cultural context of Pascual Jordan’s quantum biology, 1932–1947. Isis, 87(2): 248–273, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236063.

Strick:1996:SAT

[2310] James Strick. Swimming against the tide: Adrianus Pijper and the debate over bacterial flagella, 1946–1956. Isis, 87(2):274–305, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236064.

Cook:1996:AMH

[2311] Harold J. Cook, Deborah Fitzgerald, Keith R. Benson, and Ronald L. Numbers. Annual meeting of the History of Science Society, Minneapolis, 26–29 October 1995. Isis, 87(2):306–312, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236065.

Bray:1996:EJN

[2312] Francesca Bray. Eloge: Joseph Needham, 9 December 1900–24 March 1995. Isis, 87(2):312–317, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236066.

Burnham:1996:EED

[2313] John C. Burnham. Eloge: Emanuel David Rudolph, 9 September 1927– 22 June 1992. Isis, 87(2):317–318, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236067.

Broughton:1996:BRB

[2314] Trev Broughton. Book review: The Facts of Life: The Creation of Sexual Knowledge in Britain, 1650–1950. Isis, 87(2):319–322, June 1996.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236068.

Wise:1996:BRB

[2315] M. Norton Wise. Book review: Higher Superstition: The Academic Left and Its Quarrels with Science. Isis, 87(2):323–327, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236069.

Geison:1996:BRB

[2316] Gerald L. Geison. Book review: Lords of the Fly: Drosophila Genetics and the Experimental Life. Isis, 87(2):328–331, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236070.

Kohler:1996:BRB

[2317] Robert E. Kohler. Book review: The Private Science of Louis Pasteur by Gerald L. Geison. Isis, 87(2):331–334, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236071.

Hacking:1996:BRB

[2318] Ian Hacking. Book review: Killing Time: The Autobiography of Paul Feyerabend by Paul K. Feyerabend. Isis, 87(2):334–335, June 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236072.

Rusnock:1996:BRB

[2319] Andrea Rusnock. Book review: Interactions: Some Contacts between the Natural Sciences and the Social Sciences by I. Bernard Cohen. Isis, 87(2):335–336, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236073.

Kragh:1996:BRBb

[2320] Helge Kragh. Book review: Libres propos sur l’histoire des sciences by François Russo. Isis, 87(2):336, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236074.

Constant:1996:BRBb

[2321] Edward W. Constant, II. Book review: Exploring the Black Box: Technology, Economics, and History by Nathan Rosenberg. Isis, 87(2):337, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236075.

Flannery:1996:BRBb

[2322] Michael A. Flannery. Book review: The History of Pharmacy: A Selected Annotated Bibliography by Gregory J. Higby; Elaine C. Stroud. Isis, 87 (2):337–338, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236076.

Rainger:1996:BRB

[2323] Ronald Rainger. Book review: Silence of the Spheres: The Deaf Experience in the History of Science by Harry G. Lang. Isis, 87(2):338–339, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236077.

Smith:1996:BRBb

[2324] Norman A. F. Smith. Book review: Hommes et techniques de l’Antiquite a la Renaissance by Marie-Claire Amouretti; Georges Comet. Isis, 87(2): 339–340, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236078.

Sutton:1996:BRB

[2325] Geoffrey V. Sutton. Book review: Pythagoras’ Trousers: God, Physics, and the Gender Wars by Margaret Wertheim. Isis, 87(2):340, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236079.

Tytler:1996:BRB

[2326] Graeme Tytler. Book review: La communication non verbale avant la lettre by Anne-Marie Drouin-Hans. Isis, 87(2):340–341, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236080.

Lloyd:1996:BRBa

[2327] G. E. R. Lloyd. Book review: Studies in Greek Philosophy by Gregory Vlastos; Daniel W. Graham. Isis, 87(2):341–343, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236081.

Modrak:1996:BRB

[2328] Deborah K. W. Modrak. Book review: Animal Minds and Human Morals: The Origins of the Western Debate by Richard Sorabji. Isis, 87(2):343, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236082.

Allen:1996:BRBa

[2329] James P. Allen. Book review: Ancient Egyptian Science: A Source Book. Volume 2: Calenders, Clocks, and Astronomy by Marshall Clagett. Isis, 87(2):343–344, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236083.

Marmion:1996:BRB

[2330] Vincent J. Marmion. Book review: L’ophtalmologie dans l’Egypte greco romaine d’apres les papyrus litteraires grecs by Marie-Helene Marganne. Isis, 87(2):345, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236084.

McVaugh:1996:BRBb

[2331] Michael R. McVaugh. Book review: Avicenna Latinus: Codices by Avicenna; Marie-Therese d’Alverny; Simone van Riet; Pierre Jodogne. Isis, 87(2):345, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236085.

Grant:1996:BRB

[2332] Edward Grant. Book review: Atomtheorien des lateinischen Mittelalters by Bernhard Pabst. Isis, 87(2):345–346, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236086.

Shank:1996:BRB

[2333] Michael H. Shank. Book review: Lire dans le ciel: La bibliothèque de Simon de Phares, astrologue du XVe siècle by Jean-Patrice Boudet. Isis, 87(2):346–347, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236087.

Conrad:1996:BRB

[2334] Lawrence I. Conrad. Book review: El legado cientifico andalusi by Juan Vernet; Julio Samso. Isis, 87(2):347–348, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236088.

Horstmanshoff:1996:BRB

[2335] Manfred Horstmanshoff. Book review: Geneeskunde en humanisme: Een intellectuele Biografie van Theodericus Ulserius (c. 1460–1508) by Catrien Santing. Isis, 87(2):348, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236089.

Gingerich:1996:BRB

[2336] Owen Gingerich. Book review: Die astronomische Uhr des Strassburger Munsters: Funktion und Bedeutung eines Kosmos-Modells des 16. Jahrhunderts by Gunther Oestmann. Isis, 87(2):348–349, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236090.

Johns:1996:BRB

[2337] Adrian Johns. Book review: John Dee: The Politics of Reading and Writing in the English Renaissance by William H. Sherman. Isis, 87(2): 349–350, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236091.

Gosselin:1996:BRB

[2338] Edward A. Gosselin. Book review: Oeuvres completes. Volume 1: Chandelier by Giordano Bruno; Giovanni Aquilecchia; Yves Hersant; Oeuvres complètes. Volume 6: Cabale du cheval pegaséen by Giordano Bruno; Giovanni Aquilecchia; Tristan Dagron. Isis, 87(2):350–351, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236092.

MacLean:1996:BRB

[2339] Ian MacLean. Book review: Perilous Chastity: Women and Illness in Pre-Enlightenment Art and Medicine by Laurinda S. Dixon. Isis, 87(2): 352, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236093.

Malet:1996:BRB

[2340] Antoni Malet. Book review: Itinerari del volontarismo: Teologia e politica al tempo di Luis de Leon by Domenico Ferraro. Isis, 87(2):352–353, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236094.

Webster:1996:BRBb

[2341] Charles Webster. Book review: Alchemy and Chemistry in the Sixteenth and Seventeenth Centuries by Piyo Rattansi; Antonio Clericuzio. Isis, 87(2):353–354, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236095.

Godwin:1996:BRB

[2342] Joscelyn Godwin. Book review: Adam Haslmayr: Der erste Verkunder der Manifeste der Rosenkreuzer by Carlos Gilly. Isis, 87(2):354, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236096.

Anstey:1996:BRB

[2343] Peter Anstey. Book review: Samuel Hartlib and Universal Reformation: Studies in Intellectual Communication by Mark Greengrass; Michael Leslie; Timothy Raylor. Isis, 87(2):354–355, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236097.

Mandelbrote:1996:BRB

[2344] Scott Mandelbrote. Book review: Astrology and the Seventeenth-Century Mind: William Lilly and the Language of the Stars by Ann Geneva. Isis, 87(2):355–356, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236098.

Martin:1996:BRB

[2345] Julian Martin. Book review: Sir Matthew Hale, 1609–1676: Law, Religion, and Natural Philosophy by Alan Cromartie. Isis, 87(2):356–357, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236099.

Bates:1996:BRBa

[2346] Don Bates. Book review: English Manuscripts of Francis Glisson (1): From Anatomia hepatis (The Anatomy of the Liver), 1654 by Andrew Cunningham. Isis, 87(2):357–358, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236100.

Mahoney:1996:BRBb

[2347] Michael S. Mahoney. Book review: Penetralia motus: La fondazione relativistica della meccanica in Christiaan Huygens, con l’edizione del Codex Hugeniorum 7A by Gianfranco Mormino. Isis, 87(2):358, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236101.

Ariew:1996:BRB

[2348] Roger Ariew. Book review: L’harmonie et le chaos: Le rationalisme leibnizien et la ’nouvelle science.’ by Laurence Bouquiaux; Leibniz and the Rational Order of Nature by Donald Rutherford. Isis, 87(2):358– 360, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236102.

Schiebinger:1996:BRB

[2349] Londa Schiebinger. Book review: Women on the Margins: Three Seventeenth-Century Lives by Natalie Zemon Davis. Isis, 87(2):360–361, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236103.

Schaffer:1996:BRB

[2350] Simon Schaffer. Book review: La malle de Newton by Loup Verlet. Isis, 87(2):361–362, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236104.

Sturdy:1996:BRB

[2351] David J. Sturdy. Book review: Impolite Learning: Conduct and Community in the Republic of Letters, 1680–1750 by Anne Goldgar. Isis, 87(2): 362–363, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236105.

Verbeek:1996:BRB

[2352] Theo Verbeek. Book review: Man a Machine and Man a Plant by Julien Offray de La Mettrie; Richard A. Watson; Maya Ribalka. Isis, 87(2): 363–364, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236106.

Broman:1996:BRB

[2353] Thomas Broman. Book review: Johann Christian Polycarp Erxleben: Versuch einer Biographie und Bibliographie by Gerta Beaucamp. Isis, 87 (2):364, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236107.

Fraser:1996:BRB

[2354] Craig Fraser. Book review: La mecanique de Lagrange: Principes et methodes by Wilton Barroso Filho. Isis, 87(2):364–365, June 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236108.

Findlen:1996:BRBb

[2355] Paula Findlen. Book review: Un viaggio in Europa nel 1786: Diario di Barnaba Oriani astronomo milanese by A. Mandrino; G. Tagliaferri; P.

Tucci. Isis, 87(2):365–366, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236109.

Caneva:1996:BRB

[2356] Kenneth L. Caneva. Book review: Erganzungsband zu Werke Band 5 bis 9: Wissenschaftshistorischer Bericht zu Schellings naturphilosophischen Schriften 1797–1800 by Friedrich Wilhelm Joseph Schelling. Isis, 87(2): 366–367, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236110.

Limoges:1996:BRB

[2357] Camille Limoges. Book review: Science under Control: The French Academy of Sciences, 1795–1914 by Maurice P. Crosland. Isis, 87(2): 367–368, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236111.

Nieto-Galan:1996:BRB

[2358] Agusti Nieto-Galan. Book review: The Historical and Philosophical Background of Hauy’s Theory of Crystal Structure by R. Hooykaas. Isis, 87(2):368–369, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236112.

Schlich:1996:BRB

[2359] Thomas Schlich. Book review: Die Entwicklung der Pharmazie zur Hochschuldisziplin (1750–1875): Ein Beitrag zur Universitäts- und Sozialgeschichte by Berthold Beyerlein. Isis, 87(2):369–370, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236113.

Mohr:1996:BRB

[2360] James C. Mohr. Book review: Witnessing Insanity: Madness and Mad-Doctors in the English Court by Joel Peter Eigen. Isis, 87(2):370–371, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236114.

Sokal:1996:BRB

[2361] Michael M. Sokal. Book review: Face Value: Physiognomical Thought and the Legible Body in Marivaux, Lavater, Balzac, Gautier, and Zola by Christopher Rivers. Isis, 87(2):371–372, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236115.

Porter:1996:BRB

[2362] Dorothy Porter. Book review: Romance, Poetry, and Surgical Sleep: Literature Influences Medicine by E. M. Papper. Isis, 87(2):372, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236116.

Albury:1996:BRB

[2363] W. R. Albury. Book review: Claude Bernard: Rationalite d’une methode by Pierre Gendron. Isis, 87(2):372–373, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236117.

Rupke:1996:BRBa

[2364] Nicolaas Rupke. Book review: Biology Takes Form: Animal Morphology and the German Universities, 1800–1900 by Lynn K. Nyhart. Isis, 87(2): 373–374, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236118.

Hiebert:1996:BRB

[2365] Erwin N. Hiebert. Book review: Science and Culture: Popular and Philosophical Essays by Hermann von Helmholtz; David Cahan. Isis, 87(2): 374–375, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236119.

Rosenberg:1996:BRBa

[2366] Charles E. Rosenberg. Book review: Simon Baruch: Rebel in the Ranks of Medicine, 1840–1921 by Patricia Spain Ward. Isis, 87(2):375–376, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236120.

Kohlstedt:1996:BRB

[2367] Sally Gregory Kohlstedt. Book review: American Women Afield: Writings by Pioneering Women Naturalists by Marcia Myers Bonta. Isis, 87(2):376–377, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236121.

Alter:1996:BRB

[2368] Stephen G. Alter. Book review: Science and Religion in the Era of William James. Volume 1: The Eclipse of Certainty, 1820–1880 by Paul Jerome Croce. Isis, 87(2):377–378, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236122.

Kline:1996:BRB

[2369] Ronald Kline. Book review: The Papers of Thomas A. Edison. Volume 3: Menlo Park: The Early Years, April 1876–December 1877 by Thomas A. Edison; Robert A. Rosenberg; Paul B. Israel; Keith A. Nier; Martha J. King. Isis, 87(2):378–379, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236123.

Blaszczyk:1996:BRB

[2370] Regina Lee Blaszczyk. Book review: American Technological Sublime by David E. Nye. Isis, 87(2):379, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236124.

Hunt:1996:BRB

[2371] Bruce J. Hunt. Book review: From Obscurity to Enigma: The Work of Oliver Heaviside, 1872–1889 by Ido Yavetz. Isis, 87(2):379–380, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236125.

Hetherington:1996:BRBb

[2372] Norriss S. Hetherington. Book review: The Immortal Fire within: The Life and Work of Edward Emerson Barnard by William Sheehan. Isis, 87(2):380–381, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236126.

Rhees:1996:BRB

[2373] David J. Rhees. Book review: Crusading for Chemistry: The Professional Career of Charles Holmes Herty by Germaine M. Reed. Isis, 87 (2):381–382, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236127.

Weidman:1996:BRBb

[2374] Nadine Weidman. Book review: Brain, Mind, and Medicine: Charles Richet and the Origins of Physiological Psychology by Stewart Wolf. Isis, 87(2):382–383, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236128.

Pickren:1996:BRB

[2375] Wade Pickren. Book review: Piaget before Piaget by Fernando Vidal. Isis, 87(2):383, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236129.

Chang:1996:BRB

[2376] Hasok Chang. Book review: Einstein and Our World by David Cassidy. Isis, 87(2):383–384, June 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236130.

Heilbron:1996:BRBb

[2377] J. L. Heilbron. Book review: A Guide to the Archival Collections in the Niels Bohr Library at the American Institute of Physics. Isis, 87(2):384– 385, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236131.

Wheaton:1996:BRB

[2378] Bruce R. Wheaton. Book review: Die Atomphysiker: Eine Geschichte der theoretischen Physik am Beispiel der Sommerfeldschule by Michael Eckert. Isis, 87(2):385–386, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236132.

Pinch:1996:BRB

[2379] Trevor Pinch. Book review: Quantum Mechanics: Historical Contingency and the Copenhagen Hegemony by James T. Cushing. Isis, 87(2): 386, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236133.

Dyson:1996:BRB

[2380] Freeman J. Dyson. Book review: Ramanujan: Letters and Commentary by Bruce C. Berndt; Robert A. Rankin. Isis, 87(2):387, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236134.

Cushing:1996:BRBa

[2381] James T. Cushing. Book review: The Beat of a Different Drum: The Life and Science of Richard Feynman by Jagdish Mehra. Isis, 87(2): 387–388, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236135.

Cocks:1996:BRB

[2382] Geoffrey Cocks. Book review: Furchtbare Arzte: Medizinische Verbrechen im Dritten Reich by Till Bastian. Isis, 87(2):388–389, June 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236136.

Biddle:1996:BRB

[2383] Tami Davis Biddle. Book review: America’s Pursuit of Precision Bombing, 1910–1945 by Stephen L. McFarland. Isis, 87(2):389–390, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236137.

Beatty:1996:BRB

[2384] John Beatty. Book review: Suffering Made Real: American Science and the Survivors at Hiroshima by M. Susan Lindee. Isis, 87(2):390–391, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236138.

Doel:1996:BRB

[2385] Ronald E. Doel. Book review: The Origin of the Solar System: Soviet Research, 1925–1991 by Aleksey E. Levin; Stephen G. Brush. Isis, 87(2): 391–392, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236139.

Bowker:1996:BRB

[2386] Geoffrey C. Bowker. Book review: Opting for Oil: The Political Economy of Technological Change in the West German Chemical Industry, 1945– 1961 by Raymond G. Stokes. Isis, 87(2):392–393, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236140.

Smith:1996:BRBc

[2387] Robert W. Smith. Book review: Spaceflight Revolution: NASA Langley Research Center from Sputnik to Apollo by James R. Hansen. Isis, 87(2): 393–394, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236141.

Tatarewicz:1996:BRB

[2388] Joseph N. Tatarewicz. Book review: Powering Apollo: James E. Webb of NASA by W. Henry Lambright. Isis, 87(2):394–395, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236142.

Goldberg:1996:BRB

[2389] Steven Goldberg. Book review: Science at the Bar: Law, Science, and Technology in America by Sheila Jasanoff. Isis, 87(2):395–396, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236143.

Nickles:1996:BRB

[2390] Thomas Nickles. Book review: Criticism and the History of Science: Kuhn’s, Lakatos’s, and Feyerabend’s Criticisms of Critical Rationalism by Gunnar Andersson. Isis, 87(2):396–397, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236144.

Sargent:1996:BRBb

[2391] Rose-Mary Sargent. Book review: The Discourses of Science by Marcello Pera; Clarissa Botsford. Isis, 87(2):397, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236145.

Cantor:1996:BRB

[2392] David Cantor. Book review: A Bibliography of Medical and Biomedical Biography by Leslie T. Morton; Robert J. Moore. Isis, 87(2):397–398, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236146.

Anonymous:1996:Cb

[2393] Anonymous. Collections. Isis, 87(2):398–402, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236147.

Anonymous:1996:BMc

[2394] Anonymous. Back matter. Isis, 87(2):403–408, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236148.

Anonymous:1996:FMc

[2395] Anonymous. Front matter. Isis, 87(2):??, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236060.

Rozwadowski:1996:SWF

[2396] Helen M. Rozwadowski. Small world: Forging a scientific maritime culture for oceanography. Isis, 87(3):409–429, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235978.

Ramati:1996:HDL

[2397] Ayval Ramati. Harmony at a distance: Leibniz’s scientific academies. Isis, 87(3):430–452, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235979.

French:1996:FFA

[2398] Roger French. Foretelling the future: Arabic astrology and English medicine in the Late Twelfth Century. Isis, 87(3):453–480, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235980.

Bensaude-Vincent:1996:BHM

[2399] Bernadette Bensaude-Vincent. Between history and memory: Centennial and bicentennial images of Lavoisier. Isis, 87(3):481–499, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235981.

Hahn:1996:EDD

[2400] Roger Hahn. Eloge: Denis I. Duveen. Isis, 87(3):500–501, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235982.

Layton:1996:EMK

[2401] Edwin T. Layton. Eloge: Melvin Kranzberg, 22 November 1917–6 December 1995. Isis, 87(3):501–503, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235983.

Dear:1996:LE

[2402] Peter Dear and Mordechai Feingold. Letters to the Editor. Isis, 87(3): 504–506, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235984.

deChadarevian:1996:BRB

[2403] Soraya de Chadarevian. Book review: A Skeptical Biochemist. Isis, 87 (3):507–510, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235985.

Siegel:1996:BRB

[2404] Daniel Siegel. Book review: Maxwell on Heat and Statistical Mechanics: On “Avoiding All Personal Enquiries” of Molecules. Isis, 87(3):511–516, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235986.

Daston:1996:BRB

[2405] Lorraine Daston. Book review: The Values of Precision. Isis, 87(3): 517–519, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235987.

Ezrahi:1996:BRB

[2406] Yaron Ezrahi. Book review: Trust in Numbers: The Pursuit of Objectivity in Science and Public Life by Theodore M. Porter. Isis, 87(3): 519–521, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235988.

Slack:1996:BRB

[2407] Nancy G. Slack. Book review: Naturalist by Edward O. Wilson. Isis, 87 (3):521–524, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235989.

Boon:1996:BRB

[2408] James A. Boon. Book review: Fieldwork and Footnotes: Studies in the History of European Anthropology by Han F. Vermeulen; Arturo Alvarez Roldan. Isis, 87(3):524–526, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235990.

Rousseau:1996:BRB

[2409] G. S. Rousseau. Book review: From Faust to Strangelove: Representations of the Scientist in Western Literature by Roslynn D. Haynes. Isis, 87(3):526–527, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235991.

Inkster:1996:BRBa

[2410] Ian Inkster. Book review: Science and Technology in a Multicultural World: The Cultural Politics of Facts and Artifacts by David J. Hess. Isis, 87(3):527–528, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235992.

Hetherington:1996:BRBc

[2411] Norriss S. Hetherington. Book review: Milchstrasse, Nebel, Galaxien: Strukturen im Kosmos von Herschel bis Hubble by Gudrun Wolfschmidt.

Isis, 87(3):528, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235993.

Hannaway:1996:BRB

[2412] Caroline Hannaway. Book review: The Western Medical Tradition: 800 B.C. to A.D. 1800 by Lawrence I. Conrad; Michael Neve; Vivian Nutton; Roy Porter; Andrew Wear. Isis, 87(3):528–529, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235994.

Turner:1996:BRB

[2413] Steven Turner. Book review: Wachstum und Differenzierung der deutschen Universitäten 1830–1945 by Hartmut Titze; Hans-Georg Herrlitz; Volker Muller-Benedict; Axel Nath. Isis, 87(3):529–530, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235995.

Rosner:1996:BRBb

[2414] Lisa Rosner. Book review: Becoming a Physician: Medical Education in Great Britain, France, Germany, and the United States, 1750–1945 by Thomas Neville Bonner. Isis, 87(3):530–531, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235996.

Potter:1996:BRB

[2415] Paul Potter. Book review: Arztelexikon: Von der Antike bis zum 20. Jahrhundert by Wolfgang U. Eckart; Christoph Gradmann. Isis, 87(3): 531, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235997.

DeWitt:1996:BRB

[2416] Calvin B. DeWitt. Book review: The Earth, Humanity, and God: The Templeton Lectures, Cambridge 1993 by Colin A. Russell. Isis, 87(3): 531–532, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235998.

Molland:1996:BRB

[2417] George Molland. Book review: Concepts of Space in Greek Thought by Keimpe Algra. Isis, 87(3):532–533, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235999.

Furley:1996:BRB

[2418] David Furley. Book review: Aristotle’s Theory of Material Substance: Heat and Pneuma, Form and Soul by Gad Freudenthal. Isis, 87(3): 533–534, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236000.

MacIntosh:1996:BRB

[2419] J. J. MacIntosh. Book review: Hypatia of Alexandria by Maria Dzielska; F. Lyra. Isis, 87(3):534–535, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236001.

Lloyd:1996:BRBb

[2420] G. E. R. Lloyd. Book review: Sculptors and Physicians in Fifth-Century Greece: A Preliminary Study by Guy P. R. Metraux. Isis, 87(3):535–536, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236002.

vanderEijk:1996:BRB

[2421] Philip J. van der Eijk. Book review: Trattato sulla febbre by Ps. Alessandro d’Afrodisia; Piero Tassinari. Isis, 87(3):536, September 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236003.

Peterson:1996:BRB

[2422] Charles Peterson. Book review: Science and Civilisation in China. Volume 5: Chemistry and Chemical Technology. Part 6: Military Technology: Missiles and Sieges by Joseph Needham; Robin D. S. Yates. Isis, 87 (3):536–538, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236004.

Jacquart:1996:BRBa

[2423] Danielle Jacquart. Book review: Jews, Medicine, and Medieval Society by Joseph Shatzmiller. Isis, 87(3):538–539, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236005.

Nutton:1996:BRBb

[2424] Vivian Nutton. Book review: Die deutsche Sachliteratur des Mittelalters by William Crossgrove. Isis, 87(3):539–540, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236006.

Mancha:1996:BRB

[2425] Jose Luis Mancha. Book review: An Almanac for Trebizond for the Year 1336 by Raymond Mercier. Isis, 87(3):540–541, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236007.

Mestres:1996:BRB

[2426] Angel Mestres. Book review: The Abbreviation of the Introduction to Astrology: Together with the Medieval Latin Translation of Adelard of Bath by Abu Masar; Charles Burnett; Keiji Yamamoto; Michio Yano. Isis, 87(3):541, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236008.

Blum:1996:BRB

[2427] Paul Richard Blum. Book review: The Transformation of Natural Philosophy: The Case of Philip Melanchthon by Sachiko Kusukawa. Isis, 87 (3):541–542, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236009.

Ziggelaar:1996:BRB

[2428] August Ziggelaar. Book review: Tycho Brahes “Urania Titani”: Et digt om Sophie Brahe by Peter Zeeberg. Isis, 87(3):542–543, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236010.

Coudert:1996:BRB

[2429] Allison P. Coudert. Book review: The Word of God and the Languages of Man: Interpreting Nature in Early Modern Science and Medicine. Volume 1: Ficino to Descartes by James Bono. Isis, 87(3):543–544, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236011.

Shea:1996:BRBb

[2430] William Shea. Book review: Descartes: An Intellectual Biography by Stephen Gaukroger. Isis, 87(3):544–545, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236012.

Licoppe:1996:BRB

[2431] Christian Licoppe. Book review: Blaise Pascal: Mathematician, Physicist, and Thinker about God by Donald Adamson. Isis, 87(3):545, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236013.

Hine:1996:BRB

[2432] William L. Hine. Book review: Tommaso Campanella en France: Au XVIIe siècle by Michel-Pierre Lerner. Isis, 87(3):545–546, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236014.

Applebaum:1996:BRB

[2433] Wilbur Applebaum. Book review: Science for a Polite Society: Gender, Culture, and the Demonstration of Enlightenment by Geoffrey V. Sutton. Isis, 87(3):546–547, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236015.

Brotons:1996:BRB

[2434] Victor Navarro Brotons. Book review: Revolucion cientifica, Renacimiento e historia de la ciencia by Antonio Beltran. Isis, 87(3):547, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236016.

Ranea:1996:BRBb

[2435] Alberto Guillermo Ranea. Book review: Clasicos medicos valencianos del siglo XVII by Jose M. Lopez Piñero. Isis, 87(3):547–548, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236017.

Chemla:1996:BRB

[2436] Karine Chemla. Book review: Les mathematiques japonaises a l’epoque d’Edo: Une etude des travaux de Seki Takakazu (?-1708) et de Takebe Katahiro (1664–1739) by Annick Horiuchi. Isis, 87(3):548–549, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236018.

Jesseph:1996:BRB

[2437] Douglas M. Jesseph. Book review: Elements de la geometrie de l’infini by Bernard le Bovier de Fontenelle. Isis, 87(3):549–550, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236019.

Hankins:1996:BRBa

[2438] Thomas L. Hankins. Book review: Die Werke von Daniel Bernoulli. Volume 7: Magnetism and Technology by Daniel Bernoulli; Patricia Radeletde Grave; David Speiser; Andre Englebert. Isis, 87(3):550–551, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236020.

Roberts:1996:BRB

[2439] Lissa Roberts. Book review: The Age of Minerva by Paul Ilie. Isis, 87(3):551, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236021.

Koerner:1996:BRB

[2440] Lisbet Koerner. Book review: Linnaeus: The Man and His Work by Tore Frangsmyr. Isis, 87(3):551–552, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236022.

Frodin:1996:BRB

[2441] David G. Frodin. Book review: The Development of Biological Systematics: Antoine-Laurent de Jussieu, Nature, and the Natural System by Peter F. Stevens. Isis, 87(3):552–554, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236023.

Farber:1996:BRB

[2442] Paul Lawrence Farber. Book review: Fishes, Crayfishes, and Crabs: Louis Renard’s Natural History of the Rarest Curiosities of the Seas of the Indies by Theodore W. Pietsch. Isis, 87(3):554, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236024.

DiGregorio:1996:BRBa

[2443] Mario A. Di Gregorio. Book review: Discorso storico della mia vita naturale e civile dal 1745 al 1825 by Pietro Ferroni; Danilo Barsanti. Isis, 87(3):554–555, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236025.

Pinero:1996:BRBb

[2444] Jose M. Lopez Piñero. Book review: Quimica, botanica y farmacia en la Nueva España a finales del siglo XVIII by Patricia Aceves Pastrana.

Isis, 87(3):555–556, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236026.

Feldhay:1996:BRB

[2445] Rivka Feldhay. Book review: Copernico, Galilei e la Chiesa: Fine della controversia (1820) gli atti del Sant’Uffizio by Walter Brandmuller; Egon Johannes Greipl. Isis, 87(3):556–557, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236027.

Kim:1996:BRBa

[2446] Mi Gyung Kim. Book review: In the Shadow of Lavoisier: The Annales de Chimie and the Establishment of a New Science by Maurice Crosland. Isis, 87(3):557, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236028.

Gregory:1996:BRB

[2447] Frederick Gregory. Book review: Was ich erlebte: Aus der Erinnerung niedergeschrieben by Henrich Steffens. Isis, 87(3):557–558, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236029.

Jackson:1996:BRB

[2448] Myles W. Jackson. Book review: Goethe’s Conception of Knowledge and Science by R. H. Stephenson. Isis, 87(3):558–559, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236030.

Camerini:1996:BRB

[2449] Jane R. Camerini. Book review: The Correspondence of Charles Darwin. Volume 9: 1861 by Charles Darwin; Frederick Burkhardt; Duncan M. Porter; Joy Harvey; Marsha Richmond. Isis, 87(3):559–560, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236031.

Russett:1996:BRB

[2450] Cynthia Russett. Book review: Biology and Social Thought: 1850– 1914 by Peter J. Bowler. Isis, 87(3):560, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236032.

Rupke:1996:BRBb

[2451] Nicolaas A. Rupke. Book review: E. Ray Lankester and the Making of Modern British Biology by Joseph Lester; Peter J. Bowler. Isis, 87(3): 561, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236033.

Stamhuis:1996:BRB

[2452] Ida Stamhuis. Book review: Between Seasons and Science: Commissioned by the Royal Botanical Society of the Netherlands to Commemorate the 150-Year Jubilee by Patricia Faasse. Isis, 87(3):561–562, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236034.

Allison:1996:BRBb

[2453] Steven W. Allison. Book review: Society and Museum: A History of the Buffalo Society of Natural Sciences, 1861–1993, and the Buffalo Museum of Science, 1928–1993 by George F. Goodyear; Virginia L. Cummings; Ethel Lee Helffenstein; Joan G. Manias. Isis, 87(3):562–563, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236035.

Fitzgerald:1996:BRB

[2454] Deborah Fitzgerald. Book review: Brethren of the Net: American Entomology, 1840–1880 by W. Conner Sorensen. Isis, 87(3):563–564, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236036.

Ewan:1996:BRB

[2455] Joseph Ewan. Book review: Florida’s Pioneer Naturalist: The Life of Charles Torrey Simpson by Elizabeth Ogren Rothra. Isis, 87(3):564–565, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236037.

Richards:1996:BRB

[2456] Joan L. Richards. Book review: The Pamphlets of Lewis Carroll. Volume 2: The Mathematical Pamphlets of Charles Lutwidge Dodgson and Related Pieces by Charles Lutwidge Dodgson; Francine F. Abeles. Isis, 87(3):565, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236038.

Wilson:1996:BRB

[2457] Andrew D. Wilson. Book review: Hochgeehrter Herr Professor! Innig geliebter Louis! Ludwig Boltzmann-Henriette von Aigentler Briefwechsel by Ludwig Boltzmann; Henriette von Aigentler; Dieter Flamm. Isis, 87 (3):565–566, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236039.

Rosenberg:1996:BRBb

[2458] Robert Rosenberg. Book review: Early Science and the First Century of Physics at Union College, 1795–1895 by V. Ennis Pilcher. Isis, 87 (3):566–567, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236040.

OConnor:1996:BRB

[2459] Richard O’Connor. Book review: A Nation of Steel: The Making of Modern America, 1865–1925 by Thomas J. Misa. Isis, 87(3):567–568, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236041.

Harman:1996:BRA

[2460] P. M. Harman. Book review: Albert Einstein, Martin J. Klein, A. J. Kox, Jurgen Renn, and Robert Schulmann: The Collected Papers of Albert Einstein. Volume 4: The Swiss Years: Writings, 1912–1914. Isis, 87 (3):568–569, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236042.

Badash:1996:BRB

[2461] Lawrence Badash. Book review: Hitler’s Uranium Club: The Secret Recordings at Farm Hall by Jeremy Bernstein; David Cassidy. Isis, 87 (3):569, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236043.

Eckert:1996:BRB

[2462] Michael Eckert. Book review: Die Gesellschaft fur Schwerionenforschung: Geschichte einer Grossforschungseinrichtung fur Grundlagenforschung by Siegfried Buchhaupt. Isis, 87(3):569–570, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236044.

Cushing:1996:BRE

[2463] James T. Cushing. Book review: Erwin Schrödinger and Michel Bitbol: The Interpretation of Quantum Mechanics: Dublin Seminars (1949– 1955) and Other Unpublished Essays. Isis, 87(3):570–571, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236045.

Burkhardt:1996:BRBa

[2464] Richard W. Burkhardt, Jr. Book review: Die Entdeckung des Verhaltens: Eine Geschichte der Verhaltensforschung by Franz M. Wuketits. Isis, 87 (3):571–572, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236046.

Todd:1996:BRB

[2465] Edmund N. Todd. Book review: Konzepte von Elektrizitatsversorgung und Elektrizitatswirtschaft; Die Entstehung eines neuen Fachgebietes der Technikwissenschaften zwischen 1880 und 1945 by Norbert Gilson. Isis, 87(3):572–573, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236047.

Nye:1996:BRB

[2466] David E. Nye. Book review: Cities of Light and Heat: Domesticating Gas and Electricity in Urban America by Mark H. Rose. Isis, 87(3):573–574, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236048.

Kimmelman:1996:BRB

[2467] Barbara A. Kimmelman. Book review: Women Scientists in America: Before Affirmative Action, 1940–1972 by Margaret W. Rossiter. Isis, 87 (3):574–576, September 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236049.

Bates:1996:BRBb

[2468] Barbara Bates. Book review: Disease and Class: Tuberculosis and the Shaping of Modern North American Society by Georgina D. Feldberg. Isis, 87(3):576, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236050.

Sellers:1996:BRB

[2469] Christopher C. Sellers. Book review: Cancer Wars: How Politics Shapes What We Know and Don’t Know about Cancer by Robert N. Proctor. Isis, 87(3):576–577, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236051.

Jimenez:1996:BRB

[2470] Mary Ann Jimenez. Book review: Madness in America: Cultural and Medical Perceptions of Mental Illness before 1914 by Lynn Gamwell; Nancy Tomes. Isis, 87(3):577–578, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236052.

Harrington:1996:BRB

[2471] Anne Harrington. Book review: The Organism: A Holistic Approach to Biology Derived from Pathological Data in Man by Kurt Goldstein. Isis, 87(3):578–579, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236053.

Weizmann:1996:BRB

[2472] Fredric Weizmann. Book review: The Harmony of Illusions: Inventing Post-Traumatic Stress Disorder by Allan Young. Isis, 87(3):579–580, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236054.

Pickering:1996:BRBa

[2473] Mary Pickering. Book review: Comte after Positivism by Robert C. Scharff. Isis, 87(3):580–581, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236055.

Longhurst:1996:BRB

[2474] Brian Longhurst. Book review: Karl Mannheim and the Crisis of Liberalism: The Secret of These New Times by David Kettler; Volker Meja. Isis, 87(3):581–582, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236056.

Rouse:1996:BRBa

[2475] Joseph Rouse. Book review: Cultural Collisions: Post-Modern Technoscience by Raphael Sassower. Isis, 87(3):582–583, September 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236057.

Anonymous:1996:Cc

[2476] Anonymous. Collections. Isis, 87(3):583–588, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236058.

Anonymous:1996:BMd

[2477] Anonymous. Back matter. Isis, 87(3):589–594, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236059.

Anonymous:1996:FMd

[2478] Anonymous. Front matter. Isis, 87(3):??, September 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235977.

Brush:1996:RMP

[2479] Stephen G. Brush. The reception of Mendeleev’s periodic law in America and Britain. Isis, 87(4):595–628, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235195.

Ashworth:1996:MES

[2480] William J. Ashworth. Memory, efficiency, and symbolic analysis: Charles Babbage, John Herschel, and the industrial mind. Isis, 87(4):629–653, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235196.

Sabra:1996:SAS

[2481] A. I. Sabra. Situating Arabic science: Locality versus essence. Isis, 87 (4):654–670, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235197.

Esguerra:1996:EEE

[2482] Jorge Canizares Esguerra. Eloge: Eduardo Estrella, 1941–1996. Isis, 87 (4):671–672, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235198.

Russell:1996:EAS

[2483] G. A. Russell. Eloge: Aydin Sayili, 1913–1993. Isis, 87(4):672–675, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235199.

Simberloff:1996:LE

[2484] Daniel Simberloff, Philip J. Pauly, Wesley M. Stevens, William D. Mc-Cready, Marco Beretta, Louise Y. Palmer, Steven Shapin, and Mordechai Feingold. Letters to the Editor. Isis, 87(4):676–687, December 1996.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235200.

Melhado:1996:BRB

[2485] Evan M. Melhado. Book review: Lavoisier: Memoires d’une revolution. Isis, 87(4):688–694, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235201.

Pestre:1996:BRB

[2486] Dominique Pestre. Book review: A Matter of Choices: Memoirs of a Female Physicist. Isis, 87(4):695–700, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235202.

Westfall:1996:BRB

[2487] Richard S. Westfall. Book review: Les “Principia” de Newton. Isis, 87 (4):701–706, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235203.

Smith:1996:BRBd

[2488] Norman A. F. Smith. Book review: The Norton History of Technology by Donald Cardwell. Isis, 87(4):707, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235204.

Miller:1996:BRB

[2489] Carolyn R. Miller. Book review: The Scientific Voice by Scott L. Montgomery. Isis, 87(4):707–708, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235205.

Glick:1996:BRBb

[2490] Thomas F. Glick. Book review: Historia de la ciencia al Pais Valencia by Jose Maria Lopez Piñero; Victor Navarro Brotons. Isis, 87(4):708–709, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235206.

Cueto:1996:BRB

[2491] Marcos Cueto. Book review: Bibliographia medica hispanica, 1475–1950. Volume 2: Libros y folletos, 1601–1700 by Jose M. Lopez Piñero; Bibliographia medica hispanica, 1475–1950. Volume 4: Libros y folletos, 1801–1850 by Jose M. Lopez Piñero; Bibliographia medica hispanica, 1475–1950. Volume 8: Revistas, 1736–1950 by Jose M. Lopez Piñero; Maria Luz Lopez Terrada; Bibliographia medica hispanica, 1475–1950. Volume 9: Bibliometria de las revistas, 1736–1950 by Jose M. Lopez Piñero; Maria Luz Lopez Terrada. Isis, 87(4):709–710, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235207.

Lesch:1996:BRB

[2492] John E. Lesch. Book review: Quantification and the Quest for Medical Certainty by J. Rosser Matthews. Isis, 87(4):710–711, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235208.

Bynum:1996:BRB

[2493] W. F. Bynum. Book review: Picturing Health and Illness: Images of Identity and Difference by Sander L. Gilman. Isis, 87(4):711–712, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235209.

Bechtel:1996:BRB

[2494] William Bechtel. Book review: The Cells of the Body: A History of Somatic Cell Genetics by Henry Harris. Isis, 87(4):712–713, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235210.

Hoyrup:1996:BRB

[2495] Jens Hoyrup. Book review: Préhistoire de la géometrie: Premiers éléments d’enquête, premières conclusions by Olivier Keller. Isis, 87 (4):713–714, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235211.

Taylor:1996:BRB

[2496] C. C. W. Taylor. Book review: Démocrite: Grains de poussière dans un rayon de soleil by Jean Salem. Isis, 87(4):714–715, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235212.

McKirahan:1996:BRB

[2497] Richard McKirahan. Book review: Aristotle and Mathematics: Aporetic Method in Cosmology and Metaphysics by John J. Cleary. Isis, 87(4): 715–716, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235213.

Keyser:1996:BRB

[2498] Paul T. Keyser. Book review: Aristotle’s Physics: A Guided Study by Joe Sachs. Isis, 87(4):716–717, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235214.

Ortoleva:1996:BRB

[2499] Vincenzo Ortoleva. Book review: Pelagonius and Latin Veterinary Terminology in the Roman Empire by J. N. Adams. Isis, 87(4):717, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235215.

Sylla:1996:BRB

[2500] Edith Dudley Sylla. Book review: Thomas Bradwardine: A View of Time and a Vision of Eternity in Fourteenth-Century Thought by Edith Wilks Dolnikowski. Isis, 87(4):717–719, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235216.

Hankins:1996:BRBb

[2501] James Hankins. Book review: Nuptial Arithmetic: Marsilio Ficino’s Commentary on the Fatal Number in Book VIII of Plato’s Republic by Michael J. B. Allen. Isis, 87(4):719–720, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235217.

Riddle:1996:BRB

[2502] John M. Riddle. Book review: De vegetabilibus Buch VI, Traktat 2: Lateinisch-deutsch by Albertus Magnus; Klaus Biewer. Isis, 87(4):720, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235218.

Jacquart:1996:BRBb

[2503] Danielle Jacquart. Book review: Die Canones des Pseudo-Mesue: Eine mittelalterliche Purgantien-Lehre by Sieglinde Lieberknecht. Isis, 87(4): 720–721, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235219.

Getz:1996:BRB

[2504] Faye Getz. Book review: Manuscript Sources of Medieval Medicine: A Book of Essays by Margaret R. Schleissner. Isis, 87(4):721–722, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235220.

Maclean:1996:BRBa

[2505] Ian Maclean. Book review: Medical Ethics in the Renaissance by Winfried Schleiner. Isis, 87(4):722–723, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235221.

Donahue:1996:BRB

[2506] William H. Donahue. Book review: Documenta Copernicana: Briefe: Texte und Übersetzungen by Nicolaus Copernicus; Andreas Kuhne. Isis, 87(4):723–724, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235222.

Grene:1996:BRB

[2507] Marjorie Grene. Book review: William Harvey’s Natural Philosophy by Roger K. French; Die Mechanisierung des Herzens: Harvey und Descartes–Der vitale und der mechanische Aspekt des Kreislaufs by Thomas Fuchs. Isis, 87(4):724–725, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235223.

Gabbey:1996:BRB

[2508] Alan Gabbey. Book review: La caracteristique geometrique by Gottfried Wilhelm Leibniz; Javier Echeverria; Marc Parmentier; L’estime des apparences: 21 manuscrits de Leibniz sur les probabilités, la théorie des jeux, l’espérance de vie by Gottfried Wilhelm Leibniz; Marc Parmentier; La reforme de la dynamique: De corporum concursu (1678) et autres textes inedits by Gottfried Wilhelm Leibniz; Michel Fichant. Isis, 87 (4):725–726, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235224.

Armogathe:1996:BRB

[2509] Jean-Robert Armogathe. Book review: Die Philosophie des 17. Jahrhunderts. Part 2: Frankreich und Niederlande by Jean–Pierre Schobinger. Isis, 87(4):726–727, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235225.

Emerton:1996:BRB

[2510] Norma E. Emerton. Book review: Steno: Life, Science, Philosophy with Niels Stensen’s Prooemium and Holger Jacobaeus Niels Stensen’s Anatomical Demonstration no. XVI by Troels Kardel; Paul Maquet; Emmanuel Collins. Isis, 87(4):727–728, December 1996. CODEN ISISA4.

ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235226.

Palm:1996:BRB

[2511] Lodewijk Palm. Book review: Antoni van Leeuwenhoek, 1632–1723: L’exercise du regard by Philippe Boutibonnes. Isis, 87(4):728–729, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235227.

Leven:1996:BRB

[2512] Karl-Heinz Leven. Book review: The Collected Letters of Antoni van Leeuwenhoek. Volume 13: 1700–1701 by Antoni van Leeuwenhoek; L. C. Palm. Isis, 87(4):729–730, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235228.

Shteir:1996:BRB

[2513] Ann B. Shteir. Book review: Nature’s Body: Gender in the Making of Modern Science by Londa Schiebinger. Isis, 87(4):730–731, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235229.

Dear:1996:BRB

[2514] Peter Dear. Book review: La formation de la pratique scientifique: Le discours de l’experience en France et en Angleterre (1630–1820) by Christian Licoppe. Isis, 87(4):731–732, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235230.

vanBerkel:1996:BRB

[2515] Klaas van Berkel. Book review: Speculatie, Wetenschap en Vernuft: Fysica en astronomie volgens Wytze Foppes Dongjuma (1707–1778), instrument-maker te Leeuwarden by Huib J. Zuidervaart. Isis, 87(4): 732–733, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235231.

Cohen:1996:BRB

[2516] Claudine Cohen. Book review: El Megaterio de Bru y el Presidente Jefferson: Una relacion insospechada en los albores de la paleontologia by Jose Maria Lopez Piñero; Thomas F. Glick. Isis, 87(4):733, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235232.

Pinero:1996:BRBc

[2517] Jose M. Lopez Piñero. Book review: Tras El Dorado vegetal: Jose Celestino Mutis y la Real Expedicion Botanica del Nuevo Reino de Granada (1783–1808) by Marcelo Frias Nunez. Isis, 87(4):733–734, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235233.

Paoloni:1996:BRB

[2518] Leonello Paoloni. Book review: L’atomo fantasma: Genesi storica dell’ipotesi di Avogadro by Marco Ciardi. Isis, 87(4):734–735, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235234.

Kaiser:1996:BRB

[2519] Walter Kaiser. Book review: Weber’s Electrodynamics by Andre Koch Torres Assis. Isis, 87(4):735–736, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235235.

Williams:1996:BRB

[2520] L. Pearce Williams. Book review: Science in the Making: Scientific Development as Chronicled by the Historic Papers in the Philosophical Magazine, with Commentaries and Illustrations. Volume 1: 1798–1850 by E. A. Davis. Isis, 87(4):736–737, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235236.

Butcher:1996:BRB

[2521] Norman E. Butcher. Book review: Murchison in Moray: A Geologist on Home Ground: With the Correspondence of Roderick Impey Murchison and the Rev. Dr. George Gordon of Birnie by Michael Collie; John Diemer. Isis, 87(4):737–738, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235237.

Wermiel:1996:BRB

[2522] Sara Wermiel. Book review: Technics and Architecture: The Development of Materials and Systems for Buildings by Cecil D. Elliott. Isis, 87 (4):738–739, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235238.

Jutte:1996:BRB

[2523] Robert Jutte. Book review: Robert Remak (1815–1865): Ein judischer Arzt im Spannungsfeld von Wissenschaft und Politik by Heinz-Peter Schmiedebach. Isis, 87(4):739, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235239.

Shuttleworth:1996:BRB

[2524] Sally Shuttleworth. Book review: The Sickroom in Victorian Fiction: The Art of Being Ill by Miriam Bailin; Somatic Fictions: Imagining Illness in Victorian Culture by Athena Vrettos. Isis, 87(4):740–741, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235240.

Fritzell:1996:BRB

[2525] Peter Fritzell. Book review: Seeing New Worlds: Henry David Thoreau and Nineteenth-Century Natural Science by Laura Dassow Walls. Isis, 87 (4):741–742, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235241.

DiGregorio:1996:BRBb

[2526] Mario A. Di Gregorio. Book review: L’armonia impossibile: Alle origini del concetto di valore: Metafisica, logica e scienze della natura in R. H. Lotze dal 1838 al 1843 by Beatrice Centi. Isis, 87(4):742–743, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235242.

Barnes:1996:BRBb

[2527] Barry Barnes. Book review: Durkheim’s Philosophy of Science and the Sociology of Knowledge: Creating an Intellectual Niche by Warren Schmaus. Isis, 87(4):743–744, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235243.

Knobloch:1996:BRB

[2528] Eberhard Knobloch. Book review: Le continu et autres écrits by Hermann Weyl; Jean Largeault. Isis, 87(4):744, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235244.

DeVorkin:1996:BRB

[2529] David H. DeVorkin. Book review: A History of Astronomy from 1890 to the Present by David Leverington. Isis, 87(4):744–745, December 1996.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235245.

Hetherington:1996:BRBd

[2530] Norriss S. Hetherington. Book review: Edwin Hubble: Mariner of the Nebulae by Gale E. Christianson. Isis, 87(4):745–746, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235246.

Beyler:1996:BRL

[2531] Richard H. Beyler. Book review: Lise Meitner: A Life in Physics by Ruth Lewin Sime. Isis, 87(4):746–747, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235247.

Darrigol:1996:BRP

[2532] Olivier Darrigol. Book review: P. A. M. Dirac and R. H. Dalitz: The Collected Works of P. A. M. Dirac: 1924–1948. Isis, 87(4):747, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235248.

Tisza:1996:BRB

[2533] Laszlo Tisza. Book review: Fritz London: A Scientific Biography by Kostas Gavroglu. Isis, 87(4):748, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235249.

Richmond:1996:BRB

[2534] Marsha Richmond. Book review: Styles of Scientific Thought: The German Genetics Community, 1900–1933 by Jonathan Harwood. Isis, 87 (4):748–749, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235250.

Walker:1996:BRB

[2535] Mark Walker. Book review: Die Demontage deutscher naturwissenschaftlicher Intelligenz nach dem 2. Weltkrieg: Die Physikalisch-Technische Reichsanstalt, 1945–1948 by Lilli Peltzer. Isis, 87(4):749–750, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235251.

Inkster:1996:BRBb

[2536] Ian Inkster. Book review: Science Has No National Borders: Harry C. Kelly and the Reconstruction of Science and Technology in Postwar Japan by Hideo Yoshikawa; Joanne Kauffman; Masao Yoshida. Isis, 87 (4):750, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235252.

Kim:1996:BRBb

[2537] Dong-Won Kim. Book review: Sin-Itiro Tomonaga: Life of a Japanese Physicist by Makinosuke Matsui; Cheryl Fujimoto; Takako Sano; Hiroshi Ezawa. Isis, 87(4):750–751, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235253.

Low:1996:BRB

[2538] Morris Low. Book review: The Modern Epidemic: A History of Tuberculosis in Japan by William Johnston. Isis, 87(4):751–752, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235254.

Pyenson:1996:BRB

[2539] Lewis Pyenson. Book review: Henri Laugier en son siècle by Jean-Louis Cremieux-Brilhac; Jean-François Picard. Isis, 87(4):752–753, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235255.

Olwell:1996:BRB

[2540] Russell B. Olwell. Book review: The Plutonium Story: The Journals of Professor Glenn T. Seaborg, 1939–1946 by Glenn Theodore Seaborg; Ronald L. Kathren; Jerry B. Gough; Gary T. Benefiel; Working on the Bomb: An Oral History of World War II Hanford by S. L. Sanger; Craig Wollner. Isis, 87(4):753–754, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235256.

Owens:1996:BRB

[2541] Larry Owens. Book review: Politics on the Endless Frontier: Postwar Research Policy in the United States by Daniel Lee Kleinman. Isis, 87 (4):754–755, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235257.

Herken:1996:BRB

[2542] Gregg Herken. Book review: Impacts of the Early Cold War on the Formulation of U.S. Science Policy: Selected Memoranda of William T. Golden, October 1950–April 1951 by William T. Golden; William A. Blanpied. Isis, 87(4):755, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235258.

Pickering:1996:BRBb

[2543] Andy Pickering. Book review: The Closed World: Computers and the Politics of Discourse in Cold War America by Paul N. Edwards. Isis, 87(4):756, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235259.

Weidman:1996:BRBc

[2544] Nadine Weidman. Book review: The Alpha Males: An Early History of the Regional Primate Research Centers by W. Richard Dukelow. Isis, 87 (4):756–757, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235260.

Rutkow:1996:BRB

[2545] Ira M. Rutkow. Book review: Technology in the Hospital: Transforming Patient Care in the Early Twentieth Century by Joel D. Howell. Isis, 87 (4):757–759, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235261.

Allen:1996:BRBb

[2546] Garland E. Allen. Book review: Sex, Race, and Science: Eugenics in the Deep South by Edward J. Larson. Isis, 87(4):759–760, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235262.

Jaeger:1996:BRB

[2547] Siegfried Jaeger. Book review: Changing the Rules: Psychology in The Netherlands, 1900–1985 by Trudy Dehue. Isis, 87(4):760, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235263.

Allison:1996:BRBc

[2548] Steven W. Allison. Book review: Habitat Dioramas: Illusions of Wilderness in Museums of Natural History by Karen Wonders. Isis, 87(4): 760–761, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235264.

Burkhardt:1996:BRBb

[2549] Richard W. Burkhardt, Jr. Book review: The Natural Science of the Human Species: An Introduction to Comparative Behavioral Research:

The “Russian Manuscript” (1944–1948) by Konrad Lorenz; Agnes von Cranach; Robert D. Martin. Isis, 87(4):761–762, December 1996. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235265.

Renner:1996:BRB

[2550] Soren Renner and Arthur Fine. Book review: The Mangle of Practice: Time, Agency, and Science by Andrew Pickering. Isis, 87(4):762–764, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235266.

Rouse:1996:BRBb

[2551] Joseph Rouse. Book review: Scientific Knowledge: A Sociological Analysis by Barry Barnes; David Bloor; John Henry. Isis, 87(4):764–766, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235267.

Amann:1996:BRB

[2552] Klaus Amann. Book review: Scientific Practice and Ordinary Action: Ethnomethodology and Social Studies of Science by Michael Lynch. Isis, 87(4):766–767, December 1996. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235268.

Gieryn:1996:BRB

[2553] Thomas F. Gieryn. Book review: At the Fringes of Science by Michael W. Friedlander. Isis, 87(4):767–768, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235269.

Cuomo:1996:BRB

[2554] Chris J. Cuomo. Book review: The New Ecological Order by Luc Ferry; Carol Volk. Isis, 87(4):768–769, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235270.

Lucier:1996:BRB

[2555] Paul Lucier. Book review: History of Science in the United States: A Chronology and Research Guide by Clark A. Elliott. Isis, 87(4):769–770, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235271.

Anonymous:1996:Cd

[2556] Anonymous. Collections. Isis, 87(4):770–773, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235272.

Anonymous:1996:BMe

[2557] Anonymous. Back matter. Isis, 87(4):774–778, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235273.

Anonymous:1996:VI

[2558] Anonymous. Volume information. Isis, 87(4):779–799, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235193.

Anonymous:1996:FMe

[2559] Anonymous. Front matter. Isis, 87(4):??, December 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235194.

Neu:1997:CBH

[2560] John Neu. Current bibliography of the history of science and its cultural influences, 1997. Isis, 88(??):1–308, ???? 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237999.

Anonymous:1997:BMa

[2561] Anonymous. Back matter. Isis, 88(??):??, ???? 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/238000.

Anonymous:1997:FMa

[2562] Anonymous. Front matter. Isis, 88(??):i–xvi, ???? 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237998.

Long:1997:PPA

[2563] Pamela O. Long. Power, patronage, and the authorship of ars: From mechanical know-how to mechanical knowledge in the last scribal age. Isis, 88(1):1–41, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235824.

Ribe:1997:COM

[2564] Neil M. Ribe. Cartesian optics and the mastery of nature. Isis, 88(1): 42–61, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235825.

Cooke:1997:SPP

[2565] Kathy J. Cooke. From science to practice, or practice to science? Chickens and eggs in Raymond Pearl’s agricultural breeding research, 1907– 1916. Isis, 88(1):62–86, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235826.

Novak:1997:LBL

[2566] Steven J. Novak. LSD before Leary: Sidney Cohen’s critique of 1950s psychedelic drug research. Isis, 88(1):87–110, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235827.

Ruse:1997:BRB

[2567] Michael Ruse. Book review: Darwin et l’apres-Darwin: Une histoire de l’hypothese de selection naturelle. Isis, 88(1):111–117, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235828.

Schuster:1997:BRB

[2568] John A. Schuster. Book review: The Scientific Revolution: A Historiographical Inquiry. Isis, 88(1):118–121, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235829.

Gooding:1997:BRB

[2569] David Gooding. Book review: Scientific Practice: Theories and Stories of Doing Physics by Jed Z. Buchwald. Isis, 88(1):121–122, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235830.

Barker:1997:BRB

[2570] Peter Barker. Book review: Discipline and Experience: The Mathematical Way in the Scientific Revolution by Peter Dear. Isis, 88(1):122–124, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235831.

Gabbey:1997:BRB

[2571] Alan Gabbey. Book review: Physiologia: Natural Philosophy in Late Aristotelian and Cartesian Thought by Dennis Des Chene. Isis, 88(1): 124, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235832.

Nadler:1997:BRB

[2572] Steven Nadler. Book review: Perception and Reality: A History from Descartes to Kant by John W. Yolton. Isis, 88(1):124–125, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235833.

Paul:1997:BRBa

[2573] Harry W. Paul. Book review: L’autorite de la science: Neurosciences, espaces et temps, chaos, cosmologie by François Lurcat. Isis, 88(1):125– 126, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235834.

Cittadino:1997:BRB

[2574] Eugene Cittadino. Book review: The Norton History of the Environmental Sciences by Peter J. Bowler. Isis, 88(1):126–127, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235835.

Weidman:1997:BRBa

[2575] Nadine Weidman. Book review: A History of Psychology by John G. Benjafield. Isis, 88(1):127–128, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235836.

Misa:1997:BRBa

[2576] Thomas J. Misa. Book review: American Iron, 1607–1900 by Robert B. Gordon. Isis, 88(1):128–129, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235837.

Inkster:1997:BRB

[2577] Ian Inkster. Book review: The Technological Transformation of Japan: From the Seventeenth to the Twenty-First Century by Tessa Morris-Suzuki. Isis, 88(1):129–130, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235838.

DiCosmo:1997:BRB

[2578] Nicola Di Cosmo. Book review: Where the Roads Met: East and West in the Silk Production Processes (Seventeenth to Nineteenth Century) by Claudio Zanier. Isis, 88(1):130–131, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235839.

Dick:1997:BRB

[2579] Steven J. Dick. Book review: Ancient Astronomers by Anthony F. Aveni. Isis, 88(1):131–132, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235840.

Caldwell:1997:BRB

[2580] John Caldwell. Book review: Institution arithmetique by Boece; Jean-Yves Guillaumin. Isis, 88(1):132–133, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235841.

Heinen:1997:BRB

[2581] Anton M. Heinen. Book review: Tribiblos astronomique: Livre II by Theodore Meliteniote; Regine Leurquin; Tribiblos astronomique: Livre II: Commentaire by Theodore Meliteniote; Regine Leurquin. Isis, 88(1): 133–134, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235842.

Eastwood:1997:BRB

[2582] Bruce S. Eastwood. Book review: Byrhtferth’s Enchiridion by Byrhtferth; Peter S. Baker; Michael Lapidge. Isis, 88(1):134–136, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235843.

Moreton:1997:BRBa

[2583] Jennifer Moreton. Book review: Die Physica und der Liber de causis im 12 Jahrhundert: Zwei Studien by Thomas Ricklin. Isis, 88(1):136, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235844.

McEvoy:1997:BRB

[2584] James McEvoy. Book review: The Problem of the Rational Soul in the Thirteenth Century by Richard C. Dales. Isis, 88(1):136–137, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235845.

Seidengart:1997:BRB

[2585] Jean Seidengart. Book review: Metaphysik und Mathematik bei Giordano Bruno by Angelika Bonker-Vallon. Isis, 88(1):137–138, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235846.

Kemp:1997:BRB

[2586] Martin Kemp. Book review: La fabbrica del corpo: Libri e dissectione nel Rinascimento by Andrea Carlino. Isis, 88(1):138–139, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235847.

Carmichael:1997:BRBa

[2587] Ann G. Carmichael. Book review: De abditis nonnullis ac mirandis morborum et sanationum causis by Antonio Benivieni; Giorgio Weber. Isis, 88(1):139–140, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235848.

Malet:1997:BRB

[2588] Antoni Malet. Book review: Philosophy of Mathematics and Mathematical Practice in the Seventeenth Century by Paolo Mancosu. Isis, 88(1): 140–141, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235849.

Finocchiaro:1997:BRB

[2589] Maurice A. Finocchiaro. Book review: Contro Galileo: Alle origini dell’affaire by Massimo Bucciantini. Isis, 88(1):141–142, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235850.

DiLiscia:1997:BRB

[2590] Daniel A. Di Liscia. Book review: Nel segno di Galileo: Erudizione, filosofia e scienza a Firenze nel secolo XVII: I “Trattati accademici” di Vincenzio Capponi by Stefano Caroti. Isis, 88(1):142–143, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235851.

Chavez-Arvizo:1997:BRB

[2591] Enrique Chavez-Arvizo. Book review: Descartes’s Imagination: Proportion, Images, and the Activity of Thinking by Dennis L. Sepper. Isis, 88(1):143–144, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235852.

Mahoney:1997:BRB

[2592] Michael S. Mahoney. Book review: Un theoreme de Fermat et ses lecteurs by Catherine Goldstein. Isis, 88(1):144–145, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235853.

Grene:1997:BRBa

[2593] Marjorie Grene. Book review: The Cambridge Companion to Spinoza by Don Garrett. Isis, 88(1):145–146, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235854.

Iliffe:1997:BRB

[2594] Rob Iliffe. Book review: Newton’s Optical Writings: A Guided Study by Dannis L. Sepper. Isis, 88(1):146–147, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235855.

Gough:1997:BRB

[2595] Jerry B. Gough. Book review: Laws and Order in Eighteenth-Century Chemistry by Alistair Duncan. Isis, 88(1):147, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235856.

Bensaude-Vincent:1997:BRB

[2596] Bernadette Bensaude-Vincent. Book review: Bibliotheca Lavoisieriana: The Catalogue of the Library of Antoine Laurent Lavoisier by Marco Beretta. Isis, 88(1):147–148, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235857.

Baker:1997:BRB

[2597] Keith Michael Baker. Book review: Foundations of Social Choice and Political Theory by Condorcet; Iain McLean; Fiona Hewitt. Isis, 88(1): 148–149, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235858.

Lawrence:1997:BRBa

[2598] Christopher Lawrence. Book review: Joseph Banks and the English Enlightenment: Useful Knowledge and Polite Culture by John Gascoigne. Isis, 88(1):149–150, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235859.

Mercier:1997:BRB

[2599] Raymond Mercier. Book review: Sawai Jai Singh and His Astronomy by Virendra Nath Sharma. Isis, 88(1):150–151, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235860.

Cueto:1997:BRB

[2600] Marcos Cueto. Book review: Las heridas de la ciencia by Jose Luis Peset. Isis, 88(1):151, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235861.

Zeller:1997:BRB

[2601] Suzanne Zeller. Book review: John William Dawson: Faith, Hope, and Science by Susan Sheets-Pyenson. Isis, 88(1):151–153, March 1997. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235862.

Hannaway:1997:BRB

[2602] Caroline Hannaway. Book review: The Medical Mandarins: The French Academy of Medicine in the Nineteenth and Early Twentieth Centuries by George Weisz. Isis, 88(1):153–154, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235863.

Rosner:1997:BRB

[2603] Lisa Rosner. Book review: “To Ask the Suffrages of the Patrons”: Thomas Laycock and the Edinburgh Chair of Medicine, 1855 by Michael Barfoot. Isis, 88(1):154–155, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235864.

Borossa:1997:BRB

[2604] Julia Borossa. Book review: The Correspondence of Sigmund Freud and Sandor Ferenczi. Volume 2: 1914–1919 by Sigmund Freud; Sandor Ferenczi; Ernst Falzeder; Eva Brabant; Patricia Giampieri-Deutsch; Ingeborg Meyer-Palmedo; Peter T. Hoffer. Isis, 88(1):155–156, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235865.

Shortland:1997:BRB

[2605] Michael Shortland. Book review: Einstein Lived Here by Abraham Pais. Isis, 88(1):156–157, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235866.

Walker:1997:BRBa

[2606] Mark Walker. Book review: Physics and National Socialism: An Anthology of Primary Sources by Klaus Hentschel; Ann Hentschel. Isis, 88(1):157–158, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235867.

Leslie:1997:BRB

[2607] Stuart W. Leslie. Book review: Scientists and the Development of Nuclear Weapons: From Fission to the Limited Test Ban Treaty, 1939– 1963 by Lawrence Badash. Isis, 88(1):158–159, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235868.

Edwards:1997:BRBa

[2608] Paul N. Edwards. Book review: Refiguring Life: Metaphors of Twentieth–Century Biology by Evelyn Fox Keller. Isis, 88(1):159–160, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/235869.

Noll:1997:BRB

[2609] Mark A. Noll. Book review: Creationism in Twentieth–Century America: A Ten-Volume Anthology of Documents, 1903–1961 by Ronald L. Numbers; William Vance Trollinger,; Paul Nelson; Edward B. Davis; Mark A. Kalthoff. Isis, 88(1):160–162, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235870.

Stocking:1997:BRBa

[2610] George W. Stocking, Jr. Book review: Rethinking Race: Franz Boas and His Contemporaries by Vernon J. Williams. Isis, 88(1):162–163, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235871.

Bix:1997:BRB

[2611] Amy Sue Bix. Book review: The Rhetoric of Eugenics in Anglo– American Thought by Marouf A. Hasian. Isis, 88(1):163–164, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/235872.

Rothman:1997:BRBa

[2612] David J. Rothman. Book review: Subjected to Science: Human Experimentation in America before the Second World War by Susan E. Lederer. Isis, 88(1):164–165, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235873.

Risse:1997:BRB

[2613] Guenter B. Risse. Book review: A Summer Plague: Polio and Its Survivors by Tony Gould. Isis, 88(1):165–166, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235874.

Roll-Hansen:1997:BRB

[2614] Nils Roll-Hansen. Book review: Medizin und Wissenschaftstheorie: Die Geschichte der Wismuttherapie als wissenschaftstheoretische Fallstudie by Alfred Reckenfelderbaumer. Isis, 88(1):166–167, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235875.

Rasmussen:1997:BRB

[2615] Nicolas Rasmussen. Book review: Zur Fruhgeschichte des Elektronenmikroskops by Lin Qing. Isis, 88(1):167, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/235876.

Hevly:1997:BRB

[2616] Bruce Hevly. Book review: Science on the Run: Information Management and Industrial Geophysics at Schlumberger, 1920–1940 by Geoffrey C. Bowker. Isis, 88(1):167–169, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235877.

Rome:1997:BRB

[2617] Adam W. Rome. Book review: Lewis Mumford and the Ecological Region: The Politics of Planning by Mark Luccarelli. Isis, 88(1):169, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235878.

Segre:1997:BRB

[2618] Michael Segre. Book review: Des nobel au Vatican: La fondation de l’Academie pontificale des Sciences by Regis Ladous. Isis, 88(1):169– 170, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/235879.

Ghiselin:1997:BRB

[2619] Michael T. Ghiselin. Book review: Life Scientists: Their Convictions, Their Activities, and Their Values by Gerard M. Verschuuren. Isis, 88 (1):170–171, March 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/235880.

Gottweis:1997:BRBa

[2620] Herbert Gottweis. Book review: Frauen und Naturwissenschaften: Gesammelte Vortrage zur feministischen Wissenschaftskritik by Erika Hickel. Isis, 88(1):171–172, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235881.

Wermiel:1997:BRB

[2621] Sara Wermiel. Book review: The Tradition of Technology: Landmarks of Western Technology in the Collections of the Library of Congress by Leonard C. Bruno. Isis, 88(1):172–173, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235882.

Cassidy:1997:BRBa

[2622] James G. Cassidy. Book review: Benjamin Rush, M.D.: A Bibliographic Guide by Claire G. Fox; Gordon L. Miller; Jacquelyn C. Miller. Isis, 88 (1):173, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/235883.

Anonymous:1997:Ca

[2623] Anonymous. Collections. Isis, 88(1):173–177, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235884.

Osler:1997:ERW

[2624] Margaret J. Osler. Eloge: Richard S. Westfall, 22 April 1924–21 August 1996. Isis, 88(1):178–181, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

235885.

Fullmer:1997:EJM

[2625] June Z. Fullmer. Eloge: Jane Marion Oppenheimer, 19 September 1911– 19 March 1996. Isis, 88(1):181–183, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235886.

Fox:1997:EAC

[2626] Robert Fox. Eloge: Alistair Cameron Crombie, 4 November 1915–9 February 1996. Isis, 88(1):183–186, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/235887.

Parshall:1997:AMH

[2627] Karen V. H. Parshall, Paul H. Theerman, Keith R. Benson, and Charles C. Gillispie. Annual meeting of the History of Science Society: Atlanta, 7–10 November 1996. Isis, 88(1):187–194, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/235888.

Anonymous:1997:BMb

[2628] Anonymous. Back matter. Isis, 88(1):195–204, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235889.

Anonymous:1997:FMb

[2629] Anonymous. Front matter. Isis, 88(1):??, March 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/235823.

Todes:1997:PPF

[2630] Daniel P. Todes. Pavlov’s physiology factory. Isis, 88(2):205–246, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236572.

Harkness:1997:MEH

[2631] Deborah E. Harkness. Managing an experimental household: The Dees of Mortlake and the practice of natural philosophy. Isis, 88(2):247–262, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236573.

Davenport:1997:CCC

[2632] Anne A. Davenport. The Catholics, the Cathars, and the concept of infinity in the Thirteenth Century. Isis, 88(2):263–295, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236574.

Butrica:1997:HCJ

[2633] Andrew J. Butrica. Historical collections in jeopardy: The societe d’encouragement pour l’industrie nationale. Isis, 88(2):296–301, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236575.

Gillispie:1997:ECS

[2634] Charles C. Gillispie. Eloge: Charles Scribner, Jr., 13 July 1921–11 November 1995. Isis, 88(2):302–303, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236576.

Ferrari:1997:LE

[2635] Roberto A. Ferrari, Marcos Cueto, Daniel Siegel, M. W. Friedlander, and Thomas F. Gieryn. Letters to the Editor. Isis, 88(2):304–305, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236577.

Camerini:1997:BRB

[2636] Jane Camerini. Book review: Huxley: The Devil’s Disciple. Isis, 88(2): 306–311, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236578.

Gregory:1997:BRB

[2637] Frederick Gregory. Book review: The Flight from Science and Reason. Isis, 88(2):312–315, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236579.

Swerdlow:1997:BRB

[2638] N. M. Swerdlow. Book review: The Norton History of Astronomy and Cosmology. Isis, 88(2):316–317, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236580.

Badash:1997:BRB

[2639] Lawrence Badash. Book review: Marie Curie: A Life by Susan Quinn. Isis, 88(2):318–319, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236581.

Schweber:1997:BRB

[2640] Sam Schweber. Book review: Twentieth Century Physics by Laurie M. Brown; Abraham Pais; Brian Pippard. Isis, 88(2):319–321, June 1997.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236582.

Filner:1997:BRB

[2641] Bob Filner. Book review: The Politics of Expertise in Congress: The Rise and Fall of the Office of Technology Assessment by Bruce Bimber; Twenty Years of Science in the Public Interest: A History of the Congressional Science and Engineering Fellowship Program by Jeffrey K. Stine. Isis, 88(2):321–322, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236583.

Edwards:1997:BRBb

[2642] Paul N. Edwards. Book review: Aramis, or the Love of Technology by Bruno Latour; Catherine Porter. Isis, 88(2):322–324, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236584.

Strauss:1997:BRB

[2643] David Strauss. Book review: The Planet Mars: A History of Observation and Discovery by William Sheehan. Isis, 88(2):324–325, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236585.

Turner:1997:BRB

[2644] A. J. Turner. Book review: Instruments and the Imagination by Thomas L. Hankins; Robert J. Silverman. Isis, 88(2):325–327, June 1997. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236586.

Bennett:1997:BRB

[2645] J. A. Bennett. Book review: Instrumentos de Navegacion: Del Mediterraneo al Pacifico by Manuel Selles. Isis, 88(2):327, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236587.

Doienel:1997:BRB

[2646] Hans-Liudger Doienel. Book review: Technische Universität Braunschweig: Vom Collegium Carolinum zur Technischen Universität, 1745– 1995 by Walter Kertz. Isis, 88(2):327–328, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236588.

Baumer-Schleinkofer:1997:BRB

[2647] Anne Baumer-Schleinkofer. Book review: Vom mittelalterlichen Medizinzum modernen Biologieunterricht by Michael Freyer. Isis, 88(2):328– 329, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236589.

deMontellano:1997:BRB

[2648] Bernard Ortiz de Montellano. Book review: Medical Ethnobiology of the Highland Maya of Chiapas, Mexico: The Gastrointestinal Diseases by Elois Ann Berlin; Brent Berlin. Isis, 88(2):329–330, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236590.

McKirahan:1997:BRB

[2649] Richard McKirahan. Book review: Explaining an Eclipse: Aristotle’s Posterior Analytics 2.1-10 by Owen Goldin. Isis, 88(2):330–331, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236591.

Hine:1997:BRB

[2650] William L. Hine. Book review: Between Copernicus and Galileo: Christoph Clavius and the Collapse of Ptolemaic Cosmology by James M. Lattis. Isis, 88(2):331, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236592.

Laird:1997:BRB

[2651] W. R. Laird. Book review: The Geometry of War, 1500–1750: Catalogue of the Exhibition by Jim Bennett; Stephen Johnston. Isis, 88(2):331– 332, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236593.

Keller:1997:BRB

[2652] Alex Keller. Book review: Labour, Science, and Technology in France, 1500–1620 by Henry Heller. Isis, 88(2):332–334, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236594.

Carmichael:1997:BRBb

[2653] Ann G. Carmichael. Book review: Plague? Jesuit Accounts of Epidemic Disease in the Sixteenth Century by A. Lynn Martin. Isis, 88(2):334, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236595.

French:1997:BRB

[2654] Roger French. Book review: Sources for the History of Medicine in Late Medieval England by Carole Rawcliffe. Isis, 88(2):334–335, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236596.

Macgill:1997:BRB

[2655] Elizabeth Rawson Macgill. Book review: A New Herball: Part 1 by William Turner; George T. L. Chapman; Marilyn N. Tweedle; A New Herball: Parts 2 and 3 by William Turner; George T. L. Chapman; Frank McCombie; Anne Wesencraft. Isis, 88(2):335, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236597.

Ranea:1997:BRB

[2656] Alberto Guillermo Ranea. Book review: El mito de Vesalio by Juan Jose Barcia Goyanes. Isis, 88(2):335–336, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236598.

Osler:1997:BRB

[2657] Margaret J. Osler. Book review: The Matter of Revolution: Science, Poetry, and Politics in the Age of Milton by John Rogers. Isis, 88(2): 336–337, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236599.

DesChene:1997:BRB

[2658] Dennis Des Chene. Book review: Descartes et Regius: Autour de l’explication de l’esprit humain by Theo Verbeek. Isis, 88(2):337–338, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236600.

Ariew:1997:BRBa

[2659] Roger Ariew. Book review: Gassendi’s Ethics: Freedom in a Mechanistic Universe by Lisa T. Sarasohn. Isis, 88(2):338–339, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236601.

Grene:1997:BRBb

[2660] Marjorie Grene. Book review: Leibniz and the Kabbalah by Allison P. Coudert. Isis, 88(2):339–340, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236602.

Arthur:1997:BRB

[2661] Richard Arthur. Book review: Mathematischer, naturwissenschaftlicher und technischer Briefwechsel. Volume 4: Juli 1683–1690 by Gottfried Wilhelm Leibniz. Isis, 88(2):340–341, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236603.

Fouke:1997:BRB

[2662] Daniel C. Fouke. Book review: “Be Sober and Reasonable”: The Critique of Enthusiasm in the Seventeenth and Early Eighteenth Centuries by Michael Heyd. Isis, 88(2):341–342, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236604.

Multhauf:1997:BRB

[2663] Robert P. Multhauf. Book review: Water Which Does Not Wet Hands: The Alchemy of Michael Sendivogius by Zbigniew Szydlo. Isis, 88(2): 342–343, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236605.

Popplestone:1997:BRB

[2664] John A. Popplestone. Book review: Machine Man and Other Writings by Julien Offray de La Mettrie; Ann Thomson. Isis, 88(2):343–344, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236606.

Marshall:1997:BRB

[2665] P. J. Marshall. Book review: Visions of Empire: Voyages, Botany, and Representations of Nature by David Philip Miller; Peter Hahns Reill. Isis, 88(2):344, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236607.

Lawrence:1997:BRBb

[2666] Christopher Lawrence. Book review: Charitable Knowledge: Hospital Pupils and Practitioners in Eighteenth-Century London by Susan C. Lawrence. Isis, 88(2):344–345, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236608.

Pickering:1997:BRB

[2667] Mary Pickering. Book review: Philosophie des sciences by Auguste Comte; Juliette Grange; La philosophie d’Auguste Comte: Science, politique, religion by Juliette Grange. Isis, 88(2):345–346, June 1997. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236609.

Kring:1997:BRB

[2668] David A. Kring. Book review: The Port Orford, Oregon, Meteorite Mystery by Roy S. Clarke. Isis, 88(2):346–347, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236610.

Paul:1997:BRBb

[2669] John J. Paul. Book review: Science and the Raj, 1857–1905 by Deepak Kumar. Isis, 88(2):347–348, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236611.

Mills:1997:BRBa

[2670] Eric L. Mills. Book review: Baird’s Legacy: The History and Accomplishments of NOAA’s National Marine Fisheries Service, 1871–1996 by W. L. Hobart. Isis, 88(2):348–349, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236612.

LaBerge:1997:BRB

[2671] Ann F. La Berge. Book review: Cholera in Post-Revolutionary Paris: A Cultural History by Catherine J. Kudlick. Isis, 88(2):349–350, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236613.

Simon:1997:BRB

[2672] Linda Simon. Book review: William James on Consciousness beyond the Margin by Eugene Taylor. Isis, 88(2):350, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236614.

Tighe:1997:BRB

[2673] Janet A. Tighe. Book review: Moonlight, Magnolias, and Madness: Insanity in South Carolina from the Colonial Period to the Progressive Era by Peter McCandless. Isis, 88(2):350–351, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236615.

Cooter:1997:BRBa

[2674] Roger Cooter. Book review: Medical Services and the Hospitals in Britain, 1860–1939 by Steven Cherry. Isis, 88(2):351–352, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236616.

Misa:1997:BRBb

[2675] Thomas J. Misa. Book review: Rise of the New York Skyscraper, 1865– 1913 by Sarah Bradford Landau; Carl W. Condit. Isis, 88(2):352–353, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236617.

Morus:1997:BRB

[2676] Iwan Rhys Morus. Book review: Records of the Dawn of Photography: Talbot’s Notebooks P and Q by Larry J. Schaaf. Isis, 88(2):353–354, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236618.

Brooks:1997:BRBa

[2677] Nathan M. Brooks. Book review: The Electrification of Russia, 1880– 1926 by Jonathan Coopersmith. Isis, 88(2):354–355, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236619.

Assmus:1997:BRB

[2678] Alexi Assmus. Book review: A Memoir on The Physical Review: A History of the First Hundred Years by Paul Hartman. Isis, 88(2):355, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236620.

Grattan-Guinness:1997:BRB

[2679] Ivor Grattan-Guinness. Book review: Die Godelschen Unvollstandigkeitssatze: Zur Geschichte ihrer Entstehung und Rezeption by Christel Ketelsen. Isis, 88(2):355–356, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236621.

Dehue:1997:BRB

[2680] Trudy Dehue. Book review: De wereld van instinct: Niko Tinbergen en het ontstaan van de ethologie in Nederland (1920–1950) by D. R. Roell.

Isis, 88(2):356–357, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236622.

Berrios-Ortiz:1997:BRB

[2681] Angel Berrios-Ortiz. Book review: The Scientific Survey of Puerto Rico and the Virgin Islands: An Eighty-Year Reassessment of the Islands’ Natural History by Julio C. Figueroa Colon. Isis, 88(2):357, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236623.

Landes:1997:BRB

[2682] David S. Landes. Book review: Science, Technology, and the British Industrial “Decline,” 1870–1970 by David Edgerton. Isis, 88(2):357– 358, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236624.

Westfall:1997:BRB

[2683] Catherine Westfall. Book review: MIT Lincoln Laboratory: Technology in the National Interest by Eva C. Freeman. Isis, 88(2):358–359, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236625.

Agar:1997:BRB

[2684] Jon Agar. Book review: The History of Radio Astronomy and the National Radio Astronomy Observatory: Evolution toward Big Science by Benjamin K. Malphrus. Isis, 88(2):359–361, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236626.

Tatarewicz:1997:BRB

[2685] Joseph N. Tatarewicz. Book review: Exploring the Unknown: Selected Documents in the History of the U.S. Civil Space Program. Volume 1: Organizing for Exploration by John M. Logsdon; Linda J. Lear; Jannelle Warren-Findley; Ray A. Williamson; Dwayne A. Day. Isis, 88(2):361– 362, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236627.

Walker:1997:BRBb

[2686] J. Samuel Walker. Book review: Industrial Policy, Technology, and International Bargaining: Designing Nuclear Industries in Argentina and Brazil by Etel Solingen. Isis, 88(2):362, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236628.

Foltz:1997:BRB

[2687] Franz Foltz. Book review: A Half Century of Peer Review, 1946–1996 by Richard Mandel. Isis, 88(2):362–363, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236629.

Servos:1997:BRB

[2688] John W. Servos. Book review: The Making of an Engineer: An Illustrated History of Engineering Education in the United States and Canada by Lawrence P. Grayson. Isis, 88(2):363–364, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236630.

Haller:1997:BRB

[2689] Rudolf Haller. Book review: Ludwig Boltzmann: His Later Life and Philosophy, 1900–1906. Book 1: A Documentary History by John Blackmore. Isis, 88(2):364–365, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236631.

Hawkins:1997:BRB

[2690] Hugh Hawkins. Book review: The Strands of a Life: The Science of DNA and the Art of Education by Robert L. Sinsheimer. Isis, 88(2):365– 366, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236632.

Buxton:1997:BRB

[2691] William J. Buxton. Book review: Pitirim A. Sorokin: An Intellectual Biography by Barry V. Johnston. Isis, 88(2):366–367, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236633.

Back:1997:BRB

[2692] Kurt W. Back. Book review: Aldous Huxley and the Mysticism of Science by June Deery. Isis, 88(2):367–368, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236634.

Wiegman:1997:BRB

[2693] Nanny Wiegman. Book review: “Collegialiteit moet hier ons wachtwoord zijn”: De geschiedenis van de Onderlinge Verzekering-Maatschappij van Geneeskundigen tegen de geldelijke gevolgen van invaliditeit. “Artsen-Onderlinge,” 1896–1996. [“Collegiality Must Be Our Watchword Here”:

The History of the Mutual Insurance Company of Medical Practitioners against the Financial Consequences of Disablement, “Doctors’ Mutual,” 1896–1996.] by M. J. van Lieburg. Isis, 88(2):368–369, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236635.

Mehler:1997:BRB

[2694] Barry Mehler. Book review: Controlling Human Heredity: 1865 to the Present by Diane B. Paul. Isis, 88(2):369, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236636.

Apple:1997:BRB

[2695] Rima D. Apple. Book review: The Black Stork: Eugenics and the Death of “Defective” Babies in American Medicine and Motion Pictures since 1915 by Martin S. Pernick. Isis, 88(2):369–370, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236637.

Stocking:1997:BRBb

[2696] George W. Stocking, Jr. Book review: Scientific Racism in Modern South Africa by Saul Dubow. Isis, 88(2):370–371, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236638.

Rothman:1997:BRBb

[2697] David J. Rothman. Book review: Crossing Frontiers: Gerontology Emerges as a Science by Andrew W. Achenbaum. Isis, 88(2):371–372, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236639.

Graebner:1997:BRB

[2698] William Graebner. Book review: Disciplining Old Age: The Formation of Gerontological Knowledge by Stephen Katz. Isis, 88(2):373, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236640.

Trent:1997:BRB

[2699] James T. Trent. Book review: Feeble-Minded in Our Midst: Institutions for the Mentally Retarded in the South, 1900–1940 by Steven Noll. Isis, 88(2):373–375, June 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236641.

Brown:1997:BRB

[2700] Joanne Brown. Book review: Freud’s Paranoid Quest: Psychoanalysis and Modern Suspicion by John Farrell; Freud and the Politics of Psychoanalysis by Jose Brunner. Isis, 88(2):375–376, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236642.

Horvath:1997:BRB

[2701] Christopher D. Horvath. Book review: Queer Science: The Use and Abuse of Research into Homosexuality by Simon LeVay. Isis, 88(2):376– 377, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236643.

Russell:1997:BRB

[2702] Colin A. Russell. Book review: Religion, Science, and Naturalism by Willem B. Drees. Isis, 88(2):377–379, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236644.

Tympas:1997:BRB

[2703] Aristotle Tympas. Book review: Of Bicycles, Bakelites, and Bulbs: Toward a Theory of Sociotechnical Change by Wiebe Bijker. Isis, 88(2): 379, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236645.

Nickles:1997:BRBa

[2704] Thomas Nickles. Book review: Engaging Science: How to Understand Its Practices Philosophically by Joseph Rouse. Isis, 88(2):379–381, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236646.

Puelles:1997:BRB

[2705] Andoni Alonso Puelles. Book review: In and about the World: Philosophical Studies of Science and Technology by Hans Radder. Isis, 88(2): 381–382, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236647.

Searing:1997:BRB

[2706] Susan E. Searing. Book review: Women and Science: An Annotated Bibliography by Marilyn Bailey Ogilvie; Kerry Lynne Meek; Notable Women in the Life Sciences: A Biographical Dictionary by Benjamin F. Shearer; Barbara S. Shearer. Isis, 88(2):382–383, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236648.

Anonymous:1997:Cb

[2707] Anonymous. Collections. Isis, 88(2):383–391, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236649.

Anonymous:1997:BMc

[2708] Anonymous. Back matter. Isis, 88(2):392–396, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236650.

Anonymous:1997:FMc

[2709] Anonymous. Front matter. Isis, 88(2):??, June 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236571.

Nye:1997:ACP

[2710] Mary Jo Nye. Aristocratic culture and the pursuit of science: The de Broglies in modern France. Isis, 88(3):397–421, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236150.

Richmond:1997:LOO

[2711] Marsha L. Richmond. “A Lab of One’s Own”: The Balfour Biological Laboratory for Women at Cambridge University, 1884–1914. Isis, 88 (3):422–455, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236151.

Schaffer:1997:ETD

[2712] Simon Schaffer. Experimenters’ techniques, dyers’ hands, and the electric planetarium. Isis, 88(3):456–483, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236152.

Segre:1997:LGC

[2713] Michael Segre. Light on the Galileo case? Isis, 88(3):484–504, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236153.

Stapleton:1997:DJS

[2714] Darwin H. Stapleton. Dirk J. Struik’s Yankee science in the making: A half-century retrospective. Isis, 88(3):505–511, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236154.

Josephson:1997:EVY

[2715] Paul Josephson. Eloge: Viktor Yakovlevich Frenkel, 22 February 1930–8 February 1997. Isis, 88(3):512–513, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236155.

Sheets-Pyenson:1997:LE

[2716] Susan Sheets-Pyenson. Letters to the Editor. Isis, 88(3):514, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236156.

Williams:1997:BRBa

[2717] L. Pearce Williams. Book review: Andre-Marie Ampere: Enlightenment and Electrodynamics. Isis, 88(3):515–517, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236157.

Downes:1997:BRB

[2718] Stephen Downes. Book review: The Disunity of Science: Boundaries, Contexts, and Power by Peter Galison; David J. Stump. Isis, 88(3): 517–518, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236158.

Gates:1997:BRB

[2719] Barbara T. Gates. Book review: Open Fields: Science in Cultural Encounter by Gillian Beer. Isis, 88(3):518–519, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236159.

Rydell:1997:BRB

[2720] Robert W. Rydell. Book review: Conjuring Science: Scientific Symbols and Cultural Meanings in American Life by Christopher P. Toumey. Isis, 88(3):519, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236160.

Teich:1997:BRB

[2721] Albert H. Teich. Book review: Foundations of Futures Studies: Human Science for a New Era by Wendell Bell. Isis, 88(3):519–520, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236161.

Mukerji:1997:BRB

[2722] Chandra Mukerji. Book review: On Collecting: An Investigation into Collecting in the European Tradition by Susan M. Pearce. Isis, 88(3): 520–521, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236162.

Genovese:1997:BRB

[2723] Rosalie G. Genovese. Book review: Creative Couples in the Sciences by Helena M. Pycior; Nancy G. Slack; Pnina G. Abir-Am. Isis, 88(3): 521–522, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236163.

Margulis:1997:BRB

[2724] Lynn Margulis. Book review: Beliefs and Biology: Theories of Life and Living by Jennifer Trusted. Isis, 88(3):522–523, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236164.

Brieger:1997:BRB

[2725] Gert H. Brieger. Book review: The Cambridge Illustrated History of Medicine by Roy Porter. Isis, 88(3):523–524, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236165.

Acuna:1997:BRB

[2726] Leopoldo E. Acuna. Book review: Opera: Desire, Disease, Death by Linda Hutcheon; Michael Hutcheon. Isis, 88(3):524, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236166.

Schmaus:1997:BRB

[2727] Warren Schmaus. Book review: Essays in the History and Philosophy of Science by Pierre Duhem; Roger Ariew; Peter Barker. Isis, 88(3): 524–525, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236167.

Robinson:1997:BRB

[2728] David K. Robinson. Book review: Down from Olympus: Archaeology and Philhellenism in Germany, 1750–1970 by Suzanne L. Marchand. Isis, 88 (3):525–526, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236168.

Munns:1997:BRB

[2729] David Munns. Book review: Explorers of the Southern Sky: A History of Australian Astronomy by Raymond Haynes; Roslynn Haynes; David Malin; Richard McGee. Isis, 88(3):526–527, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236169.

Vallance:1997:BRB

[2730] J. T. Vallance. Book review: Cosmology in Antiquity by M. R. Wright. Isis, 88(3):527–528, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236170.

Taub:1997:BRB

[2731] Liba Taub. Book review: Euclid’s Phaenomena: A Translation and Study of a Hellenistic Treatise in Spherical Astronomy by Euclid; J. L. Berggren; R. S. D. Thomas. Isis, 88(3):528–529, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236171.

Fan:1997:BRB

[2732] Fa ti Fan. Book review: Astronomy and Mathematics in Ancient China: The Zhou bi suan jing by Christopher Cullen. Isis, 88(3):529–530, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236172.

Scarborough:1997:BRB

[2733] John Scarborough. Book review: Michigan Papyri XVII: The Michigan Medical Codex (P. Mich. 758 = P. Mich. inv. 21) by Louise C. Youtie. Isis, 88(3):530–531, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236173.

Grant:1997:BRB

[2734] Edward Grant. Book review: The Measure of Reality: Quantification and Western Society, 1250–1600 by Alfred W. Crosby. Isis, 88(3):531, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236174.

Molland:1997:BRB

[2735] George Molland. Book review: Before Science: The Invention of the Friars’ Natural Philosophy by Roger French; Andrew Cunningham. Isis, 88(3):532, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236175.

Tymoczko:1997:BRB

[2736] Maria Tymoczko. Book review: Understanding the Universe in Seventh-Century Ireland by Marina Smyth. Isis, 88(3):532–534, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236176.

Kuchta:1997:BRB

[2737] David Kuchta. Book review: The Body Emblazoned: Dissection and the Human Body in Renaissance Culture by Jonathan Sawday. Isis, 88(3): 534–535, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236177.

Hafter:1997:BRB

[2738] Daryl M. Hafter. Book review: Straws in the Wind: Medieval Urban Environmental Law–The Case of Northern Italy by Ronald Edward Zupko; Robert Anthony Laures. Isis, 88(3):535–536, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236178.

Feingold:1997:BRB

[2739] Mordechai Feingold. Book review: The Hartlib Papers: A Complete Text and Image Database of the Papers of Samuel Hartlib (ca. 1600– 1662) Held in Sheffield University Library, Sheffield, England by Samuel Hartlib. Isis, 88(3):536, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236179.

Snider:1997:BRB

[2740] Alvin Snider. Book review: Francis Bacon: The History of a Character Assassination by Nieves Mathews. Isis, 88(3):536–537, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236180.

Shea:1997:BRB

[2741] William Shea. Book review: Dialogo sobre los dos maximos sistemas del mundo ptolemaico y copernicano by Galileo Galilei; Antonio Beltran Mari. Isis, 88(3):537–538, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236181.

Poppi:1997:BRB

[2742] Antonino Poppi. Book review: Venetischer Aristotelismus im Ende der aristotelischen Welt: Aspekte der Welt und des Denkens des Cesare Cremonini (1550–1631) by Heinrich C. Kuhn. Isis, 88(3):538–539, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236182.

Ariew:1997:BRBb

[2743] Roger Ariew. Book review: Le monde, l’homme by Rene Descartes; Annie Bitbol-Hesperies; Jean–Pierre Verdet. Isis, 88(3):539–540, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236183.

Dhombres:1997:BRB

[2744] Jean Dhombres. Book review: Scritti di ottica: Inediti e rari (1644–1655) by Giovan Battista Hodierna; Corrado Dollo. Isis, 88(3):540, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236184.

Ford:1997:BRB

[2745] Brian J. Ford. Book review: The Microscope in the Dutch Republic: The Shaping of Discovery by Edward G. Ruestow. Isis, 88(3):540–541, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236185.

Picolet:1997:BRB

[2746] Guy Picolet. Book review: Science and Social Status: The Members of the Academie des Sciences, 1666–1750 by David J. Sturdy. Isis, 88(3): 541–542, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236186.

Liedman:1997:BRB

[2747] Sven-Eric Liedman. Book review: Zwischen Elfenbeinturm und Staatsdienst: Naturwissenschaftler des Ancien Regime im Spiegel von Condorcets Eloges des Academiciens by Martina Schmode. Isis, 88(3):542– 543, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236187.

Fleming:1997:BRB

[2748] Jim Fleming. Book review: Science and the Founding Fathers: Science in the Political Thought of Thomas Jefferson, Benjamin Franklin, John Adams, and James Madison by I. Bernard Cohen. Isis, 88(3):543–544, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236188.

Keel:1997:BRB

[2749] Othmar Keel. Book review: Calendar of the Correspondence of Dr. William Hunter, 1740–1783 by C. Helen Brock. Isis, 88(3):544–545, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236189.

Dooley:1997:BRB

[2750] Brendan Dooley. Book review: Autopsie della Terra: Illuminismo e geologia in Alberto Fortis (1741–1803) by Luca Ciancio. Isis, 88(3):545– 546, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236190.

Drayton:1997:BRB

[2751] Richard Drayton. Book review: Cultivating Women, Cultivating Science: Flora’s Daughters and Botany in England, 1760–1860 by Ann B. Shteir. Isis, 88(3):546–547, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236191.

Caneva:1997:BRB

[2752] Kenneth L. Caneva. Book review: Goethe in the History of Science by Frederick Amrine. Isis, 88(3):547–548, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236192.

Taylor:1997:BRB

[2753] Kenneth L. Taylor. Book review: North from the Hook: 150 Years of the Geological Survey of Ireland by Gordon L. Herries Davies. Isis, 88 (3):548–549, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236193.

Broman:1997:BRB

[2754] Thomas Broman. Book review: Friedrich Wilhelm Joseph Schelling und die Medizin der Romantik: Studien zu Schellings Wurzburger Periode by Werner E. Gerabek; Kunst oder Wissenschaft? Konzeptionen der Medizin in der deutschen Romantik by Urban Wiesing. Isis, 88(3):549– 550, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236194.

Kilmister:1997:BRB

[2755] C. W. Kilmister. Book review: Die Entwicklung des Tensorkalkuls: Vom absoluten Differentialkalkul zur Relativitatstheorie by Karin Reich. Isis, 88(3):550–551, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236195.

Morselli:1997:BRB

[2756] Mario Morselli. Book review: Amedeo Avogadro: Saggi e memorie sulla teorica atomica (1811–1838) by Amedeo Avogadro; Marco Ciardi; Lettere a Stanislao Cannizzaro: Scritti e carteggi, 1857–1862 by Leonello Paoloni. Isis, 88(3):551–553, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236196.

Feldenkirchen:1997:BRB

[2757] Wilfried Feldenkirchen. Book review: Technikwissenschaften: Die Entstehung der Elektrotechnik aus Industrie und Wissenschaft zwischen 1880 und 1914 by Wolfgang Konig. Isis, 88(3):553–554, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236197.

Walls:1997:BRB

[2758] Laura Dassow Walls. Book review: The Environmental Imagination: Thoreau, Nature Writing, and the Formation of American Culture by Lawrence Buell. Isis, 88(3):554–555, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236198.

Givens:1997:BRB

[2759] Douglas R. Givens. Book review: The Southwest in the American Imagination: The Writings of Sylvester Baxter, 1881–1889 by Curtis M. Hinsley; David R. Wilcox. Isis, 88(3):555–556, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236199.

Turrisi:1997:BRB

[2760] Patricia Turrisi. Book review: Peirce’s Philosophical Perspectives by Vincent G. Potter; Vincent M. Colapietro. Isis, 88(3):556–557, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236200.

Morantz-Sanchez:1997:BRB

[2761] Regina Morantz-Sánchez. Book review: Catching Babies: The Professionalization of Childbirth, 1870–1920 by Charlotte G. Borst. Isis, 88(3): 557, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236201.

Weidman:1997:BRBb

[2762] Nadine Weidman. Book review: Charcot: Constructing Neurology by Christopher Goetz; Michel Bonduelle; Toby Gelfand. Isis, 88(3):557–558, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236202.

Moulin:1997:BRB

[2763] Anne Marie Moulin. Book review: Per una filosofia della scienza sperimentale: La controversia Pasteur-Pouchet by Daniele Boccardi. Isis, 88 (3):558–560, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236203.

Anderson:1997:BRB

[2764] Warwick Anderson. Book review: Species and Specificity: An Interpretation of the History of Immunology by Pauline M. H. Mazumdar. Isis, 88(3):560, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236204.

Reisch:1997:BRB

[2765] George Reisch. Book review: Otto Neurath: Philosophy between Science and Politics by Nancy Cartwright; Jordi Cat; Lola Fleck; Thomas E. Uebel. Isis, 88(3):560–562, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236205.

Pyenson:1997:BRA

[2766] Lewis Pyenson. Book review: Albert Einstein, A. J. Kox, Martin J. Klein, Robert Schulmann, Jozsef Illy, Jean Eisenstaedt, Rita Fountain, and Annette Pringle: The Collected Papers of Albert Einstein. Volume 6: The Berlin Years: Writings, 1914–1917. Isis, 88(3):562–564, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236206.

Kauffman:1997:BRB

[2767] George B. Kauffman. Book review: The World Made New: Frederick Soddy, Science, Politics, and Environment by Linda Merricks. Isis, 88 (3):564–565, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236207.

Christianson:1997:BRB

[2768] Gale E. Christianson. Book review: Hubble’s Cosmology: A Guided Study of Selected Texts by Norriss S. Hetherington. Isis, 88(3):565, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236208.

Isham:1997:BRB

[2769] Chris J. Isham. Book review: Lemaître, Big Bang, and the Quantum Universe by Michael Heller. Isis, 88(3):565–566, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236209.

Smith:1997:BRBa

[2770] Robert W. Smith. Book review: The Very First Light: The True inside Story of the Scientific Journey Back to the Dawn of the Universe by John C. Mather; John Boslough. Isis, 88(3):566–567, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236210.

Hufbauer:1997:BRB

[2771] Karl Hufbauer. Book review: The Biological Universe: The Twentieth– Century Extraterrestrial Life Debate and the Limits of Science by Steven J. Dick. Isis, 88(3):567–568, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236211.

Olby:1997:BRB

[2772] Robert Olby. Book review: Force of Nature: The Life of Linus Pauling by Thomas Hager. Isis, 88(3):568–570, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236212.

Corn:1997:BRB

[2773] Joseph J. Corn. Book review: A Passion for Wings: Aviation and the Western Imagination, 1908–1918 by Robert Wohl. Isis, 88(3):570–571, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236213.

Matthews:1997:BRBa

[2774] Peter Hans Matthews. Book review: Keynes’s Uncertain Revolution by Bradley W. Bateman. Isis, 88(3):571–572, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236214.

Quiroga:1997:BRB

[2775] Virginia A. Metaxas Quiroga. Book review: The Machine in the Nursery: Incubator Technology and the Origins of Newborn Intensive Care by Jeffrey P. Baker. Isis, 88(3):572–573, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236215.

Seligson:1997:BRB

[2776] Rosemond D. Seligson. Book review: Occupational Therapy: The First Thirty Years, 1900 to 1930 by Virginia Anne Metaxas Quiroga. Isis, 88 (3):573–574, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236216.

Evans:1997:BRB

[2777] Hughes Evans. Book review: Screening the Body: Tracing Medicine’s Visual Culture by Lisa Cartwright. Isis, 88(3):574–575, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236217.

Howell:1997:BRB

[2778] Joel D. Howell. Book review: The Pioneers of NMR and Magnetic Resonance in Medicine: The Story of MRI by James Mattson; Merrill Simon. Isis, 88(3):575, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236218.

Gilman:1997:BRB

[2779] Sander L. Gilman. Book review: The History of Mental Symptoms: Descriptive Psychopathology since the Nineteenth Century by German E. Berrios. Isis, 88(3):575–576, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236219.

Jackson:1997:BRB

[2780] John P. Jackson. Book review: A Psychology of Difference: The American Lectures by Otto Rank; Robert Kramer. Isis, 88(3):576–577, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236220.

Gottweis:1997:BRBb

[2781] Herbert Gottweis. Book review: Molecular Politics: Developing American and British Regulatory Policy for Genetic Engineering, 1972–1982 by Susan Wright. Isis, 88(3):577–579, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236221.

Epstein:1997:BRB

[2782] Steven Epstein. Book review: AIDS and the Body Politic: Biomedicine and Sexual Difference by Catherine Waldby. Isis, 88(3):579, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236222.

Hood:1997:BRB

[2783] Clifton Hood. Book review: Horse Trails to Regional Rails: The Story of Public Transit in Greater Cleveland by James A. Toman; Blaine S. Hays. Isis, 88(3):579–580, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236223.

Post:1997:BRB

[2784] Robert C. Post. Book review: Living in a Technological Culture: Human Tools and Human Values by Mary Tiles; Hans Oberdiek. Isis, 88(3): 580–581, September 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236224.

Hamlin:1997:BRB

[2785] Christopher Hamlin. Book review: Prisoners of the Crystal Palace: Mapping and Understanding the Social and Cognitive Organization of Scientific Research Fields by Peter Stern. Isis, 88(3):581–582, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236225.

Manier:1997:BRB

[2786] Edward Manier. Book review: Inevitable Illusions: How Mistakes of Reason Rule Our Minds by Massimo Piattelli-Palmarini; Keith Botsford. Isis, 88(3):582–583, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236226.

Drake:1997:BRB

[2787] Ellen Tan Drake. Book review: Popular Physics and Astronomy: An Annotated Bibliography by Roger Smith. Isis, 88(3):583–584, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236227.

Burchsted:1997:BRB

[2788] Frederic F. Burchsted. Book review: Geologists and the History of Geology: An International Bibliography, Supplement 2, 1985–1993, and Additions by William A. S. Sarjeant. Isis, 88(3):584, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236228.

Anonymous:1997:Cc

[2789] Anonymous. Collections. Isis, 88(3):584–591, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236229.

Anonymous:1997:BMd

[2790] Anonymous. Back matter. Isis, 88(3):592–596, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236230.

Anonymous:1997:FMd

[2791] Anonymous. Front matter. Isis, 88(3):??, September 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236149.

Thompson:1997:DRL

[2792] Emily Thompson. Dead rooms and live wires: Harvard, Hollywood, and the deconstruction of architectural acoustics, 1900–1930. Isis, 88 (4):597–626, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237829.

Wright:1997:MOB

[2793] Margaret R. Wright. Marcella O’Grady Boveri (1863–1950): Her three careers in biology. Isis, 88(4):627–652, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237830.

Dehue:1997:DER

[2794] Trudy Dehue. Deception, efficiency, and random groups: Psychology and the gradual origination of the random group design. Isis, 88(4):653–673, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237831.

Pesic:1997:SSS

[2795] Peter Pesic. Secrets, symbols, and systems: Parallels between cryptanalysis and algebra, 1580–1700. Isis, 88(4):674–692, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237832.

Rosario:1997:BRB

[2796] Vernon A. Rosario. Book review: Impure Science: AIDS, Activism, and the Politics of Knowledge. Isis, 88(4):693–694, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237833.

Billington:1997:BRB

[2797] David P. Billington. Book review: Invention by Design: How Engineers Get from Thought to Thing by Henry Petroski. Isis, 88(4):694–695, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237834.

Geschwind:1997:BRB

[2798] Carl-Henry Geschwind. Book review: Thinking about the Earth: A History of Ideas in Geology by David R. Oldroyd. Isis, 88(4):695, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237835.

Borst:1997:BRB

[2799] Charlotte G. Borst. Book review: The Empty Cradle: Infertility in America from Colonial Times to the Present by Margaret Marsh; Wanda Ronner. Isis, 88(4):695–696, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237836.

Decker:1997:BRB

[2800] Hannah S. Decker. Book review: Approaching Hysteria: Disease and Its Interpretations by Mark S. Micale. Isis, 88(4):696–697, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237837.

Carson:1997:BRBa

[2801] John Carson. Book review: From Idiocy to Mental Deficiency: Historical Perspectives on People with Learning Disabilities by David Wright; Anne Digby. Isis, 88(4):697–698, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237838.

Tjossem:1997:BRB

[2802] Sara F. Tjossem. Book review: From Coastal Wilderness to Fruited Plain: A History of Environmental Change in Temperate North America, 1500–Present by Gordon G. Whitney. Isis, 88(4):698–699, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237839.

Romero:1997:BRB

[2803] Aldemaro Romero. Book review: Las plantas del mundo en la historia: Ilustraciones botanicas de cinco siglos. Isis, 88(4):699–700, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237840.

Kondratas:1997:BRB

[2804] Ramunas Kondratas. Book review: Spectacles and Other Vision Aids: A History and Guide to Collecting by J. William Rosenthal. Isis, 88 (4):700–701, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237841.

Hauger:1997:BRB

[2805] J. Scott Hauger. Book review: Molyneux’s Problem: Three Centuries of Discussion on the Perception of Forms by Marjolein Degenaar; Michael J. Collins. Isis, 88(4):701–702, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237842.

Crosby:1997:BRB

[2806] Alfred W. Crosby. Book review: The Foundations of Modern Science in the Middle Ages: Their Religious, Institutional, and Intellectual Contexts by Edward Grant. Isis, 88(4):702–703, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237843.

Pasnau:1997:BRB

[2807] Robert Pasnau. Book review: Cognitive Psychology in the Middle Ages by Simon Kemp. Isis, 88(4):703–704, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237844.

Savage-Smith:1997:BRB

[2808] Emilie Savage-Smith. Book review: Ibn al-Jazzar on Forgetfulness and Its Treatment: Critical Edition of the Arabic Text and the Hebrew Translations with Commentary and Translations into English by Gerrit Bos. Isis, 88(4):704, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237845.

Moreton:1997:BRBb

[2809] Jennifer Moreton. Book review: On the Six Days of Creation by Robert Grosseteste; C. F. J. Martin. Isis, 88(4):704–705, December 1997. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237846.

Smith:1997:BRBb

[2810] Pamela H. Smith. Book review: Darke Hierogliphicks: Alchemy in English Literature from Chaucer to the Restoration by Stanton J. Linden. Isis, 88(4):705–706, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237847.

Lattis:1997:BRB

[2811] James M. Lattis. Book review: La nave di ermete: La cosmologia di Girolamo Fracastoro by Enrico Peruzzi. Isis, 88(4):706–707, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237848.

Blair:1997:BRB

[2812] Ann Blair. Book review: La poesie du ciel en France dans la seconde moitie du seizieme siècle by Isabelle Pantin. Isis, 88(4):707–708, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237849.

MacCormack:1997:BRB

[2813] Sabine MacCormack. Book review: The Mapping of New Spain: Indigenous Cartography and the Maps of the Relaciones Geograficas by Barbara E. Mundy. Isis, 88(4):708–709, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237850.

Yoder:1997:BRB

[2814] Joella Yoder. Book review: Christiaan Huygens, 1629–1695; Een vernuftig geleerde: De technische vondsten van Christiaan Huygens by Rob H. van Gent; Anne C. van Helden; The Huygens Collection by Anne C. van Helden; Rob H. van Gent; Christiaan Huygens en de Wetenschapsrevolutie van de 17de eeuw by H. F. Cohen. Isis, 88(4):709–710, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237851.

Newcomb:1997:BRB

[2815] Sally Newcomb. Book review: Restless Genius: Robert Hooke and His Earthly Thoughts by Ellen Tan Drake. Isis, 88(4):710–711, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237852.

Kardel:1997:BRB

[2816] Troels Kardel. Book review: Chaos: Niels Stensen’s Chaos-Manuscript, Copenhagen, 1659 by Niels Stensen; August Ziggelaar. Isis, 88(4):711– 712, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237853.

Warner:1997:BRB

[2817] Deborah Jean Warner. Book review: Sympathetic Attractions: Magnetic Practices, Beliefs, and Symbolism in Eighteenth-Century England by Patricia Fara. Isis, 88(4):712, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237854.

Wilson:1997:BRB

[2818] Andrew D. Wilson. Book review: Auf dem Weg zu Kants Theorie der Zeit: Untersuchung zur Genese des Zeitbegriffs in der Philosophie Immanuel Kants by Detlev Soffler. Isis, 88(4):712–713, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237855.

Pyne:1997:BRB

[2819] Kathleen Pyne. Book review: Science and the Perception of Nature: British Landscape Art in the Late Eighteenth and Early Nineteenth Centuries by Charlotte Klonk. Isis, 88(4):713–714, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237856.

Porter:1997:BRB

[2820] Theodore M. Porter. Book review: Les maitres de l’erreur: Mesure et probabilite au XIXe siècle by Stephane Callens. Isis, 88(4):714–715, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237857.

Williams:1997:BRBb

[2821] L. Pearce Williams. Book review: Maxwell on the Electromagnetic Field: A Guided Study by Thomas K. Simpson. Isis, 88(4):715–716, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237858.

Morrell:1997:BRB

[2822] Jack Morrell. Book review: Edward Frankland: Chemistry, Controversy, and Conspiracy in Victorian England by Colin A. Russell. Isis, 88(4): 716–717, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237859.

Petroski:1997:BRB

[2823] Henry Petroski. Book review: Constructing a Bridge: An Exploration of Engineering Culture, Design, and Research in Nineteenth-Century France and America by Eda Kranakis. Isis, 88(4):717, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237860.

Casado:1997:BRB

[2824] Santos Casado. Book review: Evolucionismo y anarquismo en España, 1882–1914 by Alvaro Giron Sierra. Isis, 88(4):717–718, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237861.

Figueiroa:1997:BRB

[2825] Silvia Figueiroa. Book review: La industria del gas en Barcelona (1841– 1933): Innovacion tecnologica, territorio urbano y conflicto de intereses by Mercedes Arroyo Huguet. Isis, 88(4):718–719, December 1997.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237862.

Hawes:1997:BRB

[2826] Joseph M. Hawes. Book review: Getting the Message Through: A Branch History of the U.S. Army Signal Corps by Rebecca Robbins Raines. Isis, 88(4):719–720, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237863.

Hansen:1997:BRB

[2827] LeeAnn Hansen. Book review: Aufklarung, burgerliche Selbsterfahrung, und die “Erfindung” der Psychiatrie in Deutschland, 1770–1850 by Doris Kaufmann. Isis, 88(4):720–721, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237864.

Cooter:1997:BRBb

[2828] Roger Cooter. Book review: Small and Special: The Development of Hospitals for Children in Victorian Britain by Elizabeth M. R. Lomax. Isis, 88(4):721–722, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237865.

Carmichael:1997:BRBc

[2829] Ann G. Carmichael. Book review: Bubonic Plague in Nineteenth-Century China by Carol Benedict. Isis, 88(4):722–723, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237866.

Noelke:1997:BRB

[2830] Virginia Noelke. Book review: Bandelier: The Life and Adventures of Adolph Bandelier by Charles H. Lange; Carroll L. Riley. Isis, 88(4): 723–724, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237867.

Trimble:1997:BRB

[2831] Virginia Trimble. Book review: Cosmology and Controversy: The Historical Development of Two Theories of the Universe by Helge Kragh. Isis, 88(4):724–725, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237868.

Sime:1997:BRB

[2832] Ruth Lewin Sime. Book review: Atomic Histories by Rudolf E. Peierls. Isis, 88(4):725–726, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237869.

Cassidy:1997:BRBb

[2833] David C. Cassidy. Book review: Heisenberg Probably Slept Here: The Lives, Times, and Ideas of the Great Physicists of the Twentieth Century by Richard P. Brennan. Isis, 88(4):726, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237870.

Carson:1997:BRBb

[2834] Cathryn Carson. Book review: Wolfgang Pauli: Wissenschaftlicher Briefwechsel mit Bohr, Einstein, Heisenberg, u.a. Volume 4, Part 1: 1950–1952. [Wolfgang Pauli: Scientific Correspondence with Bohr, Einstein, Heisenberg, et al. Volume 4, Part 1: 1950–1952.] by Wolfgang Pauli; Karl von Meyenn. Isis, 88(4):726–727, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237871.

Hall:1997:BRB

[2835] Karl Hall. Book review: Yakov Ilich Frenkel: His Work, Life, and Letters by Victor Ya. Frenkel; Alexander S. Silbergleit. Isis, 88(4):727–728, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237872.

Kim:1997:BRB

[2836] Mi Gyung Kim. Book review: Arrhenius: From Ionic Theory to the Greenhouse Effect by Elisabeth Crawford. Isis, 88(4):728–729, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237873.

Fruton:1997:BRB

[2837] Joseph S. Fruton. Book review: A Lifetime of Synergy with Theory and Experiment by Andrew Streitwieser. Isis, 88(4):729–730, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237874.

Kohler:1997:BRB

[2838] Robert E. Kohler. Book review: The Eighth Day of Creation: Makers of the Revolution in Biology by Horace Freeland Judson. Isis, 88(4):

730–731, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237875.

Blaisdell:1997:BRB

[2839] Muriel L. Blaisdell. Book review: A History of the Ecosystem Concept in Ecology: More than the Sum of the Parts by Frank Benjamin Golley. Isis, 88(4):731, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237876.

Larson:1997:BRB

[2840] Edward J. Larson. Book review: The Social Meaning of Modern Biology: From Social Darwinism to Sociobiology by Howard L. Kaye. Isis, 88(4): 731–732, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237877.

Mills:1997:BRBb

[2841] Eric L. Mills. Book review: Roger: A Biography of Roger Revelle by Judith Morgan; Neil Morgan. Isis, 88(4):732–733, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237878.

Marchand:1997:BRB

[2842] Suzanne Marchand. Book review: Soziologie im Dritten Reich by Carsten Klingemann. Isis, 88(4):733–734, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237879.

Ash:1997:BRB

[2843] Mitchell G. Ash. Book review: Studien zur autoritaren Personlichkeit: Ausgewahlte Schriften by Else Frenkel-Brunswik; Dietmar Paier; Bertram F. Malle. Isis, 88(4):734–735, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237880.

Smith:1997:BRBc

[2844] Mark C. Smith. Book review: Marginality and Dissent in Twentieth– Century American Sociology: The Case of Elizabeth Briant Lee and Alfred McClung Lee by John F. Galliher; James M. Galliher. Isis, 88(4): 735–736, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237881.

Lang:1997:BRB

[2845] Harry G. Lang. Book review: Forbidden Signs: American Culture and the Campaign against Sign Language by Douglas C. Baynton. Isis, 88 (4):736, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237882.

Bullough:1997:BRB

[2846] Vern L. Bullough. Book review: Reinventing the Sexes: The Biomedical Construction of Femininity and Masculinity by Marianne van den Wijngaard. Isis, 88(4):736–737, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237883.

Matthews:1997:BRBb

[2847] J. Rosser Matthews. Book review: Rationalizing Medical Work: Decision-Support Techniques and Medical Practices by Marc Berg. Isis, 88(4):737–738, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237884.

Cartwright:1997:BRB

[2848] Lisa Cartwright. Book review: Naked to the Bone: Medical Imaging in the Twentieth Century by Bettyann Holtzmann Kevles. Isis, 88(4): 738–739, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237885.

Kevles:1997:BRB

[2849] Bettyann Holtzmann Kevles. Book review: Origins and Development of Medical Imaging by T. Doby; G. Alker. Isis, 88(4):739, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237886.

Zelizer:1997:BRB

[2850] Viviana A. Zelizer. Book review: The worth of a Child by Thomas H. Murray. Isis, 88(4):739–740, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237887.

Hacker:1997:BRB

[2851] Barton Hacker. Book review: National Military Establishments and the Advancement of Science and Technology: Studies in Twentieth–Century History by Paul Forman; Jose M. Sánchez-Ron. Isis, 88(4):740–741, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237888.

Walker:1997:BRBc

[2852] Mark Walker. Book review: The Peenemunde Wind Tunnels: A Memoir by Peter P. Wegener. Isis, 88(4):741–742, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237889.

Brooks:1997:BRBb

[2853] Nathan M. Brooks. Book review: Stalin’s Captive: Nikolaus Riehl and the Soviet Race for the Bomb by Nikolaus Riehl; Frederick Seitz. Isis, 88 (4):742–743, December 1997. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237890.

Neufeld:1997:BRB

[2854] Michael J. Neufeld. Book review: Aiming at Targets: The Autobiography of Robert C. Seamans, Jr. by Robert C. Seamans. Isis, 88(4):743–744, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237891.

Seidel:1997:BRB

[2855] Robert W. Seidel. Book review: Nuclear Rites: A Weapons Laboratory at the End of the Cold War by Hugh Gusterson. Isis, 88(4):744–745, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237892.

Norberg:1997:BRB

[2856] Arthur L. Norberg. Book review: The Supermen: The Story of Seymour Cray and the Technical Wizards behind the Supercomputer by Charles J. Murray. Isis, 88(4):745–746, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237893.

Nickles:1997:BRBb

[2857] Thomas Nickles. Book review: Beauty and Revolution in Science by James W. McAllister. Isis, 88(4):746–747, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237894.

Jardine:1997:BRB

[2858] Nick Jardine. Book review: The Elusive Synthesis: Aesthetics and Science by Alfred I. Tauber. Isis, 88(4):747–748, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237895.

Rifkin:1997:BRB

[2859] Stan Rifkin. Book review: Visual Explanations: Images and Quantities, Evidence and Narrative by Edward R. Tufte. Isis, 88(4):748–749, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237896.

Gilbert:1997:BRB

[2860] James Gilbert. Book review: Religion and Technology: A Study in the Philosophy of Culture by Jay Newman. Isis, 88(4):749–750, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237897.

Anonymous:1997:Cd

[2861] Anonymous. Collections. Isis, 88(4):750–756, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237898.

Anonymous:1997:BMe

[2862] Anonymous. Back matter. Isis, 88(4):757–760, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237899.

Anonymous:1997:VI

[2863] Anonymous. Volume information. Isis, 88(4):761–775, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237827.

Anonymous:1997:FMe

[2864] Anonymous. Front matter. Isis, 88(4):??, December 1997. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237828.

Neu:1998:CBH

[2865] John Neu. Current bibliography of the history of science and its cultural influences, 1998. Isis, 89(??):1–308, ???? 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237469.

Anonymous:1998:BMa

[2866] Anonymous. Back matter. Isis, 89(??):??, ???? 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237470.

Anonymous:1998:FMa

[2867] Anonymous. Front matter. Isis, 89(??):i–xvi, ???? 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237468.

Leonard:1998:EEM

[2868] Robert J. Leonard. Ethics and the excluded middle: Karl Menger and social science in interwar Vienna. Isis, 89(1):1–26, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236652.

Geschwind:1998:ESR

[2869] Carl-Henry Geschwind. Embracing science and research: Early Twentieth–Century Jesuits and seismology in the United States. Isis, 89(1):27–49, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236653.

Moreton:1998:DAC

[2870] Jennifer Moreton. Doubts about the calendar: Bede and the eclipse of 664. Isis, 89(1):50–65, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236654.

Debus:1998:CPC

[2871] Allen G. Debus. Chemists, physicians, and changing perspectives on the scientific revolution. Isis, 89(1):66–81, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236655.

Beltran:1998:WWE

[2872] Antonio Beltran. Wine, water, and epistemological sobriety: A note on the Koyre–MacLachlan Debate. Isis, 89(1):82–89, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236656.

MacLachlan:1998:EHS

[2873] James MacLachlan. Experimenting in the history of science. Isis, 89(1): 90–92, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236657.

Edwards:1998:BRBa

[2874] Paul N. Edwards. Book review: Computer: A History of the Information Machine. Isis, 89(1):93–99, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236658.

Montgomery:1998:BRBa

[2875] William Montgomery. Book review: Charles Darwin: Voyaging. Isis, 89(1):100–102, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236659.

Kuklick:1998:BRB

[2876] Henrika Kuklick. Book review: The Expansive Moment: The Rise of Social Anthropology in Britain and Africa, 1918–1970. Isis, 89(1):103– 111, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236660.

Nasr:1998:BRB

[2877] Seyyed Hossein Nasr. Book review: Fondateurs et commentateurs: Banu Musa, Ibn Qurra, Ibn Sinan, al-Khazin, al-Quhi, Ibn al-Samh, Ibn Hud. Isis, 89(1):112–113, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236661.

Wang:1998:BRB

[2878] Jessica Wang. Book review: James B. Conant: Harvard to Hiroshima and the Making of the Nuclear Age by James G. Hershberg. Isis, 89(1): 113–115, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236662.

Brock:1998:BRB

[2879] William H. Brock. Book review: A History of Chemistry by Bernadette Bensaude-Vincent; Isabelle Stengers; Deborah van Dam. Isis, 89(1): 116, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236663.

Goodrum:1998:BRB

[2880] Matthew R. Goodrum. Book review: History of Physical Anthropology: An Encyclopedia by Frank Spencer. Isis, 89(1):116–117, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236664.

Kidwell:1998:BRB

[2881] Peggy Aldrich Kidwell. Book review: Science in the Field by Henrika Kuklick; Robert E. Kohler. Isis, 89(1):117–118, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236665.

Stevens:1998:BRBa

[2882] Rosemary A. Stevens. Book review: The Limits of Medicine: How Science Shapes Our Hope for the Cure by Edward S. Golub. Isis, 89(1): 118–119, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236666.

Richards:1998:BRBa

[2883] Graham Richards. Book review: Inventing Ourselves: Psychology, Power, and Personhood by Nikolas Rose. Isis, 89(1):119–120, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236667.

Daemmrich:1998:BRB

[2884] Arthur Daemmrich. Book review: Geschichte der Arzneimitteltherapie by Wolf-Dieter Muller-Jahncke; Christoph Friedrich. Isis, 89(1):120– 121, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236668.

Davis:1998:BRBa

[2885] Audrey B. Davis. Book review: The Bernard Becker Collection in Ophthalmology: An Annotated Catalog by Lilla Wechsler; Christopher Hoolihan; Mark F. Weimer. Isis, 89(1):121, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236669.

Lloyd:1998:BRB

[2886] Geoffrey Lloyd. Book review: Ancient Natural History: Histories of Nature by Roger French. Isis, 89(1):121–122, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236670.

OBrien:1998:BRB

[2887] Denis O’Brien. Book review: Ancient Philosophy, Mystery, and Magic: Empedocles and Pythagorean Tradition by Peter Kingsley. Isis, 89(1): 122–124, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236671.

Durling:1998:BRB

[2888] Richard J. Durling. Book review: Le corps respirant: La pensée physiologique chez Galien by Armelle Debru. Isis, 89(1):124, March 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236672.

Sweet:1998:BRB

[2889] Victoria Sweet. Book review: Secrets of God: Writings of Hildegard of Bingen by Hildegard of Bingen; Sabina Flanagan. Isis, 89(1):124– 125, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236673.

Davenport:1998:BRB

[2890] Anne Davenport. Book review: Storia della filosofia tedesca nel medioevo–Il secolo XIII by Loris Sturlese. Isis, 89(1):125–126, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236674.

Crosby:1998:BRB

[2891] Alfred W. Crosby. Book review: The Black Death and the Transformation of the West by David Herlihy; Samuel K. Cohn. Isis, 89(1):126, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236675.

Sylla:1998:BRB

[2892] Edith Dudley Sylla. Book review: Quaestiones super de generatione et corruptione by Nicole Oresme; Stefano Caroti. Isis, 89(1):126–127, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236676.

Nutton:1998:BRB

[2893] Vivian Nutton. Book review: L’esperienza del passato: Alessandro Benedetti, filologo e medico umanista by Giovanna Ferrari. Isis, 89(1): 127–128, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236677.

Gunnoe:1998:BRB

[2894] Charles D. Gunnoe, Jr. Book review: Paracelsus: Das Werk–Die Rezeption by Volker Zimmermann; Paracelsus: Speculative Theory and the Crisis of the Early Reformation by Andrew Weeks. Isis, 89(1):128–130, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236678.

Harrie:1998:BRB

[2895] Jeanne Harrie. Book review: Oeuvres completes. Volume 2: Le souper des cendres by Giordano Bruno; Giovanni Aquilecchia; Yves Hersant; Oeuvres completes. Volume 3: De la cause, du principe et de l’un by Giordano Bruno; Giovanni Aquilecchia; Luc Hersant; Oeuvres completes. Volume 4: De l’infini, de l’univers et des mondes by Giordano Bruno; Giovanni Aquilecchia; Jean–Pierre Cavaille. Isis, 89(1):130–131, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236679.

Alexander:1998:BRB

[2896] Amir Alexander. Book review: From Indivisibles to Infinitesimals: Studies on Seventeenth-Century Mathematizations of Infinitely Small Quantities by Antoni Malet. Isis, 89(1):131–132, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236680.

Harkness:1998:BRB

[2897] Deborah Harkness. Book review: Nicolas Flamel, His Exposition of the Hieroglyphicall Figures (1624) by Nicolas Flamel; Laurinda Dixon. Isis, 89(1):132–133, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236681.

Giglioni:1998:BRBa

[2898] Guido Giglioni. Book review: A Free Enquiry into the Vulgarly Received Notion of Nature by Robert Boyle; Edward B. Davis; Michael Hunter. Isis, 89(1):133–134, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236682.

Davis:1998:BRBb

[2899] Edward B. Davis. Book review: Fallen Languages: Crises of Representation in Newtonian England, 1660–1740 by Robert Markley. Isis, 89(1):134–135, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236683.

Yoder:1998:BRB

[2900] Joella Yoder. Book review: La mécanique de Christian Huygens: La relativité du mouvement au XVIIe siècle by Christiane Vilain. Isis, 89 (1):135, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236684.

Reeds:1998:BRB

[2901] Karen Reeds. Book review: AΠOP HMA BOT ANIKON. De signaturis plantarum: Ristampa anastatica dell’edizione Norimbergae 1653 by Wolfgang Ambrosius Fabricius; Massimo Luigi Bianchi. Isis, 89(1):135– 136, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236685.

Marcum:1998:BRB

[2902] James A. Marcum. Book review: Self-Generation: Biology, Philosophy, and Literature around 1800 by Helmut Mueller-Sievers. Isis, 89(1):136– 137, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236686.

Fraser:1998:BRBa

[2903] Craig Fraser. Book review: Commentationes astronomicae: Mechanicae et astronomicae ad physicam cosmicam pertinentes by Leonhard Euler; Eric J. Aiton. Isis, 89(1):137–138, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236687.

OConnor:1998:BRB

[2904] Erin O’Connor. Book review: Imagining Monsters: Miscreations of the Self in Eighteenth-Century England by Dennis Todd. Isis, 89(1):138– 139, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236688.

Davies:1998:BRB

[2905] Gordon L. Herries Davies. Book review: A Calendar of the Correspondence of John Strange F.R.S. (1732–1799) by Luca Ciancio. Isis, 89(1): 139–140, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236689.

Galuzzi:1998:BRB

[2906] Massimo Galuzzi. Book review: Analytical Mechanics: Translated from the “Mecanique analytique,” nouvelle edition of 1811 by J. L. Lagrange; Auguste Boissonnade; Victor N. Vagliente. Isis, 89(1):140–141, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236690.

Schabas:1998:BRBa

[2907] Margaret Schabas. Book review: Early Mathematical Economics: William Whewell and the British Case by James P. Henderson. Isis, 89(1):141–142, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236691.

Schwartz:1998:BRB

[2908] Joel S. Schwartz. Book review: Darwin. [CD-ROM] by Charles Darwin; Michael T. Ghiselin. Isis, 89(1):142–143, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236692.

Dietrich:1998:BRB

[2909] Michael R. Dietrich. Book review: Gregor Mendel: The First Geneticist by Vitezslav Orel; Stephen Finn. Isis, 89(1):143–144, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236693.

Strout:1998:BRB

[2910] Cushing Strout. Book review: Science and Religion in the Era of William James. Volume 1: Eclipse of Certainty, 1820–1880 by Paul Jerome Croce. Isis, 89(1):144–145, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236694.

McLarty:1998:BRB

[2911] Colin McLarty. Book review: La notion de nombre chez Dedekind, Cantor, Frege: Theories, conceptions, et philosophie by Jean–Pierre Belna. Isis, 89(1):145–146, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236695.

Chiarappa:1998:BRB

[2912] Michael J. Chiarappa. Book review: The Artificial River: The Erie Canal and the Paradox of Progress, 1817–1862 by Carol Sheriff. Isis, 89(1):146–147, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236696.

Barrow-Green:1998:BRB

[2913] June Barrow-Green. Book review: Celestial Encounters: The Origins of Chaos and Stability by Florin Diacu; Philip Holmes. Isis, 89(1):147– 148, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236697.

Allen:1998:BRB

[2914] Ann Taylor Allen. Book review: Sauglingsfursorge zwischen sozialer Hygiene und Eugenik: Das Beispiel Berlins im Kaiserreich und in der Weimarer Republik by Sigrid Stockel. Isis, 89(1):148–149, March 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236698.

Canizares-Esguerra:1998:BRB

[2915] Jorge Canizares-Esguerra. Book review: Los primeros pasos de la ecologia en España by Santos Casado de Otaola. Isis, 89(1):149, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236699.

Junker:1998:BRB

[2916] Thomas Junker. Book review: Monad to Man: The Concept of Progress in Evolutionary Biology by Michael Ruse. Isis, 89(1):149–150, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236700.

Ghiselin:1998:BRB

[2917] Michael T. Ghiselin. Book review: This is Biology: The Science of the Living World by Ernst Mayr. Isis, 89(1):150–151, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236701.

Kragh:1998:BRB

[2918] Helge Kragh. Book review: Conceptual Developments of Twentieth– Century Field Theories by Tian Yu Cao. Isis, 89(1):151–152, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236702.

Cushing:1998:BRB

[2919] James T. Cushing. Book review: The Dawning of Gauge Theory by Lochlainn O’Raifeartaigh. Isis, 89(1):152–153, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236703.

Carson:1998:BRB

[2920] Cathryn Carson. Book review: Literatur und Quantentheorie: Die Rezeption der modernen Physik in Schriften zur Literatur und Philosophie deutschsprachiger Autoren (1925–1970) by Elisabeth Emter. Isis, 89(1):153–154, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236704.

Krige:1998:BRBa

[2921] John Krige. Book review: Launching Europe: An Ethnography of European Cooperation in Space Science by Stacia E. Zabusky. Isis, 89(1):

154, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236705.

Eghigian:1998:BRB

[2922] Greg Eghigian. Book review: Reenchanted Science: Holism in German Culture from Wilhelm II to Hitler by Anne Harrington. Isis, 89(1):154– 155, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236706.

Walker:1998:BRBa

[2923] Mark Walker. Book review: “Forschungsstelle D”: Der Schweizer Ingenieur Walter Dallenbach (1892–1990), die AEG, und die Entwicklung kernphysikalischer Grossgerate im nationalsozialistischen Deutschland by Burghard Weiss; Grossforschung in Berlin: Geschichte des Hahn-Meitner-Instituts by Burghard Weiss. Isis, 89(1):155–156, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236707.

Harwit:1998:BRB

[2924] Martin Harwit. Book review: Prompt and Utter Destruction: Truman and the Use of Atomic Bombs against Japan by J. Samuel Walker. Isis, 89(1):156–157, March 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236708.

Whittemore:1998:BRB

[2925] Gilbert Whittemore. Book review: Effects of Atomic Radiation: A Half Century of Studies from Hiroshima and Nagasaki by William J. Schull. Isis, 89(1):157–158, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236709.

Gough:1998:BRBa

[2926] Jerry B. Gough. Book review: Science, Vine, and Wine in Modern France by Harry W. Paul. Isis, 89(1):158–159, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236710.

McClellan:1998:BRBa

[2927] Michelle L. McClellan. Book review: Substance and Shadow: Women and Addiction in the United States by Stephen R. Kandall; Jennifer Petrillo. Isis, 89(1):159–160, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236711.

Westkott:1998:BRB

[2928] Marcia Westkott. Book review: Karen Horney: A Psychoanalyst’s Search for Self-Understanding by Bernard J. Paris. Isis, 89(1):160–161, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236712.

VanRiper:1998:BRB

[2929] A. Bowdoin Van Riper. Book review: Digging for Dollars: American Archaeology and the New Deal by Paul Fagette. Isis, 89(1):161–162, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236713.

Cittadino:1998:BRB

[2930] Eugene Cittadino. Book review: Ecologists and Environmental Politics: A History of Contemporary Ecology by Stephen Bocking. Isis, 89(1):162– 163, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236714.

Agar:1998:BRB

[2931] Jon Agar. Book review: Yerkes Observatory, 1892–1950: The Birth, Near Death, and Resurrection of a Scientific Research Institution by Donald E. Osterbrock. Isis, 89(1):163–164, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236715.

Stefanik:1998:BRB

[2932] Robert P. Stefanik. Book review: Planet Quest: The Epic Discovery of Alien Solar Systems by Ken Croswell. Isis, 89(1):164, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236716.

Glen:1998:BRB

[2933] William Glen. Book review: T. rex and the Crater of Doom by Walter Alvarez. Isis, 89(1):164–166, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236717.

Jeffrey:1998:BRB

[2934] Kirk Jeffrey. Book review: American Cardiology: The History of a Specialty and Its College by W. Bruce Fye. Isis, 89(1):166–167, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236718.

Marks:1998:BRB

[2935] Harry M. Marks. Book review: Wellcome Witnesses to Twentieth– Century Medicine. Volume 1 by E. M. Tansey; P. P. Catterall; D. A. Christie; S. V. Willhoft; L. A. Reynolds. Isis, 89(1):167–168, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236719.

Horrocks:1998:BRB

[2936] Sally M. Horrocks. Book review: Vitamania: Vitamins in American Culture by Rima D. Apple. Isis, 89(1):168–169, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236720.

DeVito:1998:BRB

[2937] Scott DeVito. Book review: Brute Science: Dilemmas of Animal Experimentation by Hugh LaFollette; Niall Shanks. Isis, 89(1):169–170, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236721.

Pang:1998:BRB

[2938] Alex Soojung-Kim Pang. Book review: BuckyWorks: Buckminster Fuller’s Ideas for Today by Jay Baldwin. Isis, 89(1):170–171, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236722.

Munday:1998:BRBa

[2939] Pat Munday. Book review: Mining Cultures: Men, Women, and Leisure in Butte, 1914–41 by Mary Murphy. Isis, 89(1):171–172, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236723.

Restivo:1998:BRB

[2940] Sal Restivo. Book review: On Social Structure and Science by Robert K. Merton; Piotr Sztompka. Isis, 89(1):172–173, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236724.

Hoyningen-Huene:1998:BRB

[2941] Paul Hoyningen-Huene. Book review: Changing Tools: Case Studies in the History of Scientific Methodology by Marta Feher. Isis, 89(1): 173, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236725.

Cohen:1998:BRBa

[2942] I. Bernard Cohen. Book review: The Scientific Revolution: An Annotated Bibliography by S. A. Jayawardene. Isis, 89(1):174, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236726.

Zanish-Belcher:1998:BRB

[2943] Tanya Zanish-Belcher. Book review: Notable Women in the Physical Sciences: A Biographical Dictionary by Benjamin F. Shearer; Barbara S. Shearer. Isis, 89(1):174–175, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236727.

Brown:1998:BRB

[2944] Laurie M. Brown. Book review: The Quotable Einstein by Alice Calaprice. Isis, 89(1):175, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236728.

Anonymous:1998:Ca

[2945] Anonymous. Collections. Isis, 89(1):175–180, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236729.

Cohen:1998:ERH

[2946] H. Floris Cohen. Eloge: Reijer Hooykaas, 1 August 1906–4 January 1994. Isis, 89(1):181–184, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236730.

Schabas:1998:AMH

[2947] Margaret Schabas, Keith R. Benson, and Margaret J. Osler. Annual meeting of the History of Science Society San Diego, 5–9 November 1997. Isis, 89(1):185–190, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236731.

Anonymous:1998:BMb

[2948] Anonymous. Back matter. Isis, 89(1):191–202, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236732.

Anonymous:1998:FMb

[2949] Anonymous. Front matter. Isis, 89(1):??, March 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236651.

Riskin:1998:RIR

[2950] Jessica Riskin. Rival idioms for a revolutionized science and a republican citizenry. Isis, 89(2):203–232, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237754.

Topham:1998:BCC

[2951] Jonathan R. Topham. Beyond the “common context”: The production and reading of the Bridgewater Treatises. Isis, 89(2):233–262, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237755.

Staley:1998:HRP

[2952] Richard Staley. On the histories of Relativity: The propagation and elaboration of Relativity Theory in participant histories in Germany, 1905– 1911. Isis, 89(2):263–299, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237756.

Mertens:1998:SSV

[2953] Joost Mertens. Shocks and sparks: The Voltaic pile as a demonstration device. Isis, 89(2):300–311, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237757.

Molella:1998:ESG

[2954] Arthur Molella. Eloge: Stanley Goldberg, 4 August 1934–16 October 1996. Isis, 89(2):312–315, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237758.

Hoffleit:1998:EHM

[2955] Dorrit Hoffleit. Eloge: Helen Meriwether Lewis Thomas, 21 August 1905– 6 August 1997. Isis, 89(2):316–317, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237759.

Multhauf:1998:ERW

[2956] Robert P. Multhauf. Eloge: Roderick S. Webster, 14 September 1915– 31 July 1997. Isis, 89(2):317–319, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237760.

Kassler:1998:LE

[2957] Jamie C. Kassler, Deborah E. Harkness, and Richard Arthur. Letters to the Editor. Isis, 89(2):320, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237761.

Davis:1998:BRBc

[2958] Audrey B. Davis. Book review: The Scientific 100: A Ranking of the Most Influential Scientists, Past and Present by John Simmons. Isis, 89 (2):321, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237762.

Hine:1998:BRB

[2959] William L. Hine. Book review: From Myth to Modern Mind: A Study of the Origins and Growth of Scientific Thought by Richard H. Schlagel. Isis, 89(2):321–322, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237763.

Needham:1998:BRB

[2960] Paul Needham. Book review: L’aube du savoir: Épitomé du système du monde by Pierre Duhem; Anastasios Brenner. Isis, 89(2):322–323, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237764.

Inkster:1998:BRB

[2961] Ian Inkster. Book review: Paths of Fire: An Anthropologist’s Inquiry into Western Technology by Robert McC. Adams. Isis, 89(2):323–324, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237765.

Black:1998:BRB

[2962] Jeremy Black. Book review: The Work in the World: Geographical Practice and the Written Word by Michael R. Curry. Isis, 89(2):324– 325, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237766.

Good:1998:BRB

[2963] Gregory Good. Book review: Due secoli di strumenti geomagnetici in Italia (1740–1971) by M. Basso Ricci; L. Cafarella; A. Meloni; P. Tucci. Isis, 89(2):325–326, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237767.

Henson:1998:BRB

[2964] Pamela M. Henson. Book review: Picturing Knowledge: Historical and Philosophical Problems concerning the Use of Art in Science by Brian S. Baigrie. Isis, 89(2):326, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237768.

Harris:1998:BRB

[2965] Benjamin Harris. Book review: A Pictoral History of Psychology by Wolfgang G. Bringmann; Helmut E. Luck; Rudolf Miller; Charles E. Early. Isis, 89(2):327, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237769.

Rydell:1998:BRB

[2966] Robert W. Rydell. Book review: Exhibiting Electricity by K. G. Beauchamp. Isis, 89(2):327–328, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237770.

Fleming:1998:BRB

[2967] James R. Fleming. Book review: Perils of a Restless Planet: Scientific Perspectives on Natural Disasters by Ernest Zebrowski, Jr. Isis, 89(2): 328–329, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237771.

North:1998:BRB

[2968] J. D. North. Book review: Skywatchers, Shamans, and Kings: Astronomy and the Archaeology of Power by E. C. Krupp. Isis, 89(2):329–331, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237772.

French:1998:BRB

[2969] Roger French. Book review: Adversaries and Authorities: Investigations into Ancient Greek and Chinese Science by G. E. R. Lloyd. Isis, 89(2): 331, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237773.

Staszak:1998:BRB

[2970] Jean-François Staszak. Book review: Hippocrate: Airs, eaux, lieux. Volume 2, Part 2 by Hippocrates; Jacques Jouanna. Isis, 89(2):331–332, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237774.

Waley-Cohen:1998:BRB

[2971] Joanna Waley-Cohen. Book review: Technology and Gender: Fabrics of Power in Late Imperial China by Francesca Bray. Isis, 89(2):332– 333, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237775.

Peterson:1998:BRB

[2972] Charles A. Peterson. Book review: Science and Civilisation in China. Volume 6: Biology and Biological Technology. Part 3: Agro-Industries: Sugarcane Technology. Agro-Industries and Forestry by Joseph Needham; Christian Daniels; Nicholas K. Menzies. Isis, 89(2):333–334, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237776.

Burnett:1998:BRB

[2973] Charles Burnett. Book review: Die alteste lateinische Schrist über das indische Rechnen nach al-Hwarizmi by Menso Folkerts; Paul Kunitzsch. Isis, 89(2):334–335, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237777.

Merrilees:1998:BRB

[2974] Brian Merrilees. Book review: Anglo-Norman Medicine. Volume 2: Shorter Treatises by Tony Hunt. Isis, 89(2):335–336, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237778.

Frijhoff:1998:BRB

[2975] Willem Frijhoff. Book review: De geschiedenis van de kindergeneeskunde in Nederland. Volume 1: De periode tot 1700 by M. J. Van Lieburg. Isis, 89(2):336–337, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237779.

Shapiro:1998:BRB

[2976] Alan E. Shapiro. Book review: Painting the Heavens: Art and Science in the Age of Galileo by Eileen Reeves. Isis, 89(2):337–338, June 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237780.

Jesseph:1998:BRB

[2977] Douglas M. Jesseph. Book review: Mechanismus und Subjektivitat in der Philosophie von Thomas Hobbes by Michael Esfeld. Isis, 89(2):338– 339, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237781.

Findlen:1998:BRB

[2978] Paula Findlen. Book review: “Und sie furchtet sich vor niemandem”: Die Physikerin Laura Bassi (1711–1778) by Beate Ceranski. Isis, 89(2): 339–340, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237782.

Belcastro:1998:BRB

[2979] Alessandro Belcastro. Book review: Die Werke von Daniel Bernoulli. Volume 1: Medicin und Physiologie, Mathematische Jungenschriften, Positionastronomie by Daniel Bernoulli. Isis, 89(2):340–341, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237783.

Pancaldi:1998:BRB

[2980] Giuliano Pancaldi. Book review: Marsilio Landriani: Ricerche fisiche intorno alla salubrita dell’aria by Marsilio Landriani. Isis, 89(2):341– 342, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237784.

Greenbaum:1998:BRB

[2981] Louis S. Greenbaum. Book review: Memoirs on Paris Hospitals by Jacques Tenon; Dora B. Weiner. Isis, 89(2):342–343, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237785.

Cueto:1998:BRB

[2982] Marcos Cueto. Book review: Learning to Heal: The Medical Profession in Colonial Mexico, 1767–1831 by Luz Maria Hernandez Saenz. Isis, 89 (2):343, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237786.

Schlote:1998:BRB

[2983] Karl-Heinz Schlote. Book review: Miscellanea Mathematica 10 by Bernard Bolzano; Bob van Rootselaar; Ana van der Lugt. Isis, 89(2):

343–344, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237787.

Fraser:1998:BRBb

[2984] Craig Fraser. Book review: Vorlesungen über analytische Mechanik, Berlin 1847/48 by Carl Gustav J. Jacobi; Helmut Pulte. Isis, 89(2): 344–345, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237788.

Mawhin:1998:BRB

[2985] Jean Mawhin. Book review: Poincaré and the Three-Body Problem by June Barrow-Green. Isis, 89(2):345–346, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237789.

Schneider:1998:BRB

[2986] William H. Schneider. Book review: “Almost a Man of Genius”: Clemene Royer, Feminism, and Nineteenth-Century Science by Joy Harvey. Isis, 89(2):346–347, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237790.

Weikart:1998:BRB

[2987] Richard Weikart. Book review: Charles Darwins Briefwechsel mit deutschen Naturforschern: Ein Kalendarium mit Inhaltsangaben, biographischem Register und Bibliographie/Charles Darwin’s Correspondence with German Naturalists: A Calendar with Summaries, Biographical Register, and Bibliography by Thomas Junker; Marsha Richmond. Isis, 89(2):347, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237791.

Ludecke:1998:BRB

[2988] Cornelia Ludecke. Book review: Friedrich Ratzel (1844–1904): Naturwissenschaftler, Geograph, Gelehrter by Gerhard H. Muller. Isis, 89(2): 347–348, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237792.

Benson:1998:BRB

[2989] Maxine Benson. Book review: The Natural History of the Long Expedition to the Rocky Mountains, 1819–1820 by Howard Ensign Evans. Isis, 89(2):348–350, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237793.

Rainger:1998:BRB

[2990] Ronald Rainger. Book review: Discovering Dinosaurs in the Old West: The Field Journals of Arthur Lakes by Michael F. Kohl; John S. McIntosh. Isis, 89(2):350, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237794.

Walls:1998:BRB

[2991] Laura Dassow Walls. Book review: Journal. Volume 5: 1852–1853 by Henry D. Thoreau; Patrick F. O’Connell. Isis, 89(2):350–351, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237795.

Withuhn:1998:BRB

[2992] William L. Withuhn. Book review: Early American Railroads: Franz Anton Ritter von Gerstner’s Die innern Communicationen (1842–1843) by Franz Anton Ritter von Gerstner; Frederick C. Gamst; David J. Diephouse; John C. Decker. Isis, 89(2):351–352, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237796.

Graebner:1998:BRB

[2993] William Graebner. Book review: Safety First: Technology, Labor, and Business in the Building of American Work Safety, 1870–1939 by Mark Aldrich. Isis, 89(2):352–353, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237797.

Mazumdar:1998:BRBa

[2994] Pauline M. H. Mazumdar. Book review: Das Diphtherieserum: Ein neues Therapieprinzip, seine Entwicklung und Markteinfuhrung by Carola Throm. Isis, 89(2):353–354, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237798.

Stamhuis:1998:BRB

[2995] Ida H. Stamhuis. Book review: Memories: My Life as an International Leader in Health, Suffrage, and Peace by Aletta Jacobs; Harriet Feinberg; Annie Wright. Isis, 89(2):354–355, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237799.

Ender:1998:BRB

[2996] Evelyne Ender. Book review: Nerves and Narratives: A Cultural History of Hysteria in Nineteenth-Century British Prose by Peter Melville Logan. Isis, 89(2):355, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237800.

Mancini:1998:BRB

[2997] J. M. Mancini. Book review: The First Moderns: Profiles in the Origins of Twentieth–Century Thought by William R. Everdell. Isis, 89(2):355– 356, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237801.

Gomez-Torrente:1998:BRB

[2998] Mario Gomez-Torrente. Book review: Logical Dilemmas: The Life and Work of Kurt Godel by John W. Dawson, Jr. Isis, 89(2):356–357, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237802.

Porter:1998:BRB

[2999] Theodore M. Porter. Book review: Error and the Growth of Experimental Knowledge by Deborah G. Mayo. Isis, 89(2):357–358, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237803.

Chase:1998:BRB

[3000] Gene B. Chase. Book review: Indiscrete Thoughts by Gian-Carlo Rota; Fabrizio Palombi. Isis, 89(2):358–359, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237804.

Goodstein:1998:BRB

[3001] Judith R. Goodstein. Book review: “Il Nuovo Cimento”: Novant’anni di fisica in Italia, 1855–1944 by Giuseppe Giuliani. Isis, 89(2):359–360, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237805.

MacKinnon:1998:BRB

[3002] Edward MacKinnon. Book review: Mecanique quantique: Une introduction philosophique by Michel Bitbol. Isis, 89(2):360–361, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237806.

Deichmann:1998:BRB

[3003] Ute Deichmann. Book review: Weltburger im Reich der Chemie: Prof. Dr. Walter Fuchs (1891–1957): Ein deutsches Gelehrtenschicksal im 20. Jahrhundert by Valentin Wehefritz. Isis, 89(2):361–362, June 1998. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237807.

Lewenstein:1998:BRBa

[3004] Bruce V. Lewenstein. Book review: Media Science before the Great War by Peter Broks. Isis, 89(2):362, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237808.

Hendricks:1998:BRB

[3005] Margo Hendricks. Book review: Frankenstein’s Daughters: Women Writing Science Fiction by Jane Donawerth. Isis, 89(2):362–363, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237809.

Jakab:1998:BRB

[3006] Peter L. Jakab. Book review: Airplane Stability and Control: A History of the Technologies That Made Aviation Possible by Malcolm J. Abzug; E. Eugene Larrabee. Isis, 89(2):363–364, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237810.

Dick:1998:BRB

[3007] Steven J. Dick. Book review: To See the Unseen: A History of Planetary Radar Astronomy by Andrew J. Butrica. Isis, 89(2):364–365, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237811.

Krige:1998:BRBb

[3008] John Krige. Book review: A New Force at a New Frontier: Europe’s Development in the Space Field in the Light of Its Main Actors, Policies, Law, and Activities from Its Beginnings up to the Present by Kevin Madders. Isis, 89(2):365, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237812.

Gerovitch:1998:BRB

[3009] Slava Gerovitch. Book review: Korolev: How One Man Masterminded the Soviet Drive to Beat America to the Moon by James Harford. Isis, 89(2):365–366, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237813.

Fruton:1998:BRB

[3010] Joseph S. Fruton. Book review: Toward a History of Epistemic Things: Synthesizing Proteins in the Test Tube by Hans-Jorg Rheinberger. Isis, 89(2):366–367, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237814.

Meldrum:1998:BRB

[3011] Marcia Meldrum. Book review: The Progress of Experiment: Science and Therapeutic Reform in the United States, 1900–1990 by Harry M. Marks. Isis, 89(2):367–368, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237815.

Stevens:1998:BRBb

[3012] Rosemary A. Stevens. Book review: Making Medical History: The Life and Times of Henry E. Sigerist by Elizabeth Fee; Theodore M. Brown. Isis, 89(2):368–369, June 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237816.

Larsen:1998:BRB

[3013] Oivind Larsen. Book review: A Stand against Tyranny: Norway’s Physicians and the Nazis by Maynard M. Cohen. Isis, 89(2):369–370, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237817.

Cussins:1998:BRB

[3014] Charis Cussins. Book review: The Other Machine: Discourse and Reproductive Technologies by Dion Farquhar. Isis, 89(2):370–371, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237818.

Mark:1998:BRB

[3015] Joan Mark. Book review: Elsie Clews Parsons: Inventing Modern Life by Desley Deacon. Isis, 89(2):371–372, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237819.

Magnus:1998:BRB

[3016] David Magnus and Monique Bourque. Book review: The Modeling of Nature: Philosophy of Science and Philosophy of Nature in Synthesis by William A. Wallace. Isis, 89(2):372–373, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237820.

Selin:1998:BRB

[3017] Helaine Selin. Book review: Isis Cumulative Bibliography, 1986–1995: A Bibliography of the History of Science Formed from the Annual Isis Current Bibliographies by John Neu. Isis, 89(2):373–374, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237821.

Sarjeant:1998:BRB

[3018] William A. S. Sarjeant. Book review: Scientists since 1660: A Bibliography of Biographies by Leslie Howsam. Isis, 89(2):374–375, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237822.

Appel:1998:BRB

[3019] Toby A. Appel. Book review: A Chronology of Medicine and Related Sciences by Leslie T. Morton; Robert J. Moore. Isis, 89(2):375–376, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237823.

Crosland:1998:BRB

[3020] Maurice Crosland. Book review: Histoire et memoire de l’Academie des sciences: Guide de recherches by Eric Brian; Christiane Demeulenaere-Douyere. Isis, 89(2):376–377, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237824.

Anonymous:1998:Cb

[3021] Anonymous. Collections. Isis, 89(2):377–382, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237825.

Anonymous:1998:BMc

[3022] Anonymous. Back matter. Isis, 89(2):383–386, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237826.

Anonymous:1998:FMc

[3023] Anonymous. Front matter. Isis, 89(2):??, June 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237753.

Hansen:1998:MMS

[3024] LeeAnn Hansen. Metaphors of mind and society: The origins of German psychiatry in the revolutionary era. Isis, 89(3):387–409, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237140.

Barton:1998:HLH

[3025] Ruth Barton. “Huxley, Lubbock, and half a dozen others”: Professionals and gentlemen in the formation of the X Club, 1851–1864. Isis, 89(3): 410–444, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237141.

Sengoopta:1998:GPE

[3026] Chandak Sengoopta. Glandular politics: Experimental biology, clinical medicine, and homosexual emancipation in fin-de-siècle central Europe. Isis, 89(3):445–473, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237142.

Feffer:1998:OVC

[3027] Loren Butler Feffer. Oswald Veblen and the capitalization of American mathematics: Raising money for research, 1923–1928. Isis, 89(3):474– 497, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237143.

Junker:1998:BRG

[3028] Thomas Junker. Blumenbach’s racial geometry. Isis, 89(3):498–501, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237144.

Gould:1998:MVG

[3029] Stephen Jay Gould. On mental and visual geometry. Isis, 89(3):502–504, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237145.

Heilbron:1998:TSK

[3030] J. L. Heilbron. Thomas Samuel Kuhn, 18 July 1922–17 June 1996. Isis, 89 (3):505–515, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237146.

Coudert:1998:LE

[3031] Allison P. Coudert, Marjorie Grene, Rhoda Rappaport, Brendan Dooley, and Peter Dear. Letters to the Editor. Isis, 89(3):516–517, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237147.

McLeod:1998:BRB

[3032] Mark W. McLeod. Book review: Shaping World History: Breakthroughs in Ecology, Technology, Science, and Politics by Mary Kilbourne Matossian. Isis, 89(3):518, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237148.

Kosinski:1998:BRB

[3033] Antoni A. Kosinski. Book review: Lectures in the History of Mathematics by Henk J. M. Bos. Isis, 89(3):518–519, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237149.

Sasaki:1998:BRB

[3034] Chikara Sasaki. Book review: A History of Chinese Mathematics by Jean-Claude Martzloff; Stephen S. Wilson. Isis, 89(3):519–520, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237150.

Malet:1998:BRB

[3035] Antoni Malet. Book review: The Fire within the Eye: A Historical Essay on the Nature and Meaning of Light by David Park. Isis, 89(3):520–521, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237151.

Pasachoff:1998:BRB

[3036] Jay M. Pasachoff. Book review: Historical Eclipses and Earth’s Rotation by F. Richard Stephenson. Isis, 89(3):521–522, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237152.

Oreskes:1998:BRB

[3037] Naomi Oreskes. Book review: Gender and Scientific Authority by Barbara Laslett; Sally Gregory Kohlstedt; Helen Longino; Evelynn Hammonds. Isis, 89(3):522–523, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237153.

Haber:1998:BRB

[3038] Samuel Haber. Book review: Smart Jews: The Construction of the Image of Jewish Superior Intelligence by Sander L. Gilman. Isis, 89(3):523–524, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237154.

Aieta:1998:BRB

[3039] Joseph Aieta, III. Book review: Facets of Faith and Science by Jitse M. van der Meer. Isis, 89(3):524–525, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237155.

Post:1998:BRB

[3040] Robert C. Post. Book review: Technological Change: Methods and Themes in the History of Technology by Robert Fox. Isis, 89(3):525–526, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237156.

Multhauf:1998:BRB

[3041] Robert Multhauf. Book review: Gunpowder: The History of an International Technology by Brenda J. Buchanan. Isis, 89(3):526–527, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237157.

Bonner:1998:BRB

[3042] Thomas Neville Bonner. Book review: Western Medicine: An Illustrated History by Irvine Louden. Isis, 89(3):527, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237158.

Broman:1998:BRB

[3043] Thomas Broman. Book review: The Medical World of Early Modern France by Laurence Brockliss; Colin Jones. Isis, 89(3):528–530, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237159.

Nilsson:1998:BRB

[3044] Ingemar Nilsson. Book review: The Shaping of a Profession: Physicians in Norway, Past and Present by Oivind Larsen. Isis, 89(3):530, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237160.

Loudon:1998:BRB

[3045] Irvine Loudon. Book review: Contours of Death and Disease in Early Modern England by Mary J. Dobson. Isis, 89(3):530–532, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237161.

Scull:1998:BRB

[3046] Andrew Scull. Book review: A History of Clinical Psychiatry: The Origin and History of Psychiatric Disorders by German E. Berrios; Roy Porter. Isis, 89(3):532, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237162.

Popplestone:1998:BRB

[3047] John A. Popplestone. Book review: Psychological Ideas and Society: Charles University, 1348–1998 by Josef Brožek; Jiri Hoskovec; Eunice M. Brožek. Isis, 89(3):532–533, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237163.

Todd:1998:BRB

[3048] Dennis Todd. Book review: A Cabinet of Medical Curiosities by Jan Bondeson. Isis, 89(3):533–534, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237164.

Barnes:1998:BRB

[3049] Jonathan Barnes. Book review: Die Philosophie der Antike. Volume 4: Die Hellenistische Philosophie by Michael Erler; Hellmut Flashar; Gunter Gawlick; Woldemar Gorler; Peter Steinmetz. Isis, 89(3):534–535, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237165.

Cullen:1998:BRB

[3050] Christopher Cullen. Book review: Early Chinese Work in Natural Science: A Re-Examination of the Physics of Motion, Acoustics, Astronomy, and Scientific Thoughts by Chen Cheng-Yih. Isis, 89(3):535–536, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237166.

McDermott:1998:BRB

[3051] Peter L. McDermott. Book review: The Theater of Nature: Jean Bodin and Renaissance Science by Ann Blair. Isis, 89(3):536–537, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237167.

Sawday:1998:BRB

[3052] Jonathan Sawday. Book review: The Anatomical Renaissance: The Resurrection of the Anatomical Projects of the Ancients by Andrew Cunningham. Isis, 89(3):537–538, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237168.

Edwards:1998:BRBb

[3053] John Edwards. Book review: Dialogo contra los judios by Pedro Alfonso de Huesca; Maria Jesus Lacarra; Klaus-Peter Mieth; Esperanza Ducay. Isis, 89(3):538–539, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237169.

Ashworth:1998:BRB

[3054] E. J. Ashworth. Book review: Palaestra Rationis: Discussioni su natura della copula e modalita nella filosofia “scolastica” tedesca del XVII secolo by Gino Roncaglia. Isis, 89(3):539, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237170.

Voelkel:1998:BRB

[3055] James R. Voelkel. Book review: The Harmony of the World by Johannes Kepler; E. J. Aiton; A. M. Duncan; J. V. Field. Isis, 89(3):539–540, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237171.

Knobloch:1998:BRB

[3056] Eberhard Knobloch. Book review: Elements de geometrie by G. P. de Roberval; Vincent Jullien. Isis, 89(3):540–541, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237172.

Iliffe:1998:BRB

[3057] Rob Iliffe. Book review: Telling Time: Clocks, Diaries, and English Diurnal Form, 1660–1785 by Stuart Sherman. Isis, 89(3):541–542, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237173.

Giglioni:1998:BRBb

[3058] Guido Giglioni. Book review: Robert Boyle and the Limits of Reason by Jan W. Wojcik. Isis, 89(3):542–543, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237174.

Palm:1998:BRB

[3059] Lodewijk Palm. Book review: The Fabric of Life: Microscopy in the Seventeenth Century by Marian Fournier. Isis, 89(3):543–544, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237175.

Stein:1998:BRB

[3060] Claudia Stein. Book review: Salomon Reisel (1625–1701): Barocke Naturforschung eines Leibarztes im Banne der mechanistischen Philosophie by Ralf Broer. Isis, 89(3):544–545, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237176.

Goudsblom:1998:BRB

[3061] Johan Goudsblom. Book review: A Description of Fire Engines with Water Hoses and the Method of Fighting Fires Now Used in Amsterdam by Jan van der Heyden; Jan van der Heyden,; Lettie Stibbe Multhauf. Isis, 89(3):545–546, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237177.

Grosholz:1998:BRB

[3062] Emily Grosholz. Book review: La resolution des problemes de Descartes a Kant: L’analyse a l’age de la revolution scientifique by Benoit Timmermans. Isis, 89(3):546–547, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237178.

Wilson:1998:BRBa

[3063] Andrew D. Wilson. Book review: The Embodiment of Reason: Kant on Spirit, Generation, and Community by Susan Meld Shell. Isis, 89(3):547, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237179.

Kitcher:1998:BRB

[3064] Patricia Kitcher. Book review: Prolegomena to Any Future Metaphysics That Will Be Able to Come Forward as Science: With Selections from the Critique of Pure Reason by Immanuel Kant; Gary Hatfield. Isis, 89 (3):547–548, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237180.

Lynch:1998:BRB

[3065] William T. Lynch. Book review: Cavendish by Christa Jungnickel; Russell McCormmach. Isis, 89(3):548–549, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237181.

Hahn:1998:BRB

[3066] Roger Hahn. Book review: Jean Paul Marat: Scientist and Revolutionary by Clifford D. Conner. Isis, 89(3):549–550, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237182.

Redondi:1998:BRB

[3067] Pietro Redondi. Book review: Prehistoire et histoire de la thermodynamique classique (Une histoire de la chaleur) by Robert Locqueneux. Isis, 89(3):550–551, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237183.

Grooss:1998:BRB

[3068] Kees Storm Grooss. Book review: Oefening en bespiegeling: Het verloskundig onderwijs van M. S. du Pui (1754–1834) te Leiden by M. M. Lamens-van Malenstein. Isis, 89(3):551, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237184.

Rupke:1998:BRB

[3069] Nicolaas Rupke. Book review: Victorian Science in Context by Bernard Lightman. Isis, 89(3):551–553, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237185.

Yost:1998:BRB

[3070] Robinson M. Yost. Book review: Physics in the Nineteenth Century by Robert D. Purrington. Isis, 89(3):553–554, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237186.

Rowe:1998:BRB

[3071] David E. Rowe. Book review: Mathematics and Mathematicians: Mathematics in Sweden before 1950 by Lars Garding. Isis, 89(3):554–555, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237187.

Lucas:1998:BRB

[3072] A. M. Lucas. Book review: The Scientific Savant in Nineteenth-Century Australia. (Historical Records of Australian Science, 11 [3].) by R. W. Home. Isis, 89(3):555–556, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237188.

Dean:1998:BRB

[3073] Dennis R. Dean. Book review: Theodore Henry Hittell’s The California Academy of Sciences: A Narrative History, 1853–1906 by Alan E. Leviton; Michele L. Aldrich. Isis, 89(3):556–557, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237189.

McGrath:1998:BRB

[3074] Sylvia W. McGrath. Book review: Science on the Texas Frontier: Observations of Dr. Gideon Lincecum by Jerry Bryan Lincecum; Edward Hake Phillips; Peggy A. Redshaw. Isis, 89(3):557, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237190.

Kremer:1998:BRB

[3075] Richard Kremer. Book review: Les sciences physiologiques en Europe: Analyses comparees du XIXe siècle by Christian Bonah. Isis, 89(3): 557–558, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237191.

Sheets-Pyenson:1998:BRB

[3076] Susan Sheets-Pyenson. Book review: George Gordon: An Annotated Catalogue of His Scientific Correspondence by Michael Collie; Susan Bennett. Isis, 89(3):558–559, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237192.

Vogel:1998:BRB

[3077] Morris J. Vogel. Book review: Quarantine! East European Jewish Immigrants and the New York City Epidemics of 1892 by Howard Markel. Isis, 89(3):559–560, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237193.

Fee:1998:BRB

[3078] Elizabeth Fee. Book review: Sir Arthur Newsholme and State Medicine, 1885–1935 by John M. Eyler. Isis, 89(3):560–561, September 1998. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237194.

Brookes:1998:BRB

[3079] Barbara Brookes. Book review: When Abortion Was a Crime: Women, Medicine, and Law in the United States, 1867–1973 by Leslie J. Reagan. Isis, 89(3):561–562, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237195.

Deltete:1998:BRB

[3080] Robert J. Deltete. Book review: Fritz Haber in seiner Korrespondenz mit Wilhelm Ostwald sowie in Briefen an Svante Arrhenius by Regine Zott. Isis, 89(3):562–563, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237196.

Pestre:1998:BRB

[3081] Dominique Pestre. Book review: The Quark Machines: How Europe Fought the Particle Physics War by Gordon Fraser. Isis, 89(3):563–564, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237197.

DeVorkin:1998:BRB

[3082] David H. DeVorkin. Book review: Bright Galaxies, Dark Matters by Vera Rubin. Isis, 89(3):564, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237198.

Hetherington:1998:BRB

[3083] Norriss S. Hetherington. Book review: The Life of the Cosmos by Lee Smolin. Isis, 89(3):564–565, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237199.

Rose:1998:BRB

[3084] Hilary Rose. Book review: Modest Witness@Second Millenium.Femaleman. ©Meets OncoMouse^{T M}: Feminism and Technoscience by Donna J. Haraway. Isis, 89(3):565–566, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237200.

Lewenstein:1998:BRBb

[3085] Bruce V. Lewenstein. Book review: Yes, We Have No Neutrons: An Eye-Opening Tour through the Twists and Turns of Bad Science by A. K. Dewdney. Isis, 89(3):566–567, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237201.

Freeman:1998:BRB

[3086] Donald C. Freeman. Book review: Noam Chomsky: A Life of Dissent by Robert F. Barsky. Isis, 89(3):567–568, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237202.

Guralnick:1998:BRB

[3087] Stanley M. Guralnick. Book review: The Making of the Modern University: Intellectual Transformation and the Marginalization of Morality by Julie A. Reuben. Isis, 89(3):568–569, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237203.

Gingras:1998:BRB

[3088] Yves Gingras. Book review: Les sciences pour l’ingenieur: Histoire du rendez-vous des sciences et de la societe by Girolamo Ramunni. Isis, 89 (3):569–570, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237204.

Adams:1998:BRB

[3089] Mark B. Adams. Book review: Totalitarian Science and Technology by Paul R. Josephson. Isis, 89(3):570–571, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237205.

Kojevnikov:1998:BRBa

[3090] Alexei Kojevnikov. Book review: Essays of a Soviet Scientist: A Revealing Portrait of a Life in Science and Politics by Vitalii I. Gol’danskii. Isis, 89(3):571–572, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237206.

Walker:1998:BRBb

[3091] Mark Walker. Book review: Gerhard Heberer (1901–1973): Sein Beitrag zur Biologie im 20. Jahrhundert by Uwe Hossfeld. Isis, 89(3):572–573, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237207.

Norberg:1998:BRB

[3092] Arthur L. Norberg. Book review: Fundable Knowledge: The Marketing of Defense Technology by A. D. Van Nostrand. Isis, 89(3):573–574, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237208.

Mazumdar:1998:BRBb

[3093] Pauline M. H. Mazumdar. Book review: Exquisite Specificity: The Monoclonal Antibody Revolution by Alberto Cambrosio; Peter Keating. Isis, 89 (3):574–575, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237209.

Anderson:1998:BRB

[3094] Katharine Anderson. Book review: Historical Essays on Meteorology, 1919–1995 by James Rodger Fleming. Isis, 89(3):575, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237210.

Runte:1998:BRB

[3095] Alfred Runte. Book review: Preserving Nature in the National Parks: A History by Richard West Sellars. Isis, 89(3):575–576, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237211.

Beeman:1998:BRB

[3096] Randal Beeman. Book review: To Make a Spotless Orange: Biological Control in California by Richard C. Sawyer. Isis, 89(3):576–577, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237212.

Cherniack:1998:BRB

[3097] Martin Cherniack. Book review: Hazards of the Job: From Industrial Disease to Environmental Health Science by Christopher C. Sellers. Isis, 89(3):577–578, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237213.

Lunbeck:1998:BRB

[3098] Elizabeth Lunbeck. Book review: Keeping America Sane: Psychiatry and Eugenics in the United States and Canada, 1880–1940 by Ian Robert Dowbiggin. Isis, 89(3):578–579, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237214.

Morawski:1998:BRB

[3099] Jill Morawski. Book review: Children, Race, and Power: Kenneth and Mamie Clark’s Northside Center by Gerald Markowitz; David Rosner. Isis, 89(3):579–580, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237215.

Seidel:1998:BRB

[3100] Robert Seidel. Book review: Postmodern War: The New Politics of Conflict by Chris Hables Gray. Isis, 89(3):580–581, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237216.

Hilton:1998:BRB

[3101] George W. Hilton. Book review: Cleveland Transit Vehicles: Equipment and Technology by James A. Toman; Blaine S. Hays. Isis, 89(3):581, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237217.

Pinch:1998:BRB

[3102] Trevor Pinch. Book review: Belief and Resistance: Dynamics of Contemporary Intellectual Controversy by Barbara Herrnstein Smith. Isis, 89 (3):581–583, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237218.

Ruse:1998:BRB

[3103] Michael Ruse. Book review: Philosophie de la biologie by François Duchesneau. Isis, 89(3):583–584, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237219.

Adler:1998:BRB

[3104] Kraig Adler. Book review: Biology as History: Papers from International Conferences Sponsored by the California Academy of Sciences in San Francisco and the Museo Civico di Storia Naturale in Milan. Vol. 1: Systematic Biology as an Historical Science by Giovanni Pinna; Michael T. Ghiselin; Biology as History: Papers from International Conferences Sponsored by the California Academy of Sciences in San Francisco and the Museo Civico di Storia Naturale in Milan. Vol. 2: New Perspectives on the History of Life: Essays on Systematic Biology as Historical Narrative by Michael T. Ghiselin; Giovanni Pinna. Isis, 89(3):584–585, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237220.

Scott:1998:BRB

[3105] Eugenie C. Scott. Book review: The Trouble with Science by Robin Dunbar. Isis, 89(3):585–586, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237221.

Montgomery:1998:BRBb

[3106] Scott L. Montgomery. Book review: The Necessity of Experience by Edward S. Reed. Isis, 89(3):586–587, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237222.

Edmondson:1998:BRB

[3107] Frank K. Edmondson. Book review: History of Astronomy: An Encyclopedia by John Lankford. Isis, 89(3):587, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237223.

Dunlap:1998:BRB

[3108] Thomas R. Dunlap. Book review: Biographical Dictionary of American and Canadian Naturalists and Environmentalists by Keir B. Sterling; Richard P. Harmond; George A. Cevasco; Lorne F. Hammond. Isis, 89 (3):587–588, September 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237224.

Anonymous:1998:Cc

[3109] Anonymous. Collections. Isis, 89(3):588–599, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237225.

Anonymous:1998:BMd

[3110] Anonymous. Back matter. Isis, 89(3):600–604, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237226.

Anonymous:1998:FMd

[3111] Anonymous. Front matter. Isis, 89(3):??, September 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237139.

Nyhart:1998:CEZ

[3112] Lynn K. Nyhart. Civic and economic zoology in Nineteenth-Century Germany: The “living communities” of Karl Möbius. Isis, 89(4):605– 630, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236735.

Lindsay:1998:IIW

[3113] Debra Lindsay. Intimate inmates: Wives, households, and science in Nineteenth-Century America. Isis, 89(4):631–652, December 1998. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236736.

Wright:1998:TMW

[3114] David Wright. The translation of modern Western science in Nineteenth-Century China, 1840–1895. Isis, 89(4):653–673, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236737.

Mazzotti:1998:GGM

[3115] Massimo Mazzotti. The geometers of god: Mathematics and reaction in the kingdom of Naples. Isis, 89(4):674–701, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236738.

Cohen:1998:BRBb

[3116] H. Floris Cohen. Book review: The Scientific Revolution and the Origins of Modern Science by John Henry. Isis, 89(4):702–703, December 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236739.

McCarthy:1998:BRB

[3117] Daniel McCarthy. Book review: Calendrical Calculations by Nachum Dershowitz; Edward M. Reingold. Isis, 89(4):703–704, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236740.

Goldstein:1998:BRB

[3118] Catherine Goldstein. Book review: Vom Zahlstein zum Computer: Mathematik in der Geschichte. Volume 1: Überblick und Biographien by Hans Wussing. Isis, 89(4):704–705, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236741.

Marvin:1998:BRB

[3119] Ursula B. Marvin. Book review: Comets, Popular Culture, and the Birth of Modern Cosmology by Sara Schechner Genuth. Isis, 89(4):705–706, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236742.

Schabas:1998:BRBb

[3120] Margaret Schabas. Book review: Statistical Visions in Time: A History of Time Series Analysis, 1662–1938 by Judy L. Klein. Isis, 89(4):706, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236743.

Hubbard:1998:BRB

[3121] Jennifer M. Hubbard. Book review: Scientists and the Sea, 1650–1900: A Study of Marine Science by Margaret Deacon. Isis, 89(4):706–707, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236744.

Paul:1998:BRB

[3122] Harry Paul. Book review: Histoire des sciences de la vie by Pascal Duris; Gabriel Gohau. Isis, 89(4):707–708, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236745.

MacLeod:1998:BRB

[3123] Roy MacLeod. Book review: The Furthest Shore: Images of Terra Australis from the Middle Ages to Captain Cook by William Eisler. Isis, 89 (4):708–710, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236746.

Hardy:1998:BRB

[3124] Anne Hardy. Book review: Growing Artificial Societies: Social Science from the Bottom Up by Joshua M. Epstein; Robert Axtell; Epidemics and History: Disease, Power, and Imperialism by Sheldon Watts. Isis, 89(4):710–711, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236747.

Li:1998:BRB

[3125] Alison Li. Book review: Science and Homosexualities by Vernon A. Rosario. Isis, 89(4):711–712, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236748.

Reagan:1998:BRB

[3126] Leslie J. Reagan. Book review: Fetal Positions: Individualism, Science, Visuality by Karen Newman. Isis, 89(4):712–713, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236749.

Lowood:1998:BRB

[3127] Henry Lowood. Book review: Milestones in Colour Printing, 1457–1859 by Bamber Gascoigne. Isis, 89(4):713–714, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236750.

Grabiner:1998:BRB

[3128] Judith V. Grabiner. Book review: Polyhedra by Peter R. Cromwell. Isis, 89(4):714–715, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236751.

Palmerino:1998:BRB

[3129] Carla Rita Palmerino. Book review: Consequentia mirabilis: Una regola logica tra matematica e filosofia by Fabio Bellissima; Paolo Pagli. Isis, 89 (4):715–716, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236752.

Keyser:1998:BRB

[3130] Paul T. Keyser. Book review: An Approach to Aristotle’s Physics: With Particular Attention to the Role of His Manner of Writing by David Bolotin. Isis, 89(4):716–717, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236753.

Evans:1998:BRB

[3131] James Evans. Book review: The Comet of 44 B.C. and Caesar’s Funeral Games by John T. Ramsey; A. Lewis Licht. Isis, 89(4):717–718, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236754.

Scarborough:1998:BRB

[3132] John Scarborough. Book review: Celse de la medecine. Volume 1: Livres I-II by Guy Serbat. Isis, 89(4):718–719, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/236755.

Russell:1998:BRB

[3133] G. A. Russell. Book review: Ptolemy’s Theory of Visual Perception: An English Translation of the Optics with Introduction and Commentary by A. Mark Smith. Isis, 89(4):719–720, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236756.

Webster:1998:BRB

[3134] Charles Webster. Book review: Paracelsus by Udo Benzenhofer. Isis, 89 (4):720–721, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236757.

Pumfrey:1998:BRB

[3135] Stephen Pumfrey. Book review: Francis Bacon and the Politics of Science by John E. Leary. Isis, 89(4):721–722, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236758.

Hill:1998:BRB

[3136] Katherine Hill. Book review: The History of the University of Oxford. Volume 4: Seventeenth-Century Oxford by Nicholas Tyacke. Isis, 89 (4):722–723, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236759.

Parmentier:1998:BRB

[3137] Marc Parmentier. Book review: Mathematische Schriften. Volume 2: 1672–1676. Algebra (Part 2) by Gottfried Wilhelm Leibniz. Isis, 89(4):

723–724, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236760.

Palaia:1998:BRB

[3138] Roberto Palaia. Book review: Corpo e funzioni cognitive in Leibniz by Enrico Pasini. Isis, 89(4):724–725, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236761.

Luthy:1998:BRB

[3139] Christoph Luthy. Book review: Le passioni degli atomi: Montanari e Rossetti, una polemica tra galileiani by Susana Gomez Lopez. Isis, 89 (4):725–726, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236762.

Chapman:1998:BRB

[3140] Allan Chapman. Book review: At the Sign of the Oriental Lamp: The Musschenbroek Workshop in Leiden, 1660–1750 by Peter de Clercq. Isis, 89(4):726, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236763.

Schneer:1998:BRB

[3141] Cecil J. Schneer. Book review: When Geologists Were Historians: 1665– 1750 by Rhoda Rappaport. Isis, 89(4):726–727, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236764.

Hall:1998:BRB

[3142] Michael G. Hall. Book review: The Great South Sea: English Voyages and Encounters, 1570–1750 by Glyndwr Williams. Isis, 89(4):727–728, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236765.

Richards:1998:BRBb

[3143] Joan L. Richards. Book review: Symbols, Impossible Numbers, and Geometric Entanglements: British Algebra through the Commentaries on Newton’s Universal Arithmetick by Helena M. Pycior. Isis, 89(4):728– 729, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236766.

Morello:1998:BRB

[3144] Nicoletta Morello. Book review: Histoire de la geologie. Volume 2: La grande eclosion et ses premices, 1660–1810 by François Ellenberger. Isis, 89(4):729–730, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236767.

Gessinger:1998:BRB

[3145] Joachim Gessinger. Book review: Chladni und die Entwicklung der Akustik von 1750–1860 by Dieter Ullmann. Isis, 89(4):730–731, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236768.

Gough:1998:BRBb

[3146] Jerry B. Gough. Book review: Oeuvres de Lavoisier: Correspondence. Volume 6: 1789–1791 by Antoine Lavoiser; Patrice Bret. Isis, 89(4): 731–732, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236769.

McClellan:1998:BRBb

[3147] James E. McClellan, III. Book review: Du Faubourg Montmartre au Corps des mines: L’etonnant parcours du Republicain J. H. Hassenfratz (1755–1827) by Emmanuel Grison. Isis, 89(4):732–733, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236770.

Diamond:1998:BRB

[3148] Peter J. Diamond. Book review: Thomas Reid on the Animate Creation: Papers Relating to the Life Sciences by Paul Wood. Isis, 89(4):733–734, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236771.

Wilson:1998:BRBb

[3149] Renate Wilson. Book review: Friedrich Albrecht Carl Gren in seiner Zeit, 1760–1798: Spekulant oder Selbstdenker? by Markus Seils. Isis, 89 (4):734–735, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236772.

Lindemann:1998:BRB

[3150] Mary Lindemann. Book review: Arzneimittelversorgung im 18. Jahrhundert: Die Stadt Braunschweig und die landlichen Distrikte im Herzogtum Braunschweig-Wolfenbuttel by Gabriele Beisswanger. Isis, 89(4):735–736, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236773.

Hannaway:1998:BRB

[3151] Caroline Hannaway. Book review: Health and Healing in Eighteenth-Century Germany by Mary Lindemann; The Transformation of German Academic Medicine, 1750–1820 by Thomas H. Broman. Isis, 89(4):736– 737, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236774.

King:1998:BRBa

[3152] Helen King. Book review: The Making of Man-Midwifery: Childbirth in England, 1660–1770 by Adrian Wilson. Isis, 89(4):737–738, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236775.

Ellis:1998:BRB

[3153] John H. Ellis. Book review: A Melancholy Scene of Devastation: The Public Response to the 1793 Philadelphia Yellow Fever Epidemic by J. Worth Estes; Billy G. Smith. Isis, 89(4):738–740, December 1998. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236776.

Peset:1998:BRB

[3154] Jose Luis Peset. Book review: Enfermedad y sociedad en la crisis colonial del Antiguo Regimen: Nueva Granada en el transito del siglo XVIII al XIX: Las epidemias de viruelas by Marcelo Frias Nunez. Isis, 89(4): 740–741, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236777.

Munday:1998:BRBb

[3155] Pat Munday. Book review: Justus von Liebig: The Chemical Gatekeeper by William H. Brock. Isis, 89(4):741–742, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236778.

Hurt:1998:BRB

[3156] R. Douglas Hurt. Book review: Sowing Modernity: America’s First Agricultural Revolution by Peter D. McClelland. Isis, 89(4):742, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236779.

Saint-Sernin:1998:BRB

[3157] Bertrand Saint-Sernin. Book review: Probabilites et critique philosophique selon Cournot by Thierry Martin. Isis, 89(4):742–743, December 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236780.

Greenaway:1998:BRB

[3158] Frank Greenaway. Book review: The Practice of Science in the Nineteenth Century: Teaching and Research Apparatus in the Teyler Museum by Gerard L’E. Turner. Isis, 89(4):743–744, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236781.

Fichman:1998:BRB

[3159] Martin Fichman. Book review: Where Worlds Collide: The Wallace Line by Penny van Oosterzee. Isis, 89(4):744–745, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236782.

Cateforis:1998:BRB

[3160] David Cateforis. Book review: Art and the Higher Life: Painting and Evolutionary Thought in Late Nineteenth-Century America by Kathleen Pyne. Isis, 89(4):745–746, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/236783.

Rosenband:1998:BRB

[3161] Leonard N. Rosenband. Book review: Productivity and Performance in the Paper Industry: Labour, Capital, and Technology in Britain and America, 1860–1914 by Gary Bryan Magee. Isis, 89(4):746, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236784.

Rosenkrantz:1998:BRB

[3162] Barbara Gutmann Rosenkrantz. Book review: Fevered Lives: Tuberculosis in American Culture since 1870 by Katherine Ott. Isis, 89(4): 746–748, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236785.

Daniels:1998:BRB

[3163] Peter T. Daniels. Book review: The Conquest of Assyria: Excavations in an Antique Land, 1840–1860 by Mogens Trolle Larsen; From Nineveh to New York: The Strange Story of the Assyrian Reliefs in the Metropolitan Museum and the Hidden Masterpiece at Canford School by John Malcolm Russell; Judith McKenzie; Stephanie Dalley. Isis, 89(4):748–750, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236786.

Mergen:1998:BRB

[3164] Bernard Mergen. Book review: I May Be Some Time: Ice and the English Imagination by Francis Spufford. Isis, 89(4):750–751, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236787.

Edelson:1998:BRB

[3165] Paul J. Edelson. Book review: Osler: Inspirations from a Great Physician by Charles S. Bryan. Isis, 89(4):751, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236788.

Gorelik:1998:BRB

[3166] Gennady Gorelik. Book review: The Origin of the Concept of Nuclear Forces by Laurie M. Brown; Helmut Rechenberg. Isis, 89(4):751–752, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236789.

Kojevnikov:1998:BRBb

[3167] Alexei Kojevnikov. Book review: Infinite Potential: The Life and Times of David Bohm by F. David Peat. Isis, 89(4):752–753, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236790.

Gorski:1998:BRB

[3168] William Gorski. Book review: Quantum Poetics: Yeats, Pound, Eliot, and the Science of Modernism by Daniel Albright. Isis, 89(4):753–754, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236791.

Elliott:1998:BRB

[3169] Clark A. Elliott. Book review: The Historiography of Contemporary Science and Technology by Thomas Soderqvist. Isis, 89(4):754–755, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236792.

Olwell:1998:BRB

[3170] Russell Olwell. Book review: Atomic Spaces: Living on the Manhattan Project by Peter Bacon Hales. Isis, 89(4):755–756, December 1998.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236793.

Westfall:1998:BRB

[3171] Catherine Westfall. Book review: Argonne National Laboratory, 1946– 96 by Jack M. Holl; Richard G. Hewlett; Ruth R. Harris. Isis, 89(4): 756–757, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236794.

Leslie:1998:BRB

[3172] Stuart W. Leslie. Book review: New Atlantis Revisited: Akademgorodok, the Siberian City of Science by Paul R. Josephson. Isis, 89(4):757–759, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236795.

Leverington:1998:BRB

[3173] David Leverington. Book review: AURA and Its U.S. National Observatories by Frank K. Edmondson. Isis, 89(4):759–760, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236796.

Skandalis:1998:BRB

[3174] Tina Skandalis and Lawrence Badash. Book review: A Devotion to Their Science: Pioneer Women of Radioactivity by Marelene F. Rayner-Canham; Geoffrey W. Rayner-Canham. Isis, 89(4):760–761, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236797.

Grandy:1998:BRB

[3175] David Grandy. Book review: The Voice of the Martians by George Marx. Isis, 89(4):761, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236798.

Kline:1998:BRB

[3176] Ronald Kline, William Lynch, and Jameson Wetmore. Book review: Engineering Ethics: Balancing Cost, Schedule, and Risk–Lessons Learned from the Space Shuttle by Rosa Lynn B. Pinkus; Larry J. Shuman; Norman P. Hummon; Harvey Wolfe; The Challenger Launch Decision: Risky Technology, Culture, and Deviance at NASA by Diane Vaughan. Isis, 89(4):761–763, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236799.

Samuels:1998:BRB

[3177] Warren J. Samuels. Book review: Schumpeter and the Idea of Social Science: A Metatheoretical Study by Yuichi Shionoya. Isis, 89(4):763– 764, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236800.

Allison-Bunnell:1998:BRB

[3178] Steven Allison-Bunnell. Book review: Edison in Florida: The Green Laboratory by Olav Thulesius. Isis, 89(4):764–765, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236801.

Lankton:1998:BRB

[3179] Larry Lankton. Book review: Mineheads by Bernd Becher; Hilla Becher. Isis, 89(4):765–766, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236802.

Malpas:1998:BRB

[3180] Constance Malpas. Book review: Picturing Bushmen: The Denver African Expedition of 1925 by Robert J. Gordon. Isis, 89(4):766, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/236803.

Alfonso-Goldfarb:1998:BRB

[3181] Ana Maria Alfonso-Goldfarb. Book review: Oswaldo Cruz: A construcao de um mito na ciencia brasileira by Nara Britto. Isis, 89(4):767, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236804.

Best:1998:BRB

[3182] Joel Best. Book review: Making Sense of Illness: Science, Society, and Disease by Robert A. Aronowitz. Isis, 89(4):767–768, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236805.

Gilbert:1998:BRB

[3183] Robert E. Gilbert. Book review: Eisenhower’s Heart Attack: How Ike Beat Heart Disease and Held on to the Presidency by Clarence G. Lasby. Isis, 89(4):768–769, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

236806.

Gross:1998:BRB

[3184] Alan Gross. Book review: Science by Steve Fuller. Isis, 89(4):769–770, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/236807.

Zehr:1998:BRB

[3185] Stephen Zehr. Book review: Uncertain Knowledge: An Image of Science for a Changing World by R. G. A. Dolby. Isis, 89(4):770–771, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/236808.

Mayer:1998:BRB

[3186] Andreas Mayer. Book review: Power and Invention: Situating Science by Isabelle Stengers; Paul Bains. Isis, 89(4):771–772, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236809.

Smith:1998:BRB

[3187] Barbara Herrnstein Smith. Book review: On the Origin of Objects by Brian Cantwell Smith. Isis, 89(4):772–773, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236810.

Zabell:1998:BRB

[3188] S. L. Zabell. Book review: Artificial Intelligence and Scientific Method by Donald Gillies. Isis, 89(4):773–774, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/236811.

Demaitre:1998:BRB

[3189] Luke Demaitre. Book review: Lexique de la langue scientifique (astrologie, mathematiques, medecine…): Materiaux pour le “Dictionnaire du Moyen Français” (DMF)-4 by Daniele Jacquart; Claude Thomasset. Isis, 89(4):774, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236812.

King:1998:BRBb

[3190] William M. King. Book review: Distinguished African American Scientists of the Twentieth Century by James H. Kessler; J. S. Kidd; Renee A. Kidd; Katherine A. Morin. Isis, 89(4):774–775, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236813.

Tuchman:1998:BRB

[3191] Arleen Marcia Tuchman. Book review: Doctors, Nurses, and Medical Practitioners: A Bio-Bibliographical Sourcebook by Lois N. Magner. Isis, 89(4):775–776, December 1998. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/236814.

Anonymous:1998:Cd

[3192] Anonymous. Collections. Isis, 89(4):776–778, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236815.

Anonymous:1998:BMe

[3193] Anonymous. Back matter. Isis, 89(4):779–783, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236816.

Anonymous:1998:VI

[3194] Anonymous. Volume information. Isis, 89(4):784–799, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/236733.

Anonymous:1998:FMe

[3195] Anonymous. Front matter. Isis, 89(4):??, December 1998. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/236734.

Neu:1999:CBH

[3196] John Neu. Current bibliography of the history of science and its cultural influences, 1999. Isis, 90(??):1–309, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237901.

Anonymous:1999:VIa

[3197] Anonymous. Volume information. Isis, 90(??):333–359, ???? 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/238001.

Andresen:1999:CK

[3198] Jensine Andresen. Crisis and kuhn. Isis, 90(??):S43–S67, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/238006.

Anonymous:1999:ABP

[3199] Anonymous. Appendix B: Prizes awarded by the History of Science Society. Isis, 90(??):S323–S330, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/238022.

Anonymous:1999:BMa

[3200] Anonymous. Back matter. Isis, 90(??):??, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237902.

Anonymous:1999:BMb

[3201] Anonymous. Back matter. Isis, 90(??):S331–S332, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/238023.

Anonymous:1999:FMa

[3202] Anonymous. Front matter. Isis, 90(??):i–xvii, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237900.

Anonymous:1999:FMb

[3203] Anonymous. Front matter. Isis, 90(??):S5–S319, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/238002.

Benson:1999:FCI

[3204] Keith R. Benson. Flail on, columbia: An irreverent look at HSS’s soggiest subsection, the Columbia history of science group. Isis, 90(??):S240– S245, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/238017.

Cohen:1999:BCC

[3205] I. Bernard Cohen. The Isis crises and the coming of age of the History of Science Society. Isis, 90(??):S28–S42, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/238005.

Elliott:1999:FHS

[3206] Clark A. Elliott. Forum for the history of science in America: Identity and organization. Isis, 90(??):S226–S239, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/238016.

Gariepy:1999:JFF

[3207] Thomas P. Gariepy. John Farquhar Fulton and the History of Science Society. Isis, 90(??):S7–S27, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

238004.

Gillispie:1999:APV

[3208] Charles C. Gillispie. Apologia pro vita sua. Isis, 90(??):S84–S94, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/238008.

Gleason:1999:MNY

[3209] Mary Louise Gleason. The metropolitan new York section of the History of Science Society. Isis, 90(??):S200–S218, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/238014.

Grau:1999:FND

[3210] Kevin T. Grau. Force and nature: The Department of History and Philosophy of Science at Indiana University, 1960–1998. Isis, 90(??): S295–S318, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/238020.

Hahn:1999:BHS

[3211] Roger Hahn. Berkeley’s history of science dinner club: A chronicle of fifty years of activity. Isis, 90(??):S182–S191, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/238012.

Hall:1999:RHS

[3212] Marie Boas Hall. Recollections of a history of science guinea pig. Isis, 90(??):S68–S83, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/238007.

Harvey:1999:HSH

[3213] Joy Harvey. History of science, history and science, and natural sciences: Undergraduate teaching of the history of science at Harvard, 1938–1970. Isis, 90(??):S270–S294, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

238019.

Henson:1999:OCR

[3214] Pamela M. Henson. “Objects of curious research”: The history of science and technology at the Smithsonian. Isis, 90(??):S249–S269, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/238018.

Holton:1999:SLL

[3215] Gerald Holton. Some lessons from living in the history of science. Isis, 90(??):S95–S116, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/238009.

Provine:1999:NFW

[3216] Will Provine. No free will. Isis, 90(??):S117–S132, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/238010.

Rossiter:1999:I

[3217] Margaret W. Rossiter. Introduction. Isis, 90(??):S1–S3, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/238003.

Schofield:1999:TFG

[3218] Robert E. Schofield. “Too far to go”: Early years of the Midwest Junto. Isis, 90(??):S192–S199, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

238013.

Sokal:1999:AOH

[3219] Michael M. Sokal and G. E. Erikson. Appendix A: Officers of the History of Science Society, 1924–99. Isis, 90(??):S321–S322, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/238021.

Sokal:1999:HSS

[3220] Michael M. Sokal. The History of Science Society, 1970–1999: From subscription agency to professional society. Isis, 90(??):S135–S181, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/238011.

Winsor:1999:JAS

[3221] Mary P. Winsor and Leonard G. Wilson. The joint Atlantic seminar in history of biology. Isis, 90(??):S219–S225, ???? 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/238015.

Hecht:1999:SMD

[3222] Jennifer Michael Hecht. The solvency of metaphysics: The debate over racial science and moral philosophy in France, 1890–1919. Isis, 90(1): 1–24, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237472.

Sur:1999:AAC

[3223] Abha Sur. Aesthetics, authority, and control in an Indian laboratory: The Raman–Born controversy on lattice dynamics. Isis, 90(1):25–49, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237473.

Hankins:1999:BDN

[3224] Thomas L. Hankins. Blood, dirt, and nomograms: A particular history of graphs. Isis, 90(1):50–80, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237474.

Pesic:1999:WPF

[3225] Peter Pesic. Wrestling with Proteus: Francis Bacon and the “torture” of nature. Isis, 90(1):81–94, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237475.

Germano:1999:LE

[3226] Massimo Germano, James Evans, and Mario Gomez-Torrente. Letters to the Editor. Isis, 90(1):95–97, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237476.

Root-Bernstein:1999:BRB

[3227] Robert S. Root-Bernstein. Book review: Beauty of Another Order: Photography in Science by Ann Thomas. Isis, 90(1):98–99, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237477.

Hoskin:1999:BRB

[3228] Michael Hoskin. Book review: Astronomy through the Ages: The Story of the Human Attempt to Understand the Universe by Robert Wilson. Isis, 90(1):99, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237478.

Otis:1999:BRB

[3229] Laura Otis. Book review: Science Incarnate: Historical Embodiments of Natural Knowledge by Christopher Lawrence; Steven Shapin. Isis, 90(1):99–100, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237479.

Lang:1999:BRB

[3230] Harry G. Lang. Book review: The Body and Physical Difference: Discourses of Disability by David T. Mitchell; Sharon L. Snyder. Isis, 90 (1):100–101, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237480.

Smith:1999:BRB

[3231] Susan L. Smith. Book review: Killing the Black Body: Race, Reproduction, and the Meaning of Liberty by Dorothy Roberts. Isis, 90(1): 101–102, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237481.

Flemming:1999:BRB

[3232] Rebecca Flemming. Book review: Eve’s Herbs: A History of Contraception and Abortion in the West by John M. Riddle. Isis, 90(1):102–103, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237482.

McCalman:1999:BRB

[3233] Janet McCalman. Book review: Viruses, Plagues, and History by Michael B. A. Oldstone. Isis, 90(1):103, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237483.

Hamlin:1999:BRB

[3234] Christopher Hamlin. Book review: The Search for the Ultimate Sink: Urban Pollution in Historical Perspective by Joel A. Tarr. Isis, 90(1): 103–104, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237484.

Dwyer:1999:BRB

[3235] Ellen Dwyer. Book review: A History of Psychiatry: From the Era of the Asylum to the Age of Prozac by Edward Shorter. Isis, 90(1):104– 105, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237485.

Eigen:1999:BRBa

[3236] Joel Peter Eigen. Book review: Evil or Ill? Justifying the Insanity Defense by Lawrie Reznek. Isis, 90(1):105–106, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237486.

Rudavsky:1999:BRB

[3237] Shari Rudavsky. Book review: Handwriting in America: A Cultural History by Tamara Plakins Thornton. Isis, 90(1):106–107, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237487.

Simonton:1999:BRB

[3238] Dean Keith Simonton. Book review: Strange Brains and Genius: The Secret Lives of Eccentric Scientists and Madmen by Clifford Pickover. Isis, 90(1):107–108, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237488.

Lloyd:1999:BRB

[3239] G. E. R. Lloyd. Book review: Theophrastus: Reappraising the Sources by Johannes M. van Ophuijsen; Marlein van Raalte. Isis, 90(1):108– 109, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237489.

Cosans:1999:BRB

[3240] Christopher E. Cosans. Book review: On the Elements According to Hippocrates by Galen; Phillip De Lacy. Isis, 90(1):109, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237490.

Folkerts:1999:BRB

[3241] Menso Folkerts. Book review: More mathematicorum: Rezeption und Transformation der antiken Gestalten wissenschaftlichen Wissens im 12. Jahrhundert by Mechthild Dreyer. Isis, 90(1):109–110, March 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237491.

Crozet:1999:BRB

[3242] Pascal Crozet. Book review: Al-sijzi’s Treatise on Geometrical Problem Solving by Al-Sijzi; Jan P. Hogendijk; Mohammad Bagheri. Isis, 90(1): 110–111, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237492.

Eastwood:1999:BRB

[3243] Bruce Eastwood. Book review: Astronomies and Cultures in Early Medieval Europe by Stephen C. McCluskey. Isis, 90(1):111–112, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237493.

Grant:1999:BRB

[3244] Edward Grant. Book review: Speculum Divinorum et Quorundam Naturalium. Parts XX-XXIII: On the Heavens, the Divine Movers, and the First Intellect by Henricus Bate; Carlos Steel; Guy Guldentops. Isis, 90(1):112–113, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237494.

Gosselin:1999:BRB

[3245] Edward A. Gosselin. Book review: Giordano Bruno and the Kabbalah: Prophets, Magicians, and Rabbis by Karen Silvia de Leon-Jones. Isis, 90(1):113–114, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237495.

Celeyrette:1999:BRB

[3246] Jean Celeyrette. Book review: Nicolaus Oresmes Kommentar zur Physik des Aristoteles: Kommentar mit Edition der Quaestionen zu Buch 3 und 4 der Aristotelischen Physik sowie von vier Quaestionen zu Buch 5 by Stefan Kirschner. Isis, 90(1):114–115, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237496.

Lande:1999:BRB

[3247] Nelson P. Lande. Book review: The Legacy of Scholasticism in Economic Thought: Antecedents of Choice and Power by Odd Langholm. Isis, 90 (1):115–116, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237497.

Pesenti:1999:BRB

[3248] Tiziana Pesenti. Book review: Guigonis de Caulhiaco (Guy de Chauliac), Inventarium sive Chirurgia Magna. Volume 1: Text by Michael R. McVaugh. Isis, 90(1):116–117, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237498.

Barker:1999:BRB

[3249] Peter Barker. Book review: The Word of God and the Languages of Man: Interpreting Nature in Early Modern Science and Medicine. Volume 1:

Ficino to Descartes by James J. Bono. Isis, 90(1):117, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237499.

Vickers:1999:BRBa

[3250] Brian Vickers. Book review: Philosophical Studies, c. 1611–1619. Volume 6 by Francis Bacon; Graham Rees; Michael Edwards. Isis, 90(1):117– 119, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237500.

Roberts:1999:BRB

[3251] Lissa Roberts. Book review: Pierre-Simon Laplace, 1749–1827: A Life in Exact Science by Charles Coulston Gillispie; Robert Fox; Ivor Grattan-Guinness. Isis, 90(1):119–120, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237501.

Bryden:1999:BRB

[3252] David Bryden. Book review: The Quest for Longitude: The Proceedings of the Longitude Symposium, Harvard University, Cambridge, Massachusetts, November 4–6, 1993 by William J. H. Andrewes. Isis, 90(1): 120–121, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237502.

Navarro-Brotons:1999:BRB

[3253] Victor Navarro-Brotons. Book review: Die Entwicklung des Meridiankreises, 1700–1850: Genesis eines astronomischen Hauptinstrumentes unter Berucksichtigung des Wechselverhaltnisses zwischen Astronomie, Astro-Technik, und Technik by Klaus-Dieter Herbst. Isis, 90(1):121–122, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237503.

Harley:1999:BRB

[3254] David Harley. Book review: The Letters of Dominique Chaix, Botanist-Cure by Roger L. Williams. Isis, 90(1):122–123, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237504.

Beretta:1999:BRBa

[3255] Marco Beretta. Book review: Antoine Lavoisier–The Next Crucial Year; Or, The Sources of His Quantitative Method in Chemistry by Frederic Lawrence Holmes. Isis, 90(1):123–124, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237505.

Long:1999:BRB

[3256] Pamela O. Long. Book review: Bayerisches Messing: Franz Matthias Ellmayrs “Mossing-Werkh AO. 1780”: Studien zur Geschichte, Technologie, und zum sozialen Umfeld der Messingerzeugung im vorindustriellen Bayern by Claus Priesner. Isis, 90(1):124–125, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237506.

Stewart:1999:BRB

[3257] Larry Stewart. Book review: Defoe and the New Sciences by Ilse Vickers. Isis, 90(1):125–126, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237507.

Sangwan:1999:BRB

[3258] Satpal Sangwan. Book review: Mapping an Empire: The Geographical Construction of British India, 1765–1843 by Matthew H. Edney. Isis, 90(1):126–127, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237508.

Guerrini:1999:BRBa

[3259] Anita Guerrini. Book review: “Take Time by the Forelock”: The Letters of Anthony Fothergill to James Woodforde, 1789–1813 by Christopher Lawrence; Paul Lucier; Christopher C. Booth. Isis, 90(1):127, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237509.

Hunt:1999:BRB

[3260] Bruce J. Hunt. Book review: Scientific Credibility and Technical Standards in Nineteenth- and Twentieth–Century Germany and Britain by Jed Z. Buchwald. Isis, 90(1):128–129, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237510.

Plotkin:1999:BRB

[3261] Howard Plotkin. Book review: The Tucson Meteorites: Their History from Frontier Arizona to the Smithsonian by Richard R. Willey. Isis, 90(1):129–130, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237511.

Beaver:1999:BRB

[3262] Donald deB. Beaver. Book review: Bright Paradise: Victorian Scientific Travellers by Peter Raby. Isis, 90(1):130–131, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237512.

Schwartz:1999:BRB

[3263] Joel S. Schwartz. Book review: The Correspondence of Charles Darwin. Volume 10: 1862 by Charles Darwin; Frederick Burkhardt; Duncan M. Porter; Joy Harvey; Jonathan R. Topham. Isis, 90(1):131, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237513.

Taylor:1999:BRB

[3264] Michael Taylor. Book review: Social Darwinism in European and American Thought, 1860–1945: Nature as Model and Nature as Threat by Mike Hawkins. Isis, 90(1):131–132, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237514.

Rossi:1999:BRB

[3265] William Rossi. Book review: A Wider View of the Universe: Henry Thoreau’s Study of Nature by Robert Kuhn McGregor. Isis, 90(1):132– 133, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237515.

Jacyna:1999:BRBa

[3266] Stephen Jacyna. Book review: To See with a Better Eye: A Life of R. T. H. Laennec by Jacalyn Duffin. Isis, 90(1):133–134, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237516.

Cassedy:1999:BRB

[3267] James H. Cassedy. Book review: Against the Spirit of System: The French Impulse in Nineteenth-Century American Medicine by John Harley Warner. Isis, 90(1):134–135, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237517.

Paoloni:1999:BRB

[3268] Leonello Paoloni. Book review: Il pane attossicato: Storia dell’industria dei fiammiferi in Italia, 1860–1910 by Nicoletta Nicolini. Isis, 90(1):135– 136, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237518.

Farley:1999:BRBa

[3269] John Farley. Book review: Ronald Ross: Malariologist and Polymath: A Biography by Edwin R. Nye; Mary E. Gibson. Isis, 90(1):136–137, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237519.

Fullmer:1999:BRB

[3270] June Z. Fullmer. Book review: The Life and Times of Sir Goldsworthy Gurney: Gentleman Scientist and Inventor, 1793–1875 by Dale H. Porter. Isis, 90(1):137–138, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237520.

Carlson:1999:BRB

[3271] W. Bernard Carlson. Book review: Wizard: The Life and Times of Nikola Tesla: Biography of a Genius by Marc J. Seifer. Isis, 90(1):138– 139, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237521.

Brooks:1999:BRB

[3272] John I. Brooks, III. Book review: La decouverte du social: Naissance de la sociologie en France (1870–1914) by Laurent Mucchielli. Isis, 90 (1):139–140, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237522.

Kusch:1999:BRB

[3273] Martin Kusch. Book review: Volkerpsychologie–Versuch einer Neuentdeckung: Texte von Lazarus, Steinthal, und Wundt by Georg Eckardt. Isis, 90(1):140, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237523.

Bourque:1999:BRBa

[3274] Monique Bourque. Book review: To Live in the New World: A. J. Downing and American Landscape Gardening by Judith K. Major. Isis, 90(1):140–142, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237524.

Kiefer:1999:BRBa

[3275] Geraldine Wojno Kiefer. Book review: Burning with Desire: The Conception of Photography by Geoffrey Batchen. Isis, 90(1):142–143, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237525.

Brown:1999:BRN

[3276] Laurie M. Brown. Book review: Niels Bohr and Jørgen Kalckar: Niels Bohr: Collected Works. Volume 7: Foundations of Quantum Physics II (1933–1958). Isis, 90(1):143–144, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237526.

Robotti:1999:BRB

[3277] Nadia Robotti. Book review: Nuove radiazioni, quanti e relatività in Italia, 1895–1926 by Paolantonio Marazzini. Isis, 90(1):144–145, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237527.

Johnson:1999:BRBa

[3278] Karen E. Johnson. Book review: “Um mich nicht zu vergessen”: Hertha Sponer–Ein Frauenleben fur die Physik im 20. Jahrhundert by Marie-Ann Maushart. Isis, 90(1):145–146, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237528.

Stamhuis:1999:BRB

[3279] Ida H. Stamhuis. Book review: Dijksterhuis: Een Biografie by Klaas van Berkel. Isis, 90(1):146–147, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237529.

Wang:1999:BRBa

[3280] Zuoyue Wang. Book review: Frontiers of Illusion: Science, Technology, and the Politics of Progress by Daniel Sarewitz. Isis, 90(1):147–148, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237530.

Champlin:1999:BRB

[3281] Peggy Champlin. Book review: Fieldwork: A Geologist’s Memoir of the Kalahari by Christopher Scholz. Isis, 90(1):148, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237531.

Slack:1999:BRB

[3282] Nancy G. Slack. Book review: Walking with Muir across Yosemite by Thomas R. Vale; Geraldine R. Vale. Isis, 90(1):148–149, March 1999.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237532.

Runte:1999:BRB

[3283] Alfred Runte. Book review: Laurance S. Rockefeller: Catalyst for Conservation by Robin W. Winks. Isis, 90(1):149–150, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237533.

Barrow:1999:BRB

[3284] Mark V. Barrow, Jr. Book review: Spectacular Nature: Corporate Culture and the Sea World Experience by Susan G. Davis. Isis, 90(1):150– 151, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237534.

Slaton:1999:BRB

[3285] Amy Slaton. Book review: Enough for One Lifetime: Wallace Carothers, Inventor of Nylon by Matthew E. Hermes. Isis, 90(1):151–152, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237535.

Parrott:1999:BRB

[3286] Charles Parrott. Book review: The Rational Factory: Architecture, Technology, and Work in America’s Age of Mass Production by Lindy Biggs. Isis, 90(1):152–153, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237536.

Payne:1999:BRB

[3287] Matthew Payne. Book review: Industry, State, and Society in Stalin’s Russia, 1926–1934 by David R. Shearer. Isis, 90(1):153–154, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237537.

Junod:1999:BRB

[3288] Suzanne White Junod. Book review: Our Children’s Toxic Legacy: How Science and Law Fail to Protect Us from Pesticides by John Wargo. Isis, 90(1):154–155, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237538.

Mazumdar:1999:BRB

[3289] Pauline M. H. Mazumdar. Book review: Drawing Blood: Technology and Disease Identity in Twentieth–Century America by Keith Wailoo. Isis, 90(1):155–156, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237539.

Lindee:1999:BRB

[3290] Susan Lindee. Book review: Crafting Science: A Socio-history of the Quest for the Genetics of Cancer by Joan H. Fujimura. Isis, 90(1):156– 157, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237540.

Swazey:1999:BRB

[3291] Judith P. Swazey. Book review: Between Bench and Bedside: Science, Healing, and Interleukin-2 in a Cancer Ward by Ilana Lowy. Isis, 90 (1):157–158, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237541.

Sorensen:1999:BRB

[3292] Lena Sorensen. Book review: Nurses’ Questions/Women’s Questions: The Impact of the Demographic Revolution and Feminism on the United States Working Women, 1946–1986 by Susan Rimby Leighow. Isis, 90 (1):158, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237542.

Jackson:1999:BRBa

[3293] John P. Jackson, Jr. Book review: Rebels within the Ranks: Psychologists’ Critique of Scientific Authority and Democratic Realities in New Deal America by Katherine Pandora. Isis, 90(1):158–159, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237543.

Gariepy:1999:BRB

[3294] Thomas P. Gariepy. Book review: Last Resort: Psychosurgery and the Limits of Medicine by Jack D. Pressman. Isis, 90(1):159–160, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237544.

Crenner:1999:BRB

[3295] Christopher Crenner. Book review: Surgery at New England Deaconess Hospital, 1896–1985, and Its Roots in the Harvard Surgical Service at the Boston City Hospital by William V. McDermott. Isis, 90(1):160– 161, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237545.

Sullivan:1999:BRB

[3296] Robert B. Sullivan. Book review: The Blues: A History of the Blue Cross and Blue Shield System by Robert Cunningham III; Robert M. Cunningham, Jr. Isis, 90(1):161, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237546.

Downes:1999:BRB

[3297] Stephen Downes. Book review: A Novel Defense of Scientific Realism by Jarrett Leplin. Isis, 90(1):161–162, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237547.

Goddu:1999:BRBa

[3298] Andre Goddu. Book review: Pierre Duhem et ses doctorands: Bibliographie de la litterature primaire et secondaire by Jean-François Stoffel. Isis, 90(1):162–163, March 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237548.

Buonora:1999:BRB

[3299] Paul Buonora. Book review: The History and Use of Our Earth’s Chemical Elements: A Reference Guide by Robert E. Krebs. Isis, 90(1):163– 164, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237549.

Kirsch:1999:BRB

[3300] David A. Kirsch. Book review: The Automobile: A Chronology of Its Antecedents, Development, and Impact by Clay McShane. Isis, 90(1): 164–165, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237550.

Anonymous:1999:Ca

[3301] Anonymous. Collections. Isis, 90(1):165–167, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237551.

Allen:1999:ESR

[3302] David Allen. Eloge: Susan Ruth Sheets-Pyenson, 9 September 1949–18 August 1998. Isis, 90(1):168–169, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237552.

Anonymous:1999:AMH

[3303] Anonymous. Annual meeting of the History of Science Society Kansas City, 21–25 October 1998. Isis, 90(1):170–174, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237553.

Anonymous:1999:BMc

[3304] Anonymous. Back matter. Isis, 90(1):175–180, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237554.

Anonymous:1999:FMc

[3305] Anonymous. Front matter. Isis, 90(1):??, March 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237471.

Conlin:1999:PSR

[3306] Michael F. Conlin. The popular and scientific reception of the Foucault pendulum in the United States. Isis, 90(2):181–204, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237048.

Dror:1999:AET

[3307] Otniel E. Dror. The affect of experiment: The turn to emotions in Anglo–American physiology, 1900–1940. Isis, 90(2):205–237, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237049.

Stamhuis:1999:HVH

[3308] Ida H. Stamhuis, Onno G. Meijer, and Erik J. A. Zevenhuizen. Hugo de Vries on heredity, 1889–1903: Statistics, Mendelian laws, pangenes, mutations. Isis, 90(2):238–267, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237050.

Hart:1999:RSS

[3309] Mitchell B. Hart. Racial science, social science, and the politics of Jewish assimilation. Isis, 90(2):268–297, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237051.

Cao:1999:CDB

[3310] Cong Cao. The changing dynamic between science and politics: Evolution of the highest academic honor in China, 1949–1998. Isis, 90(2): 298–324, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237052.

Clarke:1999:EMB

[3311] Adele E. Clarke. Eloge: Merriley E. Borell, 30 September 1945–27 August 1998. Isis, 90(2):325–326, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237053.

Hewlett:1999:EMG

[3312] Richard G. Hewlett. Eloge: Margaret Gowing, 26 April 1921–7 November 1998. Isis, 90(2):326–328, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237054.

Kohler:1999:BRB

[3313] Robert E. Kohler. Book review: Making Natural Knowledge: Constructivism and the History of Science. Isis, 90(2):329–331, June 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237055.

Greene:1999:BRB

[3314] Mott T. Greene. Book review: Born to Rebel: Birth Order, Family Dynamics, and Creative Lives. Isis, 90(2):332–338, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237056.

Dear:1999:BRB

[3315] Peter Dear. Book review: The Scientific Revolution: Aspirations and Achievements, 1500–1700 by James R. Jacob. Isis, 90(2):339, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237057.

Lindberg:1999:BRB

[3316] David C. Lindberg. Book review: The Bible, Protestantism, and the Rise of Natural Science by Peter Harrison. Isis, 90(2):339–341, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237058.

Apple:1999:BRB

[3317] Rima D. Apple. Book review: Women, Gender, and Science: New Directions by Sally Gregory Kohlstedt; Helen Longino. Isis, 90(2):341–342, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237059.

Grunden:1999:BRB

[3318] Walter E. Grunden. Book review: Science and Cultural Exchange in Modern History: Japan and the West by Watanabe Masao. Isis, 90(2): 342, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237060.

Wright:1999:BRB

[3319] David Wright. Book review: Science and Civilisation in China. Volume 7, Part 1: Language and Logic by Joseph Needham; Christoph Harbsmeier; Kenneth Robinson. Isis, 90(2):342–343, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237061.

Fara:1999:BRB

[3320] Patricia Fara. Book review: Time’s Pendulum: The Quest to Capture Time–From Sundials to Atomic Clocks by Jo Ellen Barnett. Isis, 90(2): 343–344, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237062.

Tiles:1999:BRB

[3321] Mary Tiles. Book review: What Is Mathematics, Really? by Reuben Hersh. Isis, 90(2):344–345, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237063.

Hunter:1999:BRB

[3322] Patti Wilger Hunter. Book review: Randomness by Deborah J. Bennett. Isis, 90(2):345–346, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237064.

Stephenson:1999:BRB

[3323] Bruce Stephenson. Book review: The Cambridge Illustrated History of Astronomy by Michael Hoskin. Isis, 90(2):346–348, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237065.

Marshall:1999:BRBa

[3324] Louise H. Marshall. Book review: Brain, Vision, Memory: Tales in the History of Neuroscience by Charles C. Gross. Isis, 90(2):348–349, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237066.

Mazur:1999:BRB

[3325] Allan Mazur. Book review: Pollution Markets in a Green Country Town: Urban Environmental Management in Transition by Roger K. Raufer. Isis, 90(2):349, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237067.

Foss-Mollan:1999:BRB

[3326] Kate Foss-Mollan. Book review: Building New York’s Sewers: Developing Mechanisms of Urban Management by Joanne Abel Goldman; Liberal Dreams and Nature’s Limits: Great Cities of North America since 1600 by James T. Lemon. Isis, 90(2):349–351, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237068.

Mendelsohn:1999:BRB

[3327] J. Andrew Mendelsohn. Book review: Die Geschichte der Infektionskrankheiten: Von der Antike bis ins 20. Jahrhundert by Karl-Heinz Leven. Isis, 90(2):351–352, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237069.

Meldrum:1999:BRB

[3328] Marcia Meldrum. Book review: The Powerful Placebo: From Ancient Priest to Modern Physician by Arthur K. Shapiro; Elaine Shapiro; The Placebo Effect: An Interdisciplinary Exploration by Anne Harrington. Isis, 90(2):352–353, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237070.

Petroski:1999:BRB

[3329] Henry Petroski. Book review: Structural Analysis: A Historical Approach by Jacques Heyman. Isis, 90(2):353–354, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237071.

Wallis:1999:BRB

[3330] Faith Wallis. Book review: Annus platonicus: A Study of World Cycles in Greek, Latin, and Arabic Sources by Godefroid de Callatay. Isis, 90 (2):354–355, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237072.

Isager:1999:BRB

[3331] Signe Isager. Book review: Hesiod’s Works and Days: A Translation and Commentary for the Social Sciences by Hesiod; David W. Tandy; Walter C. Neale. Isis, 90(2):355, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237073.

Nutton:1999:BRB

[3332] Vivian Nutton. Book review: La medicina dei Greci e dei Romani: Letteratura, lingua, scienza by Innocenzo Mazzini. Isis, 90(2):355–356, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237074.

Mark:1999:BRB

[3333] Sam Mark. Book review: The Children of Noah: Jewish Seafaring in Ancient Times by Raphael Patai; James Hornell; John M. Lundquist. Isis, 90(2):356–357, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237075.

Aveni:1999:BRB

[3334] Anthony Aveni. Book review: From Black Land to Fifth Sun: The Science of Sacred Sites by Brian Fagan. Isis, 90(2):357–358, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237076.

Takahashi:1999:BRB

[3335] Ken’ichi Takahashi. Book review: Roger Bacon and the Origins of Perspectiva in the Middle Ages: A Critical Edition and English Translation of Bacon’s Perspectiva, with Introduction and Notes by Roger Bacon; David C. Lindberg. Isis, 90(2):358–359, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237077.

Heyman:1999:BRB

[3336] Jacques Heyman. Book review: Technology and Resource Use in Medieval Europe: Cathedrals, Mills, and Mines by Elizabeth Bradford Smith; Michael Wolfe. Isis, 90(2):359–360, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237078.

Miller:1999:BRB

[3337] Timothy S. Miller. Book review: The Great Pox: The French Disease in Renaissance Europe by Jon Arrizabalaga; John Henderson; Roger French. Isis, 90(2):360–361, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237079.

Manders:1999:BRB

[3338] Kenneth Manders. Book review: Johann Faulhaber, 1580–1635: Eine Blutezeit der mathematischen Wissenschaften in Ulm by Kurt Hawlitschek. Isis, 90(2):361–362, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237080.

Ariew:1999:BRBa

[3339] Roger Ariew. Book review: Descartes: His Life and Thought by Genevieve Rodis-Lewis; Jane Marie Todd. Isis, 90(2):362–363, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237081.

Finocchiaro:1999:BRBa

[3340] Maurice A. Finocchiaro. Book review: La condamnation des livres Coperniciens et sa revocation à la lumière de documents inédits des Congregations de l’Index et de l’Inquisition by Pierre-Noel Mayaud. Isis, 90(2): 363–364, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237082.

Raeff:1999:BRB

[3341] Marc Raeff. Book review: Zeremonialwissenschaft im Furstenstaat: Studien zur juristischen und politischen Theorie absolutischer Herrschaftsreprasentation by Milos Vec. Isis, 90(2):364–365, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237083.

Palter:1999:BRB

[3342] Robert Palter. Book review: The Enlightenment of Joseph Priestley: A Study of His Life and Work from 1733 to 1773 by Robert E. Schofield. Isis, 90(2):365, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237084.

Marvin:1999:BRB

[3343] Ursula B. Marvin. Book review: Fire in the Sky: Comets and Meteors, the Decisive Centuries, in British Art and Science by Roberta J.

M. Olson; Jay M. Pasachoff. Isis, 90(2):365–366, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237085.

Ramsey:1999:BRB

[3344] Matthew Ramsey. Book review: Enlightenment and Pathology: Sensibility in the Literature and Medicine of Eighteenth-Century France by Anne C. Vila. Isis, 90(2):366–367, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237086.

Richman:1999:BRB

[3345] Allen Richman. Book review: Naval Surgeon: Life and Death at Sea in the Age of Sail by J. Worth Estes. Isis, 90(2):367–368, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237087.

Ackerman:1999:BRB

[3346] Evelyn Bernette Ackerman. Book review: The King’s Midwife: A History and Mystery of Madame du Coudray by Nina Rattner Gelbart. Isis, 90(2):368–369, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237088.

Rydell:1999:BRB

[3347] Robert W. Rydell. Book review: Science, Values, and the American West by Stephen Tchudi. Isis, 90(2):369–370, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237089.

Goetzmann:1999:BRB

[3348] William H. Goetzmann. Book review: Picturing Empire: Photography and the Visualization of the British Empire by James R. Ryan; Drawing Shadows to Stone: The Photography of the Jesup North Pacific Expedition, 1897–1902 by Laurel Kendall; Barbara Mathe; Thomas Ross Miller. Isis, 90(2):370–371, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237090.

Danahay:1999:BRB

[3349] Martin A. Danahay. Book review: J. M. W. Turner: Romantic Painter of the Industrial Revolution by William S. Rodner. Isis, 90(2):371–372, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237091.

Cassinet:1999:BRB

[3350] Jean Cassinet. Book review: Catalogo della corrispondenza di Paolo Ruffini by Francesco Barbieri; Franca Cattelani Degani. Isis, 90(2):372, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237092.

Slotten:1999:BRB

[3351] Hugh Richard Slotten. Book review: Joseph Henry: The Rise of an American Scientist by Albert E. Moyer. Isis, 90(2):372–373, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237093.

Baracca:1999:BRB

[3352] Angelo Baracca. Book review: Ludwig Boltzmann e la mecanica statistica by Carlo Cercignani. Isis, 90(2):373–374, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237094.

VanRiper:1999:BRBa

[3353] A. Bowdoin Van Riper. Book review: Geology and Religious Sentiment: The Effect of Geological Discoveries on English Society and Literature between 1829 and 1859 by J. M. I. Klaver. Isis, 90(2):374, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237095.

Eigen:1999:BRBb

[3354] Edward Eigen. Book review: Georges Cuvier, Fossil Bones, and Geological Catastrophes; New Translations and Interpretations of the Primary Texts by Martin J. S. Rudwick. Isis, 90(2):374–375, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237096.

Lederer:1999:BRB

[3355] Susan E. Lederer. Book review: Frankenstein’s Footsteps: Science, Genetics, and Popular Culture by Jon Turney. Isis, 90(2):375–376, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237097.

Cohn:1999:BRB

[3356] Helen M. Cohn. Book review: An All Consuming Passion: Origins, Modernity, and the Australian Life of Georgiana Molloy by William J. Lines. Isis, 90(2):376–377, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237098.

Vicinus:1999:BRB

[3357] Martha Vicinus. Book review: Bedside Seductions: Nursing and the Victorian Imagination, 1830–1880 by Catherine Judd. Isis, 90(2):377– 378, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237099.

Engstrom:1999:BRB

[3358] Eric J. Engstrom. Book review: Masters of Bedlam. The Transformation of the Mad-Doctoring Trade by Andrew Scull; Charlotte MacKenzie; Nicholas Hervey. Isis, 90(2):378–379, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237100.

Savitt:1999:BRB

[3359] Todd L. Savitt. Book review: Bones in the Basement: Postmortem Racism in Nineteenth-Century Medical Training by Robert L. Blakely; Judith M. Harrington. Isis, 90(2):379–380, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237101.

Helmstadter:1999:BRB

[3360] Axel Helmstadter. Book review: Von der physiologischen Chemie zur fruhen biochemischen Arzneimittelforschung: Der Apotheker und Chemiker Eugen Baumann (1846–1896) an den Universitäten Strassburg, Berlin, Freiburg, und in der pharmazeutischen Industrie by Beatrix Baumer. Isis, 90(2):380–381, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237102.

Addis:1999:BRB

[3361] Bill Addis. Book review: Building the Nineteenth Century by Tom F. Peters. Isis, 90(2):381–382, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237103.

Numbers:1999:BRB

[3362] Ronald L. Numbers. Book review: Redeeming Culture: American Religion in an Age of Science by James Gilbert. Isis, 90(2):382–383, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237104.

Noll:1999:BRB

[3363] Mark A. Noll. Book review: Summer for the Gods: The Scopes Trial and America’s Continuing Debate over Science and Religion by Edward J. Larson. Isis, 90(2):383–384, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237105.

Farrell:1999:BRB

[3364] James J. Farrell. Book review: Fallout: A Historian Reflects on America’s Half-Century Encounter with Nuclear Weapons by Paul Boyer. Isis, 90(2):384–385, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237106.

Kuznick:1999:BRB

[3365] Peter J. Kuznick. Book review: David E. Lilienthal: The Journey of an American Liberal by Steven M. Neuse. Isis, 90(2):385–386, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237107.

Wang:1999:BRBb

[3366] Zuoyue Wang. Book review: Endless Frontier: Vannevar Bush, Engineer of the American Century by G. Pascal Zachary. Isis, 90(2):386–387, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237108.

Cahan:1999:BRB

[3367] David Cahan. Book review: Institute im Bild. Part 1: Bauten der Kaiser-Wilhelm-Gesellschaft zur Forderung der Wissenschaften by Glenys Gill; Dagmar Klenke; Bibliographie zur Geschichte der Kaiser-Wilhelm-/Max-Planck-Gesellschaft zur Forderung der Wissenschaften (1911–1994) by Petra Hauke; Bibliometrische Profile von Instituten der Kaiser-Wilhelm-Gesellschaft zur Forderung der Wissenschaften (1923–1943): Institute der Chemisch-Physikalisch-Technischen und der Biologisch-Medizinischen Sektion by Heinrich Parthey; Quelleninventar Max Planck by Dirk Ullmann. Isis, 90(2):387–388, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237109.

Fleming:1999:BRB

[3368] Donald Fleming. Book review: Science at Oxford, 1914–1939: Transforming an Arts University by Jack Morrell. Isis, 90(2):388–389, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237110.

Dawson:1999:BRBa

[3369] John W. Dawson, Jr. Book review: Jacques Hadamard, a Universal Mathematician by Vladimir Maz’ya; Tatyana Shaposhnikova. Isis, 90 (2):389–390, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237111.

Feffer:1999:BRB

[3370] Loren Butler Feffer. Book review: Turing’s Legacy: A History of Computing at the National Physical Laboratory, 1945–1995 by David Yates. Isis, 90(2):390–391, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237112.

Mahoney:1999:BRB

[3371] Michael S. Mahoney. Book review: After Thought: The Computer Challenge to Human Intelligence by James Bailey. Isis, 90(2):391–392, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237113.

Heims:1999:BRB

[3372] Steve Joshua Heims. Book review: Talking Nets: An Oral History of Neural Networks by James A. Anderson; Edward Rosenfeld. Isis, 90(2): 392, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237114.

Enz:1999:BRB

[3373] Charles P. Enz. Book review: The Story of Spin by Sin-itiro Tomonaga; Takeshi Oka. Isis, 90(2):392–393, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237115.

Lagerkvist:1999:BRB

[3374] Ulf Lagerkvist. Book review: Francis Crick and James Watson and the Building Blocks of Life by Edward Edelson. Isis, 90(2):393–394, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237116.

Laubichler:1999:BRBa

[3375] Manfred D. Laubichler. Book review: Die Geschichte der genetisch orientierten Hirnforschung von Cecile und Oskar Vogt in der Zeit von 1895 bis ca. 1927 by Helga Satzinger. Isis, 90(2):394–395, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237117.

Nelkin:1999:BRB

[3376] Dorothy Nelkin. Book review: The Embryo Research Debate: Science and the Politics of Reproduction by Michael Mulkay. Isis, 90(2):395, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237118.

Pandora:1999:BRB

[3377] Katherine Pandora. Book review: Measuring Minds: Henry Herbert Goddard and the Origins of American Intelligence Testing by Leila Zenderland. Isis, 90(2):395–396, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237119.

Coutinho:1999:BRB

[3378] Marilia Coutinho. Book review: Historical Aspects of American Trypanosomiasis (Chagas’ Disease) by Matthias Perleth. Isis, 90(2):397, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237120.

Grob:1999:BRB

[3379] Gerald N. Grob. Book review: Making Us Crazy: DSM: The Psychiatric Bible and the Creation of Mental Disorders by Herb Kutchins; Stuart A. Kirk. Isis, 90(2):397–398, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237121.

Hale:1999:BRBa

[3380] Nathan G. Hale, Jr. Book review: The Emmanuel Movement (Boston, 1904–1929): The Origins of Group Treatment and the Assault on Lay Psychotherapy by Sanford Gifford. Isis, 90(2):398–399, June 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237122.

Ferguson:1999:BRB

[3381] Robert G. Ferguson. Book review: Turbulent Skies: The History of Commercial Aviation by T. A. Heppenheimer. Isis, 90(2):399–400, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237123.

Hansen:1999:BRBa

[3382] James R. Hansen. Book review: Wallops Station and the Creation of an American Space Program by Harold D. Wallace. Isis, 90(2):400–401, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237124.

Harford:1999:BRB

[3383] James Harford. Book review: Out from behind the Eight-Ball: A History of Project Echo by Donald C. Elder. Isis, 90(2):401–402, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237125.

Stranges:1999:BRB

[3384] Anthony N. Stranges. Book review: History and Overview of Solar Heat Technologies by Donald A. Beattie. Isis, 90(2):402–403, June 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237126.

Queisser:1999:BRB

[3385] Hans J. Queisser. Book review: Der “Transistor” als technisches und kulturelles Phanomen: Die Transistorisierung der Radio- und Fernsehempfanger in der Deutschen Rundfunkindustrie 1955 bis 1965 by Andreas Fickers. Isis, 90(2):403, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237127.

Marshall:1999:BRBb

[3386] Yvonne Marshall. Book review: Excavating Women: A History of Women in European Archaeology by Margarita Diaz-Andreu; Marie Louise Stig Sorensen. Isis, 90(2):403–404, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237128.

Osler:1999:BRB

[3387] Margaret J. Osler. Book review: The Cambridge History of Seventeenth-Century Philosophy by Daniel Garber; Michael Ayers. Isis, 90(2):404– 405, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237129.

Hinsley:1999:BRB

[3388] Curtis M. Hinsley. Book review: Against the Anthropological Grain by Wilcomb E. Washburn. Isis, 90(2):405–406, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237130.

Ghiselin:1999:BRB

[3389] Michael T. Ghiselin. Book review: Unifying Biology: The Evolutionary Synthesis and Evolutionary Biology by Vassiliki Betty Smocovitis. Isis, 90(2):406, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237131.

Maienschein:1999:BRB

[3390] Jane Maienschein. Book review: Unto Others: The Evolution and Psychology of Unselfish Behavior by Elliott Sober; David Sloan Wilson. Isis, 90(2):406–407, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237132.

Schroeder:1999:BRB

[3391] Ralph Schroeder. Book review: Max Weber’s Methodology: The Unification of the Cultural and Social Sciences by Fritz Ringer. Isis, 90(2): 407–408, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237133.

King:1999:BRB

[3392] William M. King. Book review: The Anatomy of Power: European Constructions of the African Body by Alexander Butchart. Isis, 90(2): 408–409, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237134.

Pratt:1999:BRB

[3393] Herbert T. Pratt. Book review: John Dalton, 1766–1844: A Bibliography of Works by and about Him with an Annotated List of His Surviving Apparatus and Personal Effects by A. L. Smyth. Isis, 90(2):409–410, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237135.

Huff:1999:BRB

[3394] Toby E. Huff. Book review: Encyclopaedia of the History of Science, Technology, and Medicine in Non-Western Cultures by Helaine Selin. Isis, 90(2):410–411, June 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237136.

Anonymous:1999:Cb

[3395] Anonymous. Collections. Isis, 90(2):411–413, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237137.

Anonymous:1999:BMd

[3396] Anonymous. Back matter. Isis, 90(2):414–420, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237138.

Anonymous:1999:FMd

[3397] Anonymous. Front matter. Isis, 90(2):??, June 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237047.

Smith:1999:STP

[3398] Pamela H. Smith. Science and taste: Painting, passions, and the new philosophy in Seventeenth-Century Leiden. Isis, 90(3):421–461, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237228.

Hopwood:1999:GBE

[3399] Nick Hopwood. “Giving body” to embryos: Modeling, mechanism, and the microtome in Late Nineteenth-Century anatomy. Isis, 90(3):462–496, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237229.

Withers:1999:RMT

[3400] Charles W. J. Withers. Reporting, mapping, trusting: Making geographical knowledge in the late Seventeenth Century. Isis, 90(3):497–521, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237230.

Zaitsev:1999:MEM

[3401] Evgeny A. Zaitsev. The meaning of Early Medieval geometry: From Euclid and surveyors’ manuals to Christian philosophy. Isis, 90(3):522– 553, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237231.

Westman:1999:EAF

[3402] Robert S. Westman. Eloge: Amos Funkenstein, 9 March 1937–11 November 1995. Isis, 90(3):554–557, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237232.

Larson:1999:BRB

[3403] Edward J. Larson. Book review: Teaching about Evolution and the Nature of Science. Isis, 90(3):558–559, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237233.

Webster:1999:BRB

[3404] Charles Webster. Book review: Wonders and the Order of Nature, 1150– 1750. Isis, 90(3):560–562, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237234.

Doel:1999:BRB

[3405] Ronald E. Doel. Book review: Nebulous Earth: The Origin of the Solar System and the Core of the Earth from Laplace to Jeffreys. Isis, 90(3): 563–568, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237235.

Garber:1999:BRB

[3406] Elizabeth Garber. Book review: To Light Such a Candle: Chapters in the History of Science and Technology by Keith J. Laidler. Isis, 90(3):569, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237236.

Hetherington:1999:BRB

[3407] Norriss S. Hetherington. Book review: The Ascent of Science by Brian L. Silver. Isis, 90(3):570, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237237.

Allison-Bunnell:1999:BRB

[3408] Steven Allison-Bunnell. Book review: Science in Public: Communication, Culture, and Credibility by Jane Gregory; Steve Miller. Isis, 90 (3):570–571, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237238.

Daston:1999:BRB

[3409] Lorraine Daston. Book review: The Footnote: A Curious History by Anthony Grafton. Isis, 90(3):571–572, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237239.

Farber:1999:BRB

[3410] Paul Lawrence Farber. Book review: Art and Science in America: Issues of Representation by Amy R. W. Meyers. Isis, 90(3):572–573, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237240.

Lestienne:1999:BRB

[3411] Remy Lestienne. Book review: Le temps by Harve Barreau. Isis, 90 (3):573–574, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237241.

Park:1999:BRB

[3412] David Park. Book review: Les figures de l’arc-en-ciel by Michel Blay. Isis, 90(3):574–575, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237242.

Greenberg:1999:BRB

[3413] Arthur Greenberg. Book review: The Atom in the History of Human Thought by Bernard Pullman; Axel Reisinger. Isis, 90(3):575–576, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237243.

Black:1999:BRB

[3414] Jeremy Black. Book review: Cartographic Encounters: Perspectives on Native American Mapmaking and Map Use by G. Malcolm Lewis. Isis, 90(3):576, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237244.

Ciancio:1999:BRB

[3415] Luca Ciancio. Book review: De la geologie a son histoire: Ouvrage edite en hommage a François Ellenberger by Gabriel Gohau. Isis, 90 (3):576–577, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237245.

Farley:1999:BRBb

[3416] John Farley. Book review: The Ovary of Eve: Egg and Sperm and Preformation by Clara Pinto-Correia. Isis, 90(3):577–578, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237246.

Luft:1999:BRB

[3417] Eric V. D. Luft. Book review: From Mesmer to Freud: Magnetic Sleep and the Roots of Psychological Healing by Adam Crabtree. Isis, 90(3): 578–579, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237247.

Bakhouche:1999:BRB

[3418] Beatrice Bakhouche. Book review: Images de la communion cosmique: Sur la terre comme au ciel by Robert Triomphe. Isis, 90(3):579–580, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237248.

Bensaude-Vincent:1999:BRB

[3419] Bernadette Bensaude-Vincent. Book review: The Substance of Civilization: Materials and Human History from the Stone Age to the Age of Silicon by Stephen L. Sass. Isis, 90(3):580, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237249.

Hardy:1999:BRB

[3420] Anne Hardy. Book review: The Burdens of Disease: Epidemics and Human Response in Western History by J. N. Hays. Isis, 90(3):580–581, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237250.

Stieb:1999:BRB

[3421] Ernst W. Stieb. Book review: The inside Story of Medicines: A Symposium by Gregory J. Higby; Elaine C. Stroud. Isis, 90(3):581–582, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237251.

Kidwell:1999:BRB

[3422] Peggy Aldrich Kidwell. Book review: Scientific Instruments, 1500–1900: An Introduction by Gerard L’E. Turner. Isis, 90(3):582–583, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237252.

Turner:1999:BRB

[3423] Gerard L’E. Turner. Book review: Western Astrolabes by Roderick Webster; Marjorie Webster. Isis, 90(3):583–584, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237253.

Coopersmith:1999:BRB

[3424] Jonathan Coopersmith. Book review: Consuming Power: A Social History of American Energies by David E. Nye. Isis, 90(3):584–585, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237254.

Rehbock:1999:BRB

[3425] Karen M. Rehbock. Book review: The Food of Paradise: Exploring Hawaii’s Culinary Heritage by Rachel Laudan. Isis, 90(3):585, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237255.

Sismondo:1999:BRB

[3426] Sergio Sismondo. Book review: Steward or Sorcerer’s Apprentice? The Evaluation of Technical Progress: A Systematic Overview of Theories and Opinions by Johan Hendrik Jacob van der Pot; Chris Turner. Isis, 90 (3):585–587, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237256.

Goldhill:1999:BRB

[3427] Simon Goldhill. Book review: Le sexe incertain: Androgynie et hermaphrodisme dans l’antiquite greco-romain by Luc Brisson. Isis, 90 (3):587, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237257.

Schiebinger:1999:BRB

[3428] Londa Schiebinger. Book review: Women, Science, and Medicine, 1500– 1700: Mothers and Sisters of the Royal Society by Lynette Hunter; Sarah Hutton. Isis, 90(3):587–589, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237258.

Goddu:1999:BRBb

[3429] Andre Goddu. Book review: Censure and Heresy at the University of Paris, 1200–1400 by J. M. M. H. Thijssen. Isis, 90(3):589, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237259.

Traister:1999:BRB

[3430] Barbara Traister. Book review: Lexicography and Physicke: The Record of Sixteenth-Century English Medical Terminology by R. W. McConchie. Isis, 90(3):589–590, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237260.

Rashed:1999:BRB

[3431] Marwan Rashed. Book review: The Liber Aristotilis of Hugo of Santalla by Charles Burnett; David Pingree. Isis, 90(3):590–591, September 1999.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237261.

vanDalen:1999:BRB

[3432] Benno van Dalen. Book review: Kusyar ibn Labban’s Introduction to Astrology by Michio Yano. Isis, 90(3):591–592, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237262.

Shank:1999:BRB

[3433] Michael H. Shank. Book review: Magic and Divination at the Courts of Burgundy and France: Text and Context of Laurens Pignon’s Contre les devineurs (1411) by Jan R. Veenstra. Isis, 90(3):592–593, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237263.

Palma:1999:BRB

[3434] Adriano Palma. Book review: Spinoza e il concetto della tradizione ebraica by Mino Chamla. Isis, 90(3):593–594, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237264.

Vickers:1999:BRBb

[3435] Brian Vickers. Book review: Objectivity in the Making: Francis Bacon and the Politics of Inquiry by Julie Robin Solomon. Isis, 90(3):594–595, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237265.

Lynch:1999:BRB

[3436] William T. Lynch. Book review: Galileo on the World Systems: A New Abridged Translation and Guide by Galileo Galilei; Maurice A. Finocchiaro. Isis, 90(3):595–596, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237266.

Finocchiaro:1999:BRBb

[3437] Maurice A. Finocchiaro. Book review: Galileo and the Church: Political Inquisition or Critical Dialogue? by Rivka Feldhay. Isis, 90(3):596–597, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237267.

Lennon:1999:BRB

[3438] Thomas M. Lennon. Book review: Treatise on the Human Mind by Louis de La Forge; Desmond M. Clarke. Isis, 90(3):597–598, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237268.

Riskin:1999:BRB

[3439] Jessica Riskin. Book review: Minerva’s Message: Stabilizing the French Revolution by Martin S. Staum. Isis, 90(3):598–599, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237269.

Gaukroger:1999:BRB

[3440] Stephen Gaukroger. Book review: La genesis de la mecanica de los fluidos (1640–1780) by Julian Simon Calero. Isis, 90(3):599–600, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237270.

Kiefer:1999:BRBb

[3441] Jurgen Kiefer. Book review: “Ich wollte, ich ware ein guter Schuhflicker… ”: Das ungluckliche Leben des bayerischen Astronomen Johann Nepomuck Fischer (1749–1805): Eine dokumentarische Biographie by Wolfgang Hernschier. Isis, 90(3):600, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237271.

Glick:1999:BRB

[3442] Thomas F. Glick. Book review: El Atlas zoologico, el megaterio y las tecnicas de pesca valencianas, 1742–1799 by Juan Bautista Bru de Ramon; Jose M. Lopez Piñero. Isis, 90(3):600–601, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237272.

Hacker:1999:BRB

[3443] Barton C. Hacker. Book review: Engineering the Revolution: Arms and Enlightenment in France, 1763–1815 by Ken Alder. Isis, 90(3):601, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237273.

Pastrana:1999:BRB

[3444] Patricia Aceves Pastrana. Book review: Catastrofe morboso de las minas mercuriales de la villa de Almaden del Azogue (1778) de Jose Pares y Franques by Alfredo Menendez Navarro. Isis, 90(3):601–602, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237274.

Archer:1999:BRB

[3445] Christon I. Archer. Book review: Barberos, cirujanos y gente de mar: La sanidad naval y la profesion quirurgica en la España ilustrada by Mikel Astrain Gallart. Isis, 90(3):602–603, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237275.

Schubring:1999:BRB

[3446] Gert Schubring. Book review: Joseph-Diez Gergonne (1771–1859): Histoire [logique] et philosophie des sciences by Mario H. Otero. Isis, 90 (3):603–604, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237276.

Brock:1999:BRB

[3447] William H. Brock. Book review: Friedrich Julius Otto (1809–1870): Pharmazeut, Chemiker, Technologe, Gesundheitsbeamter und das Collegium Carolinum in Braunschweig by Ursula Pohl. Isis, 90(3):604–605, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237277.

Eastman:1999:BRB

[3448] Joel W. Eastman. Book review: Common Lands, Common People: The Origins of Conservation in Northern New England by Richard W. Judd. Isis, 90(3):605–606, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237278.

Moore:1999:BRB

[3449] James Moore. Book review: Darwinian Myths: The Legends and Misuses of a Theory by Edward Caudill. Isis, 90(3):606–607, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237279.

Nyhart:1999:BRB

[3450] Lynn K. Nyhart. Book review: Life’s Splendid Drama: Evolutionary Biology and the Reconstruction of Life’s Ancestry, 1860–1940 by Peter J. Bowler. Isis, 90(3):607, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237280.

Camerini:1999:BRB

[3451] Jane R. Camerini. Book review: A Passion for Birds: American Ornithology after Audubon by Mark V. Barrow. Isis, 90(3):607–608, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237281.

Rosenberg:1999:BRB

[3452] Rosalind Rosenberg. Book review: Gender and American Social Science: The Formative Years by Helene Silverberg. Isis, 90(3):608–609, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237282.

Cooter:1999:BRB

[3453] Roger Cooter. Book review: A Measure of Perfection: Phrenology and the Fine Arts in America by Charles Colbert. Isis, 90(3):609–611, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237283.

Gorman:1999:BRB

[3454] Michael E. Gorman. Book review: Alexander Graham Bell: Making Connections by Naomi Pasachoff. Isis, 90(3):611–612, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237284.

Matejuk:1999:BRB

[3455] Piotr Matejuk. Book review: Dawne lunety i lornetki w zbiorach polskich by Antoni M. Piaskowski. Isis, 90(3):612, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237285.

Kilburn:1999:BRB

[3456] Kevin J. Kilburn. Book review: Victorian Telescope Makers: The Lives and Letters of Thomas and Howard Grubb by I. S. Glass. Isis, 90(3): 612–613, September 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237286.

Mawhin:1999:BRB

[3457] Jean Mawhin. Book review: Convenzione e razionalita scientifica in Henri Poincaré by Mirella Fortino. Isis, 90(3):613–614, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237287.

Montiel:1999:BRB

[3458] Luis Montiel. Book review: Carl Vogt (1817–1895): Science, philosophie et politique by Jean-Claude Pont; Daniele Bui; Françoise Dubosson; Jan Lacki. Isis, 90(3):614–615, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237288.

Walker:1999:BRB

[3459] Mark Walker. Book review: Planet Dora: A Memoir of the Holocaust and the Birth of the Space Age by Yves Beon; Michael J. Neufeld; Richard L. Fague. Isis, 90(3):615, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237289.

Bergia:1999:BRB

[3460] Silvio Bergia. Book review: Mgr. Georges Lemaître, savant et croyant: Actes du colloque tenu a Louvain-la-Neuve le 4 novembre 1994; La physique d’Einstein: Texte inedit de Georges Lemaître by Jean-François Stoffel. Isis, 90(3):615–616, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237290.

Kragh:1999:BRB

[3461] Helge Kragh. Book review: Masters of Modern Physics: The Scientific Contributions of H. A. Kramers by D. ter Haar. Isis, 90(3):616–617, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237291.

Velarde:1999:BRB

[3462] Manuel G. Velarde. Book review: The Life and Legacy of G. I. Taylor by G. K. Batchelor. Isis, 90(3):617–618, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237292.

Coletta:1999:BRB

[3463] Paolo E. Coletta. Book review: The Golden Age Remembered: U.S. Naval Aviation, 1919–1941 by E. T. Wooldridge. Isis, 90(3):618–619, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237293.

Cambrosio:1999:BRBa

[3464] Alberto Cambrosio. Book review: Molecularizing Biology and Medicine: New Practices and Alliances, 1910s-1970s by Soraya de Chadarevian; Harmke Kamminga. Isis, 90(3):619–620, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237294.

deChadarevian:1999:BRB

[3465] Soraya de Chadarevian. Book review: A History of Molecular Biology by Michel Morange; Matthew Cobb. Isis, 90(3):620–621, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237295.

Burlingame:1999:BRB

[3466] Leslie J. Burlingame. Book review: The Idea of Biodiversity: Philosophies of Paradise by David Takacs. Isis, 90(3):621–622, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237296.

Laubichler:1999:BRBb

[3467] Manfred D. Laubichler. Book review: The Camel’s Nose: Memoirs of a Curious Scientist by Knut Schmidt-Nielsen. Isis, 90(3):622–624, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237297.

Pols:1999:BRB

[3468] Hans Pols. Book review: The Wing of Madness: The Life and Work of R. D. Laing by Daniel Burston. Isis, 90(3):624–625, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237298.

Leiber:1999:BRB

[3469] Justin Leiber. Book review: Naming the Mind: How Psychology Found Its Language by Kurt Danziger. Isis, 90(3):625–626, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237299.

Nader:1999:BRB

[3470] Laura Nader. Book review: After the Fact: Two Countries, Four Decades, One Anthropologist by Clifford Geertz. Isis, 90(3):626–627, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237300.

VanRiper:1999:BRBb

[3471] A. Bowdoin Van Riper. Book review: Forbidden Archaeology’s Impact: How a Controversial Book Shocked the Scientific Community and Became an Underground Classic by Michael A. Cremo. Isis, 90(3):627, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237301.

Clarke:1999:BRB

[3472] Lee Clarke. Book review: A Hazardous Inquiry: The Rashomon Effect at Love Canal by Allan Mazur. Isis, 90(3):627–628, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237302.

Dodsworth:1999:BRB

[3473] Helen Dodsworth. Book review: Intrinsic Factors: William Bosworth Castle and the Development of Hematology and Clinical Investigation at Boston City Hospital by Anand B. Karnad. Isis, 90(3):628–629, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237303.

Schull:1999:BRB

[3474] William J. Schull. Book review: Radium Girls: Women and Industrial Health Reform, 1910–1935 by Claudia Clark. Isis, 90(3):629–630, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237304.

Arrizabalaga:1999:BRB

[3475] Jon Arrizabalaga. Book review: Curar el cancer? Los origenes de la radioterapia española en el primer tercio del siglo XX by Rosa Maria Medina Domenech. Isis, 90(3):630–631, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237305.

Pernick:1999:BRB

[3476] Martin S. Pernick. Book review: Venus Envy: A History of Cosmetic Surgery by Elizabeth Haiken. Isis, 90(3):631–632, September 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237306.

Collins:1999:BRB

[3477] Martin Collins. Book review: Dr. Strangelove’s America: Society and Culture in the Atomic Age by Margot A. Henriksen. Isis, 90(3):632–633, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237307.

Roell:1999:BRB

[3478] Craig H. Roell. Book review: Steinway and Sons by Richard K. Lieberman. Isis, 90(3):633, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237308.

Ceruzzi:1999:BRB

[3479] Paul E. Ceruzzi. Book review: Electronic Genie: The Tangled History of Silicon by Frederick Seitz; Norman G. Einspruch. Isis, 90(3):633–634, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237309.

Vinarcik:1999:BRB

[3480] Edward J. Vinarcik. Book review: Wings of Wood, Wings of Metal: Cultural and Technical Choice in American Airplane Materials, 1914– 1945 by Eric Schatzberg. Isis, 90(3):634–635, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237310.

Dawson:1999:BRBb

[3481] Virginia P. Dawson. Book review: Space and the American Imagination by Howard E. McCurdy. Isis, 90(3):635–636, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237311.

Crouch:1999:BRB

[3482] Tom D. Crouch. Book review: Wingless Flight: The Lifting Body Story by R. Dale Reed; Darlene Lister. Isis, 90(3):636–637, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237312.

Lucier:1999:BRB

[3483] Paul Lucier. Book review: Oil Baron of the Southwest: Edward L. Doheny and the Development of the Petroleum Industry in California and Mexico by Martin R. Ansell. Isis, 90(3):637–638, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237313.

Lyne:1999:BRB

[3484] John Lyne. Book review: The Rhetoric of Science by Alan G. Gross. Isis, 90(3):638–639, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237314.

Motterlini:1999:BRB

[3485] Matteo Motterlini. Book review: The Myth of the Framework: In Defence of Science and Rationality by Karl R. Popper; M. A. Notturno; Knowledge and the Body-Mind Problem: In Defence of Interaction by Karl R. Popper; M. A. Notturno. Isis, 90(3):639–641, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237315.

Koetsier:1999:BRB

[3486] Teun Koetsier. Book review: Lakatos: An Introduction by Brendan Larvor. Isis, 90(3):641, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237316.

Selin:1999:BRB

[3487] Helaine Selin. Book review: The Cambridge Dictionary of Scientists by David Millar; Ian Millar; Jon Millar; Margaret Millar. Isis, 90(3):641, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237317.

Elliott:1999:BRB

[3488] Clark A. Elliott. Book review: Biographies of Scientists: An Annotated Bibliography by Roger Smith. Isis, 90(3):641–642, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237318.

Herzenberg:1999:BRB

[3489] Caroline L. Herzenberg. Book review: Contributions of Twentieth– Century Women to Physics. Isis, 90(3):642–643, September 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237319.

Subbarayappa:1999:BRB

[3490] B. V. Subbarayappa. Book review: A Handlist of Sanskrit and Prakrit Manuscripts in the Library of the Wellcome Institute for the History of Medicine. Volume 2 by Dominic Wujastyk. Isis, 90(3):643–644, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237320.

McConchie:1999:BRB

[3491] R. W. McConchie. Book review: Names of Body Parts in English: 1400–1550 by Juhani Norri. Isis, 90(3):644, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237321.

Anonymous:1999:Cc

[3492] Anonymous. Collections. Isis, 90(3):645, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237322.

Anonymous:1999:BMe

[3493] Anonymous. Back matter. Isis, 90(3):646–652, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237323.

Anonymous:1999:FMe

[3494] Anonymous. Front matter. Isis, 90(3):??, September 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237227.

Lorcin:1999:ICI

[3495] Patricia M. E. Lorcin. Imperialism, colonial identity, and race in Algeria, 1830–1870: The role of the French Medical Corps. Isis, 90(4):653–679, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237655.

Roberts:1999:SBE

[3496] Lissa Roberts. Science becomes electric: Dutch interaction with the electrical machine during the Eighteenth Century. Isis, 90(4):680–714, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237656.

Sarasohn:1999:THD

[3497] Lisa T. Sarasohn. Thomas Hobbes and the Duke of Newcastle: A study in the mutuality of patronage before the establishment of the Royal Society. Isis, 90(4):715–737, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237657.

Olwell:1999:PIM

[3498] Russell Olwell. Physical isolation and marginalization in physics: David Bohm’s Cold War exile. Isis, 90(4):738–756, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237658.

Topper:1999:GSM

[3499] David Topper. Galileo, sunspots, and the motions of the Earth: Redux. Isis, 90(4):757–767, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/ 237659. See [8].

Lankford:1999:EHW

[3500] John Lankford. Eloge: Helen Wright, 1914–1997. Isis, 90(4):768–769, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237660.

Bondeson:1999:LE

[3501] Jan Bondeson, Dennis Todd, William C. Waterhouse, Peter Pesic, and Peter Dear. Letters to the Editor. Isis, 90(4):770–771, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237661.

Baldwin:1999:BRB

[3502] Martha Baldwin. Book review: The Aspiring Adept: Robert Boyle and His Alchemical Quest: Including Boyle’s “Lost” Dialogue on the Transmutation of Metals. Isis, 90(4):772–774, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237662.

Burnett:1999:BRB

[3503] D. Graham Burnett. Book review: Maps and Politics. Isis, 90(4):775– 780, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237663.

Jasanoff:1999:BRB

[3504] Sheila Jasanoff. Book review: The Baltimore Case: A Trial of Politics, Science, and Character. Isis, 90(4):781–783, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237664.

Tobias:1999:BRB

[3505] Sheila Tobias. Book review: Common Science? Women, Science, and Knowledge by Jean Barr; Lynda Birke. Isis, 90(4):784, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237665.

Marchand:1999:BRB

[3506] Suzanne Marchand. Book review: The Cambridge Illustrated History of Archaeology by Paul G. Bahn. Isis, 90(4):784–786, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237666.

Moreton:1999:BRB

[3507] Jennifer Moreton. Book review: Time’s Alteration: Calendar Reform in Early Modern England by Robert Poole. Isis, 90(4):786–787, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237667.

Gordon:1999:BRB

[3508] Robert B. Gordon. Book review: Traces of the Past: Unraveling the Secrets of Archaeology through Chemistry by Joseph B. Lambert. Isis, 90(4):787, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237668.

Pyenson:1999:BRB

[3509] Lewis Pyenson. Book review: Le secret de l’Occident: Du miracle passé au marasme présent by David Cosandey. Isis, 90(4):787–788, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237669.

Perkins:1999:BRB

[3510] John H. Perkins. Book review: The Living Fields: Our Agricultural Heritage by Jack R. Harlan. Isis, 90(4):788–789, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237670.

Hankins:1999:BRB

[3511] Thomas L. Hankins. Book review: Experiencing Nature: Proceedings of a Conference in Honor of Allen G. Debus by Paul H. Theerman; Karen Hunger Parshall. Isis, 90(4):789–790, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237671.

Benson:1999:BRB

[3512] Keith R. Benson. Book review: Natural Causes: Essays in Ecological Marxism by James O’Connor. Isis, 90(4):790, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237672.

Noble:1999:BRB

[3513] David W. Noble. Book review: Technology and the West: A Historical Anthology from Technology and Culture by Terry S. Reynolds; Stephen H. Cutcliffe; Technology and American History: A Historical Anthology from Technology and Culture by Stephen H. Cutcliffe; Terry S. Reynolds. Isis, 90(4):790–792, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237673.

McCluskey:1999:BRB

[3514] Stephen C. McCluskey. Book review: A Chronicle of Pre-Telescopic Astronomy by Barry Hetherington. Isis, 90(4):792–793, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237674.

Mueller:1999:BRB

[3515] Milton Mueller. Book review: Media Technology and Society: A History: From the Telegraph to the Internet by Brian Winston. Isis, 90(4):793– 794, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237675.

Carpenter:1999:BRB

[3516] Kenneth J. Carpenter. Book review: Gout: The Patrician Malady by Roy Porter; G. S. Rousseau. Isis, 90(4):794–795, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237676.

Hauger:1999:BRB

[3517] J. Scott Hauger. Book review: A Natural History of Vision by Nicholas J. Wade. Isis, 90(4):795–796, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237677.

Tjossem:1999:BRB

[3518] Sara Tjossem. Book review: Energies: An Illustrated Guide to the Biosphere and Civilization by Vaclav Smil. Isis, 90(4):796–797, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237678.

deMontellano:1999:BRB

[3519] Bernard R. Ortiz de Montellano. Book review: Stairways to the Stars: Sky-watching in Three Great Ancient Cultures by Anthony Aveni. Isis, 90 (4):797–798, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237679.

Brack-Bernsen:1999:BRB

[3520] Lis Brack-Bernsen. Book review: The Babylonian Theory of the Planets by Noel M. Swerdlow. Isis, 90(4):798–799, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237680.

Risse:1999:BRB

[3521] Guenter B. Risse. Book review: Ancient Egyptian Medicine by John F. Nunn. Isis, 90(4):800, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237681.

Findlen:1999:BRB

[3522] Paula Findlen. Book review: Knowledge, Discovery, and Imagination in Early Modern Europe: The Rise of Aesthetic Rationalism by Timothy J. Reiss. Isis, 90(4):800–801, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237682.

VanEgmond:1999:BRB

[3523] Warren Van Egmond. Book review: Summa de l’art d’aritmetica by Francesc Santcliment; Antoni Malet. Isis, 90(4):801–802, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237683.

Jones:1999:BRBa

[3524] Alexander Jones. Book review: George Gemistus Plethon: Manuel d’astronomie by Anne Tihon; Raymond Mercier. Isis, 90(4):802–803, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237684.

Zonta:1999:BRB

[3525] Mauro Zonta. Book review: Arab Roots of Gemology: Ahmad ibn Yusuf al Tifashi’s Best Thoughts on the Best of Stones by Ahmad ibn Yusuf al-Tifashi; Samar Najm Abul Huda. Isis, 90(4):803, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237685.

Shapiro:1999:BRB

[3526] Alan E. Shapiro. Book review: Archives of the Scientific Revolution: The Formation and Exchange of Ideas in Seventeenth-Century Europe by Michael Hunter. Isis, 90(4):804, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237686.

Ariew:1999:BRBb

[3527] Roger Ariew. Book review: Oeuvres de Descartes by Rene Descartes; Charles Adam; Paul Tannery; Principia philosophiae by Rene Descartes; Meditationes de prima philosophia by Rene Descartes. Isis, 90(4):804– 806, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237687.

Garrard:1999:BRB

[3528] Graeme Garrard. Book review: An Examination of the Philosophy of Bacon: Wherein Different Questions of Rational Philosophy Are Treated by Joseph de Maistre; Richard A. Lebrun. Isis, 90(4):806, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237688.

Fletcher:1999:BRB

[3529] Colin R. Fletcher. Book review: Fermat’s Enigma: The Epic Quest to Solve the World’s Greatest Mathematical Problem by Simon Singh. Isis, 90(4):806–807, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237689.

Meo:1999:BRB

[3530] M. Meo. Book review: Isaac Newton and the Scientific Revolution by Gale E. Christianson. Isis, 90(4):807–808, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237690.

Frijhoff:1999:BRB

[3531] Willem Frijhoff. Book review: De signaturis internis rerum by Oswaldus Crollius; Wilhelm Kuhlmann; Joachim Telle; Alchemomedizinische Briefe 1585 bis 1597 by Oswaldus Crollius; Wilhelm Kuhlmann; Joachim Telle. Isis, 90(4):808–809, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237691.

Sarasohn:1999:BRB

[3532] Lisa T. Sarasohn. Book review: Margaret Cavendish and the Exiles of the Mind by Anna Battigelli. Isis, 90(4):809–810, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237692.

Watkins:1999:BRB

[3533] Eric Watkins. Book review: The Skeptical Tradition around 1800: Skepticism in Philosophy, Science, and Society by Johan van der Zande; Richard Popkin. Isis, 90(4):810–811, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237693.

Fraser:1999:BRB

[3534] Craig Fraser. Book review: Opera Omnia: Series Quarta A. Volume 2: Briefwechsel mit Johann (I) Bernoulli und Nikolaus (I) Bernoulli: Commercium Epistolicum by Leonhard Euler; E. A. Fellmann; G. K. Mikhajlov. Isis, 90(4):811–812, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237694.

Schneider:1999:BRB

[3535] Ivo Schneider. Book review: Die Wurzen der Wahrscheinlichkeitsrechnung. Die Verbindung von Glucksspieltheorie und statistischer Praxis vor Laplace by Walter Hauser. Isis, 90(4):812–813, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237695.

Guerrini:1999:BRBb

[3536] Anita Guerrini. Book review: The Life Sciences in Eighteenth-Century French Thought by Jacques Roger; Keith R. Benson; Robert Ellrich. Isis, 90(4):813–814, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237696.

Bourque:1999:BRBb

[3537] Monique Bourque. Book review: Public Culture in the Early Republic: Peale’s Museum and Its Audience by David R. Brigham. Isis, 90(4): 814–815, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237697.

Holmes:1999:BRB

[3538] Frederic L. Holmes. Book review: De la situation du Tresor public au ler juin 1791: Par les Commissaires de la Tresorerie Nationale Condorcet, Lavoisier, de Vaines, Dutremblay, Rouille de l’Etang Cornut de la Fontaine by Jean–Pierre Poirier. Isis, 90(4):815–816, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237698.

Oldroyd:1999:BRB

[3539] David Oldroyd. Book review: Science in the Romantic Era by David M. Knight. Isis, 90(4):816–817, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237699.

Williams:1999:BRB

[3540] L. Pearce Williams. Book review: Science in the Making: Scientific Development as Chronicled by Historic Papers in the Philosophical Magazine: With Commentaries and Illustrations. Volume 2: 1851–1900 by E. A. Davis. Isis, 90(4):817, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237700.

Ross:1999:BRB

[3541] Sydney Ross. Book review: A Calendar of the Correspondence of Sir John Herschel by Michael J. Crowe; David R. Dyck; James J. Kevin. Isis, 90(4):817–818, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237701.

LeGall:1999:BRB

[3542] Philippe Le Gall. Book review: Bibliographie Cournotienne by Thierry Martin; Jean-Philippe Massonie. Isis, 90(4):818–819, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237702.

Caneva:1999:BRB

[3543] Kenneth L. Caneva. Book review: Selected Scientific Works of Hans Christian Orsted by Hans Christian Orsted; Karen Jelved; Andrew D. Jackson; Ole Knudsen. Isis, 90(4):819–820, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237703.

Slater:1999:BRB

[3544] Leo Slater and David C. Brock. Book review: The Making of the Chemist: The Social History of Chemistry in Europe, 1789–1914 by David Knight; Helge Kragh. Isis, 90(4):820–821, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237704.

Beretta:1999:BRBb

[3545] Marco Beretta. Book review: La “Filosofia della chimica”: Un mito scientista nella Toscana di inizio Ottocento by Alessandro Volpi. Isis, 90(4):821, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237705.

Kim:1999:BRB

[3546] Mi Gyung Kim. Book review: August Friedrich Horstmann und die physikalische Chemie by Alexander Kipnis. Isis, 90(4):821–822, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237706.

Cassidy:1999:BRB

[3547] James G. Cassidy. Book review: Joseph Leidy: The Last Man Who Knew Everything by Leonard Warren. Isis, 90(4):822–823, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237707.

Overfield:1999:BRB

[3548] Richard A. Overfield. Book review: Pioneer Naturalist on the Plains: The Diary of Elam Bartholomew, 1871 to 1934 by David M. Bartholomew. Isis, 90(4):823–824, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/237708.

Bowler:1999:BRB

[3549] Peter J. Bowler. Book review: Charles Darwin and the Evolution Revolution by Rebecca Stefoff. Isis, 90(4):824–825, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237709.

Smout:1999:BRB

[3550] Kary Doyle Smout. Book review: Darwinism Comes to America by Ronald L. Numbers. Isis, 90(4):825–826, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:/

/www.jstor.org/stable/237710.

Illi:1999:BRB

[3551] Jorg Illi. Book review: Ferdinand von Muller, 1825–1896: Spuren eines australischen Botanikers in seiner Geburtsstadt Rostock. Isis, 90(4):826, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237711.

Schroeder-Gudehus:1999:BRB

[3552] Brigitte Schroeder-Gudehus. Book review: Nationalismus und internationale Zusammenarbeit in den Naturwissenschaften: Die Deutsch-Franzosischen Wissenschaftsbeziehungen zwischen Revolution und Restauration, 1789–1832 by Kai Torsten Kanz. Isis, 90(4):826–827, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237712.

Deacon:1999:BRB

[3553] Desley Deacon. Book review: The Women Founders of the Social Sciences by Lynn McDonald. Isis, 90(4):827–828, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237713.

Metaxas:1999:BRB

[3554] Virginia A. Metaxas. Book review: A Social History of Wet Nursing in America: From Breast to Bottle by Janet Golden. Isis, 90(4):828–829, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237714.

Becquemont:1999:BRB

[3555] Daniel Becquemont. Book review: Spencer et l’evolutionnisme philosophique by Patrick Tort. Isis, 90(4):829–830, December 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237715.

Rupp-Eisenreich:1999:BRB

[3556] Britta Rupp-Eisenreich. Book review: Lecas Spencer: Religion, science, et politique by Daniel Becquemont; Laurent Mucchielli. Isis, 90(4):830– 831, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237716.

Jacyna:1999:BRBb

[3557] Stephen Jacyna. Book review: The Eclectic Legacy: Academic Philosophy and the Human Sciences in Nineteenth-Century France by John I. Brooks. Isis, 90(4):831, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237717.

Divine:1999:BRB

[3558] Robert A. Divine. Book review: The Politics and Technology of Nuclear Proliferation by Robert F. Mozley. Isis, 90(4):832, December 1999.

CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237718.

Higham:1999:BRB

[3559] Robin Higham. Book review: The Science of War: Canadian Scientists and Allied Military Technology during the Second World War by Donald H. Avery. Isis, 90(4):832–833, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237719.

Macrakis:1999:BRB

[3560] Kristie Macrakis. Book review: Stalinist Science by Nikolai Krementsov. Isis, 90(4):833–834, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237720.

Kleinman:1999:BRB

[3561] Daniel Lee Kleinman. Book review: Pasteur’s Quadrant: Basic Science and Technological Innovation by Donald E. Stokes. Isis, 90(4):834–835, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237721.

Cushing:1999:BRB

[3562] James T. Cushing. Book review: The Rise of the Standard Model: Particle Physics in the 1960s and 1970s by Lillian Hoddeson; Laurie Brown; Michael Riordan; Max Dresden. Isis, 90(4):835, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237722.

Carson:1999:BRB

[3563] Cathryn Carson. Book review: Heisenberg and the Nazi Atomic Bomb Project: A Study in German Culture by Paul Lawrence Rose. Isis, 90 (4):835–836, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237723.

Nathan:1999:BRB

[3564] Ove Nathan. Book review: Niels Bohr: Memoirs of a Working Relationship by Stefan Rozental. Isis, 90(4):836–837, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237724.

Hughes:1999:BRB

[3565] Jeff Hughes. Book review: Nevill Mott: Reminiscences and Appreciations by E. A. Davis. Isis, 90(4):837–838, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237725.

Gavroglu:1999:BRB

[3566] Kostas Gavroglu. Book review: After the Break-Through: The Emergence of High-Temperature Superconductivity as a Research Field by Helga Nowotny; Ulrike Felt. Isis, 90(4):838–839, December 1999. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237726.

Sime:1999:BRB

[3567] Ruth Lewin Sime. Book review: Wollen wir unsere Hande in Unschuld waschen? Gertrud Woker (1878–1968), Chemikerin und Internationale Frauenliga (1915–1968) by Gerit von Leitner. Isis, 90(4):839–840, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237727.

Johnson:1999:BRBb

[3568] Jeffrey Allan Johnson. Book review: Fritz Haber, 1868–1934: Eine Biographie by Margit Szollosi-Janze. Isis, 90(4):840, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237728.

Steen:1999:BRB

[3569] Kathryn Steen. Book review: Inventing Polymer Science: Staudinger, Carothers, and the Emergence of Macromolecular Chemistry by Yasu Furukawa. Isis, 90(4):841–842, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237729.

Satzinger:1999:BRB

[3570] Helga Satzinger. Book review: Otto Warburgs Beitrag zur Atmungstheorie: Das Problem der Sauerstoffaktivierung by Petra Werner. Isis, 90 (4):842–843, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237730.

Rainger:1999:BRB

[3571] Ronald Rainger. Book review: Charles Doolittle Walcott: Paleontologist by Ellis L. Yochelson. Isis, 90(4):843–844, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237731.

Yochelson:1999:BRB

[3572] Ellis L. Yochelson. Book review: Dinosaur Impressions: Postcards from a Paleontologist by Philippe Taquet; Kevin Padian. Isis, 90(4):844, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (elec-

tronic). URL http://www.jstor.org/stable/237732.

Cambrosio:1999:BRBb

[3573] Alberto Cambrosio. Book review: Correcting the Blueprint of Life: An Historical Account of the Discovery of DNA Repair Mechanisms by Errol C. Friedberg. Isis, 90(4):844–845, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237733.

Wang:1999:BRBc

[3574] Zuoyue Wang and Jedidiah Kroncke. Book review: China Voyager: Gist Gee’s Life in Science by William J. Haas. Isis, 90(4):845–846, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/237734.

Sengoopta:1999:BRB

[3575] Chandak Sengoopta. Book review: Hitler: Diagnosis of a Destructive Prophet by Fritz Redlich. Isis, 90(4):846–847, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237735.

Thayer:1999:BRB

[3576] Earl R. Thayer. Book review: The New York Academy of Medicine, 1947–1997: Enhancing the Health of the Public by Marvin Lieberman; Leon J. Warshaw. Isis, 90(4):847–848, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237736.

Castillo:1999:BRB

[3577] Edward D. Castillo. Book review: White Man’s Medicine: The Navajo and Government Doctors, 1863–1955 by Robert A. Trennert. Isis, 90 (4):848–849, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237737.

Koch:1999:BRB

[3578] Ellen Koch. Book review: Beginnings Count: The Technological Imperative in American Health Care by David J. Rothman. Isis, 90(4):849–850, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237738.

Jones:1999:BRBb

[3579] Kathleen W. Jones. Book review: Mental Ills and Bodily Cures: Psychiatric Treatment in the First Half of the Twentieth Century by Joel Braslow. Isis, 90(4):850–851, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237739.

Jones:1999:BRBc

[3580] Greta Jones. Book review: The Politics of Heredity: Essays on Eugenics, Biomedicine, and the Nature-Nurture Debate by Diane B. Paul. Isis, 90 (4):851–852, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237740.

Blaisdell:1999:BRB

[3581] John D. Blaisdell. Book review: Mad Dogs: The New Rabies Plague by Don Finley. Isis, 90(4):852, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/237741.

Hale:1999:BRBb

[3582] Nathan G. Hale, Jr. Book review: Freud and the Bolsheviks: Psychoanalysis in Imperial Russia and the Soviet Union by Martin A. Miller. Isis, 90 (4):852–853, December 1999. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/237742.

Coser:1999:BRB

[3583] Lewis A. Coser. Book review: Paul Lazarsfeld (1901–1976): La sociologie de Vienne a New York by Jacques Lautman; Bernard-Pierre Lecuyer. Isis, 90(4):853–854, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237743.

Prasch:1999:BRB

[3584] Robert E. Prasch. Book review: The Money Interest and the Public Interest: American Monetary Thought, 1920–1970 by Perry G. Mehrling. Isis, 90(4):854–855, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237744.

Jackson:1999:BRBb

[3585] John P. Jackson, Jr. Book review: From Savage to Negro: Anthropology and the Construction of Race, 1896–1954 by Lee D. Baker. Isis, 90(4): 855, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/237745.

Richmond:1999:BRB

[3586] Marsha L. Richmond. Book review: Familientreffen versus Professionselite? Vergangenheitsbewaltigung und Neustrukturierung in der deutschen Wissenschaftsgeschichte der 60er Jahre by Anke Jobmann. Isis, 90(4):855–856, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/stable/

237746.

Hansen:1999:BRBb

[3587] James R. Hansen. Book review: Countdown: A History of Space Flight by T. A. Heppenheimer. Isis, 90(4):856–857, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237747.

Krohn:1999:BRB

[3588] Wolfgang Krohn. Book review: The Sociology of Philosophies: A Global Theory of Intellectual Change by Randall Collins. Isis, 90(4):857–858, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237748.

Steigerwald:1999:BRB

[3589] Joan Steigerwald. Book review: Inscribing Science: Scientific Texts and the Materiality of Communication by Timothy Lenoir. Isis, 90(4):858– 859, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237749.

Teich:1999:BRB

[3590] Mikulas Teich. Book review: A Bio-Bibliography for the History of the Biochemical Sciences since 1800 by Joseph S. Fruton. Isis, 90(4):859, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994

(electronic). URL http://www.jstor.org/stable/237750.

Anonymous:1999:Cd

[3591] Anonymous. Collections. Isis, 90(4):859–860, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237751.

Anonymous:1999:BMf

[3592] Anonymous. Back matter. Isis, 90(4):861–866, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237752.

Anonymous:1999:VIb

[3593] Anonymous. Volume information. Isis, 90(4):867–882, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/237653.

Anonymous:1999:FMf

[3594] Anonymous. Front matter. Isis, 90(4):??, December 1999. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/237654.